A Complete Bibliography of Publications in the Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences

halieutiques et aquatiques (2000–2009)

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
      WWW URL: http://www.math.utah.edu/~beebe/

13 April 2024 Version 1.06

Title word cross-reference

+ [210, 1698, 1295, 71]. ^{12} [649]. ^{13} [981, 1719, 649, 1968, 373, 963, 1687, 1710]. ^{137} [44, 1302]. ^{14} [620, 649]. ^{15} [1088, 649, 150, 1571, 1968, 373, 1687]. ^{18} [1710]. ^{2+} [1734]. ^{3} [620]. ^{31} [1467]. 86 [813]. ^{87} [813]. ^{e} [926]. _{2} [1074, 1932, 14]. _{4} [2111, 652]. _{b} [1677]. a [1413, 1954]. ≈ 500 [944]. : [2122, 1577, 2062]. d^{2} [501]. δ [1088, 1719, 150, 1571, 1968, 373, 963, 1687, 1710]. r [612]. × [290].

-ATPase [210, 71].

/K [1295, 71]. /N [1677].

0-group [1717].

1980s [1026]. 1990s [1190, 1760]. 1A [41].

200-year [232]. 223 [20]. 23 [159]. 23-year [2131]. 24-h [236]. 2J3KL [385].

3-trifluoromethyl-4-nitrophenol [1670]. 302S [2130].

4-nonylphenol [1350].

500-year [1444].

70 [730]. 75th [926].

8 [1401].

Îles [1697, 1423]. Îles-de-la-Madeleine [1697, 1423].

abalone [855, 856, 817, 218, 1893, 1959]. ability [1822, 2037, 390]. abiotic [1865, 1508, 270, 1558, 499, 1083]. Abitibi [172]. ablation [1273, 125]. abnormally [1410]. abrupt [1561]. abrupta [1948]. absence [16]. absolute [1893]. Absorption [1878, 4]. Abundance [129, 911, 1666, 2025, 527, 632, 1797, 677, 1073, 586, 1233, 489, 891, 1177, 1665, 344, 1785, 800, 78, 1994, 59, 1657, 1530, 465, 3, 829, 418, 1787, 1095, 867, 799, 925, 949, 242, 1747, 444, 98, 698, 1043, 329, 556, 2084, 560, 1274, 1387, 77, 1208, 1959, 1248, 1444, 391, 27, 28, 29, 94, 1391, 428, 1735, 1323, 193, 693, 1232, 1062, 777, 2072, 1270, 2145, 1234]. abundances [1427, 1159, 130]. abuse [258]. abyssicola [1468]. acanthias [1999]. Acanthopagrus [941]. accelerating [2147]. accepting [1516]. access [1989]. accessible [1667, 828]. acclimation [2004, 360, 361]. Accounting [534, 329, 1845, 93, 2144, 700, 2123, 1364]. accretion [923]. accrual [550]. accumulation [116, 416, 1471, 1409, 108, 882, 1911, 1871, 33, 373, 1262]. accuracy [2025, 1858, 593, 147]. accurate [1948, 1905]. achieve [1417]. achievement [1987]. achieving [1555]. acid [1282, 1359, 205, 581, 117, 981, 643, 1602, 922, 160, 1558, 480, 313, 1883, 1276, 786, 801, 883]. acid-damaged [1276]. acidic [1086, 1524, 543, 1262, 1150, 2101]. acidification [1165, 1647, 2128, 1291, 700, 1209, 566, 1089, 1801, 2074, 1691, 706, 2002, 20, 2129, 1451]. acidified [1007, 55, 2130]. acidity [617]. acids [1256, 268, 688, 57, 981, 228]. Acipenser [1436, 1787, 2098, 1212, 2038, 1676, 360, 361]. Acoustic [1578, 231, 996, 804, 803, 1966, 756, 489, 1833, 59, 1271, 1029, 1689, 212, 598, 2055, 2060, 1554, 1772, 193, 1845, 2032, 2145, 600]. Acoustically [1903]. acoustics [2005, 1513, 1597]. across [1298, 1935, 1789, 2066, 1327, 1692, 402, 232, 1141, 2094, 1771, 1956, 1652, 106, 1751]. actions [789, 483, 484]. active [856, 1402, 1128, 305]. activities [1661, 1831, 338, 943]. activity [210, 1826, 1865, 291, 1607, 570, 2080, 1594, 2067, 1877, 897, 1490, 1698, 1643, 400, 705, 71, 1128, 1721, 496, 113, 1017, 1681]. actually [741]. aculeatus [1249, 1704, 995, 1640]. Acute [917, 71, 808, 480, 1688]. acutorostrata [853, 911]. adaptation [534]. adaptations [612]. adaptive [1127, 657, 1940, 848, 1747, 1134, 667, 1462]. added [2133]. addition [1425, 1087, 1086, 1062]. additions [896]. additive [236, 241, 2042]. additivity [554]. Adirondack [58, 722]. adjacent [785, 96, 1423]. adjust [497]. Admitting [1582, 1668]. Admixture [427]. Adriatic [360, 361, 1241]. adult [1667, 1006, 1654, 1370, 1785, 472, 1978, 247, 1044, 1143, 2118, 727, 1614, 1060, 703, 26, 850, 27, 28, 29, 666, 914, 523, 1720, 227, 513, 1245, 1410]. advantage [519]. advective [333]. aeglefinus [716, 251, 1823, 2045, 349, 2044, 1843, 1017]. aerial [2042]. aerobic [808]. aerosols [906]. aeruginosa [24]. affect [749, 1579, 285, 512, 511, 689]. affected [839, 47]. affecting [1510, 1079, 1042, 1182, 1159, 2146, 77, 1380, 122]. affects [1532, 1100, 71, 1703, 888]. affiliations [1384]. Africa [1785, 1204, 353]. African [920]. after [1460, 1190, 472, 1523, 902, 1394, 753, 132, 934, 1854, 1245]. against [1176]. Age [1411, 533, 1154, 227, 513, 1666, 146, 1278, 86, 64, 966, 1948, 857, 539, 1795, 846, 1696, 633, 1000, 1929, 1441, 174, 839, 1152, 1616, 1755, 1978, 947, 1874, 984, 569, 1622, 551, 2099, 521, 1183, 1049, 1508, 922, 987, 733, 1281, 802, 1222, 1825, 2048, 45, 508, 1516, 1868, 795, 1669, 6, 1412, 698, 422, 1304, 1862, 1596, 507, 1515, 1132, 499, 373, 1365, 1837, 1898, 1097, 364, 919, 640, 26, 2121, 532, 1821, 1653, 1517]. age-0 [64, 966, 569, 1508, 987, 45, 698]. age-composition [1696]. age-reading [1898]. Age-specific [227, 513, 1755, 1049, 26]. Age-structured [533, 1755, 1868, 6, 507, 919]. ageing [474, 1582, 1668, 728, 1372]. agent [817]. agents [820]. Ages [1983]. Aggregated [1434, 1651, 1091, 1747]. aggregating [1833, 1699]. Aggregation [187, 855, 1116, 962, 516, 1781, 2012]. aggregations [1833, 856, 1080]. Aggression [1185, 1593, 1796]. aging [372, 859]. agonistic [1792, 1355]. agricultural [1088, 1774, 31]. agriculturally [991]. Agulhas [282]. air [1756]. al [2045, 197, 177, 448, 176]. Alagnak [1444]. alalunga [1875]. alarm [622, 1524]. Alaska [509, 232, 2155, 836, 211, 1006, 614, 1629, 1540, 243, 1896, 996, 1537, 1272, 1205, 871, 649, 1396, 2000, 1847, 119, 545, 1991, 1444, 377, 930, 536, 209]. Alaskan [1988, 220]. albacares [2096, 299]. albacore [1875]. albatrosses [1186]. Albert [1400, 1399]. Alberta [1917, 591, 1178, 739, 1866, 1716, 376, 14]. alder [545]. Aldrich [1731, 1732]. Aleutian [212]. alevins [1636]. alewife [1531, 1986, 1852, 1394, 1683]. alewives [1267, 2006]. Alexander [90]. Alexandrium [541]. alga [2046]. algae [1547, 1076, 1100, 748]. Algal [636, 2138, 1073, 579, 1421, 2134, 744, 1481, 867, 1716, 946, 899, 8, 961, 2129, 543]. Algarve [1921]. algorithm [189]. algorithms [1772]. alkalinity [529, 1659, 204]. Allee [520, 52]. allegations [357]. allele [1693]. alleles [1, 1197]. allelic [501, 1841]. alleviate [907]. allocating [409]. allocation [644, 772, 733, 1334, 658, 1590, 1462]. allocations [582]. allochthonous [502, 1293]. Allometry [238, 630, 740]. Allozyme [634, 65, 879]. alluvial [1975]. Alnus [545]. alone [878]. along [477, 1785, 529, 1243, 1108, 681, 1255, 1481, 1886, 1430, 300, 1307, 748, 1486, 334, 1325, 623, 717, 173, 2075]. Alosa [1531, 1986, 1267, 1852, 1394, 1683, 2006, 1952]. alpine [1012, 174, 1206]. alpinus [1075, 347, 174, 787, 494, 1624, 144, 384, 72]. also [1111]. alteration [485]. alterations [1301, 570, 1250]. Altered [1525, 575, 1071]. altering [1572, 1573]. alternate [877, 1637]. Alternative [1046, 1162, 1498, 735, 2154, 718, 593, 158, 1372, 658, 1632, 2003, 866]. alterniflorum [1972]. alters [1530, 1965, 1208]. altitudinal [1307]. aluminium [144]. Aluminum [480, 413, 117, 1209, 1262]. Aluminum-dominated [480]. alutus [1825]. ambient [2080, 1853, 816, 601]. ambigua [1754]. America [446, 820, 1718, 1771, 1972, 69]. American [606, 562, 818, 953, 628, 982, 940, 1935, 957, 120, 738, 2008, 76, 2098, 2066, 101, 1969, 825, 380, 295, 1487, 1497, 602, 1952, 1001, 1485, 1748, 2026]. americana [2022, 136]. americanus [940, 463, 957, 582, 2008, 76, 1760, 380, 790, 1201, 1487, 602, 2026]. amino [922]. Ammodytes [2047, 1860, 981, 662]. Among [581, 781, 1850, 312, 1587, 1820, 63, 1547, 673, 416, 1501, 2111, 1880, 1830, 1675, 1787, 681, 186, 1508, 1305, 763, 2098, 1841, 733, 2118, 1414, 152, 1760, 218, 1987, 1304, 1037, 1477, 1955, 490, 1181, 1238, 40, 1125, 1391, 623, 717, 1603, 1117, 939, 2031, 73, 273, 1355, 752]. Among- [581, 1850]. among-lake [1477]. among-population [73]. Among-taxon [781]. amounts [2132]. amphibian [785]. amphibians [1722]. amphipod [977, 1393, 2084, 69]. Amphipoda [599, 186, 2084, 464]. amphipods [143, 1159, 675]. amplify [907]. ampullatus [1288]. Anadromous [2006, 210, 1075, 680, 1562, 1275, 1986, 1822, 1240, 435, 1521, 837, 973, 724, 1494, 1818, 216]. anadromy [1404]. Analyses [364, 2122, 1223, 1500, 754, 1877, 359, 1814, 1968, 1812, 939, 1932, 1120, 1019]. Analysis [95, 725, 48, 1780, 236, 1415, 316, 1317, 1, 1820, 74, 308, 776, 638, 1970, 1635, 1153, 2082, 348, 1896, 1939, 2142, 1892, 985, 1694, 427, 476, 1768, 237, 1606, 988, 1196, 1789, 1813, 1225, 1663, 844, 438, 929, 1954, 624, 163, 1476, 1068, 1922, 997, 1434, 1766, 1844, 972, 924, 533, 1881, 1142, 1783, 1053, 1218, 392, 278, 1693, 1990, 789, 1369, 483, 484, 1097, 1909, 1051, 663, 1139, 123, 99, 747, 2121, 1395, 1258, 1050, 295, 838, 1416, 125, 1613, 1821, 294, 1750, 918, 819, 1336, 164, 1517, 656, 22, 194, 736]. analyze [78]. analyzed [167, 1433]. analyzing [189, 640, 1117, 2151]. Anarhichas [1490, 662]. Ancestry [2120, 1757]. Anchovy [206, 1612, 430, 583, 2054, 1634, 1937]. angler [27]. angling [1838, 1294]. angling-induced [1838]. Anguilla [758, 1577, 166, 1327, 825, 1909, 870]. anguillid [684]. angular [2004]. animal [754, 1526, 496]. animals [612, 1660]. Annecy [1719]. Anniversaire [926]. Anniversary [926]. Announcement [1168, 1189, 1207, 1227, 1443, 1544, 926, 1617, 2058, 2059, 1729]. Annual [368, 1852, 252, 1075, 248, 1259, 1773, 1491, 1918]. annularis [1024]. annulus [640]. annulus-based [640]. anomalies [157]. Anoplopoma [129, 854]. anoxia [34, 303]. ans [1799]. Answers [1130]. Antarctic [1744, 900]. antarcticus [22]. Anthropogenic [817, 934, 1073, 116, 160, 338, 1806, 1812, 1458, 2085, 1888]. antipatharian [1983]. antipredator [192]. anurans [1965]. apex [1368]. apoptosis [315]. Appalachian [180, 466, 1551, 2065]. apparentés [893]. appearance [157]. Appel [1544]. Application [337, 74, 181, 278, 483, 484, 1278, 202, 1287, 952, 776, 231, 1441, 1456, 1896, 476, 617, 987, 229, 1997, 414, 1476, 423, 324, 2018, 661, 280, 1142, 705, 560, 1898, 2073, 1633, 692, 1416]. applications [1273, 520, 15, 1447, 1662, 656]. applied [708, 789, 297]. applies [1740]. apply [1330]. applying [465, 22]. Approach [266, 407, 644, 2077, 688, 2019, 1741, 2023, 1967, 1339, 1533, 1431, 1287, 956, 420, 638, 1007, 1275, 1155, 957, 947, 1321, 253, 1657, 459, 941, 1940, 974, 2095, 864, 225, 590, 1676, 324, 104, 85, 1963, 1099, 7, 1961, 1365, 1837, 1293, 1336, 1648, 1488, 69, 1681, 9]. Approaches [43, 1454, 1379, 1771, 909, 656]. approaching [895]. approximation [996]. Aquaculture [2134, 2137, 1941, 983, 351, 1175, 1212, 2136, 707]. aquarium [1840]. Aquatic [2051, 1540, 204, 1799, 268, 2125, 620, 586, 4, 2111, 744, 1098, 609, 1768, 1789, 2136, 343, 1897, 612, 33, 2029, 1707, 263, 1373, 262, 296, 1812, 490, 1759, 522, 136, 549, 1861, 1260, 1932, 1800]. aquatiques [1799]. aqueous [1405]. Archipelago [1959]. archival [2092, 1506]. archived [976, 721]. arctic [1391, 1075, 481, 1820, 1006, 1795, 1214, 347, 1501, 174, 787, 1459, 1479, 988, 1225, 494, 451, 764, 987, 1624, 1558, 244, 436, 1992, 2120, 144, 384, 2105, 623, 1286, 1821, 72]. arcticus [1006, 987, 2120, 1286]. Arcto [1775, 499]. Arcto-Norwegian [1775]. arctos [737]. Area [222, 2127, 2130, 1118, 557, 763, 1680, 142, 1803, 373, 963, 1590, 1224, 204, 2125, 412]. Area-dependent [412]. area-specific [373]. areas [1079, 168, 1112, 1675, 1957, 2033, 1374, 1449, 269, 1167, 1680, 1545, 2056, 1179, 540, 755, 1798, 1181, 1774, 1656, 746]. Argentina [725, 1192, 931]. argentinus [553]. aromatic [1525, 1433]. arrangement [1382]. array [2060, 496, 2032]. arrays [2008]. arrival [604, 850]. artedi [2153, 1402, 1250, 1170, 1856, 2007]. artefacts [787]. artificial [614, 764, 1990]. artisanal [1706, 1121]. Asian [1718, 313]. aspect [1028]. asper [874, 1375]. assaying [705]. assemblage [1174, 420, 1994, 840, 1685, 1309, 1352, 8, 550, 305]. assemblages [1864, 424, 1454, 1362, 1220, 371, 624, 1529, 245, 289, 1367, 300, 38, 753, 1850, 386, 461, 869, 1083, 522, 734, 2085, 2086, 732, 1751]. assess [1867, 1552, 1452, 1838, 1346, 1348]. assessed [833, 1608, 666]. Assessing [63, 1568, 1404, 1902, 1067, 855, 352, 1659, 121, 1438, 972, 1132, 857, 1833, 764, 238, 1943, 1798, 1378, 553, 1488]. Assessment [905, 815, 695, 43, 112, 1741, 1533, 1557, 74, 230, 628, 743, 983, 660, 1580, 410, 1349, 1655, 979, 427, 1940, 181, 636, 1210, 1543, 301, 302, 1476, 1427, 631, 1846, 1080, 331, 1352, 925, 147, 1814, 507, 1960, 919, 635, 2062, 854, 80, 1819, 920, 478, 927, 1232, 1288, 1724]. assessments [346, 306, 1898, 1812, 2073]. assignment [1, 524, 267, 423]. assignments [1247]. Assimilation [2136, 1320, 324]. assisted [247]. associated [1654, 1025, 1277, 246, 465, 897, 24, 2118, 896, 111, 1387, 1903, 2030, 1326, 193, 1146, 1859, 1343]. association [1057, 681, 138]. Associations [242, 175, 721]. assuming [1619]. assumption [1773]. assumptions [372, 467, 1078]. Asterias [1215]. Asterionella [1206]. asymmetry [219, 1308]. at-sea [2141]. Athabasca [1866]. Atlantic [897, 1944, 1327, 1644, 1068, 197, 906, 1593, 1158, 1282, 1359, 1620, 1947, 1407, 1031, 1734, 644, 1498, 50, 1106, 544, 1234, 1865, 1032, 1761, 179, 1054, 971, 207, 1607, 1371, 196, 857, 1191, 1489, 1377, 53, 846, 1567, 633, 1007, 1124, 891, 1977, 1572, 1573, 1661, 203, 1929, 1409, 1240, 83, 1949, 1962, 791, 1036, 1156, 1595, 1041, 878, 621, 1152, 1616, 120, 738, 2082, 1411, 1641, 403, 1939, 2142, 1874, 814, 1039, 2067, 368, 208, 815, 1479, 476, 2033, 363, 2034, 690, 2009, 1383, 923, 922, 81, 1133, 1585, 451, 330, 733, 1042]. Atlantic [1281, 802, 1044, 1642, 1061, 853, 1247, 1576, 1997, 1182, 1904, 349, 82, 1033, 414, 103, 1524, 400, 423, 1035, 1873, 901, 1446, 1314, 1315, 1942, 1028, 642, 1295, 142, 706, 518, 1038, 1403, 299, 2002, 788, 2157, 1231, 1116, 1905, 1614, 1542, 555, 327, 1969, 392, 1862, 1469, 1254, 1693, 1043, 1045, 1037, 862, 1132, 1879, 1550, 2117, 1040, 771, 1712, 159, 1767, 223, 712, 559, 473, 970, 1583, 2049, 1224, 1827, 1338, 199, 1506, 506, 911, 1344, 92, 704, 523, 1008, 1720, 240, 134, 1926, 80, 939, 2031, 1017, 72, 1895, 273, 479, 1163, 408]. Atlantic [164, 122, 655, 1063, 2149]. atlanticus [2109, 2069, 1505]. atlantique [1031]. atmosphere [2111]. atmospheric [1866, 652, 1742]. ATPase [210, 1698, 1295, 71]. atratulus [714]. atresia [2149]. attacks [89, 12, 1776]. attempt [975]. attributes [79, 972, 825, 185]. audax [1844]. augmented [1187]. auklets [1313]. aurata [1658, 1096]. auratus [1112]. auritus [1986]. Australia [1482, 1533, 546, 1111, 1685, 426, 1898, 734, 22]. auteurs [254, 491, 277, 685, 515, 1055, 1318, 1563, 1727, 2158]. Author [254, 491, 685, 1055, 1318, 1563, 1727, 2158]. Authors [276, 514]. Automated [670]. Automatic [1151]. autotrophic [140, 54, 1255]. autumnalis [1308]. aux [277, 515, 950, 1697]. Availability [962, 1971, 1016, 1527, 891, 699, 1470, 2005, 1508, 1100, 2055, 1897, 1924, 899, 549]. available [680]. avenues [1273]. averaging [1107]. aversion [409]. avian [1698]. Avoidance [542, 1356, 1855, 588, 1792, 1636, 1914]. avoiding [1142]. axis [1147]. azteca [1173]. aztecus [1177, 465].

B [306, 1053, 61]. Bacillariophyta [62]. Back [1942, 290, 422, 1777, 131]. Back-calculation [1942, 290, 422, 131]. backcrosses [1977]. backscatter [1723, 1554, 1845, 2145]. backscattering [1595]. backwater [833]. Bacteria [2030, 1897]. Bacterial [1100, 32, 586, 2134, 1560, 130, 11]. Bacterioplankton [1391]. Baffin [1220, 371]. bairdii [1973]. Balaena [220]. Balaenoptera [853, 911]. balance [1621, 2063, 1968, 71, 2090, 88, 42, 1670]. balanced [972]. balancing [1466]. Balearic [2087]. Ballast [754, 1157, 1993, 1557, 1611, 1779, 1277, 832, 613, 1778, 1269]. Ballast-mediated [754]. balteatus [286]. balthica [437, 615, 5]. Baltic [1256, 1620, 437, 1510, 241, 1962, 1957, 672, 1764, 1671, 1908, 396, 397, 676, 45, 608, 510, 1846, 85, 1254, 1879, 606, 852, 1224, 1827]. baltica [62]. band [2035]. bands [1372]. Banff [128]. Bank [2045, 2044, 797, 251, 581, 1823, 246, 171, 250, 133, 1395, 342]. Banks [355, 223]. barbatus [2097]. barcodes [1588]. Barents [1717, 1689, 864, 1043, 2031]. barium [1162, 1422, 213]. barotrauma [1842, 1838]. Barrier [146, 1880, 834, 921]. barrier-isolated [1880]. barriers [975, 974, 1384]. Base [1360, 1256, 1947, 1803]. based [1278, 1928, 2019, 1765, 905, 1, 719, 1979, 68, 1287, 420, 1860, 2124, 463, 664, 805, 1388, 588, 976, 2088, 1152, 1616, 1655, 243, 1896, 1397, 996, 1379, 1500, 298, 1951, 1435, 1560, 476, 237, 988, 1852, 841, 359, 1216, 2013, 1085, 1119, 1642, 649, 1539, 1521, 1476, 25, 1650, 1571, 1844, 1483, 1412, 422, 583, 1660, 587, 1955, 1048, 467, 1224, 640, 869, 1824, 958, 1462, 726, 1398, 920, 1681, 1019]. Baseline [1495, 1567, 427]. baselines [1369, 1915]. basic [297, 1262]. Basin [1173, 1016, 1482, 1992, 1303, 1455, 1413, 1661, 840, 1569, 402, 921, 680, 1527, 673]. Basin-scale [1016, 1527]. basins [691, 782]. basis [238, 697]. bass [2150, 1816, 1665, 1211, 1135, 56, 572, 1989, 48, 414, 2152, 289, 1486, 1528, 1721, 1771, 92, 1294, 1117, 1923]. bassanus [555]. batch [1180]. Bay [2022, 999, 34, 232, 1005, 380, 353, 370, 2011, 351, 1971, 394, 288, 1287, 1816, 1067, 1456, 243, 1272, 1205, 1612, 1220, 1182, 1904, 45, 1068, 1766, 760, 1937]. Bayes [507]. Bayesian [407, 230, 638, 1441, 1577, 1657, 941, 979, 399, 1606, 1954, 1997, 661, 1084, 429, 925, 924, 1345, 1814, 6, 637, 1151, 1218, 1963, 467, 663, 1743, 99, 2121, 406, 1050, 2103, 777]. bays [560]. be [1666, 2141, 447, 1111, 3, 781, 1689, 1961, 1248, 1819, 1361]. Beach [205]. Beacham [2161]. beaks [93]. beam [1736, 1271]. bearing [858]. bears [737, 1260]. Beaufort [220]. beaver [1396, 466]. bed [96]. been [701]. before [1190, 1962, 934]. beginning [1962, 6]. Behavior [1180, 749, 2036, 1463, 1855, 1020, 1235, 1604, 769, 1280, 1171, 1836, 1355, 1063]. Behavioral [2004, 1736, 2118, 57, 588, 1238, 286, 390]. behaviors [192, 1903]. Behaviour [1936, 904, 265, 1706, 413, 1690, 2137, 333, 117, 1341, 856, 1271, 1792, 188, 1699, 1219, 72, 1350, 1163]. Behavioural [330, 2073, 347, 2147, 233, 1877, 485, 1931]. behaviours [247]. being [832, 954]. benefits [1176, 959, 1226, 311, 954]. Benguela [583, 1398]. Benthic [2138, 2111, 1493, 2129, 140, 333, 39, 1425, 1214, 643, 579, 1421, 664, 805, 1204, 1076, 233, 1809, 228, 1381, 1255, 1169, 630, 1468, 1725, 1071, 1810, 304, 748, 994, 1847, 1393, 587, 2084, 1306, 1697, 696, 1903, 1450, 353, 734, 1148, 1187, 1546]. benthos [648, 1076]. benzoate [602]. Bering [1296, 1081, 996, 1311, 1222, 242, 525]. Bertalanffy [2162, 1582, 1668, 453, 1619, 521, 1889, 537, 692, 1077, 2028]. best [680]. betaine [896]. between [210, 146, 1549, 268, 1971, 1300, 1578, 2108, 1981, 735, 783, 1233, 1977, 1635, 1328, 60, 2097, 274, 253, 1602, 1877, 363, 501, 1461, 1953, 821, 750, 1169, 998, 1585, 494, 767, 1964, 1089, 1801, 1543, 2024, 1468, 1698, 48, 885, 1529, 1347, 1367, 2020, 242, 994, 698, 645, 2081, 724, 145, 1254, 565, 1780, 537, 876, 1139, 2037, 1575, 1982, 391, 1083, 350, 325, 717, 1302, 2054, 185, 687, 72, 779, 541, 106]. Beverton [1123, 1122, 2028]. Bias [1051, 92, 578, 1403, 1613]. biased [1144, 787, 1022]. Biche [1917]. Big [1545]. bigeye [1936]. bile [57]. bilinearis [616, 175]. Bill [90]. billfishes [659]. Bimodal [787, 1022]. bin [1197]. Bio [481, 730]. bio-indicators [730]. Bio-optical [481]. Bioaccumulation [118, 340, 1605, 143, 1512, 1173, 2153, 1108, 1169, 1353, 1824]. bioaccumulative [1079]. bioassessment [2021, 1352]. Bioavailability [111, 136, 413, 346, 1073]. biochemical [2133, 199]. biochimie [893]. biodegradable [1938]. biodeposition [1175]. biodiversity [743, 1428, 781, 1449, 1225, 2131]. bioeconomic [133, 1196]. bioeconomics [576]. Bioenergetic [2022, 44, 1877, 244, 724, 534, 1373, 425]. Bioenergetics [2064, 1484, 1046, 1024, 1023, 688, 2019, 453, 505, 2095, 923, 1908, 104, 889, 646, 1346, 293]. Bioenrichment [2052]. biofilm [413, 1425, 1474, 2066]. biofilms [54]. Biogenic [1998, 908]. biogeochemical [904, 1423, 1697]. Biogeochemistry [1079]. biogeography [459]. Biological [950, 893, 953, 457, 1588, 957, 415, 1827, 951, 954, 321, 1791, 1441, 348, 815, 465, 1612, 1538, 1476, 1541, 1737, 2030, 1845, 780, 751, 1019]. biologique [1122]. biologiques [950]. biology [271, 851, 459, 267, 779]. biomagnification [340, 1164, 2101]. biomanipulation [440]. biomarker [390]. biomarkers [981, 1013]. Biomass [142, 2138, 1584, 807, 648, 1214, 659, 231, 1471, 579, 1421, 4, 744, 1552, 1520, 1115, 1381, 310, 887, 867, 1229, 212, 510, 1846, 1893, 748, 946, 169, 8, 137, 717, 1600, 1819, 927, 2075, 1845, 61]. biomassof [1833]. biomonitor [416]. biomonitoring [387, 386]. Biophysical [672, 1980]. bioresource [1655]. biota [1057, 2136, 150, 296]. Biotic [356, 495, 2093, 270, 1352, 2081, 917, 1956]. bioturbator [2111]. biphenyl [2153, 1674]. biphenyls [118, 1169, 1373]. Biscay [1068, 1612, 1937]. bivalve [437, 503, 916]. black [986, 1251]. blacknose [714]. blacktip [557]. bladder [936]. blade [1867]. blade-strike [1867]. bleached [315]. Bleaching [574]. bloater [1390, 692]. block [212]. Blood [928, 853, 730]. bloom [2106, 1823, 2045, 24, 1716, 1558, 2044, 319]. blooms [1166, 370, 541]. blue [1431, 1287, 1377, 863, 1067, 1175, 2034, 366, 1392, 1815]. bluefin [1641, 1642, 1997, 1862, 860, 1506, 1008]. bluefish [1488]. bluegill [2153, 290, 1509, 876, 1710]. board [434]. boats [489]. Body [1784, 1309, 1713, 749, 35, 772, 1520, 2009, 624, 1783, 217, 824, 149, 1017, 1589, 1343, 575]. bog [1011]. boldness [1624]. bomb [1154, 539, 1795, 1825]. Bonavista [380]. Bonneville [1307, 1376]. Boreal [1230, 739, 569, 296, 1647, 1471, 200, 2134, 601, 967, 222, 960, 451, 2021, 1086, 1535, 691, 723, 1274, 1623, 1894, 405, 1148, 464, 445, 2139, 1089, 1801, 376, 1380, 1276, 14]. borealis [243]. Boreogadus [1459]. born [765]. both [269]. bottlenecked [201]. bottlenecks [493]. bottlenose [1288]. Bottom [1085, 1578, 615, 664, 805, 1454, 1610, 1851, 497, 1689, 1220, 355, 875, 2055, 1746, 746]. Bottom-up [1085, 1454, 1610]. bottom-water [615]. botulinum [2046]. Bou [158]. Bou-Rouch [158]. Bouclier [1799]. boulder [891, 709]. boulder-rich [891]. bound [136]. boundaries [774, 2031, 746]. boundary [1913]. bouvieri [1575]. Bow [128, 591]. bowhead [220]. box [434]. brackish [360, 361, 1009]. braided [477]. branches [1010]. Branchial [1707, 870]. Branching [1244]. Brazil [1121]. breadth [378]. breakdown [1447, 1737]. bream [1658, 1096]. breeding [332, 1299, 1092, 214, 885]. brevipes [1631]. brevipinnis [498]. brevirostrum [1676]. Bristol [288, 1456, 1272, 1205, 1766, 760]. Britannia [205]. British [205, 2160, 2, 226, 1555, 301, 666, 1985, 2161, 201, 1410, 1313, 793, 969, 375, 2070, 1722, 1194, 302, 2092, 322, 2056, 674, 845, 709, 1960, 387, 386, 1181, 1337, 1737, 27, 28, 29, 227, 513, 838, 1416, 273, 11, 1888]. broadcast [1583]. Brome [130]. Brook [1607, 1039, 1614, 722, 1165, 1299, 1967, 1339, 153, 1404, 1635, 1916, 160, 837, 730, 1126, 724, 1494, 1786, 826, 1780, 1325, 1623, 293, 149, 1551, 2065, 1343]. brown [210, 1114, 495, 1177, 1200, 465, 214, 427, 363, 823, 494, 1886, 1004, 1279, 1172, 1692, 1899, 1307, 71, 468, 1786, 1609, 2016, 737, 1636, 221, 1623, 1561, 1686]. Brownie [1365, 1837]. Browns [251, 250]. Brunswick [1607, 1039, 1614]. Buckley [197]. budget [1941]. budgets [1046, 2127]. buffer [1057]. buffering [1360]. bugensis [562, 818, 1229, 697]. Building [1122, 1123]. Bull [1274, 2025, 148, 459, 443, 1574, 124, 1325, 295, 1416]. bullheads [1370]. buoy [1271]. burbot [1935, 1953]. burial [1380, 381]. burned [1230]. burying [2047]. bycatch [668, 203, 1939, 1483]. Bythotrephes [1018, 806, 1876, 2119, 1276, 1503, 475, 2007].

C [649, 620, 981, 1719, 649, 1968, 373, 963, 1687, 1710]. C-labelled [981]. C/ [649]. CA [1476, 1734, 1047, 2122, 1577, 14]. Cache [358]. caddisfly [121]. Cadmium [54, 984, 2041, 1366, 1871, 33, 917, 213]. Cage [983, 1941]. cages [1697]. calanoid [326]. Calanus [53, 320, 319]. calcite [1998]. Calcium [441, 1405, 1804, 1065]. calculating [2145]. Calculation [310, 746, 290, 1942, 1151, 422, 131]. calibration [890]. California [1630, 1700, 1549, 1199, 95, 1052, 1568, 1547, 863, 1252, 817, 2095, 1216, 2013, 1842, 1438, 1203, 959, 100, 145, 1303, 683, 1662, 419]. californica [39]. Call [1544]. Callinectes [1431, 1287, 1067, 766, 366, 1392]. Callorhinus [976]. campechanus [1934]. camtschaticus [836]. Can [1312, 2141, 447, 1191, 1111, 3, 1851, 1689, 1130, 1248, 199, 1666, 1158, 781, 266, 1357, 694, 1819, 1361]. Canada [2160, 529, 1500, 1831, 371, 691, 2127, 2130, 130, 1423, 1697, 1878, 591, 2161, 1947, 1173, 793, 2110, 1191, 1377, 739, 743, 983, 2082, 1659, 1175, 1866, 617, 269, 1904, 1430, 832, 318, 2157, 898, 1170, 444, 845, 709, 1992, 486, 1181, 1337, 1737, 1917, 958, 1239, 1891, 80, 1985, 1927, 201, 475, 2007, 2026]. canadensis [522]. Canadian [1501, 1391, 623, 1799, 1300, 1559, 1579, 806, 1382, 1474, 774, 306, 1958, 1504, 1479, 1789, 81, 764, 1476, 1758, 1881, 1053, 1806, 262, 1274, 259, 838, 1487, 543, 1800]. canadien [1799]. Canal [2036]. canary [1154]. Cancer [1215, 1019]. candidatures [1544]. cannibalism [740, 163]. canonical [231]. canopies [1217]. canopy [1062]. Cantabrian [1004]. Canyon [666, 321]. capacities [711, 618]. capacity [1734, 975, 1949, 1594, 1641, 834, 1360, 1658, 764, 1642, 695, 1417, 392, 169]. Cape [2052]. capelin [264, 1341, 1717, 2115, 138, 1828, 516]. capensis [583]. captive [885, 1624]. captive-reared [1624]. captively [332, 749]. capture [1312, 472, 35, 1873, 25, 1753, 663, 26, 129, 428, 390, 874]. capture-recapture [25, 26]. captured [170]. capturing [1721]. Carbocentric [1777]. Carbon [1380, 992, 1579, 614, 546, 2100, 306, 1802, 1160, 985, 551, 2132, 960, 1916, 767, 1089, 1801, 1061, 1068, 653, 742, 1507, 1924, 1961, 1293, 1326, 1938, 549, 1678, 126, 32, 1807]. carbon-rich [767]. carcass [1425, 1062]. carcasses [1592, 1847, 2121]. Carcharhinus [1581, 557]. Carcharias [1785]. Carcinus [1215]. cargo [1277]. cargo-carrying [1277]. Caribbean [1759]. Carlo [2154, 1091, 2000]. Carolina [503]. carp [1718, 703, 1485]. carpio [703, 1485]. Carry [1636]. Carry-over [1636]. carrying [1277, 1641, 1642, 392]. cascade [433, 778, 70]. cascades [140]. cascading [1309]. case [1158, 447, 1948, 1533, 836, 1588, 2124, 874, 940, 1067, 1577, 1940, 611, 1612, 2116, 1764, 2053, 1516, 1699, 1447]. Caspian [607, 1104, 609, 606, 464, 610]. Cassin [1313]. Catamaran [1607, 1039, 1614]. cataractae [1442]. catastrophic [1648, 408]. Catch [1249, 1815, 1666, 1475, 2141, 1705, 2108, 1981, 288, 991, 729, 1736, 1657, 418, 310, 1374, 1183, 1466, 1842, 1222, 1995, 331, 795, 7, 1365, 1837, 547, 1884, 784, 538, 224, 92, 768, 421, 819, 1232, 2103, 1517, 2072, 595]. Catch-and-release [1815, 1842, 92, 2103]. catch-at-age [1666, 1183, 1222, 795, 1365, 1837, 1517]. catch-balancing [1466]. catch-comparison [729]. catch-per-unit-effort [418, 784]. Catch-up [1249]. catchabilities [533]. Catchability [710, 592, 399, 1849, 1664, 1084, 1487, 1161, 1517, 122]. catches [1855, 1936, 497, 1247, 1521, 1412, 1819, 576]. catching [377]. catchment [874, 960, 1803, 387, 386, 1103]. catchment-scale [1103]. catchments [1016, 2114, 2132, 1360, 1987, 742, 2127, 376, 1293, 1527]. categorical [935]. cation [1360]. cations [1803]. Catostomus [1300, 440]. Catskill [1165]. caurinus [626]. Cause [190, 346, 719, 1520, 1447, 1130]. cause-effect [719]. caused [1491]. cautionary [1655, 949]. caveats [267]. Cd [416, 759, 1108, 172]. cell [74, 315, 896, 730]. cell-associated [896]. cells [895, 1239]. Cellular [437]. centenarian [2109, 2069]. centers [496]. Central [95, 1805, 1059, 180, 659, 660, 1986, 2114, 1809, 617, 700, 763, 1192, 1105, 302, 2096, 1179, 320, 1987, 1803, 1960, 1325, 1662, 2065, 786, 69, 972]. centrarchid [762]. Centre [1799]. Centroselachus [1372]. century [446, 884, 543]. cepedianum [1725, 113]. cephalopod [93]. Cephalopoda [336]. Ceriodaphnia [480]. cernuus [1153]. cetaceans [1888]. CH [2111]. chain [1602, 568]. chalcogramma [2155, 1283, 1629, 1311, 1604, 2055, 2000]. challenge [1765, 361]. challenges [665, 1716, 1812]. Champsocephalus [1284]. Change [794, 452, 748, 1127, 1052, 1906, 2, 830, 1101, 180, 1756, 2154, 873, 744, 1552, 569, 1957, 2095, 2070, 700, 430, 822, 1357, 858, 848, 444, 695, 132, 993, 1202, 547, 1134, 970, 1248, 2148, 1771, 1733, 1030]. Change-in-ratio [794]. Changes [1018, 1455, 682, 1567, 1523, 2132, 15, 1518, 339, 853, 1810, 322, 1394, 119, 691, 963, 130, 475, 1075, 1249, 1073, 1531, 264, 1007, 983, 500, 1155, 969, 1802, 78, 808, 1659, 944, 2087, 81, 1044, 1543, 1558, 1692, 1038, 1131, 1132, 2127, 486, 1806, 19, 425, 707, 1005, 1074, 1387, 852, 326, 869, 199, 531, 1286, 651, 1678, 652, 1742, 870, 1150, 1807]. Changing [667, 1217, 1910, 1584, 1049, 2152, 731]. Channel [983, 516, 1016, 1527, 800, 1530, 974, 630, 187, 1535, 1542, 70, 573]. channel-dwelling [1530]. channels [344, 1310, 1636]. Channidae [1718]. chaoborids [998]. Chaoborus [1176, 416, 759, 1080, 33, 1102]. char [1075, 1404, 347, 174, 787, 443, 494, 1624, 132, 144, 384, 295, 72]. Characteristics [305, 312, 481, 592, 1144, 996, 960, 923, 757, 1881, 1244, 917, 556, 1759, 391, 1824, 149, 1628, 573]. Characterization [1612, 1467, 1497, 1966]. characterizes [1309]. Characterizing [398, 1367, 595]. charismatic [1337]. charr [837]. che [950]. check [945]. Chemical [1112, 1074, 622, 162, 1524, 1064, 587, 2073]. chemicals [1337]. Chemistry [893, 617, 1744, 2122, 1007, 1822, 1601, 665, 984, 388, 2128, 1329, 1810, 466, 742, 1963, 913, 945, 1703, 144, 1832, 18, 928, 786, 213]. Chemosterilization [689]. Chesapeake [2022, 1287, 1816, 1067, 2011]. chimie [893]. Chinese [351]. Chinook [2160, 177, 2161, 176, 589, 1047, 95, 1714, 1457, 1587, 1426, 411, 345, 749, 992, 63, 291, 219, 2156, 2112, 1885, 482, 1184, 1639, 168, 167, 886, 2099, 834, 1560, 501, 1437, 237, 307, 763, 1342, 1473, 889, 1525, 1695, 565, 789, 1638, 723, 483, 484, 369, 769, 2060, 1662, 989, 919, 192, 1903, 287, 185, 687, 928, 1985, 1355, 41, 1120]. Chionoecetes [1296, 1081, 1340, 1767, 678, 768]. chironomid [920]. chironomid-based [920]. chironomids [1430, 34]. Chironomus [1602]. chitobiase [2080, 198]. chloride [1093]. chlorophyll [161, 53, 1155, 1413, 2063, 594, 1954, 1687, 1700, 456]. chlorpyrifos [620, 868]. choice [37, 298, 866, 1937]. chondrichthyans [1533, 1740]. Chondrostoma [35, 79]. chromosome [1184]. chronic [1794, 284]. chronically [1147]. chronologies [1795]. chrysophytes [486]. chum [2036, 1887, 1304, 845, 1901, 96, 1480, 2076]. Chutes [1332, 1333]. cichlid [2020]. circadian [570]. circulating [1478]. circulation [558, 319]. Circumpolar [1744]. cirri [84]. cisco [2153, 1402, 1250, 1856]. Cladocera [1993, 611]. Cladoceran [793, 824, 1018, 480]. Cladophora [2046]. clam [437, 615, 773]. clams [902, 955, 503, 313]. clarity [594, 165, 1387]. clarki [821, 1217, 968, 596, 826, 1217, 77]. clarkii [2023, 1788, 1880, 1307, 2071, 1780, 1376, 1575, 2027, 2023, 1880, 1270]. class [1282, 1359, 385, 1874, 226, 605, 1854]. classes [1974, 1674]. Classical [1247, 1961]. classification [1454, 600, 804, 1352, 803, 2094]. classified [1612]. classify [1513, 1597]. clay [998]. clean [711, 437, 87, 10]. clear [1793, 301, 302, 2031, 445]. clear-cutting [301, 302]. clear-water [1793, 445]. clearcut [674]. Clearwater [1354]. Climate [744, 679, 444, 425, 1991, 1052, 1313, 2, 180, 1756, 738, 1957, 2128, 2070, 430, 822, 1357, 1281, 2024, 1386, 775, 993, 2127, 19, 1615, 970, 989, 259, 1771, 1891, 2054, 1999, 792, 1534]. Climate-forcing [679]. climate-induced [19, 989]. climate-related [1891]. climatic [2154, 1813, 1216, 2013, 1420, 1324, 1808, 605]. climatology [905]. clonal [378]. closed [146, 1920, 2033]. closely [1235, 1468]. Clostridium [2046]. closure [1762, 1522, 1959, 1937]. cloud [814, 498]. Clupea [1371, 509, 1595, 517, 450, 2115, 382, 681, 864, 1764, 142, 799, 2113, 159, 852, 536, 227, 513, 2149]. clupeaformis [1784, 2146, 294, 2001, 1912]. cluster [278]. clusters [167, 709]. Cnidaria [761]. CNRC [926]. co [2097, 1890, 1074, 1932]. co-occurring [2097, 1890]. coalescent [2124]. coalescent-based [2124]. coarse [1244]. coast [1317, 1785, 1841, 875, 301, 832, 1226, 1828, 1154, 1181, 325, 533]. Coastal [2152, 2023, 596, 394, 179, 1489, 1848, 969, 1601, 1598, 375, 1880, 880, 1115, 442, 594, 1095, 1140, 896, 871, 1529, 103, 748, 85, 782, 1486, 1166, 1217, 1348, 28, 1746, 94, 126, 1416, 2085, 2086, 1656, 2006, 675, 1888, 655, 409]. coasts [1960]. cod [236, 1158, 1282, 1359, 1620, 385, 1407, 447, 544, 1234, 179, 1054, 971, 196, 1191, 1489, 241, 846, 633, 1233, 670, 1984, 1572, 1573, 1929, 83, 1949, 1962, 699, 1152, 1616, 1459, 403, 2067, 1717, 1479, 672, 988, 923, 922, 81, 1133, 1585, 864, 802, 1334, 2024, 396, 397, 676, 1576, 82, 103, 197, 1314, 1315, 1942, 244, 1403, 788, 2157, 1231, 1116, 392, 1254, 1693, 138, 1132, 499, 1775, 816, 1550, 2117, 771, 1931, 223, 1884, 1770, 473, 66, 970, 1583, 1843, 2049, 1827, 224, 1338, 1344, 2105, 1720, 240, 134, 1926, 122, 655]. coded [209, 603]. coded-wire [209, 603]. codeficiency [969]. coexisting [2020]. cognatus [985, 1453]. cohesion [2029]. Coho [1667, 332, 1910, 2, 68, 1568, 170, 472, 434, 1523, 1536, 882, 183, 1592, 97, 442, 1792, 1841, 542, 1632, 1396, 865, 1438, 1285, 1677, 782, 1226, 556, 1753, 1790, 77, 1208, 868, 391, 292, 1171, 603]. cohort [89, 12]. cohorts [2033, 2018, 1669]. cold [1949, 1089, 1005]. cold-ocean [1005]. cold-water [1089]. collapse [1190, 2157, 554, 224, 1634]. collapsing [1132]. collected [1578, 489, 1108, 1689]. collector [912]. Colonization [232, 1667, 626, 211, 1479, 993, 1224, 1010]. colonized [1413]. colonizing [1691]. colony [2052]. color [1532, 326, 574]. Colorado [1316]. colour [529]. Columbia [2160, 793, 845, 709, 1181, 1337, 1737, 177, 1985, 2161, 201, 176, 1415, 205, 1313, 680, 63, 2, 1654, 969, 482, 168, 167, 375, 834, 226, 2070, 1722, 1555, 1143, 1950, 1194, 301, 302, 2092, 322, 2056, 674, 1060, 425, 1960, 387, 386, 27, 28, 29, 666, 227, 513, 838, 1416, 273, 11, 1888, 1410]. column [481, 265]. combination [465]. combine [729]. Combined [1932, 2077, 975, 1537, 1365, 1837]. Combining [1867, 1814, 1183]. Comment [2160, 196, 511, 469, 1418, 197, 2044, 803, 1778, 1779, 156, 2045, 804, 1419, 155, 2161, 448]. Comments [176, 449, 382, 177]. commercial [341, 1377, 170, 472, 434, 1736, 1530, 1851, 1309, 245, 7, 721, 1130, 531, 2072]. commersoni [1300, 440]. common [1578, 482, 1119, 1037, 703, 92, 1600, 1485, 736]. commonly [1433]. Communication [412, 601, 374, 600, 2029, 602, 335]. Communiqué [1443, 1544, 926, 1168, 1189, 1207, 1227, 1617, 2058, 2059, 1729]. communities [938, 195, 1018, 1455, 140, 806, 440, 739, 1834, 2100, 873, 1388, 58, 2134, 60, 601, 1809, 1500, 1768, 1682, 1794, 1255, 1225, 270, 964, 2021, 566, 1071, 1810, 232, 691, 1477, 1703, 376, 1387, 696, 1548, 1894, 1733, 1276, 961, 1148, 2129, 1972, 14, 9, 1343]. Community [1159, 2039, 383, 1906, 1559, 1811, 830, 1101, 1531, 1848, 1129, 741, 1647, 1188, 664, 805, 1204, 171, 1958, 1435, 700, 785, 1518, 1519, 1886, 1192, 1558, 402, 322, 218, 1241, 119, 1507, 278, 217, 1955, 1261, 899, 1387, 694, 2135, 1206, 1511, 717, 31, 693, 342, 1013, 1406, 2075, 445, 1030, 1109, 629, 1069, 575]. community- [1955]. Comparative [562, 818, 743, 1718, 188, 1454, 348, 1783]. compared [884]. Comparing [648, 1061, 1522, 1747]. Comparison [1388, 1152, 1616, 253, 1877, 1723, 1919, 507, 19, 1097, 386, 958, 131, 641, 407, 2019, 634, 308, 1413, 1978, 729, 718, 1449, 1169, 636, 1529, 778, 34, 1367, 1251, 1968, 376, 96, 1128, 1185, 619, 1771]. Comparisons [1461, 145, 1741, 594, 1477]. compartments [2111]. Compensation [1417, 1249, 1493, 764, 987]. Compensatory [1640, 52]. competing [616, 479]. Competition [1982, 273, 1424, 332, 1865, 1299, 1370, 21, 1340, 708, 1537, 363, 639, 1126, 468, 1141, 1590, 1991, 989, 779, 1923]. competitive [1307, 1638, 286]. complanata [112, 1504]. complete [93]. complex [589, 52, 701, 1468]. complexity [558, 2098, 289, 695, 1244, 1242, 1895]. component [523]. components [1192, 220, 1653]. composés [893]. composition [1278, 1713, 2141, 749, 1811, 2108, 1981, 1696, 1848, 1082, 741, 991, 699, 839, 450, 561, 15, 469, 785, 887, 922, 470, 272, 767, 1309, 1412, 119, 946, 217, 373, 1806, 2135, 1198, 31, 1503, 1891, 906, 2075, 445, 1472, 106, 1069]. compositions [842]. Compounds [893, 116]. compréhension [1799]. comprehensive [260]. compression [588]. concealment [639]. concentration [53, 1214, 881, 162, 508, 1913, 1700]. concentrations [1230, 84, 905, 1584, 109, 759, 1155, 984, 1602, 2132, 1405, 1801, 1908, 853, 1946, 1769, 1803, 303, 18, 1930, 1302, 1804]. concept [1330, 1618]. Conception [1005]. concepts [1896, 2028]. conceptual [257]. concerns [1545]. concordance [1811, 1956]. concurrence [1971]. concurrent [46, 794]. condition [716, 74, 1876, 1784, 1654, 633, 83, 156, 1823, 2045, 1002, 2044, 618, 2146, 1956, 155, 1017, 928]. condition-dependent [1654]. conditioning [597]. conditions [1457, 241, 770, 1671, 160, 322, 1043, 105, 1702, 1338, 1323, 995]. conduct [1019]. conducted [1836]. confidence [1301, 1119]. configuration [70]. conflict [735]. conflicts [769]. confluences [1541]. confluentus [2025, 148, 459, 443, 1574, 124, 1325, 1274, 295, 1416]. confusion [51]. congruence [781]. connect [1964]. Connecticut [1986, 423, 712]. connection [1094]. connectivity [2051, 1745, 1362, 561, 1545, 1181, 1832, 1603]. consequence [948]. Consequences [354, 1619, 1641, 886, 1516, 592, 766, 1133, 82, 1035, 1775, 1890, 311, 384, 1070, 1686, 1453]. consequent [570]. Conservation [1397, 267, 1158, 1317, 1797, 2107, 817, 459, 2043, 186, 1749, 1212, 269, 1950, 1375, 2056, 1226, 565, 1798, 2120, 1337, 1458, 2090, 968, 930]. conserve [918]. conserving [247]. consideration [1694]. considerations [2043, 2068, 656]. considering [1791]. Consistency [2053, 1578]. Constance [1600, 2017]. constituents [883]. constraining [1627]. constraints [1308, 778, 1072]. construct [97, 429]. constructed [1310]. Construire [1122]. consumer [140, 1915]. consumers [1255]. Consumption [1499, 864, 1683, 1023, 2019, 1762, 827, 453, 2102, 253, 1051, 293, 2028, 88, 42, 115]. consumption-dependent [1023]. contact [1935, 138]. contaminant [645, 1827]. contaminants [1079, 2073, 390]. contaminated [711, 1071, 111, 1707, 110]. contamination [112, 116, 1108, 1987, 618]. contemporary [626]. content [1532, 922, 438, 767, 91, 1051, 717, 126]. Contents [255, 492, 669, 1056, 1319, 1564, 1728, 2159, 236, 1033]. Context [1661, 457, 588, 923, 1266, 631, 2003, 1365, 1837, 1444]. Context-dependent [1661, 588]. Continental [1952, 1407, 877, 166, 1675, 1383, 322, 173]. Continental-scale [1952]. continua [629]. continuation [364]. continuation-ratio [364]. Continued [1393]. continuous [2023, 1793, 1342, 326]. Continuum [1633, 1200, 1330]. Contrasting [1475, 333, 177, 2037, 2054, 1628, 176, 91, 1542, 145, 1293]. contribute [2006]. contributing [1967]. Contribution [2112, 1433, 1714, 211, 1375, 1961, 723, 932]. control [735, 2088, 1823, 2045, 1694, 1940, 1560, 1351, 1360, 1944, 2044, 1541, 1991, 687, 1187, 1361]. controlled [1738, 1703, 1245]. controlling [831, 798, 1826]. Controls [632, 1570, 891, 1864, 200, 701, 270, 1514, 932, 623]. conundrum [2068]. conventional [64, 1875, 1716, 1997]. Convergence [282, 275]. conversion [1046, 1099, 223, 2028]. Copepod [1863, 1268, 559, 1548, 1439]. Copepoda [374]. copepodids [1356, 1297]. copepods [326]. Copper [284, 999, 437, 1599, 626, 570, 984, 1093, 868, 1480, 1396]. coral [1118, 873, 851, 1936]. corals [1983]. Corbicula [313]. coregonid [1646]. coregonids [1518]. coregonines [1907]. Coregonus [1390, 1784, 2153, 1719, 1385, 1308, 1402, 1250, 1170, 1856, 963, 2146, 692, 1600, 294, 2001, 2017, 1912, 2007]. coriacea [1210]. cormorant [1986]. correcting [1161]. correction [1024]. correlated [78, 1677]. correlates [910, 1559, 1448, 1363, 1764, 1384, 1136, 704, 928, 1410, 1681]. correlation [63, 2004, 231, 2123, 537]. correlations [274, 1131]. correspondence [1476]. Corrigendum [2108, 2013, 1946, 755, 2012, 2162]. corticosteroid [1688]. cortisol [172]. corvina [1316]. COSEWIC [2053]. cost [1249, 36, 570, 791, 1505, 1721, 149]. Costs [1176, 2025, 1046, 1036, 1156, 1128, 311, 44, 954, 666]. Cottus [1973, 1370, 985, 1453]. Could [1520, 1961]. counts [1885, 763, 1735, 185]. coupled [1273, 1510, 904, 125]. Coupling [1908, 1011, 1725, 1932]. covariability [63]. covariates [1743]. Covariation [1233, 1004, 322, 1649, 395, 1662]. cover [596, 1246, 1958, 814, 1263, 1839]. covered [481]. cowcod [1724]. crab [288, 836, 1431, 1287, 1067, 1296, 1081, 766, 1340, 1767, 678, 1631, 958, 768, 1215, 1019]. crabs [366, 1392]. crameri [1223]. Crangon [790, 1201]. cranial [36]. crappie [312, 1024]. Crassostrea [84]. crayfish [1292, 1811, 1020, 1351, 1953, 1924, 2130, 1429, 1346, 1389, 1030]. create [1538]. Creek [1438, 358]. creeks [1759]. crepidater [1372]. crested [1986]. CRESUS [874]. criteria [167, 181, 518, 1516, 849, 2140]. Critical [617, 194, 1741, 116, 375, 700, 1676, 1486, 1378, 1451]. croaker [203]. Cross [164, 1918]. Cross-tributary [164]. crossdating [1259, 1948]. crotonensis [1206]. crucial [268]. cruising [1046]. Crustacea [1588, 198, 599, 186, 464]. Crustacean [566, 1090, 1018, 2039, 806, 1829, 1958, 1405, 278, 1503, 1817, 1104]. crustaceans [648, 374]. Cruz [931]. Cryptic [69]. crystal [1290]. Cs [44, 1302]. Ctenophora [761]. Cu [1108, 1353, 172]. Cuban [616]. cucumber [419, 879]. cues [1524, 1828]. Culaea [1343]. cultivated [352]. Cultivation [261]. Cultivation/depensation [261]. cultural [1917]. culture [1175, 279, 350, 353]. cultured [1971, 882]. Cultus [2032]. Cumulative [1381, 1831, 937, 137, 1274, 259, 115, 1063]. Current [749, 153, 952, 1644, 1126, 714, 909, 786, 1549, 1744, 1252, 2095, 145, 1856, 683]. currents [324, 325]. curve [1278, 2121]. curves [67, 1582, 1668, 729, 590, 1208]. cutthroat [2023, 596, 1788, 1880, 821, 1307, 2071, 826, 1780, 1217, 77, 311, 968, 1376, 1575, 1982, 2027, 1270]. cutting [301, 302]. Cyanobacteria [431]. cyanobacterial [1520, 1243, 1946, 1769, 1882]. cycle [1122, 74, 116, 1867, 319, 1742]. cycles [312, 46, 2063, 320, 444]. cycling [2004, 1059, 150, 901, 909, 110]. Cynoscion [2004, 997, 1316, 720]. Cyprinid [696, 79]. Cyprinidae [1718]. Cyprinus [703, 1485]. Cystidicola [936]. cysts [157]. Cytochrome [41]. cytometric [74].

dace [1442, 1900, 714]. daily [474, 59, 1710, 1534]. dam [1654, 654]. damage [357, 928]. damaged [1276]. damming [132]. dams [1867]. danio [1238, 1238]. Danish [2047, 204]. Daphnia [383, 1176, 564, 1704, 675]. dark [1029]. darkblotched [1223]. Darling [1482]. Darwinian [1124]. data [1666, 1278, 2141, 1820, 1578, 1948, 512, 161, 1696, 367, 489, 2079, 231, 1111, 1621, 859, 1939, 1996, 990, 511, 1892, 1657, 465, 941, 427, 557, 310, 328, 1183, 1689, 988, 1789, 935, 1875, 1849, 2098, 1513, 1597, 844, 843, 1247, 1664, 1997, 590, 1091, 324, 25, 147, 924, 1814, 6, 1142, 1963, 1202, 816, 1369, 7, 1365, 1837, 1554, 721, 663, 326, 640, 538, 26, 224, 1817, 1999, 819, 1232, 1648, 2103]. data-limited [1657]. databases [1002]. date [1064, 850, 115]. dating [2109, 2069, 539, 1918]. Davis [1199, 1220]. day [411, 345, 1596]. daytime [2007]. dead [1902, 1672]. death [902]. debris [183, 654, 283, 573]. Decadal [1543, 326, 1951]. decade [30]. decades [988, 860]. decay [967]. deciduous [1217, 1737]. deciphering [2094]. decision [1415, 1694, 695, 483, 484, 99]. decisions [866]. decline [2, 1428, 190]. declines [2025, 2114, 1682, 441, 304, 1959]. Declining [2050, 446, 1630, 1945].

Declivity [30]. decomposition [1010]. deep [598, 1983, 173, 456]. deep-sea [1983]. deepwater [1372, 1821]. defense [1176]. deficiency [2047]. define [580]. defining [1113, 1676, 7, 2107]. definition [1364, 1925]. Definitive [883]. deformed [1553]. degradation [977, 198]. Degree [1759, 1236, 1596]. degree-day [1596]. dehydrogenase [1721]. del [725]. delay [776, 882]. delayed [2156, 1654, 2060]. delaying [1338]. Delta [1501, 1396, 1391, 623, 1700, 717]. demand [1214]. demersal [333, 1341, 2097, 2095, 2143, 2068, 119, 1828, 356, 1926]. demise [1623]. Demographic [825, 22, 2124, 1212, 398, 429, 1286, 1750]. demography [1407, 1137, 1574, 1045]. demonstration [1024, 72]. Denitrification [932, 1401]. Denmark [1009]. dense [1080]. densities [1029, 304]. Density [124, 737, 1615, 283, 2065, 23, 1488, 1593, 1971, 1993, 46, 1292, 615, 580, 1370, 770, 1635, 976, 2144, 1264, 2008, 1874, 1363, 886, 432, 526, 1849, 2038, 121, 48, 624, 1172, 1692, 1893, 436, 1609, 2050, 1763, 1526, 1770, 1965, 1208, 532, 1999, 1895]. density- [2038]. Density-dependent [124, 737, 2065, 23, 1488, 580, 976, 1874, 1363, 886, 121, 48, 1692, 1965]. density-independent [1692]. dentatus [188, 1015]. depend [828]. dependence [46, 1370, 770, 2144, 1172, 1897, 1999]. dependent [1023, 1587, 580, 1431, 1654, 633, 1661, 588, 1027, 976, 1152, 1616, 412, 1321, 1874, 1363, 2041, 886, 1216, 2013, 1334, 2038, 121, 48, 45, 1873, 1004, 1692, 1392, 1783, 138, 1515, 2119, 124, 737, 1965, 240, 790, 1613, 2065, 408, 23, 1488]. depensation [261]. depleted [1869]. depletion [966, 1111, 1598, 366, 415, 1526, 130, 652]. deposited [1512]. deposition [458, 410, 1773, 1257, 1695, 543, 786, 652, 1742]. depositional [108, 139]. Depth [633, 1793, 1450, 240, 1705, 1246, 1214, 964, 96, 1161]. Depth-dependent [633, 240]. Depth-integrated [1793]. Depth-specific [1450]. depuration [530, 1353, 1883]. Derivation [719]. derived [1540, 604, 1847, 1145, 1903, 693, 1872, 2006, 2072]. Deriving [1002]. Dermochelys [1210]. describe [2153]. description [1862]. design [586, 1119, 875, 1893, 986, 1483, 1412, 1815, 1702, 1462, 2086]. design- [1483, 1412]. design-based [1462]. designations [2053]. designed [434]. Designing [280, 745]. designs [1747]. despite [1959]. despotic [1763]. detect [2027, 313, 1653]. detectable [97, 1109]. detected [3]. Detecting [2025, 454, 858, 789, 1239, 1388, 2014, 554, 19, 693]. Detection [292, 1524, 1753, 605]. deteriorating [1166]. determinant [620]. determinants [1593, 347, 766, 1405, 402, 1302]. Determination [524, 1057, 354, 539, 1970, 1344, 1326]. determine [1235, 1427, 499, 2030]. determined [1047, 1251, 1507, 1506, 1008]. Determining [822, 115, 1620, 333, 956, 673, 520, 2106, 665, 1137, 561, 1198]. deterministic [1475]. deterrence [1469]. detrital [1481, 300]. Detritivorous [116, 113]. detritus [1085, 1725, 545]. detritus-based [1085]. Developing [396, 397, 1847]. Development [2019, 2147, 208, 1348, 1739, 179, 1382, 1834, 1962, 1958, 814, 1602, 1281, 875, 1716, 232, 1037, 198, 1548, 498, 1034]. Developmental [1102, 733]. developments [1274]. devices [1833, 2143, 1699]. Devon [2052]. dewatering [881]. DFADs [1699]. diadromous [2122]. diadromy [684]. Diagnostic [455]. Diagnostics [414, 307]. diapausing [1269]. Diary [1710]. Diatom [306, 1500, 371, 1476, 587, 1048, 696, 1009, 550, 1883]. diatom-based [1476, 587, 1048]. Diatom-inferred [306]. diatoms [529, 34]. Diazinon [192]. dichotomy [250]. Didymosphenia [1716]. Diel [236, 1518, 1519, 1826, 1865, 291, 2004, 1829, 1341, 2063, 1646, 1554, 325, 1781, 2012, 193]. diel-cycling [2004]. Diet [1211, 438, 1843, 268, 1645, 1970, 699, 1470, 1160, 750, 662, 1217, 519, 1389, 1070, 1870, 1015]. dietary [688, 570, 171, 15, 284, 1093, 1353, 1525]. diets [93, 1284]. differ [1949, 1242]. Differences [2108, 1981, 416, 885, 36, 857, 1964, 1468, 101, 1414, 724, 373, 2133, 943, 504, 1102, 223, 1575, 1017, 72, 779]. different [2107, 140, 440, 673, 1647, 1388, 233, 1152, 1616, 582, 1381, 1183, 2143, 922, 1353, 564, 1182, 1974, 1738]. Differential [827, 658, 775, 727, 1117, 1688, 644, 118, 482, 97, 801]. Differentiating [2156]. differentiation [1283, 2097, 1194, 65, 838, 939]. differing [806, 2132]. Difficulties [759]. difficulty [955]. Diffusion [577, 2026]. Digestion [761, 93, 1492]. Digestive [113, 253, 1643]. Dikerogammarus [599]. dilemma [134]. dimensional [337, 829, 823, 2151, 1589]. dimensions [1057]. dimethylsulfide [905, 907, 902, 900, 901, 898, 909]. Dimethylsulfoniopropionate [902, 897, 896, 901, 898, 899]. dimethylsulfoxide [899]. dimethylsulphide [903, 908]. Diminished [888]. dimorphism [810]. dinoflagellate [898, 541]. Diomedea [1186]. dioxide [992]. Diplodus [131]. diploid [992]. Diporeia [1393, 2084]. Diptera [920]. Direct [57, 461, 908, 1906, 859, 1556, 1724, 2007]. direction [531]. directly [847]. disappearance [1393]. discard [1003, 668]. discards [2141, 1078]. Discharge [1004, 1542, 1993, 1756, 168, 1614, 698, 29, 480]. Discharge-dependent [1004]. discontinuities [1538]. discriminant [1990]. discriminating [1830]. Discrimination [1824]. Disease [1591, 1560, 799]. Disentangling [1521, 532]. disk [824]. Dispersal [1891, 610, 2150, 251, 626, 820, 1984, 766, 1137, 2104, 121, 1921, 1943, 986, 1030]. dispersing [1540]. Dispersion [479, 1557, 187, 2092, 265, 504]. displaced [1325]. displacement [2004, 697]. disrupter [1350]. Disruption [1147]. disrupts [192]. dissolved [1579, 1599, 1876, 2100, 306, 1802, 960, 1089, 1801, 502, 1100, 91, 653, 1053, 742, 2140, 1938, 1678, 812, 1150, 1807]. Dissostichus [1312]. distance [1984, 1822, 718, 501, 1367, 1478, 1231, 1901, 2027]. distant [1112]. distinct [726]. distinctiveness [2014]. distinguish [1275, 1822]. Distinguishing [1571, 645]. distributed [1874, 1430]. Distribution [1474, 529, 1182, 66, 1323, 110, 1706, 1549, 341, 632, 437, 831, 798, 592, 1910, 1979, 1562, 500, 231, 1177, 1762, 758, 1297, 2010, 1736, 551, 967, 2034, 1718, 528, 1499, 1229, 1044, 1430, 778, 890, 2123, 1038, 299, 1893, 1116, 898, 265, 1693, 1763, 797, 77, 2049, 710, 220, 921, 768, 1161, 933, 2145, 1234]. Distributions [756, 787, 1022, 897, 1342, 1482, 1758, 320, 1242, 364, 1528, 575]. districts [1272, 1205]. disturbance [1593, 741, 1958, 1530, 1381, 1851, 1794, 748, 753, 358, 2081, 723, 1975]. disturbances [960]. Diurnal [497, 1029, 898]. dive [419]. diver [1893]. Divergence [221, 1211, 1437, 1603]. Divergent [1148, 1907]. diversifying [1886]. diversity [95, 1645, 634, 1568, 1756, 412, 1442, 781, 1538, 1901, 1541, 178, 2076, 726, 295, 1323, 879, 201, 69]. diving [1487]. DMS [893, 905, 904, 901, 895]. DMSP [897, 896, 901]. DNA [95, 1588, 1184, 524, 1711, 214, 1247, 1844, 1251, 931, 1224, 2062, 184, 1253, 1416, 939, 479]. Do [499, 1702, 1884, 1260, 1706, 1445, 1029, 1072, 1634]. DOC [2100]. Does [1705, 1292, 285, 512, 1655, 1823, 2045, 511, 1330, 828, 1281, 2044, 545, 1447, 1631, 102, 497, 1327, 689]. dogfish [1999]. Dolly [443, 295]. dolomieu [2150, 1665, 1211, 1135, 1989, 289, 1771, 1294, 1923]. domains [1313]. domestic [634]. domesticated [210, 1036, 330, 2061, 1238]. domestication [621, 1639, 1638, 1238]. dominance [64, 431, 1638]. dominant [1926]. dominated [376, 1097, 480]. domoic [1883]. Doppler [1554]. dormancy [1876]. Dorosoma [1725, 113]. Dorsal [1740, 1372]. Dorsal-fin [1740, 1372]. Dorset [1799, 1803, 1800, 1807]. double [1986, 1142, 80]. double-crested [1986]. down [1454, 1610, 568, 262, 1187]. Downstream [862, 210, 2147, 545, 1289, 480, 1063]. downwelling [1313]. drag [841, 1462]. drainage [205, 39, 1529, 1004, 426, 691, 480, 313, 1774]. draining [1802]. drawdown [1148]. Drawing [2079]. dredging [1241]. Dreissena [341, 562, 818, 1455, 1401, 60, 389, 1229, 1946, 1769, 101, 65, 75, 485, 811, 876, 165, 697, 916, 370, 1109]. dreissenid [1413, 1160, 1394, 697]. dreissenid-colonized [1413]. dreissenids [948]. Drift [249, 1565, 298, 250, 823, 436, 1753, 404]. drift-feeding [298, 823, 436, 1753]. drift-retention [250]. Drifters [857]. drifting [1699]. driven [979, 516, 29]. drivers [1808, 792]. drives [1863, 1610]. driving [2081]. Drought [1805, 1809]. Drought-induced [1805]. drum [15, 560]. DSc [686]. dual [22]. dubia [480]. due [885, 114, 1045, 771, 2145]. Dungeness [1019]. duration [461]. during [236, 321, 385, 1667, 903, 992, 1578, 1690, 1941, 1268, 1448, 1193, 1592, 1115, 841, 1209, 1281, 284, 1473, 1327, 853, 190, 400, 1558, 34, 358, 1060, 1440, 945, 1570, 1195, 199, 1294, 1854, 87, 10, 2131, 1069]. Dutch [1736]. dwarf [294]. dwelling [416, 1530, 1279]. Dynamic [683, 736, 37, 1713, 1007, 1937]. dynamically [1669]. Dynamics [281, 567, 1897, 742, 1666, 1127, 1455, 1906, 1463, 495, 659, 660, 643, 1129, 770, 1409, 1296, 1081, 766, 52, 1641, 352, 1726, 2142, 944, 432, 988, 1857, 81, 1357, 239, 896, 1632, 396, 1539, 163, 187, 1591, 1558, 1446, 508, 1004, 1279, 1394, 1995, 898, 731, 1420, 1324, 2071, 392, 1477, 955, 980, 124, 282, 796, 762, 547, 797, 970, 1808, 123, 1280, 516, 895, 1781, 2012, 1930, 2062, 692, 421, 1484, 1872, 2065, 927, 1945, 792, 61].

each [407]. ear [1695]. Early [1267, 872, 724, 892, 384, 1510, 1587, 1268, 1190, 649, 1752, 1491, 1433, 995, 1410]. East [1204, 1785]. eastern [1455, 716, 1191, 1190, 1665, 1413, 1962, 1296, 1081, 403, 1175, 1606, 1311, 1222, 1685, 1476, 242, 2157, 1969, 1898, 522, 525, 110, 792, 1972]. echo [1149]. echolocation [548]. Ecological [910, 2043, 2089, 1468, 510, 1733, 1681, 394, 1864, 940, 1647, 1552, 1907, 1812, 1975, 1136, 2073, 610]. ecologically [726]. Ecology [1188, 388, 630, 1071, 1170, 619]. economic [1791, 2043, 651]. economics [568, 1995]. Ecopath [1621]. Ecophysiological [1302]. Ecoregion [732, 1751]. ecoregions [2066]. Ecosystem [264, 1537, 1650, 909, 2139, 1190, 877, 659, 660, 546, 1610, 412, 1896, 1552, 246, 323, 948, 1383, 1216, 2013, 355, 631, 2068, 731, 1968, 1703, 1955, 1074, 1759, 1026, 137, 350, 1926, 1398, 2131, 792, 1110]. Ecosystem-based [1650, 1896, 1216, 2013, 1955, 1398]. Ecosystem-level [2139]. écosystèmes [1799]. ecosystems [1799, 268, 202, 365, 4, 744, 1679, 973, 1707, 1812, 1263, 1800]. ecotone [1339]. Ecotypic [1194]. edge [449, 1044]. edges [187]. Editorial [256, 1730, 256, 1730]. Editorial/Éditorial [1730]. edulis [1175, 1239]. edwardsii [488, 487, 1196]. Edwin [686]. eel [758, 1577, 166, 1327, 1447, 870]. eels [825, 684]. Effect [411, 345, 1910, 1054, 971, 1382, 117, 1816, 58, 432, 922, 1725, 1853, 1507, 144, 1934, 1127, 1066, 1320, 719, 1390, 53, 856, 791, 808, 728, 1180, 1536, 814, 526, 2055, 1521, 508, 1364, 422, 1695, 1782, 1965, 1625, 1109]. Effective [237, 1984, 1385, 1815]. Effectiveness [1331, 764]. Effects [1436, 309, 530, 716, 596, 207, 735, 1377, 846, 12, 1829, 1020, 271, 1594, 1362, 1994, 621, 882, 569, 1717, 2115, 2095, 960, 528, 1722, 1792, 998, 572, 964, 1095, 867, 1671, 1140, 1329, 79, 355, 2048, 1353, 604, 1446, 1314, 1315, 1691, 152, 706, 1478, 848, 1352, 360, 361, 1384, 1231, 1290, 49, 703, 945, 723, 618, 303, 2135, 139, 123, 377, 1496, 18, 94, 1294, 916, 293, 209, 1872, 445, 1062, 456, 1088, 535, 589, 1645, 383, 1906, 140, 1865, 615, 394, 1533, 2004, 622, 1401, 196, 180, 264, 975, 1654, 520, 1582]. effects [1668, 891, 660, 1177, 1474, 827, 482, 203, 424, 873, 410, 665, 878, 744, 35, 52, 1639, 601, 1556, 2008, 1602, 1537, 1560, 1479, 1957, 363, 2128, 393, 2087, 785, 154, 1604, 1585, 1216, 2013, 1332, 1333, 1842, 915, 1257, 1357, 679, 1989, 1685, 502, 398, 1521, 937, 1643, 163, 1033, 1309, 197, 1427, 402, 2018, 642, 1126, 835, 485, 1285, 1307, 1417, 645, 5, 540, 755, 861, 1045, 811, 468, 436, 789, 1638, 1798, 1166, 1812, 762, 1838, 1429, 1636, 1263, 1458, 259, 137, 850, 461, 404, 532, 1171, 428, 549, 1999, 2151, 1819, 261, 1350]. effects [1408, 1561, 675, 1807, 1034, 89]. efficacy [1137, 1179, 1656]. Efficiencies [1546, 1907, 1099, 223]. efficiency [604, 436, 1526, 377, 2028]. effluent [315]. effluents [137, 574]. effort [1317, 288, 1994, 582, 418, 701, 1434, 1364, 1179, 331, 980, 7, 1590, 547, 784, 1198, 768, 576, 2072, 409]. efforts [2025, 1765, 817]. Egg [53, 2102, 676, 716, 251, 1876, 1860, 1257, 79, 160, 114, 1033, 1202, 1712, 1615, 1627, 1746, 2151, 1017]. egg-to-emergence [1033]. Eggs [1203, 474, 179, 357, 1612, 1201, 1034]. Eggs-per-recruit [1203]. eight [1108, 2066]. elasmobranch [348, 2035]. elasticity [1497]. electronic [1997]. electrophoresis [879]. electroshocking [297]. eleginoides [1312]. element [1824]. Elemental [2014, 942, 1312, 1566, 561, 720]. elements [2052, 260, 125]. elevate [545]. elevated [239, 1074]. elevation [71, 1270]. elevational [300]. elimination [1347, 1674]. ELISA [1560]. ELISA-based [1560]. Ellesmere [1891]. Elliptio [112, 1504]. ellobiopsid [157]. elongatus [2019, 930]. elucidates [1162]. Elucidating [1321]. Emamectin [602]. embayment [1746]. embryo [1446, 1034]. embryonic [1281, 1422]. embryos [808, 1913, 1440, 1636]. emergence [411, 345, 1876, 1033, 1828]. emergent [943]. emerging [1221, 1716, 610]. EMG [666]. emissions [907, 1866]. emphasis [1620, 838]. Empirical [1243, 391, 1395, 1103, 161, 638, 1434, 415, 147]. empty [802]. EMS [1491]. enclosure [1512]. encrasicolus [1612, 1937]. encroachments [1521]. end [1246, 498]. end-of-season [1246]. endangered [589, 1765, 1385, 2116, 1933, 1375, 789, 918, 1001, 2032]. endemic [1468]. endocrine [1075, 1350, 199]. endocrinology [621]. endolymph [945]. ends [1902, 1672]. endurance [122]. Energetic [82, 666, 1070, 624, 327, 311, 493, 519, 294, 1625, 1410]. energetics [1784, 1604, 697]. Energy [1041, 733, 872, 1440, 771, 644, 151, 772, 2067, 833, 247, 1499, 1911, 1334, 658, 82, 1394, 1002, 1099, 2050, 723, 1551, 1187]. energy-conserving [247]. Energy-related [1041]. engineering [948]. England [285, 512, 511, 701, 1760, 1037, 1008]. Engraulis [1612, 583, 1937]. enhanced [447, 1920, 287]. enhancement [1067]. enhances [807, 2111, 1813, 1899]. Enhancing [655]. Enriched [1754, 64, 458, 575]. enrichment [650, 140, 569, 1237, 1358, 154, 339, 1514, 1117, 1796, 1652, 652]. ENSO [792]. ENSO-scale [792]. entering [1277]. entire [1026]. entry [769, 1561, 870]. environment [312, 846, 191, 529, 1147, 1418, 1049, 1576, 1524, 1038, 299, 845, 1037, 1005, 1014, 1419, 1009, 1472, 1914, 517]. Environmental [893, 1799, 465, 1255, 24, 865, 1569, 1796, 671, 831, 798, 1906, 1559, 1174, 241, 454, 367, 665, 839, 274, 1711, 569, 1586, 1835, 1764, 397, 642, 1846, 977, 1828, 892, 1072, 618, 721, 1528, 2049, 1891, 1323, 1439, 629]. environmentally [1216, 2013]. environments [711, 383, 64, 586, 1949, 1375, 1656]. environnementale [893]. environnementales [1799]. environs [249]. enzymatic [519]. enzyme [1490, 1643, 400, 113]. enzymes [1402, 943, 1721, 1315]. eos [1900]. ephemeral [1864]. epilimnetic [1532, 1534, 456]. Epilithic [1073, 137, 961, 445]. epipelagic [87, 10]. epiphyte [2100, 1406]. Epiphytic [1972]. episodes [887]. episodic [706, 2002]. epithelia [71, 1695]. epomophora [1186]. equation [2108, 1981, 22]. equilibrium [1092, 1656, 641]. equivalence [624]. ERAs [2073]. Erie [812, 1927, 571, 1018, 1455, 337, 1059, 1413, 528, 163, 38, 1070, 697, 1603, 87, 10, 370, 32, 1980]. erinacea [1411]. Erken [228]. erosional [139]. Erratum [1359, 2109, 818, 831, 411, 1054, 1101, 1527, 89, 1668, 1421, 805, 1296, 1616, 1358, 471, 1272, 1597, 1420, 1837, 513, 88, 87, 584, 780]. error [1024, 1023, 1582, 1668, 1619, 844, 843, 1954, 578, 460, 796, 1898, 927, 2145]. Errors [428, 2138, 78, 467]. escaped [1969]. Escapement [1143, 1922, 407, 68, 166, 1749, 604, 1664, 25, 99, 538, 406]. Escaping [1113, 1637]. esfenvalerate [868]. esmarkii [1595, 2024]. Esox [1199, 40]. essential [1626, 268]. Establishing [1948, 535]. Establishment [374, 1993, 1025, 1098, 121, 75]. Esteros [725]. Estimability [243]. estimate [2077, 1301, 1107, 847, 834, 212, 1247, 1664, 1997, 1893, 1345, 1412, 1526, 1548, 2028, 1019, 736]. estimated [241, 253, 427, 252, 958, 1819]. Estimates [1785, 1257, 25, 407, 1046, 1741, 1489, 1000, 794, 1793, 231, 874, 1994, 59, 167, 1492, 1877, 941, 829, 1723, 593, 604, 398, 1776, 1516, 1403, 1968, 329, 2121, 44, 92, 428, 536, 693, 1845, 115]. Estimating [1993, 2078, 463, 1929, 2144, 488, 990, 1657, 1183, 1849, 597, 1747, 554, 504, 560, 1712, 538, 439, 627, 1613, 88, 42, 342, 1804, 563, 2145, 2019, 1885, 2005, 765, 1933, 864, 537, 1128, 1051, 1743, 1915, 406, 1735, 1077, 2162, 1589, 1648, 659]. Estimation [1666, 1278, 348, 557, 795, 1491, 1483, 1202, 2042, 26, 496, 407, 677, 109, 1082, 1580, 1441, 455, 661, 1084, 925, 51, 19, 1365, 1837, 129]. estimations [2040]. estimator [1091]. estimators [308, 507, 1517, 641]. estuaries [394, 1175]. estuarine [2077, 831, 798, 615, 1745, 143, 1162, 1481, 2098, 48, 997, 142, 1494, 1141, 1737, 2037, 1626, 2062, 390, 1109]. estuarine-opportunist [1141]. Estuary [1630, 898, 1878, 205, 1920, 116, 1601, 1698, 444, 825, 1903, 916, 588, 319]. ethynylestradiol [2133]. etiological [817]. Eubalaena [1106]. Eudontomyzon [1967]. eulachon [962]. eulatero [84]. eulatero-frontal [84]. Eumetopias [1328, 1537, 962, 1589]. Euphausia [282]. Eurasian [1153, 1935, 1408, 1972]. Europe [607, 1098]. European [1407, 1370, 758, 1577, 166, 1432, 2087, 1953, 1327, 1364, 972, 849, 1072, 1136, 605, 1937]. Eutrigla [1221]. eutrophic [202, 991, 162, 130, 14]. eutrophication [394, 1998, 85, 1917, 1600]. eutrophicaton [34]. evacuation [354]. evaluate [628, 1513, 1597, 1908, 483, 484]. evaluated [2092]. Evaluating [1947, 420, 1885, 151, 2124, 1621, 1944, 2116, 2000, 1836, 1457, 558, 1522, 1435, 2056, 1002, 280, 419]. Evaluation [1317, 1193, 2154, 1749, 1642, 889, 1859, 1993, 1426, 161, 143, 1880, 141, 158, 1182, 349, 1091, 104, 1307, 2003, 1373, 1354, 1796, 1015]. event [321, 1069]. events [1802, 1252, 1809, 1648]. Evidence [650, 1164, 1298, 2038, 1680, 986, 1673, 1553, 1063, 57, 1390, 614, 820, 873, 976, 191, 1340, 1682, 885, 1105, 1393, 816, 1570, 1763, 1078, 519, 287, 961, 687, 779, 1120, 2007, 216]. evoked [868]. evolution [268, 978, 902, 534, 1130, 1627, 684, 69, 1799]. Evolutionary [612, 583, 1104, 1131, 1672, 606]. ex [1723]. Examination [2007]. examine [367]. Examining [298, 1271, 1366, 629]. example [43, 580, 1582, 1668, 1885, 1111, 941, 121, 1126, 1116, 138, 92, 1600]. examples [307, 1171]. excessive [339]. exchange [148, 1340, 2151]. exercise [992]. exhausted [472]. exhibit [46, 97]. existence [1673]. exopolymeric [1474]. exotic [1531, 62, 215, 1064, 876, 1683]. expanding [1947]. expected [2004]. expenditure [1440, 1625]. experience [1689, 508]. experienced [482]. experiment [1655, 1351, 915, 1590, 2131]. Experimental [1869, 565, 2130, 2060, 1326, 72, 1015, 1066, 1012, 383, 2137, 1301, 1941, 1647, 2111, 233, 1321, 2132, 2128, 636, 494, 1119, 355, 1028, 1307, 417, 20, 1040, 159, 2135, 1636, 405, 1148, 2129, 687, 1640, 1065, 2125, 222, 2127, 2126]. experimentally [55]. experiments [1415, 1489, 912, 765, 154, 604, 534, 1550, 1526, 1008, 155, 14]. explain [1439]. explained [1420, 1324]. Explaining [699]. explains [1980]. explanation [157, 841]. explanations [1191, 274]. explanatory [1889]. Explicit [1694, 151, 521, 359]. Exploitation [446, 1567, 794, 2078, 1370, 1962, 957, 616, 310, 2048, 1997, 343, 1126, 1345, 23]. exploited [1127, 1118, 2107, 873, 1397]. Exploration [1083]. Exploratory [1820, 1232]. Exploring [1646, 1191]. export [2114, 960, 1803, 973, 545]. exposed [570, 1147, 315, 38, 2133, 868]. exposure [39, 1193, 284, 172, 1525, 1913, 1053, 5, 934, 1480, 1827, 44]. expression [521, 97, 1490, 315, 1495, 1035, 642, 1295, 506, 870, 1688]. Extending [1889, 796]. extension [607]. extinction [267, 1136]. extirpation [1926]. extracellular [705]. extraction [1441, 1507]. extracts [1467]. extrapolation [909]. extrema [1534]. Extreme [132, 1802]. Exxon [509].

faces [667]. facilitate [1260]. facilitation [912]. Facilitative [490]. facteurs [1799]. factor [2122, 443, 1627, 1883, 736, 1063]. Factors [2023, 915, 1042, 55, 1430, 1987, 2146, 1078, 1376, 549, 126, 122, 9, 341, 1799, 831, 798, 1865, 743, 1610, 274, 465, 1789, 1508, 270, 24, 1229, 1182, 865, 1558, 1692, 5, 1139, 77, 1380, 1839, 1800]. failed [341]. Falkland [43, 248, 925]. fall [1106, 2156, 1823, 2045, 168, 167, 1396, 2044, 484]. fall-spawning [1396]. Fallopia [1679]. familiarity [384]. families [1295, 479]. family [1677, 1239]. family-correlated [1677]. fantasy [819]. Farfantepenaeus [1177]. farm [2135]. farmed [1593, 1977, 1036, 1969, 939, 1163]. farms [1469]. fasciatus [186]. fasting [1658]. fate [903, 1448, 1894, 11]. fathead [569]. Fatty [981, 228, 1558, 801, 1256, 1282, 1359, 268, 688, 581, 1602, 883]. fauna [446, 334]. faunas [838]. features [1635, 246, 1850, 933]. Fecundity [2149, 716, 2144, 1618, 676, 13, 678, 1344]. federally [1001]. Federation [194]. feed [1643]. feedback [1998, 955]. feeder [1076]. feeders [1598]. feeding [562, 818, 530, 447, 807, 89, 12, 298, 1792, 823, 1585, 853, 649, 1071, 1576, 1643, 1096, 1494, 436, 1907, 1219, 771, 1753, 1185, 516, 854]. female [1299, 53, 1296, 1081, 79, 1392, 1631, 26]. females [678]. ferruginea [2078]. Fertilization [508, 1502, 1536, 625, 456]. fidelity [1371, 985, 382, 1900, 704]. Field [135, 1548, 2019, 35, 1175, 912, 1366, 1078, 868, 1796, 125, 1361, 1021]. field-based [2019]. fifth [891]. fifth-order [891]. fill [96]. filter [1620, 1887, 1076, 760]. filters [1894]. filtration [370]. fimbria [129, 854]. fin [1372, 1740]. Finding [784]. Fine [1328, 1446, 1303, 1903, 1781, 2012]. Fine-scale [1328, 1903]. finfish [412]. fingerling [160]. fingerprints [1566]. Finland [852]. finmarchicus [53, 320, 319]. fire [1460]. Fireproof [1337]. First [1486, 210, 807, 820, 521, 358, 2127]. first-feeding [807]. first-generation [210]. first-order [2127]. First-year [1486]. Fish [433, 257, 830, 1101, 820, 1351, 757, 1699, 713, 1195, 1348, 918, 1408, 1666, 1706, 1593, 910, 1066, 1230, 535, 761, 1312, 632, 2077, 1023, 1713, 1424, 657, 1118, 1645, 1765, 1906, 474, 1559, 677, 140, 447, 275, 2051, 1811, 333, 1745, 1174, 1382, 440, 2122, 581, 1531, 260, 420, 1863, 189, 454, 520, 783, 743, 1970, 457, 1848, 116, 1082, 1025, 1370, 668, 827, 1193, 1188, 424, 1833, 873, 1388, 1204, 991, 524, 665, 1162, 2147, 453, 1441, 1362, 1867, 1630, 233, 728, 1180, 1556, 246, 772, 59, 996, 1840, 2010, 984, 1442, 378, 840, 1271, 2005, 1851]. fish [1505, 1944, 1911, 1663, 270, 315, 864, 764, 1220, 1335, 1764, 1543, 1685, 229, 585, 624, 1539, 398, 1643, 1309, 234, 890, 2123, 1942, 238, 1919, 1347, 289, 1367, 104, 1126, 1080, 1516, 1758, 597, 1669, 426, 2157, 1646, 147, 20, 695, 1417, 98, 119, 1649, 1507, 392, 1596, 993, 1244, 507, 1477, 1219, 2133, 1512, 460, 537, 1141, 1763, 169, 1128, 1242, 1261, 1850, 1051, 1743, 2073, 683, 1139, 139, 1208, 1378, 784, 1832, 123, 1458, 1511, 1626, 178, 519, 869, 531, 1198, 1746, 18, 577, 1894, 1774, 31, 1930, 1836, 916, 1302, 1926, 126, 838]. fish [1870, 193, 125, 1613, 2085, 2086, 1821, 693, 1149, 88, 42, 294, 1484, 342, 1653, 1945, 732, 2075, 933, 1839, 106, 1751, 2011, 563, 777, 1110, 1681, 752, 2145, 575, 409]. Fish-habitat [918]. fish-injury [1193]. Fish-mediated [433]. fish-stocking [238]. fished [366]. fisher [1699]. Fisheries [950, 1134, 43, 1158, 1947, 1122, 446, 1475, 1273, 447, 953, 1463, 952, 1696, 230, 2079, 940, 758, 1328, 170, 274, 1379, 1657, 1694, 1505, 1555, 1857, 81, 430, 269, 1513, 1597, 2116, 1357, 679, 1167, 929, 1140, 17, 631, 1434, 1776, 1650, 1002, 1868, 280, 1943, 1345, 1814, 637, 1151, 235, 554, 1515, 1121, 1990, 1798, 1898, 721, 1884, 1925, 1123, 951, 1331, 525, 1186, 1138, 736, 1228]. Fisheries-induced [1134]. fisheries-related [1798]. fisheries-stock [230]. fishermen [1463]. fishers [1706, 866]. fishery [43, 2155, 37, 1714, 1052, 1855, 735, 288, 628, 1124, 2079, 1762, 1111, 855, 1388, 1936, 839, 1456, 1896, 2142, 582, 1340, 1445, 2033, 2087, 266, 1466, 1849, 2048, 875, 2068, 1203, 980, 503, 707, 1365, 1837, 1898, 1097, 702, 1959, 784, 258, 419, 377, 1395, 768, 1416, 576, 1937, 2103, 780, 751, 1724, 409, 595]. fishery-independent [1849, 1724]. fishes [372, 2019, 1645, 86, 1259, 622, 1605, 1756, 1144, 1129, 344, 1428, 78, 1994, 348, 1363, 1657, 1586, 851, 1530, 978, 974, 505, 2095, 521, 1813, 267, 740, 362, 152, 2053, 2020, 215, 1384, 1116, 2035, 1072, 713, 1136, 13, 667, 696, 135, 619, 1982, 1824, 2062, 134, 1323, 1422, 1497, 1161, 115, 1472, 656]. Fishing [365, 1418, 1988, 262, 1014, 1419, 1338, 1146, 1666, 146, 236, 385, 309, 37, 2077, 1317, 1052, 1533, 2, 1113, 836, 1377, 794, 660, 855, 856, 472, 488, 1522, 582, 990, 1537, 1435, 568, 1722, 1813, 1680, 1671, 1068, 1434, 1364, 1179, 773, 235, 554, 1990, 7, 2040, 1590, 1782, 1130, 694, 710, 1631, 869, 1077, 2162, 1398, 693, 576, 1937, 261, 553, 1161, 409]. fit [17]. Fitness [311, 1092, 1119, 384]. fitting [1582, 1668, 590, 6]. five [440, 375, 1037]. fixation [2083]. fjord [1984, 1558, 1494, 1968]. fjords [1009]. flame [1337]. flame-retardant [1337]. flat [1737]. flatfish [875, 242, 525, 576]. flavescens [711, 1300, 1579, 982, 1298, 1321, 1322, 984, 1108, 253, 1711, 1892, 1147, 393, 1071, 1366, 1871, 172, 1691, 75, 485, 917, 1707, 1674, 618, 44, 1253, 1859, 1021]. flea [475]. Fleet [1463, 547, 1736, 1995, 1937]. fleets [616, 1364]. Flexible [1160, 1466, 590]. flies [1499]. floating [1271]. flood [1195, 623, 408]. flood-frequency [623]. flooded [2132]. floodgates [1529]. flooding [757]. Floodplain [287, 477, 820, 401, 833, 1608, 1529, 379, 111, 1681]. floodplains [757]. floods [114]. Florida [161, 1675, 1682, 594, 1305, 2121]. flounder [2078, 463, 239, 188, 1760, 790, 1201, 1015]. flourish [1634]. flow [1549, 1415, 2160, 74, 558, 1864, 1193, 1156, 2147, 351, 443, 974, 833, 1789, 2093, 1685, 265, 1913, 1316, 726, 1985, 2161]. flow-related [1789]. flowing [1715]. flowpaths [1864]. flows [1633, 1026, 1781, 2012]. fluctuating [219]. fluctuations [1658, 1495, 559]. fluid [366]. flume [1185]. fluminea [313]. fluorescence [1570]. fluvial [1711, 2091, 1289, 1195, 1253]. flux [601, 1045, 932]. fluxes [2111, 972, 252, 1423, 1697, 353, 1110, 908]. follicle [315]. following [1018, 1531, 1567, 1520, 887, 1309, 358, 691, 1480, 475]. Folly [819]. fontinalis [1165, 1299, 1339, 153, 1404, 1635, 1916, 160, 837, 730, 1126, 724, 1494, 1786, 826, 1780, 712, 1325, 1623, 149, 1551, 2065, 722]. Food [1204, 154, 1250, 1535, 1293, 876, 1208, 1256, 715, 1046, 1088, 1971, 118, 1236, 1820, 544, 1292, 596, 128, 1164, 614, 1902, 981, 1762, 1370, 1027, 1896, 183, 228, 1602, 568, 141, 1608, 1851, 469, 1604, 470, 272, 1085, 647, 1543, 1366, 1368, 1988, 567, 1897, 972, 731, 33, 1393, 811, 1924, 49, 262, 646, 963, 1961, 223, 1051, 2148, 493, 1125, 1450, 2028, 1652, 88, 42, 2006, 115, 2101, 809]. food-unlimited [544]. Food-web [1535, 972]. foothills [1758]. Forage [1870, 264, 996]. Foraging [1322, 1264, 1039, 1235, 1186, 298, 750, 1509, 485, 1231, 1170, 206, 619, 2037, 1280, 1721, 1796]. forced [312]. forces [2081]. forcing [323, 679, 775, 1738, 792, 1534]. Forecasting [1357, 1153, 2000, 1662]. forecasts [1272, 1205]. Forest [296, 732, 200, 2134, 1057, 1520, 691, 1274, 70, 1263, 405, 1148, 786]. forested [2114, 1360, 867]. forestry [1661]. forests [1540, 1105, 545, 522]. formation [636, 654, 2035, 945]. formed [290, 1974]. former [194]. formerly [2130]. formosa [1206]. forms [211, 726]. formulation [925]. formulations [2108, 1981]. forward [247, 1777]. forward-assisted [247]. fossil [993]. four [2132, 2133, 860]. fraction [1970]. fractionation [1326]. Fragilaria [1206]. fragmentation [1890, 1759]. fragmented [1375]. framework [257, 331, 1814, 1955, 1861]. France [352]. franciscanus [100]. Francisco [1630]. Fraser [2160, 1985, 2161, 1448, 226, 1555, 1766, 387, 386, 99, 666, 1245, 41, 201]. Frazer [211]. free [1706, 1271, 768]. free-floating [1271]. frequencies [598, 1693, 1966]. Frequency [1581, 783, 859, 1632, 1028, 280, 1905, 1202, 623]. frequentist [507]. fresh [523]. Freshwater [527, 50, 2032, 446, 413, 1173, 112, 1320, 118, 68, 207, 74, 2122, 117, 1164, 743, 1829, 1275, 599, 1162, 529, 878, 1552, 1442, 1773, 469, 186, 594, 470, 272, 270, 1405, 1933, 1964, 2136, 338, 865, 2053, 977, 426, 973, 1037, 369, 1763, 1136, 696, 1378, 1918, 29, 1891, 838, 1637, 1497, 1686, 870, 1485]. freshwaters [259]. Front [282]. frontal [84, 662]. frozen [378]. fry [1865, 857, 1757, 368, 690, 1035, 959, 943, 1185, 1796, 479]. Fuca [1928, 305]. fucensis [1928]. fullness [802]. fully [1312]. fulmar [2052]. Fulmarus [2052]. fulvescens [1787]. function [1820, 57, 1174, 1582, 1668, 453, 2018, 1703, 635, 2131]. Functional [401, 643, 750, 564, 773, 257, 56, 1027, 708, 432, 1493, 17, 417, 436, 619]. functioning [1190]. functions [500, 1481, 987, 587, 1048]. Fundulus [2041, 390]. Fundy [1904, 1182]. fungal [1010]. Funka [1971]. further [1338]. future [153, 1007, 887, 1804]. fuzzy [189, 454, 278].

Gårdsjön [1209]. gadids [246]. gadoid [1046, 844]. gadoids [1271]. Gadus [236, 1158, 1282, 1359, 1620, 1407, 544, 1234, 179, 1054, 971, 1191, 1489, 846, 633, 1233, 670, 1984, 1572, 1573, 1929, 83, 1949, 1962, 699, 1152, 1616, 403, 2067, 1717, 1479, 672, 988, 923, 922, 81, 1133, 1585, 864, 802, 1334, 2024, 676, 1576, 82, 103, 1314, 1315, 1942, 1403, 788, 2157, 1231, 1116, 1254, 1693, 138, 1132, 1775, 816, 1550, 2117, 771, 1931, 223, 1884, 1770, 66, 970, 1583, 1843, 2049, 1827, 1338, 1344, 2105, 1720, 240, 1926, 122, 655]. gag [2124, 16]. gahi [248, 942, 925]. gains [1417]. Galaxias [74, 498]. galaxiid [215]. Galeorhinus [22]. galeus [22]. galloprovincialis [353]. games [1332, 1333]. gamma [308]. Gammaracanthus [464]. Gammarus [186, 605]. gannets [555]. gap [918]. Gasterosteus [1249, 1704, 995, 1640]. gauging [1741]. GBT [896]. gear [847, 1849, 875, 890, 1526, 547, 129, 1077, 2162, 1161]. gears [519]. geminata [1716]. gender [16]. Gene [2160, 1985, 2161, 443, 1437, 97, 1295, 506, 726]. genera [1663]. general [2141, 331, 1740]. generalities [762]. generalization [721]. generalizations [67]. Generalized [1058, 236, 241, 844, 843, 2042, 692, 641, 22]. generating [504]. generation [210, 1677, 1883]. generations [1285]. generic [2028]. genes [1688]. Genetic [211, 1283, 1984, 1977, 1240, 214, 1560, 611, 1432, 1835, 1212, 1841, 1764, 929, 2084, 40, 1603, 687, 2001, 1858, 95, 1745, 1197, 1568, 189, 2124, 1935, 2097, 1442, 1223, 1880, 1385, 1479, 2104, 1964, 1644, 426, 1131, 1901, 1550, 1369, 1767, 1890, 1528, 1224, 2076, 1286, 523, 726, 295, 80, 1553, 939, 2031, 879, 1001, 201]. genetically [471, 335]. genetics [1199, 863, 65, 2120, 1238, 173]. Genome [978]. Genome-size [978]. genotype [1]. genotypes [589]. Genotypic [186]. genus [1606, 2018, 1258, 69]. geochemistry [1952]. geoduck [1948]. Geographic [146, 2085, 1328, 382, 777]. Geographical [607, 1629]. geography [700, 683]. geolocate [1620]. Geolocation [1931]. Geological [606]. geology [527]. Geometric [1663]. geometry [630, 1633]. Geomorphology [148, 1665]. georeferenced [2008]. Georges [2045, 2044, 581, 1823, 246, 171, 133, 1395, 342]. Georgia [2106, 1284]. geostatistic [829]. geostatistical [677, 1833, 718]. geostatistics [1116]. gestion [1122]. getting [955]. Ghost [1960]. giant [902]. gibbosus [36]. Gibbs [1369]. gigas [84]. gilae [1001]. gill [210, 847, 1698, 71, 159]. Gillnet [174, 489, 434, 1022, 1657, 51, 560]. gills [888, 1262]. gilthead [1658]. GIS [527, 777]. given [651]. gizzard [1725, 113]. glacial [459, 379, 221, 464]. glacialis [1106, 2052]. glacially [2070]. glaciated [69]. Glacier [232]. glass [870]. glauca [1377]. Global [677, 212, 905, 907, 909, 792]. global-warming [792]. globe [895]. GLOBEC [318]. glomerata [2046]. glucose [1670]. glycine [896]. glycolytic [943, 1721]. GnRHa [749]. goal [1922]. goals [68, 1555, 99]. gobies [1059]. gobio [1370]. goby [452]. golden [1754]. gonad [53, 1308]. gonadal [1783]. gonads [16]. gorbuscha [1420, 1324, 845, 395, 989, 666, 651, 1050, 209, 1625]. gorgonian [1983]. governs [1719]. gradient [1645, 1243, 1108, 1255, 1481, 1886, 1430, 300, 1307, 748, 232, 1486, 1217, 1195, 623, 717, 1652, 722]. gradients [1498, 151, 529, 1789, 359, 1033, 1121, 1325, 1891, 2075, 629]. Grand [355, 223, 1487]. grandparentage [423]. gravel [1016, 1527, 1975]. gravels [1303]. Gravenhurst [34]. gray [1759]. grayling [1006, 1432, 987, 849, 436, 2120, 1286]. Grazer [912]. Grazing [282, 1493, 900]. Great [1779, 754, 948, 606, 1778, 1348, 1683, 610, 1750, 933, 430, 146, 1173, 2150, 2051, 1611, 1277, 62, 191, 1362, 1694, 1830, 1940, 521, 186, 1257, 1357, 402, 613, 1514, 748, 2050, 490, 2094, 1733, 2085, 2086, 1361, 1912, 1882, 1751]. greater [1680]. greatly [1734]. green [1436, 2098, 290]. Greenland [1459, 2142, 1644, 662, 1821]. greenlip [855, 856]. Grey [1221, 1233, 1492, 26]. grid [1113]. griseus [1759]. gross [458, 1414]. ground [473, 516]. groundfish [1317, 285, 512, 1190, 171, 511, 701, 935, 679, 245, 745]. grounds [1068, 1166]. groundwater [1339, 1756, 674, 1034]. group [1717]. grouping [784]. groups [781, 733, 862, 1603]. grow [1128]. growing [1505, 1596]. grown [64, 1074]. grows [2046]. Growth [120, 1508, 175, 1265, 1770, 1821, 1249, 1114, 36, 1278, 1971, 562, 818, 688, 1713, 1587, 411, 345, 682, 64, 1300, 544, 1259, 596, 74, 1431, 196, 807, 1567, 1582, 1668, 1370, 1573, 463, 1153, 1135, 1949, 453, 976, 859, 1619, 1441, 281, 621, 1152, 1616, 2082, 1411, 1180, 1039, 1711, 1602, 1586, 708, 1213, 1622, 1951, 2033, 1115, 336, 1787, 521, 226, 1606, 97, 593, 672, 2009, 1911, 923, 1490, 1585, 881, 1933, 1281, 435, 239, 1725, 1889, 55, 502, 1853, 1576, 1624, 160, 1182, 1396, 590, 234, 197, 1035, 1172, 625, 1285, 360, 1525, 100, 422]. growth [534, 861, 1862, 468, 1037, 1992, 2035, 1924, 1856, 816, 2117, 1040, 945, 646, 1486, 460, 723, 537, 860, 223, 2061, 2146, 1991, 559, 852, 1208, 1918, 1248, 1548, 2121, 1626, 1982, 461, 1983, 1125, 1171, 287, 2062, 1817, 293, 692, 1854, 2054, 1634, 1077, 2162, 2028, 1600, 1613, 1740, 2065, 131, 1062, 995, 1640, 1453]. Grunow [62]. gryprus [26]. grypus [1233, 1492]. Guay [448]. guidelines [719, 1749]. Guinea [340]. Gulf [252, 1878, 1110, 2155, 372, 756, 581, 1605, 1566, 628, 633, 1233, 1629, 1177, 1896, 996, 1340, 76, 1182, 2000, 324, 645, 223, 852, 1316, 2049, 1026, 1926, 734]. gummy [22]. gunnari [1284]. guppies [1130]. gurnadus [1221]. gurnard [1221]. gut [354, 438, 802, 284, 113]. Gymnocephalus [1153]. gyre [2034].

h [236, 620]. Habitat [291, 2008, 183, 1115, 1899, 215, 77, 1667, 50, 1457, 1928, 1106, 148, 1426, 1317, 64, 592, 1705, 1299, 257, 1339, 1739, 1301, 1382, 1816, 891, 1665, 1635, 774, 588, 1036, 1936, 1630, 1994, 1522, 167, 1874, 814, 1039, 375, 840, 208, 815, 834, 639, 1461, 1449, 1852, 1225, 359, 828, 764, 987, 1357, 647, 1089, 1685, 1538, 1576, 1058, 858, 1676, 289, 518, 485, 2157, 959, 1417, 1452, 1244, 849, 1786, 1609, 2016, 789, 826, 1166, 2140, 1072, 356, 713, 1242, 1850, 721, 1743, 702, 311, 1208, 1378, 1633, 1458, 1626, 94, 1774, 1956, 2151]. habitat [2085, 2086, 1895, 918, 1934, 23]. Habitat- [291, 215]. Habitat-specific [1115, 1852]. Habitats [1031, 2047, 1032, 401, 558, 1381, 572, 828, 1414, 695, 1974, 545, 1141, 169, 139, 1316, 1195, 1903]. haddock [716, 251, 196, 1823, 2045, 349, 197, 2044, 1843, 1017]. haemocytes [773]. haemocytic [1239]. hake [1549, 682, 940, 616, 1723, 2087, 175, 145, 784]. half [1117]. half-life [1117]. Haliburton [1804, 486]. halibut [1644, 2092, 1821]. Halichoerus [1233, 1492, 26]. Haliotis [855, 856, 1959]. halt [1012]. Hamma [1797, 1797]. hampers [328]. handling [1734]. Hanford [168, 167]. Harbour [1239, 1173]. hard [1441, 1492, 503]. hardwater [14]. harengus [852, 1371, 1595, 517, 2115, 864, 1764, 142, 2113, 159, 2149]. Harp [475, 2007]. Harpacticoida [374]. Härsvatten [1209]. Harvest [1131, 1165, 68, 2144, 1143, 1766, 235, 955, 980, 296, 2030, 421, 1923]. harvesting [1906, 580, 1791, 200, 1520, 2087, 81, 380, 786]. hatcheries [1796]. Hatchery [1765, 1734, 1714, 64, 1797, 2036, 1757, 783, 1523, 1639, 2014, 1092, 2099, 1461, 765, 2038, 1671, 188, 885, 400, 727, 1420, 1324, 1677, 782, 1226, 1695, 1638, 1185, 1796, 779, 1355, 1021, 182]. hatchery-based [1765]. hatchery-born [765]. Hatchery-induced [1765]. hatchery-origin [2036]. hatchery-reared [1734, 2038, 1671, 188, 400, 1695, 1185, 1021, 182]. hatching [1446, 1440]. hauls [729]. Hauraki [1110]. Havre [1697]. Havre-aux-Maisons [1697]. Hawaiian [1832]. head [1384, 516]. heads [187]. headwater [1300, 1237, 1358, 880, 960, 1716, 674, 1703, 545, 522, 1451]. health [268, 934, 1859]. heart [330]. heat [1978, 1532, 1495, 642]. heat-shock [1495, 642]. heavy [1567, 44]. help [1357, 921]. hemlock [522]. hepatotoxin [1859]. herbivores [8, 1408]. herbivorous [1645]. Herbivory [2083, 383]. Heritability [219]. héritage [1122]. hermaphrodite [309]. hermaphrodites [861]. herring [1371, 509, 489, 1595, 517, 450, 2115, 382, 681, 864, 1764, 1167, 396, 142, 799, 1170, 2113, 159, 547, 852, 536, 227, 513, 548, 2149]. heteroclitus [2041, 390]. Heterodontus [1740]. Heterogeneity [1288, 477, 1197, 800, 1994, 728, 141, 1225, 964, 1535, 379, 949, 1541]. heterogeneous [550, 1817, 1684]. heterotrophy [1236]. heterozygosity [501]. Hexagenia [111]. Hg [303, 652]. Hg-polluted [303]. hidden [1931]. Hierarchical [1533, 1635, 399, 299, 663, 1743, 2121, 407, 956, 1456, 1726, 1606, 1954, 925, 924, 533, 640, 1894, 406, 1050, 563, 777]. High [481, 328, 2020, 84, 1545, 1028, 1968, 1293, 1884, 958, 722, 201, 1270, 1558]. high-elevation [1270]. high-gradient [722]. high-latitude [1968, 958]. highly [688, 1948, 409]. hilsa [813]. hippoglossoides [1644, 1821]. Hippoglossus [2092]. histocompatibility [589]. histopathological [2133]. Historic [2101, 108, 1839]. Historical [500, 1887, 1693, 1387, 394, 1155, 1962, 1682, 1521, 34, 510, 760, 406, 1232]. histories [1047, 1647, 1240, 665, 1577, 1125, 1120]. history [1122, 1498, 1587, 1118, 1292, 211, 2112, 1200, 1129, 1267, 1162, 348, 1363, 1235, 1479, 2087, 1911, 1133, 757, 1922, 1035, 1126, 788, 132, 1818, 1786, 1672, 892, 1136, 1991, 1139, 2126, 1123, 584, 552, 1934, 107, 780, 751, 1138, 1681, 1748]. hockey [67]. Hokkaido [2093]. Holland [1419]. Holothuria [879]. Holt [2028]. homarine [896]. Homarus [940, 957, 582, 2008, 76, 380, 605, 1487, 602, 2026]. home [1025, 714]. home-range [1025]. home-stream [714]. homeostasis [1093, 945, 1670]. homing [382, 1473, 473, 192]. Hommage [1465, 1400]. homogeneity [2124]. homogenization [838]. Hood [2036]. hookeri [735]. Hoplostethus [2109, 2069, 1505]. Horizontal [797, 325, 59]. hormone [97]. horse [1712]. host [1356, 1297, 1265]. hosts [1297, 939]. hot [1504]. houting [1385]. Hovvatn [1291]. Howe [205]. hoyi [1390, 692]. Hsp [730]. HSP70 [315]. HSPs [1495]. Hucho [2116]. Hudson [1779, 271, 572, 825, 916]. human [268, 1831, 1521, 259, 1859]. humic [1087, 91, 1678]. hunger [854]. Huron [1018, 983, 370, 2039, 1188, 999, 2075]. Hyalella [1173, 69]. hybrid [1254, 1355]. hybridization [1788, 821]. hybridizing [443]. hybrids [290, 2071, 1575, 1982]. Hydraulic [1633, 1016, 1527, 630, 1033, 1542, 1241]. Hydroacoustic [1080]. hydrocarbon [645]. hydrocarbons [1525, 1433]. hydrodynamic [337, 1860, 151]. Hydroelectric [498, 627]. Hydrogen [967, 1501, 1472]. hydrogeomorphic [1839]. hydrographic [246, 1901]. hydrologic [110]. Hydrological [592, 495, 753]. hydrology [1831]. hydromineral [71]. hydroperiod [1965, 18]. hydrophones [496]. Hydropsyche [39]. hydrosystem [1143, 2060]. hydrosystem-induced [2060]. hydrothermal [1928, 305]. hydroxyl [705]. hyperbenthos [1005]. hypercapnia [207]. hypereutrophic [339, 105]. Hyperoodon [1288]. hyperoxia [207]. hyperplasia [97]. hypertrophic [770]. hypolimnetic [1094, 34, 130]. Hyporheic [168, 148, 158, 1329, 379, 417, 1083, 2151, 1034]. hypothalamo [1147]. hypothalamo-pituitary-interrenal [1147]. hypotheses [510, 1646, 534]. hypothesis [1580, 336, 912, 841, 359, 1216, 2013, 28, 1634, 2007]. Hypoxia [588, 2062, 615, 2004, 1816, 1177, 1594, 1913, 1636]. Hypoxia-based [588]. Hypoxia-induced [2062].

I. [1518]. Iberá [725]. Ice [654, 481, 1246, 1459]. ice-covered [481]. icefish [1284]. Iceland [1550]. Icelandic [447, 2113, 816, 2117, 72]. ICES [159]. icon [1337]. ICPMS [1273]. Idaho [763, 1105, 1325, 185]. ideal [1706, 1763, 768]. ideally [1874]. Identification [16, 1220, 1858, 1024, 450, 598, 428, 883]. identifications [1588]. identify [341, 2091, 1064]. Identifying [1714, 657, 346, 1339, 1663, 1963]. identity [561, 1405, 832]. ignorance [235]. II [660, 1573, 1333, 1519, 397, 1315]. Iliamna [1991]. ilisha [813]. Illex [553]. Illinois [1195]. imagery [905, 553]. images [1441]. imaging [1438, 1838]. Imbalance [1256]. immune [934, 1688]. Impact [1460, 644, 487, 389, 239, 625, 1701, 353, 1276, 14, 2155, 938, 1876, 1059, 1976, 1757, 1986, 1762, 599, 1076, 352, 582, 2010, 1442, 427, 1418, 301, 302, 1427, 466, 773, 1491, 2040, 165, 694, 1014, 1419, 693, 786]. Impact-recovery [1460]. impacted [1584, 1866, 1930]. impacting [743]. Impacts [205, 1906, 230, 1214, 881, 1989, 1071, 2002, 1241, 2076, 2047, 1313, 509, 739, 628, 2080, 660, 200, 664, 805, 1388, 2134, 1522, 569, 3, 1768, 1381, 1831, 843, 900, 654, 1386, 215, 296, 1148, 610]. Impaired [1524, 1734]. impairment [2002, 1882]. impairs [1636, 1034]. imperative [1337]. impervious [1839]. implication [1772]. Implications [1477, 387, 317, 780, 751, 938, 1023, 562, 818, 1645, 1795, 89, 12, 643, 200, 1978, 1322, 561, 1835, 186, 1857, 1852, 572, 1604, 1212, 269, 679, 2021, 1950, 1543, 1632, 1529, 1569, 1776, 1403, 218, 429, 1897, 415, 986, 1304, 554, 1515, 1786, 425, 713, 547, 1074, 667, 1206, 968, 1915, 854, 1117, 293, 297, 1416, 2086, 261, 1138, 595]. Importance [965, 688, 1739, 1268, 1902, 1580, 1610, 1112, 307, 2009, 1725, 1366, 1558, 1279, 2040, 702, 1316, 1145]. important [756, 741, 1364]. imposed [1076]. impounded [222, 742, 1286]. impoundments [466]. imprecise [1621]. imprint [319]. improved [1858, 167, 600, 804, 537, 803]. improves [1845]. Improving [1364, 1462]. in-river [439]. in-season [1272, 1205, 651]. in-stream [1592, 1438, 2002]. inactivation [1209]. inadequate [1139]. inappropriate [1516]. inbreeding [589, 882]. Incentive [1379]. Incentive-based [1379]. include [796]. incomplete [427, 1369]. inconsistencies [941]. inconstans [1343]. incorporate [2121]. Incorporating [1426, 520, 1222, 1846, 1046, 367, 2078, 859, 1755, 422, 1458]. Incorporation [213, 835, 801, 397]. incorrect [1619, 1139]. Increase [860, 1445, 1851, 266, 1095, 1946, 1769, 429, 1895]. Increased [570, 1978, 1520, 97, 1689, 1760, 2061, 930]. Increases [1797, 2083, 2160, 394, 1420, 1324, 1985, 2161]. increasing [94]. increment [460]. incremental [324]. increments [1259, 260]. incubating [357, 1034]. independent [856, 1849, 890, 1692, 1651, 2131, 1724]. Index [254, 491, 685, 1055, 1318, 1563, 1727, 2158, 1348, 1339, 1476, 244, 2016, 326]. indexed [27]. indicate [2122, 1112]. indicated [1284]. indicates [1744, 2104]. indications [228, 1879, 1827]. indicator [1428, 1752, 1905, 1626, 1389]. indicators [952, 338, 730, 486, 1798, 694, 868, 1398, 693]. indices [1666, 1118, 489, 839, 814, 1689, 1002, 2003, 1702, 1819, 2072]. indigenous [984, 1871]. Indirect [811, 1906, 873, 1556, 235, 461, 1408]. Individual [1, 1607, 476, 1855, 1287, 1860, 1448, 1577, 281, 298, 2005, 681, 1466, 1449, 1852, 1305, 366, 1247, 583, 892, 537, 850, 869, 377, 2065]. Individual-based [1, 476, 1287, 1860, 298, 1852, 583, 869]. individually [1942, 704]. individuals [86, 1929]. Indonesia [409]. induced [1805, 1765, 1579, 351, 2087, 19, 2060, 1134, 1838, 989, 2062, 1150]. induces [602]. inductively [1273, 125]. industrial [1659, 1209, 1274, 543]. industrialized [1289]. infection [845, 1245]. infections [1640]. infer [997]. inference [371, 1412, 920]. inferences [2141, 1377, 1736, 467, 1750]. inferred [1106, 306, 1267, 1160, 2142, 1342, 245, 813, 931]. Inferring [1155, 947, 944, 1161]. infestation [1625, 1561]. infested [1686]. infill [1380]. Influence [715, 413, 1822, 1802, 2063, 59, 984, 1911, 1229, 1086, 1904, 2096, 150, 994, 898, 33, 1818, 1303, 646, 1687, 1423, 1697, 522, 914, 1930, 149, 421, 576, 1406, 933, 1262, 61, 1236, 1705, 1299, 241, 495, 1756, 1135, 1240, 1936, 1349, 1470, 1297, 840, 307, 2066, 1281, 802, 2024, 48, 1396, 1172, 289, 1542, 327, 1677, 1469, 782, 1043, 1901, 1924, 1609, 913, 943, 963, 1702, 1890, 1242, 1631, 2037, 1389, 734, 1600]. influenced [1016, 1527, 1644, 1495, 1913, 1274, 32]. Influences [2041, 2017, 207, 1174, 1784, 1502, 1711, 674, 1753, 356, 1528, 1583, 2085, 732, 1751]. influencing [2023, 2122, 915, 55, 1430, 865, 1987, 376, 1376, 126, 9]. inform [2068]. Information [1378, 671, 616, 1340, 147, 777]. informative [1151]. infrared [1438]. ingestion [196, 197, 33]. inhabiting [1494]. inhibited [1640]. inhibition [999, 1480]. initial [1457, 1941, 1010]. initiated [1008]. initiating [1063]. injection [749]. injury [1193]. inland [101, 1053, 968]. inner [2047, 1695]. inorganic [458, 502, 150]. inputs [715, 1088, 1621, 313]. insect [759, 781, 33]. Insecta [920]. insects [620, 1425, 1540, 1789]. inshore [489, 1383, 1944, 81]. insight [1550, 603]. insights [2008, 985]. installed [1974]. instantaneous [1548]. instream [1058]. Instructions [276, 514]. Intact [201]. intake [1551]. integrated [1793, 1580, 859, 1092, 1677, 135]. Integrating [1374, 1664, 1126, 1412, 1365, 1837, 331, 777]. intense [934]. intensity [1052, 1111, 1297, 2055, 1782, 553]. intensive [1962, 1523]. intentional [1780]. inter [21, 639, 636]. inter-river [636]. inter-specific [21, 639]. Interacting [1171]. Interaction [1585, 1093, 1027, 998, 1685, 397, 350]. Interactions [60, 1602, 1953, 1179, 712, 779, 1453, 312, 1587, 1820, 202, 1328, 191, 1792, 1508, 2093, 1222, 848, 625, 994, 1542, 662, 1159, 124, 892, 876, 490, 1083, 325]. Interactive [665, 1033, 1807]. Interannual [234, 678, 1506, 250, 76, 499, 252, 1980]. Intercohort [468, 124]. interface [106]. interference [1590]. interferes [1670]. interfering [195]. Intergenerational [1158]. interior [969, 302]. intermediary [1315, 1670]. intermediate [1519]. International [893]. interplay [365, 2024]. Interpopulation [223, 995]. interpretation [1441, 299, 1507, 460, 547]. interpretations [411, 345]. Interpreting [224, 512, 511]. Interrelations [218]. Interrenal [172, 1147]. Interspecies [2133]. interspecific [1865, 363, 1126]. interstage [1712]. interstitial [401, 1033]. intertidal [1903]. interval [115]. intervals [1301, 1119]. intestinal [1707]. Intra [21, 639, 1267, 1719, 1120]. Intra- [21, 639]. intra-otolith [1267, 1719, 1120]. intraguild [977]. Intralacustrine [1900]. intrasedimentary [437]. intraseries [63]. Intraspecific [770, 1865, 708, 1991]. intrinsic [429]. introduced [1199, 1993, 1059, 211, 1144, 424, 1211, 58, 1368, 2071, 931, 1982, 1673, 1684]. Introducing [1123]. Introduction [1098, 894, 817, 1989, 1035, 1064, 587, 296, 1511, 951, 1800]. introductions [1153, 1611, 1779, 1840, 609, 754, 1778]. introgression [1673, 184]. invader [599, 977]. invaders [831, 798, 608, 490]. invariants [780, 751]. Invasion [471, 335, 1109, 1018, 1455, 1788, 1840, 1718, 2093, 1679, 613, 1394, 1701, 826, 1780, 165, 1276, 1684, 273, 1030, 1269, 475]. invasional [490]. invasions [341, 1442, 611, 101, 606]. Invasive [1946, 1769, 1485, 2083, 1098, 1953, 374, 1229, 1716, 612, 1159, 1429, 916, 1503, 1861]. Inverse [1052, 1026]. Invertebrate [950, 1157, 1864, 741, 609, 1071, 832, 753, 2119, 1242, 1306, 386, 2135, 461, 404, 951, 522, 9]. Invertebrates [1277, 715, 333, 581, 952, 1565, 228, 186, 1169, 158, 1427, 994, 1512, 297]. invertébres [950]. inverted [1149]. investigated [879, 1013]. investigation [1226, 87, 10]. investment [1876, 1783, 1134, 1125]. ion [1047, 1710, 1670, 870]. ion-transport [870]. ionoregulation [888]. irideus [1753]. Irish [1618, 1843]. iron [899, 87, 10]. iron-stimulated [87, 10]. irradiance [1532]. irregularities [725]. irreversible [1622]. irroratus [1215]. Island [371, 301, 119, 1891, 1487, 1330, 2052, 26, 325]. Islands [43, 248, 2087, 212, 925]. isolated [1880, 842]. Isolating [1135]. isolation [1984, 1479, 526, 1901, 1780]. isolation-by-distance [1901]. isopycnic [323]. isothermal [191]. isotope [1047, 1970, 1267, 141, 505, 469, 470, 272, 438, 1068, 997, 452, 1897, 813, 1754, 1968, 373, 1961, 1293, 1790, 125, 584, 552, 1120]. isotopes [1273, 673, 546, 1160, 833, 1608, 585, 1061, 1908, 1105, 338, 1347, 1507, 1512, 1373, 1346, 2027, 1389, 1422, 1927, 1015]. Isotopic [343, 1741, 450, 15, 229, 1571, 963, 1326, 1915, 1117, 906, 1472]. issue [1800]. Isurus [747]. Itaipu [1121]. Italy [477, 1241]. Iteroparity [1950]. Ivlev [2028].

Japan [2093]. Japanese [1679, 1420, 1324, 1117, 1854]. japonica [1679]. japonicus [1117]. Jasus [488, 487, 1196]. Jean [80]. Joaquin [1700]. Johnson [2125]. Joseph [1400, 1399]. Juan [1959, 1928, 305]. jubatus [1328, 1537, 962, 1589]. just [6]. justify [1158]. juvenation [1775]. Juvenile [1592, 1310, 1436, 205, 1510, 682, 275, 291, 2004, 1607, 981, 940, 463, 1661, 1785, 791, 1036, 1156, 1041, 1978, 738, 1470, 183, 1874, 814, 2115, 639, 336, 557, 1953, 1792, 1604, 922, 526, 1490, 864, 1095, 2038, 435, 542, 1353, 885, 1366, 1396, 1524, 1509, 1035, 1873, 997, 1438, 142, 518, 1038, 218, 625, 1525, 1231, 1905, 327, 724, 845, 1452, 662, 1754, 499, 1609, 2016, 1040, 789, 1638, 771, 723, 425, 1753, 769, 2060, 1790, 1991, 559, 283, 77, 311, 1208, 868, 2037, 26, 1626, 1903, 287, 2062, 536, 2054, 790, 1553, 286, 273, 928]. juvenile [261, 1952, 1350, 1625, 213, 1934, 1355, 41, 1215, 1065]. juveniles [208, 1254, 384, 1117].

Kachemak [243]. Kajikia [1844]. Kalman [1887, 760]. kamtschatkana [1959, 1959]. Kansas [762]. Kejimkujik [2101]. Kennedy [90]. Keogh [30]. keta [2036, 1304, 845, 1901, 96, 1480, 2076]. Keweenaw [1856]. Key [668, 1221, 1881]. keys [260, 1682]. keystone [1894]. kidney [1560, 888]. kill [1097]. kill-at-age [1097]. kill-dominated [1097]. killer [1337]. kilometres [851]. kin [1879]. kinetics [2153]. king [288, 836, 1631]. Kirchhoff [996]. Kirchhoff-approximation [996]. Kirlin [804]. kisutch [1667, 332, 1910, 68, 1568, 170, 1523, 1536, 882, 1592, 97, 442, 1792, 1841, 542, 1632, 1396, 865, 1438, 1285, 1677, 782, 1226, 556, 1753, 1790, 77, 1208, 868, 391, 292, 1171, 603]. knotweed [1679]. knowledge [1947, 940, 2147, 1699]. known [846, 1929, 423]. known-age [846, 1929]. Kodiak [119]. Kohonen [1500]. Kokanee [564]. Kolmogorov [721]. Kootenai [2038]. kraft [574]. kriging [212]. Krill [1781, 2012, 187, 1219]. Kuparuk [1006].

L. [214, 184]. LA-MC-ICPMS [1273]. labelled [981]. Laboratory [336, 478, 1979, 912, 1835, 917, 1185, 155, 995, 14]. laboratory-based [1979]. laboratory-raised [1835]. laboratory-reared [917]. Labrador [365, 271, 529, 120, 1983]. Lac [130, 1917]. Lack [878, 1139, 2089]. LaCloche [983]. lactate [1721]. lacustrine [1299, 1126, 993, 1244, 1302]. ladders [1332, 1333]. laevigata [855, 856]. Lagoon [773, 1697, 352, 1678]. lagoonal [913]. Lake [1805, 211, 128, 983, 1169, 1305, 1831, 389, 1849, 999, 1514, 75, 1991, 381, 370, 475, 2007, 481, 202, 1339, 157, 1382, 1941, 161, 1784, 1795, 1793, 416, 1834, 774, 1087, 1594, 281, 2134, 314, 601, 1552, 59, 984, 1711, 944, 1504, 1998, 1787, 1351, 2070, 1953, 154, 1257, 2021, 767, 1086, 1989, 1089, 339, 101, 1569, 1367, 1386, 1776, 1290, 1170, 1803, 1064, 1477, 49, 2133, 1099, 2050, 646, 105, 963, 1961, 2140, 713, 296, 1812, 1289, 1467, 2146, 1261, 1429, 1346, 2094, 1511, 462, 1253, 294, 1148, 2001, 73, 1546, 204, 543, 2131, 1927, 1912]. lake [445, 786, 1804, 675, 1030, 456, 61, 2139, 2057, 1199, 571, 688, 650, 1455, 2039, 337, 1979, 340, 1174, 830, 1101, 1268, 1059, 1531, 1390, 1757, 1413, 1298, 1188, 1267, 1719, 1204, 78, 1321, 1556, 1892, 228, 141, 393, 1291, 528, 1852, 374, 1518, 822, 1519, 438, 964, 1265, 1266, 225, 150, 304, 567, 1394, 2020, 38, 20, 731, 1170, 1064, 1393, 278, 1856, 1570, 2130, 2084, 1289, 165, 2148, 1450, 1070, 697, 1603, 692, 126, 80, 1600, 87, 10, 2001, 370, 2017, 1927, 2075, 652, 2032, 32, 1120, 1980]. lake-dwelling [416]. lake-level [1148]. Lakes [1173, 2125, 1779, 222, 186, 2127, 2130, 1778, 2126, 732, 1230, 1012, 433, 562, 818, 1300, 1559, 1579, 793, 806, 440, 1246, 739, 725, 1474, 1155, 969, 2100, 1647, 1471, 200, 774, 306, 424, 431, 1094, 991, 529, 1501, 58, 2063, 1322, 1958, 1243, 1108, 1659, 967, 1866, 833, 2128, 1722, 1493, 700, 1360, 1532, 594, 1305, 371, 2021, 362, 441, 566, 1801, 1946, 1769, 162, 1071, 172, 937, 301, 302, 1752, 1430, 1414, 1810, 65, 91, 1691, 1386, 674, 1701, 1987, 1881, 691, 946, 1992, 217, 1806, 1090, 376, 1293, 884, 303, 1242, 1306, 667, 1808]. lakes [1009, 2094, 1206, 493, 1496, 1391, 1623, 1894, 623, 717, 1380, 405, 824, 1938, 1276, 1891, 961, 574, 464, 1187, 920, 204, 1742, 11, 1451, 106, 2101, 1807, 14, 1018, 2150, 2051, 1611, 1277, 62, 191, 754, 1694, 1830, 1940, 521, 948, 1209, 1257, 1357, 163, 402, 613, 748, 2050, 606, 490, 1348, 1733, 1683, 2085, 2086, 812, 610, 1750, 1361, 1912, 1882, 933]. Lamna [1567]. lamprey [57, 1409, 1694, 1830, 1265, 1266, 1776, 1750, 478, 1361, 1670]. lampreys [1690, 1826, 1967, 1940, 689]. lampricide [1670]. lance [662]. Land [2, 527, 1910, 1339, 1547, 1178, 673, 1958, 785, 2066, 402, 1303, 556, 1839, 573]. land-cover [1958]. landform [70]. landings [1736, 1846, 707]. landlocked [80]. Landscape [891, 200, 1880, 630, 1225, 556, 1703, 376, 921, 1751, 1057, 2010, 965, 1818, 550, 185]. Landscape-controlled [1703]. landscape-level [185]. Landscape-scale [1880]. Large [1794, 1632, 447, 539, 500, 670, 766, 412, 1322, 1504, 1530, 551, 1787, 833, 1608, 1953, 1095, 1473, 1698, 1974, 1786, 1121, 2060, 1975, 1261, 283, 137, 1506, 1331, 1746, 1926, 1253, 1487, 1684, 1927, 2032, 573, 629, 575, 809]. Large-scale [1794, 1632, 447, 670, 766, 1786, 2060, 1506, 1331, 2032]. largely [1495]. largemouth [56, 2152, 1486, 1528, 1721]. larger [1402]. larvae [179, 1745, 74, 807, 1298, 808, 1459, 1180, 1717, 2115, 15, 239, 121, 1096, 1080, 265, 1102]. Larval [1562, 196, 156, 197, 662, 178, 1510, 57, 275, 333, 251, 1599, 626, 89, 12, 1860, 1152, 1616, 1321, 1322, 246, 672, 2104, 175, 802, 1327, 1576, 1921, 835, 1943, 986, 849, 1515, 1856, 1828, 460, 504, 876, 1909, 1965, 1746, 155, 2054, 1613, 478, 1670, 1110]. laser [1273, 125]. Late [1060, 1268, 1448, 1570, 319]. late-run [1448]. Late-season [1060]. Lateolabrax [1117]. lateral [1739, 841]. latitude [1545, 402, 1968, 958]. Latitudinal [897, 1645, 529, 744, 1411, 772, 1430]. latus [941]. Laurentian [754, 948, 1348, 1683, 610, 1750, 933]. lavaretus [1719, 963, 1600, 2017]. Lawrence [831, 186, 887, 1831, 1229, 187, 898, 444, 1159, 319, 1878, 516, 136, 798, 581, 1566, 633, 1233, 1340, 324, 2049, 1026, 1926, 541]. layer [1913, 2030]. leached [653]. Lead [2109, 2069, 493]. leading [1755]. leaf [180, 1074]. learning [328, 565]. Least [635]. leatherback [1210]. leaves [1737]. LED [2088]. LED-based [2088]. legacy [1122, 2125, 2070, 2157, 1123]. Leiostomus [835, 213]. length [1278, 86, 633, 463, 859, 1655, 243, 1622, 1435, 399, 935, 2048, 1539, 1091, 152, 280, 1669, 1412, 1202, 499, 364, 2121, 850, 532, 824, 1253, 1653, 1019, 2145]. length-at-age [633, 1622, 499, 532]. length-based [463, 1655, 243, 1435, 1539, 1019]. length-frequency [280]. length-specific [399]. leniusculus [1292, 1389]. leopardus [146]. Lepeophtheirus [1356, 1297, 845, 939, 1561, 1686]. Lepetodrilus [1928]. Lepomis [36, 2153, 1020, 1509, 876, 1710]. less [1029, 1402, 1131]. lesser [2047]. lessons [1463, 1124]. lethality [1776]. leucomaenis [132]. Leucoraja [1411]. leukemia [1239]. level [1947, 1349, 1305, 887, 1519, 1643, 82, 510, 1968, 707, 1198, 185, 1148, 2139]. levels [1003, 440, 388, 16, 97, 389, 1490, 564, 703, 786, 201]. levis [1724]. lewisi [1788, 821, 2071, 1780]. lice [2088, 1297, 845, 939, 1625, 1561, 1686]. Life [1129, 1363, 1133, 320, 1786, 319, 1136, 1138, 1748, 2109, 2069, 1122, 1498, 1587, 1118, 46, 1292, 211, 2112, 1200, 1267, 1240, 1162, 1577, 1867, 348, 1235, 2087, 757, 604, 1922, 1035, 1126, 788, 132, 1818, 1672, 892, 1991, 1139, 1123, 1125, 1433, 1117, 1637, 584, 552, 1934, 780, 751, 1681]. life-history [1118, 1162, 1922, 1035, 780, 751]. life-history-based [1681]. lifetime [1202]. ligand [917]. ligands [1353]. Light [1297, 2119, 1979, 1474, 4, 191, 808, 2088, 2015, 998, 564, 2055, 1660, 8, 553]. light-based [1660]. Light-dependent [2119]. lightly [873]. like [157, 5]. Likelihood [1119, 407, 308, 590, 507, 1077, 2162]. Likelihood-based [1119]. likelihoods [1228]. Limanda [2078]. limbatus [557]. Limfjord [1009]. liming [1291]. limit [3, 1815]. limitation [1073, 191, 1085, 2066, 1393, 1992, 8, 137, 404, 1391, 2062, 1891]. limitations [1426, 1076]. limited [2083, 1984, 1657, 2104, 986, 1704, 1141]. limiting [341, 443]. limits [37, 957, 1439, 1888]. Limnodrilus [417]. Limnological [2110, 793, 85]. Limnology [2074, 1777]. limosus [1953]. line [329]. lineages [184]. linear [354, 78, 844, 843]. lines [2079, 1423, 1697]. lingcod [2019, 930]. link [268, 657, 750, 1357, 541]. Linkages [1089, 1675, 702, 1125]. linked [679, 721]. Linking [300, 980, 299, 2028]. lion [735, 982, 1328]. lions [1537, 962, 1589]. Lipid [1282, 1359, 966, 1783, 1507, 842, 1628]. lipids [1709, 1919, 1478, 1005]. literature [2029]. litre [109]. litter [1074]. litterfall [653]. little [1411, 381, 652]. Littoral [405, 1938, 1382, 1902, 1474, 1834, 314, 1504, 1169, 767, 1414, 1810, 289, 1367, 1900, 713, 1242, 1261, 1346]. littoral-spawning [1900]. lived [372, 1259, 2144, 1516]. liver [1859]. lives [585]. lividus [708]. living [1172, 519]. LLD [686]. load [109, 700]. loading [200, 49, 591]. loadings [1741]. loads [394, 617, 1451]. lobster [628, 794, 940, 957, 488, 487, 582, 2008, 1196, 1921, 280, 605, 602, 2026]. lobsters [76, 218, 380, 1487, 182]. Local [940, 1066, 1739, 1757, 1328, 1441, 840, 1112, 1663, 1521, 65, 534, 1072, 655]. local-origin [1757]. localization [478]. localized [1562, 1984]. location [1356, 179, 774, 828, 436, 2030]. locations [2143, 1061]. loci [1, 1197]. locomotion [1046]. locus [1283, 1437, 1550]. logbook [721]. logbooks [245]. logged [778]. Logging [1579, 1471, 880, 674, 691, 405]. Logging-induced [1579]. logic [189, 454, 1335]. logistic [1817]. logits [364]. lognormal [308]. Loliginidae [336]. Loligo [248, 942, 336, 1383, 1944, 1203, 925]. Long [1531, 1390, 1006, 1848, 2114, 1291, 1049, 1044, 1035, 1803, 2127, 1674, 1806, 707, 1812, 852, 2148, 404, 1915, 591, 1534, 1807, 1034, 1158, 1066, 372, 1259, 878, 2144, 1602, 1237, 1358, 465, 2128, 2087, 822, 2068, 1478, 1516, 1477, 1775, 2027, 1623, 1030]. long-chain [1602]. long-distance [1478, 2027]. long-lived [372, 1259, 2144, 1516]. Long-residence [1034]. Long-term [1531, 1390, 1006, 1848, 2114, 1291, 1049, 1044, 1035, 1803, 2127, 1674, 1806, 707, 1812, 852, 2148, 404, 1915, 591, 1534, 1807, 1158, 1066, 878, 1237, 1358, 465, 2128, 2087, 822, 2068, 1477, 1775, 1623, 1030]. longfin [942, 1383, 1944]. longimanus [1018, 806, 1876, 2119, 1503, 2007]. Longitudinal [1498, 644, 1739, 334, 1325, 1633, 1330]. longline [592, 1936, 1483, 1412, 129, 1161]. longlines [1146]. longnose [1442]. looking [887, 1777]. loss [2078, 1039, 1143, 1417, 381]. losses [1417, 771]. lota [1935, 1935, 1953]. lotic [2080, 741]. Lotka [22]. Louisiana [465]. louse [1356]. Low [1745, 1947, 1993, 1949, 2144, 1565, 1659, 432, 887, 1946, 1769, 1384, 1884, 1195, 1338, 313, 1637, 1034]. low-alkalinity [1659]. low-density [432]. low-fecundity [2144]. low-gradient [1195]. low-head [1384]. low-level [887]. low-nutrient [1946, 1769]. low-order [313]. low-trophic-level [1947]. lower [2051, 1902, 1682, 389, 1992, 319, 1982, 1878]. lowland [1025, 1329, 1974, 1010, 500]. LR [162]. lucioperca [45]. lucius [1199, 40]. Lury [925]. Lutjanus [1759, 1934]. lyase [897]. Lymnaea [413, 117].

maccoyii [860]. Mackenzie [1501, 1992, 1391, 623, 717]. mackerel [1284, 1712, 559]. Macoma [437, 615, 5]. Macquaria [1754]. macroalgal [1166]. macrobenthic [1241]. macrocephalus [1984, 23]. macrochirus [2153, 1509, 876, 1710]. macrocosm [1576]. macroenvironmental [1521]. macrofauna [355, 353]. macrofaunal [648, 1214]. Macroinvertebrate [880, 607, 1811, 39, 1902, 1834, 1809, 2021, 1810, 304, 545, 334, 1972, 14]. macroinvertebrates [1230, 643, 1482, 358, 1847, 709, 1496, 1260]. macrophyte [2100, 881, 964, 402, 717, 204, 1406]. macrophytes [591, 1408, 675, 752]. macrozooplankton [193]. maculosa [1935]. Madeleine [1697, 1423]. maenas [1215]. Magdalen [240]. magellanicus [476, 2033, 279, 954, 1215]. magister [1019]. magnetic [1467, 1838]. magnitude [531]. main [344]. Maine [756, 628, 76, 1182, 1469, 252, 223]. mainstem [1538]. maintain [1884]. maintains [1197]. maintenance [1196]. Maisons [1697]. major [589, 620, 264, 228, 1678]. make [2141]. making [695]. mako [747]. Malawi [1204, 2020]. Male [332, 1498, 1587, 1502, 226, 366, 658, 1632, 508, 1631, 689, 1163, 1019]. male-only [1631]. males [644]. Mallotus [264, 1341, 1717, 2115, 138, 1828, 516]. malma [443, 295]. mammal [934, 883]. manage [671]. managed [1253]. Management [1750, 43, 1415, 2155, 1122, 1475, 657, 2107, 680, 953, 952, 2079, 2154, 855, 1755, 1397, 459, 1694, 3, 2043, 1940, 1768, 266, 1505, 1196, 1555, 1857, 2104, 269, 679, 1642, 1529, 1250, 631, 2068, 1776, 1650, 1403, 1868, 1766, 1203, 218, 235, 1045, 1515, 565, 713, 1960, 1812, 968, 1925, 1123, 1331, 1395, 651, 1416, 1861, 1001, 1882, 780, 751, 1138, 595]. managers [1463]. Managing [836, 202]. Manan [1487]. manatee [2121]. manatus [2121]. manipulation [2128]. manipulations [183, 1592, 49]. manner [1035]. many [1, 2050]. Mapping [1454, 777, 2008, 1220, 1964, 1758, 1261]. maps [1500]. margin [322]. Marguerite [1687]. mariae [1967]. Marie [1289]. Marine [1920, 614, 269, 1105, 187, 631, 1494, 516, 1066, 535, 1312, 1424, 1118, 86, 1906, 907, 580, 2122, 581, 952, 1164, 648, 1863, 2064, 143, 1275, 546, 940, 1428, 412, 1896, 274, 1137, 1442, 1363, 1479, 1675, 1445, 1374, 133, 1418, 1449, 1813, 267, 1042, 1764, 1167, 1671, 1140, 1709, 649, 1192, 1643, 1545, 1495, 1427, 1179, 775, 1943, 986, 1847, 555, 845, 1660, 1798, 721, 13, 1181, 702, 1014, 1419, 934, 619, 1548, 1145, 419, 28, 325, 1331, 930, 1926, 1422, 1872, 909, 1656, 2006, 603, 2032, 883]. marine-derived [1847, 1145, 1872, 2006]. marine-phase [555]. marine-relict [1479]. marine-reserve [986]. marinus [1690, 1826, 57, 1860, 1409, 1694, 1830, 1940, 1265, 1266, 1776, 689, 1750, 478, 1361, 1670]. Maritimes [1487]. mark [2078, 1892, 604, 1873, 661, 1345, 663, 1288, 2103]. marked [1942]. markers [981, 1184, 214, 1437, 229, 1061, 1064]. market [1203]. marking [1754, 1422]. markings [428]. Markov [1931]. marks [290, 131]. marlin [1844, 1815]. marmorata [1909]. marshes [402]. Marys [1694]. masou [658, 1974]. mass [1273, 1621, 1565, 175, 972, 1968, 125, 88, 42, 1589]. mass-balance [1621, 1968]. mass-balanced [972]. Massachusetts [108, 107, 112, 110]. masses [662]. masu [658, 1974]. match [206]. match/mismatches [206]. Matching [1855, 1141]. mate [1583]. Maternal [393, 1916, 943, 1017, 1876, 2018, 1515, 1628]. maternal-age-dependent [1515]. mating [882, 2099, 1900, 1583]. matings [423]. matrix [105]. matter [84, 1599, 401, 1655, 102, 767, 502, 1100, 91, 997, 972, 1053, 709, 1733, 574, 1262, 1150]. matter-mediated [1053]. maturation [1587, 1567, 1572, 1135, 226, 1132, 2117, 1912]. mature [2102, 1163]. maturity [53, 1341, 1755, 2099, 16, 676, 1304, 1817]. maxima [456]. maximum [407, 308, 1949, 578, 1642, 590, 507]. maximus [634, 808]. mayflies [111]. MC [1273]. mean [1532, 1653]. means [434]. measure [1118, 694, 1248, 198]. Measurement [578, 2138, 1954, 1028, 1099, 460, 2060, 467]. measurements [976, 947, 41, 908]. measures [934]. Measuring [531, 2067, 1919, 138, 169, 1832, 1721]. mechanism [393]. Mechanisms [1370, 2105, 1895, 1193, 1630, 886, 609, 494, 485, 499]. mechanistic [274, 298]. médaille [1544]. Medal [1544]. median [635]. mediate [854]. mediated [433, 754, 1053]. medirostris [1436, 2098]. Mediterranean [1495, 707, 131]. Mediterranee [352]. meet [1271]. megalops [1533]. Meganyctiphanes [1219]. meiofauna [645]. Melanogrammus [716, 251, 1823, 2045, 349, 2044, 1843, 1017]. melanops [986, 1251]. melanostictus [1854]. melanostomus [1059]. melt [190]. meltdown [490]. membras [852]. Memoriam [686, 1731, 1732]. Mendota [165]. Menidia [1124]. mer [1544]. Mercenaria [503]. Mercury [1094, 107, 2101, 1230, 112, 1320, 340, 1605, 109, 1471, 1409, 108, 388, 2132, 222, 469, 470, 272, 647, 767, 1987, 111, 1512, 303, 1011, 18, 1930, 88, 42, 1484, 110, 1742]. merge [1369]. méridionale [1799]. Merlangius [1843, 2030]. merlangus [1843, 2030]. Merluccius [1549, 682, 616, 1723, 2087, 175, 1068, 145, 784]. meroplanktonic [265]. meso [1149]. meso-zooplankton [1149]. mesocosm [1256, 239]. mesocosms [55, 703]. mesohumic [61]. mesotrophic [713, 1511]. Meta [1768, 638, 1606, 533, 1142, 392]. Meta-analysis [1768, 1606, 533, 1142, 392]. meta-model [638]. Metabolic [1658, 400, 1245, 1075, 1114, 711, 570, 233, 2063, 1643, 1913, 1905, 618, 1120]. metabolism [275, 196, 1793, 4, 808, 284, 872, 197, 1315, 360, 771, 1406, 445, 801, 1670]. metabolites [1314]. metacommunities [2051]. Metal [39, 1108, 1805, 711, 437, 1071, 1691, 5, 1707, 618, 44, 87, 10]. metal-contaminated [711, 1071, 1707]. metal-rich [437]. metalimnetic [456]. Metallothionein [5, 172]. Metallothionein-like [5]. metals [143, 1147, 1290, 136]. metamorphic [86]. metamorphosis [48, 1447, 1843, 1854]. method [1858, 474, 677, 1197, 1833, 829, 158, 1490, 1372, 1513, 1597, 1964, 600, 804, 1352, 2091, 705, 2035, 537, 1154, 803, 1743, 135, 958, 1735, 584, 552, 1021]. methodology [1584, 229, 1997]. methods [1741, 776, 2078, 2067, 1749, 1825, 1427, 429, 1483, 1469, 1570, 19, 1128, 1462, 1146]. Methow [411, 345]. Methuselah [955]. Methyl [1230, 222, 647, 2132, 767, 303, 1011, 110]. methylation [1094]. methylmercury [1902, 1471, 2041, 1306]. metric [1596, 1103]. metrics [151, 558, 1388, 718]. Mexico [372, 1605, 1177, 645]. Michigan [1018, 650, 1973, 157, 1268, 1531, 1390, 1298, 1267, 78, 1321, 1892, 393, 1852, 374, 567, 1394, 1393, 692, 2001]. Michx. [136]. Microbial [653, 60, 1074, 1083]. microchemistry [1404, 1830, 1251, 216]. microcosm [994]. microcosms [2111, 417]. microcrustacean [1891]. microcystin [162, 1859]. microcystin-LR [162]. Microcystis [24, 370]. microgeographic [240]. microhabitat [1974]. microlepis [2124, 16]. Micromesistius [2034]. microprobe [1047, 1710]. Micropterus [2150, 1665, 1211, 1135, 56, 1989, 2152, 289, 1486, 1528, 1721, 1771, 1294, 1923]. microquantity [1919]. Microsatellite [1223, 349, 747, 1416, 95, 634, 524, 214, 1437, 237, 2098, 1247, 1844, 1251, 1224, 939]. Microsatellites [2031, 1550]. microstructural [1909]. microstructure [1714]. microzooplankton [1268]. mid [1383, 1026]. mid-1980s [1026]. mid-Atlantic [1383]. middle [444]. midwestern [65]. might [701]. migrant [1826, 974]. migrants [2147]. migrating [1761, 2078, 247, 2118, 872, 1614, 439, 914, 928, 1914, 1410]. Migration [1047, 816, 684, 1075, 682, 1690, 2156, 1654, 670, 1448, 1829, 1641, 59, 2015, 834, 1518, 841, 1875, 1519, 1841, 1281, 1044, 1473, 1642, 1698, 2022, 1327, 1182, 1904, 837, 142, 1478, 1060, 862, 1554, 1909, 325, 1781, 2012, 666, 1720, 1117, 193, 603, 2032]. migrations [1566, 1646, 898]. migratory [57, 776, 1822, 872, 1664, 343, 452, 478, 1063]. milfoils [1972]. mill [315, 137, 574]. millennium [259]. mine [205, 39, 480, 313]. minerotrophic [1011]. miniatus [2104]. minibicornis [1245]. mining [1975, 1930]. mining-impacted [1930]. minke [853, 911]. Minnesota [1757, 884]. minnow [58, 1367]. minnows [569]. minor [1540, 1490]. Miramichi [2082, 642]. mirrors [714]. mis [230]. mis-specification [230]. misclassification [1000]. misleading [1961]. mismatches [206]. Missing [1555]. Mississippi [1401, 932, 673, 344]. Missouri [965, 2120]. mitigation [1867]. mitochondrial [1967, 1844, 943, 931, 184]. Mixed [2068, 1696, 783, 729, 2142, 929, 1345, 525]. Mixed-fishery [2068]. mixed-species [525]. mixed-stock [1696, 2142, 929, 1345]. mixing [383, 1246, 191, 1569, 324, 240]. mixta [1005]. mixture [1247, 1369, 1817, 1648]. mixtures [1963]. Mizuhopecten [1971]. mobile [535, 1914]. Model [1107, 1046, 37, 1024, 1713, 2019, 411, 345, 354, 337, 907, 615, 288, 1557, 241, 189, 367, 638, 1860, 151, 463, 664, 805, 1156, 1621, 859, 1349, 1456, 1755, 488, 243, 1522, 352, 718, 1726, 253, 351, 1175, 378, 708, 298, 323, 1435, 250, 521, 181, 1723, 617, 133, 1852, 935, 823, 1272, 1205, 1933, 1222, 578, 1889, 1539, 1576, 904, 1096, 889, 1203, 429, 925, 1345, 6, 1483, 1412, 422, 583, 507, 1660, 1968, 917, 1226, 1712, 159, 1526, 467, 1238, 1009, 866, 919, 1458, 745, 869, 525, 406, 692, 790, 2028, 2151, 1740, 1819, 88, 42]. model [1103, 1937, 1656, 2103, 641, 563, 1270, 1980]. model-based [1483, 1412]. Modeling [180, 2010, 1871, 1096, 613, 1669, 1943, 692, 185, 257, 1287, 1007, 1321, 672, 700, 844, 843, 1247, 163, 1676, 1758, 6, 533, 1817, 293, 1735, 918, 23, 1519]. modelled [923]. Modelling [1407, 535, 903, 1761, 251, 1246, 1622, 1576, 1375, 114, 2123, 1072, 1269, 1951, 1291, 1192, 1921, 1743, 1026, 1484, 909]. Models [1933, 1798, 236, 1024, 1023, 1739, 1382, 161, 196, 1287, 520, 143, 1580, 2154, 1619, 78, 529, 1867, 517, 1152, 1616, 1057, 729, 1896, 772, 996, 208, 815, 1537, 399, 505, 307, 1604, 1612, 1216, 2013, 844, 843, 1357, 371, 17, 290, 396, 397, 1058, 414, 2000, 197, 1922, 1873, 104, 1868, 331, 795, 949, 415, 637, 1151, 534, 124, 1931, 796, 387, 1306, 386, 1139, 2042, 2094, 640, 258, 635, 92, 921, 1883, 1050, 920, 927, 777, 1019]. moderate [1086]. moderating [674]. modification [351]. modified [471, 335, 885, 1774]. modulate [1587]. modulation [605]. Molecular [936, 656, 69]. mollusc [503]. molt [1431]. molting [602]. Moluccas [409]. Mongolia [2116]. monitor [1452]. Monitoring [2088, 2056, 938, 1757, 670, 2080, 1111, 465, 214, 1385, 1749, 1490, 1438, 462, 1149, 194]. montane [359, 961]. Monte [2154, 1091, 2000]. moon [59]. moored [1833, 1149]. moratorium [385]. mordax [1562, 163, 1701, 444, 1070, 1683]. morhua [236, 1158, 1282, 1359, 1620, 1407, 544, 1234, 179, 1054, 971, 1191, 1489, 846, 633, 1233, 670, 1572, 1573, 1929, 83, 1949, 1962, 699, 1152, 1616, 403, 2067, 1717, 1479, 672, 988, 923, 922, 81, 1133, 1585, 864, 802, 1334, 2024, 676, 1576, 82, 103, 1314, 1315, 1942, 1403, 788, 2157, 1231, 1116, 1254, 1693, 138, 1132, 1775, 816, 1550, 2117, 771, 1931, 223, 1884, 1770, 66, 970, 1583, 1843, 2049, 1827, 1338, 1344, 2105, 1720, 240, 1926, 122, 655]. Morone [1816, 572, 2022, 48, 414, 92]. Morphological [1575, 1765, 1764, 1309, 1384, 936, 1244, 667]. morphologically [1468]. morphologies [1542]. morphology [800, 1235, 750, 965, 284, 1695, 18]. morphometric [459, 1663, 1314]. morphometrics [1929, 1589]. morphometry [700, 2021, 217]. morphotypes [2057]. mortality [1666, 2077, 1165, 966, 251, 2156, 1654, 1233, 668, 463, 482, 83, 1041, 488, 243, 582, 990, 368, 1586, 941, 1717, 2115, 102, 557, 1911, 1849, 1265, 2038, 2118, 48, 45, 238, 1491, 100, 235, 554, 1060, 504, 1712, 1615, 2060, 380, 1125, 532, 531, 536, 1735, 227, 513, 1077, 1613, 1651, 342, 1146, 1888, 1410, 2162]. Morus [555]. motility [508]. motivation [769, 854]. mottled [1973]. Mougeotia [55]. mountain [1415, 1073, 1016, 1527, 891, 1155, 424, 937, 972, 946, 1293, 456, 1165, 1758, 961]. Mountains [70, 722, 778]. Movement [813, 380, 968, 1008, 210, 910, 596, 1341, 2008, 2014, 435, 1997, 152, 2152, 597, 468, 329, 826, 1376, 2027, 1506, 704]. Movements [1667, 1973, 1025, 930, 1432, 997, 1392, 496]. moving [1565]. Moxostoma [1890]. MPA [1545]. MPAs [1545]. MRI [1838]. mRNA [642, 730]. mtDNA [726]. Mucus [1970, 2030]. Mugilidae [1663]. mullet [2097]. Mullus [2097]. multi [1677]. multi- [1677]. multibasin [314]. multicollector [1273]. Multidecadal [1630, 738, 2011]. Multifrequency [1595, 231, 1513, 1597]. multigene [1967]. multimetric [1057]. multinomial [1612]. Multiple [1593, 184, 219, 1178, 1581, 1665, 1661, 1298, 1175, 2074, 598, 290, 1664, 1814, 703, 1806, 1966, 895, 1956, 2103, 563]. multiple-pass [563]. multiple-recapture [2103]. multipopulation [1783]. Multiscale [1685]. multiseries [1883]. multisite [580]. Multispecies [1192, 43, 37, 1855, 457, 1435, 701, 1466, 1434, 1395, 595, 2068]. multispecific [1990]. multistage [766]. multistate [2078, 1873]. multistock [980]. multistrata [225]. Multivariate [1118, 997, 316, 1057, 1500, 459, 1789, 2018, 7, 9]. Multiyear [473, 488, 414, 1365, 1837]. mummichogs [2041, 390]. municipal [137, 591]. Murray [340, 1482]. Muscle [1402, 16, 97, 1490, 452, 1859]. muscles [1734]. Muskoka [486, 1804]. mussel [446, 527, 831, 798, 1401, 1175, 1933, 65, 165, 1423, 1697, 1918, 353, 370]. mussels [562, 818, 112, 60, 1160, 1773, 1835, 389, 1229, 1946, 1769, 101, 1394, 75, 485, 811, 876, 697]. Mustelus [22]. Mycteroperca [2124, 16]. Myers [1731, 1732]. mykiss [1753, 64, 1797, 530, 966, 2137, 1920, 1788, 291, 570, 1757, 783, 1822, 1978, 2014, 1115, 1461, 1722, 821, 526, 2093, 284, 1093, 1950, 2118, 1194, 1698, 1353, 604, 1664, 508, 727, 1913, 698, 2071, 1818, 1440, 917, 945, 712, 2061, 1185, 2135, 931, 850, 1575, 461, 888, 27, 28, 29, 921, 1433, 1796, 286, 273, 30, 1350, 801, 1262, 1688, 216]. Myosin [1490]. Myriophyllum [1408, 1972]. Myrophis [1447]. Mysid [1556, 675]. Mysidacea [1005]. Mysis [1979, 438, 1005, 1966]. mysticetus [220]. mystinus [863]. Mytilus [1175, 1239, 353].

N [649, 1088, 1329, 649, 150, 1571, 1677, 1968, 373, 1687]. N-rich [1329]. N.Y [722]. N/ [649]. naccarii [360, 361]. naïve [1624]. namaycush [1795, 774, 1594, 281, 1257, 1989, 1386, 1776, 1290, 1064, 1099, 2050, 646, 2140, 73]. nase [35]. nasus [1567, 35, 79]. natal [1832]. National [128, 232, 2101, 1789, 1881]. native [1788, 1020, 58, 1953, 821, 1194, 1250, 2136, 977, 215, 753, 2071, 1159, 826, 1780, 1982, 1623, 1673, 1972, 1453]. Natural [83, 1586, 102, 2021, 1064, 536, 1666, 2077, 1714, 2036, 634, 1599, 2153, 614, 783, 1233, 524, 1027, 1523, 488, 243, 471, 335, 2104, 229, 1632, 885, 1495, 238, 452, 38, 705, 2081, 1310, 5, 1712, 762, 1185, 428, 227, 513, 1253, 1796, 1077, 2162, 574, 478, 2017, 656, 1262]. natural- [2036]. natural-origin [524]. naturally [1523, 1061, 727, 1185]. Nature [1876, 1009, 1176, 759]. navigation [1530]. near [205, 1928, 2052, 1958, 645, 1149]. Nearctic [1935]. nearly [954]. nearshore [948, 374, 594, 1842, 845, 119, 1413]. necessary [587]. Nécrologie [90, 686]. needed [1417]. needs [1378]. negative [1654, 1998, 502]. nematode [1794]. nematodes [936]. Neogobius [1059]. neonate [1876]. Neotropical [1121]. Nephrops [1068, 1921]. nerka [688, 1448, 839, 1456, 1536, 944, 2070, 1049, 1555, 1857, 1272, 1205, 867, 2048, 872, 564, 1922, 1478, 1766, 625, 973, 1060, 943, 1672, 760, 737, 1991, 1444, 99, 1673, 666, 914, 1872, 1245, 2032, 201, 1410]. nervous [934]. nest [1294]. Nested [295]. nesting [1020]. net [695, 1417, 2028]. Netherlands [1700]. nets [847, 159]. netting [1309]. Network [1896, 1342, 910, 979, 972]. Network-scale [1342]. networks [2023, 1990]. neural [1990]. neurotoxicity [868]. Neuse [588]. Neustonic [76]. neutral [413, 480, 842]. Newfoundland [179, 365, 1489, 1341, 355, 103, 1828, 1554, 1005, 380, 1983, 1746]. newly [1667, 434, 828, 1512]. next [1913]. Ni [1108]. Niche [378, 494, 2020, 2057]. niches [1915]. Nikolsky [184]. Nile [1853, 1485]. niloticus [1853, 1485]. Niña [1069]. ninespine [726]. Niño [1971, 1069, 1313, 1809]. Nitrate [673, 932]. nitrification [549]. Nitrogen [579, 1421, 946, 2083, 394, 614, 546, 969, 1160, 985, 881, 1916, 1086, 896, 1061, 1105, 338, 1068, 150, 1347, 452, 1507, 1373, 2027, 549, 126]. nitrogen-limited [2083]. nitrophenol [1670]. nitzschia [1883]. No [1879, 1118, 1943, 1417, 380, 1109, 2007]. no-take [1118, 1943, 380]. nocturnal [1781, 2012]. Noel [1465, 1464]. noisy [990]. nominations [1544]. non [1356, 1533, 1191, 2153, 554]. non-additivity [554]. non-host [1356]. non-recovery [1191]. non-steady [2153]. non-target [1533]. nonanadromous [216]. nonequilibrium [641]. nonesterified [1282, 1359]. nonindigenous [806]. Nonlethal [1082, 1275]. nonlinear [17, 1091, 299, 640, 1228]. nonmetamorphic [1643]. Nonnative [1623, 821, 826, 1780, 1453]. Nonparametric [1000, 1218]. Nonrandom [1583, 882, 1900]. nonretention [170]. nonsalmonid [1894]. nontrophic [994]. nonvisual [619]. nonylphenol [1350]. norms [1977, 2117]. north [161, 1367, 713, 1960, 1181, 1808, 1030, 236, 1282, 1359, 446, 903, 562, 818, 1106, 1234, 953, 1054, 971, 1377, 53, 820, 1860, 1762, 1935, 699, 1221, 120, 738, 1939, 323, 1385, 897, 2034, 1718, 1875, 2098, 2066, 1543, 2024, 1327, 853, 1644, 101, 2068, 1483, 1412, 1969, 392, 1254, 1931, 503, 606, 1134, 1884, 1770, 1771, 173, 1854, 295, 576, 1497, 1013, 906, 1972, 69, 1748]. north-temperate [161, 1808]. Northeast [1928, 1068, 305, 1069, 1459, 988, 244, 1649, 1712, 395, 911, 2105]. northeastern [477, 2110, 1984, 4, 800, 1866, 441, 2074, 654, 1828, 824, 1439]. northern [1199, 385, 372, 2052, 1605, 982, 1111, 1094, 976, 1041, 817, 1957, 81, 2093, 1192, 190, 1414, 1900, 426, 100, 540, 755, 138, 1992, 376, 667, 66, 2049, 1026, 137, 40, 224, 1623, 1894, 1288, 1751, 732]. Northumberland [2026]. Northwest [1310, 349, 2031, 1017, 1567, 1409, 1944, 1133, 451, 858, 865, 901, 2092, 555, 1767, 1338, 1673, 134]. northwestern [1177, 1762, 1411, 1430, 691, 295]. norvegica [1219]. norvegicus [1068, 1921]. Norway [1595, 1291, 2024, 1921]. Norwegian [517, 1432, 499, 1775]. Notropis [918]. Nova [1239, 2101, 1007, 403, 1659, 1451]. novel [1155, 1676, 2035, 1099]. NRC [926]. NSS [1196]. nuclear [1467]. number [366, 240]. numbers [1669]. Numerical [1692, 351, 208, 1613]. Nunavut [371]. nurseries [1760]. nursery [944, 557, 2070, 1481, 828, 1166, 1316, 2062]. nurture [1876]. Nutrient [161, 594, 973, 1992, 408, 433, 1713, 1455, 140, 1584, 191, 569, 1237, 1358, 154, 2066, 1946, 1769, 1514, 49, 2127, 137, 550, 591, 313, 1010, 1391, 717, 353, 1652, 445]. Nutrient-chlorophyll [161]. Nutrient-dependent [408]. nutrient-impacted [1584]. nutrient-light [191]. nutrient-rich [1010]. nutrients [793, 1547, 1425, 1540, 1499, 964, 867, 965, 502, 937, 252, 376, 1387, 8, 1145, 623, 1872, 2006, 752]. nutritional [928].

O [1710, 1932]. O. [687]. obesus [1936]. Obituary [90]. objective [635]. objectives [1791]. obligate [1442]. observation [1234]. Observations [1914, 2005, 1576, 1185, 193, 1724]. observed [274, 1128]. observer [2141, 604]. obstacles [1654]. obtained [489, 1877]. OC [686]. occupation [2157]. occupied [1724]. occurrence [1274, 1771, 1839]. occurring [2097, 1061, 1890, 490]. occurs [2046]. Ocean [1313, 1544, 1714, 1448, 886, 1537, 442, 904, 540, 755, 392, 1005, 66, 895, 2054, 1634, 1149, 603, 1312, 1214, 659, 660, 1984, 897, 1606, 1944, 900, 1825, 1327, 349, 2096, 299, 1844, 1649, 173, 1854, 906]. ocean-type [886]. oceanic [905, 1644, 1043, 913]. Oceanographic [1181, 2096, 322]. oceanography [1171]. oceans [362, 908]. ocellata [1411]. ocellatus [15, 560]. odontocete [548]. Odor [868, 478]. Odor-evoked [868]. odour [1882]. off [546, 1341, 403, 2087, 245, 2092, 662, 1154, 1662, 325, 1008, 1487, 22]. offs [1129, 1519, 2118, 565, 1780]. offset [1417]. Offshore [385, 1298, 767, 645, 1413]. offspring [210, 749, 79]. OH [14]. Oil [1866, 357, 509]. Oklahoma [1665]. oleate [801]. Olfactory [1480]. olidus [74]. Oligochaeta [417]. oligotrophic [59, 767, 1045, 1206, 1927, 11]. omnidirectional [496]. omnivore [1684]. omnivorous [1645, 1292, 1924, 1389]. Oncorhynchus [1667, 938, 589, 1047, 688, 95, 1714, 1457, 2160, 1587, 1426, 64, 332, 749, 1797, 2036, 2023, 992, 530, 1910, 966, 2137, 1920, 596, 63, 1788, 68, 291, 1301, 219, 2156, 1568, 570, 1757, 783, 1448, 2112, 1885, 482, 1184, 1822, 2154, 170, 839, 1456, 1523, 1978, 1639, 1536, 882, 2102, 168, 167, 1592, 2014, 886, 944, 1880, 2099, 834, 1560, 1115, 501, 1437, 97, 1461, 2070, 1722, 821, 442, 1499, 1049, 1555, 1749, 1857, 1792, 1272, 1205, 526, 2093, 763, 1342, 1095, 1841, 867, 284, 1093, 1473, 1950, 2118, 1194, 542, 2048, 1698, 872, 1353, 649, 604]. Oncorhynchus [1664, 658, 1632, 564, 1396, 778, 865, 1922, 508, 1438, 889, 1478, 1766, 625, 1285, 1307, 727, 1525, 959, 1913, 698, 973, 1304, 1420, 1324, 1310, 845, 1452, 1677, 540, 755, 2071, 1974, 1818, 1060, 782, 1440, 917, 1901, 1226, 1695, 943, 565, 789, 1672, 945, 1638, 723, 556, 19, 760, 483, 484, 826, 1780, 369, 1753, 769, 1960, 395, 737, 1615, 712, 2060, 1790, 2061, 96, 1991, 1185, 1217, 2135, 77, 1208, 931, 1662, 989, 868, 1480, 919, 177, 1145, 1444, 99, 192, 968, 1376, 850, 1575, 1903, 2027, 391, 461, 292, 888, 2076, 27, 28, 29]. Oncorhynchus [1171, 1673, 666, 914, 921, 651, 1050, 1433, 1796, 209, 286, 1872, 687, 273, 928, 1245, 1985, 2161, 30, 1350, 1001, 584, 552, 801, 1625, 603, 2032, 1355, 41, 1262, 201, 1120, 1688, 1270, 1410, 176, 216]. one [1131]. Oneida [1552, 389, 1849, 75]. only [1631, 1819]. onset [36]. Ontario [163, 1170, 2127, 475, 2007, 1805, 1979, 1756, 2114, 1556, 1809, 617, 222, 700, 438, 441, 2074, 225, 150, 34, 304, 731, 1987, 1803, 691, 486, 1570, 1306, 1496, 812, 786, 1804, 1120, 1807]. Ontogenetic [1409, 2057, 1470, 837, 1015]. ontogeny [1436, 1235, 2037]. oocyte [2105]. opalescens [1203]. open [146, 670, 974, 587]. open-channel [974]. open-water [587]. opening [1273, 1062]. Opeongo [1170]. operation [2137]. Ophichthidae [1447]. Ophiodon [2019, 930]. opilio [1296, 1081, 1340, 1767, 678, 768]. opossum [1979]. opportunist [1141]. opportunities [1113]. Opposite [540, 755]. optical [481]. Optimal [1791, 247, 701, 1125, 651, 1634, 580, 774, 2144, 1057, 772, 572, 238, 1633]. Optimum [1003]. orange [2109, 2069, 1505]. orca [1337]. Orcinus [1337]. Orconectes [1020, 1351, 1953, 2130, 1429, 1346, 1030]. order [891, 2127, 313]. Oregon [800, 1095, 1841, 245, 782, 1226, 1154, 283, 94, 921, 573]. Oreochromis [1853, 1485]. organic [1236, 1579, 1599, 401, 2100, 306, 1802, 960, 767, 1089, 1801, 502, 1100, 91, 997, 653, 1053, 742, 709, 2073, 1938, 549, 1678, 574, 1262, 1150, 1807]. organisation [334]. organisms [1320, 1557, 471, 335, 1169, 1964, 832, 325, 1269]. organization [2081]. organizing [1500, 1261]. Organochlorine [128, 1908]. organs [84, 984]. origin [2036, 1757, 524, 1639, 160, 1904, 832, 931, 80, 1163, 1355]. originating [437]. origins [2014, 1182]. Oscillation [1054, 971, 1971]. Osmerus [1562, 163, 1701, 444, 1070, 1683]. Osmoregulation [1295]. Osmoregulatory [870, 1075]. ossification [36]. ostariophysan [622]. Ostenfeld [62]. Otago [1186]. Otaria [982]. other [793, 634, 1795, 243, 471, 335, 1332, 1333, 1309, 1747, 259]. othonopterus [1316]. Otolith [1744, 2122, 584, 552, 1312, 644, 372, 1714, 1566, 1929, 1601, 1267, 1719, 665, 1577, 728, 947, 923, 1251, 1963, 1695, 913, 945, 460, 1447, 1909, 1832, 1326, 1258, 1952, 1120, 216, 1065]. Otoliths [260, 1047, 1273, 93, 846, 1162, 1441, 1577, 450, 2014, 561, 1112, 922, 585, 871, 290, 1516, 835, 1290, 813, 422, 1754, 125, 1422, 131, 1710]. otter [1794, 355]. otter-trawl [1794]. our [2025, 1925]. outbreeding [1285]. outburst [2157]. outcome [2154]. outcomes [1857]. outflow [271]. outliers [230]. output [1680]. ova [842, 1628]. ovary [315]. Over-winter [966]. overcoming [941]. overcompensation [2064]. Overcompensatory [1923]. overdispersion [2123]. overexploitation [1133, 1989]. overexploited [1424, 1574]. overfished [1107, 1724]. overfishing [1505, 503, 66, 970]. overlap [246, 171, 2020, 1843]. overlooked [1596]. overoptimism [1655]. overturns [1781, 2012]. overwinter [1488]. Overwintering [1676, 1637, 572, 327, 2016]. ownership [573]. oxidation [1678]. oxidative [1734, 1108]. Oxygen [2047, 992, 1913, 1627, 1876, 1214, 1726, 1853, 415, 2140, 130, 353, 1927, 1034]. oxyrhynchus [1385]. oxyrinchus [747]. oxytetracycline [290]. oyster [84]. oysters [352]. ozone [61].

P [893, 904, 895, 1467]. P450 [41]. p53 [1239]. pê [950]. pêches [1122]. paced [955]. Pacifastacus [1292, 1389]. Pacific [1649, 173, 1854, 1549, 1928, 682, 1948, 509, 659, 660, 1984, 1887, 2154, 21, 450, 2102, 323, 2043, 382, 681, 1606, 1723, 1749, 1875, 1210, 1825, 649, 2096, 858, 865, 832, 2092, 1179, 799, 1844, 1483, 1412, 127, 1310, 540, 755, 145, 19, 1960, 395, 1615, 784, 1145, 1673, 536, 227, 513, 1870, 939, 305, 687, 792, 548, 1069]. pacificus [962]. pack [180]. Pagrus [1112]. PAHs [1525, 1433]. pair [1907]. paired [1000]. pairs [2020]. palaeolimnological [1048]. paleoecological [1812, 1917]. Paleoenvironmental [371]. paleolimnological [739, 983, 410, 301, 302]. Paleolimnology [1289]. Pallas [65]. pallasi [509, 450, 382, 681, 799, 536, 227, 513]. pallial [84]. palmitate [801]. Pan [1283, 1693, 1550]. panacea [1167]. Pandalid [956]. Pandalus [243]. Panopea [1948]. pantophysin [1283]. Papua [340]. Paracentrotus [708]. paradox [872]. Paralichthys [188, 1015]. Paralithodes [836, 1631]. parallels [1748]. parameter [1426, 398, 924, 19]. Parameterizing [1868]. parameters [1666, 1567, 166, 2063, 1755, 348, 243, 1586, 1314, 537, 467, 1077, 2162, 1934, 1019]. parasites [157]. parasitic [1265, 939]. Parasitism [1308]. Parastichopus [419]. Parentage [423, 1, 524]. parental [1823, 2045, 2044]. Park [232, 2101, 128, 187, 516]. Pärnu [45]. parr [891, 1240, 815, 1951, 1281, 642, 185, 1163, 164]. Parsons [1544, 1544]. part [1693, 1572, 1573]. partial [2022]. particle [1620, 1598, 897]. particle-associated [897]. Particulate [4, 899, 84]. partie [1799]. Partitioning [812, 1928, 118, 1204, 443, 1669, 1974, 2057]. parts [1492]. Parvicapsula [1245]. parvimensis [419]. pass [98, 563]. passage [975, 1654, 2147, 1492, 974, 1332, 627]. passing [1867]. passive [135]. Past [1007, 34, 153, 306, 944, 920, 30]. pasture [753]. Patagonia [931, 982, 1192]. Patagonian [942, 1312]. patch [1141]. patches [1382, 1195, 629]. patchiness [753]. paternity [1581]. path [359, 914]. pathogens [589]. pathway [116]. pathways [1713, 401, 1098, 1840, 1481, 249, 1930, 1187, 1546]. pattern [333, 1642, 138, 1763, 2075]. Patterns [86, 344, 458, 1501, 2097, 609, 1828, 1554, 704, 2086, 106, 1460, 903, 1510, 1317, 682, 607, 1910, 1865, 291, 1607, 626, 1860, 1629, 1341, 741, 482, 1135, 1162, 1630, 412, 120, 1411, 1321, 1939, 1736, 2067, 781, 1608, 442, 1789, 1383, 1342, 1841, 757, 649, 1632, 101, 103, 837, 1004, 1899, 1116, 813, 1494, 1786, 816, 1486, 797, 1242, 1217, 2094, 177, 2037, 968, 1376, 1450, 550, 1623, 1497, 1232, 603, 1030, 573, 2011, 176, 1748, 1343]. Patuxent [2022]. pay [497]. Payne [2045]. pealeii [336, 1383, 1944]. peat [1011]. peatlands [222]. pedigree [656]. Pelagic [1087, 2157, 134, 761, 1705, 1745, 1605, 539, 246, 1657, 1717, 1518, 389, 1519, 1725, 1414, 963, 1697, 683, 1280, 1746, 193, 73, 1187, 1546, 1937, 1161, 2075, 1146, 792, 1110, 409]. penaeid [421]. Penaeus [465]. penetration [306]. peninsula [1442, 1186]. Pennsylvania [964]. per-recruit [1287]. peracarids [1104]. Perca [711, 1300, 1579, 1298, 1321, 1322, 984, 1108, 253, 1711, 1892, 1147, 393, 1071, 1366, 1871, 172, 1691, 75, 485, 917, 1707, 1674, 618, 44, 1253, 1859, 1021]. perception [1641, 317]. perceptions [267, 147]. perch [711, 1300, 1579, 1298, 1321, 1322, 984, 1108, 253, 1711, 1892, 1147, 388, 393, 1825, 2022, 1071, 1366, 1871, 172, 1691, 75, 485, 1754, 917, 1707, 1674, 618, 44, 1253, 1859, 1021]. percid [874, 1375]. perennially [481]. Performance [1990, 2140, 693, 1517, 1436, 1249, 2155, 1118, 1645, 347, 2154, 1349, 974, 750, 2000, 1873, 1766, 1352, 1307, 360, 714, 618, 2061, 1772, 1245]. performing [59, 972]. period [1209]. periodicity [2035]. periods [1275, 1486]. Periphyton [1584, 1178, 1237, 1358, 1752, 1471, 1687, 8, 1496, 550, 591, 456]. permafrost [2110]. Permutation [316]. peroxide [1501, 967]. Persistence [1715]. Persistent [1236, 1530, 1079, 858, 2075]. personal [2126]. perspective [1104, 1868, 1145, 1733, 1484, 1001]. Perspectives [551, 1709, 611, 265, 1812]. perturbation [1752]. perturbations [938]. Peru [899]. Petersen [1365, 1837]. Petromyzon [1690, 1826, 57, 1409, 1694, 1830, 1940, 1265, 1266, 1776, 689, 1750, 478, 1361, 1670]. pH [622, 1876, 878, 1659, 700, 937, 190, 105, 549, 1262, 1807]. Ph.D. [1400, 1399]. Phalacrocorax [1986]. phase [59, 2087, 1265, 1760, 555, 825, 1817]. Phenological [1629]. phenology [475]. phenotypes [1632, 885]. pheromone [57, 689, 478]. pheromones [1361]. philippinarum [773]. Phocarctos [735]. phosphate [571, 1086]. Phosphorus [1941, 1209, 1045, 688, 650, 1584, 1073, 1059, 969, 200, 579, 1421, 1087, 2114, 601, 1998, 948, 154, 1954, 339, 105, 1467, 126, 1700, 1807]. photo [1678]. photo-oxidation [1678]. Photodegradation [1150]. Photodegradation-induced [1150]. photodegraded [1599]. photogrammetry [1589]. Photoinhibition [38]. photoperiod [1587, 1572, 1573, 1585, 1314, 1315]. photoreactivity [91]. Photosynthesis [1927, 1932]. photosynthetic [2063]. Phoxinus [1900]. phylogenetic [1967, 69]. phylogeographical [1748]. Phylogeography [2150, 464, 863, 173]. Physeter [23]. Physical [321, 778, 1975, 792, 1110, 1076, 162, 904, 2081, 1738, 1541, 1633, 1380, 1956, 1845]. physical-biogeochemical [904]. physicochemical [1229]. physics [1820]. Physiological [170, 697, 1561, 1686, 1410, 347, 434, 2002, 612, 1525, 534, 769]. Physiology [1448, 1572, 1573, 621, 1461, 1570]. physonect [756]. Phytoplankton [217, 1738, 2094, 1455, 383, 1811, 1268, 739, 1413, 4, 1823, 2045, 601, 528, 896, 502, 162, 999, 1514, 38, 691, 1992, 1806, 2044, 1570, 319, 376, 1878, 326, 1206, 623, 87, 10, 1013, 456, 61, 2139, 752]. Pictou [1239]. Pierre [1831]. pigment [1214, 1013]. pigments [961]. pike [1199, 40]. pikeperch [45]. Pimephales [1599, 569]. pine [1010]. pingers [1976]. pink [357, 1887, 1420, 1324, 845, 395, 1708, 989, 666, 651, 1050, 209, 1625]. pinniger [1154]. pinniped [93]. pinto [1959]. piscivore [56, 1264, 1470, 1387, 1894]. piscivores [58, 1250]. piscivorous [827, 864, 740, 1511, 1870]. piscivory [1300, 1235, 1383, 2037]. PIT [135]. pituitary [1147]. placement [283]. placements [709]. Placopecten [476, 2033, 279, 954, 1215]. plaice [1618, 1166, 1134]. Plain [1230, 739, 296, 1601, 376, 14]. Plains [569, 2094, 1362, 1751]. Planktivore [2015, 2119]. planktivorous [827, 619]. plankton [725, 969, 451, 937, 326, 123, 1930, 1715]. Planktonic [798, 1675, 571, 118, 4, 323, 748, 811, 49, 198, 1450, 1148, 831]. planning [1458]. plans [1995, 1955]. plant [887, 136, 1148]. plants [820]. plasma [1282, 1359, 1273, 945, 125]. Plastic [1805]. plasticity [1587, 643, 1909, 667, 870]. Plata [116]. platessa [1618, 1166, 1134]. platforms [645]. play [1540, 1332, 1333]. Pleasant [964]. Plectropomus [146]. Pleuronectes [1618, 1166, 1134]. Pleuronectiformes [634]. plumbeus [1581]. plume [913]. pocket [114, 2151]. Poecilia [1130]. Poeciliopsis [378]. point [577]. Points [950, 953, 68, 952, 956, 520, 457, 957, 2003, 958, 951, 954, 780, 751, 950]. polar [842]. poleward [1549]. policies [2099, 1466, 1869]. policy [1815]. pollan [1308]. pollock [2155, 1283, 1629, 1311, 1604, 2055, 2000]. pollutants [1236]. polluted [303, 493]. pollution [1460, 107]. polychlorinated [118, 2153, 1169, 1373, 1674]. polycyclic [1525, 1433]. polymorpha [562, 818, 1401, 60, 389, 1229, 1946, 1769, 101, 65, 75, 485, 811, 876, 165, 697, 916, 370, 1109]. polymorphism [929, 294]. Polynya [1459, 1013]. polynyas [1820]. polyphosphate [339]. Polyploid [2098]. polyunsaturated [1602]. Pomatomus [1488]. Pomoxis [1024, 807]. Pond [1965]. ponds [2110, 2052, 1396, 217]. Ponto [607, 1104, 609, 606, 610]. Ponto-Caspian [607, 609, 606, 610]. Pontoporeiidae [2084]. Pool [1401]. pooling [1223]. pools [551, 904, 152, 311]. poor [83]. pop [1506]. pop-up [1506]. Population [1199, 36, 626, 863, 203, 1935, 1437, 2087, 2104, 1644, 163, 65, 1844, 1420, 1324, 1767, 1224, 1286, 173, 726, 939, 216, 1127, 2025, 312, 1744, 1498, 1, 615, 1197, 308, 1377, 495, 2112, 659, 1200, 1984, 1581, 874, 982, 1111, 1929, 855, 1153, 281, 166, 174, 1022, 1041, 1523, 52, 1641, 243, 352, 944, 851, 3, 1385, 237, 1682, 988, 1857, 1305, 1133, 526, 1574, 1357, 366, 2048, 1327, 48, 624, 885, 398, 1375, 1004, 661, 1084, 799, 426, 147, 986, 1251, 392, 2133, 943, 1369, 2120, 796, 1097, 797, 1743, 1139, 1832, 1759, 747]. population [292, 531, 1673, 421, 1600, 1872, 1551, 73, 1288, 1138, 563, 1923]. population-specific [2048]. Populations [1031, 312, 1012, 2160, 1426, 1118, 1032, 63, 634, 1301, 580, 1371, 1568, 1268, 2153, 260, 783, 1184, 1211, 524, 957, 1152, 1616, 817, 882, 1397, 1874, 561, 1880, 443, 1112, 2099, 1479, 1437, 471, 335, 593, 1663, 47, 1933, 1841, 679, 1964, 1989, 2118, 1194, 896, 1618, 1632, 398, 1250, 1922, 1691, 1126, 775, 1747, 533, 20, 444, 993, 782, 1037, 1132, 1226, 1128, 683, 8, 177, 40, 2076, 1673, 1302, 240, 1416, 2031, 687, 72, 1985, 2161, 1945, 842, 1628, 656, 201, 655, 777, 176, 1748]. porbeagle [1567]. porpoises [1976]. Portugal [1921]. portusjacksoni [1740]. position [1236, 1967, 1292, 1160, 840, 298, 1924, 1125, 73]. positioning [1772]. Positive [502]. Possible [900, 1585, 1438, 1775, 1662]. post [410, 459, 323, 2032]. post-1970 [410]. post-1976 [323]. post-Wisconsinan [459]. Postconsumption [1883]. posterior [1278, 467]. postglacial [2150, 626, 1442, 1847, 667, 1224]. postlarval [76]. postrelease [1815]. postsmolts [120, 1182, 1686]. Potential [2067, 138, 1226, 786, 1888, 671, 36, 756, 157, 1287, 180, 952, 820, 1067, 200, 2088, 744, 569, 886, 1957, 15, 1944, 636, 1265, 366, 1725, 1591, 244, 1403, 2003, 1238, 1429, 377, 1344, 930, 31, 1433, 909, 273, 655, 1485]. potentials [868]. pour [950, 1544]. pout [1595, 2024]. poutassou [2034]. power [836, 1594, 765, 789, 576, 1653]. Practical [1142, 1650, 1955, 656]. prairie [1242, 2094, 9, 752]. prawn [1111, 421, 734]. pre [1659, 323]. pre- [323]. pre-industrial [1659]. prealpine [1261]. Precambrian [1811, 1246, 1810, 742]. Precaution [955]. Precautionary [266, 957]. preceding [1075]. precipitation [1998, 541]. Precipitous [1959]. precise [115]. precision [2025, 728, 829, 1689, 398, 1462]. Precocial [226]. precocious [644]. Predation [1424, 246, 690, 1886, 555, 1201, 182, 140, 365, 1986, 940, 156, 1639, 1321, 1537, 1717, 2115, 432, 1351, 1508, 1518, 330, 1222, 2024, 1698, 977, 289, 1179, 1469, 2119, 504, 737, 139, 1294, 790, 342, 1408, 675, 1215, 2007]. Predator [1027, 1021, 1343, 1127, 89, 12, 1221, 1264, 2041, 1284, 998, 1222, 1368, 848, 892, 762, 1636, 1280, 1511, 1930, 1503, 73]. Predator-dependent [1027, 2041]. predator-prey [1222]. Predator-recognition [1021]. predators [1235, 1257, 1624, 1691, 343, 2050, 425, 2037, 1651, 1807]. Predatory [599, 1046, 1018, 447, 759, 662]. predict [505, 1646, 2140]. Predicted [165, 1926, 2108, 1981, 781]. Predicting [248, 938, 1979, 774, 431, 718, 388, 85, 959, 98, 2016, 1771, 1684, 1270, 1858, 974, 1846, 1990, 386, 1894, 1103, 1656]. Prediction [263, 4, 166, 1882]. Predictions [375, 624, 1576, 1921, 768]. Predictive [1758, 1306, 1349, 387]. predictor [1114]. predictors [1255]. predicts [1759]. preference [699, 849, 1625]. preferences [1979, 1936]. preferred [327]. preindustrial [486]. preliminary [660, 1881]. premature [1686]. presence [63, 1599, 2154, 822, 1080, 235, 1982]. present [1402]. Preservation [879]. presettlement [1297, 884]. prespawning [1615, 962]. Press [926]. Presses [926]. pressure [1144]. Preston [804]. Prevalence [1969]. preventing [1861]. previous [1278, 1942]. Prey [362, 390, 1127, 759, 827, 699, 1264, 1470, 2041, 1492, 998, 1508, 1886, 1222, 1402, 343, 848, 1704, 436, 892, 1753, 1217, 1843, 2037, 1721, 962, 293, 1651, 675, 1343]. prey- [2041]. Primary [402, 2122, 528, 1414]. primiparous [678]. Prince [357, 509, 649, 536]. principles [668]. Prionace [1377]. prior [467]. prioritize [921]. prioritizing [1940]. priors [230, 429, 637, 1151]. pristine [222]. probabilistic [1475, 2117]. probabilities [1278, 1993, 2005, 26, 627]. probability [1885, 847, 266, 114]. probable [1807]. problematic [2107]. procedures [158]. Process [1951, 1431, 241, 796]. Process-based [1951]. processes [321, 1079, 333, 495, 401, 1864, 2041, 1612, 1632, 1096, 994, 1615, 1965, 350]. processing [180, 300]. Production [359, 903, 1948, 716, 68, 53, 586, 648, 1863, 2080, 2112, 1413, 1523, 352, 967, 1381, 1851, 310, 528, 2009, 636, 430, 1216, 2013, 1042, 900, 676, 225, 1396, 1414, 865, 1279, 1386, 1760, 959, 1202, 796, 169, 559, 2042, 1746, 1391, 1945, 11, 32, 641]. production-age [1948]. Productive [764, 1722, 695, 1417, 169]. Productivity [369, 1941, 783, 1762, 1887, 1537, 1538, 1179, 2003, 782, 946, 19, 760, 8, 2094, 177, 1338, 1956, 87, 10, 176]. productus [1549, 682, 1723, 145, 784]. profiles [1282, 1359, 1643]. profiling [1452, 479]. profundal [228, 1169, 2021]. progeny [1822, 1916, 1677]. progesterone [853]. program [1757, 181, 1226, 296, 1354]. programmes [238]. programming [37]. programs [1003, 1092, 1420, 1324]. project [627, 2126]. Projected [1386, 153]. promelas [1599, 569]. promote [1250]. promoted [370]. promotes [826]. Pronounced [1775]. propagation [927]. propagule [1144]. Properties [489, 2110, 1595, 79, 507, 1878]. proportional [418]. proportionality [644]. proportions [1247]. proposals [694]. proposed [1993]. propulsion [247]. Prosopium [1415, 1468]. prospective [754]. prospects [1832]. Protandry [127]. protected [1118, 1568, 1675, 1374, 1449, 269, 1167, 1680, 1545, 1179, 1798, 1181, 1656]. Protecting [1449, 875]. protection [1392]. Protective [1215]. protein [922, 642, 730, 870]. proteins [1495, 5, 1239]. protogynous [309, 851, 861]. protracted [317]. provide [306, 885, 1774, 286]. provisioning [1628]. proximate [1082]. Pseudo [1883]. Pseudo-nitzschia [1883]. pseudoharengus [1531, 1986, 1267, 1852, 1394, 1683, 2006]. Pseudomugil [426]. Pseudomugilidae [426]. Pseudopleuronectes [463, 1760, 790, 1201]. Puerto [1330]. Puget [1910, 500, 450, 989, 2076]. pulp [315, 137, 574]. pulse [461]. pulsed [1129]. pumpkinseed [36]. punctatus [1447]. punctipennis [1102]. Pungitius [184, 726]. pup [2042]. pupping [26]. Pure [1439]. purpuratus [1591].

quagga [562, 818]. Quality [346, 1460, 1299, 719, 1739, 1178, 1502, 983, 1816, 991, 1349, 814, 1602, 944, 1831, 822, 647, 1725, 1899, 402, 217, 1924, 486, 1166, 296, 884, 702, 696, 1626, 1348, 194]. Quantification [2143, 1556]. quantify [1116]. Quantifying [776, 975, 1840, 1898, 721, 1521, 343, 924, 892]. Quantitative [920, 1814]. quantity [1602]. Quasi [312]. Quasi-cycles [312]. Quebec [1831, 130, 1423, 1697, 80, 2100, 1442, 1430, 1687, 1306, 1500]. qui [1799]. quotas [1475, 1855, 1466, 1449, 377, 576].

R [1544]. raceways [1040]. radiated [504]. radiation [571, 1829, 601, 528, 1792, 542, 38, 1704, 1102, 961, 445, 32, 61, 2139]. radical [705]. radio [1143, 1060]. radio-tagged [1143, 1060]. Radiocarbon [871, 1795, 1825, 1154]. radioisotope [253]. radiometric [372, 728]. radiotelemetry [439, 627]. radiotracer [44]. radium [2109, 2069]. raft [353]. raft-culture [353]. raggedtooth [1785]. Rainbow [1788, 2093, 530, 966, 2137, 1562, 570, 1822, 1978, 2014, 1115, 1722, 821, 526, 284, 1093, 1194, 1353, 163, 508, 1701, 1913, 698, 2071, 1440, 917, 945, 2061, 2135, 931, 1575, 1982, 461, 888, 1070, 1683, 1433, 1350, 801, 1262, 1688]. rainfall [1678]. raised [1835]. Rajidae [1411]. random [1582, 1668, 856, 2099, 1819]. Range [946, 607, 1025, 1439]. ranking [1475]. Ransom [1731, 1732]. Rapid [472, 412, 601, 335, 374, 600, 1368, 602, 1069, 36, 1970, 1940]. Rapide [374, 412, 601, 335, 600, 602]. Rate [1611, 1779, 1778, 198, 1114, 544, 1489, 975, 794, 1719, 1135, 2144, 488, 1180, 1039, 1492, 310, 593, 1658, 330, 578, 1624, 429, 1913, 1905, 534, 33, 1924, 369, 1982, 819]. rates [2077, 2019, 2023, 1705, 354, 68, 1401, 53, 2078, 453, 1098, 1565, 990, 253, 1717, 2115, 102, 97, 528, 1875, 1490, 1933, 2048, 101, 1997, 300, 597, 1345, 1491, 845, 540, 755, 105, 504, 860, 223, 395, 797, 932, 1884, 177, 1983, 1380, 428, 1678, 1854, 227, 513, 88, 42, 1245, 1804, 1120, 176]. ratio [146, 1390, 794, 1711, 676, 452, 364, 678, 1253]. ration [89, 12]. ratios [648, 1267, 505, 997, 813, 1417, 169, 1790, 125, 1065]. ray [43]. Reach [168, 167, 1263, 1864, 47, 732]. Reach-scale [1263, 47]. reaches [1787]. Reaction [35, 1977, 2117]. reactions [1914]. reactive [1231]. reading [1898]. readings [1000]. realistic [1113]. really [102]. reared [1734, 332, 749, 2038, 1671, 188, 1624, 400, 917, 1037, 1695, 1185, 1021, 182]. rearing [1047, 64, 1061, 1396, 1438, 369, 94, 287]. rebuilding [1107, 1196, 1869, 429]. recapture [2078, 859, 1892, 604, 590, 1873, 25, 661, 597, 1345, 663, 26, 428, 1288, 2103, 874]. receivers [1772]. reception [1695]. receptor [1688]. recipe [979]. reciprocity [1102]. recognition [1021]. Recommandations [277, 515]. Reconciling [970]. reconstruct [760, 584, 552]. reconstructed [153, 1887]. Reconstructing [202, 659, 660, 1962, 585, 244, 1009, 1155]. Reconstruction [93, 1696, 1577, 529, 1456, 371, 1922, 510, 587, 920]. reconstructions [306]. record [1312, 993]. recorder [326]. records [1291, 34, 721, 1710]. recover [1191]. recovered [403, 1493, 1276]. recoveries [597, 603]. recovering [281, 566, 1089, 1691, 1945]. Recovery [982, 424, 315, 20, 1451, 1460, 650, 1457, 1424, 1765, 1191, 1489, 472, 434, 886, 2128, 1291, 1255, 1133, 1658, 1574, 366, 1250, 1368, 483, 484, 1955, 1480, 1338, 92, 2129, 1001, 408]. recreational [1052, 1377, 1722, 2116, 1752, 147, 1097, 1416]. recruit [1287, 1996, 990, 476, 328, 307, 244, 1403, 1203, 924, 1491, 317]. Recruitment [1510, 1284, 2024, 1649, 503, 762, 13, 248, 67, 1579, 46, 1562, 1054, 971, 285, 512, 1390, 241, 1863, 189, 454, 367, 520, 638, 659, 766, 1252, 517, 2144, 120, 1823, 2045, 1180, 1322, 511, 1363, 569, 393, 1049, 1852, 844, 843, 48, 757, 396, 397, 2096, 2000, 1279, 2018, 1846, 1921, 795, 1701, 145, 583, 1218, 1131, 1775, 980, 124, 2044, 1486, 1166, 559, 538, 317, 2105, 1050, 134, 261, 1232, 30, 164, 2011, 1980]. recurrence [1969]. recycling [433]. Red [730, 288, 836, 2097, 15, 1167, 100, 545, 560, 1934]. redbelly [1900]. redd [1797, 1885, 167, 834, 763, 1342, 1033]. redd-scale [1033]. redds [1446, 96, 2151]. redefined [2104]. redefining [354]. redfish [1566, 1258]. redhorse [1890]. redox [105]. Redside [286]. reduce [1146]. Reduced [2061, 622, 878, 3, 591, 1553, 1700]. reduces [1370]. reduction [1531, 1520, 1336]. reductions [410]. redundancy [417]. Redwood [1438]. reef [1706, 1118, 873, 851, 1516, 645, 1832, 1603, 146]. reefs [146].

Reevaluating [976]. Reference [950, 68, 637, 953, 952, 956, 520, 457, 957, 990, 1500, 871, 2003, 1707, 958, 951, 954, 780, 751, 950]. reflectance [1155]. reflects [626, 213, 1472]. refuge [1195, 286, 2007]. refuges [1215]. refugia [221, 421]. regalis [2004, 997, 720]. regarding [267]. regards [467]. regime [2155, 682, 2015, 1685, 1649, 683, 2054, 1634]. regimes [140, 454, 2091, 2081, 425, 726]. region [1910, 609, 1856, 486, 282, 252, 1289, 1804, 14]. Regional [1174, 1610, 624, 101, 1808, 550, 1653, 1052, 638, 743, 617, 1989, 301, 302, 504, 1072, 462]. regions [1257, 1090, 28]. regression [316, 17, 1091, 1817]. regular [677]. regulate [1799, 1800]. regulated [1475, 1787, 1943, 959, 698, 576]. regulates [2034]. regulating [1572, 1573, 162, 2105, 549]. regulation [1317, 1374, 1618, 946, 1138, 2139]. regulations [1857, 1815]. regulator [1678]. régulent [1799]. rehabilitation [959]. rehydration [881]. Reid [1779]. Reinhardtius [1644, 1821]. reintroducing [1835]. reintroduction [657, 2130]. Related [893, 682, 39, 1041, 1809, 1235, 1789, 2021, 1468, 1624, 215, 1525, 1798, 13, 1891, 41]. Relating [1601, 1958, 1166, 1956, 759, 1894]. relation [1073, 179, 1459, 1039, 1711, 1432, 2087, 1749, 2093, 676, 172, 82, 1810, 698, 217, 559, 899, 678, 29, 717, 1017, 204, 1628]. relations [1341]. Relationship [210, 1971, 783, 991, 1698, 1614, 31, 1065, 1549, 1300, 793, 226, 1284, 1954, 17, 244, 2084, 1139, 317, 652]. Relationships [1547, 501, 767, 1801, 573, 752, 527, 46, 1234, 161, 189, 454, 367, 1665, 1635, 579, 1421, 529, 594, 299, 1491, 1218, 1507, 1770, 1458, 538, 40, 391, 1450, 1870, 185, 2131, 1186]. relative [2036, 1299, 1739, 524, 1610, 1994, 1639, 840, 1530, 765, 2009, 828, 1119, 1366, 1995, 917, 1159, 504, 2037, 1232]. release [1805, 2100, 1842, 238, 105, 1815, 689, 92, 577, 2103, 1923]. released [57, 2137, 182]. relevance [273]. relevant [738]. reliable [1114, 1626, 2072]. reliance [1293]. relict [1479, 464]. relicta [1979, 438, 1966]. remains [793]. remediation [34]. Remote [1598, 1079, 1742]. removal [1368, 1084, 98, 1743, 1496, 722, 563]. removals [921]. reorganizations [1206]. repeat [604]. replacement [977, 1786]. replicate [1342]. replicated [1066]. Reply [512, 1779, 2045, 470, 804, 197, 1419, 177, 2161, 156, 449]. Reporting [1489, 2078, 488]. representation [927]. Reproduction [1957, 521, 1853, 82, 2017]. Reproductive [2036, 366, 749, 1299, 580, 1020, 21, 1397, 501, 16, 765, 1680, 578, 244, 1403, 727, 2003, 1134, 1583, 1631, 850, 1125, 655, 1681]. reproductively [842]. requirement [1072]. requirements [327, 1206, 1198]. requires [2147]. rerio [1238]. resampling [2000]. Research [926, 2125, 940, 399, 297, 107]. Reserve [1009, 535, 580, 1427, 986, 702]. reserves [1066, 1424, 2064, 1137, 1445, 1418, 1911, 1658, 1140, 82, 1002, 1943, 1783, 1014, 1419, 380, 678, 419, 1331, 930]. reservoir [337, 1362, 1121, 1930, 1684, 110]. reservoirs [458, 2132, 1360, 965, 217, 1090, 762]. residence [25, 1034]. residency [1275, 1162]. resident [1822, 524, 1916, 1692, 1899, 724, 1818, 1062]. residential [1834, 1958, 1511]. residents [857]. residual [1157, 1277]. resilience [753]. Resistance [589, 753, 612]. Resmoltification [1075]. resolved [901]. resolves [95]. Resolving [357, 2079, 396]. resonance [1467, 1838]. Resource [443, 987, 1141, 1206, 1100, 1847, 2037, 962, 404, 2057]. resources [1158, 228, 1260]. respiration [1719, 2063, 1414, 49, 1927, 1932]. Respiratory [620, 360]. respond [1679, 548]. Response [1012, 440, 1425, 1834, 1809, 75, 358, 1511, 1700, 571, 1165, 84, 1584, 1301, 1283, 1647, 1087, 855, 2147, 56, 1978, 183, 1237, 1358, 880, 323, 432, 15, 1518, 339, 1100, 564, 1368, 2152, 1514, 775, 993, 436, 1806, 1975, 283, 8, 989, 2049, 869, 591, 405, 1503, 1937, 722, 1742, 1839, 1534, 1688, 1923]. Responses [896, 709, 1052, 2004, 1661, 1471, 1027, 233, 1057, 1592, 1435, 1794, 330, 1535, 773, 1525, 674, 2133, 892, 1072, 1850, 1139, 2054, 543]. responsive [329]. responsiveness [89, 12]. resting [1157, 1876]. restocking [879]. restoration [1457, 1329, 1535, 423, 1609, 1429]. restricted [1973]. restrictions [1438]. resulting [1797, 145, 325]. results [1142]. retain [1327]. retardant [1460, 1337]. Retention [1254, 1737, 847, 250, 154, 1086, 1615, 797]. reticulata [1130]. Retrospective [482, 2082, 1766, 754]. return [407, 2077, 523, 1686]. returning [237, 1960]. returns [1272, 1205]. reveal [2098, 1402, 2031]. revealed [681, 438, 1909, 1870, 479]. reveals [2104, 605]. review [947, 1896, 382, 694, 909]. revisited [275, 1586, 1600]. Revisiting [1672, 1270]. revival [434]. RFSC [686]. Rhône [874]. rheophilic [79]. Rhine [71, 1700]. Rhinichthys [1442, 714]. Rhinogobius [452]. Rican [1330]. rich [437, 891, 1204, 767, 1329, 1010]. Richardsonius [286]. richness [527, 1994, 840, 1381, 785, 1841, 1503, 204]. Ricker [686, 578, 467]. Ridge [1928, 305]. riffle [1329, 152, 1083]. riffle-scale [1083]. riffles [311]. right [1106, 2090, 2090]. Rim [1870]. ring [1259, 1948, 1441, 1773]. rings [2033, 234, 1918]. Río [116]. riparian [1540, 1057, 551, 1105, 1679, 653, 545, 70, 1263, 31, 732, 1062]. riparius [1602]. Risk [1557, 1153, 616, 1053, 409, 246, 1718, 266, 1518, 267, 330, 802, 1210, 1698, 2053, 1179, 1136, 2073, 1269]. River [2160, 112, 411, 345, 1797, 680, 2, 1006, 673, 1864, 588, 108, 2082, 226, 1555, 1831, 1143, 2022, 1396, 444, 1992, 1303, 556, 2120, 1687, 1633, 931, 1444, 591, 666, 1985, 2161, 107, 201, 1410, 477, 1236, 1967, 1739, 1584, 1025, 1661, 168, 1330, 1530, 1787, 1608, 823, 636, 1473, 1521, 1004, 1038, 994, 959, 698, 1494, 1974, 111, 1890, 1975, 334, 1541, 137, 439, 914, 1603, 110, 541, 575, 809, 1415, 1587, 63, 1401, 39, 1654, 1448, 482, 344, 168, 167, 1694, 834, 307, 186, 887, 572, 1229, 2038, 1143, 1950, 423, 1766, 825, 1060, 1159]. River [565, 789, 483, 484, 425, 369, 712, 2060, 387, 386, 932, 1316, 99, 1354, 136, 916, 1700, 1245, 30, 41]. river-reservoir-wetland [110]. riverine [1088, 1482, 1850, 1743, 1824, 1772, 1198, 563, 777, 1914]. rivers [1073, 1178, 1007, 500, 109, 2063, 1432, 833, 1332, 1333, 1538, 1716, 1476, 1310, 1969, 1043, 1790, 135, 177, 176, 1681]. RNA [1711, 2062, 1253]. RNA/DNA [1711, 1253]. roach [1957, 750]. road [1989]. Robust [1492, 235, 467]. robustness [1642]. Rock [381, 652, 488, 487, 1196, 218, 280]. rockfish [1052, 626, 863, 1252, 1223, 2104, 1842, 871, 875, 2018, 2056, 533, 986, 1251, 1154, 1838, 1258, 1724]. rockfishes [173]. rocks [1834]. Rocky [1758, 961]. Role [56, 162, 799, 1356, 1690, 1826, 2106, 1572, 1573, 1540, 1264, 1137, 442, 1481, 647, 1591, 1943, 695, 1704, 460, 1815, 1590, 1346, 1836, 1678, 961, 478, 809]. roles [333, 270, 757, 1692, 977, 504, 1781, 2012]. rostrata [825]. rotational [235]. Rouch [158]. roughness [1016, 1527]. roughy [2109, 2069, 1505]. round [1059]. Route [1473, 627, 1410]. route-specific [627]. routine [275, 771]. royal [1186]. ruberrimus [871]. rubidium [1164]. rubra [545]. ruffe [1153]. Rule [1304, 624]. rules [735]. Run [1696, 589, 1587, 1448, 1456, 1523, 237, 1916, 1950, 1901, 651, 1561, 1686]. running [802]. runoff [1802, 190, 31, 541]. runs [1960]. rural [183]. Russia [2031]. Russian [616, 1303, 194]. rusticus [1351, 1346, 1030]. rusty [1351, 1346, 1030]. Rutilus [1957, 750].

Sable [26]. sablefish [129, 377, 854]. sac [808]. Sacramento [1587, 39, 1700]. safe [1223]. sagax [2043, 1216, 2013]. Saginaw [999, 370]. Saguenay [187, 516]. saida [1459]. Saint [1831, 80, 1289]. Saint-Jean [80]. Saint-Pierre [1831]. salamander [1085]. salar [1593, 1734, 644, 1498, 50, 1865, 1032, 1761, 207, 1607, 857, 1007, 891, 1977, 1661, 1240, 791, 1036, 1156, 1041, 878, 621, 738, 2082, 2142, 1874, 814, 1039, 368, 208, 815, 1951, 363, 690, 2009, 733, 1042, 1281, 1044, 1061, 1908, 1247, 1182, 1904, 1033, 1524, 400, 1035, 1873, 1446, 1028, 642, 1295, 706, 518, 1038, 2002, 1905, 1614, 1542, 555, 327, 1969, 1469, 1043, 1045, 1037, 862, 1879, 1040, 712, 1224, 704, 523, 80, 2031, 72, 1895, 273, 479, 1163, 164, 1063]. Saldanha [353]. saline [2094]. Salinity [1514, 665, 1297, 572, 1482, 2152, 835, 361, 1486, 1009, 920, 1625, 1065]. Salmo [210, 1593, 1114, 1734, 644, 1498, 50, 1865, 1032, 1761, 207, 1607, 857, 495, 1007, 891, 1200, 1977, 1661, 1240, 791, 1036, 1156, 1041, 878, 621, 738, 2082, 2142, 1874, 814, 1039, 368, 208, 815, 1951, 214, 427, 363, 690, 2009, 823, 494, 1886, 733, 1042, 1281, 1044, 1671, 1061, 1908, 1247, 1182, 1904, 1033, 1524, 400, 1035, 1873, 1446, 1028, 1004, 1279, 1172, 1692, 1899, 642, 1295, 706, 518, 1038, 2002, 1307, 1905, 1614, 1542, 555, 327, 1969, 1469, 1043, 1045, 71, 468, 1037, 1786, 862, 1609, 1879, 2016, 1040, 712, 1636, 1224]. Salmo [221, 1623, 704, 523, 80, 1637, 2031, 72, 1895, 273, 479, 1163, 1561, 1686, 164, 1063]. salmoides [56, 2152, 1486, 1528, 1721]. Salmon [1031, 1847, 1593, 407, 1667, 1734, 938, 589, 644, 1498, 1047, 1356, 95, 50, 1714, 1457, 2160, 1587, 1426, 411, 345, 332, 749, 2036, 992, 1910, 1865, 1032, 1761, 63, 2, 68, 291, 1301, 357, 207, 1607, 219, 2156, 1568, 857, 211, 1696, 1425, 1007, 1448, 2112, 891, 1885, 1977, 482, 1661, 1184, 1887, 2154, 1240, 791, 1036, 1156, 21, 170, 472, 434, 1041, 878, 621, 839, 1456, 1523, 1540, 120, 738, 2082, 1639, 1536, 882, 2102, 2142, 168, 167, 183, 1592, 1874, 814, 1039, 368, 886, 944, 208, 815, 1951, 2099, 834, 1560, 363]. salmon [226, 501, 1437, 237, 471, 335, 690, 97, 2070, 247, 307, 593, 442, 1499, 1049, 1555, 1749, 1857, 1792, 2009, 1272, 1205, 763, 1342, 1841, 330, 867, 733, 1042, 1281, 1044, 1473, 542, 2048, 1061, 872, 1908, 649, 1247, 658, 1632, 564, 1182, 1904, 1396, 1033, 858, 865, 1524, 400, 1922, 423, 1035, 1873, 1446, 1028, 642, 1295, 1438, 889, 706, 1478, 518, 1038, 1766, 625, 1285, 2002, 1525, 1905, 1614, 1542, 555, 973, 127, 327, 1304, 1420, 1324, 1969, 845, 1452, 1677, 540, 755, 1974, 1060, 1469, 782, 1043, 1045, 1037, 862, 1901, 1226, 1695, 1879, 1040]. salmon [943, 1672, 1638, 723, 556, 19, 760, 483, 484, 425, 369, 1753, 769, 1960, 395, 737, 1615, 712, 2060, 1790, 96, 1991, 77, 1208, 1662, 989, 1224, 868, 1480, 919, 177, 1145, 1444, 99, 538, 192, 1903, 391, 292, 199, 506, 2076, 1171, 287, 1673, 666, 914, 704, 523, 406, 651, 1050, 1735, 80, 209, 185, 2151, 2031, 1872, 687, 72, 1895, 273, 928, 1245, 1985, 2161, 479, 584, 552, 1625, 1163, 603, 2032, 408, 1355, 164, 41, 1062, 1260, 201, 1120, 1914, 1410, 176, 1063]. salmon-bearing [858]. salmon-derived [1540]. Salmonid [1894, 715, 1016, 1527, 1144, 800, 3, 1266, 362, 2136, 114, 661, 959, 1491, 1310, 1303, 2050, 1633, 1982, 439, 1034]. salmonids [205, 680, 1654, 298, 639, 841, 1095, 435, 1143, 1521, 2091, 1708, 1627, 283, 1171, 94, 1553, 939, 779, 1062, 1065]. salmonis [1356, 1297, 845, 939, 1561, 1686]. Salt [1514, 1637]. saltatrix [1488]. saltwater [769]. saltwater-entry [769]. Salvelinus [1075, 2025, 1165, 148, 1299, 1339, 153, 1795, 1404, 347, 1635, 774, 1594, 281, 174, 787, 459, 443, 494, 1916, 1574, 1257, 1989, 1624, 160, 837, 730, 1386, 1776, 1126, 1290, 724, 1494, 132, 1064, 1786, 1099, 2050, 124, 646, 826, 1780, 2140, 144, 712, 1325, 1274, 384, 1623, 149, 295, 1416, 1551, 2065, 72, 73, 722]. same [1634]. sample [728, 398, 1389]. sampled [1448]. sampler [318, 1369]. samplers [696]. samples [1223, 560]. Sampling [586, 109, 297, 195, 677, 1003, 420, 794, 1885, 787, 1022, 1994, 2005, 158, 1367, 1747, 98, 1198, 2086, 115, 563]. San [1700, 1630, 1959]. sanctuaries [133]. sand [1033, 1446, 662, 790, 1201]. sandbar [1581]. Sandeel [1860, 2047, 981, 1762]. Sander [1502, 847, 1508, 1849, 1402, 1701, 1783, 1248, 1603, 842, 1628, 1748, 1980]. Sands [1866, 374]. sandy [355]. Santa [931]. sapidissima [1952]. sapidus [1431, 1287, 1067, 766, 366, 1392]. sardine [2043, 430, 1216, 2013, 1854, 2054, 1634]. Sardinops [2043, 1216, 2013, 1854]. Sargasso [904]. satellite [1106, 2092, 1506, 1008, 553]. satiation [56]. Sault [1289]. saumon [1031]. saxatilis [1816, 572, 48, 414, 92]. scabra [879]. Scale [1392, 447, 1016, 670, 1404, 874, 1328, 766, 120, 167, 1880, 1794, 47, 1342, 964, 1989, 1632, 1033, 103, 1760, 788, 138, 1786, 2040, 1767, 2060, 683, 1263, 866, 1903, 1506, 1083, 1331, 2062, 1103, 1952, 792, 2032, 1527]. Scale-dependent [1392, 138]. Scaled [486]. scalefish [1898]. scales [1178, 1275, 1665, 1661, 1298, 449, 1663, 290, 1738, 1141, 895, 909, 213, 720, 448, 1751]. scaling [231]. scallop [476, 2033, 1697, 1462]. scallops [279, 954, 1215]. scattering [756]. scenarios [1457]. Schaefer [429]. Schaller [176]. schedules [1912]. schemes [882]. Schistocephalus [1640]. schlegelii [1278]. school [22]. schooling [1170]. Sciaenops [15, 560]. Science [449, 1799, 1273, 680, 1124, 2029]. sciences [1544, 1799, 263]. scientifiques [926]. Scomber [1712, 559]. scombrus [1712, 559]. scope [1594]. scope-for-activity [1594]. Scophthalmus [634, 808]. Scorpaenidae [1842, 533]. Scotia [1659, 1239, 403, 1451, 2101]. Scotian [1007, 797, 716, 581, 1191, 1190, 175, 320, 252, 326, 178, 249, 240, 1288]. Scottish [1951]. scour [114, 96]. sculpin [985, 1453]. sculpins [1973]. Sea [1266, 1241, 845, 2141, 1690, 1826, 57, 953, 735, 670, 982, 1328, 1409, 2088, 1459, 1939, 1297, 708, 1537, 1694, 1830, 1940, 476, 2033, 1284, 1658, 1916, 1265, 1671, 608, 1591, 1096, 1776, 218, 100, 279, 419, 1983, 689, 962, 954, 1870, 1637, 939, 879, 1750, 478, 1361, 1589, 1146, 1625, 1561, 1686, 1670, 1215, 236, 1256, 1620, 903, 1510, 1234, 1860, 1762, 1962, 1296, 1081, 1936, 699, 1221, 120, 996, 1717, 1385, 1957, 1689, 1311, 864, 1764, 1222, 1671, 1543, 2024, 1618, 45, 904, 510, 1846, 2068, 242, 1254, 1043, 1879, 1931, 606, 1134]. Sea [1884, 1770, 1843, 1224, 220, 1827, 525, 576, 131]. sea-run [1916, 1561, 1686]. seabed [1454, 2008, 1381, 600, 804, 803]. seabird [1762, 1483, 1662, 1870]. seagrass [1652]. seal [1233, 1492, 1469, 2042, 26]. seals [976]. search [35, 1231]. Searching [1996]. Seas [2031, 1557, 1104, 464]. season [1436, 291, 1371, 1246, 1272, 1205, 853, 1060, 1389, 651]. season- [291]. Seasonal [711, 1401, 837, 1873, 91, 2092, 1790, 1005, 1678, 652, 1249, 903, 2023, 1566, 1802, 772, 228, 1314, 1315, 1394, 1038, 773, 1571, 320, 1116, 1901, 387, 962, 651, 1017]. Seasonality [105]. Seasonally [8, 1920, 1572, 1573, 1129]. seaward [1281]. Seawater [506, 1075, 207, 878, 1561]. SeaWiFS [905]. Sebastes [1278, 1052, 626, 1566, 863, 1252, 1223, 1606, 2104, 1842, 1825, 871, 2018, 986, 1251, 1154, 1838, 1258, 1724]. Sebastolobus [173]. Secchi [824]. secondary [2080, 1585, 1861]. secretion [172]. sections [183, 1964]. sector [125]. sediment [1157, 413, 346, 719, 1401, 1214, 1536, 1565, 469, 154, 1209, 470, 272, 881, 964, 371, 994, 111, 1303, 303, 130, 334, 461, 717, 136]. sediment-associated [111]. sediment-bound [136]. Sedimentary [793, 1155, 2070]. sedimentation [339]. sediments [1173, 1474, 1834, 1611, 1779, 1277, 108, 767, 242, 417, 105, 1778, 1467, 549, 381, 1269]. see [1029]. seeded [1215]. segment [305]. segregation [1560, 841, 494, 1907]. select [2099]. selected [326]. selecting [1955, 866]. Selection [1092, 148, 1283, 729, 183, 181, 1933, 987, 1473, 2048, 885, 1576, 612, 1131, 159, 356, 1208, 1843, 914, 1934, 23, 675]. selective [368, 875, 532, 531, 370]. selectivity [174, 1755, 2144, 487, 1322, 847, 2143, 51, 1077, 2162]. selenium [1320]. self [1500, 1899, 1921, 1261]. self-organizing [1500, 1261]. self-recruitment [1921]. self-thinning [1899]. Semi [1441]. Semi-local [1441]. seminal [366]. semiochemicals [1356]. semiparametric [367]. sensing [1598]. sensitive [1384]. Sensitivity [462, 1621, 1978, 1813, 1482, 1002, 1151, 1398]. sensory [1695]. separate [407]. separated [1666]. Separating [2128, 1528, 2143]. separation [1897, 1258]. septemspinosa [790, 1201]. sequence [469, 470, 272, 726]. sequential [308, 1997, 1814, 1467]. series [2052, 871, 124, 1735, 193, 1648, 736, 1228]. Serranidae [146]. sessile [954, 734]. Seston [279, 812]. set [99]. sets [1849]. setting [1584, 452, 1772]. settlement [1112, 15, 605, 131]. seven [1806]. severely [1553]. sewage [575, 809]. Sex [861, 146, 1928, 1390, 846, 16, 676, 678, 689]. sex-based [1928]. Sex-specific [861]. Sexual [810, 378, 1572]. shad [1725, 813, 113, 1952]. shaded [1237, 1358]. shading [915]. Shaking [1565]. shallow [202, 725, 1711, 1504, 1305, 1569, 946, 703, 1242, 130, 135, 675]. Shallows [240]. shanny [504]. shape [846, 1258]. shark [1785, 1397, 1682, 1898, 22]. sharks [539, 1377, 1581, 2144, 557]. sharpen [1925]. shear [1193]. shedding [1489]. Shelf [1191, 252, 1554, 240, 1407, 1557, 877, 1177, 1383, 1842, 187, 2117, 525, 1069, 716, 581, 1190, 271, 175, 320, 797, 326, 178, 249, 1288]. shell [1773, 2033]. shellfish [350]. shells [1918]. sheltering [72]. shelves [1675]. shield [1647, 1799, 1300, 1559, 1579, 806, 1811, 1382, 1246, 1474, 774, 306, 1958, 1504, 1089, 1801, 1810, 742, 1806, 1276, 543, 1800]. Shift [887, 682, 750, 2054, 1069]. Shifting [519, 8]. Shifts [2081, 2155, 2039, 1848, 588, 1470, 1996, 15, 1649, 66, 1070, 2057]. shiner [286, 918]. ship [1611, 1779, 613, 1778]. ships [1157, 1277]. Shiraz [1458]. Shoal [1341]. shoaling [2015, 1879]. shoals [286]. shock [1978, 1495, 642]. shoreline [405, 933]. Short [1471, 1711, 866, 152, 496]. Short-term [1471, 1711, 866, 152, 496]. shortfin [747]. shortnose [1676]. show [1537]. shows [1813]. shrimp [1979, 956, 1177, 203, 465, 300, 790, 1201]. shunt [948]. sibericum [1972]. sibling [862]. side [1028, 1310]. side-aspect [1028]. sight [1385]. signal [1292]. signalling [622]. signature [1924, 963]. signatures [942, 1312, 581, 1577, 2014, 229, 1286, 720]. signifer [426]. Significance [155, 11, 2124, 432]. silicic [117]. silt [1033, 1446]. silver [530, 616, 175]. silverside [1124]. Simcoe [278]. Simeon [1465, 1464]. similar [1468, 1072]. similarity [420, 1994]. similarity-based [420]. Simple [927, 2079, 705]. Simplification [1652]. Simplified [493]. simulate [829]. simulated [1139, 548, 1332, 1333]. Simulating [1398, 657, 206, 1937]. Simulation [323, 1028, 628, 2124, 1619, 843, 745]. Simulations [856, 582, 151, 2154, 792]. Simultaneous [1024, 1561]. simultaneously [1689, 1077, 2162]. since [1455, 1659, 486]. single [781, 1539, 98, 1677, 1955, 577, 80, 563]. single- [563]. single-generation [1677]. single-pass [98]. single-species [1539, 1955]. sink [2064]. siphonulae [756]. sires [1240]. Site [985, 2101, 903, 1312, 437, 1174, 1900, 1542, 2016, 1139, 704]. site-specific [1174, 2016, 1139]. sites [1175, 1828]. siting [1469]. situ [112, 2100, 455, 1723, 160, 1219, 1099, 1548, 44, 313, 353, 125, 115, 14]. six [1809, 902]. Size [1928, 873, 1383, 45, 146, 1475, 644, 291, 857, 53, 725, 664, 805, 1388, 1610, 174, 787, 1022, 699, 35, 1152, 1616, 1655, 487, 729, 1321, 1264, 1552, 772, 1520, 368, 102, 1385, 978, 237, 810, 528, 442, 1508, 841, 526, 315, 1266, 2038, 647, 362, 624, 398, 1545, 1309, 1873, 1942, 238, 1038, 1084, 1571, 1285, 215, 1304, 1974, 1596, 1132, 1674, 1815, 159, 547, 1884, 1743, 1627, 283, 1583, 640, 2037, 869, 958, 532, 531, 790, 1613, 1017, 2065, 563, 1343, 575]. Size- [1928, 1152, 1616]. size-at-age [1152, 1616, 1596, 640]. size-based [664, 805, 1388, 841, 1571, 958]. Size-dependent [45, 1321, 2038, 1873, 790, 1613]. size-regulated [1475]. size-related [215]. size-selection [729]. size-selective [368, 532, 531]. size-specific [291, 1038]. Size-structured [1383]. sizes [1498, 1427, 651, 1017]. Skagerrak [510]. skate [43, 1411]. skeletal [976]. skin [71]. skipped [1335, 2105]. skipper [1364]. Skistodiaptomus [1439]. slices [1669]. slimy [985, 1453]. slope [47]. slopes [173]. Slow [1654, 1010, 1505, 955]. slow-growing [1505]. slow-paced [955]. slower [1128]. Small [103, 788, 1246, 739, 1129, 2134, 1322, 59, 1115, 1532, 964, 915, 435, 1801, 338, 1758, 1452, 662, 709, 683, 135, 866, 1380, 405, 2062, 294, 1888]. Small-scale [103, 788, 964, 866, 2062]. smallmouth [2150, 1665, 1211, 1135, 1989, 289, 1771, 1294, 1923]. smelt [1562, 163, 1701, 444, 1070, 1683]. smelts [998]. Smirnov [721]. smolt [2112, 1523, 2009, 1671, 1698, 865, 661, 862, 391, 199, 627, 164]. smolting [1587, 1951, 400, 199]. smolts [207, 878, 1461, 1295, 706, 2002, 71, 506, 439, 2032, 1561, 1063]. smooth [729]. snail [2083, 413, 117]. Snake [1143, 177, 176, 307, 565, 789, 483, 484, 369, 2060]. snakehead [1718]. snapper [1759, 1934]. Snohomish [556]. snorkel [185]. snow [481, 1296, 1081, 1340, 1767, 678, 768]. snowmelt [29]. snowmelt-driven [29]. Snowy [946]. soak [288]. Social [64, 508, 888]. sockeye [688, 211, 1448, 1887, 839, 1456, 1536, 944, 226, 2070, 1049, 1555, 1857, 1272, 1205, 867, 2048, 872, 1922, 1478, 1766, 625, 973, 1060, 943, 1672, 760, 737, 1991, 1444, 99, 1673, 666, 914, 1872, 1245, 2032, 201, 1410]. sodium [1093]. soil [1360]. solar [601, 1792, 542, 38, 445, 32]. somatic [644, 82, 1125, 1573]. some [1192, 1650, 1881, 1983]. Sommaire [255, 492, 669, 1056, 1319, 1564, 1728, 2159]. sorption [54]. Sound [205, 509, 649, 2076, 536, 1695, 934, 1910, 357, 450, 159, 989]. sounders [1149]. sounds [548]. source [2064, 1064]. sources [756, 673, 1183, 1908, 1664, 1366, 723, 334, 1326, 1938, 428, 1845]. South [353, 734, 1986, 2114, 1809, 1675, 617, 700, 1987, 786, 982, 1785, 1284]. south-central [1986, 2114, 1809, 617, 700, 1987, 786]. southeast [1675, 871, 930]. southeastern [1533, 614, 546, 1693, 545, 1011, 209]. Southern [1971, 1799, 581, 1547, 633, 1233, 1756, 2100, 488, 487, 1196, 1842, 1220, 1044, 1482, 1386, 1760, 1921, 280, 540, 755, 583, 860, 1898, 1926, 227, 513, 1398, 22, 1800, 1312]. Southwest [1001]. southwestern [1540, 709, 1737]. Soviet [194]. sp [452, 662]. space [874, 1308, 1692, 1345, 6, 1660, 911, 920, 1648, 1228]. spaces [667]. spacing [1140]. span [2109, 2069]. Spanish [4]. sparid [1112]. Sparus [1658, 1096]. spat [1971]. Spatial [942, 1559, 1976, 1177, 741, 1296, 1081, 766, 1252, 314, 1939, 171, 1340, 363, 1682, 1608, 1311, 1813, 1481, 245, 1172, 1038, 100, 2071, 1121, 395, 1782, 96, 1832, 220, 1050, 768, 1013, 1912, 1407, 195, 1317, 1811, 333, 1382, 1178, 558, 1665, 1661, 167, 582, 561, 1223, 388, 840, 449, 141, 442, 76, 1342, 270, 964, 1538, 396, 103, 890, 2123, 1899, 1760, 1758, 1893, 788, 138, 1486, 206, 1763, 1526, 710, 1439, 1551, 819, 2075, 448, 1751, 246]. spatially [1066, 1475, 2023, 151, 359, 1342, 1747, 980, 1439]. spatio [1711]. spatio-temporal [1711]. Spatiotemporal [2142, 1261]. spawner [67, 166, 1996, 990, 328, 307, 1846, 1763, 1583, 317]. spawner-recruit [307]. spawner-recruitment [67]. spawners [1971, 237, 765, 733, 1746]. Spawning [1371, 2113, 1163, 385, 148, 1426, 411, 345, 474, 1690, 179, 285, 512, 1016, 1527, 1287, 783, 1629, 1448, 1341, 517, 1523, 450, 168, 167, 511, 1092, 1830, 834, 2034, 867, 828, 1335, 1618, 872, 396, 82, 1396, 103, 1900, 727, 959, 1542, 2091, 2071, 1303, 1775, 1609, 1828, 369, 473, 1528, 377, 2105, 1603, 1735, 1017, 478, 720, 216, 2149]. spawning-habitat [959]. spawning-potential [1287]. special [1620, 1800]. specialization [1590]. speciation [2089, 684]. Species [1405, 598, 1890, 1258, 535, 527, 474, 2107, 1705, 607, 581, 1795, 264, 1848, 1275, 1025, 416, 874, 1153, 1611, 1779, 1204, 62, 529, 1630, 233, 1994, 1411, 2097, 1098, 996, 2010, 375, 840, 1718, 1381, 1851, 902, 2143, 785, 1663, 2116, 451, 1044, 1482, 1716, 397, 1539, 1529, 1250, 1643, 1495, 2053, 949, 1384, 2081, 119, 1064, 1693, 1786, 2133, 1738, 565, 587, 2130, 1072, 1955, 1898, 606, 1778, 866, 958, 1198, 525, 1503, 1861, 610, 131, 2075, 106, 69]. species-rich [1204]. specific [146, 291, 1174, 21, 1755, 465, 399, 1115, 639, 1049, 1852, 2048, 1038, 861, 373, 2016, 1139, 1195, 26, 1450, 627, 523, 227, 513]. specification [354, 230]. specimens [976]. speckled [1447]. spectra [725, 1610, 1552, 869]. spectrometry [1273, 125]. spectroscopy [1155, 1467]. speed [2004, 89, 12, 1219, 914]. speeds [247]. Sperm [1502, 1784, 366, 508, 678, 23]. spicatum [1408, 1972]. spill [509]. spilonotus [1468]. spine [1740]. spines [1372, 1740]. spinifera [282]. spiny [1631, 1999, 475]. spinycheek [1953]. split [1271, 1369]. split-beam [1271]. sport [424]. spot [835, 213]. spots [1504]. spotted [1490, 132]. spp [938, 1455, 1052, 1301, 807, 1566, 1020, 2154, 1252, 2102, 1499, 1749, 1842, 1095, 649, 778, 959, 936, 1304, 1310, 1452, 540, 755, 662, 1393, 19, 1960, 1615, 1429, 1145, 687, 1215]. sprat [1510, 396, 1846]. Sprattus [1510, 1846]. Spread [821, 1788, 1861]. Spring [191, 1455, 63, 1268, 1246, 2106, 482, 517, 1639, 1592, 190, 1558, 901, 565, 789, 1638, 483, 319, 2060]. spring-summer [63]. squalid [1372]. Squalus [1533, 1999]. squares [635]. squid [248, 671, 942, 336, 1383, 1944, 1203, 925]. Sr [813, 1047, 1273, 2122, 1577, 585, 813]. Sr/ [813]. Sr/Ca [1047]. SSWC [617]. St [831, 581, 1566, 633, 1233, 1340, 186, 187, 324, 1878, 2049, 1026, 516, 1926, 734]. St. [798, 1694, 887, 1831, 1229, 898, 444, 1159, 319, 136, 541]. Stability [347, 720, 2160, 245, 1693, 1985, 2161]. Stable [450, 338, 1068, 373, 1389, 1472, 673, 1970, 546, 1267, 1160, 985, 141, 505, 833, 1608, 438, 1061, 1908, 1105, 997, 1347, 452, 1897, 1754, 1507, 1968, 1512, 1373, 1961, 1293, 1763, 1790, 1346, 2027, 1422, 1927, 1015, 1120]. stable-isotope [997]. stage [1292, 465, 521, 1612, 727, 369, 356, 1195, 1819, 1488]. stage-classified [1612]. stage-explicit [521]. stage-specific [465, 1195]. stage-structured [1488]. stages [1157, 1510, 733, 160, 1433]. stagnalis [413, 117]. stained [944]. stamina [1575]. standard [1002]. standardization [331, 2072]. standards [1993]. standing [1687]. stands [1406]. star [1833, 1215]. starvation [802]. State [1587, 1334, 1660, 1228, 2153, 877, 401, 1305, 1345, 6, 1987, 1648, 617]. State-dependent [1587, 1334]. State-space [1228, 6]. States [1675, 958]. statewide [1959]. stations [1578]. statistic [721]. statistical [1939, 1736, 465, 181, 1272, 1205, 1222, 1964, 1825, 600, 804, 2035, 803, 1097, 1771, 1735, 1653, 1517]. statistics [138]. Statolith [1830]. statoliths [942]. Status [1031, 1032, 153, 2064, 1155, 170, 434, 1169, 34, 963, 1090, 1960, 888, 18]. statut [1031]. Ste. [1687]. Steady [2153, 1987, 617]. Steady-State [617]. stealthy [1623]. steelhead [64, 1797, 1920, 291, 1757, 783, 1461, 1950, 2118, 1698, 604, 1664, 727, 1753, 1185, 850, 27, 28, 29, 921, 1796, 185, 286, 273, 30]. Steller [1328, 1537, 962, 1589]. stenolepis [2092]. step [1457, 1329, 2094]. steroids [16]. stick [67]. stickleback [726, 1343]. sticklebacks [1249, 1704, 995, 1640]. Still [67]. stimulated [87, 10]. Stizostedion [810, 828, 45, 356, 762, 1021]. Stochastic [1213, 593, 1226, 892, 1712, 1336]. Stock [46, 181, 1666, 248, 1858, 447, 285, 512, 1191, 241, 846, 1696, 189, 230, 454, 367, 520, 638, 1067, 1962, 52, 403, 2142, 511, 561, 1892, 979, 1196, 2098, 844, 843, 929, 1543, 1247, 396, 1539, 1922, 2018, 1846, 1080, 244, 1403, 1921, 331, 429, 925, 924, 1345, 1491, 1814, 2003, 1218, 507, 1775, 1898, 1687, 1884, 970, 538, 224, 635, 854, 1050, 240, 209, 1819, 927, 1336, 1724, 517]. stock-recruit [244]. Stock-recruitment [46, 189, 454, 1050]. stocked [524, 690]. Stocking [1757, 1593, 1012, 427, 1266, 238, 625]. stocks [671, 1424, 63, 1107, 1190, 457, 1887, 120, 738, 450, 844, 885, 349, 2068, 1869, 1677, 540, 755, 1649, 1145, 553]. Stokes [1419]. stomach [236, 1051]. stomach-content [1051]. storage [960, 816, 1380]. stored [1002]. storm [654]. storms [1294]. straddling [553]. strains [160, 1238, 506]. Strait [2026, 2106, 271, 1220]. strategies [2155, 1457, 1797, 1885, 1867, 1522, 298, 1505, 1196, 1642, 1628, 1138, 2149]. strategy [195, 37, 620, 2000, 235, 1637]. stratification [1246, 1532, 1569, 87, 10, 61]. stratified [1781, 2012]. Stratigraphy [108]. stratospheric [61]. straying [382, 209]. Stream [1165, 1964, 1679, 674, 70, 1839, 2083, 2023, 1292, 1973, 614, 1756, 151, 558, 1665, 1370, 1635, 800, 1362, 1994, 1057, 718, 2010, 183, 1592, 1565, 1330, 1237, 1358, 840, 1951, 1115, 781, 960, 1225, 47, 359, 915, 1085, 1095, 2066, 764, 987, 624, 190, 400, 1279, 1172, 1692, 1899, 152, 1438, 25, 300, 2002, 1384, 972, 232, 1614, 1847, 709, 714, 1045, 96, 1074, 8, 139, 1263, 177, 1195, 461, 404, 1010, 1956, 297, 1323, 478, 2006, 732, 1751, 629, 176]. stream- [400]. stream-dwelling [1279]. stream-living [1172]. stream-resident [1692, 1899]. stream-type [177, 176]. streamflow [2127]. Streams [1774, 910, 140, 596, 153, 1547, 180, 614, 891, 874, 643, 579, 1421, 800, 1802, 1726, 1809, 1500, 880, 1940, 2093, 915, 867, 435, 1964, 1329, 362, 1538, 654, 858, 778, 1535, 65, 466, 1758, 215, 753, 1452, 468, 723, 1703, 545, 1325, 1185, 283, 696, 391, 522, 94, 313, 31, 1103, 722, 779, 408, 106, 573, 1270]. streamside [715]. strength [248, 385, 455, 226, 1723, 1028, 605, 1854]. strengths [1027]. Stress [769, 1108, 1591, 1691, 642, 730, 945, 928, 1120, 1688]. stressed [1493]. stresses [259]. stressor [2131]. stressor-independent [2131]. stressors [2074, 1806, 1812]. strike [1867]. Striking [2090]. striped [1816, 2097, 572, 48, 414, 1844, 1815, 92]. strongly [1038]. Strongylocentrotus [1591, 100]. strontium [1275, 229, 1061, 835, 213, 1065]. strontium-to-calcium [1065]. structural [873, 924]. Structure [1190, 146, 95, 383, 1820, 1559, 1174, 626, 241, 1984, 1929, 855, 664, 805, 1935, 1619, 174, 1958, 1520, 1892, 1223, 840, 1880, 1479, 141, 1437, 1432, 833, 1608, 154, 2104, 844, 843, 1841, 1685, 1327, 1644, 398, 402, 289, 1844, 426, 1290, 972, 986, 1251, 1244, 278, 1901, 49, 1550, 646, 2120, 2084, 1767, 899, 1387, 70, 1224, 747, 550, 717, 1891, 726, 2001, 1013, 1406, 1109, 201, 629, 1110, 216]. structured [1755, 1383, 1868, 6, 533, 507, 980, 919, 1439, 1488]. structures [1046, 2147, 1441]. structuring [1744, 961]. Stuart [872]. studied [1552]. studies [1, 586, 15, 1897, 1968, 892, 627, 313, 1487]. study [1256, 1158, 195, 447, 1948, 907, 1533, 739, 628, 2124, 874, 546, 1111, 1619, 1577, 348, 1243, 250, 1940, 1612, 494, 2116, 843, 239, 1366, 2053, 1516, 1921, 417, 994, 1803, 1699, 1512, 460, 1447, 1917, 869]. Studying [1261, 1238, 496, 1932]. sturgeon [1436, 1787, 2098, 1212, 1332, 1333, 2038, 1676, 360, 361]. subalpine [128, 1546]. subarctic [2015, 1038, 1043]. subbifurcata [504]. subclasses [2018]. Subdivision [159, 1286]. subjected [160]. sublethal [878, 284, 868]. submarine [321]. submerged [2100, 1496, 136]. submersed [964]. submersible [1724]. subnanogram [109]. suboptimal [191]. subordinate [286]. subpolar [2034]. subsequent [207, 1353, 1673]. subsequently [878]. substances [1474]. substrate [1474, 741, 233, 915, 1033, 1897]. subsurface [1549, 1864]. Subtropical [282, 725]. suburban [183]. subyearling [2156]. success [332, 2036, 1299, 1757, 1502, 1020, 21, 1180, 1536, 765, 1852, 1446, 727, 1231, 606, 1583, 1631, 850, 27, 1294, 1684, 1163, 1270]. successes [266]. Successful [434, 2147]. successfully [2020]. successive [1743]. sucker [1300, 440]. Sudbury [112, 108, 110, 107]. Sudden [157]. sufficiency [420]. suggestions [1650]. Suitability [167, 518, 2016, 1208]. suitable [635]. suite [991]. sulfate [1742]. sulfide [303]. Sulfur [381, 410, 584, 552]. Sulphur [906]. Summer [1106, 435, 1974, 2036, 63, 482, 2063, 1115, 371, 1950, 188, 2113, 565, 789, 1570, 483, 713, 1376, 2076, 1700, 1015]. summer-run [1950]. summer-spawning [2113]. summering [2092]. sunfish [1020, 290]. Superior [1174, 830, 1101, 1757, 141, 1518, 1519, 1265, 1266, 1856, 2084, 1289, 2148, 1450, 126]. supplementation [1523, 1354, 2076]. supplemented [2076]. supplementing [1226]. supply [1357, 279]. support [28]. supported [912, 1907]. supportive [214]. suprabenthic [2157]. suprabenthos [648]. Surface [1372, 1601, 617, 1284, 1532, 1870, 1149, 1103]. surfactant [113]. surficial [242]. surge [1444]. surmuletus [2097]. Surplus [430, 1945, 310, 1216, 2013]. surrogate [1248]. surrounding [51]. Survey [890, 1075, 236, 310, 497, 988, 1749, 604, 1676, 1893, 1747, 533, 1702, 224, 1819, 1914, 746]. surveys [2141, 420, 1833, 399, 935, 1342, 212, 2055, 280, 2042, 745, 1462, 1836, 216, 2145]. Survival [2137, 1459, 989, 1114, 2023, 1313, 1920, 615, 596, 807, 1006, 776, 89, 12, 524, 738, 1951, 1115, 672, 442, 2009, 1508, 572, 1842, 1095, 1671, 239, 188, 1725, 885, 160, 1182, 1904, 1033, 865, 1035, 1873, 2002, 1677, 540, 755, 1649, 468, 1515, 1040, 789, 369, 1815, 395, 177, 26, 461, 439, 627, 92, 28, 287, 687, 261, 2032, 1488, 176, 1110]. suspended [1536, 351, 279, 1423, 1697]. suspension [1598]. suspension-feeders [1598]. sustainability [855, 1140]. sustainable [1379, 1642, 848, 1925, 421]. sustained [992]. Suthers [156]. Sweden [410, 228, 1209, 190, 1011]. Swedish [468]. swim [936]. Swimming [974, 714, 1436, 1249, 1734, 2004, 570, 975, 89, 12, 791, 1156, 247, 360, 1219, 618, 2061, 1575, 149, 1245, 122]. switching [1970, 1931, 519]. sympatric [1371, 1822, 21, 2118, 2018, 727, 444, 1907, 216]. Symposium [893, 1123]. synchronicity [474]. Synchronous [2011, 1808]. synchrony [120, 763, 1760, 1043, 1870]. syndrome [817]. syndromes [1491]. Synthesis [509, 181, 107, 908]. synthetic [705]. System [1549, 1252, 145, 899, 683, 1111, 344, 1222, 485, 972, 1687, 1238, 516, 110, 777, 1688]. system-wide [485]. Systematic [1403, 1024, 1023]. systematics [1967, 173]. systems [2080, 1644, 1529, 338, 1373, 762, 334, 934, 390, 1932, 1485].

tactics [1498, 1200, 658, 1990, 7]. tag [2077, 1489, 2078, 859, 488, 1997, 590, 597, 816, 92, 603]. tag-recovery [1489, 92]. tagged [1671, 1143, 1942, 1060, 380, 704]. Tagging [382, 681, 488, 2008, 1875, 414, 2092, 1550, 1365, 1837, 1008, 209]. Tagliamento [477]. tags [1714, 1112, 1506, 2030]. Tahoe [822]. Tailoring [1048]. taimen [2116, 2116]. take [1118, 1943, 380]. tale [949]. tamarense [541]. tap [71]. Tapes [773]. target [1584, 1533, 455, 1723, 1555, 1028, 866]. targeting [1855, 856, 1736]. targets [1107]. Taste [1882]. taurus [1785]. taxa [86, 1956]. taxon [781]. taxonomic [1349, 781, 1384]. taxonomy [961]. technique [1948, 1830, 44]. techniques [1259, 465, 1919, 297, 87, 10, 131]. technology [135]. telemetry [2077, 1106, 557, 666, 2032]. teleost [74]. temperate [440, 161, 579, 1421, 2063, 1112, 1481, 1367, 1090, 713, 139, 1808, 1117, 1742, 1030]. Temperature [1645, 1431, 835, 854, 1436, 1510, 411, 345, 1826, 196, 53, 1756, 1829, 665, 808, 1152, 1616, 1459, 1586, 2095, 1284, 1532, 572, 922, 1585, 371, 239, 197, 1495, 1314, 1315, 730, 674, 327, 540, 755, 71, 816, 425, 2140, 1770, 1387, 532, 405, 1634, 149, 790, 1870, 1637, 1534, 1063]. temperature- [790]. Temperature-dependent [1431, 1152, 1616]. temperatures [1949, 314, 718, 915]. template [1225, 1633]. Temporal [1811, 969, 561, 81, 763, 1043, 2117, 1280, 1746, 1551, 942, 2160, 1317, 1976, 314, 120, 1939, 582, 1711, 363, 1682, 1608, 442, 1311, 1342, 396, 103, 245, 1172, 788, 1649, 2071, 1693, 1121, 206, 1528, 1832, 2062, 1013, 1985, 2161, 1912, 1451]. temporally [1763]. temporary [1460, 632]. ten [842]. tenets [1447]. tens [851]. Tenualosa [813]. tenuis [940]. term [1158, 1066, 1531, 1390, 1006, 1848, 1471, 2114, 878, 1711, 1237, 1358, 465, 2128, 2087, 1291, 1049, 822, 1044, 1035, 152, 2068, 1803, 1477, 1775, 2127, 1674, 1806, 707, 1812, 852, 866, 2148, 404, 496, 1915, 591, 1623, 1534, 1030, 1807]. terminating [1063]. terms [855]. Terrestrial [1961, 715, 1499]. territoriality [1796]. territory [526]. Test [839, 1066, 383, 1979, 1929, 233, 378, 1096, 1102, 135, 2131, 1361, 1021]. tested [1618]. Testing [372, 274, 1773, 823, 417, 1580, 336, 624, 534, 554, 1398]. testosterone [853]. tests [316, 359, 667, 768]. Texas [560]. TFM [1670]. Thalassiosira [62]. Thaleichthys [962]. thallium [812]. thalweg [1452]. Thau [352]. thaw [2110]. their [756, 1690, 793, 207, 952, 891, 1949, 2147, 1936, 228, 1658, 987, 1685, 322, 1002, 1491, 813, 1542, 2071, 252, 1139, 1575, 1781, 2012, 1472]. them [1964]. theoretical [123]. theories [1126]. theory [1234, 1403, 261]. Theragra [2155, 1283, 1629, 1311, 1604, 2055, 2000]. there [1690, 1504, 1393, 2050]. thereof [2089]. Thermal [477, 800, 1604, 379, 1120, 1756, 1978, 1658, 1591, 1438, 1307, 87, 10, 61]. thermoregulation [2118]. thermotolerant [1495]. thing [1131]. thinking [2068]. thinning [1899]. Third [893]. Thirty [1799]. Thompson [2, 1410]. thornyhead [173]. those [759]. Threat [1428, 1859]. threatened [657, 735, 874, 565, 483, 484, 1145, 2027, 292]. threats [613, 1780]. Three [1589, 589, 1941, 211, 739, 2112, 1087, 1630, 1432, 823, 1849, 1257, 1618, 290, 160, 1919, 691, 1674, 1496, 1683, 2151, 131, 1628]. Three-dimensional [1589, 823, 2151]. threespine [1249, 1704, 995, 1640]. threshold [1498, 1996, 700]. thresholds [588, 2144]. Throwing [553]. Thunnus [1936, 1641, 1875, 1642, 1997, 2096, 299, 1862, 860, 1506, 1008]. Thymallus [1006, 1432, 987, 849, 2120, 1286]. thynnus [1641, 1642, 1997, 1862, 1506, 1008]. Tibbo [1465, 1464]. Tidal [324, 630, 1759, 1737, 325]. tidally [516]. tightly [1510]. tilapia [1853, 1485]. timber [1165]. Time [193, 179, 288, 3, 871, 1692, 25, 1231, 1913, 2113, 1738, 124, 1743, 683, 1925, 911, 692, 1735, 1103, 920, 1648, 1517, 736, 1228]. time-series [1228]. time-varying [692, 1517]. timely [955]. times [1659, 486, 667]. Timing [1344, 1246, 2106, 1523, 1044, 1698, 604, 1664, 1035, 1417, 1901, 319, 523, 406]. Timothy [1544]. tissue [711, 1970, 16, 1658, 452, 1099, 618, 1389, 928]. tissues [1441, 1347, 1117, 801, 883, 1859]. tissular [437]. tobianus [2047, 981]. tolerance [1482, 361, 1102, 506]. too [2050]. tool [202, 2080, 1458, 1589]. tools [257, 673, 455]. toothfish [1312]. top [1454, 1610, 1930, 1187]. top-down [1454, 1610, 1187]. Topeka [918]. topographic [1339]. total [716, 1156, 1954, 1919, 303, 1011, 1253, 1700, 110, 1807]. toxic [24, 370, 541]. toxicants [1079]. Toxicity [1173, 1599, 1243, 54, 1093, 2073, 480, 1433]. toxicokinetic [143, 1099]. toxicological [509]. toxin [1946, 1769]. toxins [1882]. trace [2052, 143, 1824, 125, 87, 10]. tracers [1373, 1015]. Trachurus [1712]. tracked [2005]. Tracking [1566, 1271, 2033, 1660, 135]. tract [253]. trade [1129, 1840, 1519, 2118, 565, 1780]. trade-offs [1129, 1519, 2118, 565, 1780]. traditions [1995]. training [1021]. Trait [1789, 1824]. trait-based [1824]. traits [219, 388, 501, 2087, 1671, 1309, 1890, 1139]. trajectories [1761, 2022]. trajectory [446]. transboundary [1395]. transect [329]. transects [1047, 1577, 813]. transfer [207, 128, 1902, 2041, 1916, 587, 1048, 1260]. Transferability [518, 849, 815]. transferable [1855]. transferred [878]. transformations [949]. transforming [1799, 1800]. Transgenerational [1422]. transgenes [471, 335]. transgenic [97]. transit [1103]. Transition [877, 831, 798, 1109]. transitions [2087, 1015]. transitive [677]. Translating [1457]. translocation [1333, 498, 1270]. transmontanus [1212, 2038]. Transoceanic [1875, 1157]. transparency [623, 717, 824, 204]. transplant [534, 209]. transplanted [112]. transponder [135]. Transport [179, 1856, 1510, 992, 1079, 333, 1562, 1860, 1298, 1076, 1565, 142, 325, 110, 870]. transported [832, 1269]. trap [2088, 2008, 280]. trapping [1351, 1487]. traps [487, 1367]. trawl [37, 1317, 1578, 1855, 203, 2005, 399, 1381, 1794, 497, 1689, 875, 2055, 245, 533, 745, 746]. trawler [1736]. trawlers [547]. trawling [1111, 664, 805, 1271, 1851, 355, 734, 1934]. trawls [2143]. treatment [1389, 14]. tree [1259, 1948, 234, 1737]. tree-ring [1259, 1948]. trees [1964, 1074]. trend [310]. Trends [27, 28, 29, 1578, 1390, 1328, 1887, 1802, 1630, 738, 772, 1951, 1049, 1266, 1545, 1803, 1304, 1649, 2117, 760, 369, 884, 1633, 292, 240, 576, 1653, 1451, 1807, 736]. Trente [1799]. Triad [346]. Tributaries [334, 1174, 1830, 712]. Tributary [1538, 1541, 523, 1757, 1850, 164]. Tributary-specific [523]. Tribute [1464, 1399]. tributyltin [1173]. Trichechus [2121]. trifluoromethyl [1670]. trigger [1565]. trihalomethane [636]. trip [37, 1937]. triploid [992]. Tripping [448]. tripton [458]. Trisopterus [1595, 2024]. Troisième [893]. troll [472]. TROLS [296]. Trophic [140, 833, 1968, 1947, 433, 1236, 592, 202, 1292, 1191, 660, 1902, 981, 643, 1155, 546, 1621, 1160, 1243, 228, 2041, 388, 1169, 1305, 389, 1481, 1519, 239, 34, 510, 1571, 625, 1507, 1924, 703, 963, 707, 1090, 1346, 1026, 18, 1915, 1930, 73, 1546]. trophic-level [1968]. trophically [679]. trophodynamic [1944]. tropical [1581, 1330, 102, 1427, 300, 1516, 813, 1699, 1090, 1678, 792]. trout [210, 2025, 1114, 1012, 1165, 148, 2023, 530, 966, 1299, 2137, 1339, 596, 1788, 291, 153, 1795, 570, 495, 1200, 1635, 774, 1822, 1594, 281, 1978, 2014, 1880, 459, 443, 214, 427, 1115, 363, 1461, 1722, 821, 823, 526, 2093, 47, 494, 1916, 1886, 1574, 1257, 1671, 284, 1093, 1989, 1194, 1698, 1353, 604, 1664, 160, 778, 508, 1004, 1279, 1172, 1692, 1899, 730, 1386, 1776, 1126, 1307, 215, 727, 1290, 1913, 698, 724, 1494, 2071, 1064, 1440, 917, 71, 468, 1786, 1609, 2016, 1099, 2050, 124, 945, 646, 826, 1780, 2140, 712, 2061, 1325]. trout [1274, 1217, 2135, 77, 311, 931, 1636, 221, 968, 1376, 850, 1575, 1982, 2027, 461, 888, 1623, 1433, 293, 149, 295, 1416, 1637, 286, 1551, 2065, 73, 1350, 1001, 722, 801, 1561, 1686, 1262, 1688, 1270, 1453, 2057]. trutta [1671, 210, 1114, 495, 1200, 214, 427, 363, 823, 494, 1886, 1671, 1004, 1279, 1172, 1692, 1899, 1307, 71, 468, 1786, 1609, 2016, 1636, 221, 1623, 1637, 1561, 1686]. tshawytscha [589, 1047, 95, 1714, 1457, 2160, 1587, 1426, 749, 992, 63, 291, 219, 2156, 2112, 1885, 482, 1184, 1639, 168, 167, 886, 2099, 834, 1560, 501, 1437, 763, 1342, 1473, 889, 1525, 1695, 565, 789, 1638, 723, 483, 484, 369, 769, 2060, 1662, 989, 919, 177, 192, 1903, 687, 928, 1985, 2161, 584, 552, 1355, 41, 1120, 176]. Tsuga [522]. Tubifex [417]. Tubificidae [417]. tumor [157]. tumor-like [157]. tuna [592, 660, 1833, 1936, 1641, 1642, 1997, 2096, 299, 1699, 1862, 860, 1506, 1008]. tunas [659, 102]. tunnel [498]. turbid [2070]. turbidity [827, 998, 1231, 293]. turbine [1867]. turbot [634, 808]. turbulence [791]. turbulent [1193, 1156, 1569]. turnover [571, 343, 1738, 1099]. turtle [1939, 1146]. turtles [1210]. tussock [753]. Twentieth [884]. twice [199]. Two [1285, 372, 1797, 1820, 140, 337, 2108, 1981, 891, 725, 1025, 524, 1411, 2097, 616, 1235, 829, 833, 988, 2143, 374, 1468, 1716, 937, 466, 2020, 444, 1043, 1293, 303, 1890, 619, 958, 506, 94, 1683, 726, 1819, 1742]. two-dimensional [337, 829]. two-stage [1819]. tymensis [184]. type [1016, 1527, 2046, 1264, 886, 1643, 177, 1511, 176]. types [2112, 233, 2010, 922, 733, 33, 1818]. tyranny [1113].

U.S [875, 654, 824]. U.S. [341, 533]. U.S.A [822, 722, 107, 556]. U.S.A. [509, 108, 880, 232, 419, 732]. ultimate [1646]. ultraoligotrophic [2070]. Ultraviolet [1509, 571, 1829, 601, 528, 1792, 542, 38, 1704, 1053, 1102, 961, 445, 61, 2139]. ultraviolet-B [61]. Ulvaria [504]. unbiased [455]. uncertain [651, 409]. Uncertainties [1857, 1555, 924]. uncertainty [1620, 1426, 628, 2154, 1694, 1933, 695, 2121]. Unconfounding [1999]. uncovering [1692]. underlying [2022, 1895]. underpinnings [606]. understand [1706, 921]. Understanding [772, 2025, 1799, 202, 260, 668, 1322, 521, 1800]. Underwater [1385, 847, 1676, 1836]. undisturbed [1230]. unequal [1084]. unexploited [1581]. Union [194]. unionid [1835]. unique [57]. unit [288, 418, 1680, 331, 547, 169, 784]. United [1675, 958]. units [2107, 2079, 2104]. univariate [316]. unknown [1143, 423, 1817]. unlimited [544]. unlogged [778]. Unnatural [2099]. unresolved [1486]. unsaturated [688]. unstable [144]. unsuccessful [1654]. Unsupervised [1577]. unsupplemented [2076]. Untangling [51]. untrawlable [746]. upland [1802, 1809, 1951, 1045, 545, 376]. upland-dominated [376]. upland-draining [1802]. upon [1704]. Upper [2160, 2161, 344, 834, 510, 142, 2120, 1401, 932, 1985]. upriver [247]. upriver-migrating [247]. Upstream [826, 1690, 1826, 834, 974, 841, 1332, 1614]. uptake [571, 530, 620, 1347, 1290, 1707, 353]. Upwelling [899, 321, 1313, 516]. urban [217, 1263]. urbanization [394]. urbanized [1547]. urchin [953, 708, 1591, 100]. urchins [218]. Urophycis [940]. ursinus [976]. Ursus [737]. USA [2094, 2076, 1700, 1927, 1165, 1986, 1630, 1514, 946, 2120, 1154, 1959, 921, 1439, 2011]. usage [952]. Use [760, 169, 258, 1019, 671, 527, 64, 1910, 2, 1547, 1178, 167, 1874, 616, 1039, 2067, 388, 1461, 700, 785, 2066, 1676, 402, 1905, 1142, 1310, 1786, 1303, 556, 2140, 713, 1898, 311, 419, 1195, 1348, 496, 1149, 1839, 573]. used [1118, 2141, 2108, 1981, 874, 1111, 2091, 159, 1248, 780, 751]. useful [1361]. uses [673]. Using [1620, 1259, 1301, 580, 78, 729, 1552, 847, 505, 229, 17, 1427, 429, 1116, 147, 1452, 1346, 640, 2027, 869, 407, 236, 43, 341, 1714, 112, 1948, 161, 1566, 241, 1696, 638, 1970, 794, 1582, 1668, 1275, 1404, 2078, 151, 2124, 981, 546, 1929, 1887, 2154, 1454, 2088, 1057, 718, 1892, 1657, 1877, 298, 941, 1271, 214, 1437, 781, 833, 1608, 557, 1689, 2143, 823, 158, 1612, 1849, 844, 843, 1825, 1482, 585, 121, 1889, 598, 290, 1997, 2000, 1922, 1676, 1942, 244, 1483, 98, 2016, 1099, 1512, 1931, 1154, 1712, 1447, 2060, 1467, 1261, 721, 364]. using [135, 1918, 1009, 2042, 1721, 958, 439, 1462, 428, 2062, 692, 297, 1999, 125, 1422, 1819, 88, 42, 87, 10, 1013, 1232, 2032, 1021, 1015, 563, 1724, 1706]. Utah [1514, 1307, 968, 1376, 2027]. utilisation [1737]. utility [1388, 2014]. utilization [1036, 872, 997, 653, 1542, 1074, 2037]. Utilizing [1838]. UV [306]. UV-B [306].

Valdez [509]. validate [1259, 1539]. validated [1862]. Validating [1482, 2035]. Validation [1273, 290, 1918, 2109, 2069, 1795, 981, 856, 1175, 208, 1825, 1154, 1548, 1821, 1065]. Vallentyne [2125]. Valley [95]. Vallisneria [136]. valuable [1427]. valuation [1158]. value [1397, 828, 651]. values [580, 1719]. Vancouver [301, 325]. Varden [443, 295]. variability [195, 1510, 1607, 581, 586, 1977, 1252, 314, 1610, 561, 1622, 250, 328, 2093, 1481, 430, 822, 1140, 2118, 397, 2096, 1846, 100, 499, 892, 252, 537, 884, 96, 1078, 2049, 678, 911, 1013, 61, 595]. Variable [1347, 383, 2078, 1418, 676, 1014, 1419, 28, 1323]. variables [793, 1739, 401, 839, 1305, 965, 1889, 162, 376, 1083, 1439]. Variance [1580, 949, 316, 1311, 1653, 2145]. Variation [1184, 851, 1787, 1037, 1772, 1503, 1355, 842, 146, 236, 1122, 942, 1765, 291, 633, 1719, 1180, 1711, 228, 368, 1235, 378, 1330, 1682, 2128, 497, 1029, 442, 186, 1852, 763, 1342, 175, 1764, 2021, 2048, 1908, 234, 1035, 1873, 1172, 1899, 1002, 1571, 788, 936, 1043, 2113, 1703, 1090, 1782, 387, 1261, 13, 1238, 8, 1832, 1224, 1123, 1506, 1746, 1915, 523, 1302, 1553, 939, 1551, 73, 1952, 1912, 995, 1748]. Variations [84, 469, 470, 272, 711, 1579, 1641, 76, 1642, 773, 215, 1420, 1324, 1790, 726, 920, 1161, 1980, 1874]. various [984]. vary [1292]. varying [1518, 696, 692, 1517]. vaterite [1695]. vegetation [715]. vehicles [1836]. Vejlerne [1009]. veligers [798, 1109, 831]. velocities [1033]. velocity [749, 975, 35, 1297, 974, 714, 436, 1753, 1687, 1554]. vendace [432]. Venice [773]. vent [305]. ventilation [1905]. vents [1928]. Vera [2052]. Verification [2033, 143]. verifies [1361]. vermilion [2104]. Vermont [712]. versa [1634]. versus [407, 1734, 1741, 1234, 1865, 857, 1268, 401, 1240, 840, 1481, 188, 1247, 1558, 612, 1707, 80, 456]. vertebral [2035, 240]. Vertical [1720, 383, 1979, 1562, 1829, 1341, 2015, 897, 1518, 1519, 1646, 898, 1554, 797, 325, 1781, 2012, 1406]. Vertically [901]. very [1446]. vessel [1111, 434, 1271, 1914]. vessels [1483, 1590, 710]. via [1317, 1611, 1779, 817, 2092, 1778]. viability [79, 1375]. vice [1634]. Victoria [1482]. video [847, 1676]. videography [823]. view [66]. viewed [904]. views [1716]. villosus [264, 1341, 599, 1717, 2115, 138, 1828, 516]. Vincent [734]. viral [130]. Virginia [910, 153]. virilis [1020, 2130]. virtual [534, 1097]. viruses [32]. vision [1690, 1509]. visual [1749, 526, 619]. vitellogenin [16]. vitreum [810, 810, 828, 356, 762, 1021]. vitreus [1603, 1502, 847, 1508, 1849, 1402, 1701, 1783, 1248, 1603, 842, 1628, 1748, 1980]. vitro [508]. volcanically [305]. Volume [254, 255, 491, 492, 685, 669, 1055, 1056, 1318, 1319, 1563, 1564, 1727, 1728, 2158, 2159, 35, 2005]. Vorticity [558]. vs [699, 619]. vulnerability [156, 1639, 1651]. vulnerability-independent [1651].

walleye [2155, 1283, 1502, 1629, 524, 847, 810, 1311, 1508, 1604, 1849, 828, 1402, 2055, 2000, 1701, 1783, 373, 356, 1248, 1603, 842, 1628, 1021, 1748, 1980]. walleyes [762]. Waquoit [394]. warm [1949]. warming [907, 1386, 259, 792]. warmwater [1387]. warty [419]. Wash [556]. Washington [1959, 2076, 688, 1797, 863, 880, 435, 245, 1790, 283, 70, 391, 1354]. wastage [1921]. waste [2136]. Water [1741, 1459, 617, 1532, 1753, 623, 1013, 1460, 1993, 481, 615, 1557, 2122, 620, 1178, 983, 1214, 1793, 1822, 1601, 1611, 1779, 1277, 991, 529, 1349, 314, 35, 168, 388, 944, 1768, 2128, 594, 1831, 822, 175, 371, 1089, 598, 1366, 190, 1810, 832, 402, 613, 466, 360, 361, 265, 1913, 662, 71, 436, 217, 486, 425, 587, 296, 884, 165, 303, 1778, 1423, 1387, 696, 18, 717, 405, 523, 149, 1103, 1715, 445, 1742, 213, 1262, 475, 194]. water-body [217]. water-quality [983]. Water-yield [1741]. waterborne [530, 1093]. watermilfoil [1408, 1972]. waters [2047, 1094, 375, 617, 594, 871, 103, 705, 1053, 119, 913, 480, 1888, 1150]. Watershed [1198, 732, 527, 1174, 1471, 1274, 391, 1673, 1286, 1551]. Watershed-level [1198]. watersheds [1088, 1165, 394, 1756, 359, 29]. waterways [341]. wave [841]. wave-drag [841]. way [1113]. weak [873]. weakest [657]. weakfish [2004, 997, 720]. weakly [1524]. weather [1089, 1801]. weathering [1804]. web [1256, 1820, 1292, 128, 981, 1027, 1896, 141, 1608, 469, 154, 1604, 470, 272, 1085, 1543, 1250, 1368, 1535, 567, 1897, 972, 731, 811, 49, 646, 1293, 876, 2148, 1125, 1450, 809]. webs [715, 1088, 118, 1236, 1164, 614, 1902, 1988, 262, 963, 1961, 493, 1652, 2006, 2101]. weight [1091, 1995, 149]. weighting [147]. weights [1817]. welfare [1593]. well [1552]. well-studied [1552]. West [533, 1317, 875, 301, 325]. western [1174, 53, 1629, 1214, 1936, 1501, 897, 76, 435, 212, 2096, 662, 1862, 707, 1790, 283, 391, 1354, 249, 1391, 623, 1854, 906, 573, 797, 1380]. westslope [1788, 821, 2071, 1780]. Wetland [1348, 1805, 1129, 1802, 785, 887, 742, 376, 18, 126, 2085, 110]. wetland- [1802, 376]. wetlands [1460, 632, 2051, 1094, 748, 1348, 1253, 2086, 9, 752]. whale [375, 853, 220, 2090, 516, 23]. whales [1106, 1337, 911, 1288]. where [1706, 270, 92]. Which [694]. White [2031, 464, 1024, 1300, 440, 940, 1212, 1332, 1333, 2038, 2022, 132, 1815]. white-spotted [132]. Whitefish [1719, 1415, 1784, 2143, 1468, 963, 2146, 1600, 294, 2001, 2017, 1912]. whiting [2034, 1843, 2030]. who [270]. Whole [1429, 543, 1351, 154, 49, 2133, 2131]. Whole-lake [1429, 543, 1351, 154, 49, 2131]. wide [485]. widely [2108, 1981]. Widespread [1023, 186]. width [460]. wild [210, 1075, 1114, 1734, 1714, 332, 749, 783, 1977, 1036, 1639, 1874, 1092, 1951, 1432, 1461, 1835, 823, 330, 188, 1524, 727, 782, 917, 1707, 2133, 618, 1128, 1238, 850, 519, 27, 28, 29, 939, 779, 1163, 1561, 1686, 1355, 182]. wild-living [519]. wilderness [763]. Wildfire [739]. wildfires [358]. Will [907]. Willamette [921]. William [357, 509, 649, 536, 90, 686]. Williams [449, 177]. williamsoni [1415]. Winberg [275]. wind [2106, 1702]. Winter [1040, 589, 1587, 966, 1268, 891, 463, 1041, 1411, 1592, 363, 639, 237, 1911, 239, 45, 1760, 415, 698, 94, 790, 1201]. winter-run [589, 1587, 237]. wire [209, 603]. Wisconsin [1094, 1414, 1701, 381]. Wisconsinan [459]. withering [817]. Within [840, 426, 73, 2023, 581, 1661, 1204, 728, 52, 557, 1305, 1663, 65, 1004, 936, 1614, 145, 1254, 1303, 1850, 221, 1302, 240, 2086, 687, 69, 746]. Within- [73]. Within-basin [840]. Within-drainage [426]. within-sample [728]. within-species [581]. within-tributary [1850]. Withler [2161]. without [806, 1094, 1529, 1242, 1819]. wolffish [1490, 662]. wood [500, 1592, 551, 1095, 1974, 1496, 722]. wooded [31]. woody [183, 654, 1244, 283, 573]. worm [1447]. worry [1131]. Wyoming [946].

xanthurus [835, 213].

Y-chromosome [1184]. Year [605, 1854, 407, 385, 1339, 482, 1264, 1039, 226, 572, 360, 361, 232, 771, 1486, 356, 1444, 704, 1895, 2131, 1488]. Year-class [605, 1854, 385, 226]. year-effects [482]. years [1799, 2125, 1941, 1087, 1523, 944, 358, 145, 970, 1999]. Yellow [1892, 711, 1300, 1579, 1298, 1321, 1322, 984, 1108, 253, 1711, 1147, 388, 393, 1071, 1366, 1871, 172, 1691, 75, 485, 825, 917, 1707, 1674, 618, 44, 1253, 1859, 1021]. yellow-phase [825]. yelloweye [871]. yellowfin [2096, 299]. Yellowstone [358, 1064, 1575]. yellowtail [2078]. yessoensis [1971]. yield [1741, 447, 476, 1445, 1140, 1642, 1434, 848, 1782, 525, 930, 1019]. yield-per-recruit [476]. yields [1066]. yolk [808]. yolk-sac [808]. York [1165, 1552, 389, 1849, 75]. young [1339, 1264, 1039, 572, 1911, 1875, 360, 361, 771, 356, 223, 704, 1895, 1488]. young-of-the-year [1264, 1039, 572, 360, 361, 771, 704, 1895]. young-of-year [1339, 356, 1488]. younger [2099]. Yukon [1386].

Zabel [177]. Zealand [1505, 1196, 823, 1995, 215, 753, 687, 1110]. Zebra [485, 370, 562, 818, 831, 798, 1401, 60, 389, 1946, 1769, 101, 65, 75, 811, 876, 165, 1238]. zinc [984, 1707]. Zingel [874, 1375]. Zn [1108, 1353, 172]. zone [831, 798, 1382, 1935, 314, 1504, 1329, 289, 379, 1254, 1083, 1109]. zones [1118, 1346, 31, 1938]. zoogeographic [1416]. zoogeography [1258]. zooplanktivore [806]. zooplanktivory [1556]. Zooplankton [807, 1647, 1954, 225, 195, 2039, 806, 1811, 157, 1829, 231, 770, 424, 1610, 1552, 1958, 1520, 2132, 700, 1405, 647, 900, 1709, 566, 212, 322, 1571, 278, 1477, 1090, 206, 1554, 797, 1808, 462, 1683, 1276, 1503, 1149, 1845, 1109, 475, 1807, 1069].

References

Bernatchez:2000:IBG

[1] Louis Bernatchez and Pierre Duchesne. Individual-based genotype analysis in studies of parentage and population assignment: how many loci, how many alleles? Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57 (1):1–12, January 2000. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f99-271.

Bradford:2000:LUF

[2] Michael J. Bradford and James R. Irvine. Land use, fishing, climate change, and the decline of Thompson River, British Columbia, Coho salmon. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57(1):13–16, January 2000. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f99-283.

Ham:2000:CRS

[3] Kenneth D. Ham and Todd N. Pearsons. Can reduced salmonid population abundance be detected in time to limit management impacts? Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57(1):17–24, January 2000. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f99-175.

Duarte:2000:PLA

[4] C. M. Duarte, S. Agustí, and J. Kalff. Particulate light absorption and the prediction of phytoplankton biomass and planktonic metabolism in northeastern Spanish aquatic ecosystems. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57(1):25–33, January 2000. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f99-184.

Mouneyrac:2000:MLP

[5] C. Mouneyrac, A. Geffard, J. C. Amiard, and C. Amiard-Triquet. Metallothionein-like proteins in Macoma balthica: effects of metal exposure and natural factors. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57 (1):34–42, January 2000. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f99-183.

Millar:2000:BSS

[6] Russell B. Millar and Renate Meyer. Bayesian state-space modeling of age-structured data: fitting a model is just the beginning. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57(1):43–50, January 2000. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f99-169.

Pelletier:2000:MAD

[7] Dominique Pelletier and Jocelyne Ferraris. A multivariate approach for defining fishing tactics from commercial catch and effort data. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57(1):51–65, January 2000. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f99-176.

Rosemond:2000:SSL

[8] Amy D. Rosemond, Patrick J. Mulholland, and Susan H. Brawley. Seasonally shifting limitation of stream periphyton: response of algal populations and assemblage biomass and productivity to variation in light, nutrients, and herbivores. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57 (1):66–75, January 2000. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f99-181.

Zimmer:2000:FII

[9] Kyle D. Zimmer, Mark A. Hanson, and Malcolm G. Butler. Factors influencing invertebrate communities in prairie wetlands: a multivariate approach. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57(1):76–85, January 2000. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f99-180.

Twiss:2000:IIS

[10] Michael R. Twiss, Jean-Christian Auclair, and Murray N. Charlton. An investigation into iron-stimulated phytoplankton productivity in epipelagic Lake Erie during thermal stratification using trace metal clean techniques. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57(1):86–95, January 2000. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/f99-
189.

Weisse:2000:SFB

[11] Thomas Weisse and Erland MacIsaac. Significance and fate of bacterial production in oligotrophic lakes in British Columbia. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57(1):96–105, January 2000. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f99-187.

Chick:2000:EFR

[12] John H. Chick and Michael J. Van Den Avyle. Effects of feeding ration on larval swimming speed and responsiveness to predator attacks: implications for cohort survival. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57 (1):106–115, January 2000. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f99-185. See [89].

Rickman:2000:RVR

[13] S. J. Rickman, N. K. Dulvy, S. Jennings, and J. D. Reynolds. Recruitment variation related to fecundity in marine fishes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57(1):116–124, January 2000. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f99-205.

Yee:2000:ICO

[14] K. A. Yee, E. E. Prepas, P. A. Chambers, J. M. Culp, and G. Scrimgeour. Impact of Ca(OH)_{2} treatment on macroinvertebrate communities in eutrophic hardwater lakes in the Boreal Plain region of Alberta: in situ and laboratory experiments. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57 (1):125–136, January 2000. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f99-182.

Herzka:2000:CIC

[15] Sharon Z. Herzka and G. Joan Holt. Changes in isotopic composition of red drum (Sciaenops ocellatus) larvae in response to dietary shifts: potential applications to settlement studies. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57(1):137–147, January 2000. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f99-174.

Heppell:2000:IGR

[16] Scott A. Heppell and Craig V. Sullivan. Identification of gender and reproductive maturity in the absence of gonads: muscle tissue levels of sex steroids and vitellogenin in gag (Mycteroperca microlepis). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57(1):148–159, January 2000. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f99-188.

Kimura:2000:UNF

[17] Daniel K. Kimura. Using nonlinear functional relationship regression to fit fisheries models. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57(1): 160–170, January 2000. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f99-200.

Snodgrass:2000:ETS

[18] Joel W. Snodgrass, Charles H. Jagoe, A. Lawrence Bryan, Jr., Heather A. Brant, and J. Burger. Effects of trophic status and wetland morphology, hydroperiod, and water chemistry on mercury concentrations in fish. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57(1):171–180, January 2000. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f99-199.

Peterman:2000:CPE

[19] Randall M. Peterman, Brian J. Pyper, and Jeff A. Grout. Comparison of parameter estimation methods for detecting climate-induced changes in productivity of Pacific salmon (Oncorhynchus spp.). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57(1):181–191, January 2000. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f99-204.

Mills:2000:RFP

[20] K. H. Mills, S. M. Chalanchuk, and D. J. Allan. Recovery of fish populations in Lake 223 from experimental acidification. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57(1):192–204, January 2000. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f99-186.

Essington:2000:IIS

[21] Timothy E. Essington, Thomas P. Quinn, and Victor E. Ewert. Intraand inter-specific competition and the reproductive success of sympatric Pacific salmon. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57(1):205–213, January 2000. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f99-198.

Xiao:2000:DAG

[22] Yongshun Xiao and Terence I. Walker. Demographic analysis of gummy shark (Mustelus antarcticus) and school shark (Galeorhinus galeus) off southern Australia by applying a generalized Lotka equation and its dual equation. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57(1):214–222, January 2000. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f99-224.

Whitehead:2000:DDH

[23] Hal Whitehead. Density-dependent habitat selection and the modeling of sperm whale (Physeter macrocephalus) exploitation. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57(1):223–230, January 2000. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f99-227.

Jacoby:2000:EFA

[24] Jean M. Jacoby, Diane C. Collier, Eugene B. Welch, F. Joan Hardy, and Michele Crayton. Environmental factors associated with a toxic bloom of Microcystis aeruginosa. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57 (1):231–240, January 2000. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f99-234.

Manske:2000:ESR

[25] Micheline Manske and Carl J. Schwarz. Estimates of stream residence time and escapement based on capture-recapture data. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57(1):241–246, January 2000. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f99-203.

Schwarz:2000:EJS

[26] Carl J. Schwarz and Wayne T. Stobo. Estimation of juvenile survival, adult survival, and age-specific pupping probabilities for the female grey seal (Halichoerus gryprus) on Sable Island from capture-recapture data. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57(1):247–253, January 2000. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f99-259.

Smith:2000:TWAa

[27] Barry D. Smith, Bruce R. Ward, and David W. Welch. Trends in wild adult steelhead (Oncorhynchus mykiss) abundance in British Columbia as indexed by angler success. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57 (2):255–270, February 2000. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f99-256.

Smith:2000:TWAb

[28] Barry D. Smith and Bruce R. Ward. Trends in wild adult steelhead (Oncorhynchus mykiss) abundance for coastal regions of British Columbia support the variable marine survival hypothesis. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57(2):271–284, February 2000. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f99-254.

Smith:2000:TWAc

[29] Barry D. Smith. Trends in wild adult steelhead (Oncorhynchus mykiss) abundance for snowmelt-driven watersheds of British Columbia in relation to freshwater discharge. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57 (2):285–297, February 2000. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f99-255.

Ward:2000:DSO

[30] Bruce R. Ward. Declivity in steelhead (Oncorhynchus mykiss) recruitment at the Keogh River over the past decade. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57(2):298–306, February 2000. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f99-243.

Stauffer:2000:RWR

[31] J. C. Stauffer, R. M. Goldstein, and R. M. Newman. Relationship of wooded riparian zones and runoff potential to fish community composition in agricultural streams. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57 (2):307–316, February 2000. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f99-197.

Wilhelm:2000:BCP

[32] Steven W. Wilhelm and Ralph E. H. Smith. Bacterial carbon production in Lake Erie is influenced by viruses and solar radiation. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57(2):317–326, February 2000. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f99-202.

Munger:2000:IIR

[33] Catherine Munger and Landis Hare. Influence of ingestion rate and food types on cadmium accumulation by the aquatic insect Chaoborus. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57(2):327–332, February 2000. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f99-225.

Little:2000:PTS

[34] Joanne L. Little, Roland I. Hall, Roberto Quinlan, and John P. Smol. Past trophic status and hypolimnetic anoxia during eutrophicaton and remediation of Gravenhurst Bay, Ontario: comparison of diatoms, chironomids, and historical records. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57 (2):333–341, February 2000. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f99-235.

Flore:2000:RFC

[35] L. Flore, W. Reckendorfer, and H. Keckeis. Reaction field, capture field, and search volume of 0+ nase (Chondrostoma nasus): effects of body size and water velocity. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57(2): 342–350, February 2000. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f99-236.

Arendt:2000:PDO

[36] Jeffrey D. Arendt and David S. Wilson. Population differences in the onset of cranial ossification in pumpkinseed (Lepomis gibbosus), a potential cost of rapid growth. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57(2): 351–356, February 2000. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f99-250.

Babcock:2000:DPM

[37] Elizabeth A. Babcock and Ellen K. Pikitch. A dynamic programming model of fishing strategy choice in a multispecies trawl fishery with trip limits. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57(2):357–370, February 2000. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f99-257.

Marwood:2000:PNP

[38] Christopher A. Marwood, Ralph E. H. Smith, John A. Furgal, Murray N. Charlton, Keith R. Solomon, and Bruce M. Greenberg. Photoinhibition of natural phytoplankton assemblages in Lake Erie exposed to solar ultraviolet radiation. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57 (2):371–379, February 2000. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f99-258.

Cain:2000:MEB

[39] Daniel J. Cain, James L. Carter, Steven V. Fend, Samuel N. Luoma, Charles N. Alpers, and Howard E. Taylor. Metal exposure in a benthic macroinvertebrate, Hydropsyche californica, related to mine drainage in the Sacramento River. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57(2): 380–390, February 2000. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f99-260.

Senanan:2000:GRA

[40] Wansuk Senanan and Anne R. Kapuscinski. Genetic relationships among populations of northern pike (Esox lucius). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57(2):391–404, February 2000. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f99-261.

Wilson:2000:CPR

[41] J. Y. Wilson, R. F. Addison, D. Martens, R. Gordon, and B. Glickman. Cytochrome P450 1A and related measurements in juvenile Chinook salmon (Oncorhynchus tshawytscha) from the Fraser River. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57(2):405–413, February 2000. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f99-251.

Trudel:2000:EFC

[42] Marc Trudel, Alain Tremblay, Roger Schetagne, and Joseph B. Rasmussen. Estimating food consumption rates of fish using a mercury mass balance model. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57(2): 414–428, February 2000. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f99-262.

Agnew:2000:AAM

[43] D. J. Agnew, C. P. Nolan, J. R. Beddington, and R. Baranowski. Approaches to the assessment and management of multispecies skate and ray fisheries using the Falkland Islands fishery as an example. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57(2):429–440, February 2000. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f99-264.

Sherwood:2000:BCH

[44] Graham D. Sherwood, Joseph B. Rasmussen, David J. Rowan, Julie Brodeur, and Alice Hontela. Bioenergetic costs of heavy metal exposure in yellow perch (Perca flavescens): in situ estimates with a radiotracer (^{137}Cs) technique. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57 (2):441–450, February 2000. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f99-268.

Lappalainen:2000:SDW

[45] Jyrki Lappalainen, Vaike Erm, Jakob Kjellman, and Hannu Lehtonen. Size-dependent winter mortality of age-0 pikeperch (Stizostedion lucioperca) in Pärnu Bay, the Baltic Sea. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57(2):451–458, February 2000. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f99-270.

Bjorkstedt:2000:SRR

[46] Eric P. Bjorkstedt. Stock-recruitment relationships for life cycles that exhibit concurrent density dependence. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57(2):459–467, February 2000. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f99-263.

Isaak:2000:TPA

[47] Daniel J. Isaak and Wayne A. Hubert. Are trout populations affected by reach-scale stream slope? Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57 (2):468–477, February 2000. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f99-272.

Kimmerer:2000:AES

[48] W. J. Kimmerer, J. H. Cowan, Jr., L. W. Miller, and K. A. Rose. Analysis of an estuarine striped bass (Morone saxatilis) population: influence of density-dependent mortality between metamorphosis and recruitment. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57(2):478–486, February 2000. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f99-273.

Pace:2000:EWL

[49] Michael L. Pace and Jonathan J. Cole. Effects of whole-lake manipulations of nutrient loading and food web structure on planktonic respiration. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57(2):487–496, February 2000. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f99-279.

Bardonnet:2000:FHA

[50] Agnès Bardonnet and Jean-Luc Baglinière. Freshwater habitat of Atlantic salmon (Salmo salar). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57 (2):497–506, February 2000. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f99-226.

Millar:2000:UCS

[51] Russell B. Millar. Untangling the confusion surrounding the estimation of gillnet selectivity. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57(2): 507–511, February 2000. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f99-275.

Frank:2000:AEC

[52] Kenneth T. Frank and David Brickman. Allee effects and compensatory population dynamics within a stock complex. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57(3):513–517, March 2000. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f00-024.

Campbell:2000:EPR

[53] Robert W. Campbell and Erica J. H. Head. Egg production rates of Calanus finmarchicus in the western North Atlantic: effect of gonad maturity, female size, chlorophyll concentration, and temperature. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57(3):518–529, March 2000. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f99-278.

Hill:2000:CST

[54] Walter R. Hill, Angela T. Bednarek, and I. Lauren Larsen. Cadmium sorption and toxicity in autotrophic biofilms. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57(3):530–537, March 2000. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f99-286.

Klug:2000:FIG

[55] Jennifer L. Klug and Janet M. Fischer. Factors influencing the growth of Mougeotia in experimentally acidified mesocosms. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57(3):538–547, March 2000. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f99-274.

Essington:2000:RSF

[56] Timothy E. Essington, James R. Hodgson, and James F. Kitchell. Role of satiation in the functional response of a piscivore, largemouth bass (Micropterus salmoides). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57 (3):548–556, March 2000. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f99-289.

Bjerselius:2000:DBE

[57] Rickard Bjerselius, Weiming Li, John H. Teeter, James G. Seelye, Peter B. Johnsen, Peter J. Maniak, Gerold C. Grant, Christine N. Polkinghorne, and Peter W. Sorensen. Direct behavioral evidence that unique bile acids released by larval sea lamprey (Petromyzon marinus) function as a migratory pheromone. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57 (3):557–569, March 2000. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f99-290.

Findlay:2000:EIP

[58] C. Scott Findlay, Daniel G. Bert, and Ligang Zheng. Effect of introduced piscivores on native minnow communities in Adirondack lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57(3):570–580, March 2000. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f99-276.

Gaudreau:2000:IMP

[59] Nathalie Gaudreau and Daniel Boisclair. Influence of moon phase on acoustic estimates of the abundance of fish performing daily horizontal migration in a small oligotrophic lake. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57(3):581–590, March 2000. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f99-277.

Frischer:2000:IBZ

[60] Marc E. Frischer, Sandra A. Nierzwicki-Bauer, Robert H. Parsons, Kanda Vathanodorn, and Kelli R. Waitkus. Interactions between zebra mussels (Dreissena polymorpha) and microbial communities. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57(3):591–599, March 2000. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f00-001.

Xenopoulos:2000:IUB

[61] Marguerite A. Xenopoulos, Yves T. Prairie, and David F. Bird. Influence of ultraviolet-B radiation, stratospheric ozone variability, and thermal stratification on the phytoplankton biomass dynamics in a mesohumic lake. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57(3):600–609, March 2000. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f99-269.

Edlund:2000:TBG

[62] M. B. Edlund, C. M. Taylor, C. L. Schelske, and E. F. Stoermer. Thalassiosira baltica (Grunow) Ostenfeld (Bacillariophyta), a new exotic species in the Great Lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57 (3):610–615, March 2000. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f99-284.

Botsford:2000:ACA

[63] Louis W. Botsford and Charles M. Paulsen. Assessing covariability among populations in the presence of intraseries correlation: Columbia River spring-summer Chinook salmon (Oncorhynchus tshawytscha) stocks. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57(3):616–627, March 2000. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f00-007.

Berejikian:2000:SDG

[64] Barry A. Berejikian, E. Paul Tezak, Thomas A. Flagg, Anita L. LaRae, Eric Kummerow, and Conrad V. W. Mahnken. Social dominance, growth, and habitat use of age-0 steelhead (Oncorhynchus mykiss) grown in enriched and conventional hatchery rearing environments. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57(3):628–636, March 2000. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f99-288.

Lewis:2000:PGZ

[65] Kristin M. Lewis, Jeffrey L. Feder, and Gary A. Lamberti. Population genetics of the zebra mussel, Dreissena polymorpha (Pallas): local allozyme differentiation within midwestern lakes and streams. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57(3):637–643, March 2000. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f00-008.

Rose:2000:DSO

[66] G. A. Rose, B. deYoung, D. W. Kulka, S. V. Goddard, and G. L. Fletcher. Distribution shifts and overfishing the northern cod (Gadus morhua): a view from the ocean. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57 (3):644–663, March 2000. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f00-004.

Barrowman:2000:SMS

[67] Nicholas J. Barrowman and Ransom A. Myers. Still more spawnerrecruitment curves: the hockey stick and its generalizations. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57(4):665–676, April 2000. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f99-282.

Bradford:2000:RPC

[68] Michael J. Bradford, Ransom A. Myers, and James R. Irvine. Reference points for Coho salmon (Oncorhynchus kisutch) harvest rates and escapement goals based on freshwater production. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57(4):677–686, April 2000. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f99-281.

Witt:2000:CSD

[69] Jonathan D. S. Witt and Paul D. N. Hebert. Cryptic species diversity and evolution in the amphipod genus Hyalella within central glaciated North America: a molecular phylogenetic approach. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57(4):687–698, April 2000. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f99-285.

Rot:2000:SCC

[70] Byron W. Rot, Robert J. Naiman, and Robert E. Bilby. Stream channel configuration, landform, and riparian forest structure in the Cascade Mountains, Washington. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57 (4):699–707, April 2000. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f00-002.

Nolan:2000:ATE

[71] D. T. Nolan, R. H. Hadderingh, F. A. T. Spanings, H. A. Jenner, and S. E. Wendelaar Bonga. Acute temperature elevation in tap and Rhine water affects skin and gill epithelia, hydromineral balance, and gill Na^{+} /K^{+}-ATPase activity of brown trout (Salmo trutta) smolts. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57(4):708–718, April 2000. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f99-252.

Valdimarsson:2000:EDD

[72] Sveinn K. Valdimarsson, Neil B. Metcalfe, and Skúli Skúlason. Experimental demonstration of differences in sheltering behaviour between Icelandic populations of Atlantic salmon (Salmo salar) and Arctic char (Salvelinus alpinus). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57 (4):719–724, April 2000. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f99-253.

VanderZanden:2000:WAP

[73] M. Jake Vander Zanden, Brian J. Shuter, Nigel P. Lester, and Joseph B. Rasmussen. Within- and among-population variation in the trophic position of a pelagic predator, lake trout (Salvelinus namaycush). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57(4):725–731, April 2000. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f00-011.

Bromhead:2000:AFC

[74] Don Bromhead, John Kalish, and Paul Waring. Application of flow cytometric cell cycle analysis to the assessment of condition and growth in larvae of a freshwater teleost Galaxias olidus. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57(4):732–741, April 2000. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f00-014.

Mayer:2000:RYP

[75] C. M. Mayer, A. J. VanDeValk, J. L. Forney, L. G. Rudstam, and E. L. Mills. Response of yellow perch (Perca flavescens) in Oneida Lake, New York, to the establishment of zebra mussels (Dreissena polymorpha). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57(4):742–754, April 2000. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f00-009.

Incze:2000:NPA

[76] Lewis S. Incze, Peter Aas, Terri Ainaire, and Marcia Bowen. Neustonic postlarval American lobsters, Homarus americanus, in the western Gulf of Maine: spatial and interannual variations. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57(4):755–765, April 2000. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f00-005.

Rosenfeld:2000:HFA

[77] Jordan Rosenfeld, Marc Porter, and Eric Parkinson. Habitat factors affecting the abundance and distribution of juvenile cutthroat trout (Oncorhynchus clarki) and Coho salmon (Oncorhynchus kisutch). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57(4):766–774, April 2000. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f00-010.

Fabrizio:2000:ULM

[78] Mary C. Fabrizio, Jonathan Raz, and Rajesh Ramanath Bandekar. Using linear models with correlated errors to analyze changes in abundance of Lake Michigan fishes: 1973–1992. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57(4):775–788, April 2000. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f00-020.

Keckeis:2000:EFA

[79] Hubert Keckeis, Elisabeth Bauer-Nemeschkal, Vladimir V. Menshutkin, Hans Leo Nemeschkal, and Ewa Kamler. Effects of female attributes and egg properties on offspring viability in a rheophilic cyprinid, Chondrostoma nasus. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57(4):789–796, April 2000. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f00-006.

Tessier:2000:GAS

[80] Nathalie Tessier and Louis Bernatchez. A genetic assessment of single versus double origin of landlocked Atlantic salmon (Salmo salar) from Lake Saint-Jean, Québec, Canada. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57(4):797–804, April 2000. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f00-029.

Hutchings:2000:TCH

[81] Jeffrey A. Hutchings and Mark Ferguson. Temporal changes in harvesting dynamics of Canadian inshore fisheries for northern Atlantic cod, Gadus morhua. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57(4):805–814, April 2000. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f00-021.

Lambert:2000:ECR

[82] Yvan Lambert and Jean-Denis Dutil. Energetic consequences of reproduction in Atlantic cod (Gadus morhua) in relation to spawning level of somatic energy reserves. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57 (4):815–825, April 2000. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f00-022.

Dutil:2000:NMP

[83] Jean-Denis Dutil and Yvan Lambert. Natural mortality from poor condition in Atlantic cod (Gadus morhua). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57(4):826–836, April 2000. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f00-023.

Barille:2000:VPO

[84] Laurent Barillé, Joël Haure, Bruno Cognie, and Anne Leroy. Variations in pallial organs and eulatero-frontal cirri in response to high particulate matter concentrations in the oyster Crassostrea gigas. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57(4):837–843, April 2000. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f00-012.

Meeuwig:2000:PCE

[85] Jessica J. Meeuwig, Pirkko Kauppila, and Heikki Pitkänen. Predicting coastal eutrophication in the Baltic: a limnological approach. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57(4):844–855, April 2000. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f00-013.

Benoit:2000:PMA

[86] Hugues P. Benoît, Pierre Pepin, and Joseph A. Brown. Patterns of metamorphic age and length in marine fishes, from individuals to taxa. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57(4):856–869, April 2000. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f00-019.

Twiss:2000:EII

[87] Michael R. Twiss, Jean-Christian Auclair, and Murray N. Charlton. Erratum: an investigation into iron-stimulated phytoplankton productivity in epipelagic Lake Erie during thermal stratification using trace metal clean techniques. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57(4):870, April 2000. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f00-060.

Trudel:2000:EEF

[88] Marc Trudel, Alain Tremblay, Roger Schetagne, and Joseph B. Rasmussen. Erratum: Estimating food consumption rates of fish using a mercury mass balance model. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57 (4):871, April 2000. CODEN CJFSDX. ISSN 0706-652X (print), 1205-

7533 (electronic). URL https://cdnsciencepub.com/doi/full/10.
1139/f00-059.

Chick:2000:EEF

[89] John H. Chick and Michael J. Van Den Avyle. Erratum: Effects of feeding ration on larval swimming speed and responsiveness to predator attacks: implications for cohort survival. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57(4):872, April 2000. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f00-058. See [12].

Anonymous:2000:ONW

[90] Anonymous. Obituary / nécrologie: William Alexander (Bill) Kennedy. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57(4):873–874, April 2000. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f00-052.

Lindell:2000:SPD

[91] Måns J. Lindell, H. Wilhelm Granéli, and Stefan Bertilsson. Seasonal photoreactivity of dissolved organic matter from lakes with contrasting humic content. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57(5):875–885, May 2000. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/f00-
016.

Smith:2000:BSE

[92] David R. Smith, Kenneth P. Burnham, Desmond M. Kahn, Xi He, Cynthia J. Goshorn, Kathryn A. Hattala, and Andrew W. Kahnle. Bias in survival estimates from tag-recovery models where catch-and-release is common, with an example from Atlantic striped bass (Morone saxatilis). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57(5):886–897, May 2000. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f00-027.

Bowen:2000:RPD

[93] W. D. Bowen. Reconstruction of pinniped diets: accounting for complete digestion of otoliths and cephalopod beaks. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57(5):898–905, May 2000. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f00-032.

Solazzi:2000:EIW

[94] M. F. Solazzi, T. E. Nickelson, S. L. Johnson, and J. D. Rodgers. Effects of increasing winter rearing habitat on abundance of salmonids in two coastal Oregon streams. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57 (5):906–914, May 2000. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f00-030.

Banks:2000:AMD

[95] Michael A. Banks, Vanessa K. Rashbrook, Marco J. Calavetta, Cheryl A. Dean, and Dennis Hedgecock. Analysis of microsatellite DNA resolves genetic structure and diversity of Chinook salmon (Oncorhynchus tshawytscha) in California’s Central Valley. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57(5):915–927, May 2000. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f00-034.

Rennie:2000:SVS

[96] Colin D. Rennie and Robert G. Millar. Spatial variability of stream bed scour and fill: a comparison of scour depth in chum salmon (Oncorhynchus keta) redds and adjacent bed. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57(5):928–938, May 2000. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f00-038.

Hill:2000:CSO

[97] James A. Hill, Anders Kiessling, and Robert H. Devlin. Coho salmon (Oncorhynchus kisutch) transgenic for a growth hormone gene construct exhibit increased rates of muscle hyperplasia and detectable levels of differential gene expression. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57 (5):939–950, May 2000. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f00-015.

Mitro:2000:PFA

[98] Matthew G. Mitro and Alexander V. Zale. Predicting fish abundance using single-pass removal sampling. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57(5):951–961, May 2000. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/

doi/full/10.1139/f00-025.

Schnute:2000:BDA

[99] Jon T. Schnute, Al Cass, and Laura J. Richards. A Bayesian decision analysis to set escapement goals for Fraser River sockeye salmon (Oncorhynchus nerka). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57(5):962– 979, May 2000. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f00-017.

Morgan:2000:SVG

[100] Lance E. Morgan, Louis W. Botsford, Stephen R. Wing, and Barry D. Smith. Spatial variability in growth and mortality of the red sea urchin, Strongylocentrotus franciscanus, in northern California. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57(5):980–992, May 2000. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f00-046.

Kraft:2000:RDR

[101] Clifford E. Kraft and Ladd E. Johnson. Regional differences in rates and patterns of North American inland lake invasions by zebra mussels (Dreissena polymorpha). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57 (5):993–1001, May 2000. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f00-037.

Hampton:2000:NMR

[102] John Hampton. Natural mortality rates in tropical tunas: size really does matter. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57(5):1002–1010, May 2000. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f99-287.

Lawson:2000:SSS

[103] Gareth L. Lawson and George A. Rose. Small-scale spatial and temporal patterns in spawning of Atlantic cod (Gadus morhua) in coastal Newfoundland waters. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57 (5):1011–1024, May 2000. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f00-026.

Madenjian:2000:NAT

[104] Charles P. Madenjian, Daniel V. O’Connor, and David A. Nortrup. A new approach toward evaluation of fish bioenergetics models. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57(5):1025–1032, May 2000. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f99-280.

Penn:2000:SPR

[105] Michael R. Penn, Martin T. Auer, Susan M. Doerr, Charles T. Driscoll, Carol M. Brooks, and Steven W. Effler. Seasonality in phosphorus release rates from the sediments of a hypereutrophic lake under a matrix of pH and redox conditions. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57 (5):1033–1041, May 2000. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f00-035.

Willis:2000:PFS

[106] Theodore V. Willis and John J. Magnuson. Patterns in fish species composition across the interface between streams and lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57(5):1042–1052, May 2000. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f00-028.

Wiener:2000:MSR

[107] James G. Wiener and Pamela J. Shields. Mercury in the Sudbury River (Massachusetts, U.S.A.): pollution history and a synthesis of recent research. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57(5):1053–1061, May 2000. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f00-039.

Frazier:2000:SHA

[108] Bradley E. Frazier, James G. Wiener, Ronald G. Rada, and Daniel R. Engstrom. Stratigraphy and historic accumulation of mercury in recent depositional sediments in the Sudbury River, Massachusetts, U.S.A. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57(5):1062–1072, May 2000. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f00-044.

Colman:2000:SMS

[109] John A. Colman and Robert F. Breault. Sampling for mercury at subnanogram per litre concentrations for load estimation in rivers. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57(5):1073–1079, May 2000. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f00-040.

Waldron:2000:DHT

[110] Marcus C. Waldron, John A. Colman, and Robert F. Breault. Distribution, hydrologic transport, and cycling of total mercury and methyl mercury in a contaminated river-reservoir-wetland system (Sudbury River, eastern Massachusetts). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57 (5):1080–1091, May 2000. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f00-041.

Naimo:2000:BSA

[111] Teresa J. Naimo, James G. Wiener, W. Gregory Cope, and Nicolas S. Bloom. Bioavailability of sediment-associated mercury to Hexagenia mayflies in a contaminated floodplain river. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57(5):1092–1102, May 2000. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f00-043.

Beckvar:2000:SAM

[112] Nancy Beckvar, Sandra Salazar, Michael Salazar, and Ken Finkelstein. An in situ assessment of mercury contamination in the Sudbury River, Massachusetts, using transplanted freshwater mussels (Elliptio complanata). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57(5):1103–1112, May 2000. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f00-042.

Smoot:2000:DEG

[113] James C. Smoot and Robert H. Findlay. Digestive enzyme and gut surfactant activity of detritivorous gizzard shad (Dorosoma cepedianum). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57(6):1113–1119, June 2000. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f00-036.

Lapointe:2000:MPS

[114] Michel Lapointe, Brett Eaton, Steve Driscoll, and Christian Latulippe. Modelling the probability of salmonid egg pocket scour due to floods. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57(6):1120–1130, June 2000. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f00-033.

Whitledge:2000:DSD

[115] Gregory W. Whitledge and Robert S. Hayward. Determining sampling date interval for precise in situ estimates of cumulative food consumption by fishes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57(6):1131–1138, June 2000. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f00-003.

Colombo:2000:DFC

[116] Juan C. Colombo, Claudio Bilos, Mauricio Remes Lenicov, Dario Colautti, Patricia Landoni, and Charles Brochu. Detritivorous fish contamination in the Río de la Plata estuary: a critical accumulation pathway in the cycle of anthropogenic compounds. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57(6):1139–1150, June 2000. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f00-031.

Campbell:2000:EAS

[117] Morag M. Campbell, Keith N. White, Ravin Jugdaohsingh, Jonathan J. Powell, and Catherine R. McCrohan. Effect of aluminum and silicic acid on the behaviour of the freshwater snail Lymnaea stagnalis. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57(6):1151–1159, June 2000. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f00-053.

Berglund:2000:BDP

[118] Olof Berglund, Per Larsson, Göran Ewald, and Lennart Okla. Bioaccumulation and differential partitioning of polychlorinated biphenyls in freshwater, planktonic food webs. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57(6):1160–1168, June 2000. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f00-045.

Mueter:2000:CSC

[119] Franz J. Mueter and Brenda L. Norcross. Changes in species composition of the demersal fish community in nearshore waters of Kodiak Island, Alaska. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57(6):1169–1180, June 2000. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f00-051.

Friedland:2000:GPL

[120] Kevin D. Friedland and David G. Reddin. Growth patterns of Labrador Sea Atlantic salmon postsmolts and the temporal scale of recruitment synchrony for North American salmon stocks. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57(6):1181–1189, June 2000. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f00-069.

Kerans:2000:ADD

[121] B. L. Kerans, Peter L. Chesson, and Roy A. Stein. Assessing densitydependent establishment and dispersal: an example using caddisfly larvae. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57(6):1190–1199, June 2000. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f00-047.

Winger:2000:FAS

[122] Paul D. Winger, Pingguo He, and Stephen J. Walsh. Factors affecting the swimming endurance and catchability of Atlantic cod (Gadus morhua). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57(6):1200–1207, June 2000. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f00-049.

Scheffer:2000:EFP

[123] Marten Scheffer, Sergio Rinaldi, and Yuri A. Kuznetsov. Effects of fish on plankton dynamics: a theoretical analysis. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57(6):1208–1219, June 2000. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f00-018.

Paul:2000:DDI

[124] Andrew J. Paul, John R. Post, George L. Sterling, and Carl Hunt. Density-dependent intercohort interactions and recruitment dynamics: models and a bull trout (Salvelinus confluentus) time series. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57(6):1220–1231, June 2000. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f00-048.

Thorrold:2000:SAT

[125] Simon R. Thorrold and Stephen Shuttleworth. In situ analysis of trace elements and isotope ratios in fish otoliths using laser ablation sector field inductively coupled plasma mass spectrometry. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57(6):1232–1242, June 2000. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f00-054.

Tanner:2000:FIC

[126] Danny K. Tanner, John C. Brazner, and Valerie J. Brady. Factors influencing carbon, nitrogen, and phosphorus content of fish from a Lake Superior coastal wetland. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57 (6):1243–1251, June 2000. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f00-062.

Morbey:2000:PPS

[127] Yolanda Morbey. Protandry in Pacific salmon. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57(6):1252–1257, June 2000. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f00-064.

Campbell:2000:OTF

[128] Linda M. Campbell, David W. Schindler, Derek C. G. Muir, David B. Donald, and Karen A. Kidd. Organochlorine transfer in the food web of subalpine Bow Lake, Banff National Park. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57(6):1258–1269, June 2000. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f00-067.

Sigler:2000:AEC

[129] Michael F. Sigler. Abundance estimation and capture of sablefish (Anoplopoma fimbria) by longline gear. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57(6):1270–1283, June 2000. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f00-066.

Ricciardi-Rigault:2000:CSV

[130] Mahaut Ricciardi-Rigault, David F. Bird, and Yves T. Prairie. Changes in sediment viral and bacterial abundances with hypolimnetic oxygen depletion in a shallow eutrophic Lac Brome (Quebec, Canada). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57(6):1284–1290, June 2000. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f00-068.

Vigliola:2000:CTB

[131] Laurent Vigliola, Mireille Harmelin-Vivien, and Mark G. Meekan. Comparison of techniques of back-calculation of growth and settlement marks from the otoliths of three species of Diplodus from the Mediterranean Sea. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57(6):1291–1299, June 2000. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f00-055.

Morita:2000:ELH

[132] Kentaro Morita, Shoichiro Yamamoto, and Noboru Hoshino. Extreme life history change of white-spotted char (Salvelinus leucomaenis) after damming. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57(6):1300–1306, June 2000. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/f00-
050.

Holland:2000:BMM

[133] Daniel S. Holland. A bioeconomic model of marine sanctuaries on Georges Bank. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57(6):1307– 1319, June 2000. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f00-061.

Swain:2000:PFC

[134] D. P. Swain and A. F. Sinclair. Pelagic fishes and the cod recruitment dilemma in the Northwest Atlantic. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57(7):1321–1325, July 2000. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f00-104.

Roussel:2000:FTN

[135] J.-M. Roussel, A. Haro, and R. A. Cunjak. Field test of a new method for tracking small fishes in shallow rivers using passive integrated transponder (PIT) technology. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57 (7):1326–1329, July 2000. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f00-110.

St-Cyr:2000:BSB

[136] Louise St-Cyr and Peter G. C. Campbell. Bioavailability of sedimentbound metals for Vallisneria americana Michx., a submerged aquatic plant, in the St. Lawrence River. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57(7):1330–1341, July 2000. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f00-078.

Scrimgeour:2000:CEP

[137] Garry J. Scrimgeour and Patricia A. Chambers. Cumulative effects of pulp mill and municipal effluents on epilithic biomass and nutrient limitation in a large northern river ecosystem. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57(7):1342–1354, July 2000. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f00-057.

ODriscoll:2000:PCS

[138] Richard L. O’Driscoll, David C. Schneider, George A. Rose, and George R. Lilly. Potential contact statistics for measuring scaledependent spatial pattern and association: an example of northern cod (Gadus morhua) and capelin (Mallotus villosus). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57(7):1355–1368, July 2000. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f00-079.

Rosenfeld:2000:EFP

[139] Jordan Rosenfeld. Effects of fish predation in erosional and depositional habitats in a temperate stream. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57(7):1369–1379, July 2000. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f00-073.

Biggs:2000:TCS

[140] Barry J. F. Biggs, Steven N. Francoeur, Alexander D. Huryn, Roger Young, Christopher J. Arbuckle, and Colin R. Townsend. Trophic cascades in streams: effects of nutrient enrichment on autotrophic and consumer benthic communities under two different fish predation regimes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57(7):1380–1394, July 2000. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f00-077.

Harvey:2000:SIE

[141] Chris J. Harvey and James F. Kitchell. A stable isotope evaluation of the structure and spatial heterogeneity of a Lake Superior food web. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57(7):1395–1403, July 2000. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f00-072.

Maes:2000:BTU

[142] Joachim Maes and Frans Ollevier. Biomass transport to and from an upper estuarine area by migration of juvenile Atlantic herring Clupea harengus. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57(7):1404–1409, July 2000. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f00-083.

Clason:2000:BTM

[143] B. Clason and G.-P. Zauke. Bioaccumulation of trace metals in marine and estuarine amphipods: evaluation and verification of toxicokinetic models. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57(7):1410–1422, July 2000. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f00-063.

Poleo:2000:EUA

[144] Antonio B. S. Poléo and Frode Bjerkely. Effect of unstable aluminium chemistry on Arctic char (Salvelinus alpinus). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57(7):1423–1433, July 2000. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f00-084.

Mullin:2000:CWB

[145] Michael M. Mullin, Erica Goetze, Stace E. Beaulieu, and Jacqueline M. Lasker. Comparisons within and between years resulting in contrasting recruitment of Pacific hake (Merluccius productus) in the California Current system. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57(7):1434– 1447, July 2000. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f00-070.

Adams:2000:GVS

[146] S. Adams, B. D. Mapstone, G. R. Russ, and C. R. Davies. Geographic variation in the sex ratio, sex specific size, and age structure of Plectropomus leopardus (Serranidae) between reefs open and closed to fishing on the Great Barrier Reef. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57 (7):1448–1458, July 2000. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f00-076.

Merritt:2000:UPD

[147] Margaret F. Merritt and Terrance J. Quinn II. Using perceptions of data accuracy and empirical weighting of information: assessment of a recreational fish population. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57 (7):1459–1469, July 2000. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f00-065.

Baxter:2000:GHE

[148] Colden V. Baxter and F. Richard Hauer. Geomorphology, hyporheic exchange, and selection of spawning habitat by bull trout (Salvelinus confluentus). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57(7):1470–1481, July 2000. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f00-056.

Tang:2000:IBW

[149] Ming Tang, Daniel Boisclair, Chantal Ménard, and John A. Downing. Influence of body weight, swimming characteristics, and water temperature on the cost of swimming in brook trout (Salvelinus fontinalis). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57(7):1482–1488, July 2000. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f00-080.

Leggett:2000:IIN

[150] M. F. Leggett, O. Johannsson, R. Hesslein, D. G. Dixon, W. D. Taylor, and M. R. Servos. Influence of inorganic nitrogen cycling on the δ 15N of Lake Ontario biota. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57 (7):1489–1496, July 2000. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f00-085.

Crowder:2000:ESE

[151] David W. Crowder and Panayiotis Diplas. Evaluating spatially explicit metrics of stream energy gradients using hydrodynamic model simulations. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57(7):1497–1507, July 2000. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f00-074.

Lonzarich:2000:ERL

[152] D. G. Lonzarich, M. R. Lonzarich, and M. L. Warren, Jr. Effects of riffle length on the short-term movement of fishes among stream pools. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57(7):1508–1514, July 2000. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f00-082.

Bulger:2000:CRP

[153] Arthur J. Bulger, Bernard J. Cosby, and James R. Webb. Current, reconstructed past, and projected future status of brook trout (Salvelinus fontinalis) streams in Virginia. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57(7):1515–1523, July 2000. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f00-086.

Houser:2000:FWS

[154] Jeffrey N. Houser, Stephen R. Carpenter, and Jon J. Cole. Food web structure and nutrient enrichment: effects on sediment phosphorus retention in whole-lake experiments. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57(7):1524–1533, July 2000. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f00-075.

Suthers:2000:SLC

[155] Iain Suthers. Significance of larval condition: comment on laboratory experiments. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57(7):1534–1536, July 2000. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/f00-
098.

Elliott:2000:LCV

[156] Joel K. Elliott and William C. Leggett. Larval condition and vulnerability to predation: reply to comment by suthers. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57(7):1537–1538, July 2000. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f00-097.

Bridgeman:2000:SAC

[157] Thomas B. Bridgeman, Gretchen Messick, and Henry A. Vanderploeg. Sudden appearance of cysts and ellobiopsid parasites on zooplankton in a Michigan lake: a potential explanation of tumor-like anomalies. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57(8):1539–1544, August 2000. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f00-120.

Hunt:2000:EAP

[158] Gary W. Hunt and Emily H. Stanley. An evaluation of alternative procedures using the bou-rouch method for sampling hyporheic invertebrates. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57(8):1545–1550, August 2000. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f00-081.

Poulsen:2000:AHC

[159] Søren Poulsen, J. Rasmus Nielsen, René Holst, and Karl-Johan Stæhr. An Atlantic herring (Clupea harengus) size selection model for experimental gill nets used in the Sound (ICES Subdivision 23). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57(8):1551–1561, August 2000. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f00-087.

Lachance:2000:SSG

[160] Stephanie Lachance, Pierre Bérubé, and Michel Lemieux. In situ survival and growth of three brook trout (Salvelinus fontinalis) strains subjected to acid conditions of anthropogenic origin at the egg and fingerling stages. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57(8):1562–1573, August 2000. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f00-089.

Brown:2000:NCR

[161] Claude D. Brown, Mark V. Hoyer, Roger W. Bachmann, and Daniel E. Canfield Jr. Nutrient-chlorophyll relationships: an evaluation of empirical nutrient-chlorophyll models using Florida and north-temperate lake data. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57(8):1574–1583, August 2000. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f00-090.

Kotak:2000:RCP

[162] Brian G. Kotak, Angeline K-Y. Lam, Ellie E. Prepas, and Steve E. Hrudey. Role of chemical and physical variables in regulating microcystin-LR concentration in phytoplankton of eutrophic lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57(8):1584–1593, August 2000. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f00-091.

Lantry:2000:PDR

[163] Brian F. Lantry and Donald J. Stewart. Population dynamics of rainbow smelt (Osmerus mordax) in Lakes Ontario and Erie: a modeling analysis of cannibalism effects. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57(8): 1594–1606, August 2000. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f00-092.

Whalen:2000:CTA

[164] Kevin G. Whalen, Donna L. Parrish, Martha E. Mather, and James R. McMenemy. Cross-tributary analysis of parr to smolt recruitment of Atlantic salmon (Salmo salar). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57 (8):1607–1616, August 2000. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f00-093.

Reed-Andersen:2000:PIZ

[165] Tara Reed-Andersen, Stephen R. Carpenter, Dianna K. Padilla, and Richard C. Lathrop. Predicted impact of zebra mussel (Dreissena polymorpha) invasion on water clarity in Lake Mendota. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57(8):1617–1626, August 2000. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f00-094.

Feunteun:2000:EEA

[166] Eric Feunteun, Anthony Acou, Pascal Laffaille, and Antoine Legault. European eel (Anguilla anguilla): prediction of spawner escapement from continental population parameters. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57(8):1627–1635, August 2000. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f00-096.

Geist:2000:SCA

[167] David R. Geist, Julia Jones, Christopher J. Murray, and Dennis D. Dauble. Suitability criteria analyzed at the spatial scale of redd clusters improved estimates of fall Chinook salmon (Oncorhynchus tshawytscha) spawning habitat use in the Hanford reach, Columbia River. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57(8):1636–1646, August 2000. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f00-101.

Geist:2000:HDR

[168] David R. Geist. Hyporheic discharge of river water into fall Chinook salmon (Oncorhynchus tshawytscha) spawning areas in the Hanford reach, Columbia River. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57(8): 1647–1656, August 2000. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f00-102.

Randall:2000:UFP

[169] R. G. Randall and C. K. Minns. Use of fish production per unit biomass ratios for measuring the productive capacity of fish habitats. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57(8):1657–1667, August 2000. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f00-103.

Farrell:2000:PSC

[170] A. P. Farrell, P. Gallaugher, C. Clarke, N. DeLury, H. Kreiberg, W. Parkhouse, and R. Routledge. Physiological status of Coho salmon (Oncorhynchus kisutch) captured in commercial nonretention fisheries. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57(8):1668–1678, August 2000. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f00-116.

Garrison:2000:SDO

[171] Lance P. Garrison. Spatial and dietary overlap in the Georges Bank groundfish community. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57(8): 1679–1691, August 2000. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f00-117.

Laflamme:2000:IMC

[172] J.-S. Laflamme, Y. Couillard, P. G. C. Campbell, and A. Hontela. Interrenal metallothionein and cortisol secretion in relation to Cd, Cu, and Zn exposure in yellow perch, Perca flavescens, from abitibi lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57(8):1692–1700, August 2000. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f00-118.

Stepien:2000:PGP

[173] Carol A. Stepien, Alison K. Dillon, and Amy K. Patterson. Population genetics, phylogeography, and systematics of the thornyhead rockfishes (Sebastolobus) along the deep continental slopes of the North Pacific Ocean. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57(8):1701–1717, August 2000. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/f00-
095.

Finstad:2000:GSS

[174] Anders G. Finstad, Peder A. Jansen, and Arnfinn Langeland. Gillnet selectivity and size and age structure of an alpine Arctic char (Salvelinus alpinus) population. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57(8): 1718–1727, August 2000. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f00-071.

Jeffrey:2000:GVW

[175] Jennifer A. Jeffrey and Christopher T. Taggart. Growth variation and water mass associations of larval silver hake (Merluccius bilinearis) on the Scotian Shelf. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57(8): 1728–1738, August 2000. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f00-088.

Zabel:2000:CCP

[176] Richard W. Zabel and John G. Williams. Comments on “Contrasting patterns of productivity and survival rates for stream-type Chinook salmon (Oncorhynchus tshawytscha) populations of the Snake and Columbia rivers” by Schaller et al. (1999). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57(8):1739–1741, August 2000. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f00-099.

Schaller:2000:RZW

[177] Howard A. Schaller, Charles E. Petrosky, and Olaf P. Langness. Reply to Zabel and Williams’ comments on “Contrasting patterns of productivity and survival rates for stream-type Chinook salmon (Oncorhynchus tshawytscha) populations of the Snake and Columbia rivers” by Schaller et al. (1999). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57(8):1742– 1746, August 2000. CODEN CJFSDX. ISSN 0706-652X (print), 1205-

7533 (electronic). URL https://cdnsciencepub.com/doi/full/10.
1139/f00-100.

Shackell:2000:LFD

[178] N. L. Shackell and K. T. Frank. Larval fish diversity on the Scotian Shelf. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57(9):1747–1760, September 2000. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f00-111.

Bradbury:2000:TDE

[179] Ian R. Bradbury, Paul V. R. Snelgrove, and Sandra Fraser. Transport and development of eggs and larvae of Atlantic cod, Gadus morhua, in relation to spawning time and location in coastal Newfoundland. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57(9):1761–1772, September 2000. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f00-122.

Buzby:2000:MPE

[180] Karen M. Buzby and Sue A. Perry. Modeling the potential effects of climate change on leaf pack processing in central Appalachian streams. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57(9):1773–1783, September 2000. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f00-113.

Helu:2000:ASM

[181] S. Langitoto Helu, David B. Sampson, and Yanshui Yin. Application of statistical model selection criteria to the stock synthesis assessment program. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57(9):1784–1793, September 2000. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f00-137.

vanderMeeren:2000:PHR

[182] Gro I. van der Meeren. Predation on hatchery-reared lobsters released in the wild. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57(9):1794–1803, September 2000. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f00-134.

Giannico:2000:HSJ

[183] Guillermo Roberto Giannico. Habitat selection by juvenile Coho salmon in response to food and woody debris manipulations in suburban and rural stream sections. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57(9): 1804–1813, September 2000. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f00-132.

Takahashi:2000:MLM

[184] H. Takahashi and K. Takata. Multiple lineages of the mitochondrial DNA introgression from Pungitius pungitius (L.) to Pungitius tymensis (Nikolsky). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57(9):1814– 1833, September 2000. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f00-133.

Thompson:2000:MRB

[185] William L. Thompson and Danny C. Lee. Modeling relationships between landscape-level attributes and snorkel counts of Chinook salmon and steelhead parr in Idaho. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57(9): 1834–1842, September 2000. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f00-135.

Hogg:2000:GVA

[186] Ian D. Hogg, Yves de Lafontaine, and John M. Eadie. Genotypic variation among Gammarus fasciatus (Crustacea: Amphipoda) from the Great Lakes–St. Lawrence River: implications for the conservation of widespread freshwater invertebrates. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57(9):1843–1852, September 2000. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f00-136.

Lavoie:2000:ADK

[187] Diane Lavoie, Yvan Simard, and François J. Saucier. Aggregation and dispersion of krill at channel heads and shelf edges: the dynamics in the Saguenay–St. Lawrence Marine Park. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57(9):1853–1869, September 2000. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f00-138.

Kellison:2000:CBS

[188] G. T. Kellison, D. B. Eggleston, and J. S. Burke. Comparative behaviour and survival of hatchery-reared versus wild summer flounder (Paralichthys dentatus). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57(9): 1870–1877, September 2000. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f00-139.

Chen:2000:FLM

[189] D. G. Chen, N. B. Hargreaves, D. M. Ware, and Y. Liu. A fuzzy logic model with genetic algorithm for analyzing fish stock-recruitment relationships. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57(9):1878–1887, September 2000. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f00-141.

Laudon:2000:CPD

[190] Hjalmar Laudon, Olle Westling, and Kevin Bishop. Cause of pH decline in stream water during spring melt runoff in northern Sweden. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57(9):1888–1900, September 2000. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f00-131.

Fahnenstiel:2000:SIM

[191] G. L. Fahnenstiel, R. A. Stone, M. J. McCormick, C. L. Schelske, and S. E. Lohrenz. Spring isothermal mixing in the Great Lakes: evidence of nutrient limitation and nutrient-light interactions in a suboptimal light environment. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57(9):1901– 1910, September 2000. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f00-144.

Scholz:2000:DDA

[192] Nathaniel L. Scholz, Nathan K. Truelove, Barbara L. French, Barry A. Berejikian, Thomas P. Quinn, Edmundo Casillas, and Tracy K. Collier. Diazinon disrupts antipredator and homing behaviors in Chinook salmon (Oncorhynchus tshawytscha). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57(9):1911–1918, September 2000. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f00-147.

Thomson:2000:TSA

[193] Richard E. Thomson and Susan E. Allen. Time series acoustic observations of macrozooplankton diel migration and associated pelagic fish abundance. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57(9):1919–1931, September 2000. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f00-142.

Zhulidov:2000:CAW

[194] Alexander V. Zhulidov, Vladimir V. Khlobystov, Richard D. Robarts, and Dmitry F. Pavlov. Critical analysis of water quality monitoring in the Russian Federation and former Soviet Union. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57(9):1932–1939, September 2000. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f00-140.

Avois:2000:SSI

[195] Carol Avois, Pierre Legendre, Stéphane Masson, and Bernadette Pinel-Alloul. Is the sampling strategy interfering with the study of spatial variability of zooplankton communities? Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57(9):1940–1956, September 2000. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f00-121.

Buckley:2000:CLA

[196] Lawrence J. Buckley, R. Gregory Lough, Myron A. Peck, and Francisco E. Werner. Comment: Larval Atlantic cod and haddock growth models, metabolism, ingestion, and temperature effects. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57(9):1957–1960, September 2000. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f00-146.

Leising:2000:RBA

[197] Andrew W. Leising and Peter J. S. Franks. Reply to Buckley et al.’s “Comment: Larval Atlantic cod and haddock growth models, metabolism, ingestion, and temperature effects”. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57(9):1961–1963, September 2000. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f00-145.

Sastri:2000:RCD

[198] Akash R. Sastri and John C. Roff. Rate of chitobiase degradation as a measure of development rate in planktonic crustacea. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57(10):1965–1968, October 2000. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f00-174.

Shrimpton:2000:CAS

[199] J. Mark Shrimpton, Björn Thrandur Björnsson, and Stephen D. Mc-Cormick. Can Atlantic salmon smolt twice? Endocrine and biochemical changes during smolting. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57(10): 1969–1976, October 2000. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f00-143.

Devito:2000:LCP

[200] K. J. Devito, I. F. Creed, R. L. Rothwell, and E. E. Prepas. Landscape controls on phosphorus loading to boreal lakes: implications for the potential impacts of forest harvesting. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57(10):1977–1984, October 2000. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f00-148.

Withler:2000:IGS

[201] Ruth E. Withler, Khai D. Le, R. John Nelson, Kristina M. Miller, and Terry D. Beacham. Intact genetic structure and high levels of genetic diversity in bottlenecked sockeye salmon (Oncorhynchus nerka) populations of the Fraser River, British Columbia, Canada. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57(10):1985–1998, October 2000. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f00-149.

Bodini:2000:RTI

[202] Antonio Bodini. Reconstructing trophic interactions as a tool for understanding and managing ecosystems: application to a shallow eutrophic lake. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57(10):1999–2009, October 2000. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/f00-
153.

Diamond:2000:PES

[203] Sandra L. Diamond, Lindsay G. Cowell, and Larry B. Crowder. Population effects of shrimp trawl bycatch on Atlantic croaker. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57(10):2010–2021, October 2000. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f00-154.

Vestergaard:2000:AMR

[204] Ole Vestergaard and Kaj Sand-Jensen. Aquatic macrophyte richness in Danish lakes in relation to alkalinity, transparency, and lake area. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57(10):2022–2031, October 2000. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f00-156.

Barry:2000:IAM

[205] Karen L. Barry, Jeffrey A. Grout, Colin D. Levings, Bruce H. Nidle, and G. Elizabeth Piercey. Impacts of acid mine drainage on juvenile salmonids in an estuary near Britannia Beach in Howe Sound, British Columbia. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57(10):2032–2043, October 2000. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f00-157.

Plaganyi:2000:AFS

[206] Éva E. Plagányi, Larry Hutchings, and John G. Field. Anchovy foraging: simulating spatial and temporal match/mismatches with zooplankton. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57(10):2044–2053, October 2000. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f00-158.

Brauner:2000:EFH

[207] C. J. Brauner, M. Seidelin, S. S. Madsen, and F. B. Jensen. Effects of freshwater hyperoxia and hypercapnia and their influences on subsequent seawater transfer in Atlantic salmon (Salmo salar) smolts. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57(10):2054–2064, October 2000. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f00-161.

Guay:2000:DVN

[208] J. C. Guay, D. Boisclair, D. Rioux, M. Leclerc, M. Lapointe, and P. Legendre. Development and validation of numerical habitat models for juveniles of Atlantic salmon (Salmo salar). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57(10):2065–2075, October 2000. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f00-162.

Thedinga:2000:ESC

[209] J. F. Thedinga, A. C. Wertheimer, R. A. Heintz, J. M. Maselko, and S. D. Rice. Effects of stock, coded-wire tagging, and transplant on straying of pink salmon (Oncorhynchus gorbuscha) in southeastern Alaska. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57(10):2076–2085, October 2000. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f00-163.

Aarestrup:2000:RBG

[210] Kim Aarestrup, Christian Nielsen, and Steffen S. Madsen. Relationship between gill Na^{+}, K^{+}-ATPase activity and downstream movement in domesticated and first-generation offspring of wild anadromous brown trout (Salmo trutta). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57(10): 2086–2095, October 2000. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f00-164.

Burger:2000:GCT

[211] Carl V. Burger, Kim T. Scribner, William J. Spearman, Charles O. Swanton, and Donald E. Campton. Genetic contribution of three introduced life history forms of sockeye salmon to colonization of Frazer Lake, Alaska. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57(10):2096–2111, October 2000. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/f00-
200.

Kern:2000:GBK

[212] John W. Kern and Kenneth O. Coyle. Global block kriging to estimate biomass from acoustic surveys for zooplankton in the western Aleutian Islands. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57(10):2112–2121, October 2000. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/f00-

Wells:2000:ISC

[213] Brian K. Wells, Gretchen E. Bath, Simon R. Thorrold, and Cynthia M. Jones. Incorporation of strontium, cadmium, and barium in juvenile spot (Leiostomus xanthurus) scales reflects water chemistry. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57(10):2122–2129, October 2000. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f00-178.

Hansen:2000:GMS

[214] Michael M. Hansen, Einar E. Nielsen, Daniel E. Ruzzante, Carmen Bouza, and Karen-Lise D. Mensberg. Genetic monitoring of supportive breeding in brown trout (Salmo trutta L.), using microsatellite DNA markers. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57(10):2130–2139, October 2000. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/f00-
187.

McIntosh:2000:HSR

[215] Angus R. McIntosh. Habitat- and size-related variations in exotic trout impacts on native galaxiid fishes in New Zealand streams. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57(10):2140–2151, October 2000. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f00-188.

Zimmerman:2000:PSS

[216] Christian E. Zimmerman and Gordon H. Reeves. Population structure of sympatric anadromous and nonanadromous Oncorhynchus mykiss: evidence from spawning surveys and otolith microchemistry. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57(10):2152–2162, October 2000. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f00-192.

Olding:2000:PCC

[217] Daniel D. Olding, Johan A. Hellebust, and Marianne S. V. Douglas. Phytoplankton community composition in relation to water quality and water-body morphometry in urban lakes, reservoirs, and ponds. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57(10):2163–2174, October 2000. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f00-176.

Mayfield:2000:IAR

[218] Stephen Mayfield and George M. Branch. Interrelations among rock lobsters, sea urchins, and juvenile abalone: implications for community management. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57(11):2175– 2185, November 2000. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f00-198.

Bryden:2000:HFA

[219] Colleen A. Bryden and Daniel D. Heath. Heritability of fluctuating asymmetry for multiple traits in Chinook salmon (Oncorhynchus tshawytscha). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57(11):2186– 2192, November 2000. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f00-190.

Schick:2000:SCB

[220] R. S. Schick and D. L. Urban. Spatial components of bowhead whale (Balaena mysticetus) distribution in the Alaskan Beaufort Sea. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57(11):2193–2200, November 2000. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f00-196.

Sanz:2000:DBT

[221] Núria Sanz, José-Luis García-Marín, and Carles Pla. Divergence of brown trout (Salmo trutta) within glacial refugia. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57(11):2201–2210, November 2000. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f00-199.

Heyes:2000:MMP

[222] A. Heyes, T. R. Moore, J. W. M. Rudd, and J. J. Dugoua. Methyl mercury in pristine and impounded boreal peatlands, Experimental Lakes Area, Ontario. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57(11):2211– 2222, November 2000. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f00-197.

Purchase:2000:IDG

[223] Craig F. Purchase and Joseph A. Brown. Interpopulation differences in growth rates and food conversion efficiencies of young Grand Banks and Gulf of Maine Atlantic cod (Gadus morhua). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57(11):2223–2229, November 2000. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f00-204.

Shelton:2000:ICN

[224] P. A. Shelton and G. R. Lilly. Interpreting the collapse of the northern cod stock from survey and catch data. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57(11):2230–2239, November 2000. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f00-191.

Kuns:2000:ZPL

[225] Michelle M. Kuns and W. Gary Sprules. Zooplankton production in Lake Ontario: a multistrata approach. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57(11):2240–2247, November 2000. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f00-189.

Healey:2000:PMM

[226] M. C. Healey, M. A. Henderson, and Ingrid Burgetz. Precocial maturation of male sockeye salmon in the Fraser River, British Columbia, and its relationship to growth and year-class strength. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57(11):2248–2257, November 2000. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f00-202.

Tanasichuk:2000:ASN

[227] R. W. Tanasichuk. Age-specific natural mortality rates of adult Pacific herring (Clupea pallasi) from southern British Columbia. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57(11):2258–2266, November 2000. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f00-205.

Goedkoop:2000:FAP

[228] Willem Goedkoop, Lars Sonesten, Gunnel Ahlgren, and Merike Boberg. Fatty acids in profundal benthic invertebrates and their major food resources in Lake Erken, Sweden: seasonal variation and trophic indications. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57(11):2267–2279, November 2000. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f00-201.

Kennedy:2000:UNS

[229] Brian P. Kennedy, Joel D. Blum, Carol L. Folt, and Keith H. Nislow. Using natural strontium isotopic signatures as fish markers: methodology and application. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57(11): 2280–2292, November 2000. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f00-206.

Chen:2000:IOM

[230] Y. Chen, P. A. Breen, and N. L. Andrew. Impacts of outliers and misspecification of priors on Bayesian fisheries-stock assessment. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57(11):2293–2305, November 2000. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f00-208.

Coyle:2000:AEZ

[231] Kenneth O. Coyle. Acoustic estimates of zooplankton biomass and distribution: application of canonical correlation to scaling of multifrequency acoustic data. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57(11):2306– 2318, November 2000. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f00-155.

Milner:2000:CDS

[232] Alexander M. Milner, E. Eric Knudsen, Chad Soiseth, Anne L. Robertson, Don Schell, Ian T. Phillips, and Katrina Magnusson. Colonization and development of stream communities across a 200-year gradient in Glacier Bay National Park, Alaska, U.S.A. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57(11):2319–2335, November 2000. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f00-212.

Fischer:2000:ETM

[233] Philipp Fischer. An experimental test of metabolic and behavioural responses of benthic fish species to different types of substrate. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57(11):2336–2344, November 2000. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f00-211.

LeBreton:2000:IGV

[234] Gregory T. O. LeBreton and F. William H. Beamish. Interannual growth variation in fish and tree rings. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57(11):2345–2356, November 2000. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f00-207.

Myers:2000:FMS

[235] Ransom A. Myers, Susanna D. Fuller, and Daniel G. Kehler. A fisheries management strategy robust to ignorance: rotational harvest in the presence of indirect fishing mortality. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57(12):2357–2362, December 2000. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f00-250.

Adlerstein:2000:DVS

[236] Sara A. Adlerstein and Henny C. Welleman. Diel variation of stomach contents of North Sea cod (Gadus morhua) during a 24-h fishing survey: an analysis using generalized additive models. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57(12):2363–2367, December 2000. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f00-249.

Hedrick:2000:EPS

[237] Philip W. Hedrick, Vanessa K. Rashbrook, and Dennis Hedgecock. Effective population size of winter-run Chinook salmon based on microsatellite analysis of returning spawners. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57(12):2368–2373, December 2000. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f00-213.

Lorenzen:2000:ANM

[238] Kai Lorenzen. Allometry of natural mortality as a basis for assessing optimal release size in fish-stocking programmes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57(12):2374–2381, December 2000. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f00-215.

Keller:2000:IET

[239] Aimee A. Keller and Grace Klein-MacPhee. Impact of elevated temperature on the growth, survival, and trophic dynamics of winter flounder larvae: a mesocosm study. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57(12): 2382–2392, December 2000. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f00-217.

Swain:2000:DDV

[240] D. P. Swain and K. T. Frank. Depth-dependent vertebral number of Atlantic cod (Gadus morhua) on the Magdalen Shallows and Scotian Shelf: stock mixing or microgeographic trends within populations? Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57(12):2393–2401, December 2000. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f00-218.

Cardinale:2000:ISS

[241] M. Cardinale and F. Arrhenius. The influence of stock structure and environmental conditions on the recruitment process of Baltic cod estimated using a generalized additive model. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57(12):2402–2409, December 2000. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f00-221.

McConnaughey:2000:ABF

[242] Robert A. McConnaughey and Keith R. Smith. Associations between flatfish abundance and surficial sediments in the eastern Bering Sea. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57(12):2410–2419, December 2000. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f00-219.

Fu:2000:ENM

[243] Caihong Fu and Terrance J. Quinn II. Estimability of natural mortality and other population parameters in a length-based model: Pandalus borealis in kachemak Bay, Alaska. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57(12):2420–2432, December 2000. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f00-220.

Marshall:2000:RSR

[244] C. Tara Marshall, Nathalia A. Yaragina, Bjørn Ådlandsvik, and Andrey V. Dolgov. Reconstructing the stock-recruit relationship for Northeast Arctic cod using a bioenergetic index of reproductive potential. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57(12):2433–2442, December 2000. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f00-222.

Lee:2000:STS

[245] Yong-Woo Lee and David B. Sampson. Spatial and temporal stability of commercial groundfish assemblages off Oregon and Washington as inferred from Oregon trawl logbooks. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57(12):2443–2454, December 2000. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f00-223.

Garrison:2000:PRL

[246] Lance P. Garrison, William Michaels, Jason S. Link, and Michael J. Fogarty. Predation risk on larval gadids by pelagic fish in the Georges Bank ecosystem. I. Spatial overlap associated with hydrographic features. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57(12):2455–2469, December 2000. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f00-229.

Hinch:2000:OSS

[247] Scott G. Hinch and Peter S. Rand. Optimal swimming speeds and forward-assisted propulsion: energy-conserving behaviours of uprivermigrating adult salmon. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57(12): 2470–2478, December 2000. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f00-238.

Agnew:2000:PRS

[248] David J. Agnew, Simeon Hill, and John R. Beddington. Predicting the recruitment strength of an annual squid stock: Loligo gahi around the Falkland Islands. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57(12):2479– 2487, December 2000. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f00-240.

Shore:2000:DPW

[249] Jennifer A. Shore, Charles G. Hannah, and John W. Loder. Drift pathways on the western Scotian Shelf and its environs. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57(12):2488–2505, December 2000. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f00-214.

Hannah:2000:DRD

[250] Charles G. Hannah, Jennifer A. Shore, and John W. Loder. The driftretention dichotomy on Browns Bank: a model study of interannual variability. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57(12):2506–2518, December 2000. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f00-234.

Brickman:2000:MDM

[251] David Brickman and Kenneth T. Frank. Modelling the dispersal and mortality of Browns Bank egg and larval haddock (Melanogrammus aeglefinus). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57(12):2519–2535, December 2000. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f00-227.

Petrie:2000:AIV

[252] Brian Petrie and Philip Yeats. Annual and interannual variability of nutrients and their estimated fluxes in the Scotian Shelf–Gulf of Maine region. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57(12):2536–2546, December 2000. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f00-235.

Gingras:2000:CBC

[253] Jérôme Gingras and Daniel Boisclair. Comparison between consumption rates of yellow perch (Perca flavescens) estimated with a digestive tract model and with a radioisotope approach. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57(12):2547–2557, December 2000. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f00-230.

Anonymous:2000:AII

[254] Anonymous. Author index / index des auteurs: Volume 57, 2000. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57(12), December 2000. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfasauthor00.

Anonymous:2000:CSV

[255] Anonymous. Contents / sommaire: Volume 57, 2000. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57(12), December 2000. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfascontents00.

Anonymous:2001:EE

[256] Anonymous. Editorial / Éditorial. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 58(1), January 2001. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f00-236.

Boisclair:2001:FHM

[257] Daniel Boisclair. Fish habitat modeling: from conceptual framework to functional tools. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 58(1):1–9, January 2001. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/f00-

Schnute:2001:UAF

[258] Jon T. Schnute and Laura J. Richards. Use and abuse of fishery models. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 58(1):10–17, January 2001. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f00-150.

Schindler:2001:CEC

[259] D. W. Schindler. The cumulative effects of climate warming and other human stresses on Canadian freshwaters in the new millennium. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 58(1):18–29, January 2001. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f00-179.

Campana:2001:OIE

[260] Steven E. Campana and Simon R. Thorrold. Otoliths, increments, and elements: keys to a comprehensive understanding of fish populations? Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 58(1):30–38, January 2001. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f00-177.

Walters:2001:CDE

[261] Carl Walters and James F. Kitchell. Cultivation/depensation effects on juvenile survival and recruitment: implications for the theory of fishing. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 58(1):39–50, January 2001. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f00-160.

Pauly:2001:FCA

[262] Daniel Pauly, Ma. Lourdes Palomares, Rainer Froese, Pascualita Sa-a, Michael Vakily, David Preikshot, and Scott Wallace. Fishing down Canadian aquatic food webs. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 58 (1):51–62, January 2001. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f00-193.

Pace:2001:PAS

[263] Michael L. Pace. Prediction and the aquatic sciences. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 58(1):63–72, January 2001. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f00-151.

Carscadden:2001:ECE

[264] J. E. Carscadden, K. T. Frank, and W. C. Leggett. Ecosystem changes and the effects on capelin (Mallotus villosus), a major forage species. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 58(1):73–85, January 2001. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f00-185.

Metaxas:2001:BFP

[265] Anna Metaxas. Behaviour in flow: perspectives on the distribution and dispersion of meroplanktonic larvae in the water column. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 58(1):86–98, January 2001. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f00-159.

Hilborn:2001:PAR

[266] Ray Hilborn, Jean-Jacques Maguire, Ana M. Parma, and Andrew A. Rosenberg. The precautionary approach and risk management: can they increase the probability of successes in fishery management? Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 58(1):99–107, January 2001. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f00-225.

Hutchings:2001:CBM

[267] Jeffrey A. Hutchings. Conservation biology of marine fishes: perceptions and caveats regarding assignment of extinction risk. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 58(1):108–121, January 2001. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f00-228.

Arts:2001:EFA

[268] Michael T. Arts, Robert G. Ackman, and Bruce J. Holub. “Essential fatty acids” in aquatic ecosystems: a crucial link between diet and human health and evolution. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 58 (1):122–137, January 2001. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f00-224.

Jamieson:2001:MPA

[269] Glen S. Jamieson and Colin O. Levings. Marine protected areas in Canada — implications for both conservation and fisheries management. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 58(1):138–156, January 2001. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f00-233.

Jackson:2001:WCW

[270] Donald A. Jackson, Pedro R. Peres-Neto, and Julian D. Olden. What controls who is where in freshwater fish communities — the roles of biotic, abiotic, and spatial factors. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 58 (1):157–170, January 2001. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f00-239.

Drinkwater:2001:EHS

[271] K. F. Drinkwater and G. C. Harding. Effects of the Hudson Strait outflow on the biology of the Labrador Shelf. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 58(1):171–184, January 2001. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f00-210.

Jackson:2001:VIC

[272] Togwell A. Jackson. Variations in the isotope composition of mercury in a freshwater sediment sequence and food web. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 58(1):185–196, January 2001. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f00-186.

Volpe:2001:CAJ

[273] John P. Volpe, Bradley R. Anholt, and Barry W. Glickman. Competition among juvenile Atlantic salmon (Salmo salar) and steelhead (Oncorhynchus mykiss): relevance to invasion potential in British Columbia. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 58(1):197–207, January 2001. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f00-209.

Gargett:2001:TME

[274] A. E. Gargett, M. Li, and R. Brown. Testing mechanistic explanations of observed correlations between environmental factors and marine fisheries. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 58(1):208–219, January 2001. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f00-216.

Bochdansky:2001:WRC

[275] Alexander B. Bochdansky and William C. Leggett. Winberg revisited: convergence of routine metabolism in larval and juvenile fish. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 58(1):220–230, January 2001. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f00-226.

Anonymous:2001:IA

[276] Anonymous. Instructions to authors. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 58(1), January 2001. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas_instruct01_e.

Anonymous:2001:RA

[277] Anonymous. Recommandations aux auteurs. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 58(1), January 2001. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/cjfas_instruct01_f.

Nicholls:2001:AFC

[278] Kenneth H. Nicholls and Claudiu Tudorancea. Application of fuzzy cluster analysis to Lake Simcoe crustacean zooplankton community structure. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 58(2):231–240, February 2001. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f00-231.

Pilditch:2001:SSS

[279] C. A. Pilditch, J. Grant, and K. R. Bryan. Seston supply to sea scallops (Placopecten magellanicus) in suspended culture. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 58(2):241–253, February 2001. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f00-242.

McGarvey:2001:DEL

[280] Richard McGarvey and Michael Pennington. Designing and evaluating length-frequency surveys for trap fisheries with application to the southern rock lobster. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 58(2): 254–261, February 2001. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f00-237.

Fabrizio:2001:DIG

[281] Mary C. Fabrizio, Robert M. Dorazio, and Stephen T. Schram. Dynamics of individual growth in a recovering population of lake trout (Salvelinus namaycush). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 58(2): 262–272, February 2001. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f00-232.

Perissinotto:2001:GDE

[282] Renzo Perissinotto, Patrick Mayzaud, Jean-Philippe Labat, and Suzanne Razouls. Grazing dynamics of Euphausia spinifera in the region of the subtropical convergence and the agulhas front. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 58(2):273–281, February 2001. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f00-241.

Roni:2001:DSJ

[283] Philip Roni and Thomas P. Quinn. Density and size of juvenile salmonids in response to placement of large woody debris in western Oregon and Washington streams. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 58(2): 282–292, February 2001. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f00-246.

Kamunde:2001:CMG

[284] Collins N. Kamunde, Martin Grosell, John N. A. Lott, and Chris M. Wood. Copper metabolism and gut morphology in rainbow trout (Oncorhynchus mykiss) during chronic sublethal dietary copper exposure. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 58(2):293–305, February 2001. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f00-248.

Brodziak:2001:DSS

[285] Jon K. T. Brodziak, William J. Overholtz, and Paul J. Rago. Does spawning stock affect recruitment of New England groundfish? Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 58(2):306–318, February 2001. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f00-243.

Tinus:2001:RSR

[286] Craig A. Tinus and Gordon H. Reeves. Redside shiner (Richardsonius balteatus) shoals provide a behavioral competitive refuge for subordinate juvenile steelhead trout (Oncorhynchus mykiss). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 58(2):319–324, February 2001. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f00-244.

Sommer:2001:FRJ

[287] T. R. Sommer, M. L. Nobriga, W. C. Harrell, W. Batham, and W. J. Kimmerer. Floodplain rearing of juvenile Chinook salmon: evidence of enhanced growth and survival. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 58(2):325–333, February 2001. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f00-245.

Briand:2001:CPU

[288] Geneviève Briand, Scott C. Matulich, and Ron C. Mittelhammer. A catch per unit effort–soak time model for the Bristol Bay red king crab fishery, 1991–1997. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 58(2): 334–341, February 2001. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f00-258.

MacRae:2001:ISB

[289] Pamela S. D. MacRae and Donald A. Jackson. The influence of smallmouth bass (Micropterus dolomieu) predation and habitat complexity on the structure of littoral zone fish assemblages. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 58(2):342–351, February 2001. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f00-247.

Klumb:2001:VTB

[290] Robert A. Klumb, Michael A. Bozek, and Richard V. Frie. Validation of three back-calculation models by using multiple oxytetracycline marks formed in the otoliths and scales of bluegill × green sunfish hybrids. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 58(2):352–364, February 2001. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f00-256.

Bradford:2001:HSS

[291] Michael J. Bradford and Paul S. Higgins. Habitat-, season-, and size-specific variation in diel activity patterns of juvenile Chinook salmon (Oncorhynchus tshawytscha) and steelhead trout (Oncorhynchus mykiss). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 58(2):365–374, February 2001. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f00-253.

Shea:2001:DPT

[292] Katriona Shea and Marc Mangel. Detection of population trends in threatened Coho salmon (Oncorhynchus kisutch). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 58(2):375–385, February 2001. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f00-254.

Sweka:2001:ETP

[293] John A. Sweka and Kyle J. Hartman. Effects of turbidity on prey consumption and growth in brook trout and implications for bioenergetics modeling. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 58(2):386–393, February 2001. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f00-260.

Trudel:2001:WDF

[294] Marc Trudel, Alain Tremblay, Roger Schetagne, and Joseph B. Rasmussen. Why are dwarf fish so small? An energetic analysis of polymorphism in lake whitefish (Coregonus clupeaformis). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 58(2):394–405, February 2001. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f00-252.

Taylor:2001:NAG

[295] Eric B. Taylor, Zoë A. Redenbach, Allan B. Costello, Susan J. Pollard, and Charlie J. Pacas. Nested analysis of genetic diversity in northwestern North American char, Dolly Varden (Salvelinus malma) and bull trout (Salvelinus confluentus). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 58 (2):406–420, February 2001. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f00-262.

Prepas:2001:FHI

[296] E. E. Prepas, B. Pinel-Alloul, D. Planas, G. Méthot, S. Paquet, and S. Reedyk. Forest harvest impacts on water quality and aquatic biota on the boreal plain: introduction to the TROLS lake program. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 58(2):421–436, February 2001. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f00-259.

Taylor:2001:SSI

[297] Brad W. Taylor, Angus R. McIntosh, and Barbara L. Peckarsky. Sampling stream invertebrates using electroshocking techniques: implications for basic and applied research. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 58(3):437–445, March 2001. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f00-255.

Guensch:2001:EFS

[298] G. R. Guensch, T. B. Hardy, and R. C. Addley. Examining feeding strategies and position choice of drift-feeding salmonids using an individualbased, mechanistic foraging model. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 58(3):446–457, March 2001. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f00-257.

Maury:2001:HIN

[299] Olivier Maury, Didier Gascuel, Francis Marsac, Alain Fonteneau, and Anne-Laure De Rosa. Hierarchical interpretation of nonlinear relationships linking yellowfin tuna (Thunnus albacares) distribution to the environment in the Atlantic Ocean. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 58(3):458–469, March 2001. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f00-261.

March:2001:LSA

[300] James G. March, Jonathan P. Benstead, Catherine M. Pringle, and Mark W. Ruebel. Linking shrimp assemblages with rates of detrital processing along an elevational gradient in a tropical stream. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 58(3):470–478, March 2001. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f00-263.

Laird:2001:RPAa

[301] Kathleen Laird, Brian Cumming, and Rick Nordin. A regional paleolimnological assessment of the impact of clear-cutting on lakes from the west coast of Vancouver Island, British Columbia. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 58(3):479–491, March 2001. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f00-265.

Laird:2001:RPAb

[302] Kathleen Laird and Brian Cumming. A regional paleolimnological assessment of the impact of clear-cutting on lakes from the central interior of British Columbia. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 58 (3):492–505, March 2001. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f00-264.

Regnell:2001:EAS

[303] Olof Regnell, Tommy Hammar, Anders Helgée, and Bo Troedsson. Effects of anoxia and sulfide on concentrations of total and methyl mercury in sediment and water in two Hg-polluted lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 58(3):506–517, March 2001. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f01-001.

Lozano:2001:RDB

[304] Stephen J. Lozano, Jill V. Scharold, and Thomas F. Nalepa. Recent declines in benthic macroinvertebrate densities in Lake Ontario. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 58(3):518–529, March 2001. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f01-002.

Tsurumi:2001:CHV

[305] Maia Tsurumi and Verena Tunnicliffe. Characteristics of a hydrothermal vent assemblage on a volcanically active segment of Juan de Fuca Ridge, northeast Pacific. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 58 (3):530–542, March 2001. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f01-005.

Dixit:2001:DID

[306] Sushil S. Dixit, W. (Bill) Keller, Aruna S. Dixit, and John P. Smol. Diatom-inferred dissolved organic carbon reconstructions provide assessments of past UV-B penetration in Canadian Shield lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 58(3):543–550, March 2001. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f01-013.

Hinrichsen:2001:IID

[307] Richard A. Hinrichsen. The importance of influence diagnostics: examples from Snake River Chinook salmon spawner-recruit models. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 58(3):551–559, March 2001. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f01-014.

Cadigan:2001:CGL

[308] Noel G. Cadigan and Ransom A. Myers. A comparison of gamma and lognormal maximum likelihood estimators in a sequential population analysis. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 58(3):560–567, March 2001. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f01-003.

Armsworth:2001:EFP

[309] Paul R. Armsworth. Effects of fishing on a protogynous hermaphrodite. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 58(3):568–578, March 2001. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f01-015.

Hilborn:2001:CBT

[310] Ray Hilborn. Calculation of biomass trend, exploitation rate, and surplus production from survey and catch data. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 58(3):579–584, March 2001. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f01-018.

Rosenfeld:2001:FCH

[311] Jordan S. Rosenfeld and Shelly Boss. Fitness consequences of habitat use for juvenile cutthroat trout: energetic costs and benefits in pools and riffles. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 58(3):585–593, March 2001. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f01-019.

Allen:2001:QCC

[312] Micheal S. Allen and Leandro E. Miranda. Quasi-cycles in crappie populations are forced by interactions among population characteristics and environment. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 58(3):594–601, March 2001. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f01-020.

Soucek:2001:SSA

[313] David J. Soucek, Travis S. Schmidt, and Donald S. Cherry. In situ studies with Asian clams (Corbicula fluminea) detect acid mine drainage and nutrient inputs in low-order streams. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 58(3):602–608, March 2001. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f01-023.

Finlay:2001:STV

[314] Kerri P. Finlay, Hélène Cyr, and Brian J. Shuter. Spatial and temporal variability in water temperatures in the littoral zone of a multibasin lake. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 58(3):609–619, March 2001. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f01-017.

Janz:2001:ROS

[315] David M. Janz, Mark E. McMaster, Lynn P. Weber, Kelly R. Munkittrick, and Glen Van Der Kraak. Recovery of ovary size, follicle cell apoptosis, and HSP70 expression in fish exposed to bleached pulp mill effluent. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 58(3):620–625, March 2001. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f01-036.

Anderson:2001:PTU

[316] Marti J. Anderson. Permutation tests for univariate or multivariate analysis of variance and regression. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 58(3):626–639, March 2001. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f01-004.

Sheehy:2001:IPR

[317] M. R. J. Sheehy. Implications of protracted recruitment for perception of the spawner–recruit relationship. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 58(4):641–644, April 2001. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f01-047.

Mackas:2001:GCS

[318] D. L. Mackas and B. deYoung. GLOBEC Canada 1996–2000: a sampler. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 58(4):645–646, April 2001. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f01-032.

Plourde:2001:LCC

[319] Stéphane Plourde, Pierre Joly, Jeffrey A. Runge, Bruno Zakardjian, and Julian J. Dodson. Life cycle of Calanus finmarchicus in the lower St.

Lawrence Estuary: the imprint of circulation and late timing of the spring phytoplankton bloom. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 58 (4):647–658, April 2001. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f01-006.

McLaren:2001:LCS

[320] Ian A. McLaren, Erica Head, and D. D. Sameoto. Life cycles and seasonal distributions of Calanus finmarchicus on the central Scotian Shelf. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 58(4):659–670, April 2001. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f01-007.

Allen:2001:PBP

[321] S. E. Allen, C. Vindeirinho, R. E. Thomson, M. G. G. Foreman, and D. L. Mackas. Physical and biological processes over a submarine canyon during an upwelling event. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 58 (4):671–684, April 2001. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f01-008.

Mackas:2001:CZC

[322] D. L. Mackas, Richard E. Thomson, and Moira Galbraith. Changes in the zooplankton community of the British Columbia continental margin, 1985–1999, and their covariation with oceanographic conditions. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 58(4):685–702, April 2001. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f01-009.

Haigh:2001:SPE

[323] S. P. Haigh, K. L. Denman, and W. W. Hsieh. Simulation of the planktonic ecosystem response to pre- and post-1976 forcing in an isopycnic model of the North Pacific. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 58 (4):703–722, April 2001. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f01-010.

Lu:2001:TCM

[324] Youyu Lu, Keith R. Thompson, and Daniel G. Wright. Tidal currents and mixing in the Gulf of St. Lawrence: an application of the incremental approach to data assimilation. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 58 (4):723–735, April 2001. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f01-011.

Smith:2001:HTM

[325] C. L. Smith, A. E. Hill, M. G. G. Foreman, and M. A. Peña. Horizontal transport of marine organisms resulting from interactions between diel vertical migration and tidal currents off the west coast of Vancouver Island. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 58(4):736–748, April 2001. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f01-012.

Sameoto:2001:DCP

[326] D. Sameoto. Decadal changes in phytoplankton color index and selected calanoid copepods in continuous plankton recorder data from the Scotian Shelf. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 58(4):749–761, April 2001. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f01-033.

Morgan:2001:IER

[327] Ian J. Morgan and Neil B. Metcalfe. The influence of energetic requirements on the preferred temperature of overwintering juvenile Atlantic salmon (Salmo salar). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 58 (4):762–768, April 2001. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f01-016.

Hinrichsen:2001:HVS

[328] Richard A. Hinrichsen. High variability in spawner–recruit data hampers learning. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 58(4):769–776, April 2001. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f01-021.

Palka:2001:ARM

[329] D. L. Palka and P. S. Hammond. Accounting for responsive movement in line transect estimates of abundance. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 58(4):777–787, April 2001. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f01-024.

Johnsson:2001:BHR

[330] Jörgen I. Johnsson, Johan Höjesjö, and Ian A. Fleming. Behavioural and heart rate responses to predation risk in wild and domesticated Atlantic salmon. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 58(4):788–794, April 2001. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f01-025.

Maunder:2001:GFI

[331] Mark N. Maunder. A general framework for integrating the standardization of catch per unit of effort into stock assessment models. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 58(4):795–803, April 2001. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f01-029.

Berejikian:2001:MCB

[332] B. A. Berejikian, E. P. Tezak, L. Park, E. LaHood, S. L. Schroder, and E. Beall. Male competition and breeding success in captively reared and wild Coho salmon (Oncorhynchus kisutch). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 58(4):804–810, April 2001. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f01-039.

Bradbury:2001:CLT

[333] Ian R. Bradbury and Paul V. R. Snelgrove. Contrasting larval transport in demersal fish and benthic invertebrates: the roles of behaviour and advective processes in determining spatial pattern. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 58(4):811–823, April 2001. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f01-031.

Rice:2001:TSS

[334] S. P. Rice, M. T. Greenwood, and C. B. Joyce. Tributaries, sediment sources, and the longitudinal organisation of macroinvertebrate fauna along river systems. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 58 (4):824–840, April 2001. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f01-022.

Hedrick:2001:RCC

[335] Philip W. Hedrick. Rapid communication / communication rapide: Invasion of transgenes from salmon or other genetically modified organisms into natural populations. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 58 (5):841–844, May 2001. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f01-064.

Hatfield:2001:LTG

[336] Emma M. C. Hatfield, Roger T. Hanlon, John W. Forsythe, and Eric P. M. Grist. Laboratory testing of a growth hypothesis for juvenile squid Loligo pealeii (Cephalopoda: Loliginidae). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 58(5):845–857, May 2001. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f01-030.

Boegman:2001:ATD

[337] L. Boegman, M. R. Loewen, P. F. Hamblin, and D. A. Culver. Application of a two-dimensional hydrodynamic reservoir model to Lake Erie. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 58(5):858–869, May 2001. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f01-035.

Lake:2001:SNI

[338] James L. Lake, Richard A. McKinney, Frank A. Osterman, Richard J. Pruell, John Kiddon, Stephan A. Ryba, and Alan D. Libby. Stable nitrogen isotopes as indicators of anthropogenic activities in small freshwater systems. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 58(5):870–878, May 2001. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f01-038.

Kenney:2001:CPS

[339] William F. Kenney, Claire L. Schelske, and Andrew D. Chapman. Changes in polyphosphate sedimentation: a response to excessive phosphorus enrichment in a hypereutrophic lake. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 58(5):879–887, May 2001. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f01-040.

Bowles:2001:BBM

[340] Karl C. Bowles, Simon C. Apte, William A. Maher, Matthew Kawei, and Ross Smith. Bioaccumulation and biomagnification of mercury in Lake Murray, Papua New Guinea. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 58 (5):888–897, May 2001. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f01-042.

Allen:2001:DDC

[341] Yvonne C. Allen and Charles W. Ramcharan. Dreissena distribution in commercial waterways of the U.S.: using failed invasions to identify limiting factors. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 58(5):898–907, May 2001. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f01-043.

Tsou:2001:EPM

[342] Tien-Shui Tsou and Jeremy S. Collie. Estimating predation mortality in the Georges Bank fish community. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 58(5):908–922, May 2001. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f01-044.

MacAvoy:2001:ITA

[343] Stephen E. MacAvoy, Stephen A. Macko, and Greg C. Garman. Isotopic turnover in aquatic predators: quantifying the exploitation of migratory prey. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 58(5):923–932, May 2001. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f01-045.

Dettmers:2001:PAF

[344] John M. Dettmers, Steve Gutreuter, David H. Wahl, and Daniel A. Soluk. Patterns in abundance of fishes in main channels of the upper Mississippi River system. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 58 (5):933–942, May 2001. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f01-046.

Beer:2001:ESD

[345] W. Nicholas Beer and James J. Anderson. Effect of spawning day and temperature on salmon emergence: interpretations of a growth model for Methow River Chinook. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 58 (5):943–949, May 2001. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f01-049.

Borgmann:2001:ICS

[346] U. Borgmann, W. P. Norwood, T. B. Reynoldson, and F. Rosa. Identifying cause in sediment assessments: bioavailability and the sediment quality triad. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 58(5):950–960, May 2001. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/f01-
048.

Cutts:2001:SPB

[347] C. J. Cutts, C. E. Adams, and A. Campbell. Stability of physiological and behavioural determinants of performance in Arctic char (Salvelinus alpinus). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 58(5):961–968, May 2001. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f01-050.

Frisk:2001:EAB

[348] Michael G. Frisk, Thomas J. Miller, and Michael J. Fogarty. Estimation and analysis of biological parameters in elasmobranch fishes: a comparative life history study. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 58 (5):969–981, May 2001. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f01-051.

Lage:2001:MEH

[349] Christopher Lage, Maureen Purcell, Michael Fogarty, and Irv Kornfield. Microsatellite evaluation of haddock (Melanogrammus aeglefinus) stocks in the northwest Atlantic Ocean. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 58(5):982–990, May 2001. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f01-052.

Smaal:2001:IBS

[350] Aad Smaal, Marnix van Stralen, and Egbertha Schuiling. The interaction between shellfish culture and ecosystem processes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 58(5):991–1002, May 2001. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f01-026.

Grant:2001:NMF

[351] Jon Grant and Cedric Bacher. A numerical model of flow modification induced by suspended aquaculture in a Chinese bay. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 58(5):1003–1011, May 2001. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f01-027.

Gangnery:2001:API

[352] A. Gangnery, C. Bacher, and D. Buestel. Assessing the production and the impact of cultivated oysters in the Thau lagoon (Mediterranee, France) with a population dynamics model. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 58(5):1012–1020, May 2001. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f01-028.

Stenton-Dozey:2001:IMM

[353] Jeanie Stenton-Dozey, Trevor Probyn, and Alistair Busby. Impact of mussel (Mytilus galloprovincialis) raft-culture on benthic macrofauna, in situ oxygen uptake, and nutrient fluxes in Saldanha Bay, South Africa. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 58(5):1021–1031, May 2001. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f01-034.

Bochdansky:2001:CMS

[354] Alexander B. Bochdansky and Don Deibel. Consequences of model specification for the determination of gut evacuation rates: redefining the linear model. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 58(5):1032–1042, May 2001. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/f01-
041.

Kenchington:2001:EEO

[355] E. L. R. Kenchington, J. Prena, K. D. Gilkinson, D. C. Gordon, Jr., K. MacIsaac, C. Bourbonnais, P. J. Schwinghamer, T. W. Rowell, D. L. McKeown, and W. P. Vass. Effects of experimental otter trawling on the macrofauna of a sandy bottom ecosystem on the Grand Banks of Newfoundland. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 58(6):1043–1057, June 2001. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/f01-
053.

Pratt:2001:BIH

[356] Thomas C. Pratt and Michael G. Fox. Biotic influences on habitat selection by young-of-year walleye (Stizostedion vitreum) in the demersal stage. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 58(6):1058–1069, June 2001. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f01-054.

Brannon:2001:RAO

[357] Ernest L. Brannon, Keya C. M. Collins, Lawrence L. Moulton, and Keith R. Parker. Resolving allegations of oil damage to incubating pink salmon eggs in Prince William Sound. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 58(6):1070–1076, June 2001. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f01-055.

Minshall:2001:RCC

[358] G. Wayne Minshall, Todd V. Royer, and Christopher T. Robinson. Response of the Cache Creek macroinvertebrates during the first 10 years following disturbance by the 1988 Yellowstone wildfires. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 58(6):1077–1088, June 2001. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f01-056.

Isaak:2001:PSH

[359] Daniel J. Isaak and Wayne A. Hubert. Production of stream habitat gradients by montane watersheds: hypothesis tests based on spatially explicit path analyses. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 58(6): 1089–1103, June 2001. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f01-057.

McKenzie:2001:EABa

[360] D. J. McKenzie, E. Cataldi, P. Romano, S. F. Owen, E. W. Taylor, and P. Bronzi. Effects of acclimation to brackish water on the growth, respiratory metabolism, and swimming performance of young-of-the-year Adriatic sturgeon (Acipenser naccarii). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 58(6):1104–1112, June 2001. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f01-059.

McKenzie:2001:EABb

[361] D. J. McKenzie, E. Cataldi, P. Romano, E. W. Taylor, S. Cataudella, and P. Bronzi. Effects of acclimation to brackish water on tolerance of salinity challenge by young-of-the-year Adriatic sturgeon (Acipenser naccarii). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 58(6):1113–1121, June 2001. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f01-058.

Keeley:2001:PSS

[362] Ernest R. Keeley and James W. A. Grant. Prey size of salmonid fishes in streams, lakes, and oceans. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 58 (6):1122–1132, June 2001. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f01-060.

Harwood:2001:STE

[363] A. J. Harwood, N. B. Metcalfe, J. D. Armstrong, and S. W. Griffiths. Spatial and temporal effects of interspecific competition between Atlantic salmon (Salmo salar) and brown trout (Salmo trutta) in winter. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 58(6):1133–1140, June 2001. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f01-061.

Rindorf:2001:ALA

[364] Anna Rindorf and Peter Lewy. Analyses of length and age distributions using continuation-ratio logits. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 58(6):1141–1152, June 2001. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f01-062.

Bundy:2001:FEI

[365] Alida Bundy. Fishing on ecosystems: the interplay of fishing and predation in Newfoundland–Labrador. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 58(6):1153–1167, June 2001. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f01-063.

Kendall:2001:RPI

[366] Matthew S. Kendall, Donna L. Wolcott, Thomas G. Wolcott, and Anson H. Hines. Reproductive potential of individual male blue crabs, Callinectes sapidus, in a fished population: depletion and recovery of sperm number and seminal fluid. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 58 (6):1168–1177, June 2001. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f01-065.

Chen:2001:SME

[367] D. G. Chen and J. R. Irvine. A semiparametric model to examine stock– recruitment relationships incorporating environmental data. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 58(6):1178–1186, June 2001. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f01-037.

Good:2001:AVS

[368] S. P. Good, J. J. Dodson, M. G. Meekan, and D. A. J. Ryan. Annual variation in size-selective mortality of Atlantic salmon (Salmo salar) fry.

Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 58(6):1187–1195, June 2001. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f01-069.

Petrosky:2001:PSR

[369] C. E. Petrosky, H. A. Schaller, and P. Budy. Productivity and survival rate trends in the freshwater spawning and rearing stage of Snake River Chinook salmon (Oncorhynchus tshawytscha). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 58(6):1196–1207, June 2001. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f01-067.

Vanderploeg:2001:ZMD

[370] Henry A. Vanderploeg, James R. Liebig, Wayne W. Carmichael, Megan A. Agy, Thomas H. Johengen, Gary L. Fahnenstiel, and Thomas F. Nalepa. Zebra mussel (Dreissena polymorpha) selective filtration promoted toxic Microcystis blooms in Saginaw Bay (Lake Huron) and Lake Erie. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 58(6):1208– 1221, June 2001. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f01-066.

Joynt:2001:PIM

[371] Ernest H. Joynt III and Alexander P. Wolfe. Paleoenvironmental inference models from sediment diatom assemblages in Baffin Island lakes (Nunavut, Canada) and reconstruction of summer water temperature. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 58(6):1222–1243, June 2001. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f01-071.

Baker:2001:TAO

[372] M. Scott Baker, Jr., Charles A. Wilson, and Daniel L. VanGent. Testing assumptions of otolith radiometric aging with two long-lived fishes from the northern Gulf of Mexico. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 58 (6):1244–1252, June 2001. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f01-073.

Overman:2001:SIC

[373] Nathanael C. Overman and Donna L. Parrish. Stable isotope composition of walleye: ^{15}N accumulation with age and area-specific differences in δ 13C. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 58(6):1253–1260, June 2001. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f01-072.

Horvath:2001:RCC

[374] Thomas G. Horvath, Richard L. Whitman, and Laurel L. Last. Rapid communication / communication rapide: Establishment of two invasive crustaceans (Copepoda: Harpacticoida) in the nearshore sands of Lake Michigan. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 58(7):1261–1264, July 2001. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/f01-
089.

Gregr:2001:PCH

[375] Edward J. Gregr and Andrew W. Trites. Predictions of critical habitat for five whale species in the waters of coastal British Columbia. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 58(7):1265–1285, July 2001. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f01-078.

Prepas:2001:LVI

[376] E. E. Prepas, D. Planas, J. J. Gibson, D. H. Vitt, T. D. Prowse, W. P. Dinsmore, L. A. Halsey, P. M. McEachern, S. Paquet, G. J. Scrimgeour, W. M. Tonn, C. A. Paszkowski, and K. Wolfstein. Landscape variables influencing nutrients and phytoplankton communities in Boreal Plain lakes of northern Alberta: a comparison of wetland- and upland-dominated catchments. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 58(7):1286–1299, July 2001. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/f01-
081.

Sigler:2001:EIQ

[377] Michael F. Sigler and Chris R. Lunsford. Effects of individual quotas on catching efficiency and spawning potential in the Alaska sablefish fishery. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 58(7):1300–1312, July 2001. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f01-074.

Gray:2001:NBC

[378] Melissa M. Gray and Stephen C. Weeks. Niche breadth in clonal and sexual fish (Poeciliopsis): a test of the frozen niche variation model. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 58(7):1313–1318, July 2001. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f01-077.

Malard:2001:THH

[379] Florian Malard, Alain Mangin, Urs Uehlinger, and J. V. Ward. Thermal heterogeneity in the hyporheic zone of a glacial floodplain. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 58(7):1319–1335, July 2001. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f01-079.

Rowe:2001:MHM

[380] Sherrylynn Rowe. Movement and harvesting mortality of American lobsters (Homarus americanus) tagged inside and outside no-take reserves in Bonavista Bay, Newfoundland. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 58(7):1336–1346, July 2001. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f01-083.

Urban:2001:SBL

[381] N. R. Urban and A. E. Monte. Sulfur burial in and loss from the sediments of Little Rock Lake, Wisconsin. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 58(7):1347–1355, July 2001. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f01-085.

Hay:2001:TPH

[382] D. E. Hay, P. B. McCarter, and K. S. Daniel. Tagging of Pacific herring Clupea pallasi from 1936–1992: a review with comments on homing, geographic fidelity, and straying. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 58 (7):1356–1370, July 2001. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f01-087.

Beisner:2001:HVE

[383] Beatrix E. Beisner. Herbivory in variable environments: an experimental test of the effects of vertical mixing and Daphnia on phytoplankton community structure. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 58 (7):1371–1379, July 2001. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f01-080.

Seppa:2001:EFF

[384] Teija Seppä, Anssi Laurila, Nina Peuhkuri, Jorma Piironen, and Nicola Lower. Early familiarity has fitness consequences for Arctic char (Salvelinus alpinus) juveniles. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 58 (7):1380–1385, July 2001. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f01-084.

Anderson:2001:OSY

[385] John T. Anderson and George A. Rose. Offshore spawning and year-class strength of northern cod (2J3KL) during the fishing moratorium, 1994– 1996. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 58(7):1386–1394, July 2001. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f01-068.

Reynoldson:2001:CMP

[386] T. B. Reynoldson, D. M. Rosenberg, and V. H. Resh. Comparison of models predicting invertebrate assemblages for biomonitoring in the Fraser River catchment, British Columbia. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 58(7):1395–1410, July 2001. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f01-075.

Reece:2001:ISV

[387] Pamela F. Reece, Trefor B. Reynoldson, John S. Richardson, and David M. Rosenberg. Implications of seasonal variation for biomonitoring with predictive models in the Fraser River catchment, British Columbia. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 58(7):1411–1417, July 2001. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f01-076.

Greenfield:2001:PML

[388] Ben K. Greenfield, Thomas R. Hrabik, Chris J. Harvey, and Stephen R. Carpenter. Predicting mercury levels in yellow perch: use of water chemistry, trophic ecology, and spatial traits. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 58(7):1419–1429, July 2001. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f01-088.

Idrisi:2001:IZM

[389] Nasseer Idrisi, Edward L. Mills, Lars G. Rudstam, and Donald J. Stewart. Impact of zebra mussels (Dreissena polymorpha) on the pelagic lower trophic levels of Oneida Lake, New York. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 58(7):1430–1441, July 2001. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f01-070.

Weis:2001:PCA

[390] Judith S. Weis, Jennifer Samson, Tong Zhou, Joan Skurnick, and Peddrick Weis. Prey capture ability of mummichogs (Fundulus heteroclitus) as a behavioral biomarker for contaminants in estuarine systems. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 58(7):1442–1452, July 2001. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f01-086.

Sharma:2001:ERB

[391] Rishi Sharma and Ray Hilborn. Empirical relationships between watershed characteristics and Coho salmon (Oncorhynchus kisutch) smolt abundance in 14 western Washington streams. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 58(7):1453–1463, July 2001. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f01-091.

Myers:2001:WCC

[392] Ransom A. Myers, Brian R. MacKenzie, Keith G. Bowen, and Nicholas J. Barrowman. What is the carrying capacity for fish in the ocean? A meta-analysis of population dynamics of North Atlantic cod. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 58(7):1464–1476, July 2001. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f01-082.

Heyer:2001:MER

[393] Christopher J. Heyer, Thomas J. Miller, Fred P. Binkowski, Elaine M. Caldarone, and James A. Rice. Maternal effects as a recruitment mechanism in Lake Michigan yellow perch (Perca flavescens). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 58(7):1477–1487, July 2001. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f01-090.

Bowen:2001:EEU

[394] Jennifer L. Bowen and Ivan Valiela. The ecological effects of urbanization of coastal watersheds: historical increases in nitrogen loads and eutrophication of Waquoit Bay estuaries. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 58(8):1489–1500, August 2001. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f01-094.

Pyper:2001:SCS

[395] Brian J. Pyper, Franz J. Mueter, Randall M. Peterman, David J. Blackbourn, and Chris C. Wood. Spatial covariation in survival rates of Northeast Pacific pink salmon (Oncorhynchus gorbuscha). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 58(8):1501–1515, August 2001. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f01-096.

Koster:2001:DBCa

[396] Friedrich W. Köster, Christian Möllmann, Stefan Neuenfeldt, Michael A. St John, Maris Plikshs, and Rüdiger Voss. Developing Baltic cod recruitment models. I. Resolving spatial and temporal dynamics of spawning stock and recruitment for cod, herring, and sprat. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 58(8):1516–1533, August 2001. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f01-092.

Koster:2001:DBCb

[397] Friedrich W. Köster, Hans-Harald Hinrichsen, Michael A. St. John, Dietrich Schnack, Brian R. MacKenzie, Jonna Tomkiewicz, and Maris Plikshs. Developing Baltic cod recruitment models. II. Incorporation of environmental variability and species interaction. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 58(8):1534–1556, August 2001. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f01-093.

Kritzer:2001:CFP

[398] J. P. Kritzer, C. R. Davies, and B. D. Mapstone. Characterizing fish populations: effects of sample size and population structure on the precision of demographic parameter estimates. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 58(8):1557–1568, August 2001. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f01-098.

Harley:2001:HBM

[399] Shelton J. Harley and Ransom A. Myers. Hierarchical Bayesian models of length-specific catchability of research trawl surveys. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 58(8):1569–1584, August 2001. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f01-097.

Leonard:2001:MEA

[400] Jill B. K. Leonard and Stephen D. McCormick. Metabolic enzyme activity during smolting in stream- and hatchery-reared Atlantic salmon (Salmo salar). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 58(8):1585– 1593, August 2001. CODEN CJFSDX. ISSN 0706-652X (print), 1205-

7533 (electronic). URL https://cdnsciencepub.com/doi/full/10.
1139/f01-105.

Claret:2001:FPV

[401] C. Claret, A. J. Boulton, M.-J. Dole-Olivier, and P. Marmonier. Functional processes versus state variables: interstitial organic matter pathways in floodplain habitats. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 58 (8):1594–1602, August 2001. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f01-106.

Lougheed:2001:PDM

[402] Vanessa L. Lougheed, Barb Crosbie, and Patricia Chow-Fraser. Primary determinants of macrophyte community structure in 62 marshes across the Great Lakes basin: latitude, land use, and water quality effects. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 58(8):1603–1612, August 2001. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f01-102.

Fu:2001:WAC

[403] Caihong Fu, Robert Mohn, and L. Paul Fanning. Why the Atlantic cod (Gadus morhua) stock off eastern Nova Scotia has not recovered. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 58(8):1613–1623, August 2001. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f01-095.

Siler:2001:LTE

[404] Edward R. Siler, J. Bruce Wallace, and S. L. Eggert. Long-term effects of resource limitation on stream invertebrate drift. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 58(8):1624–1637, August 2001. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f01-101.

Steedman:2001:LWT

[405] Robert J. Steedman, Robert S. Kushneriuk, and Robert L. France. Littoral water temperature response to experimental shoreline logging around small boreal forest lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 58(8):1638–1647, August 2001. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f01-103.

Su:2001:HBM

[406] Zhenming Su, Milo D. Adkison, and Benjamin W. Van Alen. A hierarchical Bayesian model for estimating historical salmon escapement and escapement timing. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 58(8): 1648–1662, August 2001. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f01-099.

Adkison:2001:CSE

[407] Milo D. Adkison and Zhenming Su. A comparison of salmon escapement estimates using a hierarchical Bayesian approach versus separate maximum likelihood estimation of each year’s return. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 58(8):1663–1671, August 2001. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f01-100.

Weng:2001:NDR

[408] Zhongyan Weng, Nandita Mookerji, and Asit Mazumder. Nutrientdependent recovery of Atlantic salmon streams from a catastrophic flood. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 58(8):1672–1682, August 2001. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f01-107.

vanOostenbrugge:2001:RAA

[409] J. A. E. van Oostenbrugge, W. L. T. van Densen, and M. A. M. Machiels. Risk aversion in allocating fishing effort in a highly uncertain coastal fishery for pelagic fish, Moluccas, Indonesia. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 58(8):1683–1691, August 2001. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f01-115.

Ek:2001:PAE

[410] A. S. Ek and T. Korsman. A paleolimnological assessment of the effects of post-1970 reductions of sulfur deposition in Sweden. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 58(8):1692–1700, August 2001. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f01-108.

Beer:2001:EES

[411] W. Nicholas Beer and James J. Anderson. Erratum: Effect of spawning day and temperature on salmon emergence: interpretations of a growth model for Methow River Chinook. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 58(8):1701, August 2001. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f01-122.

Frank:2001:RCC

[412] Kenneth T. Frank and Nancy L. Shackell. Rapid communication / communication rapide: Area-dependent patterns of finfish diversity in a large marine ecosystem. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 58(9):1703– 1707, September 2001. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f01-123.

Ballance:2001:ISB

[413] S. Ballance, P. J. Phillips, C. R. McCrohan, J. J. Powell, R. Jugdaohsingh, and K. N. White. Influence of sediment biofilm on the behaviour of aluminum and its bioavailability to the snail Lymnaea stagnalis in neutral freshwater. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 58(9):1708– 1715, September 2001. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f01-104.

Latour:2001:DMT

[414] Robert J. Latour, John M. Hoenig, John E. Olney, and Kenneth H. Pollock. Diagnostics for multiyear tagging models with application to Atlantic striped bass (Morone saxatilis). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 58(9):1716–1726, September 2001. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f01-117.

Meding:2001:BIE

[415] Marianne E. Meding and Leland J. Jackson. Biological implications of empirical models of winter oxygen depletion. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 58(9):1727–1736, September 2001. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f01-109.

Croteau:2001:DCA

[416] Marie-Noële Croteau, Landis Hare, and André Tessier. Differences in Cd accumulation among species of the lake-dwelling biomonitor Chaoborus.

Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 58(9):1737–1746, September 2001. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f01-116.

Mermillod-Blondin:2001:TFR

[417] Florian Mermillod-Blondin, Magali Gérino, Valérie Degrange, Robert Lensi, Jean-Luc Chassé, Michael Rard, and Michel Creuzé des Châtelliers. Testing the functional redundancy of Limnodrilus and Tubifex (Oligochaeta, Tubificidae) in hyporheic sediments: an experimental study in microcosms. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 58 (9):1747–1759, September 2001. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/

doi/full/10.1139/f01-119.

Harley:2001:CPU

[418] Shelton J. Harley, Ransom A. Myers, and Alistair Dunn. Is catchper-unit-effort proportional to abundance? Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 58(9):1760–1772, September 2001. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f01-112.

Schroeter:2001:UMR

[419] Stephen C. Schroeter, Daniel C. Reed, David J. Kushner, James A. Estes, and David S. Ono. The use of marine reserves in evaluating the dive fishery for the warty sea cucumber (Parastichopus parvimensis) in California, U.S.A. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 58(9):1773– 1781, September 2001. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f01-127.

Cao:2001:ESS

[420] Yong Cao, David P. Larsen, and Robert M. Hughes. Evaluating sampling sufficiency in fish assemblage surveys: a similarity-based approach. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 58(9):1782–1793, September 2001. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f01-120.

Tanner:2001:IHR

[421] Jason E. Tanner. Influence of harvest refugia on penaeid prawn population dynamics and sustainable catch. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 58(9):1794–1804, September 2001. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f01-124.

Morita:2001:NMG

[422] Kentaro Morita and Takashi Matsuishi. A new model of growth backcalculation incorporating age effect based on otoliths. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 58(9):1805–1811, September 2001. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f01-126.

Letcher:2001:PGA

[423] Benjamin H. Letcher and Tim L. King. Parentage and grandparentage assignment with known and unknown matings: application to Connecticut River Atlantic salmon restoration. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 58(9):1812–1821, September 2001. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f01-125.

Donald:2001:RZA

[424] David B. Donald, Rolf D. Vinebrooke, R. Stewart Anderson, Jim Syrgiannis, and Mark D. Graham. Recovery of zooplankton assemblages in mountain lakes from the effects of introduced sport fish. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 58(9):1822–1830, September 2001. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f01-121.

Petersen:2001:CRW

[425] James H. Petersen and James F. Kitchell. Climate regimes and water temperature changes in the Columbia River: bioenergetic implications for predators of juvenile salmon. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 58(9):1831–1841, September 2001. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f01-111.

McGlashan:2001:WDP

[426] D. J. McGlashan, J. M. Hughes, and S. E. Bunn. Within-drainage population genetic structure of the freshwater fish Pseudomugil signifer (Pseudomugilidae) in northern Australia. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 58(9):1842–1852, September 2001. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f01-113.

Hansen:2001:AAS

[427] Michael M. Hansen, Einar E. Nielsen, Dorte Bekkevold, and Karen-Lise D. Mensberg. Admixture analysis and stocking impact assessment in brown trout (Salmo trutta), estimated with incomplete baseline data. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 58(9):1853–1860, September 2001. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f01-135.

Stevick:2001:EIU

[428] Peter T. Stevick, Per J. Palsbøll, Tim D. Smith, Mark V. Bravington, and Philip S. Hammond. Errors in identification using natural markings: rates, sources, and effects on capture–recapture estimates of abundance. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 58(9):1861–1870, September 2001. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f01-131.

McAllister:2001:UDM

[429] M. K. McAllister, E. K. Pikitch, and E. A. Babcock. Using demographic methods to construct Bayesian priors for the intrinsic rate of increase in the Schaefer model and implications for stock rebuilding. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 58(9):1871–1890, September 2001. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f01-114.

Jacobson:2001:SPV

[430] Larry D. Jacobson, Jose A. A. De Oliveira, Manuel Barange, Miguel A. Cisneros-Mata, Roberto Félix-Uraga, John R. Hunter, Jin Yeong Kim, Yasunobu Matsuura, Miguel Ñiquen, Carmela Porteiro, Brian Rothschild, Ramiro P. Sanchez, Rodolfo Serra, Andres Uriarte, and Tokio Wada. Surplus production, variability, and climate change in the great sardine and anchovy fisheries. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 58 (9):1891–1903, September 2001. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/

doi/full/10.1139/f01-110.

Downing:2001:PCD

[431] John A. Downing, Susan B. Watson, and Edward McCauley. Predicting cyanobacteria dominance in lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 58(10):1905–1908, October 2001. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f01-143.

Heikinheimo:2001:EPL

[432] Outi Heikinheimo. Effect of predation on the low-density dynamics of vendace: significance of the functional response. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 58(10):1909–1923, October 2001. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f01-132.

Attayde:2001:FMN

[433] José Luiz Attayde and Lars-Anders Hansson. Fish-mediated nutrient recycling and the trophic cascade in lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 58(10):1924–1931, October 2001. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f01-128.

Farrell:2001:SRP

[434] A. P. Farrell, P. E. Gallaugher, J. Fraser, D. Pike, P. Bowering, A. K. M. Hadwin, W. Parkhouse, and R. Routledge. Successful recovery of the physiological status of Coho salmon on board a commercial gillnet vessel by means of a newly designed revival box. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 58(10):1932–1946, October 2001. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f01-136.

Kahler:2001:SMG

[435] Thomas H. Kahler, Philip Roni, and Thomas P. Quinn. Summer movement and growth of juvenile anadromous salmonids in small western Washington streams. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 58(10): 1947–1956, October 2001. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f01-134.

OBrien:2001:FRD

[436] W. John O’Brien, Michael Barfield, and Karen Sigler. The functional response of drift-feeding Arctic grayling: the effects of prey density, water velocity, and location efficiency. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 58(10):1957–1963, October 2001. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f01-138.

Ballan-Dufrancais:2001:CTD

[437] C. Ballan-Dufrançais, A. Y. Jeantet, A. Geffard, J. C. Amiard, and C. Amiard-Triquet. Cellular and tissular distribution of copper in an intrasedimentary bivalve, the Baltic clam Macoma balthica, originating from a clean or a metal-rich site. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 58(10):1964–1974, October 2001. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f01-129.

Johannsson:2001:DMR

[438] Ora E. Johannsson, Mike F. Leggett, Lars G. Rudstam, Mark R. Servos, M. Ali Mohammadian, Gideon Gal, Ron M. Dermott, and Ray H. Hesslein. Diet of Mysis relicta in Lake Ontario as revealed by stable isotope and gut content analysis. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 58(10):1975–1986, October 2001. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f01-118.

Skalski:2001:ERS

[439] John R. Skalski, James Lady, Richard Townsend, Albert E. Giorgi, John R. Stevenson, Charles M. Peven, and Robert D. McDonald. Estimating in-river survival of migrating salmonid smolts using radiotelemetry. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 58(10):1987–1997, October 2001. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f01-133.

Brodeur:2001:RFC

[440] Philippe Brodeur, Pierre Magnan, and Michel Legault. Response of fish communities to different levels of white sucker (Catostomus commersoni) biomanipulation in five temperate lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 58(10):1998–2010, October 2001. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f01-137.

Keller:2001:CDN

[441] W. Keller, S. S. Dixit, and J. Heneberry. Calcium declines in northeastern Ontario lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 58(10): 2011–2020, October 2001. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f01-142.

Hobday:2001:RCO

[442] Alistair J. Hobday and George W. Boehlert. The role of coastal ocean variation in spatial and temporal patterns in survival and size of Coho salmon (Oncorhynchus kisutch). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 58(10):2021–2036, October 2001. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f01-145.

Hagen:2001:RPF

[443] John Hagen and Eric B. Taylor. Resource partitioning as a factor limiting gene flow in hybridizing populations of Dolly Varden char (Salvelinus malma) and bull trout (Salvelinus confluentus). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 58(10):2037–2047, October 2001. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f01-141.

Mingelbier:2001:CCA

[444] M. Mingelbier, F. Lecomte, and J. J. Dodson. Climate change and abundance cycles of two sympatric populations of smelt (Osmerus mordax) in the middle estuary of the St. Lawrence River, Canada. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 58(10):2048–2058, October 2001. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f01-140.

Watkins:2001:ESU

[445] Elise Marie Watkins, David W. Schindler, Michael A. Turner, and David Findlay. Effects of solar ultraviolet radiation on epilithic metabolism, and nutrient and community composition in a clear-water boreal lake. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 58(10):2059–2070, October 2001. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f01-150.

Anthony:2001:ETD

[446] James L. Anthony and John A. Downing. Exploitation trajectory of a declining fauna: a century of freshwater mussel fisheries in North America. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 58(10):2071–2090, October 2001. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f01-130.

Bjornsson:2001:CFY

[447] Björn Björnsson. Can fisheries yield be enhanced by large-scale feeding of a predatory fish stock? A case study of the Icelandic cod stock. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 58(10):2091–2104, October 2001. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f01-144.

Williams:2001:TSS

[448] John G. Williams. Tripping over spatial scales: a comment on guay et al. (2000). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 58(10):2105–2107, October 2001. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/f01-
148.

Guay:2001:SES

[449] J. C. Guay, D. Boisclair, M. Leclerc, M. Lapointe, and P. Legendre. Science on the edge of spatial scales: a reply to the comments of Williams (2001). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 58(10):2108–2111, October 2001. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/f01-

Gao:2001:SIC

[450] Y. W. Gao, S. H. Joner, and G. G. Bargmann. Stable isotopic composition of otoliths in identification of spawning stocks of Pacific herring (Clupea pallasi) in Puget Sound. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 58(11):2113–2120, November 2001. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f01-146.

Johns:2001:ABP

[451] D. G. Johns, M. Edwards, and S. D. Batten. Arctic boreal plankton species in the Northwest Atlantic. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 58(11):2121–2124, November 2001. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f01-156.

Maruyama:2001:CSN

[452] Atsushi Maruyama, Yoshihiro Yamada, Bosco Rusuwa, and Masahide Yuma. Change in stable nitrogen isotope ratio in the muscle tissue of a migratory goby, Rhinogobius sp., in a natural setting. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 58(11):2125–2128, November 2001. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f01-147.

Essington:2001:BGF

[453] Timothy E. Essington, James F. Kitchell, and Carl J. Walters. The von Bertalanffy growth function, bioenergetics, and the consumption rates of fish. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 58(11):2129–2138, November 2001. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f01-151.

Chen:2001:DER

[454] D. G. Chen. Detecting environmental regimes in fish stock-recruitment relationships by fuzzy logic. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 58 (11):2139–2148, November 2001. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/

doi/full/10.1139/f01-155.

Gauthier:2001:DTU

[455] Stéphane Gauthier and George A. Rose. Diagnostic tools for unbiased in situ target strength estimation. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 58(11):2149–2155, November 2001. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f01-152.

Wurtsbaugh:2001:EEV

[456] Wayne A. Wurtsbaugh, Howard P. Gross, Phaedra Budy, and Chris Luecke. Effects of epilimnetic versus metalimnetic fertilization on the phytoplankton and periphyton of a mountain lake with a deep chlorophyll maxima. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 58(11):2156– 2166, November 2001. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f01-157.

Collie:2001:BRP

[457] Jeremy S. Collie and Henrik Gislason. Biological reference points for fish stocks in a multispecies context. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 58(11):2167–2176, November 2001. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f01-158.

Effler:2001:PGD

[458] Steven W. Effler, Carol M. Matthews (Brooks), and David A. Matthews. Patterns of gross deposition in reservoirs enriched in inorganic tripton. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 58(11):2177–2188, November 2001. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f01-163.

Haas:2001:PWG

[459] Gordon R. Haas and J. D. McPhail. The post-Wisconsinan glacial biogeography of bull trout (Salvelinus confluentus): a multivariate morphometric approach for conservation biology and management. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 58(11):2189–2203, November 2001. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f01-139.

Pepin:2001:RME

[460] P. Pepin, J. F. Dower, and H. P. Benoît. The role of measurement error on the interpretation of otolith increment width in the study of growth in larval fish. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 58(11): 2204–2212, November 2001. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f01-159.

Shaw:2001:DIE

[461] E. Al Shaw and John S. Richardson. Direct and indirect effects of sediment pulse duration on stream invertebrate assemblages and rainbow trout (Oncorhynchus mykiss) growth and survival. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 58(11):2213–2221, November 2001. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f01-160.

Stemberger:2001:SZR

[462] Richard S. Stemberger, David P. Larsen, and Thomas M. Kincaid. Sensitivity of zooplankton for regional lake monitoring. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 58(11):2222–2232, November 2001. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f01-164.

DeLong:2001:EGM

[463] Allison K. DeLong, Jeremy S. Collie, Carol J. Meise, and J. Christopher Powell. Estimating growth and mortality of juvenile winter flounder, Pseudopleuronectes americanus, with a length-based model. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 58(11):2233–2246, November 2001. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f01-162.

Vainola:2001:PGR

[464] Risto Väinölä, Jouni K. Vainio, and Jukka U. Palo. Phylogeography of “glacial relict” Gammaracanthus (Crustacea, Amphipoda) from boreal lakes and the Caspian and White seas. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 58(11):2247–2257, November 2001. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f01-165.

Haas:2001:EBF

[465] Heather L. Haas, E. Conrad Lamon III, Kenneth A. Rose, and Richard F. Shaw. Environmental and biological factors associated with the stagespecific abundance of brown shrimp (Penaeus aztecus) in Louisiana: applying a new combination of statistical techniques to long-term monitoring data. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 58(11):2258– 2270, November 2001. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f01-169.

Margolis:2001:IBI

[466] Brian E. Margolis, Mark S. Castro, and Richard L. Raesly. The impact of beaver impoundments on the water chemistry of two Appalachian streams. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 58(11):2271–2283, November 2001. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f01-166.

Rivot:2001:HRB

[467] E. Rivot, E. Prévost, and E. Parent. How robust are Bayesian posterior inferences based on a Ricker model with regards to measurement errors and prior assumptions about parameters? Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 58(11):2284–2297, November 2001. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f01-167.

Nordwall:2001:ICE

[468] Fredrik Nordwall, Ingemar Näslund, and Erik Degerman. Intercohort competition effects on survival, movement, and growth of brown trout (Salmo trutta) in Swedish streams. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 58(11):2298–2308, November 2001. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f01-170.

Hintelmann:2001:CVI

[469] Holger Hintelmann, Peter Dillon, R. Douglas Evans, John W. M. Rudd, and R. A. (Drew) Bodaly. Comment: Variations in the isotope composition of mercury in a freshwater sediment sequence and food web. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 58(11):2309–2311, November 2001. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f01-153.

Jackson:2001:RVI

[470] Togwell A. Jackson. Reply: Variations in the isotope composition of mercury in a freshwater sediment sequence and food web. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 58(11):2312–2316, November 2001. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f01-154.

Hedrick:2001:EIT

[471] Philip W. Hedrick. Erratum: Invasion of transgenes from salmon or other genetically modified organisms into natural populations. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 58(11):2317, November 2001. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f01-161.

Farrell:2001:RRE

[472] Anthony P. Farrell, Patricia E. Gallaugher, and Richard Routledge. Rapid recovery of exhausted adult Coho salmon after commercial capture by troll fishing. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 58(12): 2319–2324, December 2001. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f01-188.

Robichaud:2001:MHA

[473] D. Robichaud and G. A. Rose. Multiyear homing of Atlantic cod to a spawning ground. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 58(12): 2325–2329, December 2001. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f01-190.

Bernal:2001:NAM

[474] M. Bernal, D. L. Borchers, L. Valdés, A. Lago de Lanzós, and S. T. Buckland. A new ageing method for eggs of fish species with daily spawning synchronicity. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 58(12): 2330–2340, December 2001. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f01-168.

Yan:2001:CZP

[475] Norman D. Yan, Agnes Blukacz, W. Gary Sprules, Paul K. Kindy, David Hackett, Robert E. Girard, and Bev J. Clark. Changes in zooplankton and the phenology of the spiny water flea, Bythotrephes, following its invasion of Harp Lake, Ontario, Canada. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 58(12):2341–2350, December 2001. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f01-171.

Hart:2001:IBY

[476] Deborah R. Hart. Individual-based yield-per-recruit analysis, with an application to the Atlantic sea scallop, Placopecten magellanicus. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 58(12):2351–2358, December 2001. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f01-175.

Arscott:2001:THA

[477] David B. Arscott, Klement Tockner, and J. V. Ward. Thermal heterogeneity along a braided floodplain river (Tagliamento River, northeastern Italy). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 58(12):2359–2373, December 2001. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f01-183.

Vrieze:2001:LAR

[478] Lance A. Vrieze and Peter W. Sorensen. Laboratory assessment of the role of a larval pheromone and natural stream odor in spawning stream localization by migratory sea lamprey (Petromyzon marinus). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 58(12):2374–2385, December 2001. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f01-179.

Webb:2001:DAS

[479] J. H. Webb, R. J. Fryer, J. B. Taggart, C. E. Thompson, and A. F. Youngson. Dispersion of Atlantic salmon (Salmo salar) fry from competing families as revealed by DNA profiling. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 58(12):2386–2395, December 2001. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f01-177.

Soucek:2001:ADA

[480] David J. Soucek, Donald S. Cherry, and Carl E. Zipper. Aluminumdominated acute toxicity to the cladoceran Ceriodaphnia dubia in neutral waters downstream of an acid mine drainage discharge. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 58(12):2396–2404, December 2001. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f01-185.

Belzile:2001:BOC

[481] Claude Belzile, Warwick F. Vincent, John A. E. Gibson, and Patrick Van Hove. Bio-optical characteristics of the snow, ice, and water column of a perennially ice-covered lake in the high Arctic. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 58(12):2405–2418, December 2001. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f01-187.

Deriso:2001:RPD

[482] Richard B. Deriso, David R. Marmorek, and Ian J. Parnell. Retrospective patterns of differential mortality and common year-effects experienced by spring and summer Chinook salmon (Oncorhynchus tshawytscha) of the Columbia River. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 58(12): 2419–2430, December 2001. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f01-172.

Peters:2001:ADAa

[483] Calvin N. Peters and David R. Marmorek. Application of decision analysis to evaluate recovery actions for threatened Snake River spring and summer Chinook salmon (Oncorhynchus tshawytscha). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 58(12):2431–2446, December 2001. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f01-173.

Peters:2001:ADAb

[484] Calvin N. Peters, David R. Marmorek, and Richard B. Deriso. Application of decision analysis to evaluate recovery actions for threatened Snake River fall Chinook salmon (Oncorhynchus tshawytscha). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 58(12):2447–2458, December 2001. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f01-174.

Mayer:2001:ZMD

[485] C. M. Mayer, L. G. Rudstam, E. L. Mills, S. G. Cardiff, and C. A. Bloom. Zebra mussels (Dreissena polymorpha), habitat alteration, and yellow perch (Perca flavescens) foraging: system-wide effects and behavioural mechanisms. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 58(12): 2459–2467, December 2001. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f01-176.

Paterson:2001:SCI

[486] Andrew M. Paterson, Brian F. Cumming, John P. Smol, and Roland I. Hall. Scaled chrysophytes as indicators of water quality changes since preindustrial times in the Muskoka–Haliburton region, Ontario, Canada. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 58(12):2468–2481, December 2001. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f01-180.

Frusher:2001:ILS

[487] S. D. Frusher and J. M. Hoenig. Impact of lobster size on selectivity of traps for southern rock lobster (Jasus edwardsii). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 58(12):2482–2489, December 2001. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f01-181.

Frusher:2001:ENF

[488] S. D. Frusher and J. M. Hoenig. Estimating natural and fishing mortality and tag reporting rate of southern rock lobster (Jasus edwardsii) from a multiyear tagging model. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 58(12): 2490–2501, December 2001. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f01-182.

Claytor:2001:PAI

[489] Ross R. Claytor and Jacques Allard. Properties of abundance indices obtained from acoustic data collected by inshore herring gillnet boats. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 58(12):2502–2512, December 2001. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f01-189.

Ricciardi:2001:FIA

[490] Anthony Ricciardi. Facilitative interactions among aquatic invaders: is an “invasional meltdown” occurring in the Great Lakes? Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 58(12):2513–2525, December 2001. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f01-178.

Anonymous:2001:AII

[491] Anonymous. Author index / index des auteurs: Volume 58, 2001. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 58(12), December 2001. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfasauthor01.

Anonymous:2001:CSV

[492] Anonymous. Contents / sommaire: Volume 58, 2001. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 58(12), December 2001. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfascontents01.

Sherwood:2002:SFW

[493] Graham D. Sherwood, Jennifer Kovecses, Alice Hontela, and Joseph B. Rasmussen. Simplified food webs lead to energetic bottlenecks in polluted lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 59(1):1–5, January 2002. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f01-213.

Jansen:2002:NSB

[494] Peder A. Jansen, Henning Slettvold, Anders G. Finstad, and Arnfinn Langeland. Niche segregation between Arctic char (Salvelinus alpinus) and brown trout (Salmo trutta): an experimental study of mechanisms. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 59(1):6–11, January 2002. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f01-184.

Cattaneo:2002:IHB

[495] Franck Cattanéo, Nicolas Lamouroux, Pascal Breil, and Hervé Capra. The influence of hydrological and biotic processes on brown trout (Salmo trutta) population dynamics. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 59 (1):12–22, January 2002. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f01-186.

Simpfendorfer:2002:EST

[496] Colin A. Simpfendorfer, Michelle R. Heupel, and Robert E. Hueter. Estimation of short-term centers of activity from an array of omnidirectional hydrophones and its use in studying animal movements. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 59(1):23–32, January 2002. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f01-191.

Hjellvik:2002:DVB

[497] Vidar Hjellvik, Olav Rune Godø, and Dag Tjøstheim. Diurnal variation in bottom trawl survey catches: does it pay to adjust? Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 59(1):33–48, January 2002. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f01-193.

Waters:2002:HDT

[498] Jonathan M. Waters, Michael Shirley, and Gerard P. Closs. Hydroelectric development and translocation of Galaxias brevipinnis: a cloud at the end of the tunnel? Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 59 (1):49–56, January 2002. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f01-195.

Ottersen:2002:DAM

[499] Geir Ottersen, Kristin Helle, and Bjarte Bogstad. Do abiotic mechanisms determine interannual variability in length-at-age of juvenile Arcto–Norwegian cod? Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 59 (1):57–65, January 2002. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f01-197.

Collins:2002:HCD

[500] Brian D. Collins, David R. Montgomery, and Andrew D. Haas. Historical changes in the distribution and functions of large wood in Puget Lowland rivers. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 59(1):66–76, January 2002. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f01-199.

Heath:2002:RBH

[501] Daniel D. Heath, Colleen A. Bryden, J. Mark Shrimpton, George K. Iwama, Joanne Kelly, and John W. Heath. Relationships between heterozygosity, allelic distance (d^{2}), and reproductive traits in Chinook salmon, Oncorhynchus tshawytscha. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 59(1):77–84, January 2002. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f01-192.

Klug:2002:PNE

[502] Jennifer L. Klug. Positive and negative effects of allochthonous dissolved organic matter and inorganic nutrients on phytoplankton growth. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 59(1):85–95, January 2002. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f01-194.

Peterson:2002:ROB

[503] Charles H. Peterson. Recruitment overfishing in a bivalve mollusc fishery: hard clams (Mercenaria mercenaria) in North Carolina. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 59(1):96–104, January 2002. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f01-196.

Pepin:2002:ERR

[504] P. Pepin, J. F. Dower, J. A. Helbig, and W. C. Leggett. Estimating the relative roles of dispersion and predation in generating regional differences in mortality rates of larval radiated shanny (Ulvaria subbifurcata). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 59(1):105–114, January 2002. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f01-198.

Harvey:2002:UBM

[505] Chris J. Harvey, Paul C. Hanson, Timothy E. Essington, Paul B. Brown, and James F. Kitchell. Using bioenergetics models to predict stable isotope ratios in fishes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 59(1): 115–124, January 2002. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f01-203.

Singer:2002:STG

[506] Thomas D. Singer, Koreen M. Clements, Jeffrey W. Semple, Patricia M. Schulte, Jason S. Bystriansky, Bengt Finstad, Ian A. Fleming, and R. Scott McKinley. Seawater tolerance and gene expression in two strains of Atlantic salmon smolts. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 59 (1):125–135, January 2002. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f01-205.

Nielsen:2002:CFP

[507] Anders Nielsen and Peter Lewy. Comparison of the frequentist properties of Bayes and the maximum likelihood estimators in an age-structured fish stock assessment model. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 59 (1):136–143, January 2002. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f01-204.

Liley:2002:FDR

[508] N. R. Liley, P. Tamkee, R. Tsai, and D. J. Hoysak. Fertilization dynamics in rainbow trout (Oncorhynchus mykiss): effect of male age, social experience, and sperm concentration and motility on in vitro fertilization. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 59(1):144–152, January 2002. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f01-202.

Carls:2002:STI

[509] M. G. Carls, G. D. Marty, and J. E. Hose. Synthesis of the toxicological impacts of the Exxon Valdez oil spill on Pacific herring (Clupea pallasi) in Prince William Sound, Alaska, U.S.A. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 59(1):153–172, January 2002. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f01-200.

MacKenzie:2002:EHH

[510] Brian R. MacKenzie, Jürgen Alheit, Daniel J. Conley, Poul Holm, and Carl Christian Kinze. Ecological hypotheses for a historical reconstruction of upper trophic level biomass in the Baltic Sea and Skagerrak. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 59(1):173–190, January 2002. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f01-201.

Gilbert:2002:CDS

[511] D. J. Gilbert. Comment: Does spawning stock affect recruitment of New England groundfish? Interpreting spawning stock and recruitment data in New England groundfish. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 59 (1):191–192, January 2002. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f01-207.

Brodziak:2002:RDS

[512] J. K. T. Brodziak, W. J. Overholtz, and P. J. Rago. Reply: Does spawning stock affect recruitment of New England groundfish? Interpreting spawning stock and recruitment data in New England groundfish. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 59(1):193–195, January 2002. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f01-208.

Tanasichuk:2002:EAS

[513] R. W. Tanasichuk. Erratum: Age-specific natural mortality rates of adult Pacific herring (Clupea pallasi) from southern British Columbia. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 59(1):196, January 2002. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f01-206.

Anonymous:2002:IA

[514] Anonymous. Instructions to authors. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 59(1), January 2002. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas_instruct02_e.

Anonymous:2002:RA

[515] Anonymous. Recommandations aux auteurs. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 59(1), January 2002. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/cjfas_instruct02_f.

Simard:2002:CHD

[516] Yvan Simard, Diane Lavoie, and François J. Saucier. Channel head dynamics: capelin (Mallotus villosus) aggregation in the tidally driven upwelling system of the Saguenay–St. Lawrence Marine Park’s whale feeding ground. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 59(2):197–210, February 2002. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f01-210.

Fiksen:2002:SER

[517] Øyvind Fiksen and Aril Slotte. Stock–environment recruitment models for Norwegian spring spawning herring (Clupea harengus). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 59(2):211–217, February 2002. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f02-002.

Maki-Petays:2002:THS

[518] Aki Mäki-Petäys, Ari Huusko, Jaakko Erkinaro, and Timo Muotka. Transferability of habitat suitability criteria of juvenile Atlantic salmon (Salmo salar). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 59(2):218–228, February 2002. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f01-209.

Sherwood:2002:SGE

[519] Graham D. Sherwood, Ivano Pazzia, Andrew Moeser, Alice Hontela, and Joseph B. Rasmussen. Shifting gears: enzymatic evidence for the energetic advantage of switching diet in wild-living fish. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 59(2):229–241, February 2002. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f02-001.

Chen:2002:IAE

[520] D. G. Chen, J. R. Irvine, and A. J. Cass. Incorporating Allee effects in fish stock–recruitment models and applications for determining reference points. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 59(2):242–249, February 2002. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f02-005.

He:2002:SEE

[521] Ji X. He and Donald J. Stewart. A stage-explicit expression of the von Bertalanffy growth model for understanding age at first reproduction of Great Lakes fishes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 59(2): 250–261, February 2002. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f02-008.

Snyder:2002:IEH

[522] Craig D. Snyder, John A. Young, David P. Lemarié, and David R. Smith. Influence of eastern hemlock (Tsuga canadensis) forests on aquatic invertebrate assemblages in headwater streams. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 59(2):262–275, February 2002. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f02-003.

Stewart:2002:TSV

[523] David C. Stewart, Gordon W. Smith, and Alan F. Youngson. Tributaryspecific variation in timing of return of adult Atlantic salmon (Salmo salar) to fresh water has a genetic component. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 59(2):276–281, February 2002. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f02-011.

Eldridge:2002:DRS

[524] William H. Eldridge, Marc D. Bacigalupi, Ira R. Adelman, Loren M. Miller, and Anne R. Kapuscinski. Determination of relative survival of two stocked walleye populations and resident natural-origin fish by microsatellite DNA parentage assignment. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 59(2):282–290, February 2002. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f02-007.

Spencer:2002:MSY

[525] Paul D. Spencer, Thomas K. Wilderbuer, and Chang Ik Zhang. A mixedspecies yield model for eastern Bering Sea shelf flatfish fisheries. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 59(2):291–302, February 2002. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f02-012.

Imre:2002:EVI

[526] I. Imre, J. W. A. Grant, and E. R. Keeley. The effect of visual isolation on territory size and population density of juvenile rainbow trout (Oncorhynchus mykiss). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 59(2): 303–309, February 2002. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f02-010.

Arbuckle:2002:FMA

[527] Kelly E. Arbuckle and John A. Downing. Freshwater mussel abundance and species richness: GIS relationships with watershed land use and geology. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 59(2):310–316, February 2002. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f02-006.

Hiriart:2002:EUR

[528] Véronique P. Hiriart, Bruce M. Greenberg, Stephanie J. Guildford, and Ralph E. H. Smith. Effects of ultraviolet radiation on rates and size distribution of primary production by Lake Erie phytoplankton. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 59(2):317–328, February 2002. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f01-212.

Fallu:2002:DFD

[529] Marie-Andrée Fallu, Nancie Allaire, and Reinhard Pienitz. Distribution of freshwater diatoms in 64 Labrador (Canada) lakes: species– environment relationships along latitudinal gradients and reconstruction models for water colour and alkalinity. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 59(2):329–349, February 2002. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f01-211.

Bertram:2002:EFW

[530] Barbara O. B. Bertram and Richard C. Playle. Effects of feeding on waterborne silver uptake and depuration in rainbow trout (Oncorhynchus mykiss). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 59(2):350–360, February 2002. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f02-004.

Sinclair:2002:MCD

[531] A. F. Sinclair, D. P. Swain, and J. M. Hanson. Measuring changes in the direction and magnitude of size-selective mortality in a commercial fish population. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 59(2): 361–371, February 2002. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f02-015.

Sinclair:2002:DES

[532] A. F. Sinclair, D. P. Swain, and J. M. Hanson. Disentangling the effects of size-selective mortality, density, and temperature on length-at-age. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 59(2):372–382, February 2002. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f02-014.

Millar:2002:ASM

[533] Russell B. Millar and Richard D. Methot, Jr. Age-structured metaanalysis of U.S. West Coast rockfish (Scorpaenidae) populations and hierarchical modeling of trawl survey catchabilities. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 59(2):383–392, February 2002. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f02-009.

Munch:2002:ALP

[534] Stephan B. Munch and D. O. Conover. Accounting for local physiological adaptation in bioenergetic models: testing hypotheses for growth rate evolution by virtual transplant experiments. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 59(2):393–403, February 2002. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f02-013.

Apostolaki:2002:MEE

[535] P. Apostolaki, E. J. Milner-Gulland, M. K. McAllister, and G. P. Kirkwood. Modelling the effects of establishing a marine reserve for mobile fish species. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 59(3):405–415, March 2002. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f02-018.

Stokesbury:2002:NME

[536] Kevin D. E. Stokesbury, Jay Kirsch, E. Vincent Patrick, and Brenda L. Norcross. Natural mortality estimates of juvenile Pacific herring (Clupea pallasi) in Prince William Sound, Alaska. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 59(3):416–423, March 2002. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f02-017.

Pilling:2002:IME

[537] Graham M. Pilling, Geoffrey P. Kirkwood, and Stephen G. Walker. An improved method for estimating individual growth variability in fish, and the correlation between von Bertalanffy growth parameters. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 59(3):424–432, March 2002. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f02-022.

Schnute:2002:ESS

[538] J. T. Schnute and A. R. Kronlund. Estimating salmon stock–recruitment relationships from catch and escapement data. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 59(3):433–449, March 2002. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f02-016.

Campana:2002:BDA

[539] Steven E. Campana, Lisa J. Natanson, and Sigmund Myklevoll. Bomb dating and age determination of large pelagic sharks. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 59(3):450–455, March 2002. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f02-027.

Mueter:2002:OEO

[540] Franz J. Mueter, Randall M. Peterman, and Brian J. Pyper. Opposite effects of ocean temperature on survival rates of 120 stocks of Pacific salmon (Oncorhynchus spp.) in northern and southern areas. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 59(3):456–463, March 2002. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f02-020.

Weise:2002:LBP

[541] Andréa M. Weise, Maurice Levasseur, François J. Saucier, Simon Senneville, Esther Bonneau, Suzanne Roy, Gilbert Sauvé, Sonia Michaud, and Juliette Fauchot. The link between precipitation, river runoff, and blooms of the toxic dinoflagellate Alexandrium tamarense in the St. Lawrence. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 59(3):464–473, March 2002. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f02-024.

Kelly:2002:ASU

[542] David J. Kelly and Max L. Bothwell. Avoidance of solar ultraviolet radiation by juvenile Coho salmon (Oncorhynchus kisutch). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 59(3):474–482, March 2002. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f02-023.

Vinebrooke:2002:WLA

[543] Rolf D. Vinebrooke, Sushil S. Dixit, Mark D. Graham, John M. Gunn, Yu-Wei Chen, and Nelson Belzile. Whole-lake algal responses to a century of acidic industrial deposition on the Canadian Shield. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 59(3):483–493, March 2002. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f02-025.

Bjornsson:2002:FUG

[544] Björn Björnsson and Agnar Steinarsson. The food-unlimited growth rate of Atlantic cod (Gadus morhua). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 59(3):494–502, March 2002. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f02-028.

Piccolo:2002:DRA

[545] Jack J. Piccolo and Mark S. Wipfli. Does red alder (Alnus rubra) in upland riparian forests elevate macroinvertebrate and detritus export from headwater streams to downstream habitats in southeastern Alaska? Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 59(3):503–513, March 2002. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f02-019.

Davenport:2002:TSM

[546] Stephanie R. Davenport and Nicholas J. Bax. A trophic study of a marine ecosystem off southeastern Australia using stable isotopes of carbon and nitrogen. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 59(3):514–530, March 2002. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f02-031.

Rahikainen:2002:FDH

[547] Mika Rahikainen and Sakari Kuikka. Fleet dynamics of herring trawlers — change in gear size and implications for interpretation of catch per unit effort. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 59(3):531–541, March 2002. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f02-030.

Wilson:2002:PHR

[548] Ben Wilson and Lawrence M. Dill. Pacific herring respond to simulated odontocete echolocation sounds. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 59(3):542–553, March 2002. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f02-029.

Strauss:2002:FRN

[549] Eric A. Strauss, Nicole L. Mitchell, and Gary A. Lamberti. Factors regulating nitrification in aquatic sediments: effects of organic carbon, nitrogen availability, and pH. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 59 (3):554–563, March 2002. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f02-032.

Snyder:2002:RPP

[550] Eric B. Snyder, Christopher T. Robinson, G. Wayne Minshall, and Samuel R. Rushforth. Regional patterns in periphyton accrual and diatom assemblage structure in a heterogeneous nutrient landscape. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 59(3):564–577, March 2002. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f02-021.

Guyette:2002:PAD

[551] Richard P. Guyette, William G. Cole, Daniel C. Dey, and Rose-Marie Muzika. Perspectives on the age and distribution of large wood in riparian carbon pools. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 59 (3):578–585, March 2002. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f02-026.

Weber:2002:OSI

[552] Peter K. Weber, Ian D. Hutcheon, Kevin D. McKeegan, and B. Lynn Ingram. Otolith sulfur isotope method to reconstruct salmon (Oncorhynchus tshawytscha) life history. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 59(4):587–591, April 2002. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f02-038.

Waluda:2002:TLS

[553] C. M. Waluda, P. N. Trathan, C. D. Elvidge, V. R. Hobson, and P. G. Rodhouse. Throwing light on straddling stocks of Illex argentinus: assessing fishing intensity with satellite imagery. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 59(4):592–596, April 2002. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f02-049.

Myers:2002:ETN

[554] Ransom A. Myers and Terrance J. Quinn II. Estimating and testing non-additivity in fishing mortality: implications for detecting a fisheries collapse. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 59(4):597–601, April 2002. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f02-048.

Montevecchi:2002:PMP

[555] W. A. Montevecchi, D. K. Cairns, and R. A. Myers. Predation on marinephase Atlantic salmon (Salmo salar) by gannets (Morus bassanus) in the Northwest Atlantic. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 59 (4):602–612, April 2002. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f02-033.

Pess:2002:LCL

[556] George R. Pess, David R. Montgomery, E. Ashley Steel, Robert E. Bilby, Blake E. Feist, and Harvey M. Greenberg. Landscape characteristics, land use, and Coho salmon (Oncorhynchus kisutch) abundance, Snohomish River, Wash., U.S.A. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 59 (4):613–623, April 2002. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f02-035.

Heupel:2002:EMJ

[557] M. R. Heupel and C. A. Simpfendorfer. Estimation of mortality of juvenile blacktip sharks, Carcharhinus limbatus, within a nursery area using telemetry data. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 59(4):624–632, April 2002. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/f02-
036.

Crowder:2002:VCS

[558] David W. Crowder and Panayiotis Diplas. Vorticity and circulation: spatial metrics for evaluating flow complexity in stream habitats. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 59(4):633–645, April 2002. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f02-037.

Ringuette:2002:AMS

[559] Marc Ringuette, Martin Castonguay, Jeffrey A. Runge, and François Grégoire. Atlantic mackerel (Scomber scombrus) recruitment fluctuations in relation to copepod production and juvenile growth. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 59(4):646–656, April 2002. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f02-039.

Porch:2002:EAG

[560] Clay E. Porch, Mark R. Fisher, and Lawrence W. McEachron. Estimating abundance from gillnet samples with application to red drum (Sciaenops ocellatus) in Texas bays. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 59(4):657–668, April 2002. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f02-034.

Gillanders:2002:TSV

[561] Bronwyn M. Gillanders. Temporal and spatial variability in elemental composition of otoliths: implications for determining stock identity and connectivity of populations. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 59 (4):669–679, April 2002. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f02-040.

Baldwin:2002:CGF

[562] Brad S. Baldwin, Marilyn S. Mayer, Jeffrey Dayton, Nancy Pau, Johanna Mendilla, Maura Sullivan, Aaron Moore, Aye Ma, and Edward L. Mills. Comparative growth and feeding in zebra and quagga mussels (Dreissena polymorpha and Dreissena bugensis): implications for North American lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 59(4):680–694, April 2002. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f02-043.

Wyatt:2002:ERF

[563] Robin J. Wyatt. Estimating riverine fish population size from singleand multiple-pass removal sampling using a hierarchical model. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 59(4):695–706, April 2002. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f02-041.

Koski:2002:FRK

[564] Marci L. Koski and Brett M. Johnson. Functional response of Kokanee salmon (Oncorhynchus nerka) to Daphnia at different light levels. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 59(4):707–716, April 2002. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f02-045.

Paulsen:2002:EMS

[565] Charles M. Paulsen and Richard A. Hinrichsen. Experimental management for Snake River spring–summer Chinook (Oncorhynchus tshawytscha): trade-offs between conservation and learning for a threatened species. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 59(4):717–725, April 2002. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f02-053.

Keller:2002:CZC

[566] W. Keller, N. D. Yan, K. M. Somers, and J. H. Heneberry. Crustacean zooplankton communities in lakes recovering from acidification. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 59(4):726–735, April 2002. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f02-042.

Madenjian:2002:DLM

[567] Charles P. Madenjian, Gary L. Fahnenstiel, Thomas H. Johengen, Thomas F. Nalepa, Henry A. Vanderploeg, Guy W. Fleischer, Philip J. Schneeberger, Darren M. Benjamin, Emily B. Smith, James R. Bence, Edward S. Rutherford, Dennis S. Lavis, Dale M. Robertson, David J. Jude, and Mark P. Ebener. Dynamics of the Lake Michigan food web, 1970–2000. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 59(4):736–753, April 2002. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f02-044.

Hannesson:2002:EFF

[568] Rögnvaldur Hannesson. The economics of fishing down the food chain. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 59(5):755–758, May 2002. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f02-055.

Grant:2002:ENE

[569] Susan C. H. Grant and William M. Tonn. Effects of nutrient enrichment on recruitment of age-0 fathead minnows (Pimephales promelas): potential impacts of environmental change on the boreal plains. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 59(5):759–767, May 2002. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f02-054.

Campbell:2002:IMC

[570] H. A. Campbell, R. D. Handy, and D. W. Sims. Increased metabolic cost of swimming and consequent alterations to circadian activity in rainbow trout (Oncorhynchus mykiss) exposed to dietary copper. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 59(5):768–777, May 2002. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f02-046.

Allen:2002:RPP

[571] C. D. Allen and R. E. H. Smith. The response of planktonic phosphate uptake and turnover to ultraviolet radiation in Lake Erie. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 59(5):778–786, May 2002. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f02-050.

Hurst:2002:ETS

[572] Thomas P. Hurst and David O. Conover. Effects of temperature and salinity on survival of young-of-the-year Hudson River striped bass (Morone saxatilis): implications for optimal overwintering habitats. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 59(5):787–795, May 2002. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f02-051.

Wing:2002:RCC

[573] Michael G. Wing and Arne Skaugset. Relationships of channel characteristics, land ownership, and land use patterns to large woody debris in western Oregon streams. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 59 (5):796–807, May 2002. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f02-052.

Vahatalo:2002:BCK

[574] Anssi V. Vähätalo, Kalevi Salonen, Eija Saski, and Mirja S. Salkinoja-Salonen. Bleaching of color of kraft pulp mill effluents and natural organic matter in lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 59(5):808–818, May 2002. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/f02-
047.

deBruyn:2002:ABS

[575] Adrian M. H. deBruyn, David J. Marcogliese, and Joseph B. Rasmussen. Altered body size distributions in a large river fish community enriched by sewage. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 59(5):819–828, May 2002. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f02-056.

Ulrich:2002:ITF

[576] Clara Ulrich, Sean Pascoe, Per J. Sparre, Jan-Willem De Wilde, and Paul Marchal. Influence of trends in fishing power on bioeconomics in the North Sea flatfish fishery regulated by catches or by effort quotas. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 59(5):829–843, May 2002. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f02-057.

Sparrevohn:2002:DFS

[577] Claus R. Sparrevohn, Anders Nielsen, and Josianne G. Støttrup. Diffusion of fish from a single release point. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 59(5):844–853, May 2002. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f02-059.

Kehler:2002:MEB

[578] Daniel G. Kehler, Ransom A. Myers, and Christopher A. Field. Measurement error and bias in the maximum reproductive rate for the Ricker model. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 59(5):854–864, May 2002. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f02-060.

Dodds:2002:NPR

[579] Walter K. Dodds, Val H. Smith, and Kirk Lohman. Nitrogen and phosphorus relationships to benthic algal biomass in temperate streams. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 59(5):865–874, May 2002. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f02-063.

Brooks:2002:URV

[580] Elizabeth N. Brooks. Using reproductive values to define optimal harvesting for multisite density-dependent populations: example with a marine reserve. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 59(5):875–885, May 2002. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f02-058.

Budge:2002:AWS

[581] Suzanne M. Budge, Sara J. Iverson, W. Don Bowen, and Robert G. Ackman. Among- and within-species variability in fatty acid signatures of marine fish and invertebrates on the Scotian Shelf, Georges Bank, and southern Gulf of St. Lawrence. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 59(5):886–898, May 2002. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f02-062.

Gendron:2002:SID

[582] Louise Gendron and Jean-Claude Brêthes. Simulations of the impact of different temporal and spatial allocations of fishing effort on fishing mortality in a lobster (Homarus americanus) fishery. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 59(5):899–909, May 2002. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f02-068.

Mullon:2002:EIB

[583] Christian Mullon, Philippe Cury, and Pierrick Penven. Evolutionary individual-based model for the recruitment of anchovy (Engraulis capensis) in the southern Benguela. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 59(5):910–922, May 2002. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/

doi/full/10.1139/f02-064.

Weber:2002:EOS

[584] Peter K. Weber, Ian D. Hutcheon, Kevin D. McKeegan, and B. Lynn Ingram. Erratum: Otolith sulfur isotope method to reconstruct salmon (Oncorhynchus tshawytscha) life history. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 59(5):923, May 2002. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f02-901.

Kennedy:2002:RLF

[585] Brian P. Kennedy, Andrea Klaue, Joel D. Blum, Carol L. Folt, and Keith H. Nislow. Reconstructing the lives of fish using Sr isotopes in otoliths. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 59(6):925–929, June 2002. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f02-070.

Carr:2002:SVD

[586] Geneviève M. Carr and Antoine Morin. Sampling variability and the design of bacterial abundance and production studies in aquatic environments. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 59(6):930–937, June 2002. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f02-072.

Philibert:2002:IBS

[587] Aline Philibert and Yves T. Prairie. Is the introduction of benthic species necessary for open-water chemical reconstruction in diatom-based transfer functions? Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 59(6):938–951, June 2002. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/f02-

Eby:2002:HBH

[588] Lisa A. Eby and Larry B. Crowder. Hypoxia-based habitat compression in the Neuse River Estuary: context-dependent shifts in behavioral avoidance thresholds. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 59 (6):952–965, June 2002. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f02-067.

Arkush:2002:RTP

[589] Kristen D. Arkush, Alan R. Giese, Holly L. Mendonca, Anne M. McBride, Gary D. Marty, and Philip W. Hedrick. Resistance to three pathogens in the endangered winter-run Chinook salmon (Oncorhynchus tshawytscha): effects of inbreeding and major histocompatibility complex genotypes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 59(6):966–975, June 2002. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f02-066.

Laslett:2002:FML

[590] Geoff M. Laslett, J. Paige Eveson, and Tom Polacheck. A flexible maximum likelihood approach for fitting growth curves to tag–recapture data. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 59(6):976–986, June 2002. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f02-069.

Sosiak:2002:LTR

[591] Al Sosiak. Long-term response of periphyton and macrophytes to reduced municipal nutrient loading to the Bow River (Alberta, Canada). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 59(6):987–1001, June 2002. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f02-071.

Bertrand:2002:HTC

[592] Arnaud Bertrand, Erwan Josse, Pascal Bach, Philippe Gros, and Laurent Dagorn. Hydrological and trophic characteristics of tuna habitat: consequences on tuna distribution and longline catchability. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 59(6):1002–1013, June 2002. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f02-073.

Hinrichsen:2002:AAS

[593] Richard A. Hinrichsen. The accuracy of alternative stochastic growth rate estimates for salmon populations. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 59(6):1014–1023, June 2002. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f02-065.

Hoyer:2002:NCW

[594] Mark V. Hoyer, Thomas K. Frazer, Sky K. Notestein, and Daniel E. Canfield, Jr. Nutrient, chlorophyll, and water clarity relationships in Florida’s nearshore coastal waters with comparisons to freshwater lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 59(6):1024–1031, June 2002. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f02-077.

vanOostenbrugge:2002:CCV

[595] J. A. E. van Oostenbrugge, E. J. Bakker, W. L. T. van Densen, M. A. M. Machiels, and P. A. M. van Zwieten. Characterizing catch variability in a multispecies fishery: implications for fishery management. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 59(6):1032–1043, June 2002. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f02-078.

Boss:2002:EFC

[596] Shelly M. Boss and John S. Richardson. Effects of food and cover on the growth, survival, and movement of cutthroat trout (Oncorhynchus clarki) in coastal streams. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 59 (6):1044–1053, June 2002. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f02-079.

McGarvey:2002:ERF

[597] Richard McGarvey and John E. Feenstra. Estimating rates of fish movement from tag recoveries: conditioning by recapture. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 59(6):1054–1064, June 2002. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f02-080.

Kloser:2002:SID

[598] R. J. Kloser, T. Ryan, P. Sakov, A. Williams, and J. A. Koslow. Species identification in deep water using multiple acoustic frequencies. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 59(6):1065–1077, June 2002. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f02-076.

Dick:2002:PIF

[599] Jaimie T. A. Dick, Dirk Platvoet, and David W. Kelly. Predatory impact of the freshwater invader Dikerogammarus villosus (Crustacea: Amphipoda). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 59(6):1078–1084, June 2002. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f02-074.

Legendre:2002:RCC

[600] Pierre Legendre, Kari Elsa Ellingsen, Erik Bjørnbom, and Philippe Casgrain. Rapid communication / communication rapide: Acoustic seabed classification: improved statistical method. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 59(7):1085–1089, July 2002. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f02-096.

Frost:2002:RCC

[601] Paul C. Frost and Marguerite A. Xenopoulos. Rapid communication / communication rapide: Ambient solar ultraviolet radiation and its effects on phosphorus flux into boreal lake phytoplankton communities. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 59(7):1090–1095, July 2002. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f02-100.

Waddy:2002:RCC

[602] S. L. Waddy, L. E. Burridge, M. N. Hamilton, S. M. Mercer, D. E. Aiken, and K. Haya. Rapid communication / communication rapide: Emamectin benzoate induces molting in American lobster, Homarus americanus. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 59(7):1096– 1099, July 2002. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f02-106.

Weitkamp:2002:CSO

[603] Laurie Weitkamp and Kathleen Neely. Coho salmon (Oncorhynchus kisutch) ocean migration patterns: insight from marine coded-wire tag recoveries. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 59(7):1100–1115, July 2002. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/f02-
075.

Korman:2002:EOE

[604] Josh Korman, Robert N. M. Ahrens, Paul S. Higgins, and Carl J. Walters. Effects of observer efficiency, arrival timing, and survey life on estimates of escapement for steelhead trout (Oncorhynchus mykiss) derived from repeat mark–recapture experiments. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 59(7):1116–1131, July 2002. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f02-081.

Sheehy:2002:YCD

[605] M. R. J. Sheehy and R. C. A. Bannister. Year-class detection reveals climatic modulation of settlement strength in the European lobster, Homarus gammarus. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 59 (7):1132–1143, July 2002. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f02-083.

Reid:2002:GEU

[606] David F. Reid and Marina I. Orlova. Geological and evolutionary underpinnings for the success of Ponto-Caspian species invasions in the Baltic Sea and North American Great Lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 59(7):1144–1158, July 2002. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f02-099.

BijdeVaate:2002:GPR

[607] A. Bij de Vaate, K. Jazdzewski, H. A. M. Ketelaars, S. Gollasch, and G. Van der Velde. Geographical patterns in range extension of Ponto-Caspian macroinvertebrate species in Europe. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 59(7):1159–1174, July 2002. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f02-098.

Leppakoski:2002:BSI

[608] Erkki Leppäkoski, Stephan Gollasch, Piotr Gruszka, Henn Ojaveer, Sergej Olenin, and Vadim Panov. The Baltic — a sea of invaders. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 59(7):1175–1188, July 2002. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f02-089.

Grigorovich:2002:PMA

[609] Igor A. Grigorovich, Hugh J. MacIsaac, Nikolai V. Shadrin, and Edward L. Mills. Patterns and mechanisms of aquatic invertebrate introductions in the Ponto-Caspian region. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 59(7):1189–1208, July 2002. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f02-088.

Vanderploeg:2002:DEE

[610] Henry A. Vanderploeg, Thomas F. Nalepa, David J. Jude, Edward L. Mills, Kristen T. Holeck, James R. Liebig, Igor A. Grigorovich, and Henn Ojaveer. Dispersal and emerging ecological impacts of Ponto-Caspian species in the Laurentian Great Lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 59(7):1209–1228, July 2002. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f02-087.

Hebert:2002:GPI

[611] Paul D. N. Hebert and Melania E. A. Cristescu. Genetic perspectives on invasions: the case of the cladocera. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 59(7):1229–1234, July 2002. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f02-091.

McMahon:2002:EPA

[612] Robert F. McMahon. Evolutionary and physiological adaptations of aquatic invasive animals: r selection versus resistance. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 59(7):1235–1244, July 2002. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f02-105.

MacIsaac:2002:MSB

[613] Hugh J. MacIsaac, Thomas C. Robbins, and Mark A. Lewis. Modeling ships’ ballast water as invasion threats to the Great Lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 59(7):1245–1256, July 2002. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f02-090.

Chaloner:2002:MCN

[614] Dominic T. Chaloner, Kristine M. Martin, Mark S. Wipfli, Peggy H. Ostrom, and Gary A. Lamberti. Marine carbon and nitrogen in southeastern Alaska stream food webs: evidence from artificial and natural streams. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 59(8):1257–1265, August 2002. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/f02-
084.

Borsuk:2002:SME

[615] Mark E. Borsuk, Sean P. Powers, and Charles H. Peterson. A survival model of the effects of bottom-water hypoxia on the population density of an estuarine clam (Macoma balthica). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 59(8):1266–1274, August 2002. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f02-093.

Gillis:2002:RIU

[616] D. M. Gillis and M. A. Showell. Risk and information use in two competing fleets: Russian and Cuban exploitation of silver hake (Merluccius bilinearis). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 59(8):1275–1286, August 2002. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f02-086.

Henriksen:2002:CLA

[617] A. Henriksen, P. J. Dillon, and J. Aherne. Critical loads of acidity for surface waters in south-central Ontario, Canada: regional application of the Steady-State Water Chemistry (SSWC) model. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 59(8):1287–1295, August 2002. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f02-092.

Rajotte:2002:EEM

[618] James W. Rajotte and Patrice Couture. Effects of environmental metal contamination on the condition, swimming performance, and tissue metabolic capacities of wild yellow perch (Perca flavescens). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 59(8):1296–1304, August 2002. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f02-095.

Ryer:2002:CFE

[619] Clifford H. Ryer, Angela Lawton, Ricardo J. Lopez, and Bori L. Olla. A comparison of the functional ecology of visual vs. nonvisual foraging in two planktivorous marine fishes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 59(8):1305–1314, August 2002. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f02-097.

Buchwalter:2002:RSM

[620] D. B. Buchwalter, J. J. Jenkins, and L. R. Curtis. Respiratory strategy is a major determinant of [^{3}H]water and [^{14}C]chlorpyrifos uptake in aquatic insects. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 59(8):1315–1322, August 2002. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/f02-
107.

Fleming:2002:EDG

[621] Ian A. Fleming, Thorleifur Agustsson, Bengt Finstad, Jörgen I. Johnsson, and Björn Thrandur Björnsson. Effects of domestication on growth physiology and endocrinology of Atlantic salmon (Salmo salar). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 59(8):1323–1330, August 2002. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f02-082.

Brown:2002:ERP

[622] Grant E. Brown, James C. Adrian, Jr., Michael G. Lewis, and Jon M. Tower. The effects of reduced pH on chemical alarm signalling in ostariophysan fishes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 59(8): 1331–1338, August 2002. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f02-104.

Squires:2002:WTN

[623] Margaret M. Squires and Lance F. W. Lesack. Water transparency and nutrients as controls on phytoplankton along a flood-frequency gradient among lakes of the Mackenzie Delta, western Canadian Arctic. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 59(8):1339–1349, August 2002. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f02-085.

Knouft:2002:RAB

[624] Jason H. Knouft. Regional analysis of body size and population density in stream fish assemblages: testing predictions of the energetic equivalence rule. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 59(8):1350–1360, August 2002. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/f02-
110.

Mazumder:2002:IFS

[625] Asit Mazumder and Jim A. Edmundson. Impact of fertilization and stocking on trophic interactions and growth of juvenile sockeye salmon (Oncorhynchus nerka). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 59(8): 1361–1373, August 2002. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f02-111.

Buonaccorsi:2002:PSC

[626] Vincent P. Buonaccorsi, Carol A. Kimbrell, Eric A. Lynn, and Russell D. Vetter. Population structure of copper rockfish (Sebastes caurinus) reflects postglacial colonization and contemporary patterns of larval dispersal. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 59(8):1374–1384, August 2002. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/f02-

Skalski:2002:ERS

[627] John R. Skalski, Richard Townsend, James Lady, Albert E. Giorgi, John R. Stevenson, and Robert D. McDonald. Estimating route-specific passage and survival probabilities at a hydroelectric project from smolt radiotelemetry studies. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 59(8): 1385–1393, August 2002. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f02-094.

Chen:2002:SSE

[628] Yong Chen and Carl Wilson. A simulation study to evaluate impacts of uncertainty on the assessment of American lobster fishery in the Gulf of Maine. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 59(8):1394–1403, August 2002. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/f02-
102.

Wright:2002:CPE

[629] Kristopher K. Wright and Judith L. Li. From continua to patches: examining stream community structure over large environmental gradients. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 59(8):1404–1417, August 2002. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f02-113.

Hood:2002:LAT

[630] W. Gregory Hood. Landscape allometry: from tidal channel hydraulic geometry to benthic ecology. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 59 (8):1418–1427, August 2002. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f02-109.

Link:2002:MEA

[631] Jason S. Link, Jon K. T. Brodziak, Steve F. Edwards, William J. Overholtz, David Mountain, Jack W. Jossi, Tim D. Smith, and Michael J. Fogarty. Marine ecosystem assessment in a fisheries management context. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 59(9):1429–1440, September 2002. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f02-115.

Baber:2002:CFD

[632] Matthew J. Baber, Daniel L. Childers, Kimberly J. Babbitt, and David H. Anderson. Controls on fish distribution and abundance in temporary wetlands. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 59(9): 1441–1450, September 2002. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f02-116.

Chouinard:2002:DDV

[633] G. A. Chouinard and D. P. Swain. Depth-dependent variation in condition and length-at-age of Atlantic cod (Gadus morhua) in the southern Gulf of St. Lawrence. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 59(9): 1451–1459, September 2002. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f02-103.

Bouza:2002:AMD

[634] C. Bouza, P. Presa, J. Castro, L. Sánchez, and P. Martínez. Allozyme and microsatellite diversity in natural and domestic populations of turbot (Scophthalmus maximus) in comparison with other pleuronectiformes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 59(9):1460–1473, September 2002. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f02-114.

Shertzer:2002:LMS

[635] Kyle W. Shertzer and Michael H. Prager. Least median of squares: a suitable objective function for stock assessment models? Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 59(9):1474–1481, September 2002. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f02-112.

Jack:2002:APT

[636] Jeffrey Jack, Tim Sellers, and Paul A. Bukaveckas. Algal production and trihalomethane formation potential: an experimental assessment and inter-river comparison. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 59(9): 1482–1491, September 2002. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f02-121.

Millar:2002:RPB

[637] Russell B. Millar. Reference priors for Bayesian fisheries models. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 59(9):1492–1502, September 2002. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f02-108.

Chen:2002:RMM

[638] D. G. Chen and L. Blair Holtby. A regional meta-model for stock– recruitment analysis using an empirical Bayesian approach. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 59(9):1503–1514, September 2002. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f02-118.

Harwood:2002:IIS

[639] A. J. Harwood, N. B. Metcalfe, S. W. Griffiths, and J. D. Armstrong. Intra- and inter-specific competition for winter concealment habitat in juvenile salmonids. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 59(9):1515– 1523, September 2002. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f02-119.

Schaalje:2002:UNH

[640] G. Bruce Schaalje, Jared L. Shaw, and Mark C. Belk. Using nonlinear hierarchical models for analyzing annulus-based size-at-age data. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 59(9):1524–1532, September 2002. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f02-120.

Williams:2002:CEN

[641] Erik H. Williams and Michael H. Prager. Comparison of equilibrium and nonequilibrium estimators for the generalized production model. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 59(9):1533–1552, September 2002. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f02-123.

Lund:2002:EEH

[642] Susan G. Lund, Daniel Caissie, Richard A. Cunjak, Mathilakath M. Vijayan, and Bruce L. Tufts. The effects of environmental heat stress on heat-shock mRNA and protein expression in Miramichi Atlantic salmon (Salmo salar) parr. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 59(9): 1553–1562, September 2002. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f02-117.

Dangles:2002:FPB

[643] Olivier Dangles. Functional plasticity of benthic macroinvertebrates: implications for trophic dynamics in acid streams. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 59(9):1563–1573, September 2002. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f02-122.

Aubin-Horth:2002:IDE

[644] Nadia Aubin-Horth and Julian J. Dodson. Impact of differential energy allocation in Atlantic salmon (Salmo salar) precocious males on otolith–somatic size proportionality: a longitudinal approach. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 59(10):1575–1583, October 2002. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f02-124.

Montagna:2002:DBC

[645] Paul A. Montagna, Stephen C. Jarvis, and Mahlon C. Kennicutt II. Distinguishing between contaminant and reef effects on meiofauna near offshore hydrocarbon platforms in the Gulf of Mexico. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 59(10):1584–1592, October 2002. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f02-131.

Pazzia:2002:IFW

[646] Ivano Pazzia, Marc Trudel, Mark Ridgway, and Joseph B. Rasmussen. Influence of food web structure on the growth and bioenergetics of lake trout (Salvelinus namaycush). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 59(10):1593–1605, October 2002. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f02-128.

Kainz:2002:MMZ

[647] Martin Kainz, Marc Lucotte, and Christopher C. Parrish. Methyl mercury in zooplankton — the role of size, habitat, and food quality. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 59(10):1606–1615, October 2002. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f02-125.

Cartes:2002:CPB

[648] J. E. Cartes, T. Brey, J. C. Sorbe, and F. Maynou. Comparing production–biomass ratios of benthos and suprabenthos in macrofaunal marine crustaceans. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 59(10): 1616–1625, October 2002. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f02-130.

Kline:2002:PSO

[649] Thomas C. Kline, Jr. and T. Mark Willette. Pacific salmon (Oncorhynchus spp.) early marine feeding patterns based on ^{15}N/^{14}N and ^{13}C/^{12}C in Prince William Sound, Alaska. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 59(10):1626–1638, October 2002. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f02-126.

Barbiero:2002:ERP

[650] Richard P. Barbiero, Marc L. Tuchman, Glenn J. Warren, and David C. Rockwell. Evidence of recovery from phosphorus enrichment in Lake Michigan. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 59(10):1639–1647, October 2002. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f02-132.

Su:2002:OSM

[651] Zhenming Su and Milo D. Adkison. Optimal in-season management of pink salmon (Oncorhynchus gorbuscha) given uncertain run sizes and seasonal changes in economic value. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 59(10):1648–1659, October 2002. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f02-133.

Watras:2002:SED

[652] C. J. Watras, K. A. Morrison, and T. K. Kratz. Seasonal enrichment and depletion of Hg and SO_{4} in Little Rock Lake: relationship to seasonal changes in atmospheric deposition. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 59(10):1660–1667, October 2002. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f02-134.

McArthur:2002:MUD

[653] Michael D. McArthur and John S. Richardson. Microbial utilization of dissolved organic carbon leached from riparian litterfall. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 59(10):1668–1676, October 2002. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f02-135.

Kraft:2002:ISI

[654] Clifford E. Kraft, Rebecca L. Schneider, and Dana R. Warren. Ice storm impacts on woody debris and debris dam formation in northeastern U.S. streams. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 59(10):1677–1684, October 2002. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/f02-
129.

Wroblewski:2002:ERP

[655] J. S. Wroblewski and H. W. Hiscock. Enhancing the reproductive potential of local populations of coastal Atlantic cod (Gadus morhua). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 59(10):1685–1695, October 2002. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f02-136.

Wilson:2002:MPA

[656] A. J. Wilson and M. M. Ferguson. Molecular pedigree analysis in natural populations of fishes: approaches, applications, and practical considerations. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 59(10):1696–1707, October 2002. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/f02-

Bearlin:2002:IWL

[657] Andrew R. Bearlin, E. S. G. Schreiber, Simon J. Nicol, A. M. Starfield, and Charles R. Todd. Identifying the weakest link: simulating adaptive management of the reintroduction of a threatened fish. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 59(11):1709–1716, November 2002. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f02-140.

Koseki:2002:DEA

[658] Yusuke Koseki and Koji Maekawa. Differential energy allocation of alternative male tactics in masu salmon (Oncorhynchus masou). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 59(11):1717–1723, November 2002. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f02-143.

Cox:2002:REDa

[659] Sean P. Cox, Steven J. D. Martell, Carl J. Walters, Timothy E. Essington, James F. Kitchell, Christofer Boggs, and Isaac Kaplan. Reconstructing ecosystem dynamics in the central Pacific Ocean, 1952–1998. I. Estimating population biomass and recruitment of tunas and billfishes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 59(11):1724–1735, November 2002. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f02-137.

Cox:2002:REDb

[660] Sean P. Cox, Timothy E. Essington, James F. Kitchell, Steven J. D. Martell, Carl J. Walters, Christofer Boggs, and Isaac Kaplan. Reconstructing ecosystem dynamics in the central Pacific Ocean, 1952–1998. II. A preliminary assessment of the trophic impacts of fishing and effects on tuna dynamics. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 59(11): 1736–1747, November 2002. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f02-138.

Mantyniemi:2002:BMR

[661] Samu Mäntyniemi and Atso Romakkaniemi. Bayesian mark–recapture estimation with an application to a salmonid smolt population. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 59(11):1748–1758, November 2002. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f02-146.

Munk:2002:LSL

[662] Peter Munk. Larval sand lance (Ammodytes sp.) in the diet of small juvenile wolffish (Anarhichas spp.): predatory interactions in frontal water masses off western Greenland. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 59(11):1759–1767, November 2002. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f02-147.

Rivot:2002:HBA

[663] E. Rivot and E. Prévost. Hierarchical Bayesian analysis of capture– mark–recapture data. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 59(11): 1768–1784, November 2002. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f02-145.

Duplisea:2002:SBM

[664] Daniel E. Duplisea, Simon Jennings, Karema J. Warr, and Tracy A. Dinmore. A size-based model of the impacts of bottom trawling on benthic community structure. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 59(11): 1785–1795, November 2002. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f02-148.

Elsdon:2002:IET

[665] Travis S. Elsdon and Bronwyn M. Gillanders. Interactive effects of temperature and salinity on otolith chemistry: challenges for determining environmental histories of fish. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 59(11):1796–1808, November 2002. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f02-154.

Standen:2002:ECM

[666] Emily M. Standen, Scott G. Hinch, Michael C. Healey, and Anthony P. Farrell. Energetic costs of migration through the Fraser River Canyon, British Columbia, in adult pink (Oncorhynchus gorbuscha) and sockeye (Oncorhynchus nerka) salmon as assessed by EMG telemetry. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 59(11):1809–1818, November 2002. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f02-151.

Robinson:2002:CTS

[667] Beren W. Robinson and Kevin J. Parsons. Changing times, spaces, and faces: tests and implications of adaptive morphological plasticity in the fishes of northern postglacial lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 59(11):1819–1833, November 2002. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f02-144.

Davis:2002:KPU

[668] Michael W. Davis. Key principles for understanding fish bycatch discard mortality. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 59(11):1834– 1843, November 2002. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f02-139.

Anonymous:2002:CSV

[669] Anonymous. Contents / sommaire volume 59, 2002. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 59(12), December 2002. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfascontents02.

Comeau:2002:AML

[670] Luc A. Comeau, Steven E. Campana, and Martin Castonguay. Automated monitoring of a large-scale cod (Gadus morhua) migration in the open sea. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 59(12):1845–1850, December 2002. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f02-152.

Agnew:2002:PUE

[671] D. J. Agnew, J. R. Beddington, and S. L. Hill. The potential use of environmental information to manage squid stocks. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 59(12):1851–1857, December 2002. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f02-150.

Hinrichsen:2002:BML

[672] H.-H. Hinrichsen, C. Möllmann, R. Voss, F. W. Köster, and G. Kornilovs. Biophysical modeling of larval Baltic cod (Gadus morhua) growth and survival. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 59(12):1858–1873, December 2002. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f02-149.

Chang:2002:NSI

[673] Cecily C. Y. Chang, Carol Kendall, Steven R. Silva, William A. Battaglin, and Donald H. Campbell. Nitrate stable isotopes: tools for determining nitrate sources among different land uses in the Mississippi River Basin. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 59(12):1874– 1885, December 2002. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f02-153.

Mellina:2002:STR

[674] Eric Mellina, R. Dan Moore, Scott G. Hinch, J. Stevenson Macdonald, and Greg Pearson. Stream temperature responses to clearcut logging in British Columbia: the moderating influences of groundwater and headwater lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 59(12):1886–1900, December 2002. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f02-158.

Wilhelm:2002:MPA

[675] Frank M. Wilhelm, Juliane Hamann, and Carolyn W. Burns. Mysid predation on amphipods and Daphnia in a shallow coastal lake: prey selection and effects of macrophytes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 59(12):1901–1907, December 2002. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f02-161.

Kraus:2002:EPB

[676] Gerd Kraus, Jonna Tomkiewicz, and Friedrich W. Köster. Egg production of Baltic cod (Gadus morhua) in relation to variable sex ratio, maturity, and fecundity. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 59(12): 1908–1920, December 2002. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f02-159.

Bez:2002:GFA

[677] Nicolas Bez. Global fish abundance estimation from regular sampling: the geostatistical transitive method. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 59(12):1921–1931, December 2002. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f02-155.

Sainte-Marie:2002:IVS

[678] Bernard Sainte-Marie, Jean-Marie Sévigny, and Mireille Carpentier. Interannual variability of sperm reserves and fecundity of primiparous females of the snow crab (Chionoecetes opilio) in relation to sex ratio. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 59(12):1932–1940, December 2002. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f02-162.

Jurado-Molina:2002:CFE

[679] Jesús Jurado-Molina and Patricia Livingston. Climate-forcing effects on trophically linked groundfish populations: implications for fisheries management. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 59(12):1941–1951, December 2002. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f02-160.

Bisbal:2002:BAS

[680] Gustavo A. Bisbal. The best available science for the management of anadromous salmonids in the Columbia River Basin. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 59(12):1952–1959, December 2002. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f02-157.

Hay:2002:TAA

[681] D. E. Hay and S. M. McKinnell. Tagging along: association among individual Pacific herring (Clupea pallasi) revealed by tagging. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 59(12):1960–1968, December 2002. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f02-141.

Benson:2002:CPH

[682] Ashleen J. Benson, Gordon A. McFarlane, Susan E. Allen, and John F. Dower. Changes in Pacific hake (Merluccius productus) migration patterns and juvenile growth related to the 1989 regime shift. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 59(12):1969–1979, December 2002. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f02-156.

Rodriguez-Sanchez:2002:DGS

[683] Rubén Rodríguez-Sánchez, Daniel Lluch-Belda, Héctor Villalobos, and Sofia Ortega-García. Dynamic geography of small pelagic fish populations in the California Current system on the regime time scale (1931– 1997). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 59(12):1980–1988, December 2002. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f02-142.

Tsukamoto:2002:MSE

[684] K. Tsukamoto, J. Aoyama, and M. J. Miller. Migration, speciation, and the evolution of diadromy in anguillid eels. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 59(12):1989–1998, December 2002. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f02-165.

Anonymous:2002:AII

[685] Anonymous. Author index / index des auteurs volume 59, 2002. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 59(12), December 2002. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfasauthor02.

Beamish:2003:MNW

[686] R. J. Beamish, D. J. Noakes, D. L. G. Noakes, and F. W. H. Beamish. In memoriam / nécrologie: William Edwin Ricker, OC, RFSC, LLD, DSc. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 60(1), January 2003. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f03-900.

Unwin:2003:GCS

[687] Martin J. Unwin, Michael T. Kinnison, Nelson C. Boustead, and Thomas P. Quinn. Genetic control over survival in Pacific salmon (Oncorhynchus spp.): experimental evidence between and within populations of New Zealand Chinook salmon (O. tshawytscha). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 60(1):1–11, January 2003. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f02-167.

Ballantyne:2003:IDP

[688] Ashley P. Ballantyne, Michael T. Brett, and Daniel E. Schindler. The importance of dietary phosphorus and highly unsaturated fatty acids for sockeye (Oncorhynchus nerka) growth in Lake Washington — a bioenergetics approach. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 60(1):12–22, January 2003. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f02-166.

Siefkes:2003:CMS

[689] Michael J. Siefkes, Roger A. Bergstedt, Michael B. Twohey, and Weiming Li. Chemosterilization of male sea lampreys (Petromyzon marinus) does not affect sex pheromone release. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 60(1):23–31, January 2003. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f02-169.

Henderson:2003:PSA

[690] J. Nathan Henderson and Benjamin H. Letcher. Predation on stocked Atlantic salmon (Salmo salar) fry. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 60(1):32–42, January 2003. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f03-001.

Nicholls:2003:CPC

[691] K. H. Nicholls, R. J. Steedman, and E. C. Carney. Changes in phytoplankton communities following logging in the drainage basins of three boreal forest lakes in northwestern Ontario (Canada), 1991–2000. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 60(1):43–54, January 2003. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f03-002.

Szalai:2003:MTV

[692] Emily B. Szalai, Guy W. Fleischer, and James R. Bence. Modeling time-varying growth using a generalized von Bertalanffy model with application to bloater (Coregonus hoyi) growth dynamics in Lake Michigan. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 60(1):55–66, January 2003. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f03-003.

Trenkel:2003:PID

[693] Verena M. Trenkel and Marie-Joëlle Rochet. Performance of indicators derived from abundance estimates for detecting the impact of fishing on a fish community. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 60 (1):67–85, January 2003. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f02-163.

Rochet:2003:WCI

[694] Marie-Joëlle Rochet and Verena M. Trenkel. Which community indicators can measure the impact of fishing? A review and proposals. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 60(1):86–99, January 2003. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f02-164.

Minns:2003:ANC

[695] Charles K. Minns and James E. Moore. Assessment of net change of productive capacity of fish habitats: the role of uncertainty and complexity in decision making. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 60 (1):100–116, January 2003. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f02-168.

Rosati:2003:CFS

[696] Ted C. Rosati, Jeffrey R. Johansen, and Miles M. Coburn. Cyprinid fishes as samplers of benthic diatom communities in freshwater streams of varying water quality. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 60 (2):117–125, February 2003. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f03-004.

Stoeckmann:2003:PEL

[697] Ann Stoeckmann. Physiological energetics of Lake Erie dreissenid mussels: a basis for the displacement of Dreissena polymorpha by Dreissena bugensis. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 60(2):126–134, February 2003. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f03-005.

Mitro:2003:RBA

[698] Matthew G. Mitro, Alexander V. Zale, and Bruce A. Rich. The relation between age-0 rainbow trout (Oncorhynchus mykiss) abundance and winter discharge in a regulated river. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 60(2):135–139, February 2003. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f03-006.

Floeter:2003:EDC

[699] Jens Floeter and Axel Temming. Explaining diet composition of North Sea cod (Gadus morhua): prey size preference vs. prey availability. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 60(2):140–150, February 2003. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f03-007.

Holt:2003:PTU

[700] C. A. Holt, N. D. Yan, and K. M. Somers. pH 6 as the threshold to use in critical load modeling for zooplankton community change with acidification in lakes of south-central Ontario: accounting for morphometry and geography. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 60(2):

151–158, February 2003. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f03-008.

Holland:2003:OEC

[701] Daniel S. Holland and Jean-Jacques Maguire. Optimal effort controls for the multispecies groundfish complex in New England: what might have been. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 60(2):159–170, February 2003. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f03-010.

Rodwell:2003:IHQ

[702] Lynda D. Rodwell, Edward B. Barbier, Callum M. Roberts, and Tim R. McClanahan. The importance of habitat quality for marine reserve– fishery linkages. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 60(2):171–181, February 2003. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f03-009.

Parkos:2003:EAC

[703] Joseph J. Parkos III, Victor J. Santucci, Jr., and David H. Wahl. Effects of adult common carp (Cyprinus carpio) on multiple trophic levels in shallow mesocosms. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 60(2): 182–192, February 2003. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f03-011.

Steingrimsson:2003:PCM

[704] Stefán Ó. Steingrímsson and James W. A. Grant. Patterns and correlates of movement and site fidelity in individually tagged young-of-the-year Atlantic salmon (Salmo salar). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 60 (2):193–202, February 2003. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f03-012.

Molot:2003:SMA

[705] L. A. Molot, S. A. Miller, P. J. Dillon, and C. G. Trick. A simple method for assaying extracellular hydroxyl radical activity and its application to natural and synthetic waters. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 60 (2):203–213, February 2003. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f03-014.

Magee:2003:EEA

[706] J. A. Magee, M. Obedzinski, S. D. McCormick, and J. F. Kocik. Effects of episodic acidification on Atlantic salmon (Salmo salar) smolts. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 60(2):214–221, February 2003. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f03-015.

Pinnegar:2003:LTC

[707] J. K. Pinnegar, N. V. C. Polunin, and F. Badalamenti. Long-term changes in the trophic level of western Mediterranean fishery and aquaculture landings. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 60(2):222–235, February 2003. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f03-016.

Grosjean:2003:FGM

[708] Ph. Grosjean, Ch. Spirlet, and M. Jangoux. A functional growth model with intraspecific competition applied to a sea urchin, Paracentrotus lividus. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 60(3):237–246, March 2003. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f03-017.

Negishi:2003:ROM

[709] Junjiro N. Negishi and John S. Richardson. Responses of organic matter and macroinvertebrates to placements of boulder clusters in a small stream of southwestern British Columbia, Canada. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 60(3):247–258, March 2003. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f03-013.

Salthaug:2003:CSD

[710] Are Salthaug and Sondre Aanes. Catchability and the spatial distribution of fishing vessels. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 60 (3):259–268, March 2003. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f03-018.

Audet:2003:SVT

[711] Debbie Audet and Patrice Couture. Seasonal variations in tissue metabolic capacities of yellow perch (Perca flavescens) from clean and metal-contaminated environments. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 60(3):269–278, March 2003. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f03-020.

Raffenberg:2003:IAS

[712] Matthew J. Raffenberg and Donna L. Parrish. Interactions of Atlantic salmon (Salmo salar) and trout (Salvelinus fontinalis and Oncorhynchus mykiss) in Vermont tributaries of the Connecticut River. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 60(3):279–285, March 2003. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f03-021.

Pratt:2003:FHM

[713] Thomas C. Pratt and Karen E. Smokorowski. Fish habitat management implications of the summer habitat use by littoral fishes in a north temperate, mesotrophic lake. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 60 (3):286–300, March 2003. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f03-022.

Nelson:2003:SPB

[714] Jay A. Nelson, Portia S. Gotwalt, and Joel W. Snodgrass. Swimming performance of blacknose dace (Rhinichthys atratulus) mirrors home-stream current velocity. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 60(3):301–308, March 2003. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f03-023.

Allan:2003:ISV

[715] J. David Allan, Mark S. Wipfli, John P. Caouette, Aaron Prussian, and Joanna Rodgers. Influence of streamside vegetation on inputs of terrestrial invertebrates to salmonid food webs. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 60(3):309–320, March 2003. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f03-019.

Blanchard:2003:ECF

[716] Julia L. Blanchard, Kenneth T. Frank, and James E. Simon. Effects of condition on fecundity and total egg production of eastern Scotian Shelf haddock (Melanogrammus aeglefinus). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 60(3):321–332, March 2003. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f03-024.

Squires:2003:RBS

[717] Margaret M. Squires and Lance F. W. Lesack. The relation between sediment nutrient content and macrophyte biomass and community structure along a water transparency gradient among lakes of the Mackenzie Delta. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 60(3):333–343, March 2003. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f03-027.

Gardner:2003:PST

[718] Beth Gardner, Patrick J. Sullivan, and Arthur J. Lembo, Jr. Predicting stream temperatures: geostatistical model comparison using alternative distance metrics. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 60 (3):344–351, March 2003. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f03-025.

Borgmann:2003:DCE

[719] Uwe Borgmann. Derivation of cause-effect based sediment quality guidelines. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 60(3):352–360, March 2003. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f03-026.

Wells:2003:SES

[720] Brian K. Wells, Simon R. Thorrold, and Cynthia M. Jones. Stability of elemental signatures in the scales of spawning weakfish, Cynoscion regalis. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 60(4):361–369, April 2003. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f03-028.

Reynolds:2003:QHA

[721] Julie A. Reynolds. Quantifying habitat associations in marine fisheries: a generalization of the Kolmogorov–Smirnov statistic using commercial logbook records linked to archived environmental data. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 60(4):370–378, April 2003. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f03-032.

Warren:2003:BTS

[722] Dana R. Warren and Clifford E. Kraft. Brook trout (Salvelinus fontinalis) response to wood removal from high-gradient streams of the Adirondack Mountains (N.Y., U.S.A.). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 60(4):379–389, April 2003. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f03-031.

Perry:2003:EDC

[723] Russell W. Perry, Michael J. Bradford, and Jeffrey A. Grout. Effects of disturbance on contribution of energy sources to growth of juvenile Chinook salmon (Oncorhynchus tshawytscha) in boreal streams. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 60(4):390–400, April 2003. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f03-035.

Morinville:2003:EJB

[724] Geneviève R. Morinville and Joseph B. Rasmussen. Early juvenile bioenergetic differences between anadromous and resident brook trout (Salvelinus fontinalis). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 60(4):401–410, April 2003. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/f03-
036.

Cozar:2003:APS

[725] Andrés Cózar, Carlos M. García, and José A. Gálvez. Analysis of plankton size spectra irregularities in two subtropical shallow lakes (Esteros del Iberá, Argentina). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 60 (4):411–420, April 2003. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f03-037.

Takahashi:2003:PST

[726] Hiroshi Takahashi, Tetsuya Tsuruta, and Akira Goto. Population structure of two ecologically distinct forms of ninespine stickleback, Pungitius pungitius: gene flow regimes and genetic diversity based on mtDNA sequence variations. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 60 (4):421–432, April 2003. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f03-034.

McLean:2003:DRS

[727] Jennifer E. McLean, Paul Bentzen, and Thomas P. Quinn. Differential reproductive success of sympatric, naturally spawning hatchery and wild steelhead trout (Oncorhynchus mykiss) through the adult stage. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 60(4):433–440, April 2003. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f03-040.

Francis:2003:POR

[728] R. I. C. Chris Francis. The precision of otolith radiometric ageing of fish and the effect of within-sample heterogeneity. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 60(4):441–447, April 2003. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f03-038.

Fryer:2003:UMM

[729] R. J. Fryer, A. F. Zuur, and N. Graham. Using mixed models to combine smooth size-selection and catch-comparison curves over hauls. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 60(4):448–459, April 2003. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f03-029.

Lund:2003:RBC

[730] Susan G. Lund, Mervyn E. A. Lund, and Bruce L. Tufts. Red blood cell Hsp 70 mRNA and protein as bio-indicators of temperature stress in the brook trout (Salvelinus fontinalis). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 60(4):460–470, April 2003. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f03-039.

Mills:2003:LOF

[731] E. L. Mills, J. M. Casselman, R. Dermott, J. D. Fitzsimons, G. Gal, K. T. Holeck, J. A. Hoyle, O. E. Johannsson, B. F. Lantry, J. C. Makarewicz, E. S. Millard, I. F. Munawar, M. Munawar, R. O’Gorman, R. W. Owens, L. G. Rudstam, T. Schaner, and T. J. Stewart. Lake Ontario: food web dynamics in a changing ecosystem (1970–2000). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 60(4):471–490, April 2003. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f03-033.

Wang:2003:WRR

[732] Lizhu Wang, John Lyons, Paul Rasmussen, Paul Seelbach, Thomas Simon, Michael Wiley, Paul Kanehl, Ed Baker, Scott Niemela, and Paul M. Stewart. Watershed, reach, and riparian influences on stream fish assemblages in the Northern Lakes and Forest Ecoregion, U.S.A. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 60(5):491–505, May 2003. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f03-043.

Jonsson:2003:EAA

[733] Nina Jonsson and Bror Jonsson. Energy allocation among developmental stages, age groups, and types of Atlantic salmon (Salmo salar) spawners. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 60(5):506–516, May 2003. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f03-042.

Tanner:2003:IPT

[734] Jason E. Tanner. The influence of prawn trawling on sessile benthic assemblages in Gulf St. Vincent, South Australia. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 60(5):517–526, May 2003. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f03-044.

Breen:2003:EAC

[735] Paul A. Breen, Ray Hilborn, Mark N. Maunder, and Susan W. Kim. Effects of alternative control rules on the conflict between a fishery and a threatened sea lion (Phocarctos hookeri). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 60(5):527–541, May 2003. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f03-046.

Zuur:2003:DFA

[736] A. F. Zuur, I. D. Tuck, and N. Bailey. Dynamic factor analysis to estimate common trends in fisheries time series. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 60(5):542–552, May 2003. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f03-030.

Quinn:2003:DDP

[737] T. P. Quinn, S. M. Gende, G. T. Ruggerone, and D. E. Rogers. Densitydependent predation by brown bears (Ursus arctos) on sockeye salmon (Oncorhynchus nerka). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 60 (5):553–562, May 2003. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f03-045.

Friedland:2003:MTN

[738] K. D. Friedland, D. G. Reddin, J. R. McMenemy, and K. F. Drinkwater. Multidecadal trends in North American Atlantic salmon (Salmo salar) stocks and climate trends relevant to juvenile survival. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 60(5):563–583, May 2003. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f03-047.

Charette:2003:WIP

[739] T. Charette and E. E. Prepas. Wildfire impacts on phytoplankton communities of three small lakes on the boreal plain, Alberta, Canada: a paleolimnological study. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 60 (5):584–593, May 2003. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f03-049.

Juanes:2003:ACP

[740] Francis Juanes. The allometry of cannibalism in piscivorous fishes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 60(5):594–602, May 2003. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f03-051.

Death:2003:SPL

[741] Russell G. Death. Spatial patterns in lotic invertebrate community composition: is substrate disturbance actually important? Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 60(5):603–611, May 2003. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f03-052.

Moore:2003:DCD

[742] T. R. Moore, L. Matos, and N. T. Roulet. Dynamics and chemistry of dissolved organic carbon in Precambrian Shield catchments and an impounded wetland. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 60 (5):612–623, May 2003. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f03-050.

Chu:2003:CRA

[743] Cindy Chu, Charles K. Minns, and Nicholas E. Mandrak. Comparative regional assessment of factors impacting freshwater fish biodiversity in Canada. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 60(5):624–634, May 2003. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f03-048.

Flanagan:2003:CCP

[744] Kyla M. Flanagan, Edward McCauley, Frederick Wrona, and Terry Prowse. Climate change: the potential for latitudinal effects on algal biomass in aquatic ecosystems. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 60(6):635–639, June 2003. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/

doi/full/10.1139/f03-062.

Schnute:2003:SMD

[745] J. T. Schnute and R. Haigh. A simulation model for designing groundfish trawl surveys. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 60(6):640–656, June 2003. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/f03-
041.

Zimmermann:2003:CUA

[746] Mark Zimmermann. Calculation of untrawlable areas within the boundaries of a bottom trawl survey. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 60(6):657–669, June 2003. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/

doi/full/10.1139/f03-055.

Schrey:2003:MAP

[747] Aaron W. Schrey and Edward J. Heist. Microsatellite analysis of population structure in the shortfin mako (Isurus oxyrinchus). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 60(6):670–675, June 2003. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f03-064.

McNair:2003:CBB

[748] Sheila A. McNair and Patricia Chow-Fraser. Change in biomass of benthic and planktonic algae along a disturbance gradient for 24 Great Lakes coastal wetlands. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 60(6):676–689, June 2003. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/f03-
054.

Berejikian:2003:CVI

[749] B. A. Berejikian, W. T. Fairgrieve, P. Swanson, and E. P. Tezak. Current velocity and injection of GnRHa affect reproductive behavior and body composition of captively reared offspring of wild Chinook salmon (Oncorhynchus tshawytscha). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 60 (6):690–699, June 2003. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f03-056.

Hjelm:2003:FLB

[750] Joakim Hjelm, Gerco H. van de Weerd, and Ferdinand A. Sibbing. Functional link between foraging performance, functional morphology, and diet shift in roach (Rutilus rutilus). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 60(6):700–709, June 2003. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f03-058.

Williams:2003:ILH

[751] Erik H. Williams and Kyle W. Shertzer. Implications of life-history invariants for biological reference points used in fishery management. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 60(6):710–720, June 2003. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f03-059.

Zimmer:2003:RAN

[752] Kyle D. Zimmer, Mark A. Hanson, and Malcolm G. Butler. Relationships among nutrients, phytoplankton, macrophytes, and fish in prairie wetlands. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 60(6):721–730, June 2003. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f03-060.

Melo:2003:RRP

[753] Adriano S. Melo, Dev K. Niyogi, Christoph D. Matthaei, and Colin R. Townsend. Resistance, resilience, and patchiness of invertebrate assemblages in native tussock and pasture streams in New Zealand after a hydrological disturbance. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 60 (6):731–739, June 2003. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f03-061.

Grigorovich:2003:BMA

[754] Igor A. Grigorovich, Robert I. Colautti, Edward L. Mills, Kristen Holeck, Albert G. Ballert, and Hugh J. MacIsaac. Ballast-mediated animal introductions in the Laurentian Great Lakes: retrospective and prospective analyses. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 60(6):740–756, June 2003. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f03-053.

Mueter:2003:COE

[755] Franz J. Mueter, Randall M. Peterman, and Brian J. Pyper. Corrigendum: Opposite effects of ocean temperature on survival rates of 120 stocks of Pacific salmon (Oncorhynchus spp.) in northern and southern areas. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 60(6):757, June 2003. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f03-063.

Benfield:2003:DPS

[756] Mark C. Benfield, Andone C. Lavery, Peter H. Wiebe, Charles H. Greene, Timothy K. Stanton, and Nancy J. Copley. Distributions of physonect siphonulae in the Gulf of Maine and their potential as important sources of acoustic scattering. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 60 (7):759–772, July 2003. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f03-065.

King:2003:FRF

[757] A. J. King, P. Humphries, and P. S. Lake. Fish recruitment on floodplains: the roles of patterns of flooding and life history characteristics. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 60(7):773–786, July 2003. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f03-057.

Dekker:2003:DEE

[758] Willem Dekker. On the distribution of the European eel (Anguilla anguilla) and its fisheries. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 60 (7):787–799, July 2003. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f03-066.

Croteau:2003:DRC

[759] Marie-Noële Croteau, Landis Hare, and André Tessier. Difficulties in relating Cd concentrations in the predatory insect Chaoborus to those of its prey in nature. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 60(7):800–808, July 2003. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/f03-

Peterman:2003:UKF

[760] Randall M. Peterman, Brian J. Pyper, and Brice W. MacGregor. Use of the Kalman filter to reconstruct historical trends in productivity of Bristol Bay sockeye salmon (Oncorhynchus nerka). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 60(7):809–824, July 2003. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f03-069.

Arai:2003:DPC

[761] Mary N. Arai, David W. Welch, Adrienne L. Dunsmuir, Melinda C. Jacobs, and Adrian R. Ladouceur. Digestion of pelagic Ctenophora and Cnidaria by fish. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 60(7):825–829, July 2003. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/f03-
071.

Quist:2003:RDW

[762] Michael C. Quist, Christopher S. Guy, and James L. Stephen. Recruitment dynamics of walleyes (Stizostedion vitreum) in Kansas reservoirs: generalities with natural systems and effects of a centrarchid predator. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 60(7):830–839, July 2003. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f03-067.

Isaak:2003:TVS

[763] Daniel J. Isaak, Russell F. Thurow, Bruce E. Rieman, and Jason B. Dunham. Temporal variation in synchrony among Chinook salmon (Oncorhynchus tshawytscha) redd counts from a wilderness area in central Idaho. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 60(7):840–848, July 2003. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f03-073.

Jones:2003:PCA

[764] Nicholas E. Jones, William M. Tonn, Garry J. Scrimgeour, and Chris Katopodis. Productive capacity of an artificial stream in the Canadian Arctic: assessing the effectiveness of fish habitat compensation. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 60(7):849–863, July 2003. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f03-074.

Hinrichsen:2003:PEE

[765] Richard A. Hinrichsen. The power of experiments for estimating relative reproductive success of hatchery-born spawners. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 60(7):864–872, July 2003. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f03-070.

Etherington:2003:SDL

[766] Lisa L. Etherington and David B. Eggleston. Spatial dynamics of largescale, multistage crab (Callinectes sapidus) dispersal: determinants and consequences for recruitment. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 60 (7):873–887, July 2003. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f03-072.

Kainz:2003:RBO

[767] Martin Kainz, Marc Lucotte, and Christopher C. Parrish. Relationships between organic matter composition and methyl mercury content of offshore and carbon-rich littoral sediments in an oligotrophic lake. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 60(7):888–896, July 2003. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f03-075.

Swain:2003:SDC

[768] Douglas P. Swain and Elmer J. Wade. Spatial distribution of catch and effort in a fishery for snow crab (Chionoecetes opilio): tests of predictions of the ideal free distribution. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 60 (8):897–909, August 2003. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f03-076.

Price:2003:SSE

[769] Carol Seals Price and Carl B. Schreck. Stress and saltwater-entry behavior of juvenile Chinook salmon (Oncorhynchus tshawytscha): conflicts in physiological motivation. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 60 (8):910–918, August 2003. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f03-080.

Declerck:2003:IDD

[770] Steven Declerck, Vanessa Geenens, Nicole Podoor, José Maria Conde Porcuna, and Luc De Meester. Intraspecific density dependence in the dynamics of zooplankton under hypertrophic conditions. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 60(8):919–928, August 2003. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f03-077.

Peck:2003:ELD

[771] Myron A. Peck, Lawrence J. Buckley, and David A. Bengtson. Energy losses due to routine and feeding metabolism in young-of-the-year juvenile Atlantic cod (Gadus morhua). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 60(8):929–937, August 2003. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f03-079.

Garvey:2003:ULT

[772] James E. Garvey and Elizabeth A. Marschall. Understanding latitudinal trends in fish body size through models of optimal seasonal energy allocation. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 60(8):938–948, August 2003. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f03-083.

Matozzo:2003:FRH

[773] Valerio Matozzo, Luisa Da Ros, Loriano Ballarin, Francesca Meneghetti, and Maria Gabriella Marin. Functional responses of haemocytes in the clam Tapes philippinarum from the lagoon of Venice: fishing impact and seasonal variations. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 60(8): 949–958, August 2003. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f03-084.

Dillon:2003:PLO

[774] P. J. Dillon, B. J. Clark, L. A. Molot, and H. E. Evans. Predicting the location of optimal habitat boundaries for lake trout (Salvelinus namaycush) in Canadian Shield lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 60(8):959–970, August 2003. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f03-082.

McCann:2003:DRM

[775] Kevin S. McCann, Louis W. Botsford, and Alan Hasting. Differential response of marine populations to climate forcing. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 60(8):971–985, August 2003. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f03-078.

Castro-Santos:2003:QMD

[776] Theodore Castro-Santos and Alex Haro. Quantifying migratory delay: a new application of survival analysis methods. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 60(8):986–996, August 2003. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f03-086.

Wyatt:2003:MAR

[777] Robin J. Wyatt. Mapping the abundance of riverine fish populations: integrating hierarchical Bayesian models with a geographic information system (GIS). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 60(8):997–1006, August 2003. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f03-085.

Latterell:2003:PCT

[778] Joshua J. Latterell, Robert J. Naiman, Brian R. Fransen, and Peter A. Bisson. Physical constraints on trout (Oncorhynchus spp.) distribution in the Cascade Mountains: a comparison of logged and unlogged streams. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 60(8):1007–1017, August 2003. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f03-088.

Weber:2003:IBH

[779] Edward D. Weber and Kurt D. Fausch. Interactions between hatchery and wild salmonids in streams: differences in biology and evidence for competition. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 60(8):1018– 1036, August 2003. CODEN CJFSDX. ISSN 0706-652X (print), 1205-

7533 (electronic). URL https://cdnsciencepub.com/doi/full/10.
1139/f03-087.

Williams:2003:EIL

[780] Erik H. Williams and Kyle W. Shertzer. Erratum: Implications of lifehistory invariants for biological reference points used in fishery management. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 60(8):1037, August 2003. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f03-099.

Heino:2003:ATC

[781] Jani Heino, Timo Muotka, Riku Paavola, and Lauri Paasivirta. Amongtaxon congruence in biodiversity patterns: can stream insect diversity be predicted using single taxonomic groups? Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 60(9):1039–1049, September 2003. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f03-081.

Nickelson:2003:IHC

[782] Thomas Nickelson. The influence of hatchery Coho salmon (Oncorhynchus kisutch) on the productivity of wild Coho salmon populations in Oregon coastal basins. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 60(9): 1050–1056, September 2003. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f03-091.

Chilcote:2003:RBN

[783] Mark W. Chilcote. Relationship between natural productivity and the frequency of wild fish in mixed spawning populations of wild and hatchery steelhead (Oncorhynchus mykiss). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 60(9):1057–1067, September 2003. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f03-092.

Ruttan:2003:FFG

[784] Lore M. Ruttan. Finding fish: grouping and catch-per-unit-effort in the Pacific hake (Merluccius productus) fishery. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 60(9):1068–1077, September 2003. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f03-096.

Houlahan:2003:EAL

[785] Jeff E. Houlahan and C. Scott Findlay. The effects of adjacent land use on wetland amphibian species richness and community composition. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 60(9):1078–1094, September 2003. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f03-095.

Watmough:2003:PIF

[786] S. A. Watmough, J. Aherne, and P. J. Dillon. Potential impact of forest harvesting on lake chemistry in south-central Ontario at current levels of acid deposition. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 60(9):1095– 1103, September 2003. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f03-093.

Finstad:2003:BSD

[787] Anders Gravbrøt Finstad, Peder Andreas Jansen, and Heikki Hirvonen. Bimodal size distributions in Arctic char, Salvelinus alpinus: artefacts of biased sampling. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 60(9):1104– 1110, September 2003. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f03-089.

McIntyre:2003:SST

[788] Tara M. McIntyre and Jeffrey A. Hutchings. Small-scale temporal and spatial variation in Atlantic cod (Gadus morhua) life history. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 60(9):1111–1121, September 2003. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f03-090.

Paulsen:2003:DJS

[789] Charles M. Paulsen and Timothy R. Fisher. Detecting juvenile survival effects of habitat actions: power analysis applied to endangered Snake River spring–summer Chinook (Oncorhynchus tshawytscha). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 60(9):1122–1132, September 2003. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f03-094.

Taylor:2003:TSD

[790] David L. Taylor and Jeremy S. Collie. A temperature- and size-dependent model of sand shrimp (Crangon septemspinosa) predation on juvenile winter flounder (Pseudopleuronectes americanus). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 60(9):1133–1148, September 2003. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f03-098.

Enders:2003:ETC

[791] Eva C. Enders, Daniel Boisclair, and André G. Roy. The effect of turbulence on the cost of swimming for juvenile Atlantic salmon (Salmo salar). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 60(9):1149–1160, September 2003. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f03-101.

Watters:2003:PFD

[792] George M. Watters, Robert J. Olson, Robert C. Francis, Paul C. Fiedler, Jeffrey J. Polovina, Stephen B. Reilly, Kerim Y. Aydin, Christofer H. Boggs, Timothy E. Essington, Carl J. Walters, and James F. Kitchell. Physical forcing and the dynamics of the pelagic ecosystem in the eastern tropical Pacific: simulations with ENSO-scale and global-warming climate drivers. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 60(9):1161– 1175, September 2003. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f03-100.

Bos:2003:SCR

[793] Darren G. Bos and Brian F. Cumming. Sedimentary cladoceran remains and their relationship to nutrients and other limnological variables in 53 lakes from British Columbia, Canada. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 60(10):1177–1189, October 2003. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f03-097.

Claytor:2003:CRE

[794] Ross Claytor and Jacques Allard. Change-in-ratio estimates of lobster exploitation rate using sampling concurrent with fishing. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 60(10):1190–1203, October 2003. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f03-102.

Maunder:2003:ERC

[795] Mark N. Maunder and Richard B. Deriso. Estimation of recruitment in catch-at-age models. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 60(10): 1204–1216, October 2003. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f03-104.

Punt:2003:EPM

[796] André E. Punt. Extending production models to include process error in the population dynamics. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 60(10):1217–1228, October 2003. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f03-105.

Reiss:2003:HVD

[797] Christian S. Reiss, Ian A. McLaren, and Patricia Avendaño. Horizontal and vertical distribution patterns, retention rates, and population dynamics of zooplankton on Western Bank, Scotian Shelf. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 60(10):1229–1244, October 2003. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f03-106.

Barnard:2003:PIS

[798] Christine Barnard, Jean-Jacques Frenette, and Warwick F. Vincent. Planktonic invaders of the St. Lawrence estuarine transition zone: environmental factors controlling the distribution of zebra mussel veligers. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 60(10):1245–1257, October 2003. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f03-103.

Marty:2003:RDA

[799] Gary D. Marty, Terrance J. Quinn II, Greg Carpenter, Theodore R. Meyers, and Neil H. Willits. Role of disease in abundance of a Pacific herring (Clupea pallasi) population. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 60(10):1258–1265, October 2003. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f03-109.

Ebersole:2003:THS

[800] Joseph L. Ebersole, William J. Liss, and Christopher A. Frissell. Thermal heterogeneity, stream channel morphology, and salmonid abundance in northeastern Oregon streams. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 60(10):1266–1280, October 2003. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f03-107.

Weber:2003:FAM

[801] Jean-Michel Weber, Gérard Brichon, and Georges Zwingelstein. Fatty acid metabolism in rainbow trout (Oncorhynchus mykiss) tissues: differential incorporation of palmitate and oleate. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 60(10):1281–1288, October 2003. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f03-110.

Jordaan:2003:RRE

[802] Adrian Jordaan and Joseph A. Brown. The risk of running on empty: the influence of age on starvation and gut fullness in larval Atlantic cod (Gadus morhua). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 60(10): 1289–1298, October 2003. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f03-108.

Preston:2003:CAS

[803] J. M. Preston and R. L. Kirlin. Comment on “Acoustic seabed classification: improved statistical method”. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 60(10):1299–1300, October 2003. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f03-131.

Legendre:2003:RCP

[804] Pierre Legendre. Reply to the comment by Preston and Kirlin on “Acoustic seabed classification: improved statistical method”. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 60(10):1301–1305, October 2003. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f03-132.

Duplisea:2003:ESB

[805] Daniel E. Duplisea, Simon Jennings, Karema J. Warr, and Tracy A. Dinmore. Erratum: a size-based model of the impacts of bottom trawling on benthic community structure. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 60(10):1306, October 2003. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f03-122.

Boudreau:2003:DCZ

[806] Stephanie A. Boudreau and Norman D. Yan. The differing crustacean zooplankton communities of Canadian Shield lakes with and without the nonindigenous zooplanktivore Bythotrephes longimanus. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 60(11):1307–1313, November 2003. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f03-111.

Bunnell:2003:ZBE

[807] David B. Bunnell, María J. González, and Roy A. Stein. Zooplankton biomass enhances growth, but not survival, of first-feeding Pomoxis spp. larvae. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 60(11):1314–1323, November 2003. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f03-112.

Finn:2003:EAC

[808] Roderick Nigel Finn and Ivar Rønnestad. The effect of acute changes in temperature and light on the aerobic metabolism of embryos and yolk-sac larvae of turbot (Scophthalmus maximus). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 60(11):1324–1331, November 2003. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f03-113.

deBruyn:2003:RSL

[809] Adrian M. H. deBruyn, David J. Marcogliese, and Joseph B. Rasmussen. The role of sewage in a large river food web. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 60(11):1332–1344, November 2003. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f03-114.

Henderson:2003:SSD

[810] Bryan A. Henderson, Nicholas Collins, George E. Morgan, and André Vaillancourt. Sexual size dimorphism of walleye (Stizostedion vitreum vitreum). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 60(11):1345–1352, November 2003. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f03-115.

Noonburg:2003:IEZ

[811] Erik G. Noonburg, Brian J. Shuter, and Peter A. Abrams. Indirect effects of zebra mussels (Dreissena polymorpha) on the planktonic food web. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 60(11):1353–1368, November 2003. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f03-116.

Twiss:2003:PDT

[812] Michael R. Twiss, Benjamin S. Twining, and Nicholas S. Fisher. Partitioning of dissolved thallium by seston in Lakes Erie and Ontario. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 60(11):1369–1375, November 2003. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f03-119.

Milton:2003:MPT

[813] David A. Milton and Simon R. Chenery. Movement patterns of the tropical shad hilsa (Tenualosa ilisha) inferred from transects of ^{87}Sr/ 86Sr isotope ratios in their otoliths. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 60(11):1376–1385, November 2003. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f03-133.

Girard:2003:ECC

[814] Philippe Girard, Daniel Boisclair, and Michel Leclerc. The effect of cloud cover on the development of habitat quality indices for juvenile Atlantic salmon (Salmo salar). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 60(11): 1386–1397, November 2003. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f03-118.

Guay:2003:ATB

[815] J. C. Guay, D. Boisclair, M. Leclerc, and M. Lapointe. Assessment of the transferability of biological habitat models for Atlantic salmon parr (Salmo salar). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 60(11):1398– 1408, November 2003. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f03-120.

Palsson:2003:MPA

[816] Ólafur K. Pálsson and Vilhjálmur Thorsteinsson. Migration patterns, ambient temperature, and growth of Icelandic cod (Gadus morhua): evidence from storage tag data. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 60 (11):1409–1423, November 2003. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/

doi/full/10.1139/f03-117.

Friedman:2003:AIE

[817] Carolyn S. Friedman and Carl A. Finley. Anthropogenic introduction of the etiological agent of withering syndrome into northern California abalone populations via conservation efforts. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 60(11):1424–1431, November 2003. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f03-121.

Baldwin:2003:ECG

[818] Brad S. Baldwin, Marilyn S. Mayer, Jeffrey Dayton, Nancy Pau, Johanna Mendillo, Maura Sullivan, Aaron Moore, Aye Ma, and Edward L. Mills. Erratum: Comparative growth and feeding in zebra and quagga mussels (Dreissena polymorpha and Dreissena bugensis): implications for North American lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 60 (11):1432, November 2003. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f03-150.

Walters:2003:FFA

[819] Carl Walters. Folly and fantasy in the analysis of spatial catch rate data. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 60(12):1433–1436, December 2003. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f03-152.

Chick:2003:FPD

[820] John H. Chick, Robert J. Cosgriff, and Lori S. Gittinger. Fish as potential dispersal agents for floodplain plants: first evidence in North America. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 60(12):1437–1439, December 2003. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f03-155.

Hitt:2003:SHB

[821] Nathaniel P. Hitt, Christopher A. Frissell, Clint C. Muhlfeld, and Fred W. Allendorf. Spread of hybridization between native westslope cutthroat trout, Oncorhynchus clarki lewisi, and nonnative rainbow trout, Oncorhynchus mykiss. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 60(12): 1440–1451, December 2003. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f03-125.

Jassby:2003:DLT

[822] Alan D. Jassby, John E. Reuter, and Charles R. Goldman. Determining long-term water quality change in the presence of climate variability: Lake Tahoe (U.S.A.). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 60(12): 1452–1461, December 2003. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f03-127.

Hughes:2003:TMD

[823] Nicholas F. Hughes, John W. Hayes, Karen A. Shearer, and Roger G. Young. Testing a model of drift-feeding using three-dimensional videography of wild brown trout, Salmo trutta, in a New Zealand river. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 60(12):1462–1476, December 2003. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f03-126.

Stemberger:2003:CBL

[824] Richard S. Stemberger and Eric K. Miller. Cladoceran body length and secchi disk transparency in northeastern U.S. lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 60(12):1477–1486, December 2003. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f03-124.

Morrison:2003:DAY

[825] Wendy E. Morrison and David H. Secor. Demographic attributes of yellow-phase American eels (Anguilla rostrata) in the Hudson River estuary. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 60(12):1487–1501, December 2003. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f03-129.

Peterson:2003:UMN

[826] Douglas P. Peterson and Kurt D. Fausch. Upstream movement by nonnative brook trout (Salvelinus fontinalis) promotes invasion of native cutthroat trout (Oncorhynchus clarki) habitat. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 60(12):1502–1516, December 2003. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f03-128.

DeRobertis:2003:DET

[827] Alex De Robertis, Clifford H. Ryer, Adriana Veloza, and Richard D. Brodeur. Differential effects of turbidity on prey consumption of piscivorous and planktivorous fish. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 60(12):1517–1526, December 2003. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f03-123.

Jones:2003:DVN

[828] Michael L. Jones, John K. Netto, Jason D. Stockwell, and Joseph B. Mion. Does the value of newly accessible spawning habitat for walleye (Stizostedion vitreum) depend on its location relative to nursery habitats? Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 60(12):1527–1538, December 2003. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f03-130.

Harbitz:2003:TDG

[829] Alf Harbitz and Michaela Aschan. A two-dimensional geostatistic method to simulate the precision of abundance estimates. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 60(12):1539–1551, December 2003. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f03-134.

Bronte:2003:FCC

[830] Charles R. Bronte, Mark P. Ebener, Donald R. Schreiner, David S. DeVault, Michael M. Petzold, Douglas A. Jensen, Carl Richards, and Steven J. Lozano. Fish community change in Lake Superior, 1970–2000. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 60(12):1552–1574, December 2003. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f03-136.

Barnard:2003:EPI

[831] Christine Barnard, Jean-Jacques Frenette, and Warwick F. Vincent. Erratum: Planktonic invaders of the St. Lawrence estuarine transition zone: environmental factors controlling the distribution of zebra mussel veligers. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 60(12):1575, December 2003. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f03-161.

Levings:2004:OII

[832] C. D. Levings, J. R. Cordell, S. Ong, and G. E. Piercey. The origin and identity of invertebrate organisms being transported to Canada’s Pacific coast by ballast water. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 61 (1):1–11, January 2004. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f03-135.

Herwig:2004:TSE

[833] Brian R. Herwig, Daniel A. Soluk, John M. Dettmers, and David H. Wahl. Trophic structure and energy flow in backwater lakes of two large floodplain rivers assessed using stable isotopes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 61(1):12–22, January 2004. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f03-139.

Hanrahan:2004:ECS

[834] Timothy P. Hanrahan, Dennis D. Dauble, and David R. Geist. An estimate of Chinook salmon (Oncorhynchus tshawytscha) spawning habitat and redd capacity upstream of a migration barrier in the upper Columbia River. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 61(1):23–33, January 2004. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f03-140.

Martin:2004:TSE

[835] Gretchen Bath Martin, Simon R. Thorrold, and Cynthia M. Jones. Temperature and salinity effects on strontium incorporation in otoliths of larval spot (Leiostomus xanthurus). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 61(1):34–42, January 2004. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f03-143.

Briand:2004:MFP

[836] Genevieve Briand, Thomas Heckelei, Scott C. Matulich, and Ron C. Mittelhammer. Managing fishing power: the case of Alaska red king crab (Paralithodes camtschaticus). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 61 (1):43–53, January 2004. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f03-138.

Lenormand:2004:SOP

[837] Sophie Lenormand, Julian J. Dodson, and Annie Ménard. Seasonal and ontogenetic patterns in the migration of anadromous brook charr (Salvelinus fontinalis). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 61 (1):54–67, January 2004. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f03-137.

Taylor:2004:AHD

[838] Eric B. Taylor. An analysis of homogenization and differentiation of Canadian freshwater fish faunas with an emphasis on British Columbia. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 61(1):68–79, January 2004. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f03-141.

Flynn:2004:TFI

[839] Lucy Flynn and Ray Hilborn. Test fishery indices for sockeye salmon (Oncorhynchus nerka) as affected by age composition and environmental variables. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 61(1):80–92, January 2004. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f03-142.

Grenouillet:2004:WBF

[840] G. Grenouillet, D. Pont, and C. Hérissé. Within-basin fish assemblage structure: the relative influence of habitat versus stream spatial position on local species richness. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 61 (1):93–102, January 2004. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f03-145.

Hughes:2004:WDH

[841] Nicholas F. Hughes. The wave-drag hypothesis: an explanation for sizebased lateral segregation during the upstream migration of salmonids. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 61(1):103–109, January 2004. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f03-144.

Wiegand:2004:VNP

[842] Murray D. Wiegand, Thomas A. Johnston, Jennifer Martin, and William C. Leggett. Variation in neutral and polar lipid compositions of ova in ten reproductively isolated populations of walleye (Sander vitreus). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 61(1):110–121, January 2004. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f03-146.

Jiao:2004:SSI

[843] Yan Jiao, Yong Chen, David Schneider, and Joe Wroblewski. A simulation study of impacts of error structure on modeling stock–recruitment data using generalized linear models. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 61(1):122–133, January 2004. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f03-149.

Jiao:2004:AES

[844] Yan Jiao, David Schneider, Yong Chen, and Joe Wroblewski. An analysis of error structure in modeling the stock–recruitment data of gadoid stocks using generalized linear models. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 61(1):134–146, January 2004. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f03-148.

Morton:2004:SLL

[845] Alexandra Morton, Richard Routledge, Corey Peet, and Aleria Ladwig. Sea lice (Lepeophtheirus salmonis) infection rates on juvenile pink (Oncorhynchus gorbuscha) and chum (Oncorhynchus keta) salmon in the nearshore marine environment of British Columbia, Canada. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 61(2):147–157, February 2004. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f04-016.

Cardinale:2004:ESS

[846] M. Cardinale, P. Doering-Arjes, M. Kastowsky, and H. Mosegaard. Effects of sex, stock, and environment on the shape of known-age Atlantic cod (Gadus morhua) otoliths. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 61 (2):158–167, February 2004. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f03-151.

Grant:2004:UUV

[847] Gerold C. Grant, Paul Radomski, and Charles S. Anderson. Using underwater video to directly estimate gear selectivity: the retention probability for walleye (Sander vitreus) in gill nets. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 61(2):168–174, February 2004. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f03-166.

Matsuda:2004:EPP

[848] Hiroyuki Matsuda and Peter A. Abrams. Effects of predator–prey interactions and adaptive change on sustainable yield. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 61(2):175–184, February 2004. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f03-147.

Nykanen:2004:THP

[849] M. Nykänen and A. Huusko. Transferability of habitat preference criteria for larval European grayling (Thymallus thymallus). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 61(2):185–192, February 2004. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f03-156.

Seamons:2004:EAL

[850] Todd R. Seamons, Paul Bentzen, and Thomas P. Quinn. The effects of adult length and arrival date on individual reproductive success in wild steelhead trout (Oncorhynchus mykiss). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 61(2):193–204, February 2004. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f03-158.

Gust:2004:VPB

[851] Nick Gust. Variation in the population biology of protogynous coral reef fishes over tens of kilometres. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 61 (2):205–218, February 2004. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f03-160.

Ronkkonen:2004:LTC

[852] Sanna Rönkkönen, Evald Ojaveer, Tiit Raid, and Markku Viitasalo. Long-term changes in Baltic herring (Clupea harengus membras) growth in the Gulf of Finland. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 61(2): 219–229, February 2004. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f03-167.

Kjeld:2004:CBT

[853] Matthías Kjeld, Árni Alfredsson, Örn Ólafsson, Morten Tryland, Ivar Christensen, Snorre Stuen, and Alfred Árnason. Changes in blood testosterone and progesterone concentrations of the North Atlantic minke whale (Balaenoptera acutorostrata) during the feeding season. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 61(2):230–237, February 2004. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f03-159.

Stoner:2004:THM

[854] Allan W. Stoner and Erick A. Sturm. Temperature and hunger mediate sablefish (Anoplopoma fimbria) feeding motivation: implications for stock assessment. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 61(2): 238–246, February 2004. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f03-170.

Dowling:2004:APS

[855] Natalie A. Dowling, Stephen J. Hall, and Richard McGarvey. Assessing population sustainability and response to fishing in terms of aggregation structure for greenlip abalone (Haliotis laevigata) fishery management. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 61(2):247–259, February 2004. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f03-165.

Dowling:2004:SRF

[856] Natalie A. Dowling, Stephen J. Hall, and Richard McGarvey. Simulations of random fishing behaviour as an independent validation for the effect of active targeting of greenlip abalone (Haliotis laevigata) aggregations. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 61(2):260–272, February 2004. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f03-164.

Bujold:2004:DVR

[857] Valérie Bujold, Richard A. Cunjak, Jason P. Dietrich, and David A. Courtemanche. Drifters versus residents: assessing size and age differences in Atlantic salmon (Salmo salar) fry. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 61(2):273–282, February 2004. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f03-162.

Larsen:2004:DPC

[858] David P. Larsen, Philip R. Kaufmann, Thomas M. Kincaid, and N. Scott Urquhart. Detecting persistent change in the habitat of salmon-bearing streams in the Pacific Northwest. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 61(2):283–291, February 2004. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f03-157.

Eveson:2004:IMG

[859] J. Paige Eveson, Geoff M. Laslett, and Tom Polacheck. An integrated model for growth incorporating tag–recapture, length–frequency, and direct aging data. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 61(2):292–306, February 2004. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f03-163.

Polacheck:2004:IGR

[860] Tom Polacheck, J. Paige Eveson, and Geoff M. Laslett. Increase in growth rates of southern bluefin tuna (Thunnus maccoyii) over four decades: 1960 to 2000. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 61(2): 307–322, February 2004. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f04-005.

Munday:2004:SSG

[861] Philip L. Munday, Alyson L. Hodges, J. Howard Choat, and Nick Gust. Sex-specific growth effects in protogynous hermaphrodites. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 61(3):323–327, March 2004. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f04-057.

Olsen:2004:DMA

[862] K. Håkan Olsén, Erik Petersson, Bjarne Ragnarsson, Hans Lundqvist, and Torbjörn Järvi. Downstream migration in Atlantic salmon (Salmo salar) smolt sibling groups. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 61 (3):328–331, March 2004. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f04-067.

Cope:2004:PGP

[863] Jason M. Cope. Population genetics and phylogeography of the blue rockfish (Sebastes mystinus) from Washington to California. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 61(3):332–342, March 2004. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f04-008.

Johansen:2004:CJH

[864] Geir Odd Johansen, Bjarte Bogstad, Sigbjørn Mehl, and Øyvind Ulltang. Consumption of juvenile herring (Clupea harengus) by cod (Gadus morhua) in the Barents Sea: a new approach to estimating consumption in piscivorous fish. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 61 (3):343–359, March 2004. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f03-168.

Lawson:2004:EFI

[865] P. W. Lawson, E. A. Logerwell, N. J. Mantua, R. C. Francis, and V. N. Agostini. Environmental factors influencing freshwater survival and smolt production in Pacific Northwest Coho salmon (Oncorhynchus kisutch). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 61(3):360–373, March 2004. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f04-003.

Salas:2004:STD

[866] Silvia Salas, Ussif Rashid Sumaila, and Tony Pitcher. Short-term decisions of small-scale fishers selecting alternative target species: a choice model. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 61(3):374–383, March 2004. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f04-007.

Johnston:2004:EAS

[867] N. T. Johnston, E. A. MacIsaac, P. J. Tschaplinski, and K. J. Hall. Effects of the abundance of spawning sockeye salmon (Oncorhynchus nerka) on nutrients and algal biomass in forested streams. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 61(3):384–403, March 2004. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f03-172.

Sandahl:2004:OEF

[868] Jason F. Sandahl, David H. Baldwin, Jeffrey J. Jenkins, and Nathaniel L. Scholz. Odor-evoked field potentials as indicators of sublethal neurotoxicity in juvenile Coho salmon (Oncorhynchus kisutch) exposed to copper, chlorpyrifos, or esfenvalerate. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 61 (3):404–413, March 2004. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f04-011.

Shin:2004:UIB

[869] Yunne-Jai Shin and Philippe Cury. Using an individual-based model of fish assemblages to study the response of size spectra to changes in fishing. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 61(3):414–431, March 2004. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f03-154.

Wilson:2004:OPG

[870] Jonathan Mark Wilson, José Carlos Antunes, Patricío Duro Bouça, and João Coimbra. Osmoregulatory plasticity of the glass eel of Anguilla anguilla: freshwater entry and changes in branchial ion-transport protein expression. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 61(3):432–442, March 2004. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f04-010.

Kerr:2004:ROY

[871] Lisa A. Kerr, Allen H. Andrews, Brian R. Frantz, Kenneth H. Coale, Thomas A. Brown, and Gregor M. Cailliet. Radiocarbon in otoliths of yelloweye rockfish (Sebastes ruberrimus): a reference time series for the coastal waters of southeast Alaska. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 61(3):443–451, March 2004. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f04-009.

Kiessling:2004:EUM

[872] Anders Kiessling, Kerstin Lindahl-Kiessling, and Karl-Heinz Kiessling. Energy utilization and metabolism in spawning migrating early Stuart sockeye salmon (Oncorhynchus nerka): the migratory paradox. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 61(3):452–465, March 2004. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f04-006.

Dulvy:2004:SSC

[873] N. K. Dulvy, N. V. C. Polunin, A. C. Mill, and N. A. J. Graham. Size structural change in lightly exploited coral reef fish communities: evidence for weak indirect effects. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 61(3):466–475, March 2004. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f03-169.

Danancher:2004:CRE

[874] Delphine Danancher, Jacques Labonne, Roger Pradel, and Philippe Gaudin. Capture–recapture estimates of space used in streams (CRE-SUS) at the population scale: case study on Zingel asper (percid), a threatened species of the Rhône catchment. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 61(3):476–486, March 2004. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f04-004.

King:2004:PRT

[875] Sarah E. King, Robert W. Hannah, Steven J. Parker, Keith M. Matteson, and Steven A. Berkeley. Protecting rockfish through gear design: development of a selective flatfish trawl for the U.S. west coast bottom trawl fishery. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 61(3):487–496, March 2004. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f03-153.

Raikow:2004:FWI

[876] David F. Raikow. Food web interactions between larval bluegill (Lepomis macrochirus) and exotic zebra mussels (Dreissena polymorpha). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 61(3):497–504, March 2004. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f03-171.

Choi:2004:TAS

[877] Jae S. Choi, Kenneth T. Frank, William C. Leggett, and Ken Drinkwater. Transition to an alternate state in a continental shelf ecosystem. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 61(4):505–510, April 2004. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f04-079.

Fivelstad:2004:LLT

[878] Sveinung Fivelstad, Anne Berit Olsen, Sigurd Stefansson, Sigurd Handeland, Rune Waagbø, Frode Kroglund, and John Colt. Lack of longterm sublethal effects of reduced freshwater pH alone on Atlantic salmon (Salmo salar) smolts subsequently transferred to seawater. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 61(4):511–518, April 2004. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f04-002.

Uthicke:2004:PGD

[879] Sven Uthicke and Steven Purcell. Preservation of genetic diversity in restocking of the sea cucumber Holothuria scabra investigated by allozyme electrophoresis. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 61(4):519–528, April 2004. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/f04-
013.

Haggerty:2004:MRL

[880] S. M. Haggerty, D. P. Batzer, and C. R. Jackson. Macroinvertebrate response to logging in coastal headwater streams of Washington, U.S.A. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 61(4):529–537, April 2004. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f04-014.

James:2004:ISD

[881] William F. James, John W. Barko, and Harry L. Eakin. Impacts of sediment dewatering and rehydration on sediment nitrogen concentration and macrophyte growth. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 61 (4):538–546, April 2004. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f04-018.

Gallardo:2004:ENM

[882] José A. Gallardo, Jean Paul Lhorente, Ximena García, and Roberto Neira. Effects of nonrandom mating schemes to delay the inbreeding accumulation in cultured populations of Coho salmon (Oncorhynchus kisutch). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 61(4):547–553, April 2004. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f04-012.

Wetzel:2004:DIF

[883] D. L. Wetzel and J. E. Reynolds III. Definitive identification of fatty acid constituents in marine mammal tissues. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 61(4):554–560, April 2004. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f04-021.

Ramstack:2004:TCW

[884] Joy M. Ramstack, Sherilyn C. Fritz, and Daniel R. Engstrom. Twentieth century water quality trends in Minnesota lakes compared with presettlement variability. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 61 (4):561–576, April 2004. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f04-015.

Kostow:2004:DJP

[885] Kathryn E. Kostow. Differences in juvenile phenotypes and survival between hatchery stocks and a natural population provide evidence for modified selection due to captive breeding. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 61(4):577–589, April 2004. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f04-019.

Greene:2004:CPD

[886] Correigh M. Greene and Timothy J. Beechie. Consequences of potential density-dependent mechanisms on recovery of ocean-type Chinook salmon (Oncorhynchus tshawytscha). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 61(4):590–602, April 2004. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f04-024.

Hudon:2004:SWP

[887] Christiane Hudon. Shift in wetland plant composition and biomass following low-level episodes in the St. Lawrence River: looking into the future. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 61(4):603–617, April 2004. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f04-031.

Sloman:2004:DSS

[888] Katherine A. Sloman, Graham R. Scott, D. Gordon McDonald, and Chris M. Wood. Diminished social status affects ionoregulation at the gills and kidney in rainbow trout (Oncorhynchus mykiss). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 61(4):618–626, April 2004. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f04-032.

Madenjian:2004:ECS

[889] Charles P. Madenjian, Daniel V. O’Connor, Sergei M. Chernyak, Richard R. Rediske, and James P. O’Keefe. Evaluation of a Chinook salmon (Oncorhynchus tshawytscha) bioenergetics model. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 61(4):627–635, April 2004. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f04-033.

Lewy:2004:SGC

[890] Peter Lewy, J. Rasmus Nielsen, and Holger Hovgård. Survey gear calibration independent of spatial fish distribution. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 61(4):636–647, April 2004. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f04-034.

Coulombe-Pontbriand:2004:LCB

[891] M. Coulombe-Pontbriand and Michel Lapointe. Landscape controls on boulder-rich, winter habitat availability and their effects on Atlantic salmon (Salmo salar) parr abundance in two fifth-order mountain streams. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 61(4):648–658, April 2004. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f04-023.

Pepin:2004:ELH

[892] Pierre Pepin. Early life history studies of prey–predator interactions: quantifying the stochastic individual responses to environmental variability. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 61(4):659–671, April 2004. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f04-078.

Anonymous:2004:TIS

[893] Anonymous. Third International Symposium on Biological and Environmental Chemistry of DMS(P) and Related Compounds / [Troisième symposium international sur la biochimie et la chimie environnementale du DMS(P) et des composés apparentés]. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 61(5), May 2004. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f04-904.

Levasseur:2004:II

[894] Maurice Levasseur. Introduction / introduction. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 61(5), May 2004. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f04-905.

Simo:2004:CGA

[895] Rafel Simó. From cells to globe: approaching the dynamics of DMS(P) in the ocean at multiple scales. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 61 (5):673–684, May 2004. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f04-030.

Keller:2004:RCP

[896] Maureen D. Keller, Patricia A. Matrai, Ronald P. Kiene, and Wendy K. Bellows. Responses of coastal phytoplankton populations to nitrogen additions: dynamics of cell-associated dimethylsulfoniopropionate (DMSP), glycine betaine (GBT), and homarine. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 61(5):685–699, May 2004. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f04-058.

Harada:2004:LVD

[897] Hyakubun Harada, Mary-Anne Rouse, William Sunda, and Ronald P. Kiene. Latitudinal and vertical distributions of particle-associated dimethylsulfoniopropionate (DMSP) lyase activity in the western North Atlantic Ocean. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 61(5):700–711, May 2004. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/f04-
046.

Merzouk:2004:IDD

[898] Anissa Merzouk, Maurice Levasseur, Michael Scarratt, Sonia Michaud, and Michel Gosselin. Influence of dinoflagellate diurnal vertical migrations on dimethylsulfoniopropionate and dimethylsulfide distribution and dynamics (St. Lawrence Estuary, Canada). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 61(5):712–720, May 2004. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f04-066.

Riseman:2004:PDD

[899] S. F. Riseman and G. R. DiTullio. Particulate dimethylsulfoniopropionate and dimethylsulfoxide in relation to iron availability and algal community structure in the Peru upwelling system. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 61(5):721–735, May 2004. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f04-052.

Kasamatsu:2004:PIZ

[900] Nobue Kasamatsu, So Kawaguchi, Shuichi Watanabe, Tsuneo Odate, and Mitsuo Fukuchi. Possible impacts of zooplankton grazing on dimethylsulfide production in the Antarctic Ocean. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 61(5):736–743, May 2004. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f04-072.

Levasseur:2004:VRC

[901] M. Levasseur, M. Scarratt, S. Roy, D. Laroche, S. Michaud, G. Cantin, M. Gosselin, and A. Vézina. Vertically resolved cycling of dimethylsulfoniopropionate (DMSP) and dimethylsulfide (DMS) in the Northwest Atlantic in spring. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 61(5):744–757, May 2004. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/f04-
026.

Hill:2004:DSS

[902] Richard W. Hill, John W. H. Dacey, Susan D. Hill, Ahser Edward, and Wayne A. Hicks. Dimethylsulfoniopropionate in six species of giant clams and the evolution of dimethylsulfide after death. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 61(5):758–764, May 2004. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f04-029.

Archer:2004:MSP

[903] Stephen D. Archer, Francis J. Gilbert, J. Icarus Allen, Jerry Blackford, and Philip D. Nightingale. Modelling of the seasonal patterns of dimethylsulphide production and fate during 1989 at a site in the North Sea. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 61(5):765–787, May 2004. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f04-028.

LeClainche:2004:BOD

[904] Yvonnick Le Clainche, Maurice Levasseur, Alain Vézina, John W. H. Dacey, and François J. Saucier. Behaviour of the ocean DMS(P) pools in the Sargasso Sea viewed in a coupled physical-biogeochemical ocean model. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 61(5):788–803, May 2004. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f04-027.

Belviso:2004:AGC

[905] S. Belviso, C. Moulin, L. Bopp, and J. Stefels. Assessment of a global climatology of oceanic dimethylsulfide (DMS) concentrations based on Sea-WiFS imagery (1998–2001). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 61 (5):804–816, May 2004. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f04-001.

Wadleigh:2004:SIC

[906] Moire A. Wadleigh. Sulphur isotopic composition of aerosols over the western North Atlantic Ocean. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 61(5):817–825, May 2004. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/

doi/full/10.1139/f04-073.

Bopp:2004:WMD

[907] Laurent Bopp, Olivier Boucher, Olivier Aumont, Sauveur Belviso, Jean-Louis Dufresne, Mai Pham, and Patrick Monfray. Will marine dimethylsulfide emissions amplify or alleviate global warming? A model study. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 61(5):826–835, May 2004. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f04-045.

Zemmelink:2004:DMB

[908] H. J. Zemmelink, J. W. H. Dacey, and E. J. Hintsa. Direct measurements of biogenic dimethylsulphide fluxes from the oceans: a synthesis. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 61(5):836–844, May 2004. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f04-047.

Vezina:2004:EMC

[909] Alain F. Vézina. Ecosystem modelling of the cycling of marine dimethylsulfide: a review of current approaches and of the potential for extrapolation to global scales. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 61 (5):845–856, May 2004. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f04-025.

Albanese:2004:ECF

[910] Brett Albanese, Paul L. Angermeier, and Sundar Dorai-Raj. Ecological correlates of fish movement in a network of Virginia streams. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 61(6):857–869, June 2004. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f04-096.

Skaug:2004:AMW

[911] Hans J. Skaug, Nils Øien, Tore Schweder, and Gjermund Bøthun. Abundance of minke whales (Balaenoptera acutorostrata) in the Northeast Atlantic: variability in time and space. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 61(6):870–886, June 2004. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f04-020.

Heard:2004:GCF

[912] Stephen B. Heard and Corinne K. Buchanan. Grazer–collector facilitation hypothesis supported by laboratory but not field experiments. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 61(6):887–897, June 2004. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f04-022.

Patterson:2004:IOL

[913] Heather M. Patterson, Michael J. Kingsford, and Malcolm T. McCulloch. The influence of oceanic and lagoonal plume waters on otolith chemistry. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 61(6):898–904, June 2004. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f04-036.

Standen:2004:IRS

[914] Emily M. Standen, Scott G. Hinch, and Peter S. Rand. Influence of river speed on path selection by migrating adult sockeye salmon (Oncorhynchus nerka). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 61 (6):905–912, June 2004. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f04-035.

Johnson:2004:FIS

[915] Sherri L. Johnson. Factors influencing stream temperatures in small streams: substrate effects and a shading experiment. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 61(6):913–923, June 2004. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f04-040.

Strayer:2004:EIB

[916] David L. Strayer, Kathryn A. Hattala, and Andrew W. Kahnle. Effects of an invasive bivalve (Dreissena polymorpha) on fish in the Hudson River estuary. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 61(6):924–941, June 2004. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f04-043.

Niyogi:2004:ACB

[917] S. Niyogi, P. Couture, G. Pyle, D. G. McDonald, and C. M. Wood. Acute cadmium biotic ligand model characteristics of laboratory-reared and wild yellow perch (Perca flavescens) relative to rainbow trout (Oncorhynchus mykiss). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 61 (6):942–953, June 2004. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f04-044.

Wall:2004:FHM

[918] Steven S. Wall, Charles R. Berry, Jr., Carmen M. Blausey, Jonathan A. Jenks, and Chad J. Kopplin. Fish-habitat modeling for gap analysis to conserve the endangered Topeka shiner (Notropis topeka). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 61(6):954–973, June 2004. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f04-017.

Savereide:2004:ASA

[919] James W. Savereide and Terrance J. Quinn II. An age-structured assessment model for Chinook salmon (Oncorhynchus tshawytscha). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 61(6):974–985, June 2004. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f04-039.

Verschuren:2004:QRP

[920] Dirk Verschuren, Brian F. Cumming, and Kathleen R. Laird. Quantitative reconstruction of past salinity variations in African lakes: assessment of chironomid-based inference models (Insecta: Diptera) in space and time. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 61(6):986–998, June 2004. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f04-041.

Steel:2004:LMU

[921] E. Ashley Steel, Blake E. Feist, David W. Jensen, George R. Pess, Mindi B. Sheer, Jody B. Brauner, and Robert E. Bilby. Landscape models to understand steelhead (Oncorhynchus mykiss) distribution and help prioritize barrier removals in the willamette basin, Oregon, USA. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 61(6):999–1011, June 2004. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f04-042.

Hussy:2004:EAT

[922] K. Hüssy, H. Mosegaard, and F. Jessen. Effect of age and temperature on amino acid composition and the content of different protein types of juvenile Atlantic cod (Gadus morhua) otoliths. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 61(6):1012–1020, June 2004. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f04-037.

Hussy:2004:ACG

[923] K. Hüssy and H. Mosegaard. Atlantic cod (Gadus morhua) growth and otolith accretion characteristics modelled in a bioenergetics context. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 61(6):1021–1031, June 2004. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f04-038.

Michielsens:2004:BHA

[924] Catherine G. J. Michielsens and Murdoch K. McAllister. A Bayesian hierarchical analysis of stock–recruit data: quantifying structural and parameter uncertainties. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 61 (6):1032–1047, June 2004. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f04-048.

McAllister:2004:BHF

[925] Murdoch K. McAllister, Simeon L. Hill, David J. Agnew, Geoffrey P. Kirkwood, and John R. Beddington. A Bayesian hierarchical formulation of the De lury stock assessment model for abundance estimation of Falkland Islands’ squid (Loligo gahi). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 61(6):1048–1059, June 2004. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f04-084.

Dancik:2004:AAN

[926] Bruce Dancik. Announcement: 75th Anniversary of NRC Research Press / Communiqué : 75^{e} Anniversaire des Presses scientifiques du CNRC. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 61(7), July 2004. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f04-906.

Walters:2004:SRD

[927] Carl Walters. Simple representation of the dynamics of biomass error propagation for stock assessment models. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 61(7):1061–1065, July 2004. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f04-120.

Wagner:2004:BCC

[928] Tyler Wagner and James L. Congleton. Blood chemistry correlates of nutritional condition, tissue damage, and stress in migrating juvenile Chinook salmon (Oncorhynchus tshawytscha). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 61(7):1066–1074, July 2004. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f04-050.

Kalinowski:2004:GPM

[929] Steven T. Kalinowski. Genetic polymorphism and mixed-stock fisheries analysis. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 61(7):1075–1082, July 2004. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f04-060.

Starr:2004:MLO

[930] Richard M. Starr, Victoria O’Connell, and Stephen Ralston. Movements of lingcod (Ophiodon elongatus) in southeast Alaska: potential for increased conservation and yield from marine reserves. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 61(7):1083–1094, July 2004. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f04-054.

Rossi:2004:OIR

[931] Carla M. Riva Rossi, Enrique P. Lessa, and Miguel A. Pascual. The origin of introduced rainbow trout (Oncorhynchus mykiss) in the Santa Cruz River, Patagonia, Argentina, as inferred from mitochondrial DNA. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 61(7):1095–1101, July 2004. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f04-056.

Richardson:2004:DUM

[932] William B. Richardson, Eric A. Strauss, Lynn A. Bartsch, Emy M. Monroe, Jennifer C. Cavanaugh, Lorrine Vingum, and David M. Soballe. Denitrification in the Upper Mississippi River: rates, controls, and contribution to nitrate flux. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 61 (7):1102–1112, July 2004. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f04-062.

Wei:2004:ISF

[933] A. Wei, P. Chow-Fraser, and D. Albert. Influence of shoreline features on fish distribution in the Laurentian Great Lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 61(7):1113–1123, July 2004. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f04-061.

Romano:2004:ASM

[934] T. A. Romano, M. J. Keogh, C. Kelly, P. Feng, L. Berk, C. E. Schlundt, D. A. Carder, and J. J. Finneran. Anthropogenic sound and marine mammal health: measures of the nervous and immune systems before and after intense sound exposure. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 61 (7):1124–1134, July 2004. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f04-055.

Hrafnkelsson:2004:MCL

[935] Birgir Hrafnkelsson and Gunnar Stefánsson. A model for categorical length data from groundfish surveys. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 61(7):1135–1142, July 2004. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f04-049.

Miscampbell:2004:MMV

[936] Allyson E. Miscampbell, Murray W. Lankester, and Martin L. Adamson. Molecular and morphological variation within swim bladder nematodes, Cystidicola spp. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 61 (7):1143–1152, July 2004. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f04-064.

Lafrancois:2004:CEN

[937] Brenda Moraska Lafrancois, Koren R. Nydick, Brett M. Johnson, and Jill S. Baron. Cumulative effects of nutrients and pH on the plankton of two mountain lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 61 (7):1153–1165, July 2004. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f04-053.

Arkoosh:2004:PIP

[938] M. R. Arkoosh, L. Johnson, P. A. Rossignol, and T. K. Collier. Predicting the impact of perturbations on salmon (Oncorhynchus spp.) communities: implications for monitoring. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 61(7):1166–1175, July 2004. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f04-068.

Todd:2004:PGD

[939] C. D. Todd, A. M. Walker, M. G. Ritchie, J. A. Graves, and A. F. Walker. Population genetic differentiation of sea lice (Lepeophtheirus salmonis) parasitic on Atlantic and Pacific salmonids: analyses of microsatellite DNA variation among wild and farmed hosts. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 61(7):1176–1190, July 2004. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f04-069.

Davis:2004:LEK

[940] Anthony Davis, John Mark Hanson, Hadley Watts, and Holli MacPherson. Local ecological knowledge and marine fisheries research: the case of white hake (Urophycis tenuis) predation on juvenile American lobster (Homarus americanus). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 61 (7):1191–1201, July 2004. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f04-070.

Hall:2004:BAO

[941] Norman G. Hall, S. Alex Hesp, and Ian C. Potter. A Bayesian approach for overcoming inconsistencies in mortality estimates using, as an example, data for Acanthopagrus latus. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 61(7):1202–1211, July 2004. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f04-074.

Arkhipkin:2004:STV

[942] Alexander I. Arkhipkin, Steven E. Campana, Jennifer FitzGerald, and Simon R. Thorrold. Spatial and temporal variation in elemental signatures of statoliths from the Patagonian longfin squid (Loligo gahi). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 61(7):1212–1224, July 2004. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f04-075.

Patterson:2004:MIP

[943] D. A. Patterson, H. Guderley, P. Bouchard, J. S. Macdonald, and A. P. Farrell. Maternal influence and population differences in activities of mitochondrial and glycolytic enzymes in emergent sockeye salmon (Oncorhynchus nerka) fry. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 61 (7):1225–1234, July 2004. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f04-076.

Gregory-Eaves:2004:ISS

[944] Irene Gregory-Eaves, Bruce P. Finney, Marianne S. V. Douglas, and John P. Smol. Inferring sockeye salmon (Oncorhynchus nerka) population dynamics and water quality changes in a stained nursery lake over the past ≈ 500 years. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 61 (7):1235–1246, July 2004. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f04-071.

Payan:2004:ESP

[945] P. Payan, H. De Pontual, A. Edeyer, G. Borelli, G. Boeuf, and N. Mayer-Gostan. Effects of stress on plasma homeostasis, endolymph chemistry, and check formation during otolith growth in rainbow trout (Oncorhynchus mykiss). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 61 (7):1247–1255, July 2004. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f04-059.

Nydick:2004:NRA

[946] Koren R. Nydick, Brenda Moraska Lafrancois, Jill S. Baron, and Brett M. Johnson. Nitrogen regulation of algal biomass, productivity, and composition in shallow mountain lakes, Snowy Range, Wyoming, USA. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 61(7):1256–1268, July 2004. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f04-085.

Francis:2004:IAO

[947] R. I. C. Chris Francis and Steven E. Campana. Inferring age from otolith measurements: a review and a new approach. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 61(7):1269–1284, July 2004. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f04-063.

Hecky:2004:NPS

[948] R. E. Hecky, R. E. H. Smith, D. R. Barton, S. J. Guildford, W. D. Taylor, M. N. Charlton, and T. Howell. The nearshore phosphorus shunt: a consequence of ecosystem engineering by dreissenids in the Laurentian Great Lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 61(7):1285–1293, July 2004. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f04-065.

McArdle:2004:VHT

[949] Brian H. McArdle and Marti J. Anderson. Variance heterogeneity, transformations, and models of species abundance: a cautionary tale. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 61(7):1294–1302, July 2004. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f04-051.

Anonymous:2004:BRP

[950] Anonymous. Biological reference points for invertebrate fisheries / [Points de référence biologiques pour la pê che aux invertébres]. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 61(8), August 2004. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f04-900.

Smith:2004:BRPa

[951] Stephen J. Smith and Bernard Sainte-Marie. Biological reference points for invertebrate fisheries: introduction. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 61(8):1303–1306, August 2004. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f04-135.

Caddy:2004:CUF

[952] J. F. Caddy. Current usage of fisheries indicators and reference points, and their potential application to management of fisheries for marine invertebrates. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 61(8):1307–1324, August 2004. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f04-132.

Botsford:2004:BRP

[953] Louis W. Botsford, Alan Campbell, and Robert Miller. Biological reference points in the management of North American sea urchin fisheries. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 61(8):1325–1337, August 2004. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f04-131.

Smith:2004:BRPb

[954] Stephen J. Smith and Paul Rago. Biological reference points for sea scallops (Placopecten magellanicus): the benefits and costs of being nearly sessile. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 61(8):1338–1354, August 2004. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/f04-
134.

Orensanz:2004:PHM

[955] J. M. (Lobo) Orensanz, Claudia M. Hand, Ana M. Parma, Juan Valero, and Ray Hilborn. Precaution in the harvest of Methuselah’s clams — the difficulty of getting timely feedback from slow-paced dynamics. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 61(8):1355–1372, August 2004. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f04-136.

Cadrin:2004:HAD

[956] Steven X. Cadrin, James A. Boutillier, and Josef S. Idoine. A hierarchical approach to determining reference points for pandalid shrimp. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 61(8):1373–1391, August 2004. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f04-133.

Fogarty:2004:BRP

[957] Michael J. Fogarty and Louise Gendron. Biological reference points for American lobster (Homarus americanus) populations: limits to exploitation and the precautionary approach. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 61(8):1392–1403, August 2004. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f04-140.

Siddeek:2004:CRP

[958] M. S. M. Siddeek, Bernard Sainte-Marie, Jim Boutillier, and Gretchen Bishop. Comparison of reference points estimated using a size-based method for two high-latitude crab species in the United States and Canada. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 61(8):1404–1430, August 2004. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/f04-
137.

Merz:2004:PBS

[959] Joseph E. Merz, Jose D. Setka, Gregory B. Pasternack, and Joseph M. Wheaton. Predicting benefits of spawning-habitat rehabilitation to salmonid (Oncorhynchus spp.) fry production in a regulated California river. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 61(8):1433–1446, August 2004. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/f04-
077.

Hillman:2004:ECC

[960] Graham R. Hillman, Joseph C. Feng, Cecilia C. Feng, and Yonghe Wang. Effects of catchment characteristics and disturbances on storage and export of dissolved organic carbon in a boreal headwater stream. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 61(8):1447–1460, August 2004. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f04-082.

Tank:2004:RUR

[961] Suzanne E. Tank and David W. Schindler. The role of ultraviolet radiation in structuring epilithic algal communities in Rocky Mountain montane lakes: evidence from pigments and taxonomy. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 61(8):1461–1474, August 2004. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f04-080.

Sigler:2004:ASS

[962] Michael F. Sigler, Jamie N. Womble, and Johanna J. Vollenweider. Availability to Steller sea lions (Eumetopias jubatus) of a seasonal prey resource: a prespawning aggregation of eulachon (Thaleichthys pacificus). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 61(8):1475–1484, August 2004. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f04-086.

Perga:2004:CPF

[963] Marie-Elodie Perga and Daniel Gerdeaux. Changes in the δ ^{13}C of pelagic food webs: the influence of lake area and trophic status on the isotopic signature of whitefish (Coregonus lavaretus). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 61(8):1485–1492, August 2004. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f04-089.

Johnson:2004:ESN

[964] R. K. Johnson and M. L. Ostrofsky. Effects of sediment nutrients and depth on small-scale spatial heterogeneity of submersed macrophyte communities in Lake Pleasant, Pennsylvania. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 61(8):1493–1502, August 2004. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f04-081.

Jones:2004:ILV

[965] J. R. Jones, M. F. Knowlton, D. V. Obrecht, and E. A. Cook. Importance of landscape variables and morphology on nutrients in Missouri reservoirs. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 61(8):1503–1512, August 2004. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f04-088.

Biro:2004:WLD

[966] Peter A. Biro, Ashley E. Morton, John R. Post, and Eric A. Parkinson. Over-winter lipid depletion and mortality of age-0 rainbow trout (Oncorhynchus mykiss). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 61(8): 1513–1519, August 2004. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f04-083.

Hakkinen:2004:HPD

[967] Petri J. Häkkinen, Alexandre M. Anesio, and Wilhelm Granéli. Hydrogen peroxide distribution, production, and decay in boreal lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 61(8):1520–1527, August 2004. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f04-098.

Schrank:2004:MPI

[968] Amy J. Schrank and Frank J. Rahel. Movement patterns in inland cutthroat trout (Oncorhynchus clarki utah): management and conservation implications. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 61(8): 1528–1537, August 2004. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f04-097.

Davies:2004:TCN

[969] John-Mark Davies, Weston H. Nowlin, and Asit Mazumder. Temporal changes in nitrogen and phosphorus codeficiency of plankton in lakes of coastal and interior British Columbia. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 61(8):1538–1551, August 2004. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f04-092.

Rose:2004:ROC

[970] G. A. Rose. Reconciling overfishing and climate change with stock dynamics of Atlantic cod (Gadus morhua) over 500 years. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 61(9):1553–1557, September 2004. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f04-173.

Brander:2004:ENA

[971] K. Brander and R. Mohn. Effect of the North Atlantic oscillation on recruitment of Atlantic cod (Gadus morhua). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 61(9):1558–1564, September 2004. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f04-087.

Meyer:2004:AFW

[972] Elisabeth I. Meyer and Rainer Poepperl. Assessing food-web structure, matter fluxes, and system attributes of a Central European mountain stream by performing mass-balanced network analysis. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 61(9):1565–1581, September 2004. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f04-104.

Moore:2004:NEF

[973] Jonathan W. Moore and Daniel E. Schindler. Nutrient export from freshwater ecosystems by anadromous sockeye salmon (Oncorhynchus nerka).

Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 61(9):1582–1589, September 2004. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f04-103.

Haro:2004:SPU

[974] Alex Haro, Theodore Castro-Santos, John Noreika, and Mufeed Odeh. Swimming performance of upstream migrant fishes in open-channel flow: a new approach to predicting passage through velocity barriers. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 61(9):1590–1601, September 2004. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f04-093.

Castro-Santos:2004:QCE

[975] Theodore Castro-Santos. Quantifying the combined effects of attempt rate and swimming capacity on passage through velocity barriers. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 61(9):1602–1615, September 2004. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f04-094.

Etnier:2004:RED

[976] Michael A. Etnier. Reevaluating evidence of density-dependent growth in northern fur seals (Callorhinus ursinus) based on measurements of archived skeletal specimens. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 61 (9):1616–1626, September 2004. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/

doi/full/10.1139/f04-090.

MacNeil:2004:RNF

[977] Calum MacNeil, John Prenter, Mark Briffa, Nina J. Fielding, Jaimie T. A. Dick, Gillian E. Riddell, Melanie J. Hatcher, and Alison M. Dunn. The replacement of a native freshwater amphipod by an invader: roles for environmental degradation and intraguild predation. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 61(9):1627–1635, September 2004. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f04-091.

Hardie:2004:GSE

[978] David C. Hardie and Paul D. N. Hebert. Genome-size evolution in fishes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 61(9):1636–1646, September 2004. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f04-106.

Hammond:2004:RBN

[979] T. R. Hammond. A recipe for Bayesian network driven stock assessment. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 61(9):1647–1657, September 2004. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f04-107.

Parkinson:2004:LDH

[980] Eric A. Parkinson, John R. Post, and Sean P. Cox. Linking the dynamics of harvest effort to recruitment dynamics in a multistock, spatially structured fishery. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 61(9):1658– 1670, September 2004. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f04-101.

Dalsgaard:2004:FAB

[981] Johanne Dalsgaard and Michael St John. Fatty acid biomarkers: validation of food web and trophic markers using ^{13}C-labelled fatty acids in juvenile sandeel (Ammodytes tobianus). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 61(9):1671–1680, September 2004. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f04-095.

Dans:2004:RSA

[982] Silvana Laura Dans, Enrique Alberto Crespo, Susana Noemí Pedraza, and Mariano Koen Alonso. Recovery of the South American sea lion (Otaria flavescens) population in northern Patagonia. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 61(9):1681–1690, September 2004. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f04-105.

Clerk:2004:CAW

[983] Saloni Clerk, Daniel T. Selbie, and John P. Smol. Cage aquaculture and water-quality changes in the LaCloche Channel, Lake Huron, Canada: a paleolimnological assessment. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 61 (9):1691–1701, September 2004. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/

doi/full/10.1139/f04-099.

Giguere:2004:ILC

[984] Anik Giguère, Peter G. C. Campbell, Landis Hare, D. Gordon McDonald, and Joseph B. Rasmussen. Influence of lake chemistry and fish age on cadmium, copper, and zinc concentrations in various organs of indigenous yellow perch (Perca flavescens). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 61(9):1702–1716, September 2004. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f04-100.

Gray:2004:SFS

[985] M. A. Gray, R. A. Cunjak, and K. R. Munkittrick. Site fidelity of slimy sculpin (Cottus cognatus): insights from stable carbon and nitrogen analysis. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 61(9):1717– 1722, September 2004. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f04-108.

Miller:2004:ELL

[986] J. A. Miller and A. L. Shanks. Evidence for limited larval dispersal in black rockfish (Sebastes melanops): implications for population structure and marine-reserve design. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 61(9):1723–1735, September 2004. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f04-111.

Jones:2004:RSF

[987] Nicholas E. Jones and William M. Tonn. Resource selection functions for age-0 Arctic grayling (Thymallus arcticus) and their application to stream habitat compensation. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 61(9):1736–1746, September 2004. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f04-116.

Hjermann:2004:PDN

[988] Dag Ø. Hjermann, Nils Chr. Stenseth, and Geir Ottersen. The population dynamics of Northeast Arctic cod (Gadus morhua) through two decades: an analysis based on survey data. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 61(9):1747–1755, September 2004. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f04-115.

Ruggerone:2004:SPS

[989] Gregory T. Ruggerone and Frederick A. Goetz. Survival of Puget Sound Chinook salmon (Oncorhynchus tshawytscha) in response to climateinduced competition with pink salmon (Oncorhynchus gorbuscha). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 61(9):1756–1770, September 2004. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f04-112.

Gibson:2004:ERF

[990] A. Jamie F. Gibson and Ransom A. Myers. Estimating reference fishing mortality rates from noisy spawner–recruit data. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 61(9):1771–1783, September 2004. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f04-102.

Egertson:2004:RFC

[991] Christopher J. Egertson and John A. Downing. Relationship of fish catch and composition to water quality in a suite of agriculturally eutrophic lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 61(9):1784– 1796, September 2004. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f04-109.

Bernier:2004:OCD

[992] Nicholas J. Bernier, Colin J. Brauner, John W. Heath, and David J. Randall. Oxygen and carbon dioxide transport during sustained exercise in diploid and triploid Chinook salmon (Oncorhynchus tshawytscha).

Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 61(9):1797–1805, September 2004. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f04-110.

Newbrey:2004:FRC

[993] Michael G. Newbrey and Allan C. Ashworth. A fossil record of colonization and response of lacustrine fish populations to climate change. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 61(10):1807–1816, October 2004. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f04-113.

Mermillod-Blondin:2004:INI

[994] Florian Mermillod-Blondin, Magali Gérino, Sabine Sauvage, and Michel Creuzé des Châtelliers. Influence of nontrophic interactions between benthic invertebrates on river sediment processes: a microcosm study. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 61(10):1817–1831, October 2004. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f04-114.

Wright:2004:IVE

[995] Hazel A. Wright, Robert J. Wootton, and Iain Barber. Interpopulation variation in early growth of threespine sticklebacks (Gasterosteus aculeatus) under laboratory conditions. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 61(10):1832–1838, October 2004. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f04-119.

Gauthier:2004:ACF

[996] Stéphane Gauthier and John K. Horne. Acoustic characteristics of forage fish species in the Gulf of Alaska and Bering Sea based on Kirchhoffapproximation models. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 61(10): 1839–1850, October 2004. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f04-117.

Litvin:2004:MAS

[997] Steven Y. Litvin and Michael P. Weinstein. Multivariate analysis of stable-isotope ratios to infer movements and utilization of estuarine organic matter by juvenile weakfish (Cynoscion regalis). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 61(10):1851–1861, October 2004. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f04-121.

Horppila:2004:ECT

[998] Jukka Horppila, Anne Liljendahl-Nurminen, and Tommi Malinen. Effects of clay turbidity and light on the predator–prey interaction between smelts and chaoborids. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 61(10): 1862–1870, October 2004. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f04-123.

Lehman:2004:CIP

[999] John T. Lehman, Ali Bazzi, Todd Nosher, and Jerome O. Nriagu. Copper inhibition of phytoplankton in Saginaw Bay, Lake Huron. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 61(10):1871–1880, October 2004. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f04-129.

Clark:2004:NEA

[1000] William G. Clark. Nonparametric estimates of age misclassification from paired readings. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 61(10): 1881–1889, October 2004. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f04-122.

Wares:2004:GPM

[1001] John P. Wares, Dominique Alò, and Thomas F. Turner. A genetic perspective on management and recovery of federally endangered trout (Oncorhynchus gilae) in the American southwest. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 61(10):1890–1899, October 2004. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f04-124.

Marshall:2004:DCI

[1002] C. Tara Marshall, Coby L. Needle, Nathalia A. Yaragina, Adnan M. Ajiad, and Evgeny Gusev. Deriving condition indices from standard fisheries databases and evaluating their sensitivity to variation in stored energy reserves. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 61(10):1900– 1917, October 2004. CODEN CJFSDX. ISSN 0706-652X (print), 1205-

7533 (electronic). URL https://cdnsciencepub.com/doi/full/10.
1139/f04-128.

Borges:2004:OSL

[1003] Lisa Borges, Alain F. Zuur, Emer Rogan, and Rick Officer. Optimum sampling levels in discard sampling programs. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 61(10):1918–1928, October 2004. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f04-138.

Lobon-Cervia:2004:DDC

[1004] Javier Lobón-Cerviá. Discharge-dependent covariation patterns in the population dynamics of brown trout (Salmo trutta) within a Cantabrian river drainage. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 61(10):1929– 1939, October 2004. CODEN CJFSDX. ISSN 0706-652X (print), 1205-

7533 (electronic). URL https://cdnsciencepub.com/doi/full/10.
1139/f04-118.

Richoux:2004:SCL

[1005] Nicole B. Richoux, Don Deibel, Raymond J. Thompson, and Christopher C. Parrish. Seasonal changes in the lipids of Mysis mixta (Mysidacea) from the hyperbenthos of a cold-ocean environment (Conception Bay, Newfoundland). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 61(10): 1940–1953, October 2004. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f04-139.

Buzby:2004:LTS

[1006] Karen M. Buzby and Linda A. Deegan. Long-term survival of adult Arctic grayling (Thymallus arcticus) in the Kuparuk River, Alaska. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 61(10):1954–1964, October 2004. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f04-126.

Clair:2004:PFC

[1007] Thomas A. Clair, Ian F. Dennis, Peter G. Amiro, and B. J. Cosby. Past and future chemistry changes in acidified Nova Scotian Atlantic salmon (Salmo salar) rivers: a dynamic modeling approach. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 61(10):1965–1975, October 2004. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f04-196.

Stokesbury:2004:MAB

[1008] Michael J. W. Stokesbury, Steven L. H. Teo, Andrew Seitz, Ronald K. O’Dor, and Barbara A. Block. Movement of Atlantic bluefin tuna (Thunnus thynnus) as determined by satellite tagging experiments initiated off New England. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 61(10):1976– 1987, October 2004. CODEN CJFSDX. ISSN 0706-652X (print), 1205-

7533 (electronic). URL https://cdnsciencepub.com/doi/full/10.
1139/f04-130.

Ryves:2004:RSE

[1009] David B. Ryves, Annemarie L. Clarke, Peter G. Appleby, Susanne L. Amsinck, Erik Jeppesen, Frank Landkildehus, and N. John Anderson. Reconstructing the salinity and environment of the Limfjord and Vejlerne Nature Reserve, Denmark, using a diatom model for brackish lakes and fjords. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 61(10):1988–2006, October 2004. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/f04-
127.

Spanhoff:2004:SID

[1010] Bernd Spänhoff and Mark O. Gessner. Slow initial decomposition and fungal colonization of pine branches in a nutrient-rich lowland stream. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 61(10):2007–2013, October 2004. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f04-145.

Regnell:2004:CMT

[1011] O. Regnell and T. Hammar. Coupling of methyl and total mercury in a minerotrophic peat bog in southeastern Sweden. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 61(10):2014–2023, October 2004. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f04-143.

Armstrong:2004:RTP

[1012] Trip W. Armstrong and Roland A. Knapp. Response by trout populations in alpine lakes to an experimental halt to stocking. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 61(11):2025–2037, November 2004. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f04-144.

Vidussi:2004:STV

[1013] Francesca Vidussi, Suzanne Roy, Connie Lovejoy, Marie Gammelgaard, Helge Abildhauge Thomsen, Beatrice Booth, Jean-Eric Tremblay, and Behzad Mostajir. Spatial and temporal variability of the phytoplankton community structure in the North water polynya, investigated using pigment biomarkers. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 61(11): 2038–2052, November 2004. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f04-152.

Rodwell:2004:FIM

[1014] Lynda D. Rodwell and Callum M. Roberts. Fishing and the impact of marine reserves in a variable environment. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 61(11):2053–2068, November 2004. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f04-142.

Witting:2004:EEO

[1015] David A. Witting, R. Christopher Chambers, Keith L. Bosley, and Sam C. Wainright. Experimental evaluation of ontogenetic diet transitions in summer flounder (Paralichthys dentatus), using stable isotopes as diet tracers. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 61(11):2069– 2084, November 2004. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f04-156.

Buffington:2004:BSA

[1016] John M. Buffington, David R. Montgomery, and Harvey M. Greenberg. Basin-scale availability of salmonid spawning gravel as influenced by channel type and hydraulic roughness in mountain catchments. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 61(11):2085–2096, November 2004. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f04-141.

Trippel:2004:MSD

[1017] Edward A. Trippel and Steven R. E. Neil. Maternal and seasonal differences in egg sizes and spawning activity of northwest Atlantic haddock (Melanogrammus aeglefinus) in relation to body size and condition. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 61(11):2097–2110, November 2004. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f04-125.

Barbiero:2004:CCC

[1018] Richard P. Barbiero and Marc L. Tuchman. Changes in the crustacean communities of Lakes Michigan, Huron, and Erie following the invasion of the predatory cladoceran Bythotrephes longimanus. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 61(11):2111–2125, November 2004. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f04-149.

Zhang:2004:ULB

[1019] Z. Zhang, W. Hajas, A. Phillips, and J. A. Boutillier. Use of lengthbased models to estimate biological parameters and conduct yield analyses for male Dungeness crab (Cancer magister). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 61(11):2126–2134, November 2004. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f04-155.

Dorn:2004:ENC

[1020] Nathan J. Dorn and Gary G. Mittelbach. Effects of a native crayfish (Orconectes virilis) on the reproductive success and nesting behavior of sunfish (Lepomis spp.). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 61(11): 2135–2143, November 2004. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f04-158.

Wisenden:2004:PRT

[1021] Brian D. Wisenden, Josh Klitzke, Ryan Nelson, David Friedl, and Peter C. Jacobson. Predator-recognition training of hatchery-reared walleye (Stizostedion vitreum) and a field test of a training method using yellow perch (Perca flavescens). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 61(11): 2144–2150, November 2004. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f04-164.

Finstad:2004:BPS

[1022] Anders Gravbrøt Finstad and Ole Kristian Berg. Bimodal population size distributions and biased gillnet sampling. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 61(11):2151–2157, November 2004. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f04-157.

Bajer:2004:WCD

[1023] Przemyslaw G. Bajer, Gregory W. Whitledge, and Robert S. Hayward. Widespread consumption-dependent systematic error in fish bioenergetics models and its implications. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 61(11):2158–2167, November 2004. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f04-159.

Bajer:2004:SIC

[1024] Przemyslaw G. Bajer, Robert S. Hayward, Gregory W. Whitledge, and Richard D. Zweifel. Simultaneous identification and correction of systematic error in bioenergetics models: demonstration with a white crappie (Pomoxis annularis) model. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 61 (11):2168–2182, November 2004. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/

doi/full/10.1139/f04-160.

Crook:2004:MAH

[1025] David A. Crook. Movements associated with home-range establishment by two species of lowland river fish. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 61(11):2183–2193, November 2004. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f04-151.

Savenkoff:2004:IMT

[1026] Claude Savenkoff, Martin Castonguay, Alain F. Vézina, Simon-Pierre Despatie, Denis Chabot, Lyne Morissette, and Mike O. Hammill. Inverse modelling of trophic flows through an entire ecosystem: the northern Gulf of St. Lawrence in the mid-1980s. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 61(11):2194–2214, November 2004. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f04-154.

Essington:2004:PDF

[1027] Timothy E. Essington and Sture Hansson. Predator-dependent functional responses and interaction strengths in a natural food web. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 61(11):2215–2226, November 2004. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f04-146.

Lilja:2004:SEM

[1028] Juha Lilja, Timo J. Marjomäki, Juha Jurvelius, Tuomo Rossi, and Erkki Heikkola. Simulation and experimental measurement of side-aspect target strength of Atlantic salmon (Salmo salar) at high frequency. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 61(11):2227–2236, November 2004. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f04-166.

Hjellvik:2004:DVA

[1029] Vidar Hjellvik, Olav Rune Godø, and Dag Tjøstheim. Diurnal variation in acoustic densities: why do we see less in the dark? Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 61(11):2237–2254, November 2004. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f04-161.

Wilson:2004:LTR

[1030] Karen A. Wilson, John J. Magnuson, David M. Lodge, Anna M. Hill, Timothy K. Kratz, William L. Perry, and Theodore V. Willis. A longterm rusty crayfish (Orconectes rusticus) invasion: dispersal patterns and community change in a north temperate lake. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 61(11):2255–2266, November 2004. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f04-170.

Anonymous:2004:SAS

[1031] Anonymous. The status of Atlantic salmon: Populations and habitats / [Le statut du saumon atlantique : populations et habitats]. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 61(12), December 2004. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f05-900.

Boisclair:2004:SAS

[1032] Daniel Boisclair. The status of Atlantic salmon (Salmo salar): populations and habitats. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 61(12): 2267–2270, December 2004. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f05-028.

Lapointe:2004:IES

[1033] Michel F. Lapointe, Normand E. Bergeron, F. Bérubé, M.-A. Pouliot, and P. Johnston. Interactive effects of substrate sand and silt contents, redd-scale hydraulic gradients, and interstitial velocities on eggto-emergence survival of Atlantic salmon (Salmo salar). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 61(12):2271–2277, December 2004. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f04-236.

Youngson:2004:LRG

[1034] A. F. Youngson, I. A. Malcolm, J. L. Thorley, P. J. Bacon, and C. Soulsby. Long-residence groundwater effects on incubating salmonid eggs: low hyporheic oxygen impairs embryo development. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 61(12):2278–2287, December 2004. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f04-217.

Letcher:2004:LTC

[1035] Benjamin H. Letcher, Todd Dubreuil, Matthew J. O’Donnell, Mariska Obedzinski, Kitty Griswold, and Keith H. Nislow. Long-term consequences of variation in timing and manner of fry introduction on juvenile Atlantic salmon (Salmo salar) growth, survival, and life-history expression. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 61(12):2288–2301, December 2004. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f04-214.

Enders:2004:CHU

[1036] Eva C. Enders, Daniel Boisclair, and André G. Roy. The costs of habitat utilization of wild, farmed, and domesticated juvenile Atlantic salmon (Salmo salar). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 61(12):2302– 2313, December 2004. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f04-211.

Obedzinski:2004:VFG

[1037] Mariska Obedzinski and Benjamin H. Letcher. Variation in freshwater growth and development among five New England Atlantic salmon (Salmo salar) populations reared in a common environment. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 61(12):2314–2328, December 2004. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f05-005.

Maki-Petays:2004:SDJ

[1038] Aki Mäki-Petäys, Jaakko Erkinaro, Eero Niemelä, Ari Huusko, and Timo Muotka. Spatial distribution of juvenile Atlantic salmon (Salmo salar) in a subarctic river: size-specific changes in a strongly seasonal environment. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 61(12):2329–2338, December 2004. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f04-218.

Girard:2004:FGL

[1039] Isabelle L. Girard, James W. A. Grant, and Stefán Ó. Steingrímsson. Foraging, growth, and loss rate of young-of-the-year Atlantic salmon (Salmo salar) in relation to habitat use in Catamaran Brook, New Brunswick.

Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 61(12):2339–2349, December 2004. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f04-216.

Parrish:2004:WGS

[1040] Donna L. Parrish, Ethan J. Hawes, and Kevin G. Whalen. Winter growth and survival of juvenile Atlantic salmon (Salmo salar) in experimental raceways. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 61(12):2350–2357, December 2004. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f04-212.

Finstad:2004:ERJ

[1041] Anders G. Finstad, Ola Ugedal, Torbjørn Forseth, and Tor F. Næsje. Energy-related juvenile winter mortality in a northern population of Atlantic salmon (Salmo salar). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 61 (12):2358–2368, December 2004. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/

doi/full/10.1139/f04-213.

Jonsson:2004:FAM

[1042] Bror Jonsson and Nina Jonsson. Factors affecting marine production of Atlantic salmon (Salmo salar). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 61(12):2369–2383, December 2004. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f04-215.

Niemela:2004:TSV

[1043] Eero Niemelä, Jaakko Erkinaro, J. Brian Dempson, Markku Julkunen, Alexander Zubchenko, Sergei Prusov, Martin A. Svenning, Randi Ingvaldsen, Marianne Holm, and Esa Hassinen. Temporal synchrony and variation in abundance of Atlantic salmon (Salmo salar) in two subarctic Barents Sea rivers: influence of oceanic conditions. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 61(12):2384–2391, December 2004. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f04-208.

Juanes:2004:LTC

[1044] Francis Juanes, Stephen Gephard, and Kenneth F. Beland. Long-term changes in migration timing of adult Atlantic salmon (Salmo salar) at the southern edge of the species distribution. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 61(12):2392–2400, December 2004. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f04-207.

Nislow:2004:PFD

[1045] Keith H. Nislow, John D. Armstrong, and Simon McKelvey. Phosphorus flux due to Atlantic salmon (Salmo salar) in an oligotrophic upland stream: effects of management and demography. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 61(12):2401–2410, December 2004. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f05-006.

Andersen:2004:AMS

[1046] Niels G. Andersen and Jorgen Riis-Vestergaard. Alternative model structures for bioenergetics budgets of a cruising predatory gadoid: incorporating estimates of food conversion and costs of locomotion. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 61(12):2413–2424, December 2004. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f04-153.

Bacon:2004:MRH

[1047] Charles R. Bacon, Peter K. Weber, Kimberly A. Larsen, Reginald Reisenbichler, John A. Fitzpatrick, and Joseph L. Wooden. Migration and rearing histories of Chinook salmon (Oncorhynchus tshawytscha) determined by ion microprobe Sr isotope and Sr/Ca transects of otoliths. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 61(12):2425–2439, December 2004. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f04-167.

Racca:2004:TPD

[1048] Julien M. J. Racca, Irene Gregory-Eaves, Reinhard Pienitz, and Yves T. Prairie. Tailoring palaeolimnological diatom-based transfer functions. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 61(12):2440–2454, December 2004. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f04-162.

Holt:2004:LTT

[1049] Carrie A. Holt and Randall M. Peterman. Long-term trends in agespecific recruitment of sockeye salmon (Oncorhynchus nerka) in a changing environment. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 61(12):2455– 2470, December 2004. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f04-193.

Su:2004:SHB

[1050] Zhenming Su, Randall M. Peterman, and Steven L. Haeseker. Spatial hierarchical Bayesian models for stock-recruitment analysis of pink salmon (Oncorhynchus gorbuscha). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 61(12):2471–2486, December 2004. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f04-168.

Rindorf:2004:BEF

[1051] A. Rindorf and P. Lewy. Bias in estimating food consumption of fish by stomach-content analysis. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 61 (12):2487–2498, December 2004. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/

doi/full/10.1139/f04-200.

Bennett:2004:IRR

[1052] William A. Bennett, Kimy Roinestad, Laura Rogers-Bennett, Les Kaufman, Deb Wilson-Vandenberg, and Burr Heneman. Inverse regional responses to climate change and fishing intensity by the recreational rockfish (Sebastes spp.) fishery in California. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 61(12):2499–2510, December 2004. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f04-203.

Molot:2004:RAD

[1053] L. A. Molot, W. Keller, P. R. Leavitt, R. D. Robarts, M. J. Waiser, M. T. Arts, T. A. Clair, R. Pienitz, N. D. Yan, D. K. McNicol, Y. T. Prairie, P. J. Dillon, M. Macrae, R. Bello, R. N. Nordin, P. J. Curtis, J. P.

Smol, and M. S. V. Douglas. Risk analysis of dissolved organic mattermediated ultraviolet B exposure in Canadian inland waters. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 61(12):2511–2521, December 2004. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f04-165.

Brander:2004:EEN

[1054] K. Brander and R. Mohn. Erratum: Effect of the North Atlantic oscillation on recruitment of Atlantic cod (Gadus morhua). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 61(12):2522, December 2004. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f05-019.

Anonymous:2004:AII

[1055] Anonymous. Author index / index des auteurs: Volume 61, 2004. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 61(12), December 2004. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f05-902.

Anonymous:2004:CSV

[1056] Anonymous. Contents / sommaire: Volume 61, 2004. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 61(12), December 2004. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f05-903.

Frimpong:2005:DOR

[1057] Emmanuel A. Frimpong, Trent M. Sutton, Kyoung J. Lim, Peter J. Hrodey, Bernard A. Engel, Thomas P. Simon, John G. Lee, and Dennis C. Le Master. Determination of optimal riparian forest buffer dimensions for stream biota–landscape association models using multimetric and multivariate responses. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62 (1):1–6, January 2005. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f05-020.

Lamouroux:2005:GIH

[1058] Nicolas Lamouroux and Ian G. Jowett. Generalized instream habitat models. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62(1):7–14, January 2005. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f04-163.

Bunnell:2005:IIR

[1059] David B. Bunnell, Timothy B. Johnson, and Carey T. Knight. The impact of introduced round gobies (Neogobius melanostomus) on phosphorus cycling in central Lake Erie. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62(1):15–29, January 2005. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f04-172.

Naughton:2005:LSM

[1060] George P. Naughton, Christopher C. Caudill, Matthew L. Keefer, Theodore C. Bjornn, Lowell C. Stuehrenberg, and Christopher A. Peery. Late-season mortality during migration of radio-tagged adult sockeye salmon (Oncorhynchus nerka) in the Columbia River. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62(1):30–47, January 2005. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f04-147.

Kennedy:2005:CNO

[1061] Brian P. Kennedy, C. Page Chamberlain, Joel D. Blum, Keith H. Nislow, and Carol L. Folt. Comparing naturally occurring stable isotopes of nitrogen, carbon, and strontium as markers for the rearing locations of Atlantic salmon (Salmo salar). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62 (1):48–57, January 2005. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f04-184.

Wilzbach:2005:ERC

[1062] Margaret A. Wilzbach, Bret C. Harvey, Jason L. White, and Rodney J. Nakamoto. Effects of riparian canopy opening and salmon carcass addition on the abundance and growth of resident salmonids. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62(1):58–67, January 2005. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f04-177.

Zydlewski:2005:ECT

[1063] Gayle Barbin Zydlewski, Alex Haro, and Stephen D. McCormick. Evidence for cumulative temperature as an initiating and terminating factor in downstream migratory behavior of Atlantic salmon (Salmo salar) smolts. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62(1):68–78, January 2005. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f04-179.

Munro:2005:NCM

[1064] Andrew R. Munro, Thomas E. McMahon, and James R. Ruzycki. Natural chemical markers identify source and date of introduction of an exotic species: lake trout (Salvelinus namaycush) in Yellowstone Lake. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62(1):79–87, January 2005. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f04-174.

Zimmerman:2005:ROS

[1065] Christian E. Zimmerman. Relationship of otolith strontium-to-calcium ratios and salinity: experimental validation for juvenile salmonids. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62(1):88–97, January 2005. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f04-182.

Alcala:2005:LTS

[1066] Angel C. Alcala, Garry R. Russ, Aileen P. Maypa, and Hilconida P. Calumpong. A long-term, spatially replicated experimental test of the effect of marine reserves on local fish yields. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62(1):98–108, January 2005. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f04-176.

Davis:2005:APS

[1067] Jana L. D. Davis, Alicia C. Young-Williams, Anson H. Hines, and Yonathan Zohar. Assessing the potential for stock enhancement in the case of the Chesapeake Bay blue crab (Callinectes sapidus). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62(1):109–122, January 2005. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f04-150.

LeLoch:2005:SCN

[1068] François Le Loc’h and Christian Hily. Stable carbon and nitrogen isotope analysis of Nephrops norvegicus / Merluccius merluccius fishing grounds in the Bay of Biscay (Northeast Atlantic). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62(1):123–132, January 2005. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f04-242.

Zamon:2005:RSZ

[1069] Jeannette E. Zamon and David W. Welch. Rapid shift in zooplankton community composition on the northeast Pacific shelf during the 1998– 1999 El Niño–La Niña event. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62 (1):133–144, January 2005. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f04-171.

Stetter:2005:ECD

[1070] S. L. Parker Stetter, L. D. Witzel, L. G. Rudstam, D. W. Einhouse, and E. L. Mills. Energetic consequences of diet shifts in Lake Erie rainbow smelt (Osmerus mordax). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62 (1):145–152, January 2005. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f04-175.

Kovecses:2005:IAB

[1071] J. Kövecses, G. D. Sherwood, and J. B. Rasmussen. Impacts of altered benthic invertebrate communities on the feeding ecology of yellow perch (Perca flavescens) in metal-contaminated lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62(1):153–162, January 2005. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f04-181.

Pont:2005:MHR

[1072] D. Pont, B. Hugueny, and T. Oberdorff. Modelling habitat requirement of European fishes: do species have similar responses to local and regional environmental constraints? Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62 (1):163–173, January 2005. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f04-183.

Bowman:2005:EAA

[1073] Michelle F. Bowman, Patricia A. Chambers, and David W. Schindler. Epilithic algal abundance in relation to anthropogenic changes in phosphorus bioavailability and limitation in mountain rivers. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62(1):174–184, January 2005. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f04-244.

Rier:2005:CCL

[1074] Steven T. Rier, Nancy C. Tuchman, and Robert G. Wetzel. Chemical changes to leaf litter from trees grown under elevated CO_{2} and the implications for microbial utilization in a stream ecosystem. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62(1):185–194, January 2005. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f04-148.

Aas-Hansen:2005:RWA

[1075] Øyvind Aas-Hansen, Mathilakath M. Vijayan, Helge K. Johnsen, Colin Cameron, and Even H. Jørgensen. Resmoltification in wild, anadromous Arctic char (Salvelinus alpinus): a survey of osmoregulatory, metabolic, and endocrine changes preceding annual seawater migration. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62(1):195–204, January 2005. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f04-186.

Edwards:2005:IBF

[1076] William J. Edwards, Chris R. Rehmann, Ellen McDonald, and David A. Culver. The impact of a benthic filter feeder: limitations imposed by physical transport of algae to the benthos. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62(1):205–214, January 2005. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f04-188.

Taylor:2005:NLS

[1077] Nathan G. Taylor, Carl J. Walters, and Steven J. D. Martell. A new likelihood for simultaneously estimating von Bertalanffy growth parameters, gear selectivity, and natural and fishing mortality. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62(1):215–223, January 2005. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f04-189. See corrigendum

[2162].

Rochet:2005:FVD

[1078] Marie-Joëlle Rochet and Verena M. Trenkel. Factors for the variability of discards: assumptions and field evidence. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62(1):224–235, January 2005. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f04-185.

Blais:2005:BPB

[1079] J. M. Blais. Biogeochemistry of persistent bioaccumulative toxicants: processes affecting the transport of contaminants to remote areas. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62(1):236–243, January 2005. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f04-226.

Malinen:2005:HFS

[1080] Tommi Malinen, Antti Tuomaala, and Heikki Peltonen. Hydroacoustic fish stock assessment in the presence of dense aggregations of Chaoborus larvae. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62(2):245–249, February 2005. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f05-038.

Ernst:2005:SDF

[1081] Billy Ernst, J. M. (Lobo) Orensanz, and David A. Armstrong. Spatial dynamics of female snow crab (Chionoecetes opilio) in the eastern Bering Sea. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62(2):250–268, February 2005. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f04-201.

Cox:2005:NEP

[1082] M. Keith Cox and Kyle J. Hartman. Nonlethal estimation of proximate composition in fish. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62(2): 269–275, February 2005. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f04-180.

Sliva:2005:ERS

[1083] Lucie Sliva and D. Dudley Williams. Exploration of riffle-scale interactions between abiotic variables and microbial assemblages in the hyporheic zone. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62(2):276–290, February 2005. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f04-190.

Mantyniemi:2005:BRE

[1084] Samu Mäntyniemi, Atso Romakkaniemi, and Elja Arjas. Bayesian removal estimation of a population size under unequal catchability. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62(2):291–300, February 2005. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f04-195.

Johnson:2005:BLS

[1085] Brent R. Johnson and J. Bruce Wallace. Bottom-up limitation of a stream salamander in a detritus-based food web. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62(2):301–311, February 2005. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f04-197.

Kaste:2005:IMP

[1086] Øyvind Kaste and Anne Lyche-Solheim. Influence of moderate phosphate addition on nitrogen retention in an acidic boreal lake. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62(2):312–321, February 2005. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f04-233.

Donali:2005:PRH

[1087] Espen Donali, Pål Brettum, Øyvind Kaste, Jarl Eivind Løvik, Anne Lyche-Solheim, and Tom Andersen. Pelagic response of a humic lake to three years of phosphorus addition. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62(2):322–332, February 2005. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f04-234.

Anderson:2005:RFW

[1088] Caroline Anderson and Gilbert Cabana. δ ^{15}N in riverine food webs: effects of N inputs from agricultural watersheds. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62(2):333–340, February 2005. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f04-191.

Keller:2005:LBW

[1089] W. (Bill) Keller, Jocelyne Heneberry, and Julie Leduc. Linkages between weather, dissolved organic carbon, and cold-water habitat in a Boreal Shield lake recovering from acidification. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62(2):341–347, February 2005. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f04-192.

Pinto-Coelho:2005:CZL

[1090] Ricardo Pinto-Coelho, Bernadette Pinel-Alloul, Ginette Méthot, and Karl E. Havens. Crustacean zooplankton in lakes and reservoirs of temperate and tropical regions: variation with trophic status. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62(2):348–361, February 2005. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f04-178.

Lee:2005:MCE

[1091] Yong-Woo Lee and David B. Sampson. Monte Carlo evaluation of a nonlinear regression estimator for aggregated length–weight data. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62(2):362–373, February 2005. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f04-194.

Goodman:2005:SEH

[1092] Daniel Goodman. Selection equilibrium for hatchery and wild spawning fitness in integrated breeding programs. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62(2):374–389, February 2005. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://

cdnsciencepub.com/doi/full/10.1139/f04-187.

Kamunde:2005:IDS

[1093] Collins N. Kamunde, Soumya Niyogi, and Chris M. Wood. Interaction of dietary sodium chloride and waterborne copper in rainbow trout (Oncorhynchus mykiss): copper toxicity and sodium and chloride homeostasis. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62(2):390–399, February 2005. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f04-169.

Eckley:2005:MMH

[1094] C. S. Eckley, C. J. Watras, H. Hintelmann, K. Morrison, A. D. Kent, and O. Regnell. Mercury methylation in the hypolimnetic waters of lakes with and without connection to wetlands in northern Wisconsin. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62(2):400–411, February 2005. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f04-205.

Johnson:2005:EIL

[1095] Steven L. Johnson, Jeffrey D. Rodgers, Mario F. Solazzi, and Thomas E. Nickelson. Effects of an increase in large wood on abundance and survival of juvenile salmonids (Oncorhynchus spp.) in an Oregon coastal stream. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62(2):412–424, February 2005. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f04-222.

Lika:2005:MFP

[1096] Konstadia Lika and Nikos Papandroulakis. Modeling feeding processes: a test of a new model for sea bream (Sparus aurata) larvae. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62(2):425–435, February 2005. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f04-198.

Radomski:2005:CVP

[1097] Paul Radomski, James R. Bence, and Terrance J. Quinn II. Comparison of virtual population analysis and statistical kill-at-age analysis for a recreational, kill-dominated fishery. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62(2):436–452, February 2005. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f04-228.

Garcia-Berthou:2005:IPE

[1098] E. García-Berthou, C. Alcaraz, Q. Pou-Rovira, L. Zamora, G. Coenders, and C. Feo. Introduction pathways and establishment rates of invasive aquatic species in Europe. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62 (2):453–463, February 2005. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f05-017.

Paterson:2005:SMT

[1099] Gordon Paterson, Susan Y. Huestis, D. Michael Whittle, Kenneth G. Drouillard, and G. Douglas Haffner. In situ measurement of tissue turnover and energy conversion efficiencies in lake trout (Salvelinus namaycush) using a novel toxicokinetic approach. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62(2):464–471, February 2005. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f04-206.

Klug:2005:BRD

[1100] Jennifer L. Klug. Bacterial response to dissolved organic matter affects resource availability for algae. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62 (2):472–481, February 2005. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f04-229.

Bronte:2005:EFC

[1101] Charles R. Bronte, Mark P. Ebener, Donald R. Schreiner, David S. DeVault, Michael M. Petzold, Douglas A. Jensen, Carl Richards, and Steven J. Lozano. Erratum: Fish community change in Lake Superior, 1970–2000. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62(2):482, February 2005. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f05-054.

Persaud:2005:DDT

[1102] A. D. Persaud and N. D. Yan. Developmental differences and a test for reciprocity in the tolerance of Chaoborus punctipennis larvae to ultraviolet radiation. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62(3):483–491, March 2005. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f04-204.

VanNieuwenhuyse:2005:EMP

[1103] Erwin E. Van Nieuwenhuyse. Empirical model for predicting a catchment-scale metric of surface water transit time in streams. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62(3):492–504, March 2005. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f04-202.

Cristescu:2005:CSE

[1104] Melania E. A. Cristescu and Paul D. N. Hebert. The “Crustacean Seas” — an evolutionary perspective on the Ponto–Caspian peracarids. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62(3):505–517, March 2005. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f04-210.

Koyama:2005:MNC

[1105] Akihiro Koyama, Kathleen Kavanagh, and Andrew Robinson. Marine nitrogen in central Idaho riparian forests: evidence from stable isotopes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62(3):518–526, March 2005. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f04-220.

Baumgartner:2005:SFH

[1106] Mark F. Baumgartner and Bruce R. Mate. Summer and fall habitat of North Atlantic right whales (Eubalaena glacialis) inferred from satellite telemetry. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62(3):527–543, March 2005. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f04-238.

Brodziak:2005:MAE

[1107] Jon Brodziak and Christopher M. Legault. Model averaging to estimate rebuilding targets for overfished stocks. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62(3):544–562, March 2005. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f04-199.

Giguere:2005:MBO

[1108] Anik Giguère, Peter G. C. Campbell, Landis Hare, and Carole Cossu-Leguille. Metal bioaccumulation and oxidative stress in yellow perch (Perca flavescens) collected from eight lakes along a metal contamination gradient (Cd, Cu, Zn, Ni). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62 (3):563–577, March 2005. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f04-224.

Winkler:2005:IET

[1109] Gesche Winkler, Pascal Sirois, Ladd E. Johnson, and Julian J. Dodson. Invasion of an estuarine transition zone by Dreissena polymorpha veligers had no detectable effect on zooplankton community structure. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62(3):578–592, March 2005. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f04-230.

Zeldis:2005:PFP

[1110] John R. Zeldis, John Oldman, Sira L. Ballara, and Laura A. Richards. Physical fluxes, pelagic ecosystem structure, and larval fish survival in Hauraki Gulf, New Zealand. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62 (3):593–610, March 2005. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f04-209.

Deng:2005:CVM

[1111] Roy Deng, Cathy Dichmont, David Milton, Mick Haywood, David Vance, Natasha Hall, and David Die. Can vessel monitoring system data also be used to study trawling intensity and population depletion? The example of Australia’s northern prawn fishery. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62(3):611–622, March 2005. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f04-219.

Hamer:2005:CTO

[1112] Paul A. Hamer, Gregory P. Jenkins, and Bronwyn M. Gillanders. Chemical tags in otoliths indicate the importance of local and distant settlement areas to populations of a temperate sparid, Pagrus auratus. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62(3):623–630, March 2005. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f04-221.

Branch:2005:ETG

[1113] Trevor A. Branch, Ray Hilborn, and Eugenia Bogazzi. Escaping the tyranny of the grid: a more realistic way of defining fishing opportunities. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62(3):631–642, March 2005. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f04-232.

Alvarez:2005:MRR

[1114] D. Álvarez and A. G. Nicieza. Is metabolic rate a reliable predictor of growth and survival of brown trout (Salmo trutta) in the wild? Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62(3):643–649, March 2005. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f04-223.

Harvey:2005:HSB

[1115] Bret C. Harvey, Jason L. White, and Rodney J. Nakamoto. Habitatspecific biomass, survival, and growth of rainbow trout (Oncorhynchus mykiss) during summer in a small coastal stream. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62(3):650–658, March 2005. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f04-225.

Mello:2005:UGQ

[1116] L. G. S. Mello and G. A. Rose. Using geostatistics to quantify seasonal distribution and aggregation patterns of fishes: an example of Atlantic cod (Gadus morhua). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62 (3):659–670, March 2005. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f04-227.

Suzuki:2005:DIE

[1117] Keita W. Suzuki, Akihide Kasai, Kouji Nakayama, and Masaru Tanaka. Differential isotopic enrichment and half-life among tissues in Japanese temperate bass (Lateolabrax japonicus) juveniles: implications for analyzing migration. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62(3):671–678, March 2005. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f04-231.

Begg:2005:MLH

[1118] Gavin A. Begg, Bruce D. Mapstone, Ashley J. Williams, Samantha Adams, Campbell R. Davies, and Dong C. Lou. Multivariate lifehistory indices of exploited coral reef fish populations used to measure the performance of no-take zones in a marine protected area. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62(3):679–692, March 2005. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f04-237.

Kalinowski:2005:LBC

[1119] Steven T. Kalinowski and Mark L. Taper. Likelihood-based confidence intervals of relative fitness for a common experimental design. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62(3):693–699, March 2005. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f04-239.

Wurster:2005:THS

[1120] C. M. Wurster, W. P. Patterson, D. J. Stewart, J. N. Bowlby, and T. J. Stewart. Thermal histories, stress, and metabolic rates of Chinook salmon (Oncorhynchus tshawytscha) in Lake Ontario: evidence from intra-otolith stable isotope analyses. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62(3):700–713, March 2005. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f04-241.

Okada:2005:STG

[1121] Edson Kiyoshi Okada, Angelo Antonio Agostinho, and Luiz Carlos Gomes. Spatial and temporal gradients in artisanal fisheries of a large neotropical reservoir, the Itaipu Reservoir, Brazil. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62(3):714–724, March 2005. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f05-015.

Anonymous:2005:BBL

[1122] Anonymous. Building on Beverton’s legacy: life history variation and fisheries management / [Construire sur l’héritage de Beverton : variation du cycle biologique et gestion des pêches]. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62(4), April 2005. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f05-906.

Shuter:2005:ISB

[1123] Brian J. Shuter and Peter A. Abrams. Introducing the symposium “Building on Beverton’s legacy: life history variation and fisheries management”. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62(4):725–729, April 2005. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f05-067.

Conover:2005:DFS

[1124] David O. Conover, Stephen A. Arnott, Matthew R. Walsh, and Stephan B. Munch. Darwinian fishery science: lessons from the Atlantic silverside (Menidia menidia). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62 (4):730–737, April 2005. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f05-069.

Shuter:2005:OLH

[1125] B. J. Shuter, N. P. Lester, J. LaRose, C. F. Purchase, K. Vascotto, G. Morgan, N. C. Collins, and P. A. Abrams. Optimal life histories and food web position: linkages among somatic growth, reproductive investment, and mortality. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62 (4):738–746, April 2005. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f05-070.

Magnan:2005:IEF

[1126] Pierre Magnan, Raphaël Proulx, and Michel Plante. Integrating the effects of fish exploitation and interspecific competition into current life history theories: an example with lacustrine brook trout (Salvelinus fontinalis) populations. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62 (4):747–757, April 2005. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f05-041.

Abrams:2005:EAC

[1127] Peter A. Abrams and Hiroyuki Matsuda. The effect of adaptive change in the prey on the dynamics of an exploited predator population. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62(4):758–766, April 2005. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f05-051.

Rennie:2005:CME

[1128] Michael D. Rennie, Nicholas C. Collins, Brian J. Shuter, James W. Rajotte, and Patrice Couture. A comparison of methods for estimating activity costs of wild fish populations: more active fish observed to grow slower. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62(4):767–780, April 2005. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f05-052.

DeAngelis:2005:LHT

[1129] Donald L. DeAngelis, Joel C. Trexler, and William F. Loftus. Life history trade-offs and community dynamics of small fishes in a seasonally pulsed wetland. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62(4):781–790, April 2005. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f05-050.

Reznick:2005:CCF

[1130] David N. Reznick and Cameron K. Ghalambor. Can commercial fishing cause evolution? Answers from guppies (Poecilia reticulata). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62(4):791–801, April 2005. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f05-079.

Munch:2005:HSG

[1131] S. B. Munch, M. R. Walsh, and D. O. Conover. Harvest selection, genetic correlations, and evolutionary changes in recruitment: one less thing to worry about? Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62(4):802–810, April 2005. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/f05-
078.

Olsen:2005:ACA

[1132] Esben Moland Olsen, George R. Lilly, Mikko Heino, M. Joanne Morgan, John Brattey, and Ulf Dieckmann. Assessing changes in age and size at maturation in collapsing populations of Atlantic cod (Gadus morhua). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62(4):811–823, April 2005. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f05-065.

Hutchings:2005:LHC

[1133] Jeffrey A. Hutchings. Life history consequences of overexploitation to population recovery in Northwest Atlantic cod (Gadus morhua). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62(4):824–832, April 2005. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f05-081.

Rijnsdorp:2005:FIA

[1134] A. D. Rijnsdorp, R. E. Grift, and S. B. M. Kraak. Fisheries-induced adaptive change in reproductive investment in North Sea plaice (Pleuronectes platessa)? Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62(4):833–843, April 2005. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f05-039.

Dunlop:2005:IIG

[1135] Erin S. Dunlop, Brian J. Shuter, and Mark S. Ridgway. Isolating the influence of growth rate on maturation patterns in the smallmouth bass (Micropterus dolomieu). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62 (4):844–853, April 2005. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f05-045.

Reynolds:2005:LHE

[1136] John D. Reynolds, Thomas J. Webb, and Lorraine A. Hawkins. Life history and ecological correlates of extinction risk in European freshwater fishes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62(4):854–862, April 2005. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f05-066.

Gerber:2005:RDD

[1137] Leah R. Gerber, Selina S. Heppell, Ford Ballantyne, and Enric Sala. The role of dispersal and demography in determining the efficacy of marine reserves. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62(4):863–871, April 2005. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f05-046.

Winemiller:2005:LHS

[1138] Kirk O. Winemiller. Life history strategies, population regulation, and implications for fisheries management. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62(4):872–885, April 2005. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f05-040.

Rose:2005:LRB

[1139] Kenneth A. Rose. Lack of relationship between simulated fish population responses and their life history traits: inadequate models, incorrect analysis, or site-specific factors? Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62 (4):886–902, April 2005. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f05-049.

Kaplan:2005:EVS

[1140] David M. Kaplan and Louis W. Botsford. Effects of variability in spacing of coastal marine reserves on fisheries yield and sustainability. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62(4):905–912, April 2005. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f04-243.

Polivka:2005:RMA

[1141] Karl M. Polivka. Resource matching across habitats is limited by competition at patch scales in an estuarine-opportunist fish. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62(4):913–924, April 2005. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f04-235.

Minte-Vera:2005:PAM

[1142] C. V. Minte-Vera, T. A. Branch, I. J. Stewart, and M. W. Dorn. Practical application of meta-analysis results: avoiding the double use of data. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62(4):925–929, April 2005. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f04-245.

Keefer:2005:EHU

[1143] Matthew L. Keefer, Christopher A. Peery, William R. Daigle, Michael A. Jepson, Steven R. Lee, Charles T. Boggs, Kenneth R. Tolotti, and Brian J. Burke. Escapement, harvest, and unknown loss of radio-tagged adult salmonids in the Columbia River–Snake River hydrosystem. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62(4):930–949, April 2005. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f04-246.

Colautti:2005:CIS

[1144] Robert I. Colautti. Are characteristics of introduced salmonid fishes biased by propagule pressure? Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62 (4):950–959, April 2005. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f05-002.

Scheuerell:2005:NPI

[1145] Mark D. Scheuerell, Phillip S. Levin, Richard W. Zabel, John G. Williams, and Beth L. Sanderson. A new perspective on the importance of marine-derived nutrients to threatened stocks of Pacific salmon (Oncorhynchus spp.). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62(5):961–964, May 2005. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/f05-
113.

Watson:2005:FMR

[1146] John W. Watson, Sheryan P. Epperly, Arvind K. Shah, and Daniel G. Foster. Fishing methods to reduce sea turtle mortality associated with pelagic longlines. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62(5):965–981, May 2005. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/f05-
004.

Gravel:2005:DHP

[1147] A. Gravel, P. G. C. Campbell, and A. Hontela. Disruption of the hypothalamo-pituitary-interrenal axis in 1+ yellow perch (Perca flavescens) chronically exposed to metals in the environment. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62(5):982–990, May 2005. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f05-011.

Turner:2005:DIE

[1148] Michael A. Turner, David B. Huebert, David L. Findlay, Leonard L. Hendzel, Wolfgang A. Jansen, R. A. (Drew) Bodaly, Llwellyn M. Armstrong, and Susan E. M. Kasian. Divergent impacts of experimental lake-level drawdown on planktonic and benthic plant communities in a boreal forest lake. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62 (5):991–1003, May 2005. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f05-003.

Trevorrow:2005:UMI

[1149] Mark V. Trevorrow. The use of moored inverted echo sounders for monitoring meso-zooplankton and fish near the ocean surface. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62(5):1004–1018, May 2005. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f05-013.

Wu:2005:PIC

[1150] F. C. Wu, R. B. Mills, Y. R. Cai, R. D. Evans, and P. J. Dillon. Photodegradation-induced changes in dissolved organic matter in acidic waters. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62(5):1019–1027, May 2005. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f05-009.

Millar:2005:ACS

[1151] Russell B. Millar and Wayne S. Stewart. Automatic calculation of the sensitivity of Bayesian fisheries models to informative priors. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62(5):1028–1036, May 2005. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f04-240.

Folkvord:2005:CSA

[1152] A. Folkvord. Comparison of size-at-age of larval Atlantic cod (Gadus morhua) from different populations based on size- and temperaturedependent growth models. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62 (5):1037–1052, May 2005. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f05-008.

Drake:2005:RAS

[1153] John M. Drake. Risk analysis for species introductions: forecasting population growth of Eurasian ruffe (Gymnocephalus cernuus). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62(5):1053–1059, May 2005. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f05-014.

Piner:2005:AVC

[1154] Kevin R. Piner, Owen S. Hamel, Jennifer L. Menkel, John R. Wallace, and Charles E. Hutchinson. Age validation of canary rockfish (Sebastes pinniger) from off the Oregon coast (USA) using the bomb radiocarbon method. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62(5):1060–1066, May 2005. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f05-082.

Das:2005:ISC

[1155] Biplob Das, Rolf D. Vinebrooke, Arturo Sanchez-Azofeifa, Benoit Rivard, and Alexander P. Wolfe. Inferring sedimentary chlorophyll concentrations with reflectance spectroscopy: a novel approach to reconstructing historical changes in the trophic status of mountain lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62(5):1067–1078, May 2005. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f05-016.

Enders:2005:MTS

[1156] Eva C. Enders, Daniel Boisclair, and André G. Roy. A model of total swimming costs in turbulent flow for juvenile Atlantic salmon (Salmo salar). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62(5):1079–1089, May 2005. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f05-007.

Bailey:2005:IRS

[1157] Sarah A. Bailey, Ian C. Duggan, Philip T. Jenkins, and Hugh J. MacIsaac. Invertebrate resting stages in residual ballast sediment of transoceanic ships. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62(5):1090– 1103, May 2005. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f05-024.

Ainsworth:2005:IVF

[1158] C. H. Ainsworth and U. R. Sumaila. Intergenerational valuation of fisheries resources can justify long-term conservation: a case study in Atlantic cod (Gadus morhua). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62 (5):1104–1110, May 2005. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f05-010.

Palmer:2005:CIA

[1159] M. E. Palmer and Anthony Ricciardi. Community interactions affecting the relative abundances of native and invasive amphipods in the St. Lawrence River. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62 (5):1111–1118, May 2005. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f05-012.

Garton:2005:FDT

[1160] David W. Garton, Christopher D. Payne, and Joseph P. Montoya. Flexible diet and trophic position of dreissenid mussels as inferred from stable isotopes of carbon and nitrogen. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62(5):1119–1129, May 2005. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f05-025.

Ward:2005:IDD

[1161] Peter Ward and Ransom A. Myers. Inferring the depth distribution of catchability for pelagic fishes and correcting for variations in the depth of longline fishing gear. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62 (5):1130–1142, May 2005. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f05-021.

Elsdon:2005:ALH

[1162] Travis S. Elsdon and Bronwyn M. Gillanders. Alternative life-history patterns of estuarine fish: barium in otoliths elucidates freshwater residency. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62(5):1143–1152, May 2005. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f05-029.

Weir:2005:SBS

[1163] Laura K. Weir, Jeffrey A. Hutchings, Ian A. Fleming, and Sigurd Einum. Spawning behaviour and success of mature male Atlantic salmon (Salmo salar) parr of farmed and wild origin. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62(5):1153–1160, May 2005. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f05-032.

Campbell:2005:EBR

[1164] Linda M. Campbell, Aaron T. Fisk, Xiaowa Wang, Günter Köck, and Derek C. G. Muir. Evidence for biomagnification of rubidium in freshwater and marine food webs. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62 (5):1161–1167, May 2005. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f05-027.

Baldigo:2005:SAM

[1165] Barry P. Baldigo, Peter S. Murdoch, and Douglas A. Burns. Stream acidification and mortality of brook trout (Salvelinus fontinalis) in response to timber harvest in Catskill Mountain watersheds, New York, USA. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62(5):1168–1183, May 2005. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f05-022.

Pihl:2005:RPP

[1166] Leif Pihl, Johan Modin, and Håkan Wennhage. Relating plaice (Pleuronectes platessa) recruitment to deteriorating habitat quality: effects of macroalgal blooms in coastal nursery grounds. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62(5):1184–1193, May 2005. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f05-023.

Kaiser:2005:MPA

[1167] Michel J. Kaiser. Are marine protected areas a red herring or fisheries panacea? Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62(5):1194–1199, May 2005. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f05-056.

Anonymous:2005:ACa

[1168] Anonymous. Announcement / Communiqué. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62(6), June 2005. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/

doi/full/10.1139/f05-910.

Holmqvist:2005:LTS

[1169] Niklas Holmqvist, Patrik Stenroth, Olof Berglund, Per Nyström, Wilhelm Graneli, and Per Larsson. Lake trophic status and bioaccumulation of polychlorinated biphenyls in benthic organisms: a comparison between littoral and profundal invertebrates. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62(6):1201–1209, June 2005. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f05-018.

Milne:2005:SFE

[1170] S. W. Milne, B. J. Shuter, and W. G. Sprules. The schooling and foraging ecology of lake herring (Coregonus artedi) in Lake Opeongo, Ontario, Canada. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62(6):1210–1218, June 2005. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f05-030.

Snover:2005:IEB

[1171] Melissa L. Snover, George M. Watters, and Marc Mangel. Interacting effects of behavior and oceanography on growth in salmonids with examples for Coho salmon (Oncorhynchus kisutch). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62(6):1219–1230, June 2005. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f05-058.

Lobon-Cervia:2005:STV

[1172] Javier Lobón-Cerviá. Spatial and temporal variation in the influence of density dependence on growth of stream-living brown trout (Salmo trutta). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62(6):1231–1242, June 2005. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f05-034.

Bartlett:2005:TBT

[1173] Adrienne J. Bartlett, Uwe Borgmann, D. George Dixon, Suzanne P. Batchelor, and R. James Maguire. Toxicity and bioaccumulation of tributyltin in Hyalella azteca from freshwater harbour sediments in the Great Lakes Basin, Canada. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62(6): 1243–1253, June 2005. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f05-035.

Brazner:2005:RWS

[1174] John C. Brazner, Danny K. Tanner, Naomi E. Detenbeck, Sharon L. Batterman, Stacey L. Stark, Leslie A. Jagger, and Virginia M. Snarski. Regional, watershed, and site-specific environmental influences on fish assemblage structure and function in western Lake Superior tributaries. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62(6):1254–1270, June 2005. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f05-031.

Grant:2005:MAB

[1175] Jon Grant, Peter Cranford, Barry Hargrave, Michel Carreau, Bryan Schofield, Shelley Armsworthy, Victoria Burdett-Coutts, and Diego Ibarra. A model of aquaculture biodeposition for multiple estuaries and field validation at blue mussel (Mytilus edulis) culture sites in eastern Canada. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62(6):1271–1285, June 2005. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f05-033.

Boeing:2005:CBD

[1176] Wiebke J. Boeing, Björn Wissel, and Charles W. Ramcharan. Costs and benefits of Daphnia defense against Chaoborus in nature. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62(6):1286–1294, June 2005. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f05-043.

Craig:2005:SDB

[1177] J. Kevin Craig, Larry B. Crowder, and Tyrrell A. Henwood. Spatial distribution of brown shrimp (Farfantepenaeus aztecus) on the northwestern Gulf of Mexico shelf: effects of abundance and hypoxia. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62(6):1295–1308, June 2005. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f05-036.

Carr:2005:PWQ

[1178] Geneviève M. Carr, Patricia A. Chambers, and Antoine Morin. Periphyton, water quality, and land use at multiple spatial scales in Alberta rivers. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62(6):1309–1319, June 2005. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f05-044.

Martell:2005:IPP

[1179] Steven J. D. Martell, Timothy E. Essington, Bob Lessard, James F. Kitchell, Carl J. Walters, and Christofer H. Boggs. Interactions of productivity, predation risk, and fishing effort in the efficacy of marine protected areas for the central Pacific. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62(6):1320–1336, June 2005. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f05-114.

Fuiman:2005:BRS

[1180] Lee A. Fuiman, James H. Cowan, Jr., Michael E. Smith, and Jonathan P. O’Neal. Behavior and recruitment success in fish larvae: variation with growth rate and the batch effect. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62(6):1337–1349, June 2005. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f05-053.

Robinson:2005:OCA

[1181] Clifford L. K. Robinson, John Morrison, and Michael G. G. Foreman. Oceanographic connectivity among marine protected areas on the north coast of British Columbia, Canada. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62(6):1350–1362, June 2005. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f05-088.

Lacroix:2005:DAS

[1182] Gilles L. Lacroix and Derek Knox. Distribution of Atlantic salmon (Salmo salar) postsmolts of different origins in the Bay of Fundy and Gulf of Maine and evaluation of factors affecting migration, growth, and survival. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62(6):1363–1376, June 2005. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f05-055.

Hirst:2005:ECA

[1183] David Hirst, Geir Storvik, Magne Aldrin, Sondre Aanes, and Ragnar Bang Huseby. Estimating catch-at-age by combining data from different sources. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62(6):1377–1385, June 2005. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/f05-
026.

Devlin:2005:VCD

[1184] R. H. Devlin, L. Park, D. M. Sakhrani, J. D. Baker, A. R. Marshall, E. LaHood, S. E. Kolesar, M. R. Mayo, C. A. Biagi, and M. Uh. Variation of Y-chromosome DNA markers in Chinook salmon (Oncorhynchus tshawytscha) populations. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62 (6):1386–1399, June 2005. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f05-048.

Riley:2005:AFH

[1185] Stephen C. Riley, Christopher P. Tatara, and Julie A. Scheurer. Aggression and feeding of hatchery-reared and naturally reared steelhead (Oncorhynchus mykiss) fry in a laboratory flume and a comparison with observations in natural streams. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62 (6):1400–1409, June 2005. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f05-076.

Waugh:2005:FRA

[1186] Susan Waugh, Dominique Filippi, Akira Fukuda, Makio Suzuki, Hiroyoshi Higuchi, Alvin Setiawan, and Lloyd Davis. Foraging of royal albatrosses, Diomedea epomophora, from the Otago Peninsula and its relationships to fisheries. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62 (6):1410–1421, June 2005. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f05-001.

VanderZanden:2005:PTC

[1187] M. Jake Vander Zanden, Timothy E. Essington, and Yvonne Vadeboncoeur. Is pelagic top-down control in lakes augmented by benthic energy pathways? Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62(6):1422–1431, June 2005. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/f05-
042.

Dobiesz:2005:ELH

[1188] Norine E. Dobiesz, David A. McLeish, Randy L. Eshenroder, James R. Bence, Lloyd C. Mohr, Mark P. Ebener, Thomas F. Nalepa, Aaron P. Woldt, James E. Johnson, Ray L. Argyle, and Joseph C. Makarewicz. Ecology of the Lake Huron fish community, 1970–1999. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62(6):1432–1451, June 2005. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f05-061.

Anonymous:2005:ACb

[1189] Anonymous. Announcement / Communiqué. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62(7), July 2005. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/

doi/full/10.1139/f05-911.

Bundy:2005:SFE

[1190] Alida Bundy. Structure and functioning of the eastern Scotian Shelf ecosystem before and after the collapse of groundfish stocks in the early 1990s. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62(7):1453–1473, July 2005. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f05-085.

Bundy:2005:CAC

[1191] Alida Bundy and L. Paul Fanning. Can Atlantic cod (Gadus morhua) recover? Exploring trophic explanations for the non-recovery of the cod stock on the eastern Scotian Shelf, Canada. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62(7):1474–1489, July 2005. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f05-086.

Koen-Alonso:2005:MMS

[1192] Mariano Koen-Alonso and Peter Yodzis. Multispecies modelling of some components of the marine community of northern and central Patagonia, Argentina. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62(7):1490–1512, July 2005. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/f05-
087.

Deng:2005:EFI

[1193] Zhiqun Deng, Gregory R. Guensch, Craig A. McKinstry, Robert P. Mueller, Dennis D. Dauble, and Marshall C. Richmond. Evaluation of fish-injury mechanisms during exposure to turbulent shear flow. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62(7):1513–1522, July 2005. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f05-091.

Keeley:2005:EDN

[1194] E. R. Keeley, E. A. Parkinson, and E. B. Taylor. Ecotypic differentiation of native rainbow trout (Oncorhynchus mykiss) populations from British Columbia. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62(7):1523–1539, July 2005. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/f05-
062.

Schwartz:2005:FUS

[1195] John S. Schwartz and Edwin E. Herricks. Fish use of stage-specific fluvial habitats as refuge patches during a flood in a low-gradient Illinois stream. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62(7):1540–1552, July 2005. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f05-060.

Holland:2005:BAM

[1196] Daniel S. Holland, Nokome Bentley, and Philippe Lallemand. A bioeconomic analysis of management strategies for rebuilding and maintenance of the NSS rock lobster (Jasus edwardsii) stock in southern New Zealand. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62(7):1553–1569, July 2005. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f05-059.

Bromaghin:2005:MAG

[1197] Jeffrey F. Bromaghin and Penelope A. Crane. A method to bin alleles of genetic loci that maintains population heterogeneity. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62(7):1570–1579, July 2005. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f05-093.

Smith:2005:WLS

[1198] Katherine L. Smith and Michael L. Jones. Watershed-level sampling effort requirements for determining riverine fish species composition. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62(7):1580–1588, July 2005. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f05-098.

Aguilar:2005:PGN

[1199] Andres Aguilar, James D. Banks, Kenneth F. Levine, and Robert K. Wayne. Population genetics of northern pike (Esox lucius) introduced into Lake Davis, California. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62 (7):1589–1599, July 2005. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f05-068.

Cucherousset:2005:CLH

[1200] Julien Cucherousset, Dominique Ombredane, Katia Charles, Frédéric Marchand, and Jean-Luc Baglinière. A continuum of life history tactics in a brown trout (Salmo trutta) population. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62(7):1600–1610, July 2005. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f05-057.

Taylor:2005:PWF

[1201] David L. Taylor and Donald J. Danila. Predation on winter flounder (Pseudopleuronectes americanus) eggs by the sand shrimp (Crangon septemspinosa). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62 (7):1611–1625, July 2005. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f05-047.

OFarrell:2005:ECL

[1202] Michael R. O’Farrell and Louis W. Botsford. Estimation of change in lifetime egg production from length frequency data. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62(7):1626–1639, July 2005. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f05-064.

Maxwell:2005:EPR

[1203] Michael R. Maxwell, Larry D. Jacobson, and Ramon J. Conser. Eggsper-recruit model for management of the California market squid (Loligo opalescens) fishery. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62(7):1640– 1650, July 2005. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f05-037.

Duponchelle:2005:FPW

[1204] F. Duponchelle, A. J. Ribbink, A. Msukwa, J. Mafuka, D. Mandere, and H. Bootsma. Food partitioning within the species-rich benthic fish community of Lake Malawi, East Africa. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62(7):1651–1664, July 2005. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f05-075.

Hyun:2005:SMS

[1205] Saang-Yoon Hyun, Ray Hilborn, James J. Anderson, and Billy Ernst. A statistical model for in-season forecasts of sockeye salmon (Oncorhynchus nerka) returns to the Bristol Bay districts of Alaska. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62(7):1665–1680, July 2005. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f05-071.

Saros:2005:RRA

[1206] Jasmine E. Saros, Timothy J. Michel, Sebastian J. Interlandi, and Alexander P. Wolfe. Resource requirements of Asterionella formosa and Fragilaria crotonensis in oligotrophic alpine lakes: implications for recent phytoplankton community reorganizations. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62(7):1681–1689, July 2005. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f05-077.

Anonymous:2005:ACc

[1207] Anonymous. Announcement / Communiqué. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62(8), August 2005. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f05-912.

Rosenfeld:2005:FAF

[1208] Jordan S. Rosenfeld, Thomas Leiter, Gerhard Lindner, and Lorne Rothman. Food abundance and fish density alters habitat selection, growth, and habitat suitability curves for juvenile Coho salmon (Oncorhynchus kisutch). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62(8):1691–1701, August 2005. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/f05-
072.

Huser:2005:PIA

[1209] Brian J. Huser and Emil Rydin. Phosphorus inactivation by aluminum in Lakes Gårdsjön and Härsvatten sediment during the industrial acidification period in Sweden. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62 (8):1702–1709, August 2005. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f05-083.

Kaplan:2005:RAP

[1210] I. C. Kaplan. A risk assessment for Pacific leatherback turtles (Dermochelys coriacea). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62(8): 1710–1719, August 2005. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f05-121.

Dunlop:2005:DDI

[1211] Erin S. Dunlop, Judi A. Orendorff, Brian J. Shuter, F. Helen Rodd, and Mark S. Ridgway. Diet and divergence of introduced smallmouth bass (Micropterus dolomieu) populations. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62(8):1720–1732, August 2005. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f05-089.

Jager:2005:GDI

[1212] Henriette I. Jager. Genetic and demographic implications of aquaculture in white sturgeon (Acipenser transmontanus) conservation. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62(8):1733–1745, August 2005. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f05-106.

Gudmundsson:2005:SG

[1213] Gudmundur Gudmundsson. Stochastic growth. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62(8):1746–1755, August 2005. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f05-074.

Clough:2005:IWD

[1214] Lisa M. Clough, Paul E. Renaud, and William G. Ambrose, Jr. Impacts of water depth, sediment pigment concentration, and benthic macrofaunal biomass on sediment oxygen demand in the western Arctic Ocean. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62(8):1756–1765, August 2005.

CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f05-102.

Wong:2005:PRS

[1215] Melisa C. Wong, Myriam A. Barbeau, Allan W. Hennigar, and Shawn M. C. Robinson. Protective refuges for seeded juvenile scallops (Placopecten magellanicus) from sea star (Asterias spp.) and crab (Cancer irroratus and Carcinus maenas) predation. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62(8):1766–1781, August 2005. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f05-092.

Jacobson:2005:EBH

[1216] Larry D. Jacobson, Steven J. Bograd, Richard H. Parrish, Roy Mendelssohn, and Franklin B. Schwing. An ecosystem-based hypothesis for climatic effects on surplus production in California sardine (Sardinops sagax) and environmentally dependent surplus production models. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62(8):1782–1796, August 2005. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f05-095.

Romero:2005:CPC

[1217] Nicolas Romero, Robert E. Gresswell, and Judith L. Li. Changing patterns in coastal cutthroat trout (Oncorhynchus clarki clarki) diet and prey in a gradient of deciduous canopies. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62(8):1797–1807, August 2005. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f05-099.

Munch:2005:BNA

[1218] Stephan B. Munch, Athanasios Kottas, and Marc Mangel. Bayesian nonparametric analysis of stock–recruitment relationships. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62(8):1808–1821, August 2005. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f05-073.

Onsrud:2005:SSS

[1219] M. S. R. Onsrud, S. Kaartvedt, and M. T. Breien. In situ swimming speed and swimming behaviour of fish feeding on the krill Meganyctiphanes norvegica. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62(8): 1822–1832, August 2005. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f05-090.

Jorgensen:2005:IMB

[1220] O. A. Jørgensen, C. Hvingel, P. R. Møller, and M. A. Treble. Identification and mapping of bottom fish assemblages in Davis Strait and southern Baffin Bay. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62(8): 1833–1852, August 2005. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f05-101.

Floeter:2005:GGE

[1221] Jens Floeter, Alexander Kempf, Morten Vinther, Corinna Schrum, and Axel Temming. Grey gurnard (Eutrigla gurnadus) in the North Sea: an emerging key predator? Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62 (8):1853–1864, August 2005. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f05-108.

Jurado-Molina:2005:IPI

[1222] Jesús Jurado-Molina, Patricia A. Livingston, and James N. Ianelli. Incorporating predation interactions in a statistical catch-at-age model for a predator-prey system in the eastern Bering Sea. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62(8):1865–1873, August 2005. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f05-110.

Gomez-Uchida:2005:MAS

[1223] Daniel Gomez-Uchida and Michael A. Banks. Microsatellite analyses of spatial genetic structure in darkblotched rockfish (Sebastes crameri): Is pooling samples safe? Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62 (8):1874–1886, August 2005. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f05-084.

Saisa:2005:PGS

[1224] Marjatta Säisä, Marja-Liisa Koljonen, Riho Gross, Jan Nilsson, Jaana Tähtinen, Jarmo Koskiniemi, and Anti Vasemägi. Population genetic structure and postglacial colonization of Atlantic salmon (Salmo salar) in the Baltic Sea area based on microsatellite DNA variation. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62(8):1887–1904, August 2005. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f05-094.

Huryn:2005:LHB

[1225] Alexander D. Huryn, Karie A. Slavik, Rex L. Lowe, Stephanie M. Parker, Dennis S. Anderson, and Bruce J. Peterson. Landscape heterogeneity and the biodiversity of Arctic stream communities: a habitat template analysis. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62(8):1905–1919, August 2005. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/f05-
100.

Oosterhout:2005:PBC

[1226] Gretchen R. Oosterhout, Charles W. Huntington, Thomas E. Nickelson, and Peter W. Lawson. Potential benefits of a conservation hatchery program for supplementing Oregon coast Coho salmon (Oncorhynchus kisutch) populations: a stochastic model investigation. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62(8):1920–1935, August 2005. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f05-080.

Anonymous:2005:ACd

[1227] Anonymous. Announcement / Communiqué. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62(9), September 2005. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f05-913.

deValpine:2005:SSL

[1228] Perry de Valpine and Ray Hilborn. State-space likelihoods for nonlinear fisheries time-series. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62(9):1937– 1952, September 2005. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f05-116.

Jones:2005:IPF

[1229] Lisa A. Jones and Anthony Ricciardi. Influence of physicochemical factors on the distribution and biomass of invasive mussels (Dreissena polymorpha and Dreissena bugensis) in the St. Lawrence River. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62(9):1953–1962, September 2005. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f05-096.

Allen:2005:MMC

[1230] Erik W. Allen, Ellie E. Prepas, Stephan Gabos, William M. J. Strachan, and Weiping Zhang. Methyl mercury concentrations in macroinvertebrates and fish from burned and undisturbed lakes on the boreal plain. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62(9):1963–1977, September 2005. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f05-103.

Meager:2005:ETR

[1231] Justin J. Meager, Turid Solbakken, Anne C. Utne-Palm, and Tina Oen. Effects of turbidity on the reactive distance, search time, and foraging success of juvenile Atlantic cod (Gadus morhua). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62(9):1978–1984, September 2005. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f05-104.

Walters:2005:EAH

[1232] Carl J. Walters and Ray Hilborn. Exploratory assessment of historical recruitment patterns using relative abundance and catch data. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62(9):1985–1990, September 2005. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f05-105.

Chouinard:2005:CBG

[1233] G. A. Chouinard, D. P. Swain, M. O. Hammill, and G. A. Poirier. Covariation between grey seal (Halichoerus grypus) abundance and natural mortality of cod (Gadus morhua) in the southern Gulf of St. Lawrence.

Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62(9):1991–2000, September 2005. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f05-107.

Blanchard:2005:DAR

[1234] Julia L. Blanchard, Craig Mills, Simon Jennings, Clive J. Fox, Brian D. Rackham, Paul D. Eastwood, and Carl M. O’Brien. Distribution– abundance relationships for North Sea Atlantic cod (Gadus morhua): observation versus theory. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62(9): 2001–2009, September 2005. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f05-109.

Graeb:2005:FBM

[1235] Brian D. S. Graeb, Tracy Galarowicz, David H. Wahl, John M. Dettmers, and Mathew J. Simpson. Foraging behavior, morphology, and life history variation determine the ontogeny of piscivory in two closely related predators. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62(9):2010– 2020, September 2005. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f05-112.

Berglund:2005:POP

[1236] Olof Berglund, Per Nyström, and Per Larsson. Persistent organic pollutants in river food webs: influence of trophic position and degree of heterotrophy. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62(9):2021– 2032, September 2005. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f05-115.

Greenwood:2005:PRL

[1237] Jennifer L. Greenwood and Amy D. Rosemond. Periphyton response to long-term nutrient enrichment in a shaded headwater stream. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62(9):2033–2045, September 2005. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f05-117.

Robison:2005:PMS

[1238] Barrie D. Robison and William Rowland. A potential model system for studying the genetics of domestication: behavioral variation among wild and domesticated strains of zebra danio (Danio rerio). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62(9):2046–2054, September 2005. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f05-118.

St-Jean:2005:DGP

[1239] S. D. St-Jean, R. E. Stephens, S. C. Courtenay, and C. L. Reinisch. Detecting p53 family proteins in haemocytic leukemia cells of Mytilus edulis from Pictou Harbour, Nova Scotia, Canada. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62(9):2055–2066, September 2005. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f05-119.

Duston:2005:GIP

[1240] J. Duston, T. Astatkie, and P. F. MacIsaac. Genetic influence of parr versus anadromous sires on the life histories of Atlantic salmon (Salmo salar). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62(9):2067–2075, September 2005. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f05-120.

Morello:2005:IHD

[1241] E. B. Morello, C. Froglia, R. J. A. Atkinson, and P. G. Moore. Impacts of hydraulic dredging on a macrobenthic community of the Adriatic Sea, Italy. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62(9):2076– 2087, September 2005. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f05-122.

Rennie:2005:IHC

[1242] Michael D. Rennie and Leland J. Jackson. The influence of habitat complexity on littoral invertebrate distributions: patterns differ in shallow prairie lakes with and without fish. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62(9):2088–2099, September 2005. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f05-123.

Giani:2005:ESC

[1243] A. Giani, D. F. Bird, Y. T. Prairie, and J. F. Lawrence. Empirical study of cyanobacterial toxicity along a trophic gradient of lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62(9):2100–2109, September 2005. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f05-124.

Newbrey:2005:BCM

[1244] Michael G. Newbrey, Michael A. Bozek, Martin J. Jennings, and James E. Cook. Branching complexity and morphological characteristics of coarse woody structure as lacustrine fish habitat. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62(9):2110–2123, September 2005. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f05-125.

Wagner:2005:MRS

[1245] G. N. Wagner, S. G. Hinch, L. J. Kuchel, A. Lotto, S. R. M. Jones, D. A. Patterson, J. S. Macdonald, G. Van Der Kraak, M. Shrimpton, K. K. English, S. Larsson, S. J. Cooke, M. C. Healey, and A. P. Farrell. Metabolic rates and swimming performance of adult Fraser River sockeye salmon (Oncorhynchus nerka) after a controlled infection with Parvicapsula minibicornis. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62(9):2124– 2133, September 2005. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f05-126.

Cahill:2005:MIC

[1246] Kendra L. Cahill, John M. Gunn, and Martyn N. Futter. Modelling ice cover, timing of spring stratification, and end-of-season mixing depth in small Precambrian Shield lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62(9):2134–2142, September 2005. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f05-127.

Koljonen:2005:CIA

[1247] Marja-Liisa Koljonen, Jerome J. Pella, and Michele Masuda. Classical individual assignments versus mixture modeling to estimate stock proportions in Atlantic salmon (Salmo salar) catches from DNA microsatellite data. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62(9):2143– 2158, September 2005. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f05-128.

Sass:2005:CGU

[1248] Greg G. Sass and James F. Kitchell. Can growth be used as a surrogate measure of walleye (Sander vitreus) abundance change? Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62(9):2159–2168, September 2005. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f05-129.

Alvarez:2005:CGS

[1249] D. Álvarez and N. B. Metcalfe. Catch-up growth and swimming performance in threespine sticklebacks (Gasterosteus aculeatus): seasonal changes in the cost of compensation. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62(9):2169–2176, September 2005. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f05-130.

Krueger:2005:FWA

[1250] Damon M. Krueger and Thomas R. Hrabik. Food web alterations that promote native species: the recovery of cisco (Coregonus artedi) populations through management of native piscivores. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62(10):2177–2188, October 2005. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f05-132.

Miller:2005:CPS

[1251] J. A. Miller, M. A. Banks, D. Gomez-Uchida, and A. L. Shanks. A comparison of population structure in black rockfish (Sebastes melanops) as determined with otolith microchemistry and microsatellite DNA. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62(10):2189–2198, October 2005. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f05-133.

Field:2005:SVR

[1252] John C. Field and Stephen Ralston. Spatial variability in rockfish (Sebastes spp.) recruitment events in the California Current system. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62(10):2199–2210, October 2005. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f05-134.

Tardif:2005:RDR

[1253] D. Tardif, H. Glémet, P. Brodeur, and M. Mingelbier. RNA/DNA ratio and total length of yellow perch (Perca flavescens) in managed and natural wetlands of a large fluvial lake. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62(10):2211–2218, October 2005. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f05-137.

Nielsen:2005:RJW

[1254] Einar E. Nielsen, Peter Grønkjær, Dorte Meldrup, and Helge Paulsen. Retention of juveniles within a hybrid zone between North Sea and Baltic Sea Atlantic cod (Gadus morhua). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62(10):2219–2225, October 2005. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f05-139.

Hogsden:2005:EPB

[1255] Kristy L. Hogsden and Rolf D. Vinebrooke. Environmental predictors of benthic consumers and autotrophic communities along a recovery gradient. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62(10):2226–2239, October 2005. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f05-138.

Ahlgren:2005:IFA

[1256] Gunnel Ahlgren, Lies Van Nieuwerburgh, Ingrid Wänstrand, Marianne Pedersén, Merike Boberg, and Pauli Snoeijs. Imbalance of fatty acids in the base of the Baltic Sea food web — a mesocosm study. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62(10):2240–2253, October 2005. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f05-140.

Jonas:2005:EED

[1257] Jory L. Jonas, Randall M. Claramunt, John D. Fitzsimons, J. Ellen Marsden, and Brian J. Ellrott. Estimates of egg deposition and effects of lake trout (Salvelinus namaycush) egg predators in three regions of the Great Lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62(10):2254–2264, October 2005. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/f05-
141.

Stransky:2005:SSZ

[1258] Christoph Stransky and Shayne E. MacLellan. Species separation and zoogeography of redfish and rockfish (genus Sebastes) by otolith shape analysis. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62(10):2265–2276, October 2005. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/f05-
143.

Black:2005:UTR

[1259] Bryan A. Black, George W. Boehlert, and Mary M. Yoklavich. Using tree-ring crossdating techniques to validate annual growth increments in long-lived fishes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62(10): 2277–2284, October 2005. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f05-142.

Winder:2005:DBF

[1260] Monika Winder, Daniel E. Schindler, Jonathan W. Moore, Susan P. Johnson, and Wendy J. Palen. Do bears facilitate transfer of salmon resources to aquatic macroinvertebrates? Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62(10):2285–2293, October 2005. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://

cdnsciencepub.com/doi/full/10.1139/f05-136.

Reyjol:2005:SSV

[1261] Yorick Reyjol, Philipp Fischer, Sovan Lek, Roland Rösch, and Reiner Eckmann. Studying the spatiotemporal variation of the littoral fish community in a large prealpine lake, using self-organizing mapping. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62(10):2294–2302, October 2005. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f05-097.

Winter:2005:IAB

[1262] Anna R. Winter, Joel W. Nichols, and Richard C. Playle. Influence of acidic to basic water pH and natural organic matter on aluminum accumulation by gills of rainbow trout (Oncorhynchus mykiss). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62(10):2303–2311, October 2005. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f05-063.

Roy:2005:RSE

[1263] Allison H. Roy, Christina L. Faust, Mary C. Freeman, and Judith L. Meyer. Reach-scale effects of riparian forest cover on urban stream ecosystems. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62(10):2312–2329, October 2005. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f05-135.

Galarowicz:2005:FYY

[1264] Tracy L. Galarowicz and David H. Wahl. Foraging by a young-of-the-year piscivore: the role of predator size, prey type, and density. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62(10):2330–2342, October 2005. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f05-148.

Jorgensen:2005:GPH

[1265] Jeffrey C. Jorgensen and James F. Kitchell. Growth potential and host mortality of the parasitic phase of the sea lamprey (Petromyzon marinus) in Lake Superior. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62(10): 2343–2353, October 2005. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f05-144.

Jorgensen:2005:SLP

[1266] Jeffrey C. Jorgensen and James F. Kitchell. Sea lamprey (Petromyzon marinus) size trends in a salmonid stocking context in Lake Superior. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62(10):2354–2361, October 2005. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f05-145.

Dufour:2005:ELH

[1267] Elise Dufour, William P. Patterson, Tomas O. Höök, and Edward S. Rutherford. Early life history of Lake Michigan alewives (Alosa pseudoharengus) inferred from intra-otolith stable isotope ratios. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62(10):2362–2370, October 2005. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f05-147.

Bundy:2005:IMV

[1268] Marie H. Bundy, Henry A. Vanderploeg, Peter J. Lavrentyev, and Paul A. Kovalcik. The importance of microzooplankton versus phytoplankton to copepod populations during late winter and early spring in Lake Michigan. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62(10):2371–2385, October 2005. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/f05-
111.

Wonham:2005:MIR

[1269] Marjorie J. Wonham, Sarah A. Bailey, Hugh J. MacIsaac, and Mark A. Lewis. Modelling the invasion risk of diapausing organisms transported in ballast sediments. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62(10): 2386–2398, October 2005. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f05-146.

Young:2005:PCT

[1270] Michael K. Young, Paula M. Guenther-Gloss, and Ashley D. Ficke. Predicting cutthroat trout (Oncorhynchus clarkii) abundance in highelevation streams: revisiting a model of translocation success. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62(10):2399–2408, October 2005. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f05-149.

Handegard:2005:WFM

[1271] N. O. Handegard and D. Tjøstheim. When fish meet a trawling vessel: examining the behaviour of gadoids using a free-floating buoy and acoustic split-beam tracking. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62(10): 2409–2422, October 2005. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f05-131.

Hyun:2005:ESM

[1272] Saang-Yoon Hyun, Ray Hilborn, James J. Anderson, and Billy Ernst. Erratum: a statistical model for in-season forecasts of sockeye salmon (Oncorhynchus nerka) returns to the Bristol Bay districts of Alaska. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62(10):2423, October 2005. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f05-186.

Barnett-Johnson:2005:VSI

[1273] Rachel Barnett-Johnson, Frank C. Ramos, Churchill B. Grimes, and R. Bruce MacFarlane. Validation of Sr isotopes in otoliths by laser ablation multicollector inductively coupled plasma mass spectrometry (LA-MC-ICPMS): opening avenues in fisheries science applications. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62(11):2425–2430, November 2005. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f05-194.

Ripley:2005:BTS

[1274] Travis Ripley, Garry Scrimgeour, and Mark S. Boyce. Bull trout (Salvelinus confluentus) occurrence and abundance influenced by cumulative industrial developments in a Canadian boreal forest watershed. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62(11):2431–2442, November 2005. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f05-150.

Courtemanche:2005:NAU

[1275] David A. Courtemanche, Frederick G. Whoriskey, Jr., Valerie Bujold, and R. Allen Curry. A nonlethal approach using strontium in scales to distinguish periods of marine and freshwater residency of anadromous species. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62(11):2443–2449, November 2005. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f05-162.

Strecker:2005:IBI

[1276] Angela L. Strecker and Shelley E. Arnott. Impact of Bythotrephes invasion on zooplankton communities in acid-damaged and recovered lakes on the Boreal Shield. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62(11): 2450–2462, November 2005. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f05-152.

Duggan:2005:IAR

[1277] Ian C. Duggan, Colin D. A. van Overdijk, Sarah A. Bailey, Philip T. Jenkins, Helene Limén, and Hugh J. MacIsaac. Invertebrates associated with residual ballast water and sediments of cargo-carrying ships entering the Great Lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62(11): 2463–2474, November 2005. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f05-160.

Baba:2005:EAC

[1278] Katsuhisa Baba, Masayoshi Sasaki, and Noriyuki Mitsutani. Estimation of age composition from length data by posterior probabilities based on a previous growth curve: application to Sebastes schlegelii. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62(11):2475–2483, November 2005. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f05-158.

Lobon-Cervia:2005:IRP

[1279] Javier Lobón-Cerviá. The importance of recruitment for the production dynamics of stream-dwelling brown trout (Salmo trutta). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62(11):2484–2493, November 2005. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f05-166.

Scheuerell:2005:TDF

[1280] Jennifer M. Scheuerell, Daniel E. Schindler, Mark D. Scheuerell, Kurt L. Fresh, Thomas H. Sibley, Arni H. Litt, and Judith H. Shepherd. Temporal dynamics in foraging behavior of a pelagic predator. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62(11):2494–2501, November 2005. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f05-164.

Jonsson:2005:DCD

[1281] Nina Jonsson, Bror Jonsson, and Lars Petter Hansen. Does climate during embryonic development influence parr growth and age of seaward migration in Atlantic salmon (Salmo salar)? Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62(11):2502–2508, November 2005. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f05-154.

Alkanani:2005:LCN

[1282] T. Alkanani, C. C. Parrish, K. J. Rodnick, and A. K. Gamperl. Lipid class and nonesterified fatty acid profiles in plasma of North Atlantic cod (Gadus morhua). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62(11): 2509–2518, November 2005. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f05-151.

Canino:2005:GDW

[1283] Michael F. Canino, Patrick T. O’Reilly, Lorenz Hauser, and Paul Bentzen. Genetic differentiation in walleye pollock (Theragra chalcogramma) in response to selection at the pantophysin (Pan i) locus. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62(11):2519–2529, November 2005. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f05-155.

Hill:2005:RMI

[1284] Simeon L. Hill, Keith Reid, and Anthony W. North. Recruitment of mackerel icefish (Champsocephalus gunnari) at South Georgia indicated by predator diets and its relationship with sea surface temperature. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62(11):2530–2537, November 2005. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f05-157.

McClelland:2005:TGO

[1285] Erin K. McClelland, James M. Myers, Jeffrey J. Hard, Linda K. Park, and Kerry A. Naish. Two generations of outbreeding in Coho salmon (Oncorhynchus kisutch): effects on size and growth. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62(11):2538–2547, November 2005. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f05-159.

Stamford:2005:PSG

[1286] M. D. Stamford and E. B. Taylor. Population subdivision and genetic signatures of demographic changes in Arctic grayling (Thymallus arcticus) from an impounded watershed. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62(11):2548–2559, November 2005. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f05-156.

Bunnell:2005:IBM

[1287] David B. Bunnell and Thomas J. Miller. An individual-based modeling approach to spawning-potential per-recruit models: an application to blue crab (Callinectes sapidus) in Chesapeake Bay. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62(11):2560–2572, November 2005. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f05-153.

Whitehead:2005:HMR

[1288] Hal Whitehead and Tonya Wimmer. Heterogeneity and the mark– recapture assessment of the Scotian Shelf population of northern bottlenose whales (Hyperoodon ampullatus). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62(11):2573–2585, November 2005. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f05-178.

Reavie:2005:PFL

[1289] Euan D. Reavie, John A. Robbins, Eugene F. Stoermer, Marianne S. V. Douglas, Gail E. Emmert, Nancy R. Morehead, and Alena Mudroch. Paleolimnology of a fluvial lake downstream of Lake Superior and the industrialized region of sault Saint marie. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62(11):2586–2608, November 2005. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f05-170.

Melancon:2005:ECS

[1290] Sonia Melancon, Brian J. Fryer, Stuart A. Ludsin, Joel E. Gagnon, and Zhaoping Yang. Effects of crystal structure on the uptake of metals by lake trout (Salvelinus namaycush) otoliths. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62(11):2609–2619, November 2005. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f05-161.

Hindar:2005:LTR

[1291] Atle Hindar and Richard F. Wright. Long-term records and modelling of acidification, recovery, and liming at Lake Hovvatn, Norway. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62(11):2620–2631, November 2005. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f05-165.

Bondar:2005:DTP

[1292] Carin A. Bondar, K. Bottriell, K. Zeron, and John S. Richardson. Does trophic position of the omnivorous signal crayfish (Pacifastacus leniusculus) in a stream food web vary with life history stage or density? Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62(11):2632–2639, November 2005. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f05-167.

Pulido-Villena:2005:FWR

[1293] Elvira Pulido-Villena, Isabel Reche, and Rafael Morales-Baquero. Food web reliance on allochthonous carbon in two high mountain lakes with contrasting catchments: a stable isotope approach. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62(11):2640–2648, November 2005. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f05-169.

Steinhart:2005:ESA

[1294] Geoffrey B. Steinhart, Nancy J. Leonard, Roy A. Stein, and Elizabeth A. Marschall. Effects of storms, angling, and nest predation during angling on smallmouth bass (Micropterus dolomieu) nest success. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62(11):2649–2660, November 2005. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f05-171.

Mackie:2005:OGE

[1295] P. Mackie, P. A. Wright, B. D. Glebe, and J. S. Ballantyne. Osmoregulation and gene expression of Na^{+} /K^{+} ATPase in families of Atlantic salmon (Salmo salar) smolts. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62 (11):2661–2672, November 2005. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/

doi/full/10.1139/f05-168.

Ernst:2005:ESD

[1296] Billy Ernst, J. M. (Lobo) Orensanz, and David A. Armstrong. Erratum: Spatial dynamics of female snow crab (Chionoecetes opilio) in the eastern Bering Sea. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62 (11):2673, November 2005. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f05-197.

Genna:2005:LIS

[1297] R. L. Genna, W. Mordue, A. W. Pike, and A. J. Mordue (Luntz). Light intensity, salinity, and host velocity influence presettlement intensity and distribution on hosts by copepodids of sea lice, Lepeophtheirus salmonis. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62(12):2675–2682, December 2005. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f05-163.

Dettmers:2005:EAM

[1298] John M. Dettmers, John Janssen, Bernard Pientka, Richard S. Fulford, and David J. Jude. Evidence across multiple scales for offshore transport of yellow perch (Perca flavescens) larvae in Lake Michigan. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62(12):2683–2693, December 2005. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f05-173.

Blanchfield:2005:RIB

[1299] Paul J. Blanchfield and Mark S. Ridgway. The relative influence of breeding competition and habitat quality on female reproductive success in lacustrine brook trout (Salvelinus fontinalis). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62(12):2694–2705, December 2005. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f05-176.

Bertolo:2005:RBP

[1300] Andrea Bertolo and Pierre Magnan. The relationship between piscivory and growth of white sucker (Catostomus commersoni) and yellow perch (Perca flavescens) in headwater lakes of the Canadian Shield. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62(12):2706–2715, December 2005. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f05-174.

Bradford:2005:UCI

[1301] Michael J. Bradford, Josh Korman, and Paul S. Higgins. Using confidence intervals to estimate the response of salmon populations (Oncorhynchus spp.) to experimental habitat alterations. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62(12):2716–2726, December 2005. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f05-179.

Sundbom:2005:EDV

[1302] Marcus Sundbom and Markus Meili. Ecophysiological determinants of the variation in ^{137}Cs concentrations between and within lacustrine fish populations. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62(12):2727– 2739, December 2005. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f05-172.

Opperman:2005:ILU

[1303] Jeff J. Opperman, Kathleen A. Lohse, Colin Brooks, N. Maggi Kelly, and Adina M. Merenlender. Influence of land use on fine sediment in salmonid spawning gravels within the Russian River Basin, California. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62(12):2740–2751, December 2005. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f05-187.

Morita:2005:RAS

[1304] Kentaro Morita, Shoko H. Morita, Masa aki Fukuwaka, and Hiroyuki Matsuda. Rule of age and size at maturity of chum salmon (Oncorhynchus keta): implications of recent trends among Oncorhynchus spp. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62(12):2752–2759, December 2005. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f05-182.

Hoyer:2005:LLT

[1305] Mark V. Hoyer, Christine A. Horsburgh, Daniel E. Canfield, Jr., and Roger W. Bachmann. Lake level and trophic state variables among a population of shallow Florida lakes and within individual lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62(12):2760–2769, December 2005. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f05-177.

Rennie:2005:PMB

[1306] M. D. Rennie, N. C. Collins, C. F. Purchase, and A. Tremblay. Predictive models of benthic invertebrate methylmercury in Ontario and Quebec lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62(12):2770– 2783, December 2005. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f05-181.

McHugh:2005:EEC

[1307] Peter McHugh and Phaedra Budy. An experimental evaluation of competitive and thermal effects on brown trout (Salmo trutta) and Bonneville cutthroat trout (Oncorhynchus clarkii utah) performance along an altitudinal gradient. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62(12): 2784–2795, December 2005. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f05-184.

Harrod:2005:PSC

[1308] Chris Harrod and David Griffiths. Parasitism, space constraints, and gonad asymmetry in the pollan (Coregonus autumnalis). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62(12):2796–2801, December 2005. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f05-175.

Layman:2005:BSO

[1309] C. A. Layman, R. B. Langerhans, and K. O. Winemiller. Body size, not other morphological traits, characterizes cascading effects in fish assemblage composition following commercial netting. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62(12):2802–2810, December 2005. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f05-183.

Morley:2005:JSO

[1310] Sarah A. Morley, Patricia S. Garcia, Todd R. Bennett, and Philip Roni. Juvenile salmonid (Oncorhynchus spp.) use of constructed and natural side channels in Pacific northwest rivers. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62(12):2811–2821, December 2005. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f05-185.

Horne:2005:STV

[1311] John K. Horne and Paul D. Walline. Spatial and temporal variance of walleye pollock (Theragra chalcogramma) in the eastern Bering Sea. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62(12):2822–2831, December 2005. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f05-192.

Ashford:2005:COE

[1312] J. R. Ashford, C. M. Jones, E. Hofmann, I. Everson, C. Moreno, G. Duhamel, and R. Williams. Can otolith elemental signatures record the capture site of Patagonian toothfish (Dissostichus eleginoides), a fully marine fish in the Southern Ocean? Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62(12):2832–2840, December 2005. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f05-191.

Bertram:2005:OCN

[1313] Douglas F. Bertram, Anne Harfenist, and Barry D. Smith. Ocean climate and El Niño impacts on survival of Cassin’s auklets from upwelling and downwelling domains of British Columbia. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62(12):2841–2853, December 2005. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f05-190.

Levesque:2005:ESTa

[1314] H. M. Levesque, C. Short, T. W. Moon, J. S. Ballantyne, and W. R. Driedzic. Effects of seasonal temperature and photoperiod on Atlantic cod (Gadus morhua). I. Morphometric parameters and metabolites. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62(12):2854–2863, December 2005. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f05-188.

Levesque:2005:ESTb

[1315] H. M. Levesque, J. Bondy, C. Short, J. S. Ballantyne, W. R. Driedzic, and T. W. Moon. Effects of seasonal temperature and photoperiod on Atlantic cod (Gadus morhua). II. Enzymes of intermediary metabolism. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62(12):2864–2873, December 2005. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f05-189.

Rowell:2005:ICR

[1316] Kirsten Rowell, Karl W. Flessa, David L. Dettman, and Martha Román. The importance of Colorado River flow to nursery habitats of the Gulf corvina (Cynoscion othonopterus). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62(12):2874–2885, December 2005. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f05-193.

Bellman:2005:EUW

[1317] Marlene A. Bellman, Scott A. Heppell, and Chris Goldfinger. Evaluation of a US west coast groundfish habitat conservation regulation via analysis of spatial and temporal patterns of trawl fishing effort. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62(12):2886–2900, December 2005. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f05-180.

Anonymous:2005:AII

[1318] Anonymous. Author index / index des auteurs: Volume 62, 2005. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62(12), December 2005. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f05-916.

Anonymous:2005:CSV

[1319] Anonymous. Contents / sommaire: Volume 62, 2005. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 62(12), December 2005. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f05-915.

Belzile:2006:ESM

[1320] Nelson Belzile, Yu-Wei Chen, John M. Gunn, Jian Tong, Yves Alarie, Tania Delonchamp, and Chun-Yan Lang. The effect of selenium on mercury assimilation by freshwater organisms. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 63(1):1–10, January 2006. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f05-202.

Fulford:2006:EPS

[1321] Richard S. Fulford, James A. Rice, Thomas J. Miller, and Fred P. Binkowski. Elucidating patterns of size-dependent predation on larval yellow perch (Perca flavescens) in Lake Michigan: an experimental and modeling approach. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 63 (1):11–27, January 2006. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f05-195.

Fulford:2006:FSL

[1322] Richard S. Fulford, James A. Rice, Thomas J. Miller, Fred P. Binkowski, John M. Dettmers, and Brian Belonger. Foraging selectivity by larval yellow perch (Perca flavescens): implications for understanding recruitment in small and large lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 63 (1):28–42, January 2006. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f05-196.

Taylor:2006:DAD

[1323] Christopher M. Taylor, Thomas L. Holder, Riccardo A. Fiorillo, Lance R. Williams, R. Brent Thomas, and Melvin L. Warren, Jr. Distribution, abundance, and diversity of stream fishes under variable environmental conditions. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 63(1):43–54, January 2006. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/f05-
203.

Morita:2006:PDJ

[1324] Kentaro Morita, Shoko H. Morita, and Masa aki Fukuwaka. Population dynamics of Japanese pink salmon (Oncorhynchus gorbuscha): are recent increases explained by hatchery programs or climatic variations? Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 63(1):55–62, January 2006. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f05-207.

Rieman:2006:BTS

[1325] Bruce E. Rieman, James T. Peterson, and Deborah L. Myers. Have brook trout (Salvelinus fontinalis) displaced bull trout (Salvelinus confluentus) along longitudinal gradients in central Idaho streams? Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 63(1):63–78, January 2006. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f05-206.

Solomon:2006:EDS

[1326] Christopher T. Solomon, Peter K. Weber, Joseph J. Cech, Jr., B. Lynn Ingram, Mark E. Conrad, Madhav V. Machavaram, Asya R. Pogodina, and Rebecca L. Franklin. Experimental determination of the sources of otolith carbon and associated isotopic fractionation. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 63(1):79–89, January 2006. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f05-200.

Kettle:2006:HDE

[1327] A. James Kettle and Keith Haines. How does the European eel (Anguilla anguilla) retain its population structure during its larval migration across the North Atlantic Ocean? Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 63 (1):90–106, January 2006. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f05-198.

Dillingham:2006:FSG

[1328] Peter W. Dillingham, John R. Skalski, and Kristen E. Ryding. Fine-scale geographic interactions between Steller sea lion (Eumetopias jubatus) trends and local fisheries. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 63 (1):107–119, January 2006. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f05-208.

Kasahara:2006:ERS

[1329] Tamao Kasahara and Alan R. Hill. Effects of riffle–step restoration on hyporheic zone chemistry in N-rich lowland streams. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 63(1):120–133, January 2006. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f05-199.

Greathouse:2006:DRC

[1330] Effie A. Greathouse and Catherine M. Pringle. Does the river continuum concept apply on a tropical island? Longitudinal variation in a Puerto Rican stream. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 63(1): 134–152, January 2006. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f05-201.

Smith:2006:EMR

[1331] Martin D. Smith, Junjie Zhang, and Felicia C. Coleman. Effectiveness of marine reserves for large-scale fisheries management. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 63(1):153–164, January 2006. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f05-205.

Jager:2006:CLOa

[1332] Henriette I. Jager. Chutes and ladders and other games we play with rivers. I. Simulated effects of upstream passage on white sturgeon. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 63(1):165–175, January 2006. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f05-226.

Jager:2006:CLOb

[1333] Henriette I. Jager. Chutes and ladders and other games we play with rivers. II. Simulated effects of translocation on white sturgeon. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 63(1):176–185, January 2006. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f05-225.

Jorgensen:2006:SDE

[1334] Christian Jørgensen and Øyvind Fiksen. State-dependent energy allocation in cod (Gadus morhua). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 63 (1):186–199, January 2006. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f05-209.

Jorgensen:2006:LSS

[1335] Christian Jørgensen, Bruno Ernande, Øyvind Fiksen, and Ulf Dieckmann. The logic of skipped spawning in fish. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 63(1):200–211, January 2006. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f05-210.

Walters:2006:SAS

[1336] Carl J. Walters, Steven J. D. Martell, and Josh Korman. A stochastic approach to stock reduction analysis. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 63(1):212–223, January 2006. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f05-213.

Ross:2006:FKW

[1337] Peter S. Ross. Fireproof killer whales (Orcinus orca): flame-retardant chemicals and the conservation imperative in the charismatic icon of British Columbia, Canada. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 63 (1):224–234, January 2006. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f05-244.

Shelton:2006:FUL

[1338] Peter A. Shelton, Alan F. Sinclair, Ghislain A. Chouinard, Robert Mohn, and Daniel E. Duplisea. Fishing under low productivity conditions is further delaying recovery of Northwest Atlantic cod (Gadus morhua). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 63(2):235–238, February 2006. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f05-253.

Borwick:2006:TIA

[1339] Jason Borwick, Jim Buttle, and Mark S. Ridgway. A topographic index approach for identifying groundwater habitat of young-of-year brook trout (Salvelinus fontinalis) in the land–lake ecotone. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 63(2):239–253, February 2006. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f05-212.

Gillis:2006:SEI

[1340] Darren M. Gillis, Elmer Wade, and Douglas P. Swain. Spatial evidence for information exchange and competition in the Gulf of St. Lawrence snow crab (Chionoecetes opilio) fishery. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 63(2):254–267, February 2006. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f05-211.

Davoren:2006:SBM

[1341] Gail K. Davoren, John T. Anderson, and William A. Montevecchi. Shoal behaviour and maturity relations of spawning capelin (Mallotus villosus) off Newfoundland: demersal spawning and diel vertical movement patterns. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 63(2):268–284, February 2006. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f05-204.

Isaak:2006:NSS

[1342] Daniel J. Isaak and Russell F. Thurow. Network-scale spatial and temporal variation in Chinook salmon (Oncorhynchus tshawytscha) redd distributions: patterns inferred from spatially continuous replicate surveys. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 63(2):285–296, February 2006.

CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f05-214.

Zimmerman:2006:PCA

[1343] Mara S. Zimmerman. Predator communities associated with brook stickleback (Culaea inconstans) prey: patterns in body size. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 63(2):297–309, February 2006. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f05-216.

Skjaeraasen:2006:TDP

[1344] Jon Egil Skjæraasen, Trygve Nilsen, and Olav S. Kjesbu. Timing and determination of potential fecundity in Atlantic cod (Gadus morhua). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 63(2):310–320, February 2006. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f05-218.

Michielsens:2006:BSS

[1345] Catherine G. J. Michielsens, Murdoch K. McAllister, Sakari Kuikka, Tapani Pakarinen, Lars Karlsson, Atso Romakkaniemi, Ingemar Perä, and Samu Mäntyniemi. A Bayesian state–space mark–recapture model to estimate exploitation rates in mixed-stock fisheries. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 63(2):321–334, February 2006. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f05-215.

Roth:2006:UBS

[1346] Brian M. Roth, Catherine L. Hein, and M. Jake Vander Zanden. Using bioenergetics and stable isotopes to assess the trophic role of rusty crayfish (Orconectes rusticus) in lake littoral zones. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 63(2):335–344, February 2006. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f05-217.

MacNeil:2006:VUE

[1347] M. Aaron MacNeil, Ken G. Drouillard, and Aaron T. Fisk. Variable uptake and elimination of stable nitrogen isotopes between tissues in fish. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 63(2):345–353, February 2006.

CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f05-219.

Seilheimer:2006:DUW

[1348] T. S. Seilheimer and P. Chow-Fraser. Development and use of the wetland fish index to assess the quality of coastal wetlands in the Laurentian Great Lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 63(2): 354–366, February 2006. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f05-220.

Feio:2006:ITL

[1349] Maria J. Feio, Trefor B. Reynoldson, and Manuel A. S. Graça. The influence of taxonomic level on the performance of a predictive model for water quality assessment. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 63 (2):367–376, February 2006. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f05-221.

Ward:2006:EED

[1350] Ashley J. W. Ward, Alison J. Duff, and Suzanne Currie. The effects of the endocrine disrupter 4-nonylphenol on the behaviour of juvenile rainbow trout (Oncorhynchus mykiss). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 63(2):377–382, February 2006. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f05-223.

Hein:2006:FPT

[1351] Catherine L. Hein, Brian M. Roth, Anthony R. Ives, and M. Jake Vander Zanden. Fish predation and trapping for rusty crayfish (Orconectes rusticus) control: a whole-lake experiment. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 63(2):383–393, February 2006. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f05-229.

Mazor:2006:EBA

[1352] Raphael D. Mazor, Trefor B. Reynoldson, David M. Rosenberg, and Vincent H. Resh. Effects of biotic assemblage, classification, and assessment method on bioassessment performance. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 63(2):394–411, February 2006. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f05-222.

Kjoss:2006:EDL

[1353] Victoria A. Kjoss, Chris M. Wood, and D. Gordon McDonald. Effects of different ligands on the bioaccumulation and subsequent depuration of dietary Cu and Zn in juvenile rainbow trout (Oncorhynchus mykiss). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 63(2):412–422, February 2006. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f05-230.

Sharma:2006:ECR

[1354] Rishi Sharma, Gary Morishima, Shizhen Wang, André Talbot, and Larry Gilbertson. An evaluation of the Clearwater River supplementation program in western Washington. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 63 (2):423–437, February 2006. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f05-228.

Wessel:2006:VAB

[1355] Maria Lang Wessel, William W. Smoker, Robert M. Fagen, and John Joyce. Variation of agonistic behavior among juvenile Chinook salmon (Oncorhynchus tshawytscha) of hatchery, hybrid, and wild origin. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 63(2):438–447, February 2006. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f05-227.

Bailey:2006:RSH

[1356] Roberta J. E. Bailey, Michael A. Birkett, Anna Ingvarsdóttir, A. Jennifer Mordue (Luntz), William Mordue, Bríd O’Shea, John A. Pickett, and Lester J. Wadhams. The role of semiochemicals in host location and non-host avoidance by salmon louse (Lepeophtheirus salmonis) copepodids. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 63(2):448–456, February 2006. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f05-231.

Jones:2006:FEC

[1357] Michael L. Jones, Brian J. Shuter, Yingming Zhao, and Jason D. Stockwell. Forecasting effects of climate change on Great Lakes fisheries: models that link habitat supply to population dynamics can help. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 63(2):457–468, February 2006. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f05-239.

Greenwood:2006:EPR

[1358] Jennifer L. Greenwood and Amy D. Rosemond. Erratum: Periphyton response to long-term nutrient enrichment in a shaded headwater stream. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 63(2):469, February 2006. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f05-255.

Alkanani:2006:ELC

[1359] T. Alkanani, C. C. Parrish, K. J. Rodnick, and A. K. Gamperl. Erratum: Lipid class and nonesterified fatty acid profiles in plasma of North Atlantic cod (Gadus morhua). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 63 (2):470, February 2006. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f05-254.

Houle:2006:BCR

[1360] Daniel Houle, Rock Ouimet, Suzanne Couture, and Christian Gagnon. Base cation reservoirs in soil control the buffering capacity of lakes in forested catchments. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 63 (3):471–474, March 2006. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f06-007.

Wagner:2006:FTV

[1361] C. Michael Wagner, Michael L. Jones, Michael B. Twohey, and Peter W. Sorensen. A field test verifies that pheromones can be useful for sea lamprey (Petromyzon marinus) control in the Great Lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 63(3):475–479, March 2006. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f06-008.

Falke:2006:ERC

[1362] Jeffrey A. Falke and Keith B. Gido. Effects of reservoir connectivity on stream fish assemblages in the Great Plains. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 63(3):480–493, March 2006. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f05-233.

Goodwin:2006:LHC

[1363] Nicholas B. Goodwin, Alastair Grant, Allison L. Perry, Nicholas K. Dulvy, and John D. Reynolds. Life history correlates of densitydependent recruitment in marine fishes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 63(3):494–509, March 2006. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f05-234.

Marchal:2006:IDF

[1364] Paul Marchal, Bo Andersen, David Bromley, Ane Iriondo, Stéphanie Mahévas, Floor Quirijns, Brian Rackham, Marina Santurtún, Nicola Tien, and Clara Ulrich. Improving the definition of fishing effort for important European fleets by accounting for the skipper effect. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 63(3):510–533, March 2006. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f05-238.

Polacheck:2006:ICA

[1365] Tom Polacheck, J. Paige Eveson, Geoff M. Laslett, Kenneth H. Pollock, and William S. Hearn. Integrating catch-at-age and multiyear tagging data: a combined Brownie and Petersen estimation approach in a fishery context. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 63(3):534–548, March 2006. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f05-232.

Kraemer:2006:FSE

[1366] Lisa D. Kraemer, Peter G. C. Campbell, Landis Hare, and Jean-Christian Auclair. A field study examining the relative importance of food and water as sources of cadmium for juvenile yellow perch (Perca flavescens). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 63(3):549–557, March 2006. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f05-236.

MacRae:2006:CNT

[1367] Pamela S. D. MacRae and Donald A. Jackson. Characterizing north temperate lake littoral fish assemblages: a comparison between distance sampling and minnow traps. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 63 (3):558–568, March 2006. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f05-240.

Lepak:2006:RFW

[1368] Jesse M. Lepak, Clifford E. Kraft, and Brian C. Weidel. Rapid food web recovery in response to removal of an introduced apex predator. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 63(3):569–575, March 2006. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f05-248.

Pella:2006:GSM

[1369] Jerome Pella and Michele Masuda. The Gibbs and split–merge sampler for population mixture analysis from genetic data with incomplete baselines. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 63(3):576–596, March 2006. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f05-224.

Davey:2006:MDD

[1370] Andrew J. H. Davey, George F. Turner, Stephen J. Hawkins, and C. Patrick Doncaster. Mechanisms of density dependence in stream fish: exploitation competition for food reduces growth of adult European bullheads (Cottus gobio). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 63 (3):597–606, March 2006. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f05-246.

Brophy:2006:SSF

[1371] Deirdre Brophy, Bret S. Danilowicz, and Pauline A. King. Spawning season fidelity in sympatric populations of Atlantic herring (Clupea harengus). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 63(3):607–616, March 2006. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f05-235.

Irvine:2006:SBD

[1372] Sarah B. Irvine, John D. Stevens, and Laurie J. B. Laurenson. Surface bands on deepwater squalid dorsal-fin spines: an alternative method for ageing Centroselachus crepidater. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 63(3):617–627, March 2006. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f05-237.

Paterson:2006:ESN

[1373] Gordon Paterson, Kenneth G. Drouillard, and G. Douglas Haffner. An evaluation of stable nitrogen isotopes and polychlorinated biphenyls as bioenergetic tracers in aquatic systems. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 63(3):628–641, March 2006. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f05-241.

Hilborn:2006:IMP

[1374] Ray Hilborn, Fiorenza Micheli, and Giulio A. De Leo. Integrating marine protected areas with catch regulation. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 63(3):642–649, March 2006. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f05-243.

Labonne:2006:MPV

[1375] Jacques Labonne and Philippe Gaudin. Modelling population viability in fragmented environments: contribution to the conservation of an endangered percid (Zingel asper). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 63 (3):650–659, March 2006. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f05-245.

Schrank:2006:FIS

[1376] Amy J. Schrank and Frank J. Rahel. Factors influencing summer movement patterns of Bonneville cutthroat trout (Oncorhynchus clarkii utah). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 63(3):660–669, March 2006. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f05-250.

Campana:2006:ERC

[1377] Steven E. Campana, Linda Marks, Warren Joyce, and Nancy E. Kohler. Effects of recreational and commercial fishing on blue sharks (Prionace glauca) in Atlantic Canada, with inferences on the North Atlantic population. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 63(3):670–682, March 2006. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f05-251.

Rosenfeld:2006:INA

[1378] Jordan S. Rosenfeld and Todd Hatfield. Information needs for assessing critical habitat of freshwater fish. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 63(3):683–698, March 2006. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f05-242.

Grafton:2006:IBA

[1379] R. Quentin Grafton, Ragnar Arnason, Trond Bjørndal, David Campbell, Harry F. Campbell, Colin W. Clark, Robin Connor, Diane P. Dupont, Rögnvaldur Hannesson, Ray Hilborn, James E. Kirkley, Tom Kompas, Daniel E. Lane, Gordon R. Munro, Sean Pascoe, Dale Squires, Stein Ivar Steinshamn, Bruce R. Turris, and Quinn Weninger. Incentive-based approaches to sustainable fisheries. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 63(3):699–710, March 2006. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f05-247.

Squires:2006:CBI

[1380] Margaret M. Squires, David Mazzucchi, and Kevin J. Devito. Carbon burial and infill rates in small Western Boreal lakes: physical factors affecting carbon storage. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 63 (4):711–720, April 2006. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f05-252.

Hiddink:2006:CIS

[1381] J. G. Hiddink, S. Jennings, M. J. Kaiser, A. M. Queirós, D. E. Duplisea, and G. J. Piet. Cumulative impacts of seabed trawl disturbance on benthic biomass, production, and species richness in different habitats. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 63(4):721–736, April 2006. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f05-266.

BrindAmour:2006:ESA

[1382] Anik Brind’Amour and Daniel Boisclair. Effect of the spatial arrangement of habitat patches on the development of fish habitat models in the littoral zone of a Canadian Shield lake. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 63(4):737–753, April 2006. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f05-249.

Hunsicker:2006:SSP

[1383] Mary E. Hunsicker and Timothy E. Essington. Size-structured patterns of piscivory of the longfin inshore squid (Loligo pealeii) in the mid-Atlantic continental shelf ecosystem. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 63(4):754–765, April 2006. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f05-258.

McLaughlin:2006:ELH

[1384] Robert L. McLaughlin, Louise Porto, David L. G. Noakes, Jeffrey R. Baylis, Leon M. Carl, Hope R. Dodd, Jon D. Goldstein, Daniel B. Hayes, and Robert G. Randall. Effects of low-head barriers on stream fishes: taxonomic affiliations and morphological correlates of sensitive species. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 63(4):766–779, April 2006. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f05-256.

Hansen:2006:USG

[1385] Michael M. Hansen, Einar E. Nielsen, and Karen-Lise D. Mensberg. Underwater but not out of sight: genetic monitoring of effective population size in the endangered North Sea houting (Coregonus oxyrhynchus). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 63(4):780–787, April 2006. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f05-260.

Mackenzie-Grieve:2006:PIC

[1386] Jody L. Mackenzie-Grieve and John R. Post. Projected impacts of climate warming on production of lake trout (Salvelinus namaycush) in southern Yukon lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 63 (4):788–797, April 2006. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f05-257.

Robillard:2006:HCA

[1387] M. M. Robillard and M. G. Fox. Historical changes in abundance and community structure of warmwater piscivore communities associated with changes in water clarity, nutrients, and temperature. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 63(4):798–809, April 2006. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f05-259.

Duplisea:2006:CUD

[1388] Daniel E. Duplisea and Martin Castonguay. Comparison and utility of different size-based metrics of fish communities for detecting fishery impacts. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 63(4):810–820, April 2006. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f05-261.

Stenroth:2006:SII

[1389] Patrik Stenroth, Niklas Holmqvist, Per Nyström, Olof Berglund, Per Larsson, and Wilhelm Granéli. Stable isotopes as an indicator of diet in omnivorous crayfish (Pacifastacus leniusculus): the influence of tissue, sample treatment, and season. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 63(4):821–831, April 2006. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f05-265.

Bunnell:2006:LTT

[1390] David B. Bunnell, Charles P. Madenjian, and Thomas E. Croley II. Long-term trends of bloater (Coregonus hoyi) recruitment in Lake Michigan: evidence for the effect of sex ratio. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 63(4):832–844, April 2006. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f05-271.

Spears:2006:BPA

[1391] Bryan M. Spears and Lance F. W. Lesack. Bacterioplankton production, abundance, and nutrient limitation among lakes of the Mackenzie Delta (western Canadian arctic). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 63 (4):845–857, April 2006. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f05-264.

Medici:2006:SDM

[1392] David A. Medici, Thomas G. Wolcott, and Donna L. Wolcott. Scaledependent movements and protection of female blue crabs (Callinectes sapidus). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 63(4):858–871, April 2006. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f05-263.

Nalepa:2006:CDB

[1393] Thomas F. Nalepa, David L. Fanslow, Andrew J. Foley III, Gregory A. Lang, Brian J. Eadie, and Michael A. Quigley. Continued disappearance of the benthic amphipod Diporeia spp. in Lake Michigan: is there evidence for food limitation? Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 63 (4):872–890, April 2006. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f05-262.

Madenjian:2006:CSE

[1394] Charles P. Madenjian, Steven A. Pothoven, John M. Dettmers, and Jeffrey D. Holuszko. Changes in seasonal energy dynamics of alewife (Alosa pseudoharengus) in Lake Michigan after invasion of dreissenid mussels. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 63(4):891–902, April 2006. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f06-017.

Soboil:2006:EAT

[1395] Mark Lucas Soboil and Jon G. Sutinen. Empirical analysis and transboundary management for Georges Bank multispecies fishery. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 63(4):903–916, April 2006. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f05-269.

Lang:2006:IFS

[1396] Dirk W. Lang, Gordon H. Reeves, James D. Hall, and Mark S. Wipfli. The influence of fall-spawning Coho salmon (Oncorhynchus kisutch) on growth and production of juvenile Coho salmon rearing in beaver ponds on the Copper River Delta, Alaska. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 63(4):917–930, April 2006. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f05-268.

Gallucci:2006:CME

[1397] Vincent F. Gallucci, Ian G. Taylor, and Karim Erzini. Conservation and management of exploited shark populations based on reproductive value. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 63(4):931–942, April 2006. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f05-267.

Travers:2006:STS

[1398] Morgane Travers, Yunne-Jai Shin, Lynne Shannon, and Philippe Cury. Simulating and testing the sensitivity of ecosystem-based indicators to fishing in the southern Benguela ecosystem. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 63(4):943–956, April 2006. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f06-003.

Purchase:2006:TJA

[1399] C. F. Purchase, H. P. Benoît, D. L. Boyce, G. E. Brown, V. Gotceitas, B. J. Laurel, V. Puvanendran, and R. J. Thompson. Tribute: Joseph Albert Brown, Ph.D. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 63(5), May 2006. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f06-041.

Purchase:2006:HJA

[1400] C. F. Purchase, H. P. Benoît, D. L. Boyce, G. E. Brown, V. Gotceitas, B. J. Laurel, V. Puvanendran, and R. J. Thompson. Hommage: Joseph Albert Brown, Ph.D. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 63(5), May 2006. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/f06-
051.

Bruesewitz:2006:SEZ

[1401] Denise A. Bruesewitz, Jennifer L. Tank, Melody J. Bernot, William B. Richardson, and Eric A. Strauss. Seasonal effects of the zebra mussel (Dreissena polymorpha) on sediment denitrification rates in Pool 8 of the Upper Mississippi River. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 63 (5):957–969, May 2006. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f06-002.

Kaufman:2006:MER

[1402] Scott D. Kaufman, John M. Gunn, George E. Morgan, and Patrice Couture. Muscle enzymes reveal walleye (Sander vitreus) are less active when larger prey (cisco, Coregonus artedi) are present. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 63(5):970–979, May 2006. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f06-004.

Marshall:2006:SBE

[1403] C. Tara Marshall, Coby L. Needle, Anders Thorsen, Olav Sigurd Kjesbu, and Nathalia A. Yaragina. Systematic bias in estimates of reproductive potential of an Atlantic cod (Gadus morhua) stock: implications for stock–recruit theory and management. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 63(5):980–994, May 2006. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f05-270.

Courtemanche:2006:AAB

[1404] David A. Courtemanche, Frederick G. Whoriskey, Jr., Valerie Bujold, and R. Allen Curry. Assessing anadromy of brook char (Salvelinus fontinalis) using scale microchemistry. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 63(5):995–1006, May 2006. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f06-009.

Jeziorski:2006:SIA

[1405] Adam Jeziorski and Norman D. Yan. Species identity and aqueous calcium concentrations as determinants of calcium concentrations of freshwater crustacean zooplankton. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 63(5):1007–1013, May 2006. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f06-019.

Vis:2006:IVS

[1406] Chantal Vis, Christiane Hudon, and Richard Carignan. Influence of the vertical structure of macrophyte stands on epiphyte community metabolism. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 63(5):1014– 1026, May 2006. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f06-021.

Andrews:2006:MSD

[1407] Jessica M. Andrews, William S. C. Gurney, Michael R. Heath, Alejandro Gallego, Carl M. O’Brien, Christopher Darby, and Graham Tyldesley. Modelling the spatial demography of Atlantic cod (Gadus morhua) on the European continental shelf. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 63 (5):1027–1048, May 2006. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f06-006.

Ward:2006:FPE

[1408] Darren M. Ward and Raymond M. Newman. Fish predation on Eurasian watermilfoil (Myriophyllum spicatum) herbivores and indirect effects on macrophytes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 63(5):1049– 1057, May 2006. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f06-010.

Drevnick:2006:ODM

[1409] Paul E. Drevnick, Martin J. Horgan, James T. Oris, and Boyd E. Kynard. Ontogenetic dynamics of mercury accumulation in Northwest Atlantic sea lamprey (Petromyzon marinus). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 63(5):1058–1066, May 2006. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f06-012.

Young:2006:PEC

[1410] Jeffery L. Young, Scott G. Hinch, Steven J. Cooke, Glenn T. Crossin, David A. Patterson, Anthony P. Farrell, Glen Van Der Kraak, Andrew G. Lotto, Andrea Lister, Michael C. Healey, and Karl K. English. Physiological and energetic correlates of en route mortality for abnormally early migrating adult sockeye salmon (Oncorhynchus nerka) in the Thompson River, British Columbia. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 63 (5):1067–1077, May 2006. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f06-014.

Frisk:2006:AGL

[1411] Michael G. Frisk and Thomas J. Miller. Age, growth, and latitudinal patterns of two Rajidae species in the northwestern Atlantic: little skate (Leucoraja erinacea) and winter skate (Leucoraja ocellata). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 63(5):1078–1091, May 2006. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f06-005.

Miller:2006:IDM

[1412] T. J. Miller and J. R. Skalski. Integrating design- and model-based inference to estimate length and age composition in North Pacific longline catches. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 63(5):1092–1114, May 2006. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f06-022.

Depew:2006:NOC

[1413] David C. Depew, Stephanie J. Guildford, and Ralph E. H. Smith. Nearshore–offshore comparison of chlorophyll a and phytoplankton production in the dreissenid-colonized eastern basin of Lake Erie. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 63(5):1115–1129, May 2006. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f06-016.

Lauster:2006:GPP

[1414] George H. Lauster, Paul C. Hanson, and Timothy K. Kratz. Gross primary production and respiration differences among littoral and pelagic habitats in northern Wisconsin lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 63(5):1130–1141, May 2006. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f06-018.

Alexander:2006:DAF

[1415] Clint A. D. Alexander, Calvin N. Peters, David R. Marmorek, and Paul Higgins. A decision analysis of flow management experiments for Columbia River mountain whitefish (Prosopium williamsoni) management. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 63(5):1142–1156, May 2006. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f06-032.

Taylor:2006:MDA

[1416] Eric B. Taylor and A. B. Costello. Microsatellite DNA analysis of coastal populations of bull trout (Salvelinus confluentus) in British Columbia: zoogeographic implications and its application to recreational fishery management. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 63(5):1157– 1171, May 2006. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f06-026.

Minns:2006:CRN

[1417] Charles K. Minns. Compensation ratios needed to offset timing effects of losses and gains and achieve no net loss of productive capacity of fish habitat. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 63(5):1172–1182, May 2006. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f06-025.

Holland:2006:CFI

[1418] Daniel S. Holland and T. Kevin Stokes. Comment on “Fishing and the impact of marine reserves in a variable environment”. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 63(5):1183–1185, May 2006. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f06-023.

Rodwell:2006:RCH

[1419] Lynda D. Rodwell and Callum M. Roberts. Reply to the comment by Holland and Stokes on “Fishing and the impact of marine reserves in a variable environment”. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 63 (5):1186–1188, May 2006. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f06-024.

Morita:2006:EPD

[1420] Kentaro Morita, Shoko H. Morita, and Masa aki Fukuwaka. Erratum: Population dynamics of Japanese pink salmon (Oncorhynchus gorbuscha): are recent increases explained by hatchery programs or climatic variations? Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 63(5):1189, May 2006. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f06-027.

Dodds:2006:ENP

[1421] Walter K. Dodds, Val H. Smith, and Kirk Lohman. Erratum: Nitrogen and phosphorus relationships to benthic algal biomass in temperate streams. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 63(5):1190–1191, May 2006. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f06-040.

Thorrold:2006:TME

[1422] Simon R. Thorrold, Geoffrey P. Jones, Serge Planes, and Jonathan A. Hare. Transgenerational marking of embryonic otoliths in marine fishes using barium stable isotopes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 63 (6):1193–1197, June 2006. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f06-048.

Richard:2006:ISM

[1423] Marion Richard, Philippe Archambault, Gérard Thouzeau, and Gaston Desrosiers. Influence of suspended mussel lines on the biogeochemical fluxes in adjacent water in the Îles-de-la-Madeleine (Quebec, Canada). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 63(6):1198–1213, June 2006. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f06-030.

Baskett:2006:PCR

[1424] Marissa L. Baskett, Mary Yoklavich, and Milton S. Love. Predation, competition, and the recovery of overexploited fish stocks in marine reserves. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 63(6):1214–1229, June 2006. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f06-013.

Claeson:2006:RNB

[1425] Shannon M. Claeson, Judith L. Li, Jana E. Compton, and Peter A. Bisson. Response of nutrients, biofilm, and benthic insects to salmon carcass addition. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 63(6):1230–1241, June 2006. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f06-029.

Beechie:2006:IPU

[1426] Timothy J. Beechie, Correigh M. Greene, Lisa Holsinger, and Eric M. Beamer. Incorporating parameter uncertainty into evaluation of spawning habitat limitations on Chinook salmon (Oncorhynchus tshawytscha) populations. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 63(6):1242– 1250, June 2006. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f06-015.

Lincoln-Smith:2006:UIA

[1427] M. P. Lincoln-Smith, K. A. Pitt, J. D. Bell, and B. D. Mapstone. Using impact assessment methods to determine the effects of a marine reserve on abundances and sizes of valuable tropical invertebrates. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 63(6):1251–1266, June 2006. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f06-033.

Dulvy:2006:TDF

[1428] Nicholas K. Dulvy, Simon Jennings, Stuart I. Rogers, and David L. Maxwell. Threat and decline in fishes: an indicator of marine biodiversity. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 63(6):1267–1275, June 2006. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f06-035.

Rosenthal:2006:WLE

[1429] Sadie K. Rosenthal, Samantha S. Stevens, and David M. Lodge. Wholelake effects of invasive crayfish (Orconectes spp.) and the potential for restoration. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 63(6):1276–1285, June 2006. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/f06-
037.

Larocque:2006:FID

[1430] Isabelle Larocque, Reinhard Pienitz, and Nicolas Rolland. Factors influencing the distribution of chironomids in lakes distributed along a latitudinal gradient in northwestern Quebec, Canada. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 63(6):1286–1297, June 2006. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f06-020.

Brylawski:2006:TDG

[1431] Bryce J. Brylawski and Thomas J. Miller. Temperature-dependent growth of the blue crab (Callinectes sapidus): a molt process approach. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 63(6):1298–1308, June 2006. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f06-011.

Heggenes:2006:GSR

[1432] Jan Heggenes, Tore Qvenild, Michael D. Stamford, and Eric B. Taylor. Genetic structure in relation to movements in wild European grayling (Thymallus thymallus) in three Norwegian rivers. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 63(6):1309–1319, June 2006. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f06-028.

Sundberg:2006:CCA

[1433] Henrik Sundberg, Rasha Ishaq, Ulla Tjärnlund, Gun Åkerman, Kerstin Grunder, Cecilia Bandh, Dag Broman, and Lennart Balk. Contribution of commonly analyzed polycyclic aromatic hydrocarbons (PAHs) to potential toxicity in early life stages of rainbow trout (Oncorhynchus mykiss). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 63(6):1320–1333, June 2006. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f06-034.

Lorenzen:2006:AYF

[1434] Kai Lorenzen, Oriana Almeida, Robert Arthur, Caroline Garaway, and Sophie Nguyen Khoa. Aggregated yield and fishing effort in multispecies fisheries: an empirical analysis. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 63(6):1334–1343, June 2006. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f06-038.

Hall:2006:LBM

[1435] Stephen J. Hall, Jeremy S. Collie, Daniel E. Duplisea, Simon Jennings, Mark Bravington, and Jason Link. A length-based multispecies model for evaluating community responses to fishing. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 63(6):1344–1359, June 2006. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f06-039.

Allen:2006:EOS

[1436] Peter J. Allen, Brian Hodge, Inge Werner, and Joseph J. Cech, Jr. Effects of ontogeny, season, and temperature on the swimming performance of juvenile green sturgeon (Acipenser medirostris). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 63(6):1360–1369, June 2006. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f06-031.

Heath:2006:PSD

[1437] Daniel D. Heath, J. Mark Shrimpton, Russell I. Hepburn, Sara K. Jamieson, Sarah K. Brode, and Margaret F. Docker. Population structure and divergence using microsatellite and gene locus markers in Chinook salmon (Oncorhynchus tshawytscha) populations. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 63(6):1370–1383, June 2006. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f06-044.

Madej:2006:APT

[1438] M. A. Madej, C. Currens, V. Ozaki, J. Yee, and D. G. Anderson. Assessing possible thermal rearing restrictions for juvenile Coho salmon (Oncorhynchus kisutch) through thermal infrared imaging and in-stream monitoring, Redwood Creek, California. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 63(6):1384–1396, June 2006. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f06-043.

Thum:2006:PSS

[1439] Ryan A. Thum and Richard S. Stemberger. Pure spatial and spatially structured environmental variables explain Skistodiaptomus copepod range limits in the northeastern USA. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 63(6):1397–1404, June 2006. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f06-046.

Ninness:2006:EED

[1440] Marcie M. Ninness, E. Don Stevens, and Patricia A. Wright. Energy expenditure during hatching in rainbow trout (Oncorhynchus mykiss) embryos. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 63(6):1405–1413, June 2006. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f06-036.

Fablet:2006:SLE

[1441] Ronan Fablet. Semi-local extraction of ring structures in images of biological hard tissues: application to the Bayesian interpretation of fish otoliths for age and growth estimation. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 63(6):1414–1428, June 2006. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f06-001.

Girard:2006:IPM

[1442] Philippe Girard and Bernard Angers. The impact of postglacial marine invasions on the genetic diversity of an obligate freshwater fish, the longnose dace (Rhinichthys cataractae), on the Quebec peninsula. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 63(6):1429–1438, June 2006. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f06-047.

Anonymous:2006:AC

[1443] Anonymous. Announcement / communiqué. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 63(7), July 2006. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/

doi/full/10.1139/f06-900.

Schindler:2006:YCR

[1444] Daniel E. Schindler, Peter R. Leavitt, Susan P. Johnson, and Curtis S. Brock. A 500-year context for the recent surge in sockeye salmon (Oncorhynchus nerka) abundance in the Alagnak River, Alaska. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 63(7):1439–1444, July 2006. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f06-069.

Hart:2006:WDM

[1445] Deborah R. Hart. When do marine reserves increase fishery yield? Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 63(7):1445–1449, July 2006. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f06-071.

Levasseur:2006:ESV

[1446] Mylène Levasseur, Normand E. Bergeron, Michel F. Lapointe, and Francis Bérubé. Effects of silt and very fine sand dynamics in Atlantic salmon (Salmo salar) redds on embryo hatching success. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 63(7):1450–1459, July 2006. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f06-050.

Powles:2006:DEM

[1447] Perce M. Powles, Jonathan A. Hare, Elisabeth H. Laban, and Stanley M. Warlen. Does eel metamorphosis cause a breakdown in the tenets of otolith applications? A case study using the speckled worm eel (Myrophis punctatus, ophichthidae). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 63 (7):1460–1468, July 2006. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f06-052.

Cooke:2006:PIL

[1448] Steven J. Cooke, Scott G. Hinch, Glenn T. Crossin, David A. Patterson, Karl K. English, J. Mark Shrimpton, Glen Van Der Kraak, and Anthony P. Farrell. Physiology of individual late-run Fraser River sockeye salmon (Oncorhynchus nerka) sampled in the ocean correlates with fate during spawning migration. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 63 (7):1469–1480, July 2006. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f06-042.

Holland:2006:PMB

[1449] Daniel S. Holland and Kurt E. Schnier. Protecting marine biodiversity: a comparison of individual habitat quotas and marine protected areas. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 63(7):1481–1495, July 2006. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f06-049.

Sierszen:2006:DSP

[1450] Michael E. Sierszen, Gregory S. Peterson, and Jill V. Scharold. Depthspecific patterns in benthic–planktonic food web relationships in Lake Superior. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 63(7):1496–1503, July 2006. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f06-057.

Whitfield:2006:RAN

[1451] C. J. Whitfield, J. Aherne, S. A. Watmough, P. J. Dillon, and T. A. Clair. Recovery from acidification in Nova Scotia: temporal trends and critical loads for 20 headwater lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 63 (7):1504–1514, July 2006. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f06-053.

Mossop:2006:UTP

[1452] Brent Mossop and Michael J. Bradford. Using thalweg profiling to assess and monitor juvenile salmon (Oncorhynchus spp.) habitat in small streams. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 63(7):1515–1525, July 2006. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f06-060.

Zimmerman:2006:ISS

[1453] Julie K. H. Zimmerman and Bruce Vondracek. Interactions of slimy sculpin (Cottus cognatus) with native and nonnative trout: consequences for growth. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 63(7):1526–1535, July 2006. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/f06-
054.

Eastwood:2006:MSA

[1454] Paul D. Eastwood, Sami Souissi, Stuart I. Rogers, Roger A. Coggan, and Craig J. Brown. Mapping seabed assemblages using comparative top-down and bottom-up classification approaches. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 63(7):1536–1548, July 2006. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f06-058.

Barbiero:2006:CSP

[1455] Richard P. Barbiero, David C. Rockwell, Glenn J. Warren, and Marc L. Tuchman. Changes in spring phytoplankton communities and nutrient dynamics in the eastern basin of Lake Erie since the invasion of Dreissena spp. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 63(7):1549–1563, July 2006. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f06-059.

Flynn:2006:HMS

[1456] Lucy Flynn, André E. Punt, and Ray Hilborn. A hierarchical model for salmon run reconstruction and application to the Bristol Bay sockeye salmon (Oncorhynchus nerka) fishery. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 63(7):1564–1577, July 2006. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f06-045.

Bartz:2006:TRS

[1457] Krista K. Bartz, Kerry M. Lagueux, Mark D. Scheuerell, Tim Beechie, Andrew D. Haas, and Mary H. Ruckelshaus. Translating restoration scenarios into habitat conditions: an initial step in evaluating recovery strategies for Chinook salmon (Oncorhynchus tshawytscha). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 63(7):1578–1595, July 2006. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f06-055.

Scheuerell:2006:SMT

[1458] Mark D. Scheuerell, Ray Hilborn, Mary H. Ruckelshaus, Krista K. Bartz, Kerry M. Lagueux, Andrew D. Haas, and Kit Rawson. The shiraz model: a tool for incorporating anthropogenic effects and fish–habitat relationships in conservation planning. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 63(7):1596–1607, July 2006. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f06-056.

Fortier:2006:SAC

[1459] Louis Fortier, Pascal Sirois, Josée Michaud, and David Barber. Survival of Arctic cod larvae (Boreogadus saida) in relation to sea ice and temperature in the Northeast Water Polynya (Greenland Sea). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 63(7):1608–1616, July 2006. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f06-064.

Angeler:2006:IRP

[1460] David G. Angeler and José M. Moreno. Impact-recovery patterns of water quality in temporary wetlands after fire retardant pollution. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 63(7):1617–1626, July 2006. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f06-062.

Hill:2006:CBH

[1461] Megan S. Hill, Gayle Barbin Zydlewski, and William L. Gale. Comparisons between hatchery and wild steelhead trout (Oncorhynchus mykiss) smolts: physiology and habitat use. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 63(7):1627–1638, July 2006. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f06-061.

Smith:2006:IPD

[1462] Stephen J. Smith and Mark J. Lundy. Improving the precision of designbased scallop drag surveys using adaptive allocation methods. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 63(7):1639–1646, July 2006. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f06-063.

Branch:2006:FDF

[1463] Trevor A. Branch, Ray Hilborn, Alan C. Haynie, Gavin Fay, Lucy Flynn, Jennifer Griffiths, Kristin N. Marshall, Jeffrey K. Randall, Jennifer M. Scheuerell, Eric J. Ward, and Mark Young. Fleet dynamics and fishermen behavior: lessons for fisheries managers. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 63(7):1647–1668, July 2006. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f06-072.

Anonymous:2006:TSN

[1464] Anonymous. Tribute: Simeon Noel Tibbo (1914–2003). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 63(8), August 2006. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f06-901.

Anonymous:2006:HSN

[1465] Anonymous. Hommage: Simeon Noel Tibbo (1914–2003). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 63(8), August 2006. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f06-902.

Holland:2006:FCB

[1466] Daniel S. Holland and Guillermo E. Herrera. Flexible catch-balancing policies for multispecies individual fishery quotas. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 63(8):1669–1685, August 2006. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f06-066.

Reitzel:2006:CPS

[1467] Kasper Reitzel, Joakim Ahlgren, Adolf Gogoll, Henning S. Jensen, and Emil Rydin. Characterization of phosphorus in sequential extracts from lake sediments using^{31}P nuclear magnetic resonance spectroscopy. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 63(8):1686–1699, August 2006. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f06-070.

Kennedy:2006:EDB

[1468] Benjamen M. Kennedy, Brett W. Thompson, and Chris Luecke. Ecological differences between two closely related morphologically similar benthic whitefish (Prosopium spilonotus and Prosopium abyssicola) in an endemic whitefish complex. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 63 (8):1700–1709, August 2006. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f06-065.

Nelson:2006:ISD

[1469] Marcy L. Nelson, James R. Gilbert, and Kevin J. Boyle. The influence of siting and deterrence methods on seal predation at Atlantic salmon (Salmo salar) farms in Maine, 2001–2003. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 63(8):1710–1721, August 2006. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f06-067.

Galarowicz:2006:IPA

[1470] Tracy L. Galarowicz, Julie A. Adams, and David H. Wahl. The influence of prey availability on ontogenetic diet shifts of a juvenile piscivore. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 63(8):1722–1733, August 2006. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f06-073.

Desrosiers:2006:STR

[1471] Mélanie Desrosiers, Dolors Planas, and Alfonso Mucci. Short-term responses to watershed logging on biomass mercury and methylmercury accumulation by periphyton in boreal lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 63(8):1734–1745, August 2006. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f06-077.

Whitledge:2006:SHI

[1472] Gregory W. Whitledge, Brett M. Johnson, and Patrick J. Martinez. Stable hydrogen isotopic composition of fishes reflects that of their environment. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 63(8):1746–1751, August 2006. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/f06-
076.

Keefer:2006:RSL

[1473] Matthew L. Keefer, Christopher C. Caudill, Christopher A. Peery, and Theodore C. Bjornn. Route selection in a large river during the homing migration of Chinook salmon (Oncorhynchus tshawytscha). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 63(8):1752–1762, August 2006. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f06-068.

Cyr:2006:DBE

[1474] Hélène Cyr and Kathryn E. Morton. Distribution of biofilm exopolymeric substances in littoral sediments of Canadian Shield lakes: the effects of light and substrate. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 63(8): 1763–1776, August 2006. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f06-079.

Apostolaki:2006:CDP

[1475] P. Apostolaki, E. A. Babcock, and M. K. McAllister. Contrasting deterministic and probabilistic ranking of catch quotas and spatially and size-regulated fisheries management. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 63(8):1777–1792, August 2006. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f06-078.

Lavoie:2006:DBI

[1476] Isabelle Lavoie, Stéphane Campeau, Martine Grenier, and Peter J. Dillon. A diatom-based index for the biological assessment of eastern Canadian rivers: an application of correspondence analysis (CA). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 63(8):1793–1811, August 2006. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f06-084.

Olden:2006:ILT

[1477] Julian D. Olden, Olaf P. Jensen, and M. Jake Vander Zanden. Implications of long-term dynamics of fish and zooplankton communities for among-lake comparisons. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 63 (8):1812–1821, August 2006. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f06-082.

Magnoni:2006:ELD

[1478] Leonardo J. Magnoni, David A. Patterson, Anthony P. Farrell, and Jean-Michel Weber. Effects of long-distance migration on circulating lipids of sockeye salmon (Oncorhynchus nerka). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 63(8):1822–1829, August 2006. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f06-083.

Hardie:2006:EIC

[1479] David C. Hardie, Roxanne M. Gillett, and Jeffrey A. Hutchings. The effects of isolation and colonization history on the genetic structure of marine-relict populations of Atlantic cod (Gadus morhua) in the Canadian Arctic. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 63(8):1830–1839, August 2006. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f06-085.

Sandahl:2006:OIR

[1480] Jason F. Sandahl, Go Miyasaka, Nobuhisa Koide, and Hiroshi Ueda. Olfactory inhibition and recovery in chum salmon (Oncorhynchus keta) following copper exposure. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 63 (8):1840–1847, August 2006. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f06-074.

Islam:2006:SVN

[1481] Md. Shahidul Islam and Masaru Tanaka. Spatial variability in nursery functions along a temperate estuarine gradient: role of detrital versus algal trophic pathways. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 63(8): 1848–1864, August 2006. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f06-086.

Kefford:2006:VSS

[1482] Ben J. Kefford, Dayanthi Nugegoda, Leon Metzeling, and Elizabeth J. Fields. Validating species sensitivity distributions using salinity tolerance of riverine macroinvertebrates in the southern Murray–Darling Basin (Victoria, Australia). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 63(8): 1865–1877, August 2006. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f06-080.

Miller:2006:ESB

[1483] T. J. Miller and J. R. Skalski. Estimation of seabird bycatch for North Pacific longline vessels using design- and model-based methods. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 63(8):1878–1889, August 2006. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f06-075.

Trudel:2006:BMD

[1484] Marc Trudel and Joseph B. Rasmussen. Bioenergetics and mercury dynamics in fish: a modelling perspective. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 63(8):1890–1902, August 2006. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f06-081.

Zambrano:2006:IPC

[1485] Luis Zambrano, Enrique Martínez-Meyer, Naercio Menezes, and A. Townsend Peterson. Invasive potential of common carp (Cyprinus carpio) and Nile tilapia (Oreochromis niloticus) in American freshwater systems. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 63(9):1903–1910, September 2006. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f06-088.

Peer:2006:FYG

[1486] Adam C. Peer, Dennis R. DeVries, and Russell A. Wright. Firstyear growth and recruitment of coastal largemouth bass (Micropterus salmoides): spatial patterns unresolved by critical periods along a salinity gradient. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 63(9):1911– 1924, September 2006. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f06-089.

Tremblay:2006:CLA

[1487] M. J. Tremblay, S. J. Smith, D. A. Robichaud, and P. Lawton. The catchability of large American lobsters (Homarus americanus) from diving and trapping studies off Grand Manan Island, Canadian maritimes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 63(9):1925–1933, September 2006. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f06-090.

Wiedenmann:2006:DDO

[1488] John Wiedenmann and Timothy E. Essington. Density-dependent overwinter survival in young-of-year bluefish (Pomatomus saltatrix)? A new approach for assessing stage-structured survival. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 63(9):1934–1943, September 2006. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f06-094.

Cadigan:2006:RSR

[1489] Noel G. Cadigan and John Brattey. Reporting and shedding rate estimates from tag-recovery experiments on Atlantic cod (Gadus morhua) in coastal Newfoundland. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 63(9): 1944–1958, September 2006. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f06-096.

Imsland:2006:MEL

[1490] Albert K. Imsland, Nathalie R. Le François, Simon G. Lamarre, Delphine Ditlecadet, Sindri Sigurdsson, and Atle Foss. Myosin expression levels and enzyme activity in juvenile spotted wolffish (Anarhichas minor) muscle: a method for monitoring growth rates. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 63(9):1959–1967, September 2006. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f06-091.

Michielsens:2006:EAM

[1491] Catherine G. J. Michielsens, Samu Mäntyniemi, and Pekka J. Vuorinen. Estimation of annual mortality rates caused by early mortality syndromes (EMS) and their impact on salmonid stock–recruit relationships. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 63(9):1968–1981, September 2006. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f06-095.

Grellier:2006:RDP

[1492] Kate Grellier and Philip S. Hammond. Robust digestion and passage rate estimates for hard parts of grey seal (Halichoerus grypus) prey. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 63(9):1982–1998, September 2006. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f06-092.

Hogsden:2006:BGF

[1493] Kristy L. Hogsden and Rolf D. Vinebrooke. Benthic grazing and functional compensation in stressed and recovered lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 63(9):1999–2010, September 2006. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f06-098.

Morinville:2006:MFP

[1494] Geneviève R. Morinville and Joseph B. Rasmussen. Marine feeding patterns of anadromous brook trout (Salvelinus fontinalis) inhabiting an estuarine river fjord. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 63(9): 2011–2027, September 2006. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f06-097.

Lejeusne:2006:BEH

[1495] Christophe Lejeusne, Thierry Pérez, Vanessa Sarrazin, and Pierre Chevaldonné. Baseline expression of heat-shock proteins (HSPs) of a “thermotolerant” Mediterranean marine species largely influenced by natural temperature fluctuations. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 63(9):2028–2037, September 2006. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f06-102.

Smokorowski:2006:EPM

[1496] Karen E. Smokorowski, Thomas C. Pratt, William G. Cole, Laurie J. McEachern, and Elaine C. Mallory. Effects on periphyton and macroinvertebrates from removal of submerged wood in three Ontario lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 63(9):2038–2049, September 2006. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f06-104.

Velez-Espino:2006:CEP

[1497] Luis A. Vélez-Espino, Michael G. Fox, and Robert L. McLaughlin. Characterization of elasticity patterns of North American freshwater fishes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 63(9):2050–2066, September 2006. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f06-093.

Aubin-Horth:2006:LGT

[1498] Nadia Aubin-Horth, Jean-Francois Bourque, Gaetan Daigle, Richard Hedger, and Julian J. Dodson. Longitudinal gradients in threshold sizes for alternative male life history tactics in a population of Atlantic salmon (Salmo salar). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 63(9):2067– 2075, September 2006. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f06-103.

Hocking:2006:CDS

[1499] Morgan D. Hocking and Thomas E. Reimchen. Consumption and distribution of salmon (Oncorhynchus spp.) nutrients and energy by terrestrial flies. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 63(9):2076– 2086, September 2006. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f06-110.

Grenier:2006:DRC

[1500] Martine Grenier, Stéphane Campeau, Isabelle Lavoie, Young Seuk Park, and Sovan Lek. Diatom reference communities in Québec (Canada) streams based on Kohonen self-organizing maps and multivariate analyses. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 63(9):2087–2106, September 2006. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f06-101.

Febria:2006:PHP

[1501] Catherine M. Febria, Lance F. W. Lesack, Jolie A. L. Gareis, and Max L. Bothwell. Patterns of hydrogen peroxide among lakes of the Mackenzie Delta, western Canadian Arctic. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 63(9):2107–2118, September 2006. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f06-106.

Casselman:2006:SQI

[1502] Stephen J. Casselman, Albrecht I. Schulte-Hostedde, and Robert Montgomerie. Sperm quality influences male fertilization success in walleye (Sander vitreus). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 63(9):2119– 2125, September 2006. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f06-108.

Strecker:2006:VRC

[1503] Angela L. Strecker, Shelley E. Arnott, Norman D. Yan, and Robert Girard. Variation in the response of crustacean zooplankton species richness and composition to the invasive predator Bythotrephes longimanus. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 63(9):2126–2136, September 2006. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f06-105.

Griffiths:2006:THS

[1504] Natalie A. Griffiths and Hélène Cyr. Are there hot spots for Elliptio complanata in the shallow littoral zone of a large Canadian Shield lake? Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 63(9):2137–2147, September 2006. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f06-100.

Hilborn:2006:COM

[1505] Ray Hilborn, John Annala, and Daniel S. Holland. The cost of overfishing and management strategies for new fisheries on slow-growing fish: orange roughy (Hoplostethus atlanticus) in New Zealand. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 63(10):2149–2153, October 2006. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f06-115.

Sibert:2006:IVL

[1506] John R. Sibert, Molly E. Lutcavage, Anders Nielsen, Richard W. Brill, and Steven G. Wilson. Interannual variation in large-scale movement of Atlantic bluefin tuna (Thunnus thynnus) determined from pop-up satellite archival tags. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 63(10): 2154–2166, October 2006. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f06-114.

Murry:2006:ELE

[1507] Brent A. Murry, John M. Farrell, Mark A. Teece, and Peter M. Smyntek. Effect of lipid extraction on the interpretation of fish community trophic relationships determined by stable carbon and nitrogen isotopes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 63(10):2167–2172, October 2006. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f06-116.

Hoxmeier:2006:GSA

[1508] R. John H. Hoxmeier, David H. Wahl, Ronald C. Brooks, and Roy C. Heidinger. Growth and survival of age-0 walleye (Sander vitreus): interactions among walleye size, prey availability, predation, and abiotic factors. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 63(10):2173–2182, October 2006. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f06-087.

Leech:2006:UVF

[1509] Dina M. Leech and Sönke Johnsen. Ultraviolet vision and foraging in juvenile bluegill (Lepomis macrochirus). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 63(10):2183–2190, October 2006. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f06-107.

Baumann:2006:RVB

[1510] H. Baumann, H.-H. Hinrichsen, C. Möllmann, F. W. Köster, A. M. Malzahn, and A. Temming. Recruitment variability in Baltic Sea sprat (Sprattus sprattus) is tightly coupled to temperature and transport patterns affecting the larval and early juvenile stages. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 63(10):2191–2201, October 2006. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f06-112.

Schulze:2006:RRP

[1511] Torsten Schulze, Ulrich Baade, Hendrik Dörner, Reiner Eckmann, Susanne S. Haertel-Borer, Franz Hölker, and Thomas Mehner. Response of the residential piscivorous fish community to introduction of a new predator type in a mesotrophic lake. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 63(10):2202–2212, October 2006. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f06-099.

Paterson:2006:BND

[1512] Michael J. Paterson, Paul J. Blanchfield, Cheryl Podemski, Holger H. Hintelmann, Cynthia C. Gilmour, Reed Harris, Nives Ogrinc, John W. M. Rudd, and Ken A. Sandilands. Bioaccumulation of newly deposited mercury by fish and invertebrates: an enclosure study using stable mercury isotopes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 63(10): 2213–2224, October 2006. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f06-118.

Jech:2006:MMC

[1513] J. Michael Jech and William L. Michaels. A multifrequency method to classify and evaluate fisheries acoustics data. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 63(10):2225–2235, October 2006. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f06-126.

Marcarelli:2006:SCP

[1514] Amy M. Marcarelli, Wayne A. Wurtsbaugh, and Olivia Griset. Salinity controls phytoplankton response to nutrient enrichment in the Great Salt Lake, Utah, USA. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 63(10): 2236–2248, October 2006. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f06-113.

OFarrell:2006:FMI

[1515] Michael R. O’Farrell and Louis W. Botsford. The fisheries management implications of maternal-age-dependent larval survival. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 63(10):2249–2258, October 2006. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f06-130.

Marriott:2006:CIC

[1516] Ross J. Marriott and Bruce D. Mapstone. Consequences of inappropriate criteria for accepting age estimates from otoliths, with a case study for a long-lived tropical reef fish. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 63(10):2259–2274, October 2006. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f06-109.

Wilberg:2006:PTV

[1517] Michael J. Wilberg and James R. Bence. Performance of time-varying catchability estimators in statistical catch-at-age analysis. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 63(10):2275–2285, October 2006. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f06-111.

Hrabik:2006:DVM

[1518] Thomas R. Hrabik, Olaf P. Jensen, Steven J. D. Martell, Carl J. Walters, and James F. Kitchell. Diel vertical migration in the Lake Superior pelagic community. I. changes in vertical migration of coregonids in response to varying predation risk. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 63(10):2286–2295, October 2006. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f06-124.

Jensen:2006:DVM

[1519] Olaf P. Jensen, Thomas R. Hrabik, Steven J. D. Martell, Carl J. Walters, and James F. Kitchell. Diel vertical migration in the Lake Superior pelagic community. II. Modeling trade-offs at an intermediate trophic level. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 63(10):2296–2307, October 2006. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/f06-
125.

Ghadouani:2006:CIC

[1520] Anas Ghadouani, Bernadette Pinel-Alloul, and Ellie E. Prepas. Could increased cyanobacterial biomass following forest harvesting cause a reduction in zooplankton body size structure? Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 63(10):2308–2317, October 2006. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f06-117.

LAbee-Lund:2006:DLM

[1521] Jan Henning L’Abée-Lund, Thrond O. Haugen, and Leif Asbjørn Vøllestad. Disentangling local from macroenvironmental effects: quantifying the effect of human encroachments based on historical river catches of anadromous salmonids. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 63 (10):2318–2329, October 2006. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/

doi/full/10.1139/f06-123.

Fujioka:2006:MEF

[1522] Jeffrey T. Fujioka. A model for evaluating fishing impacts on habitat and comparing fishing closure strategies. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 63(10):2330–2342, October 2006. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f06-120.

Ford:2006:CRT

[1523] Michael J. Ford, Howard Fuss, Brant Boelts, Eric LaHood, Jeffrey Hard, and Jason Miller. Changes in run timing and natural smolt production in a naturally spawning Coho salmon (Oncorhynchus kisutch) population after 60 years of intensive hatchery supplementation. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 63(10):2343–2355, October 2006. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f06-119.

Leduc:2006:IDC

[1524] Antoine Olivier Henri Claude Leduc, Ellie Roh, Mark Charles Harvey, and Grant Edwin Brown. Impaired detection of chemical alarm cues by juvenile wild Atlantic salmon (Salmo salar) in a weakly acidic environment. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 63(10):2356–2363, October 2006. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/f06-
128.

Meador:2006:AGR

[1525] J. P. Meador, F. C. Sommers, G. M. Ylitalo, and C. A. Sloan. Altered growth and related physiological responses in juvenile Chinook salmon (Oncorhynchus tshawytscha) from dietary exposure to polycyclic aromatic hydrocarbons (PAHs). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 63(10):2364–2376, October 2006. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f06-127.

Rago:2006:SME

[1526] Paul J. Rago, James R. Weinberg, and Christopher Weidman. A spatial model to estimate gear efficiency and animal density from depletion experiments. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 63(10):2377–2388, October 2006. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f06-121.

Buffington:2006:EBS

[1527] John M. Buffington, David R. Montgomery, and Harvey M. Greenberg. Erratum: Basin-scale availability of salmonid spawning gravel as influenced by channel type and hydraulic roughness in mountain catchments. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 63(10):2389, October 2006. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f06-143.

Rogers:2006:SGE

[1528] Mark W. Rogers, Micheal S. Allen, and Wesley F. Porak. Separating genetic and environmental influences on temporal spawning distributions of largemouth bass (Micropterus salmoides). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 63(11):2391–2399, November 2006. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f06-122.

Kroon:2006:CSA

[1529] Frederieke J. Kroon and Dean H. Ansell. A comparison of species assemblages between drainage systems with and without floodgates: implications for coastal floodplain management. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 63(11):2400–2417, November 2006. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f06-134.

Gutreuter:2006:PDC

[1530] Steve Gutreuter, Jon M. Vallazza, and Brent C. Knights. Persistent disturbance by commercial navigation alters the relative abundance of channel-dwelling fishes in a large river. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 63(11):2418–2433, November 2006. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f06-129.

Bunnell:2006:LTC

[1531] David B. Bunnell, Charles P. Madenjian, and Randall M. Claramunt. Long-term changes of the Lake Michigan fish community following the reduction of exotic alewife (Alosa pseudoharengus). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 63(11):2434–2446, November 2006. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f06-132.

Houser:2006:WCA

[1532] Jeffrey N. Houser. Water color affects the stratification, surface temperature, heat content, and mean epilimnetic irradiance of small lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 63(11):2447–2455, November 2006. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f06-131.

Braccini:2006:HAA

[1533] J. Matías Braccini, Bronwyn M. Gillanders, and Terence I. Walker. Hierarchical approach to the assessment of fishing effects on non-target chondrichthyans: case study of Squalus megalops in southeastern Australia. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 63(11):2456–2466, November 2006. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f06-141.

Wilhelm:2006:LTR

[1534] Susann Wilhelm, Thomas Hintze, David M. Livingstone, and Rita Adrian. Long-term response of daily epilimnetic temperature extrema to climate forcing. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 63(11): 2467–2477, November 2006. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f06-140.

Lepori:2006:FWR

[1535] Fabio Lepori, Dirk Gaul, Daniel Palm, and Björn Malmqvist. Foodweb responses to restoration of channel heterogeneity in boreal streams. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 63(11):2478–2486, November 2006. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f06-142.

Galbraith:2006:ESS

[1536] Ryan V. Galbraith, Erland A. MacIsaac, J. Stevenson Macdonald, and Anthony P. Farrell. The effect of suspended sediment on fertilization success in sockeye (Oncorhynchus nerka) and Coho (Oncorhynchus kisutch) salmon. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 63(11):2487–2494, November 2006. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f06-133.

Guenette:2006:EMS

[1537] Sylvie Guénette, Sheila J. J. Heymans, Villy Christensen, and Andrew W. Trites. Ecosystem models show combined effects of fishing, predation, competition, and ocean productivity on Steller sea lions (Eumetopias jubatus) in Alaska. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 63(11): 2495–2517, November 2006. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f06-136.

Kiffney:2006:TSC

[1538] P. M. Kiffney, C. M. Greene, J. E. Hall, and J. R. Davies. Tributary streams create spatial discontinuities in habitat, biological productivity, and diversity in mainstem rivers. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 63(11):2518–2530, November 2006. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f06-138.

Kristensen:2006:HVL

[1539] Kasper Kristensen, Peter Lewy, and Jan E. Beyer. How to validate a length-based model of single-species fish stock dynamics. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 63(11):2531–2542, November 2006. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f06-135.

Francis:2006:AIP

[1540] Tessa B. Francis, Daniel E. Schindler, and Jonathan W. Moore. Aquatic insects play a minor role in dispersing salmon-derived nutrients into riparian forests in southwestern Alaska. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 63(11):2543–2552, November 2006. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f06-144.

Rice:2006:TCP

[1541] Stephen P. Rice, Robert I. Ferguson, and Trevor B. Hoey. Tributary control of physical heterogeneity and biological diversity at river confluences. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 63(11):2553–2566, November 2006. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f06-145.

Moir:2006:DHI

[1542] Hamish J. Moir, Chris N. Gibbins, Chris Soulsby, and John H. Webb. Discharge and hydraulic interactions in contrasting channel morphologies and their influence on site utilization by spawning Atlantic salmon (Salmo salar). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 63(11):2567– 2585, November 2006. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f06-137.

Kempf:2006:DCN

[1543] Alexander Kempf, Jens Floeter, and Axel Temming. Decadal changes in the North Sea food web between 1981 and 1991 — implications for fish stock assessment. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 63(11):

2586–2602, November 2006. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f06-147.

Anonymous:2006:ACN

[1544] Anonymous. Announcement: Call for nominations for the Timothy R. Parsons Medal in ocean sciences / communiqué : Appel de candidatures pour la médaille Timothy R. Parsons (sciences de la mer). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 63(12), December 2006. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f06-903.

Laurel:2006:BCH

[1545] Benjamin J. Laurel and Ian R. Bradbury. “Big” concerns with high latitude marine protected areas (MPAs): trends in connectivity and MPA size. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 63(12):2603–2607, December 2006. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f06-151.

VanderZanden:2006:EBP

[1546] M. Jake Vander Zanden, Sudeep Chandra, Sang-Kyu Park, Yvonne Vadeboncoeur, and Charles R. Goldman. Efficiencies of benthic and pelagic trophic pathways in a subalpine lake. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 63(12):2608–2620, December 2006. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f06-148.

Busse:2006:RAN

[1547] Lilian B. Busse, Juliet C. Simpson, and Scott D. Cooper. Relationships among nutrients, algae, and land use in urbanized southern California streams. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 63(12):2621–2638, December 2006. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f06-146.

Sastri:2006:FVI

[1548] Akash R. Sastri and John F. Dower. Field validation of an instantaneous estimate of in situ development and growth for marine copepod communities. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 63(12):2639–2647, December 2006. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f06-149.

Agostini:2006:RBP

[1549] Vera N. Agostini, Robert C. Francis, Anne B. Hollowed, Stephen D. Pierce, Chris Wilson, and Albert N. Hendrix. The relationship between Pacific hake (Merluccius productus) distribution and poleward subsurface flow in the California Current system. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 63(12):2648–2659, December 2006. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f06-139.

Pampoulie:2006:GSA

[1550] Christophe Pampoulie, Daniel E. Ruzzante, Valérie Chosson, Thóra Dögg Jörundsdóttir, Lorna Taylor, Vilhjálmur Thorsteinsson, Anna Kristín Daníelsdóttir, and Gudrún Marteinsdóttir. The genetic structure of Atlantic cod (Gadus morhua) around Iceland: insight from microsatellites, the Pan I locus, and tagging experiments. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 63(12):2660–2674, December 2006. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f06-150.

Utz:2006:TSV

[1551] Ryan M. Utz and Kyle J. Hartman. Temporal and spatial variation in the energy intake of a brook trout (Salvelinus fontinalis) population in an Appalachian watershed. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 63(12): 2675–2686, December 2006. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f06-152.

Gamble:2006:UZB

[1552] Allison E. Gamble, Russ Lloyd, John Aiken, Ora E. Johannsson, and Edward L. Mills. Using zooplankton biomass size spectra to assess ecological change in a well-studied freshwater lake ecosystem: Oneida Lake, New York. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 63(12):2687–2699, December 2006. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f06-153.

Tiira:2006:ERG

[1553] Katriina Tiira, Jorma Piironen, and Craig R. Primmer. Evidence for reduced genetic variation in severely deformed juvenile salmonids. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 63(12):2700–2707, December 2006. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f06-154.

Record:2006:PDV

[1554] Nicholas R. Record and Brad de Young. Patterns of diel vertical migration of zooplankton in acoustic Doppler velocity and backscatter data on the Newfoundland shelf. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 63(12): 2708–2721, December 2006. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f06-157.

Holt:2006:MTU

[1555] Carrie A. Holt and Randall M. Peterman. Missing the target: uncertainties in achieving management goals in fisheries on Fraser River, British Columbia, sockeye salmon (Oncorhynchus nerka). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 63(12):2722–2733, December 2006. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f06-155.

Gal:2006:MFZ

[1556] Gideon Gal, Lars G. Rudstam, Edward L. Mills, Jana R. Lantry, Ora E. Johannsson, and Charles H. Greene. Mysid and fish zooplanktivory in Lake Ontario: quantification of direct and indirect effects. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 63(12):2734–2747, December 2006. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f06-156.

Brickman:2006:RAM

[1557] David Brickman. Risk assessment model for dispersion of ballast water organisms in shelf seas. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 63(12): 2748–2759, December 2006. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f06-158.

Leu:2006:FAD

[1558] Eva Leu, Stig Falk-Petersen, Slawomir Kwaśniewski, Angela Wulff, Kåre Edvardsen, and Dag O. Hessen. Fatty acid dynamics during the spring bloom in a High Arctic fjord: importance of abiotic factors versus community changes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 63(12):2760– 2779, December 2006. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f06-159.

Bertolo:2006:SEC

[1559] Andrea Bertolo and Pierre Magnan. Spatial and environmental correlates of fish community structure in Canadian Shield lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 63(12):2780–2792, December 2006. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f06-161.

Hard:2006:GEE

[1560] Jeffrey J. Hard, Diane G. Elliott, Ronald J. Pascho, Dorothy M. Chase, Linda K. Park, James R. Winton, and Donald E. Campton. Genetic effects of ELISA-based segregation for control of bacterial kidney disease in Chinook salmon (Oncorhynchus tshawytscha). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 63(12):2793–2808, December 2006. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f06-163.

Wells:2006:PES

[1561] Alan Wells, Christal E. Grierson, Monique MacKenzie, Iain J. Russon, Helena Reinardy, Claire Middlemiss, Pål A. Bjørn, Bengt Finstad, Sjoerd E. Wendelaar Bonga, Christopher D. Todd, and Neil Hazon. Physiological effects of simultaneous, abrupt seawater entry and sea lice (Lepeophtheirus salmonis) infestation of wild, sea-run brown trout (Salmo trutta) smolts. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 63(12):2809–2821, December 2006. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f06-160.

Bradbury:2006:LTV

[1562] I. R. Bradbury, K. Gardiner, P. V. R. Snelgrove, S. E. Campana, P. Bentzen, and L. Guan. Larval transport, vertical distribution, and localized recruitment in anadromous rainbow smelt (Osmerus mordax). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 63(12):2822–2836, December 2006. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f06-164.

Anonymous:2006:AII

[1563] Anonymous. Author index / index des auteurs: Volume 63, 2006. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 63(12), December 2006. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f06-906.

Anonymous:2006:CSV

[1564] Anonymous. Contents / sommaire: Volume 63, 2006. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 63(12), December 2006. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f06-907.

Gibbins:2007:SML

[1565] Chris Gibbins, Damià Vericat, Ramon J. Batalla, and Carlos M. Gomez. Shaking and moving: low rates of sediment transport trigger mass drift of stream invertebrates. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 64 (1):1–5, January 2007. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f06-181.

Campana:2007:TSM

[1566] Steven E. Campana, Alexandra Valentin, Jean-Marie Sévigny, and Don Power. Tracking seasonal migrations of redfish (Sebastes spp.) in and around the Gulf of St. Lawrence using otolith elemental fingerprints. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 64(1):6–18, January 2007. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f06-162.

Cassoff:2007:CBG

[1567] Rachel M. Cassoff, Steven E. Campana, and Sigmund Myklevoll. Changes in baseline growth and maturation parameters of Northwest Atlantic porbeagle, Lamna nasus, following heavy exploitation. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 64(1):19–29, January 2007. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f06-167.

Bucklin:2007:AGD

[1568] Katherine A. Bucklin, Michael A. Banks, and Dennis Hedgecock. Assessing genetic diversity of protected Coho salmon (Oncorhynchus kisutch) populations in California. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 64 (1):30–42, January 2007. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f06-171.

Loewen:2007:EIS

[1569] Mark R. Loewen, Josef Daniel Ackerman, and Paul F. Hamblin. Environmental implications of stratification and turbulent mixing in a shallow lake basin. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 64(1):43–57, January 2007. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/f06-
165.

Pemberton:2007:CPP

[1570] Katharine L. Pemberton, Ralph E. H. Smith, Greg M. Silsbe, Todd Howell, and Susan B. Watson. Controls on phytoplankton physiology in Lake Ontario during the late summer: evidence from new fluorescence methods. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 64(1):58–73, January 2007. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f06-166.

Matthews:2007:DTV

[1571] Blake Matthews and Asit Mazumder. Distinguishing trophic variation from seasonal and size-based isotopic (δ ^{15}N) variation of zooplankton. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 64(1):74–83, January 2007. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f06-168.

Davie:2007:RSAa

[1572] Andrew Davie, Mark J. R. Porter, Niall R. Bromage, and Herve Migaud. The role of seasonally altering photoperiod in regulating physiology in Atlantic cod (Gadus morhua). Part I. Sexual maturation. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 64(1):84–97, January 2007. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f06-169.

Davie:2007:RSAb

[1573] Andrew Davie, Mark J. R. Porter, Niall R. Bromage, and Herve Migaud. The role of seasonally altering photoperiod in regulating physiology in Atlantic cod (Gadus morhua). Part II. Somatic growth. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 64(1):98–112, January 2007. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f06-170.

Johnston:2007:DRO

[1574] Fiona D. Johnston, John R. Post, Craig J. Mushens, Jim D. Stelfox, Andrew J. Paul, and Brian Lajeunesse. The demography of recovery of an overexploited bull trout, Salvelinus confluentus, population. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 64(1):113–126, January 2007. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f06-172.

Seiler:2007:MSS

[1575] Steven M. Seiler and Ernest R. Keeley. Morphological and swimming stamina differences between Yellowstone cutthroat trout (Oncorhynchus clarkii bouvieri), rainbow trout (Oncorhynchus mykiss), and their hybrids. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 64(1):127–135, January 2007. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f06-175.

Kristiansen:2007:MFG

[1576] Trond Kristiansen, Øyvind Fiksen, and Arild Folkvord. Modelling feeding, growth, and habitat selection in larval Atlantic cod (Gadus morhua): observations and model predictions in a macrocosm environment. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 64(1):136–151, January 2007. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f06-176.

Fablet:2007:UBR

[1577] Ronan Fablet, Françoise Daverat, and Hélène De Pontual. Unsupervised Bayesian reconstruction of individual life histories from otolith signatures: case study of Sr : Ca transects of European eel (Anguilla anguilla) otoliths. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 64(1):152–165, January 2007. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f06-173.

Bez:2007:ADC

[1578] Nicolas Bez, David Reid, Suzanna Neville, Yves Vérin, Vidar Hjellvik, and Hans D. Gerritsen. Acoustic data collected during and between bottom trawl stations: consistency and common trends. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 64(1):166–180, January 2007. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f06-174.

Bertolo:2007:LIV

[1579] Andrea Bertolo and Pierre Magnan. Logging-induced variations in dissolved organic carbon affect yellow perch (Perca flavescens) recruitment in Canadian Shield lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 64 (2):181–186, February 2007. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f07-004.

Deriso:2007:VEI

[1580] Richard B. Deriso, Mark N. Maunder, and John R. Skalski. Variance estimation in integrated assessment models and its importance for hypothesis testing. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 64(2):187–197, February 2007. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f06-178.

Daly-Engel:2007:FMP

[1581] Toby S. Daly-Engel, R. Dean Grubbs, Brian W. Bowen, and Robert J. Toonen. Frequency of multiple paternity in an unexploited tropical population of sandbar sharks (Carcharhinus plumbeus). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 64(2):198–204, February 2007. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f07-005.

Cope:2007:AAE

[1582] Jason M. Cope and André E. Punt. Admitting ageing error when fitting growth curves: an example using the von Bertalanffy growth function with random effects. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 64(2): 205–218, February 2007. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f06-179.

Rowe:2007:NMB

[1583] Sherrylynn Rowe, Jeffrey A. Hutchings, and Jon Egil Skjæraasen. Nonrandom mating in a broadcast spawner: mate size influences reproductive success in Atlantic cod (Gadus morhua). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 64(2):219–226, February 2007. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f06-182.

Bowes:2007:PBR

[1584] Michael J. Bowes, Jim T. Smith, John Hilton, Michael M. Sturt, and Patrick D. Armitage. Periphyton biomass response to changing phosphorus concentrations in a nutrient-impacted river: a new methodology for phosphorus target setting. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 64 (2):227–238, February 2007. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f06-180.

Imsland:2007:IBT

[1585] Albert Kjartansson Imsland, Atle Foss, Thomas Alvseike, Arild Folkvord, Sigurd Olav Stefansson, and Thor Magne Jonassen. Interaction between temperature and photoperiod on growth and feeding of Atlantic cod (Gadus morhua): possible secondary effects. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 64(2):239–248, February 2007. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f07-006.

Griffiths:2007:NMG

[1586] David Griffiths and Chris Harrod. Natural mortality, growth parameters, and environmental temperature in fishes revisited. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 64(2):249–255, February 2007. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f07-002.

Beckman:2007:SDL

[1587] Brian R. Beckman, Brad Gadberry, Paul Parkins, Kathleen A. Cooper, and Kristen D. Arkush. State-dependent life history plasticity in Sacramento River winter-run Chinook salmon (Oncorhynchus tshawytscha): interactions among photoperiod and growth modulate smolting and early male maturation. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 64(2):256–271, February 2007. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f07-001.

Costa:2007:BIT

[1588] Filipe O. Costa, Jeremy R. deWaard, James Boutillier, Sujeevan Ratnasingham, Robert T. Dooh, Mehrdad Hajibabaei, and Paul D. N. Hebert. Biological identifications through DNA barcodes: the case of the crustacea. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 64(2):272–295, February 2007. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f07-008.

Waite:2007:TDP

[1589] Jason N. Waite, Wendy J. Schrader, Jo-Ann E. Mellish, and Markus Horning. Three-dimensional photogrammetry as a tool for estimating morphometrics and body mass of Steller sea lions (Eumetopias jubatus). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 64(2):296–303, February 2007. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f07-014.

Poos:2007:EEA

[1590] Jan-Jaap Poos and Adriaan D. Rijnsdorp. An “experiment” on effort allocation of fishing vessels: the role of interference competition and area specialization. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 64(2): 304–313, February 2007. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f06-177.

Lester:2007:DDP

[1591] Sarah E. Lester, Elizabeth D. Tobin, and Michael D. Behrens. Disease dynamics and the potential role of thermal stress in the sea urchin, Strongylocentrotus purpuratus. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 64(2): 314–323, February 2007. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f07-010.

Giannico:2007:JCS

[1592] Guillermo R. Giannico and Scott G. Hinch. Juvenile Coho salmon (Oncorhynchus kisutch) responses to salmon carcasses and in-stream wood manipulations during winter and spring. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 64(2):324–335, February 2007. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f07-011.

Adams:2007:MDW

[1593] C. E. Adams, J. F. Turnbull, A. Bell, J. E. Bron, and F. A. Huntingford. Multiple determinants of welfare in farmed fish: stocking density, disturbance, and aggression in Atlantic salmon (Salmo salar). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 64(2):336–344, February 2007. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f07-018.

Evans:2007:EHS

[1594] David O. Evans. Effects of hypoxia on scope-for-activity and power capacity of lake trout (Salvelinus namaycush). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 64(2):345–361, February 2007. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f07-007.

Fassler:2007:MBP

[1595] Sascha M. M. Fässler, Rita Santos, Norma García-Núñez, and Paul G. Fernandes. Multifrequency backscattering properties of Atlantic herring (Clupea harengus) and Norway pout (Trisopterus esmarkii). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 64(2):362–374, February 2007. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f07-013.

Neuheimer:2007:GDD

[1596] Anna B. Neuheimer and Christopher T. Taggart. The growing degreeday and fish size-at-age: the overlooked metric. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 64(2):375–385, February 2007. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f07-003.

Jech:2007:EMM

[1597] J. Michael Jech and William L. Michaels. Erratum: a multifrequency method to classify and evaluate fisheries acoustics data. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 64(2):386, February 2007. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f07-009.

Grant:2007:RSP

[1598] Jon Grant, Gary Bugden, Edward Horne, Marie-Claude Archambault, and Michel Carreau. Remote sensing of particle depletion by coastal suspension-feeders. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 64(3):387– 390, March 2007. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f07-021.

Brooks:2007:TCL

[1599] Marjorie L. Brooks, Joseph S. Meyer, and Connie J. Boese. Toxicity of copper to larval Pimephales promelas in the presence of photodegraded natural dissolved organic matter. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 64(3):391–401, March 2007. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f07-012.

Thomas:2007:IEP

[1600] Gregor Thomas and Reiner Eckmann. The influence of eutrophication and population biomass on common whitefish (Coregonus lavaretus) growth — the Lake Constance example revisited. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 64(3):402–410, March 2007. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f07-019.

Dorval:2007:ROC

[1601] Emmanis Dorval, Cynthia M. Jones, Robyn Hannigan, and Jacques van Montfrans. Relating otolith chemistry to surface water chemistry in a coastal plain estuary. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 64 (3):411–424, March 2007. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f07-015.

Goedkoop:2007:IBF

[1602] Willem Goedkoop, Marnie Demandt, and Gunnel Ahlgren. Interactions between food quantity and quality (long-chain polyunsaturated fatty acid concentrations) effects on growth and development of Chironomus riparius. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 64(3):425–436, March 2007. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f07-016.

Strange:2007:GDC

[1603] Rex Meade Strange and Carol A. Stepien. Genetic divergence and connectivity among river and reef spawning groups of walleye (Sander vitreus vitreus) in Lake Erie. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 64 (3):437–448, March 2007. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f07-022.

Hurst:2007:TEB

[1604] Thomas P. Hurst. Thermal effects on behavior of juvenile walleye pollock (Theragra chalcogramma): implications for energetics and food web models. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 64(3):449–457, March 2007. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f07-025.

Cai:2007:BMP

[1605] Yan Cai, Jay R. Rooker, Gary A. Gill, and Jason P. Turner. Bioaccumulation of mercury in pelagic fishes from the northern Gulf of Mexico. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 64(3):458–469, March 2007. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f07-017.

Helser:2007:BHM

[1606] Thomas E. Helser, Ian J. Stewart, and Han-Lin Lai. A Bayesian hierarchical meta-analysis of growth for the genus Sebastes in the eastern Pacific Ocean. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 64(3):470–485, March 2007. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f07-024.

Breau:2007:IVA

[1607] Cindy Breau, Laura K. Weir, and James W. A. Grant. Individual variability in activity patterns of juvenile Atlantic salmon (Salmo salar) in Catamaran Brook, New Brunswick. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 64(3):486–494, March 2007. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f07-026.

Herwig:2007:STP

[1608] Brian R. Herwig, David H. Wahl, John M. Dettmers, and Daniel A. Soluk. Spatial and temporal patterns in the food web structure of a large floodplain river assessed using stable isotopes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 64(3):495–508, March 2007. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f07-023.

Palm:2007:ISH

[1609] Daniel Palm, Eva Brännäs, Fabio Lepori, Kjell Nilsson, and Stefan Stridsman. The influence of spawning habitat restoration on juvenile brown trout (Salmo trutta) density. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 64(3):509–515, March 2007. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f07-027.

Finlay:2007:REV

[1610] Kerri Finlay, Beatrix E. Beisner, Alain Patoine, and Bernadette Pinel-Alloul. Regional ecosystem variability drives the relative importance of bottom-up and top-down factors for zooplankton size spectra. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 64(3):516–529, March 2007. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f07-028.

Drake:2007:RSI

[1611] John M. Drake and David M. Lodge. Rate of species introductions in the Great Lakes via ships’ ballast water and sediments. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 64(3):530–538, March 2007. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f07-029.

Ibaibarriaga:2007:CSC

[1612] Leire Ibaibarriaga, Miguel Bernal, Lorenzo Motos, Andrés Uriarte, David L. Borchers, Mike E. Lonergan, and Simon N. Wood. Characterization of stage-classified biological processes using multinomial models: a case study of anchovy (Engraulis encrasicolus) eggs in the Bay of Biscay. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 64(3):539–553, March 2007. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f07-030.

Tian:2007:ELF

[1613] Tian Tian, Øyvind Fiksen, and Arild Folkvord. Estimating larval fish growth under size-dependent mortality: a numerical analysis of bias. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 64(3):554–562, March 2007. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f07-031.

Mitchell:2007:RUM

[1614] Sean C. Mitchell and Richard A. Cunjak. Relationship of upstream migrating adult Atlantic salmon (Salmo salar) and stream discharge within Catamaran Brook, New Brunswick. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 64(3):563–573, March 2007. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f07-032.

Quinn:2007:DCP

[1615] T. P. Quinn, D. M. Eggers, J. H. Clark, and H. B. Rich, Jr. Density, climate, and the processes of prespawning mortality and egg retention in Pacific salmon (Oncorhynchus spp.). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 64(3):574–582, March 2007. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f07-035.

Folkvord:2007:ECS

[1616] A. Folkvord. Erratum: Comparison of size-at-age of larval Atlantic cod (Gadus morhua) from different populations based on size- and temperature-dependent growth models. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 64(3):583–585, March 2007. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f07-045.

Dancik:2007:AC

[1617] Bruce P. Dancik. Announcement / Communiqué. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 64(4), April 2007. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f07-901.

Kennedy:2007:CFR

[1618] J. Kennedy, P. R. Witthames, and R. D. M. Nash. The concept of fecundity regulation in plaice (Pleuronectes platessa) tested on three Irish Sea spawning populations. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 64 (4):587–601, April 2007. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f07-034.

Eveson:2007:CAI

[1619] J. Paige Eveson, Tom Polacheck, and Geoff M. Laslett. Consequences of assuming an incorrect error structure in von Bertalanffy growth models: a simulation study. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 64 (4):602–617, April 2007. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f07-036.

Andersen:2007:UPF

[1620] K. H. Andersen, A. Nielsen, U. H. Thygesen, H.-H. Hinrichsen, and S. Neuenfeldt. Using the particle filter to geolocate Atlantic cod (Gadus morhua) in the Baltic Sea, with special emphasis on determining uncertainty. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 64(4):618–627, April 2007. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f07-037.

Essington:2007:EST

[1621] Timothy E. Essington. Evaluating the sensitivity of a trophic massbalance model (Ecopath) to imprecise data inputs. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 64(4):628–637, April 2007. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f07-042.

Gurney:2007:MLA

[1622] William S. C. Gurney, Graham Tyldesley, Simon N. Wood, Philip J. Bacon, Michael R. Heath, Alan Youngson, and Anton Ibbotson. Modelling length-at-age variability under irreversible growth. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 64(4):638–653, April 2007. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f07-039.

Spens:2007:NBT

[1623] Johan Spens, Anders Alanärä, and Lars-Ove Eriksson. Nonnative brook trout (Salvelinus fontinalis) and the demise of native brown trout (Salmo trutta) in northern boreal lakes: stealthy, long-term patterns? Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 64(4):654–664, April 2007. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f07-040.

Laakkonen:2007:BTP

[1624] Mika V. M. Laakkonen and Heikki Hirvonen. Is boldness towards predators related to growth rate in naïve captive-reared Arctic char (Salvelinus alpinus)? Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 64(4):665–671, April 2007. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f07-041.

Webster:2007:ESL

[1625] Sandra J. Webster, Lawrence M. Dill, and Kevin Butterworth. The effect of sea lice infestation on the salinity preference and energetic expenditure of juvenile pink salmon (Oncorhynchus gorbuscha). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 64(4):672–680, April 2007. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f07-043.

Searcy:2007:GRI

[1626] Steven P. Searcy, David B. Eggleston, and Jonathan A. Hare. Is growth a reliable indicator of habitat quality and essential fish habitat for a juvenile estuarine fish? Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 64 (4):681–691, April 2007. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f07-038.

Rombough:2007:OCF

[1627] Peter J. Rombough. Oxygen as a constraining factor in egg size evolution in salmonids. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 64 (4):692–699, April 2007. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f07-047.

Wiegand:2007:CSO

[1628] M. D. Wiegand, T. A. Johnston, W. C. Leggett, K. E. Watchorn, A. J. Ballevona, L. R. Porteous, and J. M. Casselman. Contrasting strategies of ova lipid provisioning in relation to maternal characteristics in three walleye (Sander vitreus) populations. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 64(4):700–712, April 2007. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f07-033.

Ciannelli:2007:PGP

[1629] Lorenzo Ciannelli, Kevin M. Bailey, Kung-Sik Chan, and Nils Chr. Stenseth. Phenological and geographical patterns of walleye pollock (Theragra chalcogramma) spawning in the western Gulf of Alaska. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 64(4):713–722, April 2007. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f07-049.

Feyrer:2007:MTT

[1630] Frederick Feyrer, Matthew L. Nobriga, and Ted R. Sommer. Multidecadal trends for three declining fish species: habitat patterns and mechanisms in the San Francisco Estuary, California, USA. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 64(4):723–734, April 2007. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f07-048.

Sato:2007:DMO

[1631] Taku Sato, Masakazu Ashidate, Tadao Jinbo, and Seiji Goshima. Does male-only fishing influence reproductive success of the female spiny king crab, Paralithodes brevipes? Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 64 (4):735–742, April 2007. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f07-044.

Koseki:2007:LSF

[1632] Yusuke Koseki and Ian A. Fleming. Large-scale frequency dynamics of alternative male phenotypes in natural populations of Coho salmon (Oncorhynchus kisutch): patterns, processes, and implications. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 64(4):743–753, April 2007. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f07-046.

Rosenfeld:2007:HGP

[1633] Jordan S. Rosenfeld, John Post, Geneva Robins, and Todd Hatfield. Hydraulic geometry as a physical template for the river continuum: application to optimal flows and longitudinal trends in salmonid habitat. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 64(5):755–767, May 2007. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f07-020.

Takasuka:2007:OGT

[1634] Akinori Takasuka, Yoshioki Oozeki, and Ichiro Aoki. Optimal growth temperature hypothesis: Why do anchovy flourish and sardine collapse or vice versa under the same ocean regime? Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 64(5):768–776, May 2007. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f07-052.

Deschenes:2007:HAR

[1635] Julie Deschênes and Marco A. Rodríguez. Hierarchical analysis of relationships between brook trout (Salvelinus fontinalis) density and stream habitat features. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 64(5):777–785, May 2007. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/f07-
053.

Roussel:2007:CEB

[1636] Jean-Marc Roussel. Carry-over effects in brown trout (Salmo trutta): hypoxia on embryos impairs predator avoidance by alevins in experimental channels. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 64(5):786–792, May 2007. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f07-055.

Thomsen:2007:OST

[1637] Dennis S. Thomsen, Anders Koed, Christian Nielsen, and Steffen S. Madsen. Overwintering of sea trout (Salmo trutta) in freshwater: escaping salt and low temperature or an alternate life strategy? Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 64(5):793–802, May 2007. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f07-059.

Pearsons:2007:EHD

[1638] Todd N. Pearsons, Anthony L. Fritts, and Jennifer L. Scott. The effects of hatchery domestication on competitive dominance of juvenile spring Chinook salmon (Oncorhynchus tshawytscha). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 64(5):803–812, May 2007. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f07-058.

Fritts:2007:EDR

[1639] Anthony L. Fritts, Jennifer L. Scott, and Todd N. Pearsons. The effects of domestication on the relative vulnerability of hatchery and wild origin spring Chinook salmon (Oncorhynchus tshawytscha) to predation. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 64(5):813–818, May 2007. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f07-057.

Wright:2007:CGT

[1640] Hazel A. Wright, Robert J. Wootton, and Iain Barber. Compensatory growth in threespine sticklebacks (Gasterosteus aculeatus) inhibited by experimental Schistocephalus infections. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 64(5):819–826, May 2007. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f07-056.

Fromentin:2007:CVC

[1641] Jean-Marc Fromentin and Laurence T. Kell. Consequences of variations in carrying capacity or migration for the perception of Atlantic bluefin tuna (Thunnus thynnus) population dynamics. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 64(5):827–836, May 2007. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f07-050.

Kell:2007:ERM

[1642] Laurence T. Kell and Jean-Marc Fromentin. Evaluation of the robustness of maximum sustainable yield based management strategies to variations in carrying capacity or migration pattern of Atlantic bluefin tuna (Thunnus thynnus). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 64 (5):837–847, May 2007. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f07-051.

Lamarre:2007:DME

[1643] Simon Gaston Lamarre, Nathalie Rose Le François, Hélène Lemieux, Inger-Britt Falk-Petersen, and Pierre Ulrich Blier. The digestive and metabolic enzyme activity profiles of a nonmetamorphic marine fish species: effects of feed type and feeding level. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 64(6):849–856, June 2007. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f07-054.

Knutsen:2007:PGS

[1644] Halvor Knutsen, Per Erik Jorde, Ole Thomas Albert, A. Rus Hoelzel, and Nils Chr. Stenseth. Population genetic structure in the North Atlantic Greenland halibut (Reinhardtius hippoglossoides): influenced by oceanic current systems? Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 64 (6):857–866, June 2007. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f07-070.

Behrens:2007:TDE

[1645] Michael D. Behrens and Kevin D. Lafferty. Temperature and diet effects on omnivorous fish performance: implications for the latitudinal diversity gradient in herbivorous fishes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 64(6):867–873, June 2007. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f07-063.

Mehner:2007:EUH

[1646] Thomas Mehner, Peter Kasprzak, and Franz Hölker. Exploring ultimate hypotheses to predict diel vertical migrations in coregonid fish. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 64(6):874–886, June 2007. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f07-067.

Derry:2007:ZCR

[1647] Alison M. Derry and Shelley E. Arnott. Zooplankton community response to experimental acidification in boreal shield lakes with different ecological histories. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 64 (6):887–898, June 2007. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f07-061.

Ward:2007:SSM

[1648] Eric J. Ward, Ray Hilborn, Rod G. Towell, and Leah Gerber. A state– space mixture approach for estimating catastrophic events in time series data. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 64(6):899–910, June 2007. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f07-060.

Mueter:2007:RSN

[1649] Franz J. Mueter, Jennifer L. Boldt, Bernard A. Megrey, and Randall M. Peterman. Recruitment and survival of Northeast Pacific Ocean fish stocks: temporal trends, covariation, and regime shifts. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 64(6):911–927, June 2007. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f07-069.

Marasco:2007:EBF

[1650] Richard J. Marasco, Daniel Goodman, Churchill B. Grimes, Peter W. Lawson, André E. Punt, and Terrance J. Quinn II. Ecosystem-based fisheries management: some practical suggestions. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 64(6):928–939, June 2007. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f07-062.

Torgersen:2007:APV

[1651] Thomas Torgersen. Aggregated predators and vulnerability-independent mortality of prey. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 64(7):941–955, July 2007. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/f07-
066.

Tewfik:2007:SSF

[1652] Alexander Tewfik, Joseph B. Rasmussen, and Kevin S. McCann. Simplification of seagrass food webs across a gradient of nutrient enrichment. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 64(7):956–967, July 2007. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f07-071.

Wagner:2007:RTF

[1653] Tyler Wagner, James R. Bence, Mary T. Bremigan, Daniel B. Hayes, and Michael J. Wilberg. Regional trends in fish mean length at age: components of variance and the statistical power to detect trends. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 64(7):968–978, July 2007. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f07-068.

Caudill:2007:SDP

[1654] Christopher C. Caudill, William R. Daigle, Matthew L. Keefer, Charles T. Boggs, Michael A. Jepson, Brian J. Burke, Richard W. Zabel, Theodore C. Bjornn, and Christopher A. Peery. Slow dam passage in adult Columbia River salmonids associated with unsuccessful migration: delayed negative effects of passage obstacles or condition-dependent mortality? Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 64(7):979–995, July 2007. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f07-065.

Fonseca:2007:DSM

[1655] Duane B. Fonseca and Matt R. J. Sheehy. Does size matter? A cautionary experiment on overoptimism in length-based bioresource assessment. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 64(7):996–1008, July 2007. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f07-076.

Walters:2007:EMP

[1656] Carl J. Walters, Ray Hilborn, and Richard Parrish. An equilibrium model for predicting the efficacy of marine protected areas in coastal environments. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 64(7):1009–1018, July 2007. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f07-072.

Griffiths:2007:EAP

[1657] Shane P. Griffiths, Petra M. Kuhnert, William N. Venables, and Stephen J. M. Blaber. Estimating abundance of pelagic fishes using gillnet catch data in data-limited fisheries: a Bayesian approach. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 64(7):1019–1033, July 2007. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f07-073.

Ibarz:2007:MRT

[1658] Antoni Ibarz, Josefina Blasco, Mònica Sala-Rabanal, Ángeles Gallardo, Anna Redondo, and Jaume Fernández-Borràs. Metabolic rate and tissue reserves in gilthead sea bream (Sparus aurata) under thermal fluctuations and fasting and their capacity for recovery. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 64(7):1034–1042, July 2007. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f07-079.

Ginn:2007:APC

[1659] Brian K. Ginn, Brian F. Cumming, and John P. Smol. Assessing pH changes since pre-industrial times in 51 low-alkalinity lakes in Nova Scotia, Canada. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 64(8):1043– 1054, August 2007. CODEN CJFSDX. ISSN 0706-652X (print), 1205-

7533 (electronic). URL https://cdnsciencepub.com/doi/full/10.
1139/f07-078.

Nielsen:2007:SSM

[1660] Anders Nielsen and John R. Sibert. State–space model for light-based tracking of marine animals. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 64 (8):1055–1068, August 2007. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f07-064.

Deschenes:2007:CDR

[1661] Julie Deschênes, Marco A. Rodríguez, and Pierre Bérubé. Contextdependent responses of juvenile Atlantic salmon (Salmo salar) to forestry activities at multiple spatial scales within a river basin. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 64(8):1069–1079, August 2007. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f07-077.

Roth:2007:CSO

[1662] Jennifer E. Roth, Kyra L. Mills, and William J. Sydeman. Chinook salmon (Oncorhynchus tshawytscha) — seabird covariation off central California and possible forecasting applications. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 64(8):1080–1090, August 2007. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f07-081.

Ibanez:2007:GMA

[1663] Ana L. Ibañez, Ian G. Cowx, and Paul O’Higgins. Geometric morphometric analysis of fish scales for identifying genera, species, and local populations within the Mugilidae. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 64 (8):1091–1100, August 2007. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f07-075.

Korman:2007:IMS

[1664] Josh Korman, Caroline C. Melville, and Paul S. Higgins. Integrating multiple sources of data on migratory timing and catchability to estimate escapement for steelhead trout (Oncorhynchus mykiss). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 64(8):1101–1115, August 2007. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f07-083.

Dauwalter:2007:GSH

[1665] Daniel C. Dauwalter, Dale K. Splinter, William L. Fisher, and Richard A. Marston. Geomorphology and stream habitat relationships with smallmouth bass (Micropterus dolomieu) abundance at multiple spatial scales in eastern Oklahoma. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 64(8): 1116–1129, August 2007. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f07-085.

Aanes:2007:EPF

[1666] Sondre Aanes, Steinar Engen, Bernt-Erik Sæther, and Ronny Aanes. Estimation of the parameters of fish stock dynamics from catch-at-age data and indices of abundance: can natural and fishing mortality be separated? Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 64(8):1130–1142, August 2007. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/f07-
074.

Anderson:2007:MAC

[1667] Joseph H. Anderson and Thomas P. Quinn. Movements of adult Coho salmon (Oncorhynchus kisutch) during colonization of newly accessible habitat. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 64(8):1143–1154, August 2007. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/f07-
087.

Cope:2007:EAA

[1668] Jason M. Cope and André E. Punt. Erratum: Admitting ageing error when fitting growth curves: an example using the von Bertalanffy growth function with random effects. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 64 (8):1155, August 2007. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f07-095.

McGarvey:2007:MFN

[1669] Richard McGarvey, John E. Feenstra, and Qifeng Ye. Modeling fish numbers dynamically by age and length: partitioning cohorts into “slices”. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 64(9):1157–1173, September 2007. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f07-080.

Wilkie:2007:LTN

[1670] Michael P. Wilkie, John A. Holmes, and John H. Youson. The lampricide 3-trifluoromethyl-4-nitrophenol (TFM) interferes with intermediary metabolism and glucose homeostasis, but not with ion balance, in larval sea lamprey (Petromyzon marinus). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 64(9):1174–1182, September 2007. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f07-082.

Kallio-Nyberg:2007:EMC

[1671] Irma Kallio-Nyberg, Irma Saloniemi, Eero Jutila, and Ari Saura. Effects of marine conditions, fishing, and smolt traits on the survival of tagged, hatchery-reared sea trout (Salmo trutta trutta) in the Baltic Sea. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 64(9):1183–1198, September 2007. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f07-084.

Pavey:2007:RED

[1672] Scott A. Pavey, Troy R. Hamon, and Jennifer L. Nielsen. Revisiting evolutionary dead ends in sockeye salmon (Oncorhynchus nerka) life history. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 64(9):1199–1208, September 2007. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f07-091.

Spies:2007:EEN

[1673] Ingrid B. Spies, Eric C. Anderson, Kerry Naish, and Paul Bentzen. Evidence for the existence of a native population of sockeye salmon (Oncorhynchus nerka) and subsequent introgression with introduced populations in a Pacific Northwest watershed. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 64(9):1209–1221, September 2007. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f07-086.

Paterson:2007:LTP

[1674] Gord Paterson, Ken G. Drouillard, Todd A. Leadley, and G. Doug Haffner. Long-term polychlorinated biphenyl elimination by three size classes of yellow perch (Perca flavescens). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 64(9):1222–1233, September 2007. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f07-088.

Hare:2007:PLA

[1675] Jonathan A. Hare and Harvey J. Walsh. Planktonic linkages among marine protected areas on the south Florida and southeast United States continental shelves. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 64(9):1234– 1247, September 2007. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f07-089.

Li:2007:OHU

[1676] Xinhai Li, Matthew K. Litvak, and John E. Hughes Clarke. Overwintering habitat use of shortnose sturgeon (Acipenser brevirostrum): defining critical habitat using a novel underwater video survey and modeling approach. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 64(9):1248– 1257, September 2007. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f07-093.

Moyer:2007:IFC

[1677] G. R. Moyer, M. S. Blouin, and M. A. Banks. The influence of family-correlated survival on N_{b} /N for progeny from integrated multiand single-generation hatchery stocks of Coho salmon (Oncorhynchus kisutch). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 64(9):1258–1265, September 2007. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f07-099.

Suhett:2007:SCD

[1678] Albert Luiz Suhett, André Megali Amado, Alex Enrich-Prast, Francisco de Assis Esteves, and Vinicius Fortes Farjalla. Seasonal changes of dissolved organic carbon photo-oxidation rates in a tropical humic lagoon: the role of rainfall as a major regulator. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 64(9):1266–1272, September 2007. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f07-103.

Lecerf:2007:SER

[1679] Antoine Lecerf, Denise Patfield, Anatole Boiché, Miira P. Riipinen, Eric Chauvet, and Michael Dobson. Stream ecosystems respond to riparian invasion by Japanese knotweed (Fallopia japonica). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 64(9):1273–1283, September 2007. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f07-092.

Kaiser:2007:EGR

[1680] Michel J. Kaiser, Robert E. Blyth-Skyrme, Paul J. B. Hart, Gareth Edwards-Jones, and David Palmer. Evidence for greater reproductive output per unit area in areas protected from fishing. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 64(9):1284–1289, September 2007. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f07-090.

Zeug:2007:ECF

[1681] Steven C. Zeug and Kirk O. Winemiller. Ecological correlates of fish reproductive activity in floodplain rivers: a life-history-based approach. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 64(10):1291–1301, October 2007. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f07-094.

Heithaus:2007:STV

[1682] Michael R. Heithaus, Derek Burkholder, Robert E. Hueter, Linda I. Heithaus, Harold L. Pratt, Jr., and Jeffrey C. Carrier. Spatial and temporal variation in shark communities of the lower Florida Keys and evidence for historical population declines. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 64(10):1302–1313, October 2007. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f07-098.

Storch:2007:CTE

[1683] Adam J. Storch, Kimberly L. Schulz, Carla E. Cáceres, Peter M. Smyntek, John M. Dettmers, and Mark A. Teece. Consumption of two exotic zooplankton by alewife (Alosa pseudoharengus) and rainbow smelt (Osmerus mordax) in three Laurentian Great Lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 64(10):1314–1328, October 2007. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f07-096.

Vatland:2007:PIS

[1684] Shane Vatland and Phaedra Budy. Predicting the invasion success of an introduced omnivore in a large, heterogeneous reservoir. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 64(10):1329–1345, October 2007. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f07-100.

Kennard:2007:MEF

[1685] Mark J. Kennard, Julian D. Olden, Angela H. Arthington, Bradley J. Pusey, and N. LeRoy Poff. Multiscale effects of flow regime and habitat and their interaction on fish assemblage structure in eastern Australia. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 64(10):1346–1359, October 2007. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f07-108.

Wells:2007:PCP

[1686] Alan Wells, Christal E. Grierson, Laura Marshall, Monique MacKenzie, Iain J. Russon, Helena Reinardy, Rolf Sivertsgård, Pål A. Bjørn, Bengt Finstad, Sjoerd E. Wendelaar Bonga, Christopher D. Todd, and Neil Hazon. Physiological consequences of “premature freshwater return” for wild sea-run brown trout (Salmo trutta) postsmolts infested with sea lice (Lepeophtheirus salmonis). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 64(10):1360–1369, October 2007. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f07-107.

Rasmussen:2007:IVC

[1687] Joseph B. Rasmussen and Veronique Trudeau. Influence of velocity and chlorophyll standing stock on periphyton δ ^{13}C and δ ^{15}N in the Ste. Marguerite River system, Quebec. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 64(10):1370–1381, October 2007. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f07-109.

Yada:2007:DEC

[1688] Takashi Yada, Teruo Azuma, Susumu Hyodo, Tetsuya Hirano, E. Gordon Grau, and Carl B. Schreck. Differential expression of corticosteroid receptor genes in rainbow trout (Oncorhynchus mykiss) immune system in response to acute stress. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 64(10): 1382–1389, October 2007. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f07-110.

Hjellvik:2007:CPB

[1689] Vidar Hjellvik, Dag Tjøstheim, and Olav Rune Godø. Can the precision of bottom trawl indices be increased by using simultaneously collected acoustic data? The Barents Sea experience. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 64(10):1390–1402, October 2007. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f07-101.

Binder:2007:TRV

[1690] Thomas R. Binder and D. Gordon McDonald. Is there a role for vision in the behaviour of sea lampreys (Petromyzon marinus) during their upstream spawning migration? Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 64(10):1403–1412, October 2007. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f07-102.

Lippert:2007:ECP

[1691] Kelly A. Lippert, John M. Gunn, and George E. Morgan. Effects of colonizing predators on yellow perch (Perca flavescens) populations in lakes recovering from acidification and metal stress. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 64(10):1413–1428, October 2007. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f07-097.

Lobon-Cervia:2007:NCS

[1692] Javier Lobón-Cerviá. Numerical changes in stream-resident brown trout (Salmo trutta): uncovering the roles of density-dependent and density-independent factors across space and time. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 64(10):1429–1447, October 2007. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f07-111.

Nielsen:2007:HAP

[1693] Einar Eg Nielsen, Brian R. MacKenzie, Eydfinn Magnussen, and Dorte Meldrup. Historical analysis of Pan I in Atlantic cod (Gadus morhua): temporal stability of allele frequencies in the southeastern part of the species distribution. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 64(10): 1448–1455, October 2007. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f07-104.

Haeseker:2007:ECU

[1694] Steven L. Haeseker, Michael L. Jones, Randall M. Peterman, James R. Bence, Wenjing Dai, and Gavin C. Christie. Explicit consideration of uncertainty in Great Lakes fisheries management: decision analysis of sea lamprey (Petromyzon marinus) control in the St. Marys River. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 64(10):1456–1468, October 2007. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f07-114.

Oxman:2007:EVD

[1695] Dion S. Oxman, Rachel Barnett-Johnson, Michael E. Smith, Allison Coffin, Diane L. Miller, Ron Josephson, and Arthur N. Popper. The effect of vaterite deposition on sound reception, otolith morphology, and inner ear sensory epithelia in hatchery-reared Chinook salmon (Oncorhynchus tshawytscha). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 64(11):1469– 1478, November 2007. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f07-106.

Chasco:2007:RRM

[1696] Brandon Chasco, Ray Hilborn, and André E. Punt. Run reconstruction of mixed-stock salmon fisheries using age-composition data. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 64(11):1479–1490, November 2007. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f07-105.

Richard:2007:ISS

[1697] Marion Richard, Philippe Archambault, Gérard Thouzeau, Chris W. McKindsey, and Gaston Desrosiers. Influence of suspended scallop cages and mussel lines on pelagic and benthic biogeochemical fluxes in Havreaux-Maisons Lagoon, Îles-de-la-Madeleine (Quebec, Canada). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 64(11):1491–1505, November 2007. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f07-116.

Kennedy:2007:RBS

[1698] Benjamen M. Kennedy, William L. Gale, and Kenneth G. Ostrand. Relationship between smolt gill Na^{+}, K^{+} ATPase activity and migration timing to avian predation risk of steelhead trout (Oncorhynchus mykiss) in a large estuary. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 64(11):1506– 1516, November 2007. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f07-117.

Moreno:2007:FBF

[1699] Gala Moreno, Laurent Dagorn, Gorka Sancho, and David Itano. Fish behaviour from fishers’ knowledge: the case study of tropical tuna around drifting fish aggregating devices (DFADs). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 64(11):1517–1528, November 2007. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f07-113.

VanNieuwenhuyse:2007:RSC

[1700] Erwin E. Van Nieuwenhuyse. Response of summer chlorophyll concentration to reduced total phosphorus concentration in the Rhine River (Netherlands) and the Sacramento–San Joaquin Delta (California, USA). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 64(11):1529–1542, November 2007. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f07-121.

Mercado-Silva:2007:IRS

[1701] Norman Mercado-Silva, Greg G. Sass, Brian M. Roth, Stephen Gilbert, and M. Jake Vander Zanden. Impact of rainbow smelt (Osmerus mordax) invasion on walleye (Sander vitreus) recruitment in Wisconsin lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 64(11):1543–1550, November 2007. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f07-112.

Poulard:2007:DSD

[1702] Jean-Charles Poulard and Verena M. Trenkel. Do survey design and wind conditions influence survey indices? Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 64(11):1551–1562, November 2007. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f07-123.

Petrin:2007:LCC

[1703] Zlatko Petrin, Brendan McKie, Ishi Buffam, Hjalmar Laudon, and Björn Malmqvist. Landscape-controlled chemistry variation affects communities and ecosystem function in headwater streams. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 64(11):1563–1572, November 2007. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f07-118.

Modarressie:2007:LRU

[1704] Ricarda Modarressie and Theo C. M. Bakker. A limited role for ultraviolet radiation when threespine sticklebacks (Gasterosteus aculeatus) prey upon Daphnia. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 64(11): 1573–1580, November 2007. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f07-120.

Bigelow:2007:DHD

[1705] Keith A. Bigelow and Mark N. Maunder. Does habitat or depth influence catch rates of pelagic species? Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 64(11):1581–1594, November 2007. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f07-115.

Abernethy:2007:WDF

[1706] Kirsten E. Abernethy, Edward H. Allison, Philip P. Molloy, and Isabelle M. Côté. Why do fishers fish where they fish? Using the ideal free distribution to understand the behaviour of artisanal reef fishers. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 64(11):1595–1604, November 2007. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f07-125.

Niyogi:2007:BVI

[1707] Soumya Niyogi, Gregory G. Pyle, and Chris M. Wood. Branchial versus intestinal zinc uptake in wild yellow perch (Perca flavescens) from reference and metal-contaminated aquatic ecosystems. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 64(11):1605–1613, November 2007. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f07-124.

Rajasingh:2007:WSP

[1708] Hannah Rajasingh, Dag Inge Våge, Scott A. Pavey, and Stig W. Omholt. Why are salmonids pink? Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 64 (11):1614–1627, November 2007. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/

doi/full/10.1139/f07-119.

Kattner:2007:PMZ

[1709] G. Kattner, W. Hagen, R. F. Lee, R. Campbell, D. Deibel, S. Falk-Petersen, M. Graeve, B. W. Hansen, H. J. Hirche, S. H. Jónasdóttir, M. L. Madsen, P. Mayzaud, D. Müller-Navarra, P. D. Nichols, G.-A. Paffenhöfer, D. Pond, H. Saito, D. Stübing, and P. Virtue. Perspectives on marine zooplankton lipids. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 64(11):1628–1639, November 2007. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f07-122.

Weidel:2007:DBL

[1710] Brian C. Weidel, Takayuki Ushikubo, Stephen R. Carpenter, Noriko T. Kita, Jonathan J. Cole, James F. Kitchell, Michael L. Pace, and John W.

Valley. Diary of a bluegill (Lepomis macrochirus): daily δ ^{13}C and δ ^{18}O records in otoliths by ion microprobe. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 64(12):1641–1645, December 2007. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f07-157.

Glemet:2007:STG

[1711] Hélène Glémet and Marco A. Rodríguez. Short-term growth (RNA/DNA ratio) of yellow perch (Perca flavescens) in relation to environmental influences and spatio-temporal variation in a shallow fluvial lake. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 64(12):1646–1655, December 2007. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f07-126.

Portilla:2007:ENI

[1712] Enrique Portilla, Eddie McKenzie, Doug Beare, and Dave Reid. Estimating natural interstage egg mortality of Atlantic mackerel (Scomber scombrus) and horse mackerel (Trachurus trachurus) in the Northeast Atlantic using a stochastic model. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 64(12):1656–1668, December 2007. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f07-128.

Bar:2007:DMN

[1713] Nadav S. Bar, Trygve Sigholt, Karl D. Shearer, and Åshild Krogdahl. A dynamic model of nutrient pathways, growth, and body composition in fish. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 64(12):1669–1682, December 2007. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f07-127.

Barnett-Johnson:2007:ICW

[1714] Rachel Barnett-Johnson, Churchill B. Grimes, Chantell F. Royer, and Christopher J. Donohoe. Identifying the contribution of wild and hatchery Chinook salmon (Oncorhynchus tshawytscha) to the ocean fishery using otolith microstructure as natural tags. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 64(12):1683–1692, December 2007. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f07-129.

Walks:2007:PPF

[1715] D. J. Walks. Persistence of plankton in flowing water. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 64(12):1693–1702, December 2007. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f07-131.

Kirkwood:2007:DGT

[1716] Andrea E. Kirkwood, Troina Shea, Leland J. Jackson, and Edward Mc-Cauley. Didymosphenia geminata in two Alberta headwater rivers: an emerging invasive species that challenges conventional views on algal bloom development. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 64(12): 1703–1709, December 2007. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f07-152.

Hallfredsson:2007:EPP

[1717] Elvar H. Hallfredsson and Torstein Pedersen. Effects of predation from pelagic 0-group cod (Gadus morhua) on mortality rates of capelin (Mallotus villosus) larvae in the Barents Sea. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 64(12):1710–1722, December 2007. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f07-137.

Herborg:2007:CDI

[1718] Leif-Matthias Herborg, Nicholas E. Mandrak, Becky C. Cudmore, and Hugh J. MacIsaac. Comparative distribution and invasion risk of snakehead (Channidae) and Asian carp (Cyprinidae) species in North America. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 64(12):1723–1735, December 2007. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f07-130.

Dufour:2007:WCL

[1719] Elise Dufour, Daniel Gerdeaux, and Christopher M. Wurster. Whitefish (Coregonus lavaretus) respiration rate governs intra-otolith variation of δ ^{13}C values in Lake Annecy. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 64(12):1736–1746, December 2007. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f07-132.

Strand:2007:VMA

[1720] Espen Strand and Geir Huse. Vertical migration in adult Atlantic cod (Gadus morhua). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 64(12):1747– 1760, December 2007. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f07-135.

Selch:2007:CCP

[1721] Trevor M. Selch and Steven R. Chipps. The cost of capturing prey: measuring largemouth bass (Micropterus salmoides) foraging activity using glycolytic enzymes (lactate dehydrogenase). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 64(12):1761–1769, December 2007. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f07-133.

Hirner:2007:ERT

[1722] Joanna Lynne McGarvie Hirner and Sean P. Cox. Effects of rainbow trout (Oncorhynchus mykiss) on amphibians in productive recreational fishing lakes of British Columbia. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 64(12):1770–1780, December 2007. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f07-139.

Henderson:2007:CSE

[1723] Mark J. Henderson and John K. Horne. Comparison of in situ, ex situ, and backscatter model estimates of Pacific hake (Merluccius productus) target strength. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 64(12):1781– 1794, December 2007. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f07-134.

Yoklavich:2007:FIA

[1724] Mary M. Yoklavich, Milton S. Love, and Karin A. Forney. A fisheryindependent assessment of an overfished rockfish stock, cowcod (Sebastes levis), using direct observations from an occupied submersible. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 64(12):1795–1804, December 2007. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f07-145.

Kim:2007:EDQ

[1725] Gene W. Kim, Alpa P. Wintzer, Trisha K. Menker, Roy A. Stein, John M. Dettmers, Russell A. Wright, and Dennis R. DeVries. Effect of detritus quality on growth and survival of gizzard shad (Dorosoma cepedianum): potential importance to benthic–pelagic coupling. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 64(12):1805–1815, December 2007. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f07-143.

Garvey:2007:HMO

[1726] James E. Garvey, Matt R. Whiles, and Dennis Streicher. A hierarchical model for oxygen dynamics in streams. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 64(12):1816–1827, December 2007. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f07-144.

Anonymous:2007:AII

[1727] Anonymous. Author index / index des auteurs: Volume 64, 2007. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 64(12), December 2007. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f07-903.

Anonymous:2007:CSV

[1728] Anonymous. Contents / sommaire: Volume 64, 2007. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 64(12), December 2007. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f07-902.

Macdonald:2008:AC

[1729] Cameron Macdonald. Announcement / Communiqué. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 65(1), January 2008. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f07-910.

Jackson:2008:EE

[1730] Donald A. Jackson. Editorial/Éditorial. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 65(1), January 2008. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f07-908.

Hutchings:2008:RAMa

[1731] Jeffrey A. Hutchings. Ransom Aldrich Myers (1952–2007): In memoriam. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 65(1), January 2008. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f07-906.

Hutchings:2008:RAMb

[1732] Jeffrey A. Hutchings. Ransom Aldrich Myers (1952–2007): In memoriam. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 65(1), January 2008. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f07-907.

Sprules:2008:ECG

[1733] W. Gary Sprules. Ecological change in Great Lakes communities — a matter of perspective. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 65 (1):1–9, January 2008. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f07-136.

Anttila:2008:CHO

[1734] Katja Anttila, Matti Järvilehto, and Satu Mänttäri. Ca^{2+} handling and oxidative capacity are greatly impaired in swimming muscles of hatchery-reared versus wild Atlantic salmon (Salmo salar). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 65(1):10–16, January 2008. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f07-142.

Szerlong:2008:SMM

[1735] R. Glenn Szerlong and David E. Rundio. A statistical modeling method for estimating mortality and abundance of spawning salmon from a time series of counts. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 65(1):17–26, January 2008. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f07-141.

Gillis:2008:BIS

[1736] Darren M. Gillis, Adriaan D. Rijnsdorp, and Jan Jaap Poos. Behavioral inferences from the statistical distribution of commercial catch: patterns of targeting in the landings of the Dutch beam trawler fleet. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 65(1):27–37, January 2008. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f07-147.

Sakamaki:2008:RBB

[1737] Takashi Sakamaki and John S. Richardson. Retention, breakdown, and biological utilisation of deciduous tree leaves in an estuarine tidal flat of southwestern British Columbia, Canada. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 65(1):38–46, January 2008. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f07-151.

Pannard:2008:PST

[1738] Alexandrine Pannard, Myriam Bormans, and Yvan Lagadeuc. Phytoplankton species turnover controlled by physical forcing at different time scales. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 65(1):47–60, January 2008. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f07-149.

Bouchard:2008:RIL

[1739] Judith Bouchard and Daniel Boisclair. The relative importance of local, lateral, and longitudinal variables on the development of habitat quality models for a river. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 65 (1):61–73, January 2008. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f07-140.

Tovar-Avila:2008:DFS

[1740] Javier Tovar-Ávila, Christopher Izzo, Terence I. Walker, J. Matias Braccini, and Robert W. Day. Dorsal-fin spine growth of Heterodontus portusjacksoni: a general model that applies to dorsal-fin spines of chondrichthyans? Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 65(1):74–82, January 2008. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f07-146.

Bennett:2008:WYE

[1741] K. E. Bennett, J. J. Gibson, and P. M. McEachern. Water-yield estimates for critical loadings assessment: comparisons of gauging methods versus an isotopic approach. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 65 (1):83–99, January 2008. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f07-155.

Watras:2008:RTR

[1742] C. J. Watras and K. A. Morrison. The response of two remote, temperate lakes to changes in atmospheric mercury deposition, sulfate, and the water cycle. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 65(1):100–116, January 2008. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f07-159.

Rivot:2008:HBM

[1743] Etienne Rivot, Etienne Prévost, Anne Cuzol, Jean-Luc Baglinière, and Eric Parent. Hierarchical Bayesian modelling with habitat and time covariates for estimating riverine fish population size by successive removal method. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 65(1):117–133, January 2008. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f07-153.

Ashford:2008:OCI

[1744] Julian R. Ashford, Cynthia M. Jones, Eileen E. Hofmann, Inigo Everson, Carlos A. Moreno, Guy Duhamel, and Richard Williams. Otolith chemistry indicates population structuring by the Antarctic Circumpolar Current. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 65(2):135–146, February 2008. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f07-158.

Bradbury:2008:LGC

[1745] I. R. Bradbury, S. E. Campana, and P. Bentzen. Low genetic connectivity in an estuarine fish with pelagic larvae. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 65(2):147–158, February 2008. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f07-154.

Snelgrove:2008:TVF

[1746] Paul V. R. Snelgrove, Ian R. Bradbury, Brad deYoung, and Sandra Fraser. Temporal variation in fish egg and larval production by pelagic and bottom spawners in a large Newfoundland coastal embayment. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 65(2):159–175, February 2008. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f07-148.

Mier:2008:EAS

[1747] Kathryn L. Mier and Susan J. Picquelle. Estimating abundance of spatially aggregated populations: comparing adaptive sampling with other survey designs. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 65(2):176–197, February 2008. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f07-138.

Zhao:2008:LHV

[1748] Yingming Zhao, Brian J. Shuter, and Donald A. Jackson. Life history variation parallels phylogeographical patterns in North American walleye (Sander vitreus) populations. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 65 (2):198–211, February 2008. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f07-162.

Holt:2008:EVS

[1749] Kendra R. Holt and Sean P. Cox. Evaluation of visual survey methods for monitoring Pacific salmon (Oncorhynchus spp.) escapement in relation to conservation guidelines. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 65 (2):212–226, February 2008. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f07-160.

Velez-Espino:2008:MID

[1750] Luis Antonio Vélez-Espino, Robert L. McLaughlin, and Thomas C. Pratt. Management inferences from a demographic analysis of sea lamprey (Petromyzon marinus) in the Laurentian Great Lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 65(2):227–244, February 2008. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f07-166.

Wilson:2008:LIS

[1751] Henry F. Wilson and Marguerite A. Xenopoulos. Landscape influences on stream fish assemblages across spatial scales in a northern Great Plains ecoregion. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 65(2):245–257, February 2008. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f07-165.

Lambert:2008:PEI

[1752] Daniel Lambert, Antonella Cattaneo, and Richard Carignan. Periphyton as an early indicator of perturbation in recreational lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 65(2):258–265, February 2008. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f07-168.

Piccolo:2008:WVI

[1753] John J. Piccolo, Nicholas F. Hughes, and Mason D. Bryant. Water velocity influences prey detection and capture by drift-feeding juvenile Coho salmon (Oncorhynchus kisutch) and steelhead (Oncorhynchus mykiss irideus). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 65(2):266–275, February 2008. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f07-172.

Munro:2008:ESI

[1754] Andrew R. Munro, Bronwyn M. Gillanders, Travis S. Elsdon, David A. Crook, and Andrew C. Sanger. Enriched stable isotope marking of juvenile golden perch (Macquaria ambigua) otoliths. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 65(2):276–285, February 2008. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f08-010.

Forrest:2008:ASM

[1755] Robyn E. Forrest, Steven J. D. Martell, Michael C. Melnychuk, and Carl J. Walters. An age-structured model with leading management parameters, incorporating age-specific selectivity and maturity. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 65(2):286–296, February 2008. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f07-156.

Chu:2008:IAT

[1756] Cindy Chu, Nicholas E. Jones, Nicholas E. Mandrak, Andrew R. Piggott, and Charles K. Minns. The influence of air temperature, groundwater discharge, and climate change on the thermal diversity of stream fishes in southern Ontario watersheds. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 65 (2):297–308, February 2008. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f08-007.

Caroffino:2008:SSL

[1757] David C. Caroffino, Loren M. Miller, Anne R. Kapuscinski, and Joseph J. Ostazeski. Stocking success of local-origin fry and impact of hatchery ancestry: monitoring a new steelhead (Oncorhynchus mykiss) stocking program in a Minnesota tributary to Lake Superior. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 65(2):309–318, February 2008. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f07-167.

McCleary:2008:PMS

[1758] Richard J. McCleary and Marwan A. Hassan. Predictive modeling and spatial mapping of fish distributions in small streams of the Canadian Rocky Mountain foothills. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 65 (2):319–333, February 2008. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f07-161.

Rypel:2008:DAE

[1759] Andrew L. Rypel and Craig A. Layman. Degree of aquatic ecosystem fragmentation predicts population characteristics of gray snapper (Lutjanus griseus) in Caribbean tidal creeks. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 65(3):335–339, March 2008. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f07-192.

Manderson:2008:SSP

[1760] John P. Manderson. The spatial scale of phase synchrony in winter flounder (Pseudopleuronectes americanus) production increased among southern New England nurseries in the 1990s. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 65(3):340–351, March 2008. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f07-169.

Booker:2008:MTM

[1761] Douglas J. Booker, Neil C. Wells, and I. Philip Smith. Modelling the trajectories of migrating Atlantic salmon (Salmo salar). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 65(3):352–361, March 2008. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f07-173.

Daunt:2008:ISF

[1762] Francis Daunt, Sarah Wanless, Simon P. R. Greenstreet, Henrik Jensen, Keith C. Hamer, and Michael P. Harris. The impact of the sandeel fishery closure on seabird food consumption, distribution, and productivity in the northwestern North Sea. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 65 (3):362–381, March 2008. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f07-164.

Purchase:2008:TSS

[1763] Craig F. Purchase and Jeffrey A. Hutchings. A temporally stable spatial pattern in the spawner density of a freshwater fish: evidence for an ideal despotic distribution. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 65 (3):382–388, March 2008. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f07-171.

Jorgensen:2008:GEC

[1764] Hanne B. H. Jørgensen, Cino Pertoldi, Michael M. Hansen, Daniel E. Ruzzante, and Volker Loeschcke. Genetic and environmental correlates of morphological variation in a marine fish: the case of Baltic Sea herring (Clupea harengus). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 65 (3):389–400, March 2008. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f07-177.

Belk:2008:HIM

[1765] Mark C. Belk, Lisa J. Benson, Josh Rasmussen, and Steven L. Peck. Hatchery-induced morphological variation in an endangered fish: a challenge for hatchery-based recovery efforts. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 65(3):401–408, March 2008. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f07-176.

Martell:2008:RAH

[1766] Steven J. D. Martell, Carl J. Walters, and Ray Hilborn. Retrospective analysis of harvest management performance for Bristol Bay and Fraser River sockeye salmon (Oncorhynchus nerka). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 65(3):409–424, March 2008. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f07-170.

Puebla:2008:PGS

[1767] Oscar Puebla, Jean-Marie Sévigny, Bernard Sainte-Marie, Jean-Claude Brêthes, AnnDorte Burmeister, Earl G. Dawe, and Mikio Moriyasu. Population genetic structure of the snow crab (Chionoecetes opilio) at the Northwest Atlantic scale. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 65 (3):425–436, March 2008. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f07-163.

Haxton:2008:MAI

[1768] Tim J. Haxton and C. Scott Findlay. Meta-analysis of the impacts of water management on aquatic communities. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 65(3):437–447, March 2008. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f07-175.

Knoll:2008:IZM

[1769] Lesley B. Knoll, Orlando Sarnelle, Stephen K. Hamilton, Carrie E. H. Kissman, Alan E. Wilson, Joan B. Rose, and Mechelle R. Morgan. Invasive zebra mussels (Dreissena polymorpha) increase cyanobacterial toxin concentrations in low-nutrient lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 65(3):448–455, March 2008. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f07-181.

Rindorf:2008:GTD

[1770] Anna Rindorf, Henrik Jensen, and Corinna Schrum. Growth, temperature, and density relationships of North Sea cod (Gadus morhua). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 65(3):456–470, March 2008. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f07-150.

Sharma:2008:PSB

[1771] Sapna Sharma and Donald A. Jackson. Predicting smallmouth bass (Micropterus dolomieu) occurrence across North America under climate change: a comparison of statistical approaches. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 65(3):471–481, March 2008. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f07-178.

Simpfendorfer:2008:VPA

[1772] Colin Ashley Simpfendorfer, Michelle RaNae Heupel, and Angela B. Collins. Variation in the performance of acoustic receivers and its implication for positioning algorithms in a riverine setting. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 65(3):482–492, March 2008. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f07-180.

Haag:2008:TAA

[1773] Wendell R. Haag and Amy M. Commens-Carson. Testing the assumption of annual shell ring deposition in freshwater mussels. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 65(3):493–508, March 2008. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f07-182.

Stammler:2008:SMD

[1774] Katie L. Stammler, Robert L. McLaughlin, and Nicholas E. Mandrak. Streams modified for drainage provide fish habitat in agricultural areas.

Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 65(3):509–522, March 2008. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f07-183.

Ottersen:2008:PLT

[1775] Geir Ottersen. Pronounced long-term juvenation in the spawning stock of arcto-Norwegian cod (Gadus morhua) and possible consequences for recruitment. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 65(3):523–534, March 2008. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f07-185.

Madenjian:2008:NEL

[1776] Charles P. Madenjian, Brian D. Chipman, and J. Ellen Marsden. New estimates of lethality of sea lamprey (Petromyzon marinus) attacks on lake trout (Salvelinus namaycush): implications for fisheries management. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 65(3):535–542, March 2008. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f07-187.

Prairie:2008:CLL

[1777] Yves T. Prairie. Carbocentric limnology: looking back, looking forward. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 65(3):543–548, March 2008. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f08-011.

Reid:2008:CRS

[1778] Janet W. Reid and Patrick L. Hudson. Comment on “Rate of species introductions in the Great Lakes via ships’ ballast water and sediments”. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 65(3):549–553, March 2008. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f08-018.

Drake:2008:RCR

[1779] John M. Drake and David M. Lodge. Reply to the comment by Reid and Hudson on “Rate of species introductions in the Great Lakes via ships’ ballast water and sediments”. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 65 (3):554–555, March 2008. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f08-019.

Peterson:2008:ATO

[1780] Douglas P. Peterson, Bruce E. Rieman, Jason B. Dunham, Kurt D. Fausch, and Michael K. Young. Analysis of trade-offs between threats of invasion by nonnative brook trout (Salvelinus fontinalis) and intentional isolation for native westslope cutthroat trout (Oncorhynchus clarkii lewisi). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 65(4):557–573, April 2008. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/f07-
184.

Sourisseau:2008:KDV

[1781] M. Sourisseau, Y. Simard, and F. J. Saucier. Krill diel vertical migration fine dynamics, nocturnal overturns, and their roles for aggregation in stratified flows. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 65 (4):574–587, April 2008. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f07-179.

Ralston:2008:SVF

[1782] Stephen Ralston and Michael R. O’Farrell. Spatial variation in fishing intensity and its effect on yield. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 65 (4):588–599, April 2008. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f07-174.

Moles:2008:GIW

[1783] Matthew D. Moles, Thomas A. Johnston, Beren W. Robinson, William C. Leggett, and John M. Casselman. Is gonadal investment in walleye (Sander vitreus) dependent on body lipid reserves? A multipopulation comparative analysis. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 65 (4):600–614, April 2008. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f07-186.

Burness:2008:BCI

[1784] Gary Burness, Albrecht I. Schulte-Hostedde, and Robert Montgomerie. Body condition influences sperm energetics in lake whitefish (Coregonus clupeaformis). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 65(4):615–620, April 2008. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/f07-
188.

Dicken:2008:EJA

[1785] Matthew Laurence Dicken, Anthony John Booth, and Malcolm John Smale. Estimates of juvenile and adult raggedtooth shark (Carcharias taurus) abundance along the east coast of South Africa. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 65(4):621–632, April 2008. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f07-190.

Ohlund:2008:LHL

[1786] Gunnar Öhlund, Fredrik Nordwall, Erik Degerman, and Torleif Eriksson. Life history and large-scale habitat use of brown trout (Salmo trutta) and brook trout (Salvelinus fontinalis) — implications for species replacement patterns. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 65 (4):633–644, April 2008. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f08-003.

Haxton:2008:VLS

[1787] Tim J. Haxton and C. Scott Findlay. Variation in lake sturgeon (Acipenser fulvescens) abundance and growth among river reaches in a large regulated river. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 65 (4):645–657, April 2008. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f08-005.

Boyer:2008:RTO

[1788] Matthew C. Boyer, Clint C. Muhlfeld, and Fred W. Allendorf. Rainbow trout (Oncorhynchus mykiss) invasion and the spread of hybridization with native westslope cutthroat trout (Oncorhynchus clarkii lewisi). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 65(4):658–669, April 2008. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f08-001.

Horrigan:2008:TPA

[1789] Nelli Horrigan and Donald J. Baird. Trait patterns of aquatic insects across gradients of flow-related factors: a multivariate analysis of Canadian national data. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 65 (4):670–680, April 2008. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f07-191.

Reichert:2008:SVS

[1790] William L. Reichert, Correigh M. Greene, and Robert E. Bilby. Seasonal variations in stable isotope ratios of juvenile Coho salmon (Oncorhynchus kisutch) from western Washington rivers. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 65(4):681–690, April 2008. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f07-189.

Bue:2008:OHC

[1791] Brian G. Bue, Ray Hilborn, and Michael R. Link. Optimal harvesting considering biological and economic objectives. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 65(4):691–700, April 2008. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f08-009.

Holtby:2008:ESU

[1792] L. Blair Holtby and Max L. Bothwell. Effects of solar ultraviolet radiation on the behaviour of juvenile Coho salmon (Oncorhynchus kisutch): avoidance, feeding, and agonistic interactions. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 65(4):701–711, April 2008. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f08-013.

Coloso:2008:DIC

[1793] James J. Coloso, Jonathan J. Cole, Paul C. Hanson, and Michael L. Pace. Depth-integrated, continuous estimates of metabolism in a clear-water lake. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 65(4):712–722, April 2008. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f08-006.

Hinz:2008:LSR

[1794] Hilmar Hinz, Jan G. Hiddink, James Forde, and Michel J. Kaiser. Largescale responses of nematode communities to chronic otter-trawl disturbance. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 65(4):723–732, April 2008. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f08-002.

Campana:2008:BRC

[1795] Steven E. Campana, John M. Casselman, and Cynthia M. Jones. Bomb radiocarbon chronologies in the Arctic, with implications for the age validation of lake trout (Salvelinus namaycush) and other Arctic species. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 65(4):733–743, April 2008. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f08-012.

Tatara:2008:EES

[1796] Christopher P. Tatara, Stephen C. Riley, and Julie A. Scheurer. Environmental enrichment in steelhead (Oncorhynchus mykiss) hatcheries: field evaluation of aggression, foraging, and territoriality in natural and hatchery fry. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 65(4):744–753, April 2008. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f08-004.

Berejikian:2008:ISO

[1797] Barry A. Berejikian, Thom Johnson, Richard S. Endicott, and Joy Lee-Waltermire. Increases in steelhead (Oncorhynchus mykiss) redd abundance resulting from two conservation hatchery strategies in the Hamma Hamma River, Washington. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 65 (4):754–764, April 2008. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f08-014.

Pelletier:2008:MIA

[1798] Dominique Pelletier, Joachim Claudet, Jocelyne Ferraris, Lisandro Benedetti-Cecchi, and José Antonio Garcìa-Charton. Models and indicators for assessing conservation and fisheries-related effects of marine protected areas. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 65(4):765–779, April 2008. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/f08-
026.

Anonymous:2008:TYA

[1799] Anonymous. Thirty years of aquatic science at the Dorset Environmental Science Centre: transforming understanding of factors that regulate aquatic ecosystems on the southern Canadian Shield / [Trente ans de sciences aquatiques au Centre des sciences environnementales de Dorset : l’évolution de la compréhension des facteurs qui régulent les écosystèmes aquatiques de la partie méridionale du Bouclier canadien]. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 65(5), May 2008. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f08-900.

Yan:2008:IDS

[1800] Norman D. Yan, Andrew M. Paterson, Keith M. Somers, and Wolfgang A. Scheider. An introduction to the Dorset special issue: transforming understanding of factors that regulate aquatic ecosystems on the southern Canadian Shield. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 65 (5):781–785, May 2008. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f08-077.

Keller:2008:RBD

[1801] W. (Bill) Keller, Andrew M. Paterson, Keith M. Somers, Peter J. Dillon, Jocelyne Heneberry, and Andrea Ford. Relationships between dissolved organic carbon concentrations, weather, and acidification in small Boreal Shield lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 65(5):786–795, May 2008. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/f07-
193.

Eimers:2008:ISC

[1802] M. Catherine Eimers, Jim Buttle, and Shaun A. Watmough. Influence of seasonal changes in runoff and extreme events on dissolved organic carbon trends in wetland- and upland-draining streams. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 65(5):796–808, May 2008. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f07-194.

Molot:2008:LTT

[1803] L. A. Molot and Peter J. Dillon. Long-term trends in catchment export and lake concentrations of base cations in the Dorset study area, central Ontario. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 65(5):809–820, May 2008. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f08-035.

Watmough:2008:ECW

[1804] Shaun A. Watmough and Julian Aherne. Estimating calcium weathering rates and future lake calcium concentrations in the Muskoka–Haliburton region of Ontario. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 65(5):821–833, May 2008. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/f07-
196.

Adkinson:2008:DIM

[1805] Angela Adkinson, Shaun A. Watmough, and Peter J. Dillon. Droughtinduced metal release from a wetland at Plastic Lake, central Ontario. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 65(5):834–845, May 2008. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f07-195.

Paterson:2008:LTC

[1806] Andrew M. Paterson, Jennifer G. Winter, Kenneth H. Nicholls, Bev J. Clark, Charles W. Ramcharan, Norman D. Yan, and Keith M. Somers. Long-term changes in phytoplankton composition in seven Canadian Shield lakes in response to multiple anthropogenic stressors. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 65(5):846–861, May 2008. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f08-022.

Yan:2008:LTT

[1807] Norman D. Yan, Keith M. Somers, Robert E. Girard, Andrew M. Paterson, W. (Bill) Keller, Charles W. Ramcharan, James A. Rusak, Ron Ingram, George E. Morgan, and John M. Gunn. Long-term trends in zooplankton of Dorset, Ontario, lakes: the probable interactive effects of changes in pH, total phosphorus, dissolved organic carbon, and predators. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 65(5):862–877, May 2008. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f07-197.

Rusak:2008:RCD

[1808] James A. Rusak, Norman D. Yan, and Keith M. Somers. Regional climatic drivers of synchronous zooplankton dynamics in north-temperate lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 65(5):878–889, May 2008. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f08-043.

Gilbert:2008:RBM

[1809] Beth Gilbert, Peter J. Dillon, Keith M. Somers, Ron A. Reid, and Lem Scott. Response of benthic macroinvertebrate communities to El Niño related drought events in six upland streams in south-central Ontario. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 65(5):890–905, May 2008. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f08-057.

Lento:2008:CLB

[1810] Jennifer Lento, Peter J. Dillon, Keith M. Somers, and Ron A. Reid. Changes in littoral benthic macroinvertebrate communities in relation to water chemistry in 17 Precambrian Shield lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 65(5):906–918, May 2008. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f08-033.

Bowman:2008:TSC

[1811] Michelle F. Bowman, Ron Ingram, Ron A. Reid, Keith M. Somers, Norman D. Yan, Andrew M. Paterson, George E. Morgan, and John M. Gunn. Temporal and spatial concordance in community composition of phytoplankton, zooplankton, macroinvertebrate, crayfish, and fish on the Precambrian Shield. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 65 (5):919–932, May 2008. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f08-034.

Quinlan:2008:LTA

[1812] Roberto Quinlan, Roland I. Hall, Andrew M. Paterson, Brian F. Cumming, and John P. Smol. Long-term assessments of ecological effects of anthropogenic stressors on aquatic ecosystems from paleoecological analyses: challenges to perspectives of lake management. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 65(5):933–944, May 2008. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f08-027.

Hsieh:2008:SAS

[1813] Chih hao Hsieh, Christian S. Reiss, Roger P. Hewitt, and George Sugihara. Spatial analysis shows that fishing enhances the climatic sensitivity of marine fishes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 65(5):947–961, May 2008. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/f08-
017.

Michielsens:2008:CMB

[1814] Catherine G. J. Michielsens, Murdoch K. McAllister, Sakari Kuikka, Samu Mäntyniemi, Atso Romakkaniemi, Tapani Pakarinen, Lars Karlsson, and Laura Uusitalo. Combining multiple Bayesian data analyses in a sequential framework for quantitative fisheries stock assessment. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 65(5):962–974, May 2008. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f08-015.

Pine:2008:CRS

[1815] William E. Pine III, Steven J. D. Martell, Olaf P. Jensen, Carl J. Walters, and James F. Kitchell. Catch-and-release and size limit regulations for blue, white, and striped marlin: the role of postrelease survival in effective policy design. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 65 (5):975–988, May 2008. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f08-020.

Costantini:2008:EHH

[1816] Marco Costantini, Stuart A. Ludsin, Doran M. Mason, Xinsheng Zhang, William C. Boicourt, and Stephen B. Brandt. Effect of hypoxia on habitat quality of striped bass (Morone saxatilis) in Chesapeake Bay. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 65(5):989–1002, May 2008. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f08-021.

Surette:2008:RML

[1817] Tobie Surette and Jacques Allard. A regression mixture with logistic weights for modeling heterogeneous crustacean growth data when the maturity phase is unknown. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 65 (5):1003–1011, May 2008. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f08-023.

Narum:2008:ILR

[1818] Shawn R. Narum, Joseph S. Zendt, David Graves, and William R. Sharp. Influence of landscape on resident and anadromous life history types of Oncorhynchus mykiss. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 65 (6):1013–1023, June 2008. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/F08-025.

Trenkel:2008:TSB

[1819] Verena M. Trenkel. A two-stage biomass random effects model for stock assessment without catches: What can be estimated using only biomass survey indices? Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 65(6):1024– 1035, June 2008. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
F08-028.

Berreville:2008:EDA

[1820] Olivier F. Berreville, Alain F. Vézina, Keith R. Thompson, and Bert Klein. Exploratory data analysis of the interactions among physics, food web structure, and function in two Arctic polynyas. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 65(6):1036–1046, June 2008. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/F08-016.

Treble:2008:GAA

[1821] Margaret A. Treble, Steven E. Campana, Rick J. Wastle, Cynthia M. Jones, and Jesper Boje. Growth analysis and age validation of a deepwater Arctic fish, the Greenland halibut (Reinhardtius hippoglossoides). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 65(6):1047–1059, June 2008. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/F08-030.

Donohoe:2008:IWC

[1822] Christopher J. Donohoe, Peter B. Adams, and Chantell F. Royer. Influence of water chemistry and migratory distance on ability to distinguish progeny of sympatric resident and anadromous rainbow trout (Oncorhynchus mykiss). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 65(6): 1060–1075, June 2008. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/F08-029.

Friedland:2008:DFP

[1823] Kevin D. Friedland, Jonathan A. Hare, Grayson B. Wood, Laurel A. Col, Lawrence J. Buckley, David G. Mountain, Joseph Kane, Jon Brodziak, R. Gregory Lough, and Cynthia H. Pilskaln. Does the fall phytoplankton bloom control recruitment of Georges Bank haddock, Melanogrammus aeglefinus, through parental condition? Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 65(6):1076–1086, June 2008. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/F08-040.

Short:2008:DTB

[1824] Terry M. Short, Lawrence R. Deweese, and Neil M. Dubrovsky. Discrimination of trait-based characteristics by trace element bioaccumulation in riverine fishes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 65(6): 1087–1100, June 2008. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/F08-036.

Kastelle:2008:BRA

[1825] Craig R. Kastelle, Daniel K. Kimura, and Betty J. Goetz. Bomb radiocarbon age validation of Pacific Ocean perch (Sebastes alutus) using new statistical methods. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 65 (6):1101–1112, June 2008. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/F08-038.

Binder:2008:RTC

[1826] Thomas R. Binder and D. Gordon McDonald. The role of temperature in controlling diel activity in upstream migrant sea lampreys (Petromyzon marinus). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 65(6):1113–1121, June 2008. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/F08-
070.

Schnell:2008:BIC

[1827] Sabine Schnell, Doris Schiedek, Rolf Schneider, Lennart Balk, Pekka J. Vuorinen, Heta Karvinen, and Thomas Lang. Biological indications of contaminant exposure in Atlantic cod (Gadus morhua) in the Baltic Sea. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 65(6):1122–1134, June 2008. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/F08-042.

Penton:2008:PLE

[1828] Paulette M. Penton and Gail K. Davoren. Patterns of larval emergence of capelin (Mallotus villosus) and environmental cues at demersal spawning sites on the northeastern coast of Newfoundland. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 65(6):1135–1143, June 2008. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/F08-037.

Cooke:2008:ETU

[1829] Sandra L. Cooke, Craig E. Williamson, Dina M. Leech, Wiebke J. Boeing, and Lisette Torres. Effects of temperature and ultraviolet radiation on diel vertical migration of freshwater crustacean zooplankton. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 65(6):1144–1152, June 2008. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/F08-039.

Hand:2008:SMT

[1830] Carrol P. Hand, Stuart A. Ludsin, Brian J. Fryer, and J. Ellen Marsden. Statolith microchemistry as a technique for discriminating among Great Lakes sea lamprey (Petromyzon marinus) spawning tributaries. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 65(6):1153–1164, June 2008. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/F08-045.

Hudon:2008:CIH

[1831] Christiane Hudon and Richard Carignan. Cumulative impacts of hydrology and human activities on water quality in the St. Lawrence River (Lake Saint-Pierre, Quebec, Canada). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 65(6):1165–1180, June 2008. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/F08-069.

Ruttenberg:2008:STV

[1832] Benjamin I. Ruttenberg, Scott L. Hamilton, and Robert R. Warner. Spatial and temporal variation in the natal otolith chemistry of a Hawaiian reef fish: prospects for measuring population connectivity. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 65(6):1181–1192, June 2008. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/F08-052.

Doray:2008:GMA

[1833] Mathieu Doray, Pierre Petitgas, and Erwan Josse. A geostatistical method for assessing biomassof tuna aggregations around moored fish aggregating devices with star acoustic surveys. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 65(6):1193–1205, June 2008. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/F08-050.

DeSousa:2008:RLM

[1834] Simon De Sousa, Bernadette Pinel-Alloul, and Antonia Cattaneo. Response of littoral macroinvertebrate communities on rocks and sediments to lake residential development. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 65(6):1206–1216, June 2008. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/F08-031.

Hoftyzer:2008:GEI

[1835] Eileen Hoftyzer, Josef Daniel Ackerman, Todd J. Morris, and Gerald L. Mackie. Genetic and environmental implications of reintroducing laboratory-raised unionid mussels to the wild. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 65(6):1217–1229, June 2008. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/F08-024.

Stoner:2008:ERF

[1836] Allan W. Stoner, Clifford H. Ryer, Steven J. Parker, Peter J. Auster, and W. Waldo Wakefield. Evaluating the role of fish behavior in surveys conducted with underwater vehicles. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 65(6):1230–1243, June 2008. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/F08-032.

Polacheck:2008:EIC

[1837] Tom Polacheck, J. Paige Eveson, Geoff M. Laslett, Kenneth H. Pollock, and William S. Hearn. Erratum: Integrating catch-at-age and multiyear tagging data: a combined Brownie and Petersen estimation approach in a fishery context. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 65(6):1244, June 2008. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/F08-
072.

Rogers:2008:UMR

[1838] Bonnie L. Rogers, Christopher G. Lowe, Esteban Fernández-Juricic, and Lawrence R. Frank. Utilizing magnetic resonance imaging (MRI) to assess the effects of angling-induced barotrauma on rockfish (Sebastes). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 65(7):1245–1249, July 2008. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/F08-102.

Wenger:2008:SFO

[1839] Seth J. Wenger, James T. Peterson, Mary C. Freeman, Byron J. Freeman, and D. David Homans. Stream fish occurrence in response to impervious cover, historic land use, and hydrogeomorphic factors. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 65(7):1250–1264, July 2008. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/F08-046.

Gertzen:2008:QIP

[1840] Erin Gertzen, Oriana Familiar, and Brian Leung. Quantifying invasion pathways: fish introductions from the aquarium trade. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 65(7):1265–1273, July 2008. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/F08-056.

Johnson:2008:GSM

[1841] Marc A. Johnson and Michael A. Banks. Genetic structure, migration, and patterns of allelic richness among Coho salmon (Oncorhynchus kisutch) populations of the Oregon coast. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 65(7):1274–1285, July 2008. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/F08-044.

Jarvis:2008:EBC

[1842] Erica T. Jarvis and Christopher G. Lowe. The effects of barotrauma on the catch-and-release survival of southern California nearshore and shelf rockfish (Scorpaenidae, Sebastes spp.). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 65(7):1286–1296, July 2008. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/F08-071.

Rowlands:2008:DOP

[1843] William Ll. Rowlands, Mark Dickey-Collas, Audrey J. Geffen, and Richard D. M. Nash. Diet overlap and prey selection through metamorphosis in Irish Sea cod (Gadus morhua), haddock (Melanogrammus aeglefinus), and whiting (Merlangius merlangus). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 65(7):1297–1306, July 2008. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/F08-041.

McDowell:2008:PSS

[1844] Jan R. McDowell and John E. Graves. Population structure of striped marlin (Kajikia audax) in the Pacific Ocean based on analysis of microsatellite and mitochondrial DNA. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 65(7):1307–1320, July 2008. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/F08-054.

Warren:2008:ABP

[1845] Joseph D. Warren and Peter H. Wiebe. Accounting for biological and physical sources of acoustic backscatter improves estimates of zooplankton biomass. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 65(7):1321– 1333, July 2008. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
F08-047.

MacKenzie:2008:IEV

[1846] Brian R. MacKenzie, Jan Horbowy, and Fritz W. Köster. Incorporating environmental variability in stock assessment: predicting recruitment, spawner biomass, and landings of sprat (Sprattus sprattus) in the Baltic Sea. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 65(7):1334–1341, July 2008. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/F08-051.

Monaghan:2008:SCM

[1847] Kieran A. Monaghan and Alexander M. Milner. Salmon carcasses as a marine-derived resource for benthic macroinvertebrates in a developing postglacial stream, Alaska. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 65 (7):1342–1351, July 2008. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/F08-058.

Collie:2008:LTS

[1848] Jeremy S. Collie, Anthony D. Wood, and H. Perry Jeffries. Long-term shifts in the species composition of a coastal fish community. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 65(7):1352–1365, July 2008. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/F08-048.

Irwin:2008:EWS

[1849] Brian J. Irwin, Theodore J. Treska, Lars G. Rudstam, Patrick J. Sullivan, James R. Jackson, Anthony J. VanDeValk, and John L. Forney. Estimating walleye (Sander vitreus) density, gear catchability, and mortality using three fishery-independent data sets for Oneida Lake, New York. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 65(7):1366–1378, July 2008. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/F08-062.

Reyjol:2008:AWT

[1850] Yorick Reyjol, Marco A. Rodríguez, Nathalie Dubuc, Pierre Magnan, and Réjean Fortin. Among- and within-tributary responses of riverine fish assemblages to habitat features. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 65 (7):1379–1393, July 2008. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/F08-060.

Hiddink:2008:CBT

[1851] Jan Geert Hiddink, Adriaan D. Rijnsdorp, and Gerjan Piet. Can bottom trawling disturbance increase food production for a commercial fish species? Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 65(7):1393–1401, July 2008. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/F08-064.

Hook:2008:AVH

[1852] Tomas O. Höök, Edward S. Rutherford, Thomas E. Croley, Doran M. Mason, and Charles P. Madenjian. Annual variation in habitat-specific recruitment success: implications from an individual-based model of Lake Michigan alewife (Alosa pseudoharengus). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 65(7):1402–1412, July 2008. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/F08-066.

Kolding:2008:EAO

[1853] Jeppe Kolding, Lise Haug, and Sigurd Stefansson. Effect of ambient oxygen on growth and reproduction in Nile tilapia (Oreochromis niloticus). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 65(7):1413–1424, July 2008. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/F08-059.

Takahashi:2008:YCS

[1854] Motomitsu Takahashi, Hiroshi Nishida, Akihiko Yatsu, and Yoshiro Watanabe. Year-class strength and growth rates after metamorphosis of Japanese sardine (Sardinops melanostictus) in the western North Pacific Ocean during 1996–2003. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 65 (7):1425–1434, July 2008. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/F08-063.

Branch:2008:MCQ

[1855] Trevor A. Branch and Ray Hilborn. Matching catches to quotas in a multispecies trawl fishery: targeting and avoidance behavior under individual transferable quotas. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 65 (7):1435–1446, July 2008. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/F08-065.

Oyadomari:2008:TGL

[1856] Jason K. Oyadomari and Nancy A. Auer. Transport and growth of larval cisco (Coregonus artedi) in the Keweenaw Current region of Lake Superior. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 65(7):1447–1458, July 2008. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/F08-068.

Holt:2008:UPD

[1857] Carrie A. Holt and Randall M. Peterman. Uncertainties in population dynamics and outcomes of regulations in sockeye salmon (Oncorhynchus nerka) fisheries: implications for management. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 65(7):1459–1474, July 2008. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/F08-053.

Anderson:2008:IMP

[1858] Eric C. Anderson, Robin S. Waples, and Steven T. Kalinowski. An improved method for predicting the accuracy of genetic stock identification. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 65(7):1475–1486, July 2008. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/F08-049.

Wilson:2008:EHH

[1859] Alan E. Wilson, Duane C. Gossiaux, Tomas O. Höök, John P. Berry, Peter F. Landrum, Julianne Dyble, and Stephanie J. Guildford. Evaluation of the human health threat associated with the hepatotoxin microcystin in the muscle and liver tissues of yellow perch (Perca flavescens). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 65(7):1487–1497, July 2008. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/F08-067.

Christensen:2008:SAM

[1860] Asbjørn Christensen, Henrik Jensen, Henrik Mosegaard, Mike St. John, and Corinna Schrum. Sandeel (Ammodytes marinus) larval transport patterns in the North Sea from an individual-based hydrodynamic egg and larval model. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 65(7):1498– 1511, July 2008. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
F08-073.

VanderZanden:2008:MFP

[1861] M. Jake Vander Zanden and Julian D. Olden. A management framework for preventing the secondary spread of aquatic invasive species. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 65(7):1512–1522, July 2008. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/F08-099.

Neilson:2008:VDA

[1862] John D. Neilson and Steven E. Campana. A validated description of age and growth of western Atlantic bluefin tuna (Thunnus thynnus). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 65(8):1523–1527, August 2008. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/F08-127.

Castonguay:2008:CPD

[1863] Martin Castonguay, Stéphane Plourde, Dominique Robert, Jeffrey A. Runge, and Louis Fortier. Copepod production drives recruitment in a marine fish. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 65(8):1528–1531, August 2008. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
F08-126.

Datry:2008:RFC

[1864] Thibault Datry and Scott T. Larned. River flow controls ecological processes and invertebrate assemblages in subsurface flowpaths of an ephemeral river reach. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 65(8): 1532–1544, August 2008. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/F08-075.

Blanchet:2008:EAF

[1865] Simon Blanchet, Géraldine Loot, Louis Bernatchez, and Julian J. Dodson. The effects of abiotic factors and intraspecific versus interspecific competition on the diel activity patterns of Atlantic salmon (Salmo salar) fry. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 65(8):1545–1553, August 2008. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/F08-079.

Hazewinkel:2008:AEA

[1866] Roderick R. O. Hazewinkel, Alexander P. Wolfe, Sergi Pla, Chris Curtis, and Kris Hadley. Have atmospheric emissions from the athabasca oil sands impacted lakes in northeastern Alberta, Canada? Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 65(8):1554–1567, August 2008. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/F08-074.

Ferguson:2008:CTB

[1867] John W. Ferguson, Gene R. Ploskey, Kjell Leonardsson, Richard W. Zabel, and Hans Lundqvist. Combining turbine blade-strike and life cycle models to assess mitigation strategies for fish passing dams. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 65(8):1568–1585, August 2008. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/F08-078.

Martell:2008:PAS

[1868] Steven J. D. Martell, William E. Pine, and Carl J. Walters. Parameterizing age-structured models from a fisheries management perspective. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 65(8):1586–1600, August 2008. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/F08-055.

Martell:2008:EPR

[1869] Steven J. D. Martell and Carl J. Walters. Experimental policies for rebuilding depleted stocks. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 65 (8):1601–1609, August 2008. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/F08-061.

Thayer:2008:FFP

[1870] Julie A. Thayer, Douglas F. Bertram, Scott A. Hatch, Mark J. Hipfner, Leslie Slater, William J. Sydeman, and Yutaka Watanuki. Forage fish of the Pacific rim as revealed by diet of a piscivorous seabird: synchrony and relationships with sea surface temperature. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 65(8):1610–1622, August 2008. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/F08-076.

Kraemer:2008:MCA

[1871] Lisa D. Kraemer, Peter G. C. Campbell, and Landis Hare. Modeling cadmium accumulation in indigenous yellow perch (Perca flavescens). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 65(8):1623–1634, August 2008. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/F08-081.

Uchiyama:2008:EMD

[1872] Tadayasu Uchiyama, Bruce P. Finney, and Milo D. Adkison. Effects of marine-derived nutrients on population dynamics of sockeye salmon (Oncorhynchus nerka). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 65(8): 1635–1648, August 2008. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/F08-085.

Letcher:2008:SVS

[1873] Benjamin H. Letcher and Gregg E. Horton. Seasonal variation in sizedependent survival of juvenile Atlantic salmon (Salmo salar): performance of multistate capture–mark–recapture models. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 65(8):1649–1666, August 2008. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/F08-083.

Gibson:2008:WPI

[1874] A. Jamie F. Gibson, Heather D. Bowlby, and Peter G. Amiro. Are wild populations ideally distributed? Variations in density-dependent habitat use by age class in juvenile Atlantic salmon (Salmo salar). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 65(8):1667–1680, August 2008. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/F08-087.

Ichinokawa:2008:TMR

[1875] Momoko Ichinokawa, Atilio L. Coan, and Yukio Takeuchi. Transoceanic migration rates of young North Pacific albacore, Thunnus alalunga, from conventional tagging data. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 65 (8):1681–1691, August 2008. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/F08-095.

Brown:2008:NND

[1876] Meghan Elizabeth Brown. Nature and nurture in dormancy: dissolved oxygen, pH, and maternal investment impact Bythotrephes longimanus resting egg emergence and neonate condition. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 65(8):1692–1704, August 2008. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/F08-082.

Guenard:2008:CBA

[1877] Guillaume Guénard, Daniel Boisclair, Ola Ugedal, Torbjørn Forseth, and Bror Jonsson. Comparison between activity estimates obtained using bioenergetic and behavioural analyses. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 65(8):1705–1720, August 2008. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://

cdnsciencepub.com/doi/full/10.1139/F08-080.

Roy:2008:APP

[1878] Suzanne Roy, Frédérick Blouin, André Jacques, and Jean-Claude Therriault. Absorption properties of phytoplankton in the Lower Estuary and Gulf of St. Lawrence (Canada). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 65(8):1721–1737, August 2008. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/F08-089.

Palm:2008:NIA

[1879] Stefan Palm, Johan Dannewitz, Torbjörn Järvi, Marja-Liisa Koljonen, Tore Prestegaard, and K. Håkan Olsén. No indications of Atlantic salmon (Salmo salar) shoaling with kin in the Baltic Sea. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 65(8):1738–1748, August 2008. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/F08-088.

Guy:2008:LSE

[1880] Troy J. Guy, Robert E. Gresswell, and Michael A. Banks. Landscapescale evaluation of genetic structure among barrier-isolated populations of coastal cutthroat trout, Oncorhynchus clarkii clarkii. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 65(8):1749–1762, August 2008. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/F08-090.

Minns:2008:PNA

[1881] Charles K. Minns, James E. Moore, Brian J. Shuter, and Nicholas E. Mandrak. A preliminary national analysis of some key characteristics of Canadian lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 65(8): 1763–1778, August 2008. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/F08-110.

Watson:2008:TOC

[1882] Susan B. Watson, Jeff Ridal, and Gregory L. Boyer. Taste and odour and cyanobacterial toxins: impairment, prediction, and management in the Great Lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 65(8): 1779–1796, August 2008. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/F08-084.

Stewart:2008:PDA

[1883] James E. Stewart. Postconsumption domoic acid generation by the diatom Pseudo-nitzschia multiseries as a factor in depuration models.

Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 65(9):1797–1799, September 2008. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/F08-136.

Rindorf:2008:DNS

[1884] Anna Rindorf and Bo Sølgaard Andersen. Do North Sea cod (Gadus morhua) fisheries maintain high catch rates at low stock size? Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 65(9):1800–1813, September 2008. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/F08-086.

Courbois:2008:EPS

[1885] Jean-Yves Courbois, Stephen L. Katz, Daniel J. Isaak, E. Ashley Steel, Russell F. Thurow, A. Michelle Wargo Rub, Tony Olsen, and Chris E. Jordan. Evaluating probability sampling strategies for estimating redd counts: an example with Chinook salmon (Oncorhynchus tshawytscha). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 65(9):1814–1830, September 2008. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/F08-092.

Jensen:2008:PBT

[1886] Hallvard Jensen, Kimmo K. Kahilainen, Per-Arne Amundsen, Karl Øystein Gjelland, Antti Tuomaala, Tommi Malinen, and Thomas Bøhn. Predation by brown trout (Salmo trutta) along a diversifying prey community gradient. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 65(9):1831– 1841, September 2008. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/F08-096.

Dorner:2008:HTP

[1887] Brigitte Dorner, Randall M. Peterman, and Steven L. Haeseker. Historical trends in productivity of 120 Pacific pink, chum, and sockeye salmon stocks reconstructed by using a Kalman filter. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 65(9):1842–1866, September 2008. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/F08-094.

Williams:2008:PLA

[1888] Rob Williams, Anna Hall, and Arliss Winship. Potential limits to anthropogenic mortality of small cetaceans in coastal waters of British Columbia. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 65(9):1867–1878, September 2008. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
F08-098.

Kimura:2008:EBG

[1889] Daniel K. Kimura. Extending the von Bertalanffy growth model using explanatory variables. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 65(9): 1879–1891, September 2008. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/F08-091.

Reid:2008:STI

[1890] Scott M. Reid, Chris C. Wilson, Leon M. Carl, and Troy G. Zorn. Species traits influence the genetic consequences of river fragmentation on two co-occurring redhorse (Moxostoma) species. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 65(9):1892–1904, September 2008. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/F08-093.

Strecker:2008:DLC

[1891] Angela L. Strecker, Rebecca Milne, and Shelley E. Arnott. Dispersal limitation and climate-related environmental gradients structure microcrustacean composition in freshwater lakes, Ellesmere Island, Canada. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 65(9):1905–1918, September 2008. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/F08-097.

Glover:2008:YPP

[1892] David C. Glover, John M. Dettmers, David H. Wahl, and David F. Clapp. Yellow perch (Perca flavescens) stock structure in Lake Michigan: an analysis using mark–recapture data. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 65(9):1919–1930, September 2008. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/F08-100.

McGarvey:2008:DSD

[1893] Richard McGarvey, Stephen Mayfield, Karen Byth, Thor Saunders, Rowan Chick, Brian Foureur, John E. Feenstra, Peter Preece, and Alan Jones. A diver survey design to estimate absolute density, biomass, and spatial distribution of abalone. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 65(9):1931–1944, September 2008. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/F08-101.

Spens:2008:SNL

[1894] Johan Spens and John P. Ball. Salmonid or nonsalmonid lakes: predicting the fate of northern boreal fish communities with hierarchical filters relating to a keystone piscivore. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 65(9):1945–1955, September 2008. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/F08-103.

Venter:2008:MUI

[1895] Oscar Venter, James W. A. Grant, Michelle V. Noël, and Jae woo Kim. Mechanisms underlying the increase in young-of-the-year Atlantic salmon (Salmo salar) density with habitat complexity. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 65(9):1956–1964, September 2008. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/F08-106.

Gaichas:2008:NME

[1896] Sarah K. Gaichas and Robert C. Francis. Network models for ecosystembased fishery analysis: a review of concepts and application to the Gulf of Alaska marine food web. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 65(9): 1965–1982, September 2008. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/F08-104.

McGoldrick:2008:DBS

[1897] Daryl J. McGoldrick, David R. Barton, Michael Power, Ryan W. Scott, and Barbara J. Butler. Dynamics of bacteria–substrate stable isotope separation: dependence on substrate availability and implications for aquatic food web studies. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 65(9):

1983–1990, September 2008. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/F08-109.

Punt:2008:QAR

[1898] André E. Punt, David C. Smith, Kyne KrusicGolub, and Simon Robertson. Quantifying age-reading error for use in fisheries stock assessments, with application to species in Australia’s southern and eastern scalefish and shark fishery. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 65(9):1991– 2005, September 2008. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/F08-111.

Lobon-Cervia:2008:HQE

[1899] Javier Lobón-Cerviá. Habitat quality enhances spatial variation in the self-thinning patterns of stream-resident brown trout (Salmo trutta). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 65(9):2006–2015, September 2008. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/F08-105.

Massicotte:2008:ISF

[1900] Rachel Massicotte, Pierre Magnan, and Bernard Angers. Intralacustrine site fidelity and nonrandom mating in the littoral-spawning northern redbelly dace (Phoxinus eos). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 65(9):2016–2025, September 2008. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/F08-114.

Olsen:2008:IHS

[1901] Jeffrey B. Olsen, Blair G. Flannery, Terry D. Beacham, Jeffrey F. Bromaghin, Penelope A. Crane, Charles F. Lean, Karen M. Dunmall, and John K. Wenburg. The influence of hydrographic structure and seasonal run timing on genetic diversity and isolation-by-distance in chum salmon (Oncorhynchus keta). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 65(9): 2026–2042, September 2008. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/F08-108.

Cremona:2008:AIM

[1902] Fabien Cremona, Dolors Planas, and Marc Lucotte. Assessing the importance of macroinvertebrate trophic dead ends in the lower transfer of methylmercury in littoral food webs. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 65(9):2043–2052, September 2008. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/F08-116.

Semmens:2008:ADF

[1903] Brice Xavier Semmens. Acoustically derived fine-scale behaviors of juvenile Chinook salmon (Oncorhynchus tshawytscha) associated with intertidal benthic habitats in an estuary. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 65(9):2053–2062, September 2008. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/F08-107.

Lacroix:2008:IOM

[1904] Gilles L. Lacroix. Influence of origin on migration and survival of Atlantic salmon (Salmo salar) in the Bay of Fundy, Canada. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 65(9):2063–2079, September 2008. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/F08-119.

Millidine:2008:UVF

[1905] K. J. Millidine, N. B. Metcalfe, and J. D. Armstrong. The use of ventilation frequency as an accurate indicator of metabolic rate in juvenile Atlantic salmon (Salmo salar). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 65(10):2081–2087, October 2008. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/F08-118.

Benoit:2008:IEC

[1906] Hugues P. Benoît and Douglas P. Swain. Impacts of environmental change and direct and indirect harvesting effects on the dynamics of a marine fish community. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 65(10): 2088–2104, October 2008. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/F08-112.

Ohlberger:2008:ESP

[1907] Jan Ohlberger, Thomas Mehner, Georg Staaks, and Franz Hölker. Is ecological segregation in a pair of sympatric coregonines supported by divergent feeding efficiencies? Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 65(10): 2105–2113, October 2008. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/F08-120.

Kiljunen:2008:CSI

[1908] Mikko Kiljunen, Heikki Peltonen, Roger I. Jones, Hannu Kiviranta, Pekka J. Vuorinen, Matti Verta, and Juha Karjalainen. Coupling stable isotopes with bioenergetics to evaluate sources of variation in organochlorine concentrations in Baltic salmon (Salmo salar). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 65(10):2114–2126, October 2008. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/F08-121.

Reveillac:2008:AML

[1909] Elodie Réveillac, Eric Feunteun, Patrick Berrebi, Pierre-Alexandre Gagnaire, Raymonde Lecomte-Finiger, Pierre Bosc, and Tony Robinet. Anguilla marmorata larval migration plasticity as revealed by otolith microstructural analysis. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 65(10): 2127–2137, October 2008. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/F08-122.

Bilby:2008:ECL

[1910] Robert E. Bilby and Lauren A. Mollot. Effect of changing land use patterns on the distribution of Coho salmon (Oncorhynchus kisutch) in the Puget Sound region. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 65(10): 2138–2148, October 2008. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/F08-113.

Huss:2008:IGH

[1911] Magnus Huss, Pär Byström, Åsa Strand, Lars-Ove Eriksson, and Lennart Persson. Influence of growth history on the accumulation of energy reserves and winter mortality in young fish. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 65(10):2149–2156, October 2008. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/F08-115.

Wang:2008:STV

[1912] Hui-Yu Wang, Tomas O. Höök, Mark P. Ebener, Lloyd C. Mohr, and Philip J. Schneeberger. Spatial and temporal variation of maturation schedules of lake whitefish (Coregonus clupeaformis) in the Great Lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 65(10):2157–2169, October 2008. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/F08-124.

Miller:2008:OCW

[1913] Silvana C. Miller, Shannon E. Reeb, Patricia A. Wright, and Todd E. Gillis. Oxygen concentration in the water boundary layer next to rainbow trout (Oncorhynchus mykiss) embryos is influenced by hypoxia exposure time, metabolic rate, and water flow. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 65(10):2170–2177, October 2008. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/F08-123.

Xie:2008:OAR

[1914] Yunbo Xie, Catherine G. J. Michielsens, Andrew P. Gray, Fiona J. Martens, and Jacqueline L. Boffey. Observations of avoidance reactions of migrating salmon to a mobile survey vessel in a riverine environment. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 65(10):2178–2190, October 2008. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/F08-128.

Solomon:2008:LTV

[1915] Christopher T. Solomon, Stephen R. Carpenter, James A. Rusak, and M. Jake Vander Zanden. Long-term variation in isotopic baselines and implications for estimating consumer trophic niches. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 65(10):2191–2200, October 2008. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/F08-125.

Jardine:2008:MTC

[1916] Timothy D. Jardine, Eric Chernoff, and R. Allen Curry. Maternal transfer of carbon and nitrogen to progeny of sea-run and resident brook trout (Salvelinus fontinalis). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 65(10): 2201–2210, October 2008. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/F08-132.

Schindler:2008:CEL

[1917] D. W. Schindler, Alexander P. Wolfe, Rolf Vinebrooke, Angela Crowe, Jules M. Blais, Brenda Miskimmin, Rina Freed, and Bianca Perren. The cultural eutrophication of Lac la Biche, Alberta, Canada: a paleoecological study. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 65(10):2211–2223, October 2008. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/F08-
117.

Rypel:2008:VAG

[1918] Andrew L. Rypel, Wendell R. Haag, and Robert H. Findlay. Validation of annual growth rings in freshwater mussel shells using cross dating. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 65(10):2224–2232, October 2008. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/F08-129.

Lu:2008:CTM

[1919] Yuehan Lu, Stuart A. Ludsin, David L. Fanslow, and Steven A. Pothoven. Comparison of three microquantity techniques for measuring total lipids in fish. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 65(10):2233–2241, October 2008. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/F08-
135.

Bond:2008:MSS

[1920] Morgan H. Bond, Sean A. Hayes, Chad V. Hanson, and R. Bruce Mac-Farlane. Marine survival of steelhead (Oncorhynchus mykiss) enhanced by a seasonally closed estuary. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 65(10):2242–2252, October 2008. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/F08-131.

Marta-Almeida:2008:MSN

[1921] Martinho Marta-Almeida, Jesus Dubert, Álvaro Peliz, Antonina dos Santos, and Henrique Queiroga. A modelling study of Norway lobster (Nephrops norvegicus) larval dispersal in southern Portugal: predictions of larval wastage and self-recruitment in the Algarve stock. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 65(10):2253–2268, October 2008. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/F08-138.

Lessard:2008:EGA

[1922] Robert B. Lessard, Ray Hilborn, and Brandon E. Chasco. Escapement goal analysis and stock reconstruction of sockeye salmon populations (Oncorhynchus nerka) using life-history models. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 65(10):2269–2278, October 2008. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/F08-130.

Zipkin:2008:ORS

[1923] Elise F. Zipkin, Patrick J. Sullivan, Evan G. Cooch, Clifford E. Kraft, Brian J. Shuter, and Brian C. Weidel. Overcompensatory response of a smallmouth bass (Micropterus dolomieu) population to harvest: release from competition? Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 65(10): 2279–2292, October 2008. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/F08-133.

Olsson:2008:IFQ

[1924] Karin Olsson, Per Nyström, Patrik Stenroth, Erika Nilsson, Marie Svensson, and Wilhelm Granéli. The influence of food quality and availability on trophic position, carbon signature, and growth rate of an omnivorous crayfish. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 65(10):2293–2304, October 2008. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/F08-

Shelton:2008:TSO

[1925] Peter A. Shelton and Alan F. Sinclair. It’s time to sharpen our definition of sustainable fisheries management. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 65(10):2305–2314, October 2008. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/F08-151.

Swain:2008:PED

[1926] Douglas P. Swain and Ghislain A. Chouinard. Predicted extirpation of the dominant demersal fish in a large marine ecosystem: Atlantic cod (Gadus morhua) in the southern Gulf of St. Lawrence. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 65(11):2315–2319, November 2008. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/F08-175.

Wang:2008:PRS

[1927] Xin Wang, David Depew, Sherry Schiff, and Ralph E. H. Smith. Photosynthesis, respiration, and stable isotopes of oxygen in a large oligotrophic lake (Lake Erie, USA–Canada). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 65(11):2320–2331, November 2008. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/F08-134.

Bates:2008:SSB

[1928] Amanda E. Bates. Size- and sex-based habitat partitioning by lepetodrilus fucensis near hydrothermal vents on the Juan de Fuca Ridge, Northeast Pacific. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 65(11):2332– 2341, November 2008. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/F08-139.

Doering-Arjes:2008:EPA

[1929] P. Doering-Arjes, M. Cardinale, and H. Mosegaard. Estimating population age structure using otolith morphometrics: a test with knownage Atlantic cod (Gadus morhua) individuals. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 65(11):2342–2350, November 2008. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/F08-143.

Stewart:2008:IPM

[1930] A. Robin Stewart, Michael K. Saiki, James S. Kuwabara, Charles N. Alpers, Mark Marvin-DiPasquale, and David P. Krabbenhoft. Influence of plankton mercury dynamics and trophic pathways on mercury concentrations of top predator fish of a mining-impacted reservoir. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 65(11):2351–2366, November 2008. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/F08-140.

Pedersen:2008:GNS

[1931] M. W. Pedersen, D. Righton, U. H. Thygesen, K. H. Andersen, and H. Madsen. Geolocation of North Sea cod (Gadus morhua) using hidden Markov models and behavioural switching. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 65(11):2367–2377, November 2008. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/F08-144.

Wissel:2008:CAC

[1932] Björn Wissel, Zoraida J. Quiñones-Rivera, and Brian Fry. Combined analyses of O_{2} and CO_{2} for studying the coupling of photosynthesis and respiration in aquatic systems. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 65(11):2378–2388, November 2008. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/F08-146.

Jiao:2008:MMS

[1933] Yan Jiao, Richard Neves, and Jess Jones. Models and model selection uncertainty in estimating growth rates of endangered freshwater mussel populations. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 65(11):2389– 2398, November 2008. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/F08-141.

Wells:2008:ETJ

[1934] R. J. David Wells, James H. Cowan, William F. Patterson, and Carl J. Walters. Effect of trawling on juvenile red snapper (Lutjanus campechanus) habitat selection and life history parameters. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 65(11):2399–2411, November 2008. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/F08-145.

Elmer:2008:PGS

[1935] Kathryn R. Elmer, Jeroen K. J. Van Houdt, Axel Meyer, and Filip A. M. Volckaert. Population genetic structure of North American burbot (Lota lota maculosa) across the Nearctic and at its contact zone with Eurasian burbot (Lota lota lota). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 65(11): 2412–2426, November 2008. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/F08-173.

Evans:2008:BHP

[1936] Karen Evans, Adam Langley, Naomi P. Clear, Peter Williams, Toby Patterson, John Sibert, John Hampton, and John S. Gunn. Behaviour and habitat preferences of bigeye tuna (Thunnus obesus) and their influence on longline fishery catches in the western Coral Sea. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 65(11):2427–2443, November 2008. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/F08-148.

Vermard:2008:DMB

[1937] Youen Vermard, Paul Marchal, Stéphanie Mahévas, and Olivier Thébaud. A dynamic model of the Bay of Biscay pelagic fleet simulating fishing trip choice: the response to the closure of the European anchovy (Engraulis encrasicolus) fishery in 2005. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 65(11):2444–2453, November 2008. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/F08-147.

Stets:2008:LZS

[1938] E. G. Stets and J. B. Cotner. Littoral zones as sources of biodegradable dissolved organic carbon in lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 65(11):2454–2460, November 2008. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/F08-142.

Gardner:2008:STS

[1939] Beth Gardner, Patrick J. Sullivan, Stephen J. Morreale, and Sheryan P. Epperly. Spatial and temporal statistical analysis of bycatch data: patterns of sea turtle bycatch in the North Atlantic. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 65(11):2461–2470, November 2008. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/F08-152.

Hansen:2008:RAA

[1940] Gretchen J. A. Hansen and Michael L. Jones. A rapid assessment approach to prioritizing streams for control of Great Lakes sea lampreys (Petromyzon marinus): a case study in adaptive management. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 65(11):2471–2484, November 2008. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/F08-153.

Bristow:2008:PBP

[1941] Corben E. Bristow, Antoine Morin, Ray H. Hesslein, and Cheryl L. Podemski. Phosphorus budget and productivity of an experimental lake during the initial three years of cage aquaculture. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 65(11):2485–2495, November 2008. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/F08-155.

Li:2008:BCP

[1942] L. Li, H. Høie, A. J. Geffen, E. Heegaard, J. Skadal, and A. Folkvord. Back-calculation of previous fish size using individually tagged and marked Atlantic cod (Gadus morhua). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 65(11):2496–2508, November 2008. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/F08-157.

McGilliard:2008:MNT

[1943] Carey R. McGilliard and Ray Hilborn. Modeling no-take marine reserves in regulated fisheries: assessing the role of larval dispersal. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 65(11):2509–2523, November 2008. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/F08-150.

Hunsicker:2008:EPT

[1944] Mary E. Hunsicker and Timothy E. Essington. Evaluating the potential for trophodynamic control of fish by the longfin inshore squid (Loligo pealeii) in the Northwest Atlantic Ocean. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 65(11):2524–2535, November 2008. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/F08-154.

Walters:2008:SPD

[1945] Carl J. Walters, Ray Hilborn, and Villy Christensen. Surplus production dynamics in declining and recovering fish populations. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 65(11):2536–2551, November 2008. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/F08-170.

Knoll:2008:CIZ

[1946] Lesley B. Knoll, Orlando Sarnelle, Stephen K. Hamilton, Carrie E. H. Kissman, Alan E. Wilson, Joan B. Rose, and Mechelle R. Morgan. Corrigendum: Invasive zebra mussels (Dreissena polymorpha) increase cyanobacterial toxin concentrations in low-nutrient lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 65(11):2552, November 2008. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/F08-182.

Anderson:2008:EKB

[1947] Sean C. Anderson, Heike K. Lotze, and Nancy L. Shackell. Evaluating the knowledge base for expanding low-trophic-level fisheries in Atlantic Canada. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 65(12):2553–2571, December 2008. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
F08-156.

Black:2008:EHA

[1948] Bryan A. Black, Darlene C. Gillespie, Shayne E. MacLellan, and Claudia M. Hand. Establishing highly accurate production-age data using the tree-ring technique of crossdating: a case study for Pacific geoduck (panopea abrupta). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 65(12):

2572–2578, December 2008. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/F08-158.

Dutil:2008:ACG

[1949] Jean-Denis Dutil, Coraline Jabouin, Richard Larocque, Gaston Desrosiers, and Pierre U. Blier. Atlantic cod (Gadus morhua) from cold and warm environments differ in their maximum growth capacity at low temperatures. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 65(12): 2579–2591, December 2008. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/F08-159.

Keefer:2008:ICR

[1950] Matthew L. Keefer, Robert H. Wertheimer, Allen F. Evans, Charles T. Boggs, and Christopher A. Peery. Iteroparity in Columbia River summerrun steelhead (Oncorhynchus mykiss): implications for conservation. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 65(12):2592–2605, December 2008. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/F08-160.

Gurney:2008:PBM

[1951] William S. C. Gurney, Philip J. Bacon, Graham Tyldesley, and Alan F. Youngson. Process-based modelling of decadal trends in growth, survival, and smolting of wild salmon (Salmo salar) parr in a Scottish upland stream. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 65(12):2606–2622, December 2008. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
F08-149.

Walther:2008:CSV

[1952] Benjamin D. Walther and Simon R. Thorrold. Continental-scale variation in otolith geochemistry of juvenile American shad (Alosa sapidissima). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 65(12):2623–2635, December 2008. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/F08-164.

Hirsch:2008:IBN

[1953] Philipp Emanuel Hirsch and Philipp Fischer. Interactions between native juvenile burbot (Lota lota) and the invasive spinycheek crayfish (Orconectes limosus) in a large European lake. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 65(12):2636–2643, December 2008. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/F08-162.

Kamarainen:2008:ZTP

[1954] Amy M. Kamarainen, Freya E. Rowland, Reinette Biggs, and Stephen R. Carpenter. Zooplankton and the total phosphorus–chlorophyll a relationship: hierarchical Bayesian analysis of measurement error. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 65(12):2644–2655, December 2008. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/F08-161.

Poos:2008:PFS

[1955] Mark S. Poos, Nicholas E. Mandrak, and Robert L. McLaughlin. A practical framework for selecting among single-species, community-, and ecosystem-based recovery plans. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 65(12):2656–2666, December 2008. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/F08-166.

Sullivan:2008:RSP

[1956] S. Mažeika P. Sullivan and Mary C. Watzin. Relating stream physical habitat condition and concordance of biotic productivity across multiple taxa. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 65(12):2667–2677, December 2008. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
F08-165.

Harma:2008:RAR

[1957] Meri Härmä, Antti Lappalainen, and Lauri Urho. Reproduction areas of roach (Rutilus rutilus) in the northern Baltic Sea: potential effects of climate change. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 65(12):2678– 2688, December 2008. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/F08-167.

Gelinas:2008:RCZ

[1958] Malorie Gélinas and Bernadette Pinel-Alloul. Relating crustacean zooplankton community structure to residential development and landcover disturbance near Canadian Shield lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 65(12):2689–2702, December 2008. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/F08-163.

Rothaus:2008:PDP

[1959] Don P. Rothaus, Brent Vadopalas, and Carolyn S. Friedman. Precipitous declines in pinto abalone (Haliotis kamtschatkana kamtschatkana) abundance in the San Juan Archipelago, Washington, USA, despite statewide fishery closure. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 65(12):2703– 2711, December 2008. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/F08-168.

Price:2008:GRM

[1960] M. H. H. Price, C. T. Darimont, N. F. Temple, and S. M. MacDuffee. Ghost runs: management and status assessment of Pacific salmon (Oncorhynchus spp.) returning to British Columbia’s central and north coasts. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 65(12):2712–2718, December 2008. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
F08-174.

Perga:2008:TCC

[1961] Marie-Elodie Perga, Martin Kainz, and Asit Mazumder. Terrestrial carbon contribution to lake food webs: could the classical stable isotope approach be misleading? Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 65(12): 2719–2727, December 2008. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/F08-172.

Eero:2008:RHS

[1962] Margit Eero, Friedrich W. Köster, and Brian R. MacKenzie. Reconstructing historical stock development of Atlantic cod (Gadus morhua) in the eastern Baltic Sea before the beginning of intensive exploitation. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 65(12):2728–2741, December 2008. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/F08-176.

Munch:2008:BAI

[1963] S. B. Munch and L. M. Clarke. A Bayesian approach to identifying mixtures from otolith chemistry data. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 65(12):2742–2751, December 2008. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/F08-169.

Kalinowski:2008:STS

[1964] Steven T. Kalinowski, Michael H. Meeuwig, Shawn R. Narum, and Mark L. Taper. Stream trees: a statistical method for mapping genetic differences between populations of freshwater organisms to the sections of streams that connect them. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 65(12):2752–2760, December 2008. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/F08-171.

Rogers:2008:PHA

[1965] Tracy Noel Rogers and David R. Chalcraft. Pond hydroperiod alters the effect of density-dependent processes on larval anurans. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 65(12):2761–2768, December 2008. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/F08-177.

Rudstam:2008:ACM

[1966] Lars G. Rudstam, Frank R. Knudsen, Helge Balk, Gideon Gal, Brent T. Boscarino, and Thomas Axenrot. Acoustic characterization of Mysis relicta at multiple frequencies. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 65 (12):2769–2779, December 2008. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/

doi/full/10.1139/F08-179.

Blank:2008:MMA

[1967] Miriam Blank, Karl Jürss, and Ralf Bastrop. A mitochondrial multigene approach contributing to the systematics of the brook and river lampreys and the phylogenetic position of eudontomyzon mariae. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 65(12):2780–2790, December 2008. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/F08-178.

Nilsen:2008:TSH

[1968] Marianne Nilsen, Torstein Pedersen, Einar Magnus Nilssen, and Stein Fredriksen. Trophic studies in a high-latitude fjord ecosystem — a comparison of stable isotope analyses (δ ^{13}C and δ ^{15}N) and trophic-level estimates from a mass-balance model. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 65(12):2791–2806, December 2008. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/F08-180.

Morris:2008:PRE

[1969] Matthew R. J. Morris, Dylan J. Fraser, Anthony J. Heggelin, Frederick G. Whoriskey, Jonathan W. Carr, Shane F. O’Neil, and Jeffrey A. Hutchings. Prevalence and recurrence of escaped farmed Atlantic salmon (Salmo salar) in eastern North American rivers. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 65(12):2807–2826, December 2008. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/F08-181.

Church:2009:MNT

[1970] M. Robbins Church, Joseph L. Ebersole, Kirk M. Rensmeyer, Ryan B. Couture, Frederic T. Barrows, and David L. G. Noakes. Mucus: a new tissue fraction for rapid determination of fish diet switching using stable isotope analysis. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 66(1):1–5, January 2009. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/F08-
206.

Baba:2009:RBS

[1971] Katsuhisa Baba, Rieko Sugawara, Hisashi Nitta, Kiyoshi Endou, and Akira Miyazono. Relationship between spat density, food availability, and growth of spawners in cultured mizuhopecten yessoensis in Funka Bay: concurrence with El Niño Southern Oscillation. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 66(1):6–17, January 2009. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/F08-183.

Wilson:2009:EMC

[1972] Sarah J. Wilson and Anthony Ricciardi. Epiphytic macroinvertebrate communities on Eurasian watermilfoil (Myriophyllum spicatum) and native milfoils Myriophyllum sibericum and Myriophyllum alterniflorum in eastern North America. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 66 (1):18–30, January 2009. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/F08-187.

Breen:2009:MMS

[1973] Matthew J. Breen, Carl R. Ruetz, Kurt J. Thompson, and Steven L. Kohler. Movements of mottled sculpins (Cottus bairdii) in a Michigan stream: how restricted are they? Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 66(1):31–41, January 2009. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/F08-189.

Nagayama:2009:SMP

[1974] Shigeya Nagayama, Yôichi Kawaguchi, Daisuke Nakano, and Futoshi Nakamura. Summer microhabitat partitioning by different size classes of masu salmon (Oncorhynchus masou) in habitats formed by installed large wood in a large lowland river. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 66 (1):42–51, January 2009. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/F08-191.

Rempel:2009:PER

[1975] Laura L. Rempel and Michael Church. Physical and ecological response to disturbance by gravel mining in a large alluvial river. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 66(1):52–71, January 2009. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/F08-184.

Carlstrom:2009:STI

[1976] Julia Carlström, Per Berggren, and Nick J. C. Tregenza. Spatial and temporal impact of pingers on porpoises. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 66(1):72–82, January 2009. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/F08-186.

Darwish:2009:GVR

[1977] Tamara L. Darwish and Jeffrey A. Hutchings. Genetic variability in reaction norms between farmed and wild backcrosses of Atlantic salmon (Salmo salar). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 66(1):83–90, January 2009. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
F08-185.

Fowler:2009:CHS

[1978] Stephanie L. Fowler, Diana Hamilton, and Suzanne Currie. A comparison of the heat shock response in juvenile and adult rainbow trout (Oncorhynchus mykiss) — implications for increased thermal sensitivity with age. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 66(1):91–100, January 2009. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/F08-192.

Boscarino:2009:PVD

[1979] Brent T. Boscarino, Lars G. Rudstam, Ellis R. Loew, and Edward L. Mills. Predicting the vertical distribution of the opossum shrimp, Mysis relicta, in Lake Ontario: a test of laboratory-based light preferences. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 66(1):101–113, January 2009. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/F08-190.

Zhao:2009:BML

[1980] Yingming Zhao, Michael L. Jones, Brian J. Shuter, and Edward F. Roseman. A biophysical model of Lake Erie walleye (Sander vitreus) explains interannual variations in recruitment. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 66(1):114–125, January 2009. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/F08-188.

Branch:2009:DPC

[1981] Trevor A. Branch. Differences in predicted catch composition between two widely used catch equation formulations. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 66(1):126–132, January 2009. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/F08-196.

Seiler:2009:CBN

[1982] Steven M. Seiler and Ernest R. Keeley. Competition between native and introduced salmonid fishes: cutthroat trout have lower growth rate in the presence of cutthroat–rainbow trout hybrids. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 66(1):133–141, January 2009. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/F08-194.

Sherwood:2009:AGR

[1983] Owen A. Sherwood and Evan N. Edinger. Ages and growth rates of some deep-sea gorgonian and antipatharian corals of Newfoundland and Labrador. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 66(1):142–152, January 2009. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/F08-
195.

Cunningham:2009:GID

[1984] Kathryn Maja Cunningham, Michael Francis Canino, Ingrid Brigette Spies, and Lorenz Hauser. Genetic isolation by distance and localized fjord population structure in Pacific cod (Gadus macrocephalus): limited effective dispersal in the northeastern Pacific Ocean. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 66(1):153–166, January 2009. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/F08-199.

Walter:2009:GFI

[1985] Ryan P. Walter, Tutku Aykanat, David W. Kelly, J. Mark Shrimpton, and Daniel D. Heath. Gene flow increases temporal stability of Chinook salmon (Oncorhynchus tshawytscha) populations in the Upper Fraser River, British Columbia, Canada. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 66(2):167–176, February 2009. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub. com/doi/full/10.1139/F08-201. See comment [2160] and reply [2161].

Dalton:2009:IDC

[1986] Christopher M. Dalton, David Ellis, and David M. Post. The impact of double-crested cormorant (Phalacrocorax auritus) predation on anadromous alewife (Alosa pseudoharengus) in south-central Connecticut, USA. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 66(2):177–186, February 2009. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/F08-198.

Mills:2009:FIA

[1987] R. Brad Mills, Andrew M. Paterson, Jules M. Blais, David R. S. Lean, John P. Smol, and Greg Mierle. Factors influencing the achievement of steady state in mercury contamination among lakes and catchments of south-central Ontario. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 66 (2):187–200, February 2009. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/F08-204.

Litzow:2009:FTA

[1988] Michael A. Litzow and Daniel Urban. Fishing through (and up) Alaskan food webs. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 66(2):201–211, February 2009. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
F08-207.

Kaufman:2009:IRA

[1989] Scott D. Kaufman, Ed Snucins, John M. Gunn, and Wayne Selinger. Impacts of road access on lake trout (Salvelinus namaycush) populations: regional scale effects of overexploitation and the introduction of smallmouth bass (Micropterus dolomieu). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 66(2):212–223, February 2009. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/F08-205.

Palmer:2009:PAN

[1990] Miquel Palmer, Antoni Quetglas, Beatriz Guijarro, Joan Moranta, Francesc Ordines, and Enric Massutí. Performance of artificial neural networks and discriminant analysis in predicting fishing tactics from multispecific fisheries. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 66(2):

224–237, February 2009. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/F08-208.

Rich:2009:CIC

[1991] Harry B. Rich, Thomas P. Quinn, Mark D. Scheuerell, and Daniel E. Schindler. Climate and intraspecific competition control the growth and life history of juvenile sockeye salmon (Oncorhynchus nerka) in Iliamna Lake, Alaska. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 66(2):238–246, February 2009. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
F08-210.

Ogbebo:2009:NLP

[1992] Fortune E. Ogbebo, Marlene S. Evans, Marley J. Waiser, Vijay P. Tumber, and Jonathan J. Keating. Nutrient limitation of phytoplankton growth in Arctic lakes of the lower Mackenzie River Basin, northern Canada. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 66(2):247–260, February 2009. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/F08-202.

Bailey:2009:EEP

[1993] Sarah A. Bailey, Luis Antonio Vélez-Espino, Ora E. Johannsson, Marten A. Koops, and Chris J. Wiley. Estimating establishment probabilities of cladocera introduced at low density: an evaluation of the proposed ballast water discharge standards. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 66(2):261–276, February 2009. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/F08-200.

Fischer:2009:ESE

[1994] Jesse R. Fischer and Craig P. Paukert. Effects of sampling effort, assemblage similarity, and habitat heterogeneity on estimates of species richness and relative abundance of stream fishes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 66(2):277–290, February 2009. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/F08-209.

Marchal:2009:RWT

[1995] Paul Marchal, Philippe Lallemand, and Kevin Stokes. The relative weight of traditions, economics, and catch plans in New Zealand fleet dynamics. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 66(2):291–311, February 2009. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/F08-193.

Garvey:2009:STS

[1996] James E. Garvey, Russell A. Wright, and Elizabeth A. Marschall. Searching for threshold shifts in spawner–recruit data. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 66(2):312–320, February 2009. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/F08-212.

Kurota:2009:SBM

[1997] Hiroyuki Kurota, Murdoch K. McAllister, Gareth L. Lawson, Jacob I. Nogueira, Steven L. H. Teo, and Barbara A. Block. A sequential Bayesian methodology to estimate movement and exploitation rates using electronic and conventional tag data: application to Atlantic bluefin tuna (Thunnus thynnus). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 66(2): 321–342, February 2009. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/F08-197.

Hamilton:2009:BCP

[1998] Stephen K. Hamilton, Denise A. Bruesewitz, Geoffrey P. Horst, David B. Weed, and Orlando Sarnelle. Biogenic calcite–phosphorus precipitation as a negative feedback to lake eutrophication. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 66(2):343–350, February 2009. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/F09-003.

Taylor:2009:UEC

[1999] Ian G. Taylor and Vincent F. Gallucci. Unconfounding the effects of climate and density dependence using 60 years of data on spiny dogfish (Squalus acanthias). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 66 (3):351–366, March 2009. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/F08-211.

Lee:2009:EPG

[2000] Yong-Woo Lee, Bernard A. Megrey, and S. Allen Macklin. Evaluating the performance of Gulf of Alaska walleye pollock (Theragra chalcogramma) recruitment forecasting models using a Monte Carlo resampling strategy. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 66(3):367–381, March 2009. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/F08-203.

VanDeHey:2009:GSL

[2001] Justin A. VanDeHey, Brian L. Sloss, Paul J. Peeters, and Trent M. Sutton. Genetic structure of lake whitefish (Coregonus clupeaformis) in Lake Michigan. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 66(3):382–393, March 2009. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
F08-213.

McCormick:2009:IEA

[2002] Stephen D. McCormick, Amanda Keyes, Keith H. Nislow, and Michelle Y. Monette. Impacts of episodic acidification on in-stream survival and physiological impairment of Atlantic salmon (Salmo salar) smolts. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 66(3):394–403, March 2009. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/F09-002.

Morgan:2009:ERP

[2003] M. J. Morgan, H. Murua, G. Kraus, Y. Lambert, G. Marteinsdóttir, C. T. Marshall, L. O’Brien, and J. Tomkiewicz. The evaluation of reference points and stock productivity in the context of alternative indices of stock reproductive potential. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 66 (3):404–414, March 2009. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/F09-009.

Brady:2009:BRJ

[2004] Damian C. Brady, Timothy E. Targett, and Danielle M. Tuzzolino. Behavioral responses of juvenile weakfish (Cynoscion regalis) to diel-cycling hypoxia: swimming speed, angular correlation, expected displacement, and effects of hypoxia acclimation. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 66(3):415–424, March 2009. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/F09-007.

Handegard:2009:SVT

[2005] Nils Olav Handegard and Dag Tjøstheim. The sampling volume of trawl and acoustics: estimating availability probabilities from observations of tracked individual fish. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 66 (3):425–437, March 2009. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/F09-004.

Walters:2009:AAA

[2006] Annika W. Walters, Rebecca T. Barnes, and David M. Post. Anadromous alewives (Alosa pseudoharengus) contribute marine-derived nutrients to coastal stream food webs. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 66 (3):439–448, March 2009. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/F09-008.

Young:2009:EDD

[2007] Joelle D. Young, Ellis R. Loew, and Norman D. Yan. Examination of direct daytime predation by Coregonus artedi on Bythotrephes longimanus in Harp Lake, Ontario, Canada: no evidence for the refuge hypothesis. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 66(3):449–459, March 2009. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/F09-006.

Geraldi:2009:HEA

[2008] Nathan R. Geraldi, Richard A. Wahle, and Michael Dunnington. Habitat effects on American lobster (Homarus americanus) movement and density: insights from georeferenced trap arrays, seabed mapping, and tagging. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 66(3):460–470, March 2009. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/F09-011.

Horton:2009:ASS

[2009] Gregg E. Horton, Benjamin H. Letcher, Michael M. Bailey, and Michael T. Kinnison. Atlantic salmon (Salmo salar) smolt production: the relative importance of survival and body growth. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 66(3):471–483, March 2009. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/F09-005.

Gevrey:2009:MIL

[2010] Muriel Gevrey, Frédéric Sans-Piché, Gaël Grenouillet, Loïc Tudesque, and Sovan Lek. Modeling the impact of landscape types on the distribution of stream fish species. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 66 (3):484–495, March 2009. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/F09-001.

Wood:2009:SMF

[2011] Robert J. Wood and Herbert M. Austin. Synchronous multidecadal fish recruitment patterns in Chesapeake Bay, USA. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 66(3):496–508, March 2009. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/F09-013.

Sourisseau:2009:CKD

[2012] M. Sourisseau, Y. Simard, and F. J. Saucier. Corrigendum: Krill diel vertical migration fine dynamics, nocturnal overturns, and their roles for aggregation in stratified flows. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 66(3):509, March 2009. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/F09-017.

Jacobson:2009:CEB

[2013] Larry D. Jacobson, Steven J. Bograd, Richard H. Parrish, Roy Mendelssohn, and Franklin B. Schwing. Corrigendum: an ecosystembased hypothesis for climatic effects on surplus production in California sardine (Sardinops sagax) and environmentally dependent surplus production models. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 66(3):510–511, March 2009. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
F09-018.

Gibson-Reinemer:2009:ESO

[2014] Daniel K. Gibson-Reinemer, Brett M. Johnson, Patrick J. Martinez, Dana L. Winkelman, Alan E. Koenig, and Jon D. Woodhead. Elemental signatures in otoliths of hatchery rainbow trout (Oncorhynchus mykiss): distinctiveness and utility for detecting origins and movement. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 66(4):513–524, April 2009. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/F09-015.

Gjelland:2009:PVM

[2015] Karl Øystein Gjelland, Thomas Bøhn, John K. Horne, Ingrid Jensvoll, Frank Reier Knudsen, and Per-Arne Amundsen. Planktivore vertical migration and shoaling under a subarctic light regime. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 66(4):525–539, April 2009. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/F09-014.

Palm:2009:PSS

[2016] Daniel Palm, Eva Brännäs, and Kjell Nilsson. Predicting site-specific overwintering of juvenile brown trout (Salmo trutta) using a habitat suitability index. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 66(4):540–546, April 2009. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/F09-
016.

Wahl:2009:INR

[2017] Bernd Wahl and Herbert Löffler. Influences on the natural reproduction of whitefish (Coregonus lavaretus) in Lake Constance. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 66(4):547–556, April 2009. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/F09-019.

Lucero:2009:MSR

[2018] Yasmin Lucero. A multivariate stock–recruitment function for cohorts with sympatric subclasses: application to maternal effects in rockfish (genus Sebastes). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 66(4):557–564, April 2009. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/F09-
012.

Beaudreau:2009:DNF

[2019] Anne H. Beaudreau and Timothy E. Essington. Development of a new field-based approach for estimating consumption rates of fishes and comparison with a bioenergetics model for lingcod (Ophiodon elongatus). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 66(4):565–578, April 2009. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/F09-021.

Martin:2009:HNO

[2020] Christopher H. Martin and Martin J. Genner. High niche overlap between two successfully coexisting pairs of Lake Malawi cichlid fishes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 66(4):579–588, April 2009. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/F09-023.

Jyvasjarvi:2009:NVP

[2021] Jussi Jyväsjärvi, Kimmo T. Tolonen, and Heikki Hämäläinen. Natural variation of profundal macroinvertebrate communities in boreal lakes is related to lake morphometry: implications for bioassessment. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 66(4):589–601, April 2009. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/F09-025.

Kerr:2009:BTU

[2022] Lisa A. Kerr and David H. Secor. Bioenergetic trajectories underlying partial migration in Patuxent River (Chesapeake Bay) white perch (Morone americana). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 66(4):602–612, April 2009. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/F09-
027.

Berger:2009:FIC

[2023] Aaron M. Berger and Robert E. Gresswell. Factors influencing coastal cutthroat trout (Oncorhynchus clarkii clarkii) seasonal survival rates: a spatially continuous approach within stream networks. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 66(4):613–632, April 2009. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/F09-029.

Kempf:2009:RNS

[2024] Alexander Kempf, Jens Floeter, and Axel Temming. Recruitment of North Sea cod (Gadus morhua) and Norway pout (Trisopterus esmarkii) between 1992 and 2006: the interplay between climate influence and predation. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 66(4):633–648, April 2009. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/F09-022.

Al-Chokhachy:2009:DDA

[2025] Robert Al-Chokhachy, Phaedra Budy, and Mary Conner. Detecting declines in the abundance of a bull trout (Salvelinus confluentus) population: understanding the accuracy, precision, and costs of our efforts. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 66(4):649–658, April 2009. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/F09-026.

denHeyer:2009:DAL

[2026] Cornelia Elizabeth den Heyer, E. Michael P. Chadwick, and Jeffrey A. Hutchings. Diffusion of American lobster (Homarus americanus) in Northumberland Strait, Canada. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 66(4):659–671, April 2009. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/

doi/full/10.1139/F09-024.

Sepulveda:2009:UNS

[2027] Adam J. Sepulveda, Warren T. Colyer, Winsor H. Lowe, and Mark R. Vinson. Using nitrogen stable isotopes to detect long-distance movement in a threatened cutthroat trout (Oncorhynchus clarkii utah). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 66(4):672–682, April 2009. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/F09-020.

Temming:2009:GME

[2028] Axel Temming and Jens P. Herrmann. A generic model to estimate food consumption: linking von Bertalanffy’s growth model with Beverton and Holt’s and Ivlev’s concepts of net conversion efficiency. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 66(4):683–700, April 2009. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/F09-028.

Neff:2009:CCA

[2029] Margaret R. Neff and Donald A. Jackson. Communication and cohesion in aquatic science literature. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 66 (4):701–712, April 2009. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/F09-031.

Smith:2009:BAM

[2030] Cindy J. Smith, Bret S. Danilowicz, and Wim G. Meijer. Bacteria associated with the mucus layer of Merlangius merlangus (whiting) as biological tags to determine harvest location. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 66(5):713–716, May 2009. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/F09-035.

Tonteri:2009:MRC

[2031] Anni Tonteri, Alexei Je. Veselov, Alexander V. Zubchenko, Jaakko Lumme, and Craig R. Primmer. Microsatellites reveal clear genetic boundaries among Atlantic salmon (Salmo salar) populations from the Barents and White Seas, northwest Russia. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 66(5):717–735, May 2009. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/F09-010.

Welch:2009:FMM

[2032] David W. Welch, Michael C. Melnychuk, Erin R. Rechisky, Aswea D. Porter, Melinda C. Jacobs, Adrian Ladouceur, R. Scott McKinley, and George D. Jackson. Freshwater and marine migration and survival of endangered Cultus Lake sockeye salmon (Oncorhynchus nerka) smolts using POST, a large-scale acoustic telemetry array. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 66(5):736–750, May 2009. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/F09-032.

Hart:2009:VAS

[2033] Deborah R. Hart and Antonie S. Chute. Verification of Atlantic sea scallop (Placopecten magellanicus) shell growth rings by tracking cohorts in fishery closed areas. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 66 (5):751–758, May 2009. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/F09-033.

Hatun:2009:NAS

[2034] Hjálmar Hátún, Mark R. Payne, and Jan Arge Jacobsen. The North Atlantic subpolar gyre regulates the spawning distribution of blue whiting (Micromesistius poutassou). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 66 (5):759–770, May 2009. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/F09-037.

Okamura:2009:NSM

[2035] Hiroshi Okamura and Yasuko Semba. A novel statistical method for validating the periodicity of vertebral growth band formation in elasmobranch fishes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 66(5):771–780, May 2009. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/F09-
039.

Berejikian:2009:RBR

[2036] Barry A. Berejikian, Donald M. Van Doornik, Julie A. Scheurer, and Richard Bush. Reproductive behavior and relative reproductive success of natural- and hatchery-origin Hood canal summer chum salmon (Oncorhynchus keta). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 66(5):781–789, May 2009. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/F09-
041.

Scharf:2009:CPR

[2037] Frederick S. Scharf, Jeffrey A. Buckel, and Francis Juanes. Contrasting patterns of resource utilization between juvenile estuarine predators: the influence of relative prey size and foraging ability on the ontogeny of piscivory. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 66(5):790–801, May 2009. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/F09-030.

Justice:2009:EDS

[2038] Casey Justice, Brian J. Pyper, Raymond C. P. Beamesderfer, Vaughn L. Paragamian, Pete J. Rust, Matthew D. Neufeld, and Susan C. Ireland. Evidence of density- and size-dependent mortality in hatchery-reared juvenile white sturgeon (Acipenser transmontanus) in the Kootenai River. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 66(5):802–815, May 2009. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/F09-034.

Barbiero:2009:RSC

[2039] Richard P. Barbiero, Mary Balcer, David C. Rockwell, and Marc L. Tuchman. Recent shifts in the crustacean zooplankton community of Lake Huron. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 66(5):816–828, May 2009. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/F09-036.

Piet:2009:ISF

[2040] G. J. Piet and F. J. Quirijns. The importance of scale for fishing impact estimations. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 66(5):829–835, May 2009. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/F09-042.

Goto:2009:IPP

[2041] Daisuke Goto and William G. Wallace. Influences of prey- and predatordependent processes on cadmium and methylmercury trophic transfer to mummichogs (Fundulus heteroclitus). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 66(5):836–846, May 2009. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/F09-038.

Salberg:2009:ESP

[2042] Arnt-Børre Salberg, Tor Arne Øigård, Garry B. Stenson, Tore Haug, and Kjell T. Nilssen. Estimation of seal pup production from aerial surveys using generalized additive models. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 66(5):847–858, May 2009. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/F09-040.

Hannesson:2009:EEC

[2043] Rögnvaldur Hannesson, Samuel Herrick, and John Field. Ecological and economic considerations in the conservation and management of the Pacific sardine (Sardinops sagax). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 66(5):859–868, May 2009. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/F09-045.

Payne:2009:CDF

[2044] Mark R. Payne, Mark Dickey-Collas, and Richard D. M. Nash. Comment on “Does the fall phytoplankton bloom control recruitment of Georges Bank haddock, Melanogrammus aeglefinus, through parental condition?”. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 66(5):869–872, May 2009. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/F09-
043.

Friedland:2009:RCP

[2045] Kevin D. Friedland, Jonathan A. Hare, Grayson B. Wood, Laurel A. Col, Lawrence J. Buckley, David G. Mountain, Joseph Kane, Jon Brodziak, R. Gregory Lough, and Cynthia H. Pilskaln. Reply to the comment by Payne et al. on “Does the fall phytoplankton bloom control recruitment of Georges Bank haddock, Melanogrammus aeglefinus, through parental condition?”. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 66(5):873–877, May 2009. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/F09-
044.

Byappanahalli:2009:CBT

[2046] Muruleedhara N. Byappanahalli and Richard L. Whitman. Clostridium botulinum type E occurs and grows in the alga Cladophora glomerata. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 66(6):879–882, June 2009. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/F09-052.

Behrens:2009:ODI

[2047] Jane W. Behrens, Gunni Ærtebjerg, Jens Kjerulf Petersen, and Jacob Carstensen. Oxygen deficiency impacts on burying habitats for lesser sandeel, Ammodytes tobianus, in the inner Danish waters. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 66(6):883–895, June 2009. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/F09-046.

Kendall:2009:EPS

[2048] Neala W. Kendall and Thomas P. Quinn. Effects of population-specific variation in age and length on fishery selection and exploitation rates of sockeye salmon (Oncorhynchus nerka). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 66(6):896–908, June 2009. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/F09-047.

Ruppert:2009:ACG

[2049] Jonathan L. W. Ruppert, Marie-Josée Fortin, George A. Rose, and Rodolphe Devillers. Atlantic cod (Gadus morhua) distribution response to environmental variability in the northern Gulf of St. Lawrence. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 66(6):909–918, June 2009. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/F09-049.

Paterson:2009:DLT

[2050] Gordon Paterson, D. Michael Whittle, Kenneth G. Drouillard, and G. Douglas Haffner. Declining lake trout (Salvelinus namaycush) energy density: are there too many salmonid predators in the Great Lakes? Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 66(6):919–932, June 2009. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/F09-048.

Bouvier:2009:ACF

[2051] Lynn D. Bouvier, Karl Cottenie, and Susan E. Doka. Aquatic connectivity and fish metacommunities in wetlands of the lower Great Lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 66(6):933–948, June 2009. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/F09-050.

Brimble:2009:BTE

[2052] Samantha K. Brimble, Jules M. Blais, Linda E. Kimpe, Mark L. Mallory, Bronwyn E. Keatley, Marianne S. V. Douglas, and John P. Smol. Bioenrichment of trace elements in a series of ponds near a northern fulmar (Fulmarus glacialis) colony at Cape Vera, Devon Island. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 66(6):949–958, June 2009. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/F09-053.

Lukey:2009:CCS

[2053] James R. Lukey and Stephen S. Crawford. Consistency of COSEWIC species at risk designations: freshwater fishes as a case study. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 66(6):959–971, June 2009. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/F09-054.

Takahashi:2009:CRL

[2054] Motomitsu Takahashi, Yoshiro Watanabe, Akihiko Yatsu, and Hiroshi Nishida. Contrasting responses in larval and juvenile growth to a climate–ocean regime shift between anchovy and sardine. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 66(6):972–982, June 2009. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/F09-051.

Kotwicki:2009:ELI

[2055] Stan Kotwicki, Alex De Robertis, Paul von Szalay, and Rick Towler. The effect of light intensity on the availability of walleye pollock (Theragra chalcogramma) to bottom trawl and acoustic surveys. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 66(6):983–994, June 2009. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/F09-055.

Marliave:2009:MER

[2056] Jeff Marliave and Wendell Challenger. Monitoring and evaluating rockfish conservation areas in British Columbia. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 66(6):995–1006, June 2009. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/F09-056.

Zimmerman:2009:ONS

[2057] Mara S. Zimmerman, Stephanie N. Schmidt, Charles C. Krueger, M. Jake Vander Zanden, and Randy L. Eshenroder. Ontogenetic niche shifts and resource partitioning of lake trout morphotypes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 66(6):1007–1018, June 2009. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/F09-060.

Dancik:2009:AC

[2058] Bruce Dancik. Announcement / Communiqué. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 66(7), July 2009. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f09-902.

Jackson:2009:AC

[2059] Don Jackson. Announcement / Communiqué. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 66(7), July 2009. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f09-903.

Rechisky:2009:EMH

[2060] Erin L. Rechisky, David W. Welch, Aswea D. Porter, Melinda C. Jacobs, and Adrian Ladouceur. Experimental measurement of hydrosysteminduced delayed mortality in juvenile Snake River spring Chinook salmon (Oncorhynchus tshawytscha) using a large-scale acoustic array. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 66(7):1019–1024, July 2009. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/F09-078.

Reinbold:2009:RSP

[2061] Drew Reinbold, Gary H. Thorgaard, and Patrick A. Carter. Reduced swimming performance and increased growth in domesticated rainbow trout, Oncorhynchus mykiss. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 66 (7):1025–1032, July 2009. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/F09-064.

Stierhoff:2009:HIG

[2062] Kevin L. Stierhoff, Timothy E. Targett, and James H. Power. Hypoxiainduced growth limitation of juvenile fishes in an estuarine nursery: assessment of small-scale temporal dynamics using RNA:DNA. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 66(7):1033–1047, July 2009. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/F09-066.

Forget:2009:IDC

[2063] Marie-Hélène Forget, Richard Carignan, and Christiane Hudon. Influence of diel cycles of respiration, chlorophyll, and photosynthetic parameters on the summer metabolic balance of temperate lakes and rivers. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 66(7):1048–1058, July 2009. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/F09-058.

Claessen:2009:BOS

[2064] David Claessen, Anneke S. de Vos, and André M. de Roos. Bioenergetics, overcompensation, and the source–sink status of marine reserves. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 66(7):1059–1071, July 2009. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/F09-061.

Utz:2009:DDI

[2065] Ryan M. Utz and Kyle J. Hartman. Density-dependent individual growth and size dynamics of central Appalachian brook trout (Salvelinus fontinalis). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 66(7):1072–1080, July 2009. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/F09-063.

Johnson:2009:ILU

[2066] Laura T. Johnson, Jennifer L. Tank, and Walter K. Dodds. The influence of land use on stream biofilm nutrient limitation across eight North American ecoregions. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 66 (7):1081–1094, July 2009. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/F09-065.

Gollock:2009:PMM

[2067] Matthew J. Gollock, Kristopher J. Hunter, Douglas A. Syme, Marcus Freeman, R. Scott McKinley, and A. Kurt Gamperl. Potential methods for measuring the activity patterns and energy use of Atlantic cod (Gadus morhua). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 66(7):1095–1106, July 2009. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/F09-062.

Mackinson:2009:MFE

[2068] Steven Mackinson, Barrie Deas, Doug Beveridge, and John Casey. Mixed-fishery or ecosystem conundrum? Multispecies considerations inform thinking on long-term management of North Sea demersal stocks. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 66(7):1107–1129, July 2009. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/F09-057.

Andrews:2009:LRD

[2069] Allen H. Andrews, Dianne M. Tracey, and Matthew R. Dunn. Lead– radium dating of orange roughy (Hoplostethus atlanticus): validation of a centenarian life span. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 66 (7):1130–1140, July 2009. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/F09-059.

Hill:2009:RSL

[2070] Aaron C. Hill, Jack A. Stanford, and Peter R. Leavitt. Recent sedimentary legacy of sockeye salmon (Oncorhynchus nerka) and climate change in an ultraoligotrophic, glacially turbid British Columbia nursery lake. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 66(7):1141–1152, July 2009. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/F09-067.

Muhlfeld:2009:STS

[2071] Clint C. Muhlfeld, Thomas E. McMahon, Durae Belcer, and Jeffrey L. Kershner. Spatial and temporal spawning dynamics of native westslope cutthroat trout, Oncorhynchus clarkii lewisi, introduced rainbow trout, Oncorhynchus mykiss, and their hybrids. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 66(7):1153–1168, July 2009. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/F09-073.

Ye:2009:HRA

[2072] Yimin Ye and Darren Dennis. How reliable are the abundance indices derived from commercial catch–effort standardization? Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 66(7):1169–1178, July 2009. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/F09-070.

Robinson:2009:BTO

[2073] Peter D. Robinson. Behavioural toxicity of organic chemical contaminants in fish: application to ecological risk assessments (ERAs). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 66(7):1179–1188, July 2009. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/F09-069.

Keller:2009:LNO

[2074] W. (Bill) Keller. Limnology in northeastern Ontario: from acidification to multiple stressors. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 66 (7):1189–1198, July 2009. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/F09-080.

Warner:2009:LHP

[2075] David M. Warner, Jeffrey S. Schaeffer, and Timothy P. O’Brien. The Lake Huron pelagic fish community: persistent spatial pattern along biomass and species composition gradients. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 66(8):1199–1215, August 2009. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/F09-072.

Small:2009:ISG

[2076] Maureen P. Small, Ken Currens, Thom H. Johnson, Alice E. Frye, and Jennifer F. Von Bargen. Impacts of supplementation: genetic diversity in supplemented and unsupplemented populations of summer chum salmon (Oncorhynchus keta) in Puget Sound (Washington, USA). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 66(8):1216–1229, August 2009. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/F09-068.

Bacheler:2009:CTT

[2077] Nathan M. Bacheler, Jeffrey A. Buckel, Joseph E. Hightower, Lee M. Paramore, and Kenneth H. Pollock. A combined telemetry–tag return approach to estimate fishing and natural mortality rates of an estuarine fish. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 66(8):1230–1244, August 2009. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/F09-
076.

Cowen:2009:EER

[2078] Laura Cowen, Stephen John Walsh, Carl James Schwarz, Noel Cadigan, and Joanne Morgan. Estimating exploitation rates of migrating yellowtail flounder (Limanda ferruginea) using multistate mark–recapture methods incorporating tag loss and variable reporting rates. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 66(8):1245–1255, August 2009. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/F09-082.

Cope:2009:DLR

[2079] Jason M. Cope and André E. Punt. Drawing the lines: resolving fishery management units with simple fisheries data. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 66(8):1256–1273, August 2009. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/F09-084.

Conley:2009:ACA

[2080] Justin M. Conley, Mark S. Schorr, Mark L. Hanson, Steven J. Symes, and Sean M. Richards. Is ambient chitobiase activity a monitoring tool for impacts on secondary production in lotic systems? Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 66(8):1274–1281, August 2009. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/F09-092.

Monti:2009:SBB

[2081] Dominique Monti and Pierre Legendre. Shifts between biotic and physical driving forces of species organization under natural disturbance regimes.

Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 66(8):1282–1293, August 2009. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/F09-075.

Friedland:2009:RGA

[2082] Kevin D. Friedland, David Moore, and Fiona Hogan. Retrospective growth analysis of Atlantic salmon (Salmo salar) from the Miramichi River, Canada. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 66(8):1294– 1308, August 2009. CODEN CJFSDX. ISSN 0706-652X (print), 1205-

7533 (electronic). URL https://cdnsciencepub.com/doi/full/10.
1139/F09-077.

Arango:2009:HIS

[2083] Clay Porter Arango, Leslie Anne Riley, Jennifer Leah Tank, and Robert Ogden Hall. Herbivory by an invasive snail increases nitrogen fixation in a nitrogen-limited stream. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 66(8):1309–1317, August 2009. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/F09-079.

Pilgrim:2009:GSB

[2084] Erik M. Pilgrim, Jill V. Scharold, John A. Darling, and John R. Kelly. Genetic structure of the benthic amphipod Diporeia (Amphipoda: Pontoporeiidae) and its relationship to abundance in Lake Superior. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 66(8):1318–1327, August 2009. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/F09-086.

Trebitz:2009:GAH

[2085] Anett S. Trebitz, John C. Brazner, Nicholas P. Danz, Mark S. Pearson, Gregory S. Peterson, Danny K. Tanner, Debra L. Taylor, Corlis W. West, and Thomas P. Hollenhorst. Geographic, anthropogenic, and habitat influences on Great Lakes coastal wetland fish assemblages. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 66(8):1328–1342, August 2009. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/F09-089.

Trebitz:2009:PHF

[2086] Anett S. Trebitz, John C. Brazner, Mark S. Pearson, Gregory S. Peterson, Danny K. Tanner, and Debra L. Taylor. Patterns in habitat and fish assemblages within Great Lakes coastal wetlands and implications for sampling design. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 66(8): 1343–1354, August 2009. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/F09-090.

Hidalgo:2009:PEC

[2087] Manuel Hidalgo, Enric Massutí, Beatriz Guijarro, Joan Moranta, Lorenzo Ciannelli, Josep Lloret, Pere Oliver, and Nils Christian Stenseth. Population effects and changes in life history traits in relation to phase transitions induced by long-term fishery harvesting: European hake (Merluccius merluccius) off the Balearic Islands. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 66(8):1355–1370, August 2009. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/F09-081.

Flamarique:2009:MPC

[2088] Iñigo Novales Flamarique, Christina Gulbransen, Moira Galbraith, and Dario Stucchi. Monitoring and potential control of sea lice using an LED-based light trap. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 66(8): 1371–1382, August 2009. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/F09-094.

Hendry:2009:ESL

[2089] Andrew P. Hendry. Ecological speciation! Or the lack thereof? Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 66(8):1383–1398, August 2009. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/F09-074.

Schick:2009:SRB

[2090] Robert S. Schick, Patrick N. Halpin, Andrew J. Read, Christopher K. Slay, Scott D. Kraus, Bruce R. Mate, Mark F. Baumgartner, Jason J. Roberts, Benjamin D. Best, Caroline P. Good, Scott R. Loarie, and James S. Clark. Striking the right balance in right whale conservation.

Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 66(9):1399–1403, September 2009. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/F09-115.

Moir:2009:NMI

[2091] Hamish J. Moir, Christopher N. Gibbins, John M. Buffington, John H. Webb, Chris Soulsby, and Mark J. Brewer. A new method to identify the fluvial regimes used by spawning salmonids. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 66(9):1404–1408, September 2009. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/F09-136.

Loher:2009:SDP

[2092] Timothy Loher and Calvin L. Blood. Seasonal dispersion of Pacific halibut (Hippoglossus stenolepis) summering off British Columbia and the US Pacific Northwest evaluated via satellite archival tagging. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 66(9):1409–1422, September 2009. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/F09-093.

Inoue:2009:RTO

[2093] Mikio Inoue, Hiroshi Miyata, Yousuke Tange, and Yoshinori Taniguchi. Rainbow trout (Oncorhynchus mykiss) invasion in Hokkaido streams, northern Japan, in relation to flow variability and biotic interactions. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 66(9):1423–1434, September 2009. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/F09-088.

Salm:2009:PPA

[2094] Courtney R. Salm, Jasmine E. Saros, Sherilyn C. Fritz, Christopher L. Osburn, and David M. Reineke. Phytoplankton productivity across prairie saline lakes of the Great Plains (USA): a step toward deciphering patterns through lake classification models. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 66(9):1435–1448, September 2009. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/F09-083.

Harvey:2009:ETC

[2095] Chris J. Harvey. Effects of temperature change on demersal fishes in the California Current: a bioenergetics approach. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 66(9):1449–1461, September 2009. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/F09-087.

Langley:2009:IOV

[2096] Adam Langley, Karine Briand, David Seán Kirby, and Raghu Murtugudde. Influence of oceanographic variability on recruitment of yellowfin tuna (Thunnus albacares) in the western and central Pacific Ocean. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 66(9):1462–1477, September 2009. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/F09-096.

Galarza:2009:PGD

[2097] J. A. Galarza, G. F. Turner, E. Macpherson, and C. Rico. Patterns of genetic differentiation between two co-occurring demersal species: the red mullet (Mullus barbatus) and the striped red mullet (Mullus surmuletus). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 66(9):1478–1490, September 2009. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/F09-098.

Israel:2009:PMD

[2098] Joshua A. Israel, K. Jun Bando, Eric C. Anderson, and Bernie May. Polyploid microsatellite data reveal stock complexity among estuarine North American green sturgeon (Acipenser medirostris). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 66(9):1491–1504, September 2009. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/F09-091.

Hankin:2009:URM

[2099] David G. Hankin, Jacqueline Fitzgibbons, and Yaming Chen. Unnatural random mating policies select for younger age at maturity in hatchery Chinook salmon (Oncorhynchus tshawytscha) populations. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 66(9):1505–1521, September 2009. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/F09-085.

Demarty:2009:SDO

[2100] M. Demarty and Y. T. Prairie. In situ dissolved organic carbon (DOC) release by submerged macrophyte–epiphyte communities in southern Quebec lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 66(9):1522–1531, September 2009. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
F09-099.

Wyn:2009:MBF

[2101] Brianna Wyn, Karen A. Kidd, Neil M. Burgess, and R. Allen Curry. Mercury biomagnification in the food webs of acidic lakes in Kejimkujik National Park and National Historic Site, Nova Scotia. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 66(9):1532–1545, September 2009. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/F09-097.

Garner:2009:ECM

[2102] Shawn R. Garner, John W. Heath, and Bryan D. Neff. Egg consumption in mature Pacific salmon (Oncorhynchus spp.). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 66(9):1546–1553, September 2009. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/F09-103.

Whitlock:2009:BMR

[2103] Rebecca Whitlock and Murdoch McAllister. A Bayesian mark–recapture model for multiple-recapture data in a catch-and-release fishery. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 66(9):1554–1568, September 2009. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/F09-100.

Hyde:2009:PGS

[2104] John R. Hyde and Russell D. Vetter. Population genetic structure in the redefined vermilion rockfish (Sebastes miniatus) indicates limited larval dispersal and reveals natural management units. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 66(9):1569–1581, September 2009. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/F09-104.

Skjaeraasen:2009:MRO

[2105] Jon Egil Skjæraasen, James Kennedy, Anders Thorsen, Merete Fonn, Bente Njøs Strand, Ian Mayer, and Olav Sigurd Kjesbu. Mechanisms regulating oocyte recruitment and skipped spawning in Northeast Arctic cod (Gadus morhua). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 66(9): 1582–1596, September 2009. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/F09-102.

Collins:2009:RWD

[2106] A. Kathleen Collins, Susan E. Allen, and Rich Pawlowicz. The role of wind in determining the timing of the spring bloom in the Strait of Georgia. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 66(9):1597–1616, September 2009. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/F09-071.

Bernard:2009:HDD

[2107] Andrea M. Bernard, Moira M. Ferguson, David L. G. Noakes, Bruce J. Morrison, and Chris C. Wilson. How different is different? Defining management and conservation units for a problematic exploited species. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 66(9):1617–1630, September 2009. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/F09-106.

Branch:2009:CDP

[2108] Trevor A. Branch. Corrigendum: Differences in predicted catch composition between two widely used catch equation formulations. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 66(9):1631, September 2009. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/F09-107.

Andrews:2009:ELR

[2109] Allen H. Andrews, Dianne M. Tracey, and Matthew R. Dunne. Erratum: Lead–radium dating of orange roughy (Hoplostethus atlanticus): validation of a centenarian life span. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 66(9):1632–1633, September 2009. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/F09-135.

Breton:2009:LPP

[2110] Julie Breton, Catherine Vallières, and Isabelle Laurion. Limnological properties of permafrost thaw ponds in northeastern Canada. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 66(10):1635–1648, October 2009. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/F09-108.

Figueiredo-Barros:2009:BBE

[2111] Marcos Paulo Figueiredo-Barros, Adriano Caliman, João J. F. Leal, Reinaldo L. Bozelli, Vinicius F. Farjalla, and Francisco A. Esteves. Benthic bioturbator enhances CH_{4} fluxes among aquatic compartments and atmosphere in experimental microcosms. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 66(10):1649–1657, October 2009. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/F09-111.

Copeland:2009:CTL

[2112] Timothy Copeland and David A. Venditti. Contribution of three life history types to smolt production in a Chinook salmon (Oncorhynchus tshawytscha) population. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 66(10): 1658–1665, October 2009. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/F09-110.

Oskarsson:2009:STV

[2113] Gudmundur J. Óskarsson and Christopher T. Taggart. Spawning time variation in Icelandic summer-spawning herring (Clupea harengus). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 66(10):1666–1681, October 2009. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/F09-095.

Eimers:2009:LTD

[2114] M. Catherine Eimers, Shaun A. Watmough, Andrew M. Paterson, Peter J. Dillon, and Huaxia Yao. Long-term declines in phosphorus export from forested catchments in south-central Ontario. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 66(10):1682–1692, October 2009. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/F09-101.

Hallfredsson:2009:EPJ

[2115] Elvar H. Hallfredsson and Torstein Pedersen. Effects of predation from juvenile herring (Clupea harengus) on mortality rates of capelin (Mallotus villosus) larvae. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 66(10): 1693–1706, October 2009. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/F09-105.

Jensen:2009:ERF

[2116] Olaf P. Jensen, David J. Gilroy, Zeb Hogan, Brant C. Allen, Thomas R. Hrabik, Brian C. Weidel, Sudeep Chandra, and M. Jake Vander Zanden. Evaluating recreational fisheries for an endangered species: a case study of taimen, Hucho taimen, in Mongolia. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 66(10):1707–1718, October 2009. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/F09-109.

Pardoe:2009:TTP

[2117] Heidi Pardoe, Anssi Vainikka, Gudmundur Thórdarson, Gudrun Marteinsdóttir, and Mikko Heino. Temporal trends in probabilistic maturation reaction norms and growth of Atlantic cod (Gadus morhua) on the Icelandic shelf. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 66(10): 1719–1733, October 2009. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/F09-132.

Keefer:2009:BTA

[2118] Matthew L. Keefer, Christopher A. Peery, and Brett High. Behavioral thermoregulation and associated mortality trade-offs in migrating adult steelhead (Oncorhynchus mykiss): variability among sympatric populations. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 66(10):1734–1747, October 2009. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/F09-
131.

Pangle:2009:LDP

[2119] Kevin L. Pangle and Scott D. Peacor. Light-dependent predation by the invertebrate planktivore Bythotrephes longimanus. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 66(10):1748–1757, October 2009. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/F09-133.

Peterson:2009:APS

[2120] Douglas P. Peterson and William R. Ardren. Ancestry, population structure, and conservation genetics of Arctic grayling (Thymallus arcticus) in the upper Missouri River, USA. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 66(10):1758–1774, October 2009. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/F09-113.

Schwarz:2009:HBA

[2121] Lisa K. Schwarz and Michael C. Runge. Hierarchical Bayesian analysis to incorporate age uncertainty in growth curve analysis and estimates of age from length: Florida manatee (Trichechus manatus) carcasses. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 66(10):1775–1789, October 2009. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/F09-117.

Brown:2009:OCA

[2122] Randy J. Brown and Kenneth P. Severin. Otolith chemistry analyses indicate that water Sr : Ca is the primary factor influencing otolith Sr : Ca for freshwater and diadromous fish but not for marine fish. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 66(10):1790–1808, October 2009. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/F09-112.

Lewy:2009:MDF

[2123] Peter Lewy and Kasper Kristensen. Modelling the distribution of fish accounting for spatial correlation and overdispersion. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 66(10):1809–1820, October 2009. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/F09-114.

Cushman:2009:EDS

[2124] Elizabeth L. Cushman, Nathaniel K. Jue, Allan E. Strand, and Erik E. Sotka. Evaluating the demographic significance of genetic homogeneity using a coalescent-based simulation: a case study with gag (Mycteroperca microlepis). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 66(10): 1821–1830, October 2009. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/F09-140.

Blanchfield:2009:JVL

[2125] Paul J. Blanchfield, Michael J. Paterson, John A. Shearer, and David W. Schindler. Johnson and Vallentyne’s legacy: 40 years of aquatic research at the Experimental Lakes Area. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 66(11):1831–1836, November 2009. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/F09-148.

Schindler:2009:PHE

[2126] D. W. Schindler. A personal history of the Experimental Lakes Project. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 66(11):1837–1847, November 2009. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/F09-134.

Parker:2009:LTC

[2127] Brian R. Parker, David W. Schindler, Ken G. Beaty, Michael P. Stainton, and Susan E. M. Kasian. Long-term changes in climate, streamflow, and nutrient budgets for first-order catchments at the Experimental Lakes Area (Ontario, Canada). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 66(11): 1848–1863, November 2009. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/F09-149.

Hesslein:2009:SEW

[2128] Raymond H. Hesslein, Michael A. Turner, Douglas Guss, and Mark Lyng. Separating the effects on water chemistry of climate variation and experimental manipulation in the long-term acidification and recovery of lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 66(11):1864–1874, November 2009. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/F09-124.

Turner:2009:BAC

[2129] Michael A. Turner, David L. Findlay, Helen M. Baulch, Llwellyn M. Armstrong, Susan E. M. Kasian, Donald K. McNicol, and Rolf D. Vinebrooke.

Benthic algal communities: recovery from experimental acidification. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 66(11):1875–1891, November 2009. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/F09-126.

Phillips:2009:ERC

[2130] Iain D. Phillips, Rolf D. Vinebrooke, and Michael A. Turner. Experimental reintroduction of the crayfish species Orconectes virilis into formerly acidified Lake 302S (Experimental Lakes Area, Canada). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 66(11):1892–1902, November 2009. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/F09-118.

Vinebrooke:2009:SIT

[2131] Rolf D. Vinebrooke, Michael A. Turner, David L. Findlay, and Michael J. Paterson. A stressor-independent test for biodiversity–ecosystem function relationships during a 23-year whole-lake experiment. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 66(11):1903–1909, November 2009. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/F09-120.

Hall:2009:CMM

[2132] Britt D. Hall, Katharine A. Cherewyk, Michael J. Paterson, and R. (Drew) A. Bodaly. Changes in methyl mercury concentrations in zooplankton from four experimental reservoirs with differing amounts of carbon in the flooded catchments. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 66(11):1910–1919, November 2009. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/F09-123.

Palace:2009:IDB

[2133] V. P. Palace, R. E. Evans, K. G. Wautier, K. H. Mills, P. J. Blanchfield, B. J. Park, C. L. Baron, and K. A. Kidd. Interspecies differences in biochemical, histopathological, and population responses in four wild fish species exposed to ethynylestradiol added to a whole lake. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 66(11):1920–1935, November 2009. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/F09-125.

Findlay:2009:AIA

[2134] David L. Findlay, Cheryl L. Podemski, and Susan E. M. Kasian. Aquaculture impacts on the algal and bacterial communities in a small boreal forest lake. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 66(11):1936– 1948, November 2009. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/F09-121.

Rooney:2009:EER

[2135] Rebecca C. Rooney and Cheryl L. Podemski. Effects of an experimental rainbow trout (Oncorhynchus mykiss) farm on invertebrate community composition. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 66(11):1949– 1964, November 2009. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/F09-130.

Kullman:2009:AFS

[2136] Marilynn A. Kullman, Karen A. Kidd, Cheryl L. Podemski, Michael J. Paterson, and Paul J. Blanchfield. Assimilation of freshwater salmonid aquaculture waste by native aquatic biota. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 66(11):1965–1975, November 2009. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/F09-128.

Blanchfield:2009:SBR

[2137] Paul J. Blanchfield, Lori S. Tate, and Cheryl L. Podemski. Survival and behaviour of rainbow trout (Oncorhynchus mykiss) released from an experimental aquaculture operation. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 66(11):1976–1988, November 2009. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/F09-127.

Baulch:2009:BAB

[2138] Helen M. Baulch, Michael A. Turner, David L. Findlay, Rolf D. Vinebrooke, and William F. Donahue. Benthic algal biomass — measurement and errors. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 66(11):1989– 2001, November 2009. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/F09-122.

Xenopoulos:2009:ELR

[2139] Marguerite A. Xenopoulos, Peter R. Leavitt, and David W. Schindler. Ecosystem-level regulation of boreal lake phytoplankton by ultraviolet radiation. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 66(11):2002–2010, November 2009. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
F09-119.

Plumb:2009:PTD

[2140] John M. Plumb and Paul J. Blanchfield. Performance of temperature and dissolved oxygen criteria to predict habitat use by lake trout (Salvelinus namaycush). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 66(11):2011– 2023, November 2009. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/F09-129.

Benoit:2009:CDS

[2141] Hugues P. Benoît and Jacques Allard. Can the data from at-sea observer surveys be used to make general inferences about catch composition and discards? Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 66(12): 2025–2039, December 2009. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/F09-116.

Gauthier-Ouellet:2009:SDA

[2142] Marika Gauthier-Ouellet, Mélanie Dionne, François Caron, Tim L. King, and Louis Bernatchez. Spatiotemporal dynamics of the Atlantic salmon (Salmo salar) Greenland fishery inferred from mixed-stock analysis. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 66(12):2040–2051, December 2009. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/F09-147.

Holst:2009:QSS

[2143] René Holst, Richard S. T. Ferro, Ludvig Ahm Krag, Robert J. Kynoch, and Niels Madsen. Quantification of species selectivity by using separating devices at different locations in two whitefish demersal trawls. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 66(12):2052–2061, December 2009. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/F09-145.

Forrest:2009:ETO

[2144] Robyn E. Forrest and Carl J. Walters. Estimating thresholds to optimal harvest rate for long-lived, low-fecundity sharks accounting for selectivity and density dependence in recruitment. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 66(12):2062–2080, December 2009. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/F09-137.

Zwolinski:2009:EFA

[2145] Juan Zwolinski, Paul G. Fernandes, Vítor Marques, and Yorgos Stratoudakis. Estimating fish abundance from acoustic surveys: calculating variance due to acoustic backscatter and length distribution error. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 66(12):2081–2095, December 2009. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/F09-138.

Rennie:2009:FAG

[2146] Michael D. Rennie, W. Gary Sprules, and Timothy B. Johnson. Factors affecting the growth and condition of lake whitefish (Coregonus clupeaformis). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 66(12):2096– 2108, December 2009. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/F09-139.

Enders:2009:DSF

[2147] Eva C. Enders, Michael H. Gessel, and John G. Williams. Development of successful fish passage structures for downstream migrants requires knowledge of their behavioural response to accelerating flow. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 66(12):2109–2117, December 2009. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/F09-141.

Schmidt:2009:LTF

[2148] Stephanie N. Schmidt, M. Jake Vander Zanden, and James F. Kitchell. Long-term food web change in Lake Superior. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 66(12):2118–2129, December 2009. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/F09-151.

vanDamme:2009:FAS

[2149] Cindy J. G. van Damme, Mark Dickey-Collas, Adriaan D. Rijnsdorp, and Olav S. Kjesbu. Fecundity, atresia, and spawning strategies of Atlantic herring (Clupea harengus). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 66(12):2130–2141, December 2009. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/F09-153.

Borden:2009:PPD

[2150] W. Calvin Borden and Robert A. Krebs. Phylogeography and postglacial dispersal of smallmouth bass (Micropterus dolomieu) into the Great Lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 66(12):2142– 2156, December 2009. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/F09-155.

Tonina:2009:TDM

[2151] Daniele Tonina and John M. Buffington. A three-dimensional model for analyzing the effects of salmon redds on hyporheic exchange and egg pocket habitat. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 66(12): 2157–2173, December 2009. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/F09-146.

Lowe:2009:CLB

[2152] Michael R. Lowe, Dennis R. DeVries, Russell A. Wright, Stuart A. Ludsin, and Brian J. Fryer. Coastal largemouth bass (Micropterus salmoides) movement in response to changing salinity. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 66(12):2174–2188, December 2009. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/F09-152.

Burtnyk:2009:SNS

[2153] Michael D. Burtnyk, Gordon Paterson, Kenneth G. Drouillard, and G. Douglas Haffner. Steady and non-steady state kinetics describe polychlorinated biphenyl bioaccumulation in natural populations of bluegill (Lepomis macrochirus) and cisco (Coregonus artedi). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 66(12):2189–2198, December 2009. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/F09-150.

Dorner:2009:EPA

[2154] Brigitte Dorner, Randall M. Peterman, and Zhenming Su. Evaluation of performance of alternative management models of Pacific salmon (Oncorhynchus spp.) in the presence of climatic change and outcome uncertainty using Monte Carlo simulations. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 66(12):2199–2221, December 2009. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/F09-144.

Amar:2009:IRS

[2155] Z. Teresa A’mar, André E. Punt, and Martin W. Dorn. The impact of regime shifts on the performance of management strategies for the Gulf of Alaska walleye pollock (Theragra chalcogramma) fishery. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 66(12):2222–2242, December 2009. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/F09-142.

Buchanan:2009:DMD

[2156] Rebecca A. Buchanan, John R. Skalski, and Geoffrey A. McMichael. Differentiating mortality from delayed migration in subyearling fall Chinook salmon (Oncorhynchus tshawytscha). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 66(12):2243–2255, December 2009. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/F09-154.

McQuinn:2009:PFO

[2157] Ian H. McQuinn. Pelagic fish outburst or suprabenthic habitat occupation: legacy of the Atlantic cod (Gadus morhua) collapse in eastern Canada. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 66(12):2256–2262, December 2009. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
F09-143.

Anonymous:2009:AII

[2158] Anonymous. Author index / index des auteurs: Volume 66, 2009. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 66(12), December 2009. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f09-905.

Anonymous:2009:CSV

[2159] Anonymous. Contents / sommaire: Volume 66, 2009. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 66(12), December 2009. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f09-906.

Beacham:2010:CGF

[2160] Terry D. Beacham and Ruth E. Withler. Comment on “Gene flow increases temporal stability of Chinook salmon (Oncorhynchus tshawytscha) populations in the Upper Fraser River, British Columbia, Canada”. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 67(1):202–205, January 2010. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/F09-177. See

[1985, 2161].

Walter:2010:RCB

[2161] Ryan P. Walter, J. Mark Shrimpton, and Daniel D. Heath. Reply to the comment by Beacham and Withler on “Gene flow increases temporal stability of Chinook salmon (Oncorhynchus tshawytscha) populations in the Upper Fraser River, British Columbia, Canada”. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 67(1):206–208, January 2010. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/F09-178. See [1985, 2160].

Taylor:2011:CNL

[2162] Nathan G. Taylor, Carl J. Walters, and Steven J. D. Martell. Corrigendum: A new likelihood for simultaneously estimating von Bertalanffy growth parameters, gear selectivity, and natural and fishing mortality. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 68(8):1507, August 2011. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/

f2011-100. See [1077].